A Complete Bibliography of Publications in the British Journal for the History of Science: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.33

Title word cross-reference

$1 [Dea02]. $10.95 [Mac00a]. $100.00 [Web08]. $102.50 [Aga04]. $103.00 [Lüt04a, Man04]. $105.95 [Han07b]. $110.00 [Mac09a]. $112.00 [Gui04]. $12.00 [Kra07]. $120.00 [Kus07, Mor01c]. $122.00 [Cuo02, Nut05, Ter08]. $124.00 [Chi06]. $124.95 [Hig07]. $127.00 [Tso05]. $129.00 [Smi09a]. $13.50 [Tho03c]. $130.00 [Joh05c]. $132.00 [Gol05]. $14.00 [Hug07b]. $14.95 [Goo09b]. $143.00 [Hen03a]. $149.00 [Elm09]. $15.00 [Gor02]. $15.95 [Aga01b]. $150 [Ric01b]. $150.00 [Sho03]. $152.00 [Aug09]. $16.00 [Hut04]. $16.95 [Aga01a, Tim02]. $169.00 [Kim05]. $17.00 [Ful03]. $17.95 [Hal04a]. $179.00 [Gra06]. $18.00 [Hal02a, Wal03]. $18.95 [Man01a, Sch03a, Sta03]. $189.00 [Mos08b]. $19.00 [Bor03, Boh01]. $19.50 [Cha03]. $19.95 [Bar01b, Boo07, Fri03a, Har08, Lam02, Mun08, Rol06, Tim05]. $195.00 [Hun05b]. $199.00 [Elm09]. $20.00 [Can02b, Phi02, Qui04, Yam04b, Yeo04, Goo02a]. $21.95 [Sha02]. $217.00 [Plo05]. $22.00 [Dix06, Nes06b, Nut04, Suz05]. $22.50 [Ade04, Gat05, Jen01, Joh06c]. $23.00 [And03a, Dea06a, Mül01]. $24.00 [Jam03, Lip08, Smi05b]. $24.95 [Dol02c, Ful01b, Lau06, Mar05, Mor03b, Seg05, Smi02b, Tos09, Wil09d]. $24.99 [Coo07a, Gau08]. $25.00 [App05a, Bar01a, Goo02b, Hei07b, Hen05b, Hen04e, Jon04, Out04, Pem07, Ric01a, Sch01, She07, Siv08, Sta06b, Ste09b, Tur06b]. $25.95 [Bru06, Mac07]. $26.00 [Far06, Fre04d, Ful03, Jor02, Mor05b, Sta07b, Tim03c]. $26.95 [Biv06]. $26.99 [Joh05b, Tso07]. $27.00 [Biv06, Gar03b]. $27.50 [Bro05a, Chr05, Edd05b, Gil08, Hei07a, Rob05, Sha07]. $27.95 [Hes04, Kas02, Tim03a]. $28.00 [Goo08a, Gou09a, Hay03, Smi06f]. $280.00 [Nea00b, Nea00b]. $285.00 [Hun05b]. $29.00 [Bud09, Dew07, Dix05, Ful09b, Gra05]. $29.95 [Art04, Bee01, Ben02, Ber08, Dol02b, Edg08, Fan02, Fra05, Kni03, Mer08, Sta07a, Stu06, Por01]. $29.99 [Biv08, Pal06]. $30.00 [Alb03b, Dol01, Han07a, Hay06, Hei08, Hen05b, Hig04a, Kra08a, Pag01, Sta03, Van04a]. $315.00 [Hun05b]. $32.50 [Chi04a]. $33.00 [Por03]. $34.00 [Tot04]. $34.50 [App05b]. $345.00 [D’A03]. $35.00 [Big07, Boo07, Bra00, Bra07, Coo06a, Del08, Dod02, Fox07, GG06a, Hag04, Han07b, Hei08, Hen06b, Hou07, Hut09, Joh07b, Jor02, Kni05, Mag06b, Nay09, Shi03b, Tim04, Top02, Cor00, Dix00]. $36.50 [Fin01]. $37.50 [Aga01a, Bar08, Bro09, Goo09a, Har02c, Hug09, Jor06, Roo08]. $38.00 [DiM07]. $39.00 [Aga06, GG04, Ric01a]. $39.50 [Tyb04]. $39.95 [Haf04, Joh07a, Jor01, Lin07, Log08, Mor06c, Nay09, Spa06, WM08, Mac01a]. $40.00 [Brö02b, Edd05b, For05, Hen05b, Hug08, Jar08a, Sle02, Smi07b, Web06, Web09]. $43.00 [Web09]. $44.50 [Eds04]. $45.00 [Ade06, Dix08a, Edd07b, Gol09, Har05a, Kni04, Kni08, Log08, Mor06c, Mun09, Ric09b, Sch08a, Smi08d, Smi09a, Web09, Whi04, Dan01, Jam01a, Mor01b]. $47.50 [Mor06b]. $49.50 [Aga02, Har00a, Mun07, Nes02]. $49.95 [Dig08, Mcc05, Roo08, Tur09a, Col01b]. $50.00 [Ade08, Bro02e, Das07, Dea07, Han07a, Qui02, Siv08, Ben00]. $50.50 [Bre05]. $52.50 [Goo06a]. $53.00 [Goo08b]. $54.95 [Ful01a, Ras06, Har01a]. $55.00 [Bax09, Del06a, Mel01, Ste07, Emm01, Smi02a]. $56.00 [Par01a]. $56.95 [Aga00].

$58.00 [Kus03, Lea05]. $59.95 [Daw06, Kip04, Mac08c, Man01a, Olb02, Høy01]. $60.00 [Hen07a, Lip08, Pet07b, Smi06g]. $62.00 [Coo02b]. $630.00 [Rad03b]. $64.95 [Col00, Goo01, Mac00b]. $65.00 [Can08b, Coo07a, Goo06b, Jen02, Jor08, Rei07a, Sha07]. $67.50 [Spa02]. $69.00 [GG09]. $69.95 [Wei02a]. $70.00 [Can04a, Gau08, Goo08a]. $72.50 [Ful09b]. $75.00 [Bow07, Bra00, Eag05b, Edd09, Kir05, Kre06, Mac09b, Mir07, Mor08]. $79.95 [Jam05]. $80.00 [Dew07, Eag06b, Sta07a, Ste04b, Urr07]. $82.00 [Aga04]. $84.00 [Kle04]. $84.95 [Wor07]. $85.00 [But06, Hul07, Lin05, Smi08f, Spa05, Twe09]. $86.00 [DM05b]. $88.00 [GG02]. $88.50 [Jen01]. $89.00 [GG06b, Hen03b]. $89.95 [Tim08]. $9.95 [CK02]. $90.00 [Elm02, Lev07, Tur08, Ren00]. $94.00 [Edd07a]. $94.95 [Chi08]. $95.00 [Hun02, Joh06a, Sta06c]. $99.00 [Cla09, Joh02d, Maz09]. $99.95 [Bro01b, Eag08]. $995.00 [Sha09]. ≫ [Spr06]. ≪ [Spr06].

/a [TBbTc90].

0 [Mac05]. 0-00-710189-9 [Dea05].

0-0674-01699-8 [Mac08c]. 0-09-974971-8 [Bow04]. 0-12-088459-3 [Web07]. 0-13-1777390-9 [Goo08b]. 0-140-16735-8 [Smi06e]. 0-140-43631-6 [Ric06]. 0-16-061774-X [Aga04]. 0-16-066599-X [Aga04]. 0-19-262927-1 [Smi03a]. 0-19-263285-X [Pen04]. 0-19-280266-6 [Suz05]. 0-19-280404-9 [Dol04b]. 0-19-286207-3 [Aga01b]. 0-19-507870-5 [Ris01]. 0-19-511229-6 [Fyf05]. 0-19-5113993 [Har00b]. 0-19-511557-0 [Fyf02a]. 0-19-511763-8 [Jam03]. 0-19-511987-8 [Ful02]. 0-19-512137-6 [Cha01]. 0-19-513050-2 [Alb03b]. 0-19-514326-4 [Dol02b]. 0-19-514694-8 [Van05]. 0-19-514922-X [Rob08]. 0-19-517127-6 [Kni06a]. 0-19-517198-5 [Big07]. 0-19-517408-9 [Vic07]. 0-19-517433-X [Sch08a]. 0-19-517717-7 [Pet07b]. 0-19-517735-5 [Maz07a]. 0-19-564194-9 [Mac00a]. 0-19-726326-7 [Fyf07]. 0-19-812283-7 [Smi01b]. 0-19-812290-X [Man00a]. 0-19-814687-6 [Doo04]. 0-19-818470-0 [Hen02a]. 0-19-818640-1 [Sle03]. 0-19-820608-9 [Out01]. 0-19-820657-7 [Hug01]. 0-19-825079-7 [Hod06b]. 0-19-825080-0 [Hod06b]. 0-19-850297-4 [Rav01]. 0-19-850584-1 [Har03c]. 0-19-850651-1 [Maz02b]. 0-19-850658-9 [Kai05a]. 0-19-850791-7 [Smi07a]. 0-19-850862-X [Hen04d]. 0-19-850872-7 [Bla03]. 0-19-850877-8 [Eag06a]. 0-19-850953-7 [Joh03b]. 0-19-851053-5 [Lüt04b]. 0-19-851530-8 [Bro04a]. 0-19-851538-3 [Brö04b]. 0-19-851544-8 [Hug07a]. 0-19-852675-X [Gui05]. 0-19-853069-2 [Hig07]. 0-19-855210-6 [Yam00]. 0-19-856484-8 [Hut07]. 0-19-856566-6 [Hig01]. 0-19-856792-8 [Sta07d]. 0-19-857122-4 [Aga08a]. 0-19-858409-1 [Das07]. 0-19-924144-9 [Abr08]. 0-19-924639-4 [Tho03a]. 0-19-924748-X [Liv05a]. 0-19-924792-7 [Man06]. 0-19-924962-8 [Noa04]. 0-19-925320-X [Jon06a]. 0-19-925323-4 [Doo05]. 0-19-925921-6 [Sum05]. 0-19-926897-5 [Kus07]. 0-19-927016-3 [Hsu06]. 0-19-927629-2 [Rei07b]. 0-19-927668-4 [Sta07c]. 0-19-927824-5 [Sta09a]. 0-19-9279055 [Tra07]. 0-19-928530-6 [Leh08]. 0-19-928878-X [Smi08c]. 0-19-929644-8 [Bro08b]. 0-19-96667-0 [Sle05]. 0-1981-8700-9 [Wal06b]. 0-226-00092-3 [Ric01a]. 0-226-00093-1 [Ric01a]. 0-226-019680-3 [Edd07b]. 0-226-02420-2 [Goo08a]. 0-226-02421-0 [Goo08a]. 0-226-04264-2 [App05a]. 0-226-04561-7 [She07]. 0-226-0681-7 [Hei07b]. 0-226-07371-8 [Bre05]. 0-226-08089-7 [Dew07]. 0-226-08090-0 [Dew07]. 0-226-09276-3 [Lip08]. 0-226-09277-1 [Lip08]. 0-226-09907-5 [Phi02]. 0-226-10225-4 [Hag04]. 0-226-1091-6 [Kni08]. 0-226-10939-9 [Ril02]. 0-226-11292-6 [Van04a]. 0-226-11366-3 [Pem07]. 0-226-11378-7 [Aga06]. 0-226-13672-8 [Boh01]. 0-226-13681-7 [Far06]. 0-226-13948-4 [Hei07a]. 0-226-16472-1 [Siv08]. 0-226-18487-0 [Cor00]. 0-226-20639-4 [Hal02a]. 0-226-24613-2 [Web06]. 0-226-26896-9 [Gat05]. 0-226-27648-1 [Tur06b]. 0-226-27916-2 [Ren00]. 0-226-28446-8 [Bro05a]. 0-226-31976-8 [Nes06b]. 0-226-32287-4 [Yeo04]. 0-226-32733-7 [Hut04]. 0-226-33387-6 [Goo02b]. 0-226-35420-2 [Dod02]. 0-226-35933-6 [Joh06c]. 0-226-40954-6 [Sha07]. 0-226-40955-4 [Sha07]. 0-226-42267-4 [Kra08a]. 0-226-45284-0 [Tim03c]. 0-226-45798-2 [Bar01a]. 0-226-46723-6 [Ful03]. 0-226-47536-0 [Whi04]. 0-226-48214-6 [Dix05]. 0-226-48733-4 [Smi08a]. 0-226-53857-5 [Por03]. 0-226-54202-5 [Sta06b]. 0-226-57711-2 [Edd05b]. 0-226-57714-7 [Edd05b]. 0-226-66795-2 [Chr05]. 0-226-71210-9 [Kni05]. 0-226-72079-9 [Out04]. 0-226-73111-1 [Mor06c]. 0-226-73530-3 [Tot04]. 0-226-73569-9 [Smi07b]. 0-226-74410-8 [Top02]. 0-226-74863-4 [Log08]. 0-226-75754-4 [Qui04]. 0-226-76733-7 [Dix08a]. 0-226-76863-5 [Sch01]. 0-226-79845-3 [Hug09]. 0-226-81140-9 [Pag01]. 0-226-81740-7 [Bor03]. 0-226-87158-4 [Del08]. 0-226-87375-7 [Gra05]. 0-226-89436-3 [Can02b]. 0-231-12004-4 [Dol02c]. 0-231-13010-4 [Web06]. 0-231-13038-4 [Mun07]. 0-252-02601-2 [Dol02b]. 0-26-76399-4 [Hen06b]. 0-262-02494-2 [Hun05a]. 0-262-02589-2 [Smo08]. 0-262-04215-0 [Sir06]. 0-262-06254-2 [Can08a]. 0-262-07193-2 [Van02b]. 0-262-07232-7 [Cla06]. 0-262-07245-9 [Ner05]. 0-262-08282-9 [Ram02]. 0-262-08285-3 [Aga03a]. 0-262-08298-5 [Hem03]. 0-262-10084-3 [Mor02c]. 0-262-10116-5 [Pan08]. 0-262-11288-4 [Hun07a]. 0-262-11297-3 [Hug08]. 0-262-13363-6 [Hay02b]. 0-262-13369-5 [Jon02a]. 0-262-13393-8 [Col02]. 0-262-13397-0 [Joh04c]. 0-262-13421-7 [Aga05]. 0-262-14075-6 [Pag03]. 0-262-14093-4 [Mac08a]. 0-262-16229-6 [De 07]. 0-262-18204-1 [Car01a]. 0-262-18226-2 [Sum06]. 0-262-18232-7 [Wil06]. 0-262-19449-X [Fin02]. 0-262-20138-0 [Pem04]. 0-262-23223-5 [Dol04c]. 0-262-51124-X [Goo03]. 0-268-02201-1 [Hei08]. 0-268-03071-5 [Dea07]. 0-268-03484-2 [Han07a]. 0-292-75257-1 [Hal04b]. 0-295-98581-X [Siv08]. 0-300-05976-0 [Sch02b]. 0-300-07016-0 [Aga01a]. 0-300-07383-6 [Dix00]. 0-300-07815-3 [Aga01a]. 0-300-08252-5 [Dew01]. 0-300-08536-2 [For03]. 0-300-08813-2 [Har06]. 0-300-10334-4 [Mor07c]. 0-300-10899-0 [Web09]. 0-300-11263-7 [Hut09]. 0-306-47385-2 [Joh06a]. 0-309-06987-4 [Sta03]. 0-309-08408-3 [Hes04]. 0-316-64847-7 [Cro04]. 0-316-85989-3 [Goo04a]. 0-333-71574-8 [Hen04b]. 0-333-77278-4 [Rei07a]. 0-333-90735-3 [Hig03a]. 0-333-97002-0 [Hen04c]. 0-333-97006-3 [Hen04c]. 0-335-21548-3 [Top08]. 0-375-40128-8 [Coo06a]. 0-387-95136-9 [Mcc05]. 0-393-04124-7 [Aga00]. 0-415-10140-7 [Smi00a]. 0-415-10141-7 [Smi00a]. 0-415-20638-3 [Sec08]. 0-415-26905-9 [Til04]. 0-415-28603-4 [Smi05a]. 0-415-286065-9 [Tur06a]. 0-415-28922-X [Bow06a]. 0-415-30174-2 [Dig08]. 0-415-39142-3 [dC08]. 0-415-92112-0 [Gou03c]. 0-415-96688-4 [Ree07]. 0-431-10443-3 [Jam03]. 0-434-01315-3 [Hen06a]. 0-444-50871-6 [Pet07a]. 0-465-01859-9 [Sta03]. 0-470-86088-X [Hun06]. 0-472-11132-9 [Eds04]. 0-485-11431-3 [Yos00]. 0-485-11468-2 [Hen04a]. 0-485-11575-1 [For04]. 0-485-12145-X [Yos00]. 0-520-08816-6 [Bra00]. 0-520-08817-4 [Bra00]. 0-520-22441-8 [Rob05]. 0-520-23106-6 [Ste04b]. 0-520-23249-6 [Goo06b]. 0-520-23531-2 [Seg05]. 0-520-23793-5 [Mar05]. 0-520-24135-5 [Ste07]. 0-520-24261-0 [Han07a]. 0-520-24580-6 [Dig08]. 0-521-00000-9 [Smi02b]. 0-521-00062-9 [Yam04b]. 0-521-00611-2 [Dea06a]. 0-521-25625-9 [Hun05b]. 0-521-25626-7 [Hun05b]. 0-521-25627-5 [Hun05b]. 0-521-33053 [Mac05]. 0-521-35371-8 [Col01b]. 0-521-43098-4 [But06]. 0-521-44373-3 [Col00]. 0-521-44913-8 [Edd04a]. 0-521-46569-9 [Ric01b]. 0-521-47053-6 [Ful01a]. 0-521-52094-0 [Jon04]. 0-521-53026-1 [Biv06]. 0-521-53909-9 [Smi06f]. 0-521-546681-8 [Coo07a]. 0-521-54689-3 [Tso07]. 0-521-55069-6 [Mac00b]. 0-521-55372-5 [Tyb02]. 0-521-55827-1 [Kas02]. 0-521-56399-2 [Man01a]. 0-521-57199-5 [Sta06c]. 0-521-57244-4 [Wee08]. 0-521-59032-9 [Hod01]. 0-521-59034-5 [Hod06a]. 0-521-59146-5 [Mor03a]. 0-521-62361-8 [Goo01]. 0-521-62672-2 [And03a]. 0-521-63149-1 [Ash01]. 0-521-64202-7 [Edd03a]. 0-521-64211-6 [Høy01]. 0-521-64967-6 [Dix06]. 0-521-65187-5 [Har01a]. 0-521-65191-3 [Joh03a]. 0-521-65270-7 [Wag05]. 0-521-65700-8 [Bar01b]. 0-521-65939-6 [Por01]. 0-521-66160-9 [Ber06a]. 0-521-67231-7 [Mac09b]. 0-521-77187-0 [Olb02]. 0-521-77302-4 [Kip04]. 0-521-78561-8 [Tim05]. 0-521-78612-6 [Yam04b]. 0-521-79067-0 [Leh08]. 0-521-79143-X [Ben02]. 0-521-79662-8 [Smi05b]. 0-521-79937-6 [Lam02]. 0-521-8049-7 [GG06a]. 0-521-80536-8 [Sha02]. 0-521-80699-2 [Bow07]. 0-521-81008-6 [Yam04a]. 0-521-81516-9 [Ren06]. 0-521-81599-1 [Joh05c]. 0-521-82413-3 [Hod06a]. 0-521-82712-4 [Mir07]. 0-521-82729-9 [Smi06g]. 0-521-82873-2 [Lev07]. 0-521-82967-4 [Mac08b]. 0-521-83031-1 [Eag06b]. 0-521-83591-7 [Urr07]. 0-521-84459-2 [Hod06a]. 0-521-84836-9 [Ste06]. 0-521-85439-3 [Dep07]. 0-521-85690-6 [Daw07a]. 0-521-86609-X [Win08]. 0-521-89461-1 [Nut04]. 0-5214-7565-5 [Biv08]. 0-565-09164-6 [Cla05]. 0-566-05481-7 [Fyf02b]. 0-618-34151-X [Sta07b]. 0-631-20112-2 [Bee01]. 0-674-00292-X [Rad03b]. 0-674-00372-1 [Pal02]. 0-674-00412-4 [Rad02]. 0-674-00433-7 [Sta03]. 0-674-00435-3 [Tho02b]. 0-674-00455-8 [Sen02b]. 0-674-00595-3 [Hal03]. 0-674-00647-X [Fre04d]. 0-674-00652-6 [Ful03]. 0-674-01143-0 [Dod06]. 0-674-01157-0 [Hal06]. 0-674-01286-0 [Goo06a]. 0-674-01487-1 [Rob06]. 0-674-01495-2 [Hen07a]. 0-674-01497-9 [Maz07a]. 0-674-01548-7 [Mac07]. 0-674-01685-8 [Bra07]. 0-674-01688-2 [Rob06]. 0-674-01691 [Haf08]. 0-674-01876-2 [Edg08]. 0-674-01892-3 [Fox07]. 0-674-01896-6 [Log08]. 0-674-02299-8 [Ber08]. 0-674-02306-4 [Bra07]. 0-679-42932-8 [Gli06]. 0-691-00261-4 [Pal00]. 0-691-00413-7 [Dan01]. 0-691-01042-0 [Tyb04]. 0-691-01715-8 [Ree08]. 0-691-04918-1 [Aga02]. 0-691-05952-7 [Har05a]. 0-691-07021-0 [Sta03]. 0-691-07051-2 [Tim02]. 0-691-09085-8 [Bru06]. 0-691-09550-7 [Dol04a]. 0-691-09552-3 [Fre04c]. 0-691-10850-0 [Har00a]. 0-691-11445-5 [Mag06b]. 0-691-11541-9 [Cor06]. 0-691-11857-4 [Gan07]. 0-691-12310-1 [War08]. 0-7123-0894-6 [Mok06]. 0-7141-2794-9 [Cha05]. 0-7185-0106-3 [Ver04]. 0-7188-3017-2 [GG03b]. 0-7190-5592-X [Fis04]. 0-7190-5812-0 [Gre02]. 0-7190-5911-9 [Mus03]. 0-7190-5919-4 [Jor02]. 0-7190-5966-6 [Mur02]. 0-7190-5989-5 [Cla01]. 0-7190-6734-0 [Kha06b]. 0-7190-6737-5 [Kha06b]. 0-7190-7373-1 [Hoc08]. 0-7190-7376-6 [Ham08b]. 0-7195-6283-X [Hor03]. 0-7294-0747-0 [Spa02]. 0-7294-0872-8 [Ter08]. 0-7382-0602-4 [Sen06a]. 0-7456-2810-9 [Biv06]. 0-7499-22-397 [Tur02a]. 0-7503-01473-3 [Nea00b]. 0-7503-0391-3 [Nea00b]. 0-7503-0454-5 [Ben00]. 0-7503-0693-9 [Aga01c]. 0-7503-0754-4 [Kre06]. 0-7503-0866-4 [Lea05]. 0-7509-2667-8 [Hes03]. 0-7524-2766-0 [Joh05b]. 0-754-60239-7 [Luc05]. 0-7546-0078-5 [Hen01]. 0-7546-0106-4 [Cla04a]. 0-7546-0139-0 [Gou06]. 0-7546-0202-8 [Fyf03b]. 0-7546-0313-X [Coo03]. 0-7546-0332-6 [Rit02]. 0-7546-0506-X [Alb08]. 0-7546-0528-0 [Bro04c]. 0-7546-0938-3 [Hob06]. 0-7546-3069-2 [Cla07a]. 0-7546-3177-X [Bro05b]. 0-7546-3352-7 [Mor06d]. 0-7546-3404-3 [Coo06b]. 0-7546-3408-6 [Smi06d]. 0-7546-3558-9 [Bre06]. 0-7546-3562-7 [Siv06]. 0-7546-3619-4 [Wil09a]. 0-7546-3671-2 [Far07a]. 0-7546-3679-8 [Dou06]. 0-7546-3933-9 [Pic07]. 0-7546-3996-7 [Bar07]. 0-7546-4102-3 [Eag08]. 0-7546-5044-8 [Rob07]. 0-7546-5296-3 [Kem09]. 0-7546-5359-5 [Wee07]. 0-75546-0397-0 [Mor06e]. 0-7914-6679-5 [Mor08]. 0-7923-6037-0 [Hem02b]. 0-7923-6067-2 [Ghi01]. 0-7923-6298-5 [Loc02]. 0-7923-6349-3 [Joh02d]. 0-7923-6363-9 [D’A03]. 0-7923-6457-0 [Hen03a]. 0-7923-6481-3 [Plo05]. 0-7923-6782-0 [Lüt04a]. 0-7923-6820-7 [Kim05]. 0-7923-6933-5 [GG02]. 0-7923-7022-8 [Kle04]. 0-8014-3764-4 [Sut03]. 0-8014-4127-7 [Mor06b]. 0-8014-4348-2 [Mor06c]. 0-8014-8661-0 [Tho03c]. 0-8018-4011-1 [Lin05]. 0-8018-5868-2 [Ril01a]. 0-8018-5869-0 [Ril01a]. 0-8018-5882-8 [Rad00a]. 0-8018-6390-2 [Sle01]. 0-8018-6391-0 [Chi02b]. 0-8018-6450-X [Hem02a]. 0-8018-6579-4 [Tim03b]. 0-8018-6606-5 [ETV04]. 0-8018-6610-3 [Kni02]. 0-8018-6753-3 [Spa03]. 0-8018-6789-4 [Gar03b]. 0-8018-6884-X [Kni04]. 0-8018-6898-X [Aga04]. 0-8018-7344-4 [Whi06]. 0-8018-7991-4 [Pro08]. 0-8018-7992-2 [Tre06]. 0-8018-8171-4 [Web07]. 0-8018-8291-5 [Gra08b]. 0-8018-8349-0 [Ham07a]. 0-8018-8389-7 [Ade08]. 0-8018-8426-8 [Dea08]. 0-8018-8427-6 [Dea08]. 0-8018-8445-4 [Bur08]. 0-8020-4471-9 [Jen02]. 0-8020-4699-1 [Mel01]. 0-8047-3646-4 [Sch03a]. 0-8047-3874-2 [Mor01b]. 0-8047-3926-9 [Big06]. 0-8047-3972-2 [Lau06]. 0-8047-4359-2 [Car06]. 0-8047-5376-8 [Har08]. 0-80473-260-4 [Ste05a]. 0-8061-3201-9 [Fri03a, Fri03b]. 0-8061-3413-5 [Eag05b]. 0-8078-5488-3 [Rol06]. 0-8087-2791-6 [Stu06]. 0-8122-3598-3 [Har02c]. 0-8122-3827-3 [Del06a]. 0-8126-9425-2 [Ful01b]. 0-8133-3907-3 [Chi04a]. 0-8133-3981-2 [Shi03b]. 0-8135-2809 [Goo02a]. 0-8135-3003-2 [Hay03]. 0-8135-3157-8 [Tim04]. 0-8135-3200-0 [Hig04a]. 0-8135-3662-6 [Lin07]. 0-8176-2741-3 [Hug00]. 0-8176-4030-4 [Sta03]. 0-8204-4108-2 [Wei02a]. 0-8218-2623-9 [GG06b]. 0-8223-2871-2 [GG03a]. 0-8223-3412-7 [Wor07]. 0-8223-3423-2 [Wor07]. 0-8223-3723-1 [Smi08f]. 0-8223-3814-9 [Ren07]. 0-8229-4191-0 [Goo06b]. 0-8229-5733-7 [Hal04a].

0-8229-5795-7 [Fra05]. 0-8248-2366-4 [Fan02]. 0-8264-7186-2 [Mac06b]. 0-8387-5538-0 [Ben06]. 0-85115-706-8 [Nea00a]. 0-85115-801-3 [Elm02]. 0-85115-864-1 [Can04a]. 0-85244-666-7 [Hur07]. 0-85272-467-5 [Mor07b]. 0-85296-792-6 [Hun02]. 0-85296-795-0 [Emm01]. 0-85296-797-7 [Joh02c]. 0-85331-821-2 [Bro03c]. 0-85404-599-6 [Mor01a]. 0-85484-069-9 [Bro02e]. 0-85484-072-9 [Qui02]. 0-85967-992-6 [Sch02a]. 0-86078-633-1 [Ril00]. 0-86078-669-2 [Ril01b]. 0-86078-842-3 [Edd03b]. 0-86078-862-8 [Gou03a]. 0-86078-863-6 [End04a]. 0-86078-865-2 [Bre04]. 0-86078-876-8 [Edd05a]. 0-86078-895-4 [Ste05b]. 0-86078-903-9 [Hig07]. 0-86078-924-1 [Bek06]. 0-86078-952-7 [Sam07]. 0-86078-958-6 [Mor06c]. 0-86078-959-4 [Mor06c]. 0-86078-964-0 [Han07b]. 0-86078-993-4 [Cam08]. 0-87169-237-6 [Dol01]. 0-87169-902-8 [Goo02c]. 0-87169-912-5 [Wal03]. 0-87169-924-9 [Kle05]. 0-87338-809-7 [Daw06]. 0-87580-340-7 [DiM07]. 0-88-135360-4 [Joh07a]. 0-88135-274-8 [Par01a]. 0-88135-283-7 [WM08]. 0-88135-285-3 [Jor01]. 0-88135-288-8 [For05]. 0-88135-294-2 [Kni03]. 0-88135-296-9 [Tim08]. 0-88135-299-3 [Mac01a]. 0-88135-356-6 [Hei08]. 0-88135-365-5 [Ras06]. 0-88135-367-1 [Nay09]. 0-88135-372-8 [Haf08]. 0-88135-385-X [Haf04]. 0-88920-324-5 [Art04]. 0-902205-58-7 [Fri04]. 0-902459-10-5 [Jor02]. 0-904040-56-9 [Can04c]. 0-905193-05-9 [Ash04b]. 0-905193-06-7 [Ash04b]. 0-906271-18-5 [Alb03a]. 0-906271-21-5 [Hig06]. 0-906271-22-3 [Hig06]. 0-906450-13-6 [Jor02]. 0-906450-14-4 [Mor05b]. 0-908011-69-5 [Pre03]. 0-94-305625-X [Gor02]. 0-940228-48-3 [Sle02]. 0-941690-85-7 [Bro03b].

0-941901-29-7 [Mor03b]. 0-942299-90-6 [Fin01]. 0-9504066-7-8 [Wal06c]. 0-9504066-8-6 [Ren06]. 0-9537868-4-6 [Bro06a]. 0-9546484-1-2 [Hen07a]. 0-9583706-7-2 [Mor02a]. 0-9672617-2-4 [Kra07]. 0-9672617-3-2 [Hug07b]. 0-9683689-7-2 [Hen04e]. 00004-2439 [Wer04].

1-124 [Noa09]. 1-4000-7696-X [Goo09b]. 1-4020-0047-2 [Man04]. 1-4020-0100-2 [Hen03b]. 1-4020-0233-5 [Gui04]. 1-4020-0495-8 [Chi06]. 1-4020-0509-1 [Tso05]. 1-4020-1109-1 [Spa05]. 1-4020-1151-2 [DM05b]. 1-4020-1969-6 [Maz07a]. 1-4020-2777-X [Gra06]. 1-4020-4245-0 [Mos08b]. 1-4020-4246-9 [Mos08b]. 1-4039-0488-X [Kha06a]. 1-4039-0643-2 [Edg04]. 1-4039-2170-9 [Can07]. 1-4039-86444-4 [Led08]. 1-55849-221-6 [Cla04b]. 1-56098-533-X [Jam01b]. 1-56098-776-6 [Jam01a]. 1-56098-891-6 [Jam01b]. 1-56881-141-1 [GG04]. 1-57586-292-1 [Cha03]. 1-57607-090-5 [Gou03b]. 1-57607-199-5 [Sho03]. 1-57607-887-6 [Coo07a]. 1-58046-120-4 [Kir05]. 1-58046-195-6 [Hul07]. 1-58465-602-6 [Jor08]. 1-59102-262-2 [Ham07b]. 1-59110-051-8 [Hen05b]. 1-84014-239-1 [Rud07]. 1-84014-625-7 [Alb01]. 1-84046-136-5 [Ful01c]. 1-84046-250-7 [CK02]. 1-84115-060-6 [Ske04]. 1-84115-185-8 [Cro05a]. 1-84169-234-4 [Smi03b]. 1-84306-027-2 [Coo04]. 1-84306-045-0 [Mil04]. 1-84371-113-3 [Maz07a]. 1-84383-133-3 [Maz07a]. 1-85-196778-8 [Maz07a]. 1-851170-31-6 [Can01b]. 1-85182-813-3 [Ade06]. 1-85196-109-7 [Dea02]. 1-85285-309-3 [Goo06c]. 1-85285-331-X [Har03a]. 1-85285-451-0 [Eag07]. 1-85315-547-0 [Spa06]. 1-85410-961-8 [Smi07a]. 1-85433-332-1 [Dix03a]. 1-85506-454-5 [Smi00a].

1-85506-455-3 [Smi00a]. 1-85506-872-9 [Rad03b]. 1-85506-928-8 [Rob03]. 1-85506-999-7 [Mor06a]. 1-85702-959-3 [Pad01]. 1-85928-361-6 [Bro01b]. 1-85928-397-7 [Kus00]. 1-85973-421-9 [Har04c]. 1-85973-571-1 [Ben04b]. 1-85984-652-1 [Gar03b]. 1-85984-687-4 [Pal06]. 1-86094-204-0 [Mil02]. 1-86094-709-3 [Aug09]. 1-8618-9285-0 [Boo07]. 1-86189-084-2 [Dri03]. 1-86189-224-1 [Mun08]. 1-86189-238-1 [Mun08]. 1-86189-255-1 [Boo07]. 1-86189-256-X [Mun08]. 1-86197-296-2 [Far07b]. 1-86239-174-2 [Mor07a]. 1-872501-91-5 [Fin03a]. 1-874045-88-7 [Lea04]. 1-88135-363-9 [Jam05]. 1-884964-29-X [Fyf03a]. 1-888009-09-8 [Mor01c]. 1-888009-10-1 [Mor01c]. 1-888009-11-X [Mor01c]. 1-888009-12-8 [Smi02a]. 1-890951-43-9 [Jor06]. 1-891788-02-7 [Hen05b]. 1-891788-12-4 [Hen05b]. 1-891788-13-2 [Hen05b]. 1-891788-14-0 [Hen05b]. 1-891788-97-3 [Han07b]. 1-900747-24-3 [Dol02a]. 1-900747-31-6 [Div01]. 1-900747-47-2 [Coo02b]. 1-900747-53-7 [Coo06c]. 1-900755-32-7 [Nes02]. 1-901663-76-0 [Joh04b]. 1-901663-88-4 [Edd06]. 1-903985-12-9 [Tho03b]. 1-905383-00-2 [Noa09]. 100.00 [Ash01]. 1000 [Hör78]. 112 [LHL00]. 1137-0238 [Fox05]. 123 [GWW04]. 140.00 [NG01b]. 1431-7613 [Her05b]. 1477-3082 [Pye06]. 15.99 [Pad01]. 1520 [Wit01]. 1600 [vC02]. 166 [Man00b]. 168.00 [Har01c]. 1702 [AH05]. 1798 [Dol00]. 1830/1885 [Sch06b]. 1862 [Hun05b]. 1870s [Sta01b, Tho09]. 1873 [Hun05b]. 1879 [Hun05b]. 19.95 [Aga01c]. 1930 [Gos00]. 1930s [Chi03a, Nor07]. 1970s [Hor07]. 1983 [Deb84]. 1999 [Daw00, Fri01]. 19th [Coo03, Cro01b]. 19th-Century [Coo03, Cro01b].

2 [BRW04]. 2-02-063249-7 [Spr06].

2-13-051976-8 [Bur07a]. 2-13-052659-4 [Chi03c]. 2-226-11537-4 [NG01b]. 2-503-51218-6 [Hig05]. 2-7071-3616-6 [Chi03b]. 2-7117-5353-0 [Cro05b]. 2-7355-0487-5 [Fox04]. 2-7475-5100-8 [Vid05]. 2-8257-0740-6 [Rud04b]. 2-8257-0832-1 [Coo05a]. 2-8321-0047-3 [Rud04b]. 2-84-299-210-5 [Maz04a]. 2-84324-141-3 [Chi01a]. 2-85939-624-1 [Chi01b]. 2-86847-681-3 [Chi04b]. 2-88182-481-1 [Rud04b]. 2-901057-40-3 [Big01]. 2-908327-45-7 [Cor05]. 2.99 [Ful01c]. 20-22 [Hig08]. 2001 [Liv02]. 2008 [Mai09]. 20th [Aga03b, SS01]. 22.87 [Big01]. 23.50 [Smi08d]. 2nd [Ric06].

3 [Fra09]. 3-03910-299-0 [Kim07]. 3-05-003770-9 [Nes06a]. 3-205-99239-3 [Har01b]. 3-406-46638-9 [Smi02b]. 3-486-56551-6 [Hop04]. 3-506-72045-7 [Har01c]. 3-515-07640-9 [Eag05a]. 3-515-07778-2 [Nes01]. 3-515-07975-0 [Ste04a]. 3-598-11603-9 [Smi06c]. 3-631-53932-0 [Far07c]. 3-7643-2434-1 [Fie05]. 3-7643-2741-3 [Hug00]. 3-7643-5312-0 [Sta03]. 3-7643-6167-0 [Ric04]. 3-7643-6280-4 [Maz02a]. 3-7643-6468-8 [Aga03b]. 3-7643-6476-9 [Set04]. 3-7643-6584-6 [Høy05]. 3-7890-5367-8 [Kra02]. 3-7965-1975-X [Smi04]. 3-8171-1630-6 [Joh02b]. 3-86135-382-2 [Har02d]. 3-86135-481-0 [Wei02b]. 3-86135-483-7 [Zig04]. 3-88570-505-2 [Sha03b]. 3-89244-629-6 [Nes06c]. 3-906757-09-9 [End04b]. 3-925347-59-3 [Hop01]. 3-928186-49-3 [Sta04]. 3-928186-51-5 [Par01b]. 3-928186-76-0 [Nes07]. 30.00 [Dew01]. 33053 [Mac05]. 34.00 [Par01b]. 35.00 [Har00a, Har00b]. 3rd [Bra00].

4-938424-73-8 [tF00]. 45.00 [Kus00, Nea00a]. 495.00 [Dea02].

50.00 [Hoc08, Hod01]. 50th [Bla03, Dje01]. 52.50 [Ril00]. 55.00 [Hug01]. 57.50 [Ril01b].

60.00 [Yam00]. 60th [Win08].

75.00 [Mac06b, Rav01].

80-200-1322-9 [Mor07c]. 80-223-1425-0 [Man01b]. 80-239-4390-1 [Mor07c]. 80.00 [Man00a]. 84-370-2349-1 [Med00]. 84-7306-626-X [Goo02d]. 84-95599-10-4 [Edg05]. 88-222-5238-1 [Pan06]. 88-430-2321-7 [Cor04b]. 88-900162-2-1 [Goo05].

9-782752-602305 [Hoc09]. 90-04-11466-1 [Gol05]. 90-04-11776-8 [Chu02]. 90-04-11942-6 [Nut05]. 90-04-11977-9 [Cuo02]. 90-04-12045-9 [Kus03]. 90-351-16941 [Rav00]. 90-420-0516-5 [Jen01]. 90-420-0526-2 [Jen01]. 90-420-0851-2 [Kno01]. 90-420-1031-2 [Ade04]. 90-420-1578-0 [App05b]. 90-420-1852-6 [Edd07a]. 90-5702-481-0 [Coo02b]. 90-5742-0341 [Mag06a]. 90-5823-057-0 [Mül01]. 90-6292-129-9 [Hig02]. 90-6984-340-4 [Gin04]. 90-6984-342-0 [Iro04]. 906292-158-2 [Hig08]. 91-628-3837-7 [Hes01]. 91-7305-147-0 [Whi05]. 91-7844300-8 [Hes01]. 92-9063-095-7 [Brö02b]. 94.00 [Chi01a]. 978-0-19-514922-7 [Rob08]. 978-0-19-726426-3 [Tur09b]. 978-0-19-920566-0 [Gor09]. 978-0-19-921461-7 [Smi09b]. 978-0-19-921543-0 [Far09a]. 978-0-19-921887-5 [Duf09]. 978-0-19-922689-4 [Gil09]. 978-0-19-923920-7 [Ful09a]. 978-0-19-925013-4 [Gou09b]. 978-0-19-929551-7 [Far09a]. 978-0-19-956625-9 [Smi09b]. 978-0-199-29663-7 [Poo09].

978-0-203-52549-8 [Smi08e]. 978-0-226-11172-8 [Mun09]. 978-0-226-11360-9 [Goo09a]. 978-0-226-13645-5 [Hou07]. 978-0-226-2-791-9 [McO09]. 978-0-226-30205-8 [Nay09]. 978-0-226-31656-7 [Ful09b]. 978-0-226-48118-0 [Bar08]. 978-0-226-60856-3 [Roo08]. 978-0-226-61668-1 [Rup08]. 978-0-226-64077-8 [Can09]. 978-0-226-70224-7 [Smi08d]. 978-0-226-70932-1 [And09]. 978-0-226-72081-4 [Ste09b]. 978-0-226-73128-5 [Coh09]. 978-0-226-73784-3 [Jar08a]. 978-0-226-74945-7 [Bax09]. 978-0-226-75024-8 [Bud09]. 978-0-226-77057-4 [Sta09b]. 978-0-226-77097-0 [Bro09]. 978-0-226-77935-5 [Web09]. 978-0-226-90405-4 [Gol09]. 978-0-22631656-7 [Ful09b]. 978-0-230-50708 [Das09]. 978-0-230-53562-6 [Sle09a]. 978-0-230-54240-2 [Rit09]. 978-0-252-07433-2 [Wil09b]. 978-0-262-02620-8 [Wee09a]. 978-0-262-08364-5 [Ric08]. 978-0-262-11306-6 [Edd08]. 978-0-262-19557-7 [Ack08]. 978-0-300-11035-7 [Stu09]. 978-0-300-11196-5 [Wee09b]. 978-0-300-11796-7 [Kno08]. 978-0-313-33053-7 [Can08b]. 978-0-313-33576-1 [Can08b]. 978-0-325-07108-4 [Chi08]. 978-0-415-96127-1 [Rob09b]. 978-0-520-24751-2 [Wil09c]. 978-0-520-25364-3 [Leh09]. 978-0-521-00903-4 [Rob09c]. 978-0-521-68957-1 [Gau08]. 978-0-521-81012-8 [Smi08f]. 978-0-521-83866-5 [Smi08b].

978-0-521-84077-4 [MW07]. 978-0-521-85575-4 [Sta08b]. 978-0-521-85822-4 [Qui08]. 978-0-521-86931-7 [Gau08]. 978-0-521-87087-0 [Edd09]. 978-0-521-87249-2 [Tur08]. 978-0-521-87559-2 [Gau09]. 978-0-5214-4502-1 [Del09]. 978-0-631-22600-0 [Shr08]. 978-0-674-02316-1 [Mer08]. 978-0-674-02321-5 [Pet08]. 978-0-674-02499-1 [MW09b]. 978-0-674-02828-9 [MW09a]. 978-0-691-12539-2 [Joh07b]. 978-0-691-12950-1 [Joh09]. 978-0-7188-3060-1 [Gil08]. 978-0-7190-6928-4 [Rut08]. 978-0-7190-7498-1 [Ric09a]. 978-0-7190-7770-8 [Jac09]. 978-0-7546-4863-5 [Gis09]. 978-0-7546-5181-9 [Nye09]. 978-0-7546-5758-3 [Dun08]. 978-0-7546-5843-6 [Mos08a]. 978-0-7546-5913-6 [Gil08]. 978-0-7546-6322-5 [Nes09]. 978-0-7546-6370-6 [Wer09]. 978-0-7546-6414-7 [Coe09]. 978-0-8018-8553-2 [Gra08a]. 978-0-8018-8599-0 [Dix09]. 978-0-8018-8645-4 [Web09]. 978-0-8018-8709-3 [Pet09]. 978-0-8018-8813-7 [Lyn09]. 978-0-8018-8825-0 [Dig09]. 978-0-8020-9359-2 [Pau09]. 978-0-8047-5901-4 [Tos09]. 978-0-8135-4220-1 [Tur09a]. 978-0-8218-4343-7 [GG09]. 978-0-8223-4068-3 [Ste09a]. 978-0-86341-823-5 [Mor09b]. 978-0-87071-196-1 [Wil09d]. 978-0-88135-395-7 [Roo08]. 978-0-88135-396-9 [Smi09a]. 978-0-906450-01-7 [Kle09]. 978-0684313207 [Sha09].

978-1-84465-086-6 [Ren09]. 978-1-84631-140-6 [Twe09]. 978-1-85196-889-3 [Del09]. 978-1-85196-906-7 [Maz09]. 978-1-85196-920-3 [Cla09]. 978-1-86189-333-8 [Stu09]. 978-1-905-472-07-9 [Jon09]. 978-1-905674-37-4 [Sle09b]. 978-1-906129-02-6 [Raj09]. 978-184-71843-2 [Pye09]. 978-184046804-5 [Dix09]. 978-1845-45210-0 [Shr09]. 978-2-296-02389-5 [Ful09a]. 978-2-8031-0231-0 [McG08]. 978-226-76329-3 [Gou09a]. 978-3-05-004496-5 [Bru09]. 978-3-7930-9448-7 [Coe09]. 978-3791336640 [Web08]. 978-3803151735 [Web08]. 978-88-339-1802-0 [Fin09]. 978-88-8292-363-1 [Pug09]. 978-9-0041-70506 [Dea09]. 978-90-04-16063-7 [Elm09]. 978-90-04-16176-4 [Smi09a]. 978-90-6984-483-1 [Mac09a]. 981-277-109-3 [Mor09b].

= [HS98].

A. [Dun08, Maz04a, Mor07a, Sta07e]. Aaboe [Mcc05]. Åkerberg [Whi05]. Aant [WM08]. Abbey [Noa09]. ABC [Coo07a, Gou03b, Sho03]. ABC-CLIO [Coo07a, Gou03b]. Abigail [Ren06]. Abingdon [Dig08, Sec08]. Abir [Ric01a]. Abir-Am [Ric01a]. abolitionism [Mil08]. Abraham [Goo02c, Smi02b, CG00a]. absolute [GO09]. Academia [Har08, Mor07c]. Academic [Can01b, Chi06, Coo02b, D’A03, DM05b, Ghi01, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Joh02d, Joh06a, Kim05, Kle04, Kni08, Loc02, Lüt04a, Man04, Mül01, Plo05, Spa05, Tso05, Web07].

Academic/Science [Coo02b]. Académie [Rud04b]. Academy [Hig05, Lea04, Mor07c, Rud04b, Sle02, Tur09b]. Accidents [TDNS08]. account [Cav08, Som04]. Accounts [Mor05b]. Accumulation [Qui08]. Accuracy [But06, Goo04b]. Accursed [Coo06a, Sil03]. achieved [Tay06]. Achilles [Col02]. Ackermann [Eag06a]. Acoustics [Pem04, Tho02a]. Across [D’A03, Cav08, Plo05, Rei07b, SS00, Sel00]. Acta [Hop01, Joh02b]. Activism [Chi06, AM01]. actor [Tho09]. acts [Cav08]. Acumen [Ren09]. Adam [Coh09, Hig06, Lyn09, Liv08, Rud08, Smi08b]. Adas [Edg08]. Addenda [Ste04a]. address [Bow06b, Bro05d, Jor00]. Adelard [Ste04a, EoToBB01]. Adler [Ben02]. Administration [Aga04, Bro03a]. Adolphe [Car01a, Pap06, Roc01]. Adrian [Ric06]. Advancement [SM05b, Wee07, WHF08]. Adventures [Bro03b, Dea05, Gre99, Key02]. Adversaries [Fre04d, Wei01]. Advertising [Mar09, Wig08]. Advice [Sch03a, Hil00, Var00]. advisory [Plo05]. aerocompass [Dav09b]. Aeronautics [Aga04]. Aesthetic [Mor06b, Her04]. Aetna [Wil09d, Tau08]. affair [Rou07]. affaire [Ful09a, Rou07]. Africa [Gre02, Luc05, Poo09, Dub06, Rel02]. Aftalion [Mor03b]. After [Aga03a, Bor03, Dol00, HH00, Kno01, Tur02b]. Afterlife [MW09b, Som07]. Aftermath [Gor09, Hen07b]. Afterword [Hal04a, Liv05b]. AG [Kim07]. Against [Pen04, Sta03, Tim04, Bar06, Ham07a, Hol00, TF02, Wei02d]. Agar [CK02]. Agatha [Goo03, Aga03a, HHAH01]. Age [Aga00, Ben06, Coh09, Dig09, Gan07, Gas98a, Gol09, Hod06b, Joh05b, Kni05, Kno08, Mac00b, Mag06b, Mor06c, Mun09, Out04, Siv08, Smi06e, Tur09a, Cla08, Coe07, DM05c, Ham08a, Hue03, Mor00, Por04, Rea05, Ric02b, RH97, Ris02, Rud04a, Rud05a, Rud05b, Rud08, Wit07, Coo07b]. Ages [Eag05a, Nut05, Kin01]. Agnesi [Pet09, Maz07b]. agnosticism [Lig02]. Agnotology [Tos09, PS08]. Agrarian [Har06]. Agricultural [Kim07, Har05b]. Agriculture [Hal04b, Har06, Bra08, Fit03, McC02]. Agustí [Kle04]. Aileen [Ric09b, Tur06b]. Air [Cla04b, Atm04, Bru08, CKM99]. air-pump [Bru08]. Akademie [Bru09, Gin04, Iro04, Mac09a, Nes06a]. Akihito [Dig08]. al [EB07, Fox05, Hod06a, Med00]. al-Furqan [EB07]. Alain [Fyf02b, Kem09]. Alan [Ful09b, Sle02, Car01a, Ful09a, Mar05]. Albany [Mor08]. Albert [Hen05b, Rud04b, Jam01a]. Albin [NG01b]. Alchemical [Smi02b]. Alchemie [Smi02b, Sch00]. Alchemiehandbuch [Smi04]. Alchemist [Edd05b, Tra07, New03]. alchemistischen [Smi06c]. Alchemy [Edd05b, Hen07a, Kni02, NP02, Num07, Pag03, Roo08, Smi05b, Smi09a, Lev01, Lin03, Mor05a, NG01a, Pri07, Sch00]. Alder [Goo04a]. Alderhot [Nes09]. Aldershot [Alb01, Alb08, Bar07, Bek06, Bre04, Bre06, Bro05b, Bro03c, Bro01b, Bro04c, Cam08, Cla04a, Cla07a, Coe09, Coo03, Coo06b, Dou06, Dun08, Eag08, Edd03b, Edd05a, End04a, Far07a, Fyf02b, Fyf03b, Gil08, Gis09, Gou03a, Gou06, Han07b, Hen01, Hig07, Hob06, Kem09, Kus00, Luc05, Mor06c, Mor06d, Mor06e, Mos08a, Nye09, Pic07, Ril00, Ril01b, Rit02, Rob07, Rud07, Sam07, Sch02a, Siv06, Smi06d, Ste05b, Wee07, Wer09, Wil09a]. Aleksandr [Yam00]. Alessandro [Pug09]. Alex [Hem02b, Big06, Hay06, Sum06]. Alexander [Eag08, Far07c, Ste05a, Tyb04]. Alexandria [Høy01, Cuo00]. Alexis [Sel08]. Alfred [Can08a, Coo06a, Gar03b, Ghi01, GG06a, Sim04, Web06, Coo05b, Fic04, Jon02b, Slo04, WC02, WB02]. Algebra [GG09, GP07].

Alice [Sta03, Twe09]. Alisa [Sir06]. Alison [Mor06d, Kha06a]. Alix [Edd09]. Allan [Goo06b, Bro06a, Sta07a]. Allard [McG08]. Allen [Chi06, End04a, Goo03, Hen07a, Tim08, Deb06, Hig04c, Tim04, Tim08]. Allergies [Stu09, Mit07]. Allergy [Stu09]. allgemeinen [App05b]. Almanacs [Rut08]. Alpes [Rud04b]. alten [Man01b]. Alter [Rad00a]. Alternative [Biv07, Duf09]. Alternatives [Vic07, Sta06a]. Altruism [Tur09b, Dix08b]. Am [Ric01a, Far07c, Joh02b]. amateur [Hoc09]. amateurs [Ogi00, Ver03]. Amazing [Ber08, Del06b]. Ambitions [Bre06, Hen04b, Nut04, Dea01, RH04]. Ambix [Hen07a]. America [Ber08, Boo07, Edd03b, Edg08, Fyf02a, Fyf03b, Goo06a, Jon02a, Jon04, Lin07, Log08, Mor01b, Pem04, Ast01, Bar00b, Cow01, Del06b, Gol04, Hou06, Mur00, Nor07, Tho02a]. American [Aga02, Aga04, Col00, Daw06, Dol01, Edd05b, Goo02c, GG06b, GG09, Har06, Hug08, Jam01a, Kle05, Lin05, Pet08, Ree08, Tim03b, Wal03, Web07, Bek01, Ben04a, Car07, DeV00, Fit03, Igo07, Jef01, Joh02a, Kin05, Kri06, Moy97, New03, Swe01, Zen98]. Amherst [Cla04b]. Amir [Ste05a]. Amiria [Urr07]. Amsteldijk [Mül01]. Amsterdam [Ade04, App05b, Coo02b, Edd07a, Gin04, Iro04, Jen01, Kno01, Mac09a, Mag06a, Maz04a, Pet07a, Rav00, Ver03]. Amsterdamska [Ric08]. Amundson [Bow07]. amusing [Som04]. analogical [Ver05]. analogy [Smi08g, Smi08h]. Analysis [AH02, GG06b, Sim02, Sta09a, FK06, Hee08, Jah03, Mcg06]. Anatomical [Bro02e, Daw00, Car99, Mar09]. Ancestors [Goo09b, Lyn09, Liv08]. Ancient [Cam08, Gau08, Hsu06, Llo04, Nut04, Tot04, Wil09d, Gra07b, Llo06, Sas01, Tau08]. Anders [Par01a]. Andersen [Sta08b]. Anderson [Edd07b, Led08]. André [Hoc09, McG08]. Andrea [Bro02e, Daw00, Dep07]. Andreas [Roo08, Hop04]. Andrew [Aga05, Goo02b, Hun05a, Lea05, Ver04, Fin03a, Fyf02b, Gar03b, Gra05, Hug07a, Kas02, Tyb02]. Angel [Man00b]. Angela [Kir05, Ste09a]. Angeles [Han07a, Mar05, Rob05, Ste07, Ste04b]. Angelique [Wal06b]. Angerstein [Dol02a, ABB01]. Angus [Har03a]. Animal [Elt01, Hal02a, Kir05, MW07, Mun08, Smi08d, Bla07, CJ02, Rad07, Smi06b]. animal-human [CJ02]. Animal/Human [Kir05]. Animals [Alb03b, Mun07, Rit02, Spa03, Whi06, Asm01, DM05a, KJ01, Rob02]. Anita [Fri03a, GG06a]. Anker [Hal03]. Ankete [Yeo04]. Ann [Eds04, Jam03, Sta03, Yeo04, Goo09b, Jor08]. Anna [Ben02, Smi09a, Mac06b]. Annals [Tim03c, Kre02]. Anne [Biv08, Bra00, Daw00]. Annette [Kus03]. Annie [Ske04, Key01, Ogi00]. Anniversary [Win08]. Annotated [Gol05, Tyb04, BJ00, Gin02]. Answers [Sta07a]. Antarctica [TDNS08]. Anthology [Bro05a, Gar03b, Mor09b, Sta03, Gat02, Hen96, J^{+}07, WB02]. Anthony [Eag06a, Pag03, Van02b]. anthropological [Kuk06]. anthropologists [Cha02b, GO09]. Anthropologues [Cor05, Cha02b]. Anthropology [Mar05, Shr08, Urr07, GHL03, Hen05a, Kuk08]. Anthropomorphism [Mun07, DM05a]. Anti [Can07, Tim03a, Tim03d, Wor07, Dur05]. Anti-Cancer [Tim03a, Tim03d]. Anti-Semitism [Can07]. Anti-vaccination [Wor07, Dur05]. Antievolutionism [NS00]. Antipodes [NS00]. Antiquaries [Goo06c, Swe04]. Antiquity [ETV04, Høy01, Leh09, Rob09c, Cuo00, Cuo07, Lon01, Sed07]. antisemitism [Fro05]. Antoine [Kni03, Ber01]. Antonio [Lüt04a]. Antony [Joh04b]. Apart [Kra08a, Kai05b]. Apeiron [Hen04e]. Apollo [Aga04, Joh02a]. Apollonius [Cuo02, FU01]. Apparatus [Ste04a]. Appenzeller [Smi04]. Application [Hun02]. Applied [Coo02b, BG02]. Apprenticeship [Shr09]. Approach [Aga03a, Goo08a, Ara06, HH00]. Approaches [Coo07a]. Appropriation [Pal06, Bow03]. Ära [Man01b]. Arabatzis [Goo08a]. Arabian [Whi06, Der03]. Arabic [Sam07, Fre05, Mcg06]. Arago [Lev00]. Aratus [Har01a, Gee00]. Arbor [Eds04]. Archaeologists [Sho03, Mur99]. Archaeology [Can04a, Div01, Sho03, Cos00, Goo02e, Hay02a, Mur99]. Archer [Lev07]. Archimedes [Ber06b, Ber06a, Gra06, Hem02b, Loc02, oAEN04]. Architectural [Pem04, Tho02a]. Architecture [For03, Ren05, Ste00b]. archival [Elz06]. Archives [Big01, Cla04a, DM05b, For05, Kim05, Lüt04a, Man04, Maz07a, Mos08b, Spa05, RP02b]. Arctic [Haf04, BS02]. Areson [Dig09]. argument [Bul04]. arguments [Man03]. Arikha [Bro02e]. Aristides [Ful02]. Aristotle [Dep07, Leh08, Fal05, Joh05a]. Armour [Wer04]. Arms [Bas04b, Coo06b, Wer04]. Armstrong [Coo06b, Bas04b]. Arne [Fyf03a]. Arnold [Alb08, Cha05, Siv08, Sta04]. Arnott [Far07a]. Arrangement [Mor06e]. Art [Ade05, Brö02b, Eag06b, For03, For04, Hen04a, Hen06b, Jor02, Jor06, Kem09, Sch02a, Web08, Das04, Giv05, HW01a, NFJ97]. Artefacts [Mül01, Goo02e]. Artefects [Coo06c]. Arthur [Bru06, Sta03, Mus09, Sta07b]. Artibus [Hig05]. Artificial [Cop04, Ste09b, Wee09a, BVN07, Ris07]. Artigas [Ade08, Hei08]. Artisan [Hen06b, Twe09]. artist [HP02]. Artists [Jor02, Ham01a]. Arts [ETV04, Hug01, Joh07a, Jor01, Kha06b, Cro01c, Lon01, Mor97a, UU07]. Asger [Mcc05]. Ash [Tre06]. Ashbourne [Coo04, Mil04]. Ashgate [Alb01, Alb08, Bar07, Bek06, Bre04, Bre06, Bro05b, Bro01b, Bro04c, Cam08, Cla04a, Cla07a, Coe09, Coo03, Coo06b, Dou06, Dun08, Eag08, Edd03b, Edd05a, End04a, Far07a, Fyf02b, Fyf03b, Gil08, Gis09, Gou03a, Gou06, Han07b, Hen01, Hig07, Hob06, Kem09, Kus00, Luc05, Mor06c, Mor06d, Mor06e, Mos08a, Nes09, Nye09, Pic07, Ril00, Ril01b, Rit02, Rob07, Rud07, Sam07, Sch02a, Siv06, Smi06d, Ste05b, Wee07, Wer09, Wil09a]. Ashworth [Sum05]. Asia [Das09]. Askant [Mag06a]. Asma [Alb03b]. Aspects [Cha01, Fet00]. assistance [Jam05]. Assistant [Sta03]. associate [Mac01a, Mül01]. Associated [Ben06, Man06, BRW04]. Association [Chi04b, Jon04, Lin05, Mac05, WHF08, Bow06a, Bro03c, Kha06b, Nea00a]. Assumptions [Cla09]. Astore [Fyf03b]. Astrolabe [Eag05b, CE02]. Astrolabes [Hig07, vCA^{+}05, Tur03b]. Astrologer [Tra07]. Astrological [Pag01, Tra01]. Astrology [Pag03, Rut08, Har00c, NG01a]. Astronomer [Aga01c, Nea00b, Nea00a, Moo99, Hom09, Wil97]. Astronomers [Aga02, Loc02, DeV00, MCS09, Ogi00, Ste00a]. Astronomiae [Joh02b]. Astronomical [Big01, Smi08b, Mos07, Mus09]. Astronomico [Big01]. Astronomy [App05b, Ben02, Big07, CG00a, D’A03, Gin04, Goo02c, Har01a, Hes01, Joh02b, Mcc05, SS00, Smi02a, Aab01, BZ04, Can01a, Gee00, Har00c, Hei05, Hof07, Hol99, KD00, Ogi00, Rei05a, Sta01a, Ver02, Rei05a]. astrophysical [Sta07f]. Asylum [For03, Tho03a, Ste00b, Wri01]. Athanasius [Gor02, Row00]. Athlone [For04, Hen04a, Yos00]. Atlanta [Jen01, Kno01]. Atlantic [Rob06, Rob09b, Bai05, DD08, Osl02, Sch04a]. Atmography [Gol01]. Atmospheric [GR00, Hem02b]. Atomic [Tur03a, Her06]. Atomism [Lüt04a, Cle00, Sta03]. Atoms [Kle05, Sie02]. AU$49.95 [Pre03]. Aube [Hoc09]. Audiences [Bar08, Par01a, Lig07, LBV00]. Audion [Hem03, Hon01]. Aufklärung [Nes06c, Hoc03]. Augmentés [Rud04b]. Augstein [Smi00a]. Augustí [Edg05]. Augustin [Coo05a]. Augustin-Pyramus [Coo05a]. Augustus [Har01a, Gee00]. Aurum [Smi07a]. Austin [Hal04b]. Australia [Bar01b, Wal06c, Dun99]. Australians [Kuk06]. Author [Can02b]. Authority [Far06, Roo08, Wil09a, Cro01c, DV04, McT05, Num07, Sta01b, Van02a]. authors [Ano09o, Joh00]. Authorship [ETV04, Top02, Lon01, Sec00]. Autobiographical [Bar01a, KCH00]. Autobiographies [Ash04b, HH03]. Autobiography [Jen02, NM00]. autour [Chi01b, Hul00, Rud04b]. Available [Mac09a]. avancement [Chi04b]. Averaged [Pet08, Igo07]. Avicenna [McG08]. Avicennan [Mcg06]. AVISTA [Kem09]. Axel [Joh02b]. Axiomatization [Cor04a, Gra06, Smi08g, Smi08h]. Ayval [DM05b].

B [Aga04, Bax09, Coo06a, Dea02, Edd07a, Fin03a, Jor08, Liv05a, Ner05, Pau09, Pet07a, Rud04b, Sil03, Spa06, Ste04b, Tim04]. B. [Hod06b, Mül01]. Babbage [Cro04]. baby [Sch05b]. Bachelard [Til04]. Back [Ano09a, Ano09c, Ano09e, Ano09g, Chi02a, Coo02a, Cou04, Fre04b, Sen02a, Shi03a, Alb08, Arn06b]. Bacon [Dea09, Hen02a, Man00a, Man01a, Man06, Sha02, Wee07, Gau01, Man03, SM05b, Zag01]. Baconian [Hun07b]. Bad [Boo07]. Baden [Kra02]. Baden-Baden [Kra02]. bag [Dea02]. Bailey [Dou06]. Bailyn [Rob06]. Baird [Emm01, Goo06b, Joh02c, Joh04b, Bur00, KB02, McL00]. Bakker [Rav00]. Ball [Mor09b, Hom09]. Ballantyne [Siv06].

Baltas [Ful02]. Baltimore [Ade08, Aga04, Bur08, Chi02b, Dea08, Dig09, Dix09, ETV04, Gar03b, Gra08a, Gra08b, Ham07a, Hem02a, Kni02, Kni04, Lyn09, Pet09, Pro08, Rad00a, Sle01, Spa03, Tim03b, Tre06, Web07, Web09, Whi06]. Band [Smi06c, Sta04, Wei02b]. Banks [Mac00b, Mil02, BC00, Gas98a]. Barbara [Coo07a, Gou03b, Sho03, App05a, Bro05a, Pag01]. Barber [Jor02]. Barcelona [Fox05, Goo02d, NG04]. Barcelonesa [Fox05]. Bardeen [Hes04, HD02]. Barker [Sta08b, Tyb02]. Barlow [Hig04b]. Barnes [Gis09, Ham07a, Maz08, Val05]. barometers [Dry09]. Baroque [Gor02, Row00]. barrel [Sum01]. Barry [Gis09, Maz08]. Bart [Hig08, Mac08b]. barter [End00]. Bartky [Mor01b]. Bartusiak [Sta03]. based [Loc07]. Basel [Aga03b, Fie05, Gol05, Høy05, Hug00, Maz02a, Ric04, Set04, Smi04, Sta03, Gin02]. Bashford [Kha06a]. Basic [Sta03]. Basil [Hun06]. Basilisken [Hop01]. Basilisken-Presse [Hop01]. Basingstoke [Can07, Edg04, Hen04c, Hen04b, Hig03a, Kha06a, Led08, Rit09, Sle09a, Smi08e]. Bastable [Coo06b]. battery [UU07]. Battle [Car01a, Tim04, Roc01, Wei02d]. Bauer [Dol02b]. Beach [Haf08, Hei08, Joh07a, Nay09, Ras06, Roo08, Smi09a, Tim08, WM08]. Beagle [Dea05, Ric01b, Key00, Key02]. Bearing [Mos07, Smi08b]. Beatrix [Web07]. Beauchamp [Hun02]. Beautiful [Bro03c, HH01]. Beauty [Nes06b, Sta03, Hau03, Mil01]. Became [Gor07, GG03a, Sta06b, Mor05d, Wei02c]. Beck [Smi02b]. Becker [Smi08f]. Becoming [Jon02a, Mur00]. Bednarek [Rol06]. Bee [Mun08]. Beer [Sle05, Tei00]. Before [Can02b, Coh09, Hol08, Hor07, Ken07, Rud08]. Beginning [Bax09, Sha08a]. Beginnings [Gou06, KG04]. Behavior [Dew07, Bur05]. Behind [Ben04b, BI06, Hei08, Mac02, Elz06, WM08]. Being [Ric09a]. beings [Nor07]. Beisner [Web07]. Beitrag [Har01b, Tei00]. Beiträge [Hör78]. belated [Edd04a]. Belief [Hei08, Mor02c, Jac00a, May05a, SA06]. Beliefs [Bar07, KE05]. Belknap [Edg08, Haf08, Mer08]. Ben [Rei07a]. Bender [Daw06]. Benedict [App05a, Rud04b]. Benedikt [Rei07b]. benefit [RN09]. Benfey [Mor03b]. Benjamin [Bra07, Gou09a, Hen05b, Bra07, GG04]. Bensaude [Chi01a, Chi03b, Par01a, Vid05, Wee09a, Wer09]. Bensaude-Vincent [Chi01a, Chi03b, Par01a, Vid05, Wee09a, Wer09]. Beretta [Joh07a, Kni03, Pan06]. Berg [Ben04b, Dol02a, Har04c]. Berggren [Tyb04]. Berghahn [Shr09]. Berit [Cla04a]. Berkel [Rav00]. Berkeley [Bra00, Dig08, Goo06b, Han07a, Hug07b, Kra07, Leh09, Mar05, Rob05, Seg05, Ste07, Ste04b, Wil09c]. Berlin [Aga03b, Bru09, Fie05, Gos00, Har02d, Høy05, Hug00, Maz02a, Nes01, Nes06a, Par01b, Ric04, Set04, Sta03, Sta04, Web08, Wei02b, Zig04, Gos00]. Berman [Nor07]. Bern [Alb03a, End04b, Kim07]. Bernadette [Par01a, Chi01a, Chi03b, Vid05, Wee09a, Wer09]. Bernal [Hug07a, Bro05c]. Bernard [Bra00, Goo02c, Hen05b, Jor08, Maz04a, Ric09b, Rob06, Bar08, Coo03, Mor06a]. Bernardi [Vid05]. Berry [Gar03b]. Bert [Rav00, Daw06]. Bertoloni [Dea08]. Bertucci [Fin09, Goo05]. Bessel [Hof07]. Besson [Fyf02b]. Betrayal [Sta07b, Mil05]. Better [Bur08, Sle07]. Between [Bou02, Shr09, Aga03b, Bra08, Gui04, Har05b, Kim07, Kra07, Mae06, SS01]. Beurton [Olb02]. Beyond [Del09, Ful09a, Pye09, Sok08, GHL03, Mar05, Maz07a]. Bezzola [App05b]. Biagioli [She07]. bible [Edd04a]. biblical [Liv03]. Bibliographia [Med00]. Bibliographic [Lin05]. Bibliographie [Smi06c]. Bibliography [Fyf02b, Lin05, TBbTc90]. Bibliopolis [Man00b]. Biblioteca [Pan06]. Bibliothèque [Coo05a, Cor05, Rud04b]. Bicks [Hob06]. Bier [Har01b, Tei00]. Bild [BSD08, Bru09]. Bilder [Bru09, BSD08]. Bildung [Hop04, Zig04]. Bill [Dig08]. Bindman [Jor02]. Biodiversity [Joh07b]. biografie [Rav00]. Biographical [Goo08a, Gou03b, Ara06, Hai01]. Biographies [And03a, Boh01, Das00, Dix06, Maz07a, IH05]. Biography [Gli06, Kai05a, Mac08b, Maz07a, Maz09, Nye09, Sha09, Söd07, Bro02d, MM00, Sch04b]. Biohistorica [Hop01]. Biological [Mac08c, Wag05]. Biologie [Har02d]. Biology [Bow07, Ghi01, Hag04, MW07, Wag05, BHR00, CW02, GWT00]. Bioprospecting [Rob06, Sch04a]. Biosciences [Her05b]. Biotech [Qui08]. Biotechnology [Pal06, Bow03]. Biowissenschaften [Wei02b, Zig04]. Bipartisan [Wig08]. Birgit [Bru09]. Birkhauser [Hug00, Aga03b, Fie05, Høy05, Maz02a, Ric04, Set04, Sta03]. Birmingham [Jac09, Jor02, Jon08]. Birstein [Chi04a]. Birth [Aga00, Hod06b, Hun05a, Wei02a, BW01, Fau03, Gas98b, Ili00, RH97]. Birthday [Bla03, Dje01]. Birthing [DiM07, Zin05]. bis [BHR00, Har02d, Smi06c]. Bitter [Stu06]. Bivins [Duf09, Smi08e]. BJH [Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h]. BJHS [AAH00a, Ano00a, Top00]. Black [Dix09, Hem03, Sta07b, Com07, Hon01, Mil05]. Black-Box [Hem03, Hon01]. Blackett [Mac07, Nye04]. Blackwell [Bee01, Hei08, Shr08]. Blondel [Chi03b, Wer09]. Blood [Lee02, Sen06a]. Boas [Lüt04b]. Bob [Cla05]. Boden [Abr08]. Bodies [Bro02e, Goo05, Smi06e, BP01, Car99]. Bodily [Dur05, Wor07]. Body [Daw00, Hen06b, Mor08, Qui04, Che06, Mus06, Sil01].

Boerhaave [Hig02, Hig08, Iro04, Kno02]. Boethius [Eag05a, Ste04a]. Böhlau [Har01b]. Bohr [Kra07]. Bois [Fin03b]. Bois-Reymond [Fin03b]. Bollati [Fin09]. Bologna [Goo05, Kni03]. Bolt [Ben02]. bomb [Hug04a]. Bonaparte [Cor05, Cha02b]. Boner [Hig06]. Bones [MW09b, Som07, Tho09]. Book [Abr08, Ack08, Ade04, Ade06, Ade08, Aga00, Aga01a, Aga01b, Aga01c, Aga02, Aga03a, Aga03b, Aga04, Aga05, Aga06, Aga08a, Alb01, Alb03a, Alb03b, Alb08, And03a, And09, App05a, App05b, Art04, Ash01, Ash04b, Aug09, Bar01a, Bar01b, Bar07, Bar08, Bax09, Bee01, Bek06, Ben00, Ben02, Ben04b, Ben06, Ber06b, Ber06a, Ber08, Big01, Big06, Big07, Biv06, Biv08, Bla03, Boh01, Boo07, Bor03, Bow04, Bow06a, Bow07, Bra00, Bra07, Bre04, Bre05, Bre06, Bro04a, Bro05a, Bro05b, Bro06a, Brö02b, Brö04b, Bro08b, Bro09, Bro03b, Bro03c, Bro01b, Bro02e, Bro04c, Bru06, Bru09, Bud09, Bur07a, Bur08, But06, CK02, Cam08, Can09, Can08a, Can01b, Can02b, Can04a, Can07, Can08b, Can04c, Car01a, Car06, Cha05, Cha01, Cha03, Chi08, Chi02b, Chi04a, Chi06, Chi01a]. Book [Chi01b, Chi03b, Chi03c, Chi04b, Chr05, Chu02, Cla01, Cla04a, Cla05, Cla06, Cla07a, Cla09, Cla04b, Coe09, Coh09, Col00, Col01b, Col02, Coo05a, Coo07a, Coo03, Coo04, Coo06a, Coo06b, Coo06c, Coo02b, Cor00, Cor04b, Cor05, Cor06, Cro04, Cro05a, Cro05b, Cuo02, D’A03, Dan01, Das07, Das09, Daw00, Daw06, Daw07a, De 07, Dea05, Dea02, Dea06a, Dea07, Dea08, Dea09, Del06a, Del08, Del09, DM05b, Dep07, Dew01, Dew07, DiM07, Dig08, Dig09, Div01, Dix00, Dix03a, Dix05, Dix06, Dix08a, Dix09, Dod02, Dod06, Dol01, Dol02a, Dol02b, Dol02c, Dol04a, Dol04b, Dol04c, Doo04, Doo05, Dou06, Dri03, Duf09, Dun08, ETV04, Eag05a, Eag05b, Eag06a, Eag06b, Eag07, Eag08, Edd03a, Edd03b, Edd04a, Edd05a, Edd05b, Edd06, Edd07a, Edd07b, Edd08, Edd09, Edg04, Edg05, Edg08]. Book [Eds04, EB07, Elm09, Elm02, Emm01, End04a, End04b, tF00, Fan02, Far07a, Far07b, Far07c, Far09a, Far06, Fie05, Fin01, Fin03a, Fin09, Fin02, Fis04, For03, For04, For05, Fox04, Fox05, Fox07, Fra09, Fra05, Fre04c, Fre04d, Fri03a, Fri03b, Fri04, Ful01a, Ful01b, Ful01c, Ful02, Ful03, Ful09a, Ful09b, Fyf02b, Fyf02a, Fyf03a, Fyf03b, Fyf05, Fyf07, Gan07, Gar03b, Gat05, Gau08, Gau09, Ghi01, Gil08, Gil09, Gin04, Gis09, Gli06, Gol05, Gol09, Goo01, Goo02a, Goo02b, Goo03, Goo04a, Goo05, Goo06a, Goo08a, Goo09a, Goo06b, Goo02c, Goo02d, Goo06c, Goo08b, Goo09b, Gor07, Gor09, Gor02, Gou03a, Gou06, Gou09a, Gou09b, Gou03b, Gou03c, Gra08a, GG02, GG03b, GG03a, GG04, GG06a, GG06b, GG09, Gra05, Gra06, Gra08b, Gre02, Gui04, Gui05]. Book [Haf04, Haf08, Hag04, Hal02a, Hal03, Hal04a, Hal04b, Hal06, Ham07a, Ham08b, Ham07b, Han07a, Han07b, Har01a, Har00a, Har00b, Har05a, Har03a, Har02c, Har01b, Har01c, Har02d, Har03c, Har04c, Har06, Har08, Hay02b, Hay03, Hay06, Hei07b, Hei07a, Hei08, Hem02a, Hem02b, Hem03, Hen01, Hen02a, Hen03a, Hen04c, Hen04b, Hen04a, Hen05b, Hen06a, Hen06b, Hen07a, Hen03b, Hen04d, Hen04e, Her05b, Hes01, Hes03, Hes04, Hig04a, Hig03a, Hig01, Hig02, Hig04b, Hig05, Hig06, Hig07, Hig08, Hob06, Hoc08, Hoc09, Hod01, Hod06a, Hod06b, Hop01, Hop04, Hor03, Hou07, Høy01, Høy05, Hsu06, Hug00, Hug01, Hug07a, Hug07b, Hug08, Hug09, Hul07, Hun02, Hun05a, Hun05b, Hun06, Hun07a, Hur07, Hut04, Hut07, Hut09, Iro04, Jac09, Jam01a, Jam01b, Jam03, Jam05, Jar08a, Jen01]. Book [Jen02, Joh03a, Joh07a, Joh07b, Joh09, Joh02b, Joh02c, Joh02d, Joh03b, Joh04b, Joh04c, Joh05b, Joh05c, Joh06a, Joh06c, Jon02a, Jon04, Jon06a, Jon09, Jor01, Jor02, Jor06, Jor08, Kai05a, Kas02, Kem09, Kha06a, Kha06b, Kim05, Kim07, Kip04, Kir05, Kle04, Kle05, Kle09, Kni02, Kni03, Kni04, Kni05, Kni06a, Kni08, Kno08, Kno01, Kra02, Kra07, Kra08a, Kre06, Kus00, Kus03, Kus07, Lam02, Lau06, Lea04, Lea05, Led08, Leh08, Leh09, Lev07, Lin05, Lin07, Lip08, Liv05a, Loc02, Log08, Luc05, Lüt04a, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac08a, Mac09b, Mac09a, Mac01a, Mac05, Mac06b, Mac07, Mac08c, Mag06a, Mag06b, Mai09, Man04, Man00a, Man00b, Man01a, Man01b, Man06, Mar05, MW09a, Maz04a, Maz02b, Maz02a, Maz07a, Maz09, McG08, McO09]. Book [Mcc05, Med00, Mel01, Mer08, Mil02, Mil04, Mir07, Mok06, Mor03a, Mor05b, Mor06a, Mor06c, Mor06b, Mor07a, Mor07b, Mor07c, Mor01a, Mor02a, Mor03b, Mor06d, Mor06e, Mor01b, Mor01c, Mor02c, Mor08, Mor09b, Mos08a, Mos08b, MW07, MW09b, Mül01, Mun07, Mun08, Mun09, Mur02, Mus03, Nay09, Nea00b, Nea00a, Ner05, Nes01, Nes02, Nes06a, Nes06b, Nes06c, Nes07, Nes09, NG01b, Noa04, Noa09, Nut04, Nut05, Nye09, Olb02, Out01, Out04, Pad01, Pag01, Pag03, Pal00, Pal02, Pal06, Pan06, Pan08, Par01a, Par01b, Pau09, Pem04, Pem07, Pen04, Pet07a, Pet07b, Pet09, Pet08, Phi02, Pic07, Plo05, Poo09, Por01, Por03, Pre03, Pro08, Pug09, Pye06, Pye09, Qui02, Qui04, Qui08, Rad00a, Rad02, Rad03b, Raj09, Ram02, Ras06, Rav00, Rav01, Ree08, Ree07]. Book [Rei07a, Rei07b, Ren00, Ren06, Ren07, Ren09, Ric04, Ric06, Ric09a, Ric09b, Ric01a, Ric01b, Ric08, Ril00, Ril01a, Ril01b, Ril02, Ris01, Rit02, Rit09, Rob03, Rob05, Rob07, Rob06, Rob08, Rob09b, Rob09c, Rol06, Roo08, Rud04b, Rud07, Rup08, Rut08, Sam07, Sch08a, Sch03a, Sch02a, Sch01, Sch02b, Sec08, Seg05, Sen02b, Sen06a, Set04, Sha02, Sha03b, Sha07, Sha09, She07, Shi03b, Sho03, Shr08, Shr09, Sir06, Sir09, Siv06, Siv08, Ske04, Sle01, Sle02, Sle03, Sle05, Sle09a, Sle09b, Smi00a, Smi01b, Smi02a, Smi02b, Smi03b, Smi03a, Smi04, Smi05a, Smi05b, Smi06d, Smi06e, Smi06f, Smi06g, Smi06c, Smi07a, Smi07b, Smi08b, Smi08a, Smi08d, Smi08f, Smi08e, Smi08c, Smi09b, Smi09a, Smo08, Spa02, Spa03, Spa05, Spa06, Spr06, Sta03, Sta04, Sta07a, Sta09a, Sta08b]. Book [Sta06b, Sta06c, Sta07b, Sta07c, Sta07d, Sta09b, Ste04a, Ste05a, Ste05b, Ste09a, Ste09b, Ste07, Ste04b, Ste06, Stu06, Stu09, Sum05, Sum06, Sut03, Suz05, Ter08, Tho03a, Tho02b, Tho03b, Tho03c, Til04, Tim02, Tim03a, Tim03b, Tim03c, Tim04, Tim05, Tim08, Top02, Top08, Tos09, Tot04, Tra07, Tre06, Tso05, Tso07, Tur02a, Tur06a, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Van02b, Ver04, Vic07, Vid05, Wag05, Wal03, Wal06b, Wal06c, WM08, War08, Web06, Web07, Web08, Web09, Wee07, Wee08, Wee09a, Wee09b, Wei02a, Wei02b, Wer04, Wer09, Whi05, Whi06, Whi04, Wil06, Wil09a, Wil09d, Wil09c, Wil09b, Win08, Wor07, Yam00, Yam04a, Yam04b, Yeo04, Yos00, Zig04, dC08, EG04, How00, Loz00, TBbTc90]. book [Top00]. Books [Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ber06a, Boo07, Bro03b, CK02, Daw00, Dix09, Dri03, Far07b, Fin01, FSJ00, Ful01c, Fyf02b, Gar03b, Ham07b, Jor06, Mor09b, Mun08, Sta03, Stu09, Tur02a, Por01]. Books/MIT [Jor06]. Boon [Sle09b]. Borderlands [Dol02b, She01]. Borders [Pye09]. Boringhieri [Fin09]. Boris [Hig06, Yam00]. Born [Aga01b, GC00]. Bosch [Fin02, Smi01a]. Boston [Aga03b, Chi06, DM05b, Elm09, Fie05, Ghi01, Gol05, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Høy05, Hug00, Joh02d, Joh06a, Kim05, Kle04, Lüt04a, Man04, Maz02a, Maz07a, Ric04, Set04, Smi09a, Spa05, Sta03, Sta07b, Tso05]. Botanic [Raj09]. Botanical [Raj09, End00, Nol07]. Botany [Del06a, Spa05, SS05, Wil03]. Bothered [Log08, Hou06]. Boudri [Gui04]. Bought [Por03, MS02]. Boulder [Chi04a, Shi03b]. Boundary [Kir05, CJ02]. Bowden [Mor07a]. Bowker [Smo08]. Bowler [Hei07b]. Bowring [Pal06]. Box [Dix09, Hem03, Ske04, Yeo04, Com07, Hon01, Key01, Sch05b, tH02]. boxed [Raj09]. Boyd [Can04a]. Boydell [Can04a, Elm02, Maz07a, Nea00a]. Boyle [Dea02, Edd05b, Dea02, HD00, Hun07b, NP02]. Bragg [Mor09b]. Brahe [Man00b, Smi08b, Mos07]. Brain [Cro04, Dan01, Van05, Jac00b]. Brains [Bor03, Tur02b]. Brains/Practices/Relativism [Bor03, Tur02b]. Bratislava [Man01b]. Bravo [Haf04]. Bray [Elm09]. Breathing [Mit07, Stu09]. Bred [Der03, Whi06]. Bredekamp [Bru09, Web08]. Breeding [Har03c, WO01]. Breidbach [Her05b, Web08]. Brendan [Lea04]. Brepols [Hig05]. Brey [Aga05]. Brian [Hen05b, Kno01]. Bridget [Can04c]. Brief [Suz05, Por02]. Briefwechsel [Sta04, SEM04]. Bright [Ade06]. Brill [Chu02, Cuo02, Dea09, Elm09, Gol05, Kus03, Nut05, Smi09a]. Bristol [Aga01c, Ben00, Bro06a, Kre06, Lea05, Maz07a, Mor06a, Nea00b, Rad03b, Rob03, Smi00a]. Britain [Can09, Cla07a, Coo02b, Dix03a, Edd03b, End04a, Fyf03b, Fyf07, Gil08, Goo06c, Hou07, Kip04, Mac09b, Mac05, Maz07a, Mor08, Mur02, Rei07a, Ril00, Rit09, Sch02b, Sta07c, Tur06b, Tur09b, WHF08, Yos00, AA01, Ast01, Bow06b, BG02, BLR01, Can05, Car01b, Cow01, Cro07, Dar05, Dau05, Dav09b, Dix08b, Dra00, Edg06, Fre01, Fyf04, GS02, Hor00, Hor07, Mar04, MS05b, Mor97b, Mor00, Mus06, Ott08, Sch06a, Swe04, Tay04, Wor00]. British [Bre04, Bro03a, Cha05, Cla09, Coo06b, Cor00, Del09, Dix03a, Dod06, Edd05a, For03, Hal03, Hay06, Jon09, Jor02, Kle09, Mac00b, Mok06, Mor05b, Mor06a, Mor07b, Mor01a, Nay09, Smi01b, Tur09b, Ank01, Ano00e, Bas04b, Edn97, tF04, Gas98a, Gol07, Har00c, Hig03b, Mco01, Ogi00, Ryl00, Sch05b, Ste00b, Ste09c, Tor02, WHF08, Yeo01b]. Broadcast [Hem02a, Slo00]. Brock [Nes09]. Brockliss [Liv05a]. Broks [Top08]. Broman [Bre05]. Brömer [Har02d]. Brooke [Kus07]. Brooks [Lin05]. Brotóns [Goo02d]. Brouwer [Rav01, vD05]. Brown [Cro04, Ful03, Goo04a, Hes03, Hug07a]. Browne [Gli06]. Bruce [Roo08, Ben02, Hen07a]. Brunel [Buc02, Har03a, Buc02, Mil08]. Brüning [Smi06c]. Bruno [Man00b, Vid05]. Brunswick [Goo02a, Hay03, Hig04a, Lin07, Tim04, Tur09a]. Brutes [Rit02, KJ01]. Bryce [Mor02a]. BSHS [Jor02, Kle09, Mor05b]. Buchanan [Har03a]. Buchdruckerkunst [Smi06c]. Buchwald [Hun05a]. Buckland [Som04]. Bucknell [Ben06]. Buckyball [Kni02, Lev01]. Bud [Coo02b, Mül01]. Budapest [Vid05]. Buffalo [Jen02, Mel01, Pau09]. Buffon [Spa02, Lov01]. Build [Cro04]. Building [Joh04c, Mor02b]. Built [Cla07a, Joh06c, Hug04b, Tay04]. Buisseret [Rob08]. Buklijas [Mai09]. Bunn [Dix03a]. Burden [Goo02a]. Burdman [GG06a]. Burek [Mor07a]. Bürgerliche [Hop04]. Burkhardt [Dew07, Hod01, Hod06a]. Burnett [Joh09]. Burney [Ham08b]. Burns [Coo07a]. Burroughs [Jon09]. Bursting [Mor06c, Rud05a]. Burt [Mun08]. Busard [Ste04a]. Business [Rud07, Mor05c]. Bynum [Biv08, Das07, Far09a].

C [Aga01a, Aga03a, Bro03b, Daw00, Dea09, Dix03a, Dix05, Dix09, Fin03a, Gou03a, Gou06, Gou03b, HW01a, Hag04, Jam01b, Maz07a, Mor05b, Mor07a, Sch02a, Sch01, Sen06a, Smi02b, Smo08, BR96, Gil08, Har00c, Man00a, Rat07, WHF08]. C$39.00 [Pal06]. C. [Far07a, Smi00b]. c.1700 [Cro07]. c.1870 [Cro07]. C.S.I.C [Med00]. CA [Cha03, Haf08, Han07b, Kra07, Eds04, Loz00]. Cabinets [Alb08, Arn06b]. cadavers [Mar09]. Cadbury [Pad01]. Cailliau [Aga01b]. Calaprice [Sta03]. calculating [Rat07]. Calendar [Ash01, CDK98]. Califonia [Han07a]. California [Bra00, Dig08, Goo06b, Hug07b, Leh09, Mar05, Rob05, Seg05, Sle02, Ste07, Ste04b, Wil09c]. Called [Cro05a, EoToBB01, Fer03, Ste04a]. caloric [Lev00]. Calvet [Liv05a, Bro02a]. Calvinist [Gin04, Iro04, Kno02, Ver02]. cambio [Fox05]. Cambridge [Ack08, Aga03a, Aga05, Alb03a, And03a, Ash01, Ash04b, Bar01b, Ben02, Ber06b, Ber06a, Ber08, Biv06, Biv08, Bow07, Bra07, But06, Can08a, Car01a, Cla06, Col00, Col01b, Col02, Coo07a, Daw07a, De 07, Dea06a, Del09, Dep07, Dix06, Dix09, Dod06, Dol04c, Eag06b, Edd03a, Edd04a, Edd08, Edd09, Edg08, Fin02, Fox07, Fre04d, Fri04, Ful01a, Ful01c, Ful03, Gau08, Gau09, Gil08, Goo01, Goo03, Goo06a, GG03b, GG06a, Gra05, Haf08, Hal03, Hal06, Har01a, Hay02b, Hem03, Hen07a, Hig06, Hod01, Hod06a, Høy01, Hug08, Hun05a, Hun05b, Hun07a, Joh03a, Joh04c, Joh05c, Jon02a, Jon04, Kas02, Kip04, Lam02, Leh08, Lev07, Log08, Mac00b, Mac08b, Mac08a, Mac09b, Mac05, Mac07, Mac08c, Mai09, Man01a, Maz07a, Mer08, Mir07, Mor03a, Mor01a, Mor02c, MW07, MW09b]. Cambridge [Ner05, Nut04, Olb02, Pag03, Pal02, Pan08, Pem04, Pet08, Por01, Pye09, Qui08, Rad02, Rad03b, Ram02, Ren06, Ric01b, Ric08, Rob06, Rob09c, Sen02b, Sha02, Sir06, Smi02b, Smi05b, Smi06f, Smi06g, Smi08b, Smi08f, Smo08, Sta03, Sta07a, Sta08b, Sta06c, Ste06, Sum06, Tho02b, Tim03a, Tim03d, Tim05, Tso07, Tur08, Tyb02, Urr07, Van02b, Wag05, Wee08, Wee09a, Wil06, Win08, Yam04a, Yam04b, AH05, Bla07, Ham03, HM02, Kip03, MMMM02, O’C05, Por09, War03]. Cambrosio [Maz04a]. Camerini [Gar03b]. Camerota [Han07b]. campaign [Spo02]. Campi [Woo06]. CAN$44.95 [Sta03]. Canada [Bar01b, Dun99]. Cancer [Tim02, Tim03a, Tim03c, Tim03d, Kre02]. Candaux [Coo05a]. Candler [Ter08]. Candolle [Coo05a]. canon [Rad00b]. Canopus [Bro06a]. Canton [For05, Haf04, Jam05, Jor01, Kni03, Mac01a, Par01a]. Cantor [Bro04c, Lip08, Sta07c]. Cantrell [Joh05b]. Cape [Bre06, Gli06, RH04]. Captain [Ade04]. Cardiac [Tim03b, Tim04, Jef01, Wei02d]. Care [Tim03b, Jef01]. Career [Kni03, Ber01, Edd02, Edd04b, GB02, Rat07]. Carey [Del09]. Caribbean [Hal04b, McC02]. Carl [Cha01, Fin02, Her05b, Bla03, Fet00, Kim05, Smi01a]. Carlino [Bro02e, Daw00]. Carlo [Pan06]. Carnap [Ful01b, Fri00]. Carnot [Chi04b]. Carocci [Cor04b]. Carolina [Rol06]. Caroline [Ash04b, Stu06, HH03, Hob06, Hos03]. Carolyn [Pre03]. Carozzi [Rud04b]. Carrel [Sel08]. Carson [Hug07b, Ree08]. Carsten [Led08, Nes06a]. Carte [Big01]. Carter [Mur02]. Cartographic [Luc05, Rel02]. cartographies [Nay05]. Cartography [Dod02, Hos01]. Case [Joh09, Qui08, Rob05, Bur07b, Dun09, Man03, Moo02, Nai06, Tur03a, Ver03]. Cases [Phi02, Ste09a, Cec01, CLW07]. Casper [Can00]. Cassiano [Fin03a]. Cassirer [Ful01b, Fri00]. Castell [Pye09]. Catalan [Goo02d]. Catalans [Goo02d]. Catalogue [Bro02e, Eag06a, Hig05, Hig07, Sch02a, Car99, HW01a, NHA^{+}02, vC02, vCA^{+}05].

Cataloguing [Mco01]. Catalunya [Fox05]. Category [Smi06g, Dix03b]. Catharine [Gou03b]. Catherine [Hig06, Kle09, Wee07, Hig06]. Catholic [Dea07, Bru01, Hel05]. Cathryn [Hug07b]. Cattle [Whi06, Der03]. Caudemos [Med00]. Cause [Wal03, Swe01]. Causes [Sta03, Mil01]. Causeway [Ken08]. cave [Som04]. Cavendish [Bro05b, Hut07, MS05a]. CD [Han07b, Hen05b, Hig08]. CD-ROM [Han07b, Hen05b, Hig08]. cease [Wer01]. Ceccarelli [Phi02]. Celebrate [Win08]. celestes [Goo02d]. Cengage [Sha09]. Census [Gol05, Gin02]. centenaries [MT07]. Centennial [For05, Hug07b, CH05, Lar01]. Center [CC07, Gar03b]. Centre [Brö02b, For03, Nes02, Pye06, Qui02]. centuries [Che06]. Century [Ade04, Aga03b, Bro05b, Bro03c, Can08a, Can04a, Car06, Cla07a, Coo03, Dea08, Del08, Dew01, Dix00, Edd08, Elm02, Fre04c, Gau08, Goo06c, Ham07a, Har05a, Har08, Hem02b, Hoc08, Joh09, Kus03, Liv05a, Lüt04a, Mac07, Maz07a, Maz09, Men00, Mor06a, Mor01b, Mus03, Pal02, Pan08, Rad00a, Rei07a, Ric09b, Rit02, Rol06, Smi05a, Smi08b, Spa02, Spa03, Tho03b, Til04, Tim04, Tur06a, Wal06b, Whi05, Wil09b, Yeo04, Alb02, Bar06, Bed03, Ber06c, Bos02, Bow06b, Bur07b, Can01a, Dav09b, Die06, ED06, Ene09, Fre01, FN06, Goo02e, Gou99, Gra07a, HH01, Hay02a, Hee08, Hof07, How00, Jam08, Kel00, Ken08, Maz00, Mus09, Nay06, NG04, Pap06, Ren05, Ric03, Sch06a, Sim02, Smi08g, Smi08h, Sta03, Sta01a, Zui05, Zui06, tH02, Alt99, Bar00b, Ben03]. Century [Bro02a, Bro03a, Cle00, Cro01b, GR00, Gra07b, HW01b, HP00, Hol00, HK01, Jac06, KLG03, LM02, Man03, MS05b, Mil00, Mor98, Mos07, Nye04, Ols08, Rob02, SS01, Swe04, Tay04, TS01, Wal06a, Wei02d]. ceraunia [Goo02e]. Certain [Aga01a, Sco98]. Cervin [Rud04b]. Cesare [Smi08f]. Cesi [Jar08b]. Ceylon [Win09a]. Châtelet [Ter08, ZH06]. Chabás [Goo02c]. chacun [BV00, Chi01a]. Chadarevian [Lau06]. Chair [AH05, Lev07, O’C05]. Challenge [Dix09, Lip08, Maz02b, CS06, Gra00]. Challenged [Joh09, Bur07b]. challenges [Cro01c]. Chambers [Mil02]. Chandak [Log08]. Chandrasekhar [Gui05]. Chang [Kle09, Kni06a]. Change [Chi02b, Dol04c, Lev07, Ris01, Web07, AH05, Fle98, Sto00, Wil02]. Changed [Hun06, Mah03]. changes [Kre03]. Changing [Bow07, Gou03c, Amu05]. Chaouli [Kni04]. Chapel [Rol06, Stu06]. Chapman [Bro06a]. Chappey [Cor05]. Chapters [Tyb04, BJ00]. Charcot [Sir06, DHC03]. Charisma [Kni08]. Charles [Cro04, Daw07a, Dea05, Gli06, Hod01, Hod06a, Hut07, Joh06a, MS05a, Mor06c, Pic07, Ric01b, Sha09, Ske04, Smi08a, Wad03, Bro02d, Cor06, Edd03a, Gol09, Hal02a, Har00a, Her05a, Hug09, Key00, Key01, Key02, Lea04, May05a, Ric06, Sim04, Smi06a, WPM05]. Charleston [Joh05b]. Charlotte [Bur08, Kus00]. Charvolin [Hoc09]. Chase [Mer08]. Chasing [Hen06a]. Chatelier [Fox04]. Chatto [Cla09, Dea02, Del09, Maz07a, Maz09]. Chaucer [Eag05b]. Chazdon [Hag04]. Chemical [Kle05, Mor01a, Mor02a, Mor03b, Mor06e, Roo08, Tim08, Aft01, Deb06, Sie02]. Chemicals [Lam02, Rus01]. Chemist [Dol01, Iro04, Ful00b, GS02, Kno02, Sim04]. Chemistry [Bro04a, Car01a, Car06, Chu02, Hen07a, Kni02, Kni04, Lev07, Lüt04a, Mor01a, Mor02a, Par01a, Ram02, Rob07, Sim05, Smi09a, Cha02a, Cle00, Cro03, Edd02, Edd04b, Fau03, GS02, HL00, Hor00, KLG03, Lev01, LTGS02, LBV00, Mor05a, NG04, Pri07, Roc01, Sim02, Tay06, Wri00, AH05]. chemists [GO09]. Chen [Joh02d, Sta08b]. Chene [Sut03]. Chester [Kir05]. Cheyne [Fri03a, Fri03b, Gue00]. Chicago [Aga06, And09, App05a, Bar01a, Bar08, Bax09, Ben02, Boh01, Boo07, Bor03, Bre05, Bro05a, Bro09, Bud09, Can09, Can02b, Chr05, Coh09, Cor00, Daw00, Del08, Dew07, Dix05, Dix08a, Dod02, Dol02b, Edd05b, Edd07b, Far06, Ful01b, Ful03, Ful09b, Fyf03a, Gat05, Gol09, Goo02b, Goo08a, Goo09a, Gor02, Gou09a, Gra05, Hag04, Hal02a, Hay06, Hei07b, Hei07a, Hen06b, Hou07, Hug09, Hut04, Jar08a, Joh06c, Kni05, Kni08, Kra08a, Lip08, Log08, Mac09a, McO09, Mor06c, Mun09, Nay09, Nes06b, Out04, Pag01, Pem07, Phi02, Por03, Qui04, Ren00, Ric09b, Ric01a, Ril02, Roo08, Rup08, Sch01, Sha07, She07, Siv08, Smi07b, Smi08a, Smi08d, Sta06b, Sta09b, Ste09b, Tim03c, Top02, Tot04, Tur06b, Van04a, Web06, Web09, Whi04, Wil09b, Yeo04, Yos00]. Chichester [Hun06]. child [Ste09c]. Childbirth [McT05, Wil09a]. Children [Har05a, Mor98]. chimie [Cro05b, Fau03]. Chimisso [Til04]. China [Bra07, Chu02, Dod02, Dod06, Elm09, Fan02, Nut04, Wag05, BDLM07, Dod01, Elm05, Elm06, Fan03, tF04, Hos01, NW08, Wri00]. Chinese [Cam08, Hsu06, Bra08, Llo04, Llo06]. Chinnici [Big01]. Chipley [Goo03, HHAH01]. Chisholm [GG03b]. Chlorine [Kle09]. Chris [Can09, Stu06]. Christchurch [Mor02a]. Christiaan [Gui04]. Christianity [Dix05, LN03]. Christine [Chi03b, Wer09, For03, Rit02]. Christmas [Mor09b, J^{+}07]. Christoph [Ric04, Par01b]. Christopher [Biv08, Lev07, Boo07, Hul07, Hut04]. chronology [GO09]. chronometers [Rob09a]. chrysalis [Kuk06]. Church [Han07a, Jon09, Bru01, McM05]. Chymistry [Edd05b, Smi09a, NP02, Pri07, Roo07].

Chymists [Pri07, Smi09a]. ci [Sta03]. ciel [Big01]. Ciencia [Med00, Goo02d]. Cinema [Boo07]. Ciphers [Eag05a, Kin01]. Circle [Tso05, Twe09]. circulating [Woo06]. Circulation [Das09]. Circumcision [Hou07, Dar05]. Citizen [Tho02b, Rig00]. Citizens [Pet08, Igo07]. citoyennes [Hoc09]. Città [Pug09]. civ [Sta03]. civic [Fin05a]. Civil [Aga04, Bre05, NB02]. Civilisation [Brö02b, Wag05, NW08, çG00b]. Civilizing [Edg08]. claims [Ste09c]. Claire [Mun08]. Clare [Jon06a]. Claremont [Bro03b]. Clarendon [Bro08b, Gui05, Hen02a, Hod06b, Hsu06, Hug01, Man00a, Man06, Rav01, Smi09b, Tho03a, Tra07, Yam00]. Clark [Dig09, Gou09b, Kni08, Ric01a, Ril02]. Clarke [Har00b]. Class [GG04, Bur01, Yan02]. Classic [Hag04, CW02]. Classical [EB07, Har01a, Høy01]. Classics [Ric06, Goo02d]. classification [Edd02, Edd04b, Mae06]. Claude [Maz04a]. Claudia [Del06a]. Claudine [Van04a]. Claus [Dea09]. Clearance [Tho03c, Pol02]. Cleempoel [Hig05, Hig07]. Clements [Sch02a, HW01a]. Clencia [Med00]. Clercq [Hig02]. Clerestory [Mor02a]. Clericuzio [Lüt04a]. Clerk [Har02a, Hun05b, Hun06, Mah03]. Clifford [GG03b, Chi02c]. Climate [Nay09, Ris01, Fle98, FJC06, Gol07]. Climbing [Kai05a, MM00]. CLIO [Coo07a, Gou03b, Sho03, Ade04, Edd07a, Jen01, Kno01]. Clockmaker [Eag07, Nor05]. Clocks [Sut03, Des01]. Closed [WM08, Elz06]. cloth [Bra00, Dol02b, Mor02c, Ram02, Sta03]. Club [Mor05b]. Clucas [Hen04a, Mos08b]. Co [Jon09]. cobbler [GB02]. Coca [Spa06, KNRW03]. Cocaine [Spa06, KNRW03]. coelestium [Hen05b]. Coen [Mun09, Nay09]. Coffee [Bro04a, LTGS02]. Cognitive [Abr08, Bor03, CC07, Llo07, Smi09b, Sta08b, ABC06, Bod06, Tur02b]. Cogwheel [Cro04]. Cohen [Bra00, Gra08a, Hen03a, Hen05b, Van04a]. Cohesion [Row02, Yam04a]. Cold [Coo02b, Cue07, Pen04, Tim03c, Tso07, TF02, Web09, BG02, Kre02, Rei05c]. Cole [Sen02b]. Colin [Jam03, Mor01a, Ver04]. collaboration [Har00a]. collaborations [Hom09]. Collected [Bek06, Bre04, Cam08, Edd03b, Edd05a, End04a, Gou03a, Han07b, Ril00, Ril01b, Sam07, Ste05b]. Collecting [Alb01, Alb08, Ham01b, Kne00, Nai06]. Collection [Chi04b, Hig02, Hig05, Sch02a, Spr06, HW01a]. Collections [Can01b, Joh07a, Pan06, Alb02, Ber05, Ham01b]. Collective [Ric01a, Cav08]. Collector [Coo04, Mil04]. Collectors [Fyf02b, Joh07b]. College [Aug09, Gay07, Jor08, Mil02, Sta07d, FG05]. Colleges [Kim07, Har05b]. Collie [Mor07b]. Collies [Whi06, Der03]. Collins [Aga06, Ful03, Goo09a, Pem07]. Colonial [Del06a, Dod02, Rob06, SS05, Bro03a, Hos01, Kum07, Nai06, Sch04a]. Colonialism [Kha06a, Bas04a]. Colour [Kre06, Joh01]. Colouring [Kle04, NG01c]. Columbia [Dol02c, Mun07, Web06]. Combe [Fra09]. Come [Brö04b, Har02b]. Comenius [Man01b]. Comenius-Universität [Man01b]. Comfort [Dix09]. Commemorate [Wee07, SM05b]. Commemorating [Bro05d]. commemoration [AAH00a, AAH00b]. Commemorative [Ric01a]. Comment [Var00]. Commentaries [Ber06a, Hag04, oAEN04, Ber06b, CW02, Hen02a, Man06]. commentarii [Hen05b]. Commentary [Ber06a, Hen05b, oAEN04, Ber06b]. Comments [Gou06, KG04]. Commerce [Del06a, Kno08, Coo07b, SS05]. commercial [Hor07]. Commercialization [Nes09, Bro08a]. Commission [For04].

Common [Gui05, Pen04, TF02]. Commonwealth [Elm02, Poo09, Dub06]. Communicating [LBV00, Par01a, Sec08, Kni06b]. Community [Smi08b, May05b, Mos07]. Companion [Dea06a, Fyf05, Rob08, Smi05a, Tur06a, Bui07, Ham03, HBB^{+}03]. Company [Aga00, Cro04, Win09a]. Comparative [Kir05, Nes02]. Compasso [Han07b]. competent [Mco01]. competition [O’C05]. complete [Can01b, Noa09]. Composition [Kle05, Sie02]. Computer [CK02, Cro04, Cro05a, Hod06b, Fer03]. Computers [Aga03a, HH00]. Computing [Col02, Cop04, Sum06, Mac01b, RS02]. con [Fox05]. Conant [Bar01a]. Concept [Bou02, Edd07a, Gui04, Olb02, Bai05, BFR00, Hof07, Sch05b, Zag01]. Conception [Kni05, Ric02b]. Concepts [Kre03, Rob06]. Condition [Aga01a, Sco98]. Conditions [Mor08, Mus06]. Conference [Hig08]. Confessions [Kus03, HW01b]. Conflict [Coo03, Cro01b]. conflicto [Fox05]. Confronts [Ade08, Dol04c, Wil02]. Confucian [Fan02, Dod01]. Conica [Cuo02, FU01]. Conjunctions [Jor02, Ham01a]. Connor [Edd06]. Conquest [Har02c, Web09, Hay01, Sno06]. Conservation [Hal04a, Wei00]. conservée [Big01]. consimilibus [McG08]. Conspectus [Ste04a]. Constance [Dig09]. Constant [Hof07]. Constantin [Nes01]. Constructing [Chi02b, Spo02, Sto00]. Construction [Cor00, Das09, Rad02, Edn97, Set07, Sum01]. Constructivism [Edd04a, Gol98]. constructivist [Edd04a]. Consumption [Sum05, Ash03]. contemporains [Maz04a]. Contemporaries [Dea09]. Contemporary [Hun07a, Maz04a, dC08, DS06, Kai05c]. Contesting [Del09]. Context [Bro04c, Cuo02, Nut04, Spa02, And03b, BZ04, FU01, Hor07, Loc07, Lov01].

Contexts [Coe09, Eds04, Kra08b, Nay05]. Continuum [Fra09, Mac06b, Maz07a, Mor06a]. Contours [Rob06, Bai05, Bor07]. contribution [Tei00]. Contributions [Hör78, Bro04a, Eag06a]. Controlling [Dod01, Fan02]. Controversies [Ful02, MPB00]. Controversy [Bro05b, Dix09, Edd07a, Kle09, Maz07a, Com07]. conventions [Mor09a]. Conversation [Ful03, LC01]. converses [Pat07]. Conway [Cla04b]. Cook [Kno08]. Cookson [Joh05b]. Cooper [Edd09]. Cooter [Smi05a]. Copeland [Hod06b]. Copenhagen [Kra07]. Copèrnic [Goo02d]. Copernican [BZ04]. Copernicana [Gol05]. Copernicanism [Nay03]. Copernicans [Gin04, Ver02]. Copernicu [Gol05]. Copernicus [Hen05b, Hen06a, Gin02]. Cornelis [Hen04d]. Cornell [Mor06c, Mor06b, Sut03, Tho03c]. Cornwall [Nay06]. Corpuscles [Lüt04a, Cle00]. Correspondance [Big01]. Correspondence [Ash01, End04b, Har00b, Hod01, Hod06a, Hut07, Mor09b, Nea00b, SEM04, Chi03a, CDK98, FJ08, MS05a]. Corry [Gra06]. Corsi [Bur07a]. Corvallis [Wil09d]. Cosmographia [Mos08a, McL07]. Cosmographic [Luc05, Rel02]. cosmológico [Man00b]. Cosmology [Coe09, Kra08b]. Cosmos [Aga04]. Cossons [Div01]. Coulson [Sim04]. Coulston [Cor06, Har00a]. Council [Ano00e]. Country [Hul07, Law05]. Cours [Rud04b]. Court [Ful01b, Joh09, Web06, Bos02, Bur07b, Mae06, Slo04]. courtier [Rat07]. courtier-inventor [Rat07]. Courts [Ade06]. Cover [Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h]. Cowen [Edd03b]. Cowper [Mus09]. Cows [Seg05, Sch03b]. Craft [Mor02c, Jac00a]. crafting [Sim04]. creación [Fox05]. Creager [Kir05, Ste09a]. Creating [Can04b, Jon02a, Mur00]. Creation [Bro03c, Coe09, Ric09a, Smi06g, Top02, Dix03b, HH01, Kra08b, Ram01, Sec00]. Creationism [Leh09, Lyn02, Rob03, Rob05, Moo02, NS00, Sed07]. creations [Nor07]. Creativity [For05, Lar01]. Creatures [Joh07b]. Credit [She07, Bia06]. Creed [Ril01b, Mac00c]. Crelinsten [War08]. crime [Bro03a]. Crimes [Sen06a, Lee02]. Criminal [LGR06, Sen02b, Smi08f, Col01a]. criminality [Bro03a]. Criminals [Smi08f, BW06]. Criminology [Smi08f, BW06]. Crisis [Set07]. Cristina [Shr09, Til04]. Critas [Joh04a]. Critias [Leh08]. Critic [Til04]. Critical [Ber06a, Jor02, Kha06a, Smi03b, oAEN04, Bas04a, Ber06b, Ric02a]. Critics [Can02b, Leh09, Sed07]. Criticus [Ste04a]. Crocker [Man04]. Cronin [Coo03]. Crookes [Nes09, Bro08a]. Crosbie [Rei07a, Yos00]. Crosland [Gil08]. Crowe [Ash01]. Crowther [Chi03a]. Crucible [Jon09]. crusades [Loc07]. Crystal [Aga00, For04, RH97]. CS567 [Ril00]. CS598 [Ril01b]. CS700 [Edd03b]. CS713 [Gou03a]. CS724 [End04a]. CS726 [Bre04]. CS736 [Edd05a]. CS751 [Ste05b]. CS773 [Bek06]. CS803 [Sam07]. CSLI [Cha03]. CT [Can08b]. CTHS [Fox04]. Cuddington [Web07]. Cueto [Web09]. Cullen [Tay06]. Cult [Nes06b, Hau03]. Cultivation [Sha07, Jon06b, Smi00b]. Cultural [App05a, Bra07, Bur08, Cla01, Dod06, tF00, Fre04d, Gou03c, GG03a, Haf04, Hal04a, Rei07a, Ste06, Yos00, BZ04, BS02, Cro01c, Elm06, tF04, Jan00, MS05b, Sle07, Wat97, Wei01, Wei00]. Culture [Alb01, Alb03b, Bar07, Ben04b, Bre04, Bre06, Bro05b, Brö02b, Bro08b, Daw07a, ETV04, End04a, Fis04, Ful03, Ful09a, Gal97, Gou09b, Gou03c, Hay06, Hsu06, Jac09, Joh03a, Joh06c, Jor08, Mar05, Maz07a, Mur02, Mus03, Pem04, Ren00, Ril00, Rob07, Rob09c, Smi01b, Smi06d, Spa03, Whi06, AA01, Asm01, Car01b, Cla07b, Cuo07, Dun09, ED06, Fei04, Fin04, Gau06, GHL03, Hug04b, Jon08, Kne00, LC01, Llo04, Lon01, LM02, Maz04b, Mor97b, RT00, Ryl00, SL06, Smi06a, Sok08, Sta03, Sta01a, Tho02a, Yeo01a, Yeo01b, Zic00]. Cultures [Bee01, Car06, D’A03, For05, GL00, Mac08c, Plo05, Roo08, Sle02, Wee07, Dri01, KLG03, SS00, Sel00, Lar01]. Cuomo [Høy01, Rob09c]. Cure [Tim03c, Kre02]. curiosità [Ber07, Fin09]. Curiosity [App05a, Eag08, EM06, Nut04]. Curious [Alb08, MW09b, Arn06b, Som07]. Currents [GG03b, Chi02c]. curse [Ken07, Ken07]. Curth [Rut08]. Cushing [Sch02a]. Cushing/John [HW01a, Sch02a]. Customs [Ash03, Sum05]. cxvi [Man00a]. cxxviii [Man06]. Cybernetics [Cla06, CC08, Ger02]. Cyberspeak [Cla06, Ger02]. Cyborgs [Pal06, Bow03]. Cylinder [Ber06a]. Cynthia [Del08]. Czech [Mor07c].

D [Bar07, Ben06, Bro05c, Col01b, Dix03a, Dun08, Edd06, End04b, Fyf02a, Gor02, Hug07a, Joh02b, Nes02, Rol06, Sir09]. D. [Joh09, Rav01]. d’Aigues [Hoc09]. Daitch [Hes04]. dal [Fin03a]. Dalen [Rav01]. Daley [Rol06]. D’Ambrosio [Plo05]. Dame [Dea07, Han07a, Hei08]. Damned [Bro03c, HH01]. Dando [Mac08c]. Dangerous [Boo07]. Daniel [For05, Jen01, Coo05a, Del09, Gra08a, Wei02a]. Danielle [Cha05, Cro05b]. dans [Big01, Hoc09, Rud04b]. Darby [Hou07]. DARPA [Sum06, RS02]. Dartmouth [Jor08]. Darwin [Ade08, BHR00, Bro03b, Daw07a, Dea05, Far07a, Gli06, Har02d, Hod01, Hod06a, Mor03a, Mor06c, Ric06, Ric01b, Ske04, Tur09a, Tur08, Web06, BHR00, Bro02d, Bro05d, Bul04, Daw07b, Gre99, Her05a, Key00, Key01, Key02, Rud05b, Slo04, SA05, Smi06a, WS01a]. Darwinian [Ren06, LRR04]. Darwinism [Daw06, Dix09, Lip08, Rad00a, Rad03b, Ade05, Alt99, Ben04a, CS06, Str00]. Darwins [Web08]. d’Ascq [Chi01b]. Daston [Boh01, Far06, Fin01, Jor06, Mun07, Wee08]. dating [GO09]. Dauben [Ric04]. Daughter [Ske04, Key01]. Daum [Hop04]. Daunton [Fyf07]. David [Ash01, Fin03a, Gar03b, Goo02b, Gra06, Haf08, Hug07b, Jor02, Mor07c, Siv08, Tur02a, Yam04b, Aga02, Bar07, Bow06a, Bro05b, Bro04c, Can04a, Chu02, Cor04a, Dix05, Eag05a, Edd03b, End04a, Far07b, Fra09, Fyf02a, Goo02a, Ham07a, Har08, Hun07a, Kra08a, Leh09, Lyn09, Mac08a, Mac09b, Pen04, Rob08, Sec08, Sha03b, Smi08a, Tho03a]. Davies [Smi07a]. Davis [Dea02, Lea04, Goo06b]. Davy [Dol01, Ful00b]. Dawn [Tur09a, Ham08a, Ogi00]. Dawning [Rav01]. Dawson [Tur08, Pac03]. Day [Biv06, Mor03b, Har04b]. Days [Pag01, Tra01]. DC [Aga04, Hes04, Smi08f, Sta03]. Deadly [Mac08c, Web09, Cue07]. Dean [Aga02, DeV00]. Dear [Hei07a, Hen04b, Mac09a]. Dearborn [Fyf03a]. Death [Ade04]. Debate [Bow06a, Dix08a, Kra07, Maz07a, Rad03b, Smi08d, BS01, Ken07, Man00b, Mcg06, Rad07, Sny06, Sta07f, Tho09]. Debates [Maz04a, Rad03b, Str00]. Debating [Bro03b, Goo08b, Gre99]. débats [Maz04a]. Deborah [Jam01b, Jam05, Mar05, Nay09, Har06, Mun09, Pad01, Wee09b]. Debus [Hen07a, Tim08, Deb06]. Débuts [Fox04]. December [Jam01b, Jam05]. Décio [Sta09a]. deCODE [Wil09c, For08]. Deconstructing [Hug04a]. découverte [Chi03b]. Dee [Mos08b, Clu06]. Deep [Aga04, Haf08, Mud01]. Deepak [Mac00a]. Defence [Coo02b, BG02]. Defibrillator [Tim03b, Jef01]. degree [Cai00]. del [Ber07, Fin09, Han07b, Pug09]. Delbourgo [Ber08, Rob09b]. Delft [Ben06, Hue03]. Delhi [Mac00a]. delineating [Mco01].

Delisle [Goo08b]. della [Man00b]. dels [Goo02d]. deluge [Bur07c]. Democracy [Pau09, PP07]. Demography [Ren07, Sch06b]. Demonization [Hou07, Dar05]. demonstration [Ver05]. d’Enginyeria [Fox05]. Denis [Mor02a]. Dennis [Sut03]. Denver [Gou03b]. Department [Sha03b, Whi05, Dry09]. dépossédé [Hig03b]. Depression [Fri03a, Fri03b, Gue00]. Derek [Wat00]. Derry [Dol04a, Whi06]. Descartes [Sha07, Sut03, Des01, Jon06b, Sha00, Wer01]. Descent [Dix09, Ken07, Ric06]. Describing [Mos08a, McL07]. Description [Del08, Wal06a]. Descriptions [Ste04a]. Design [Dix09, Edg08, Com07]. Designing [Bar08, Ren05, Lig07]. Designs [Led08]. Desmond [Ric06]. Destined [Joh06a, Wad03]. d’Estudis [Goo02d]. Detection [Ham08b]. Determined [Fox07, Hah05]. Detroit [Sha09]. Deutsch [Joh02b]. deutsche [Hop04]. deutschen [Har01b, Hoc03, Nes06c, Tei00]. Deutsches [Sta04]. Deutschland [Har01b, Har01c, Tei00]. developed [Mai09]. Development [Bru06, Coo03, Maz02a, Olb02, Pie00, Ric04, BFR00, Cro01b, DS02, DGS03, Hor07]. Devices [Led08]. Devo [Bow07, Amu05]. devotion [UU07]. Dew [Rob09b]. Dfl [Hig02]. d’histoire [Coo05a, Cor05, Kim05, Rud04b]. Diabetes [Stu06]. Diagrams [Ber06a, Kra08a, oAEN04, Ber06b, Kai05b]. Dialogue [Can02b, Mel01, ML00]. Dialogues [Maz04a]. Diana [Wal06c, Chi08, Wal06c]. Diary [Dol02a, ABB01, Gol01]. Diaspora [Bar01b, Dun99]. Dibner [Aga03a, Hun05a, Ram02, Van02b]. Dick [Fyf03b, Joh05c, Ast01]. dictionaries [Yeo01a]. Dictionary [Biv06, Das07, Gou03b, Mor06a, Sha09, Smi02b, Yam04b, BP05, Hai01, Kip03, MMMM02]. Did [Bul04]. Diderick [Hen03a]. Diderot [HP02]. Didi [Sir06]. Didi-Huberman [Sir06]. Diego [Pet07a]. Diemer [Mor07b]. Diepholz [Nes07, Sta04]. differences [Hof07]. Different [Mor02a]. Digby [Hed08]. Digital [Hen05b]. Dijksterhuis [Rav00]. Dilemma [Kim07, Har05b]. dilemmas [Chi03a]. Dimension [Lau06, dCH04]. Dinosaur [Pad01, Cad00]. Dioptrics [EB07]. Dipartimento [Goo05]. Diputación [Fox05]. Disability [Tho03a, Wri01]. disciplinary [CC07]. discipline [Ren05]. Disciplines [Van02b, GS99]. Disciplining [Ren07, Sch06b]. Discord [Goo06b, Fra02]. Discourse [Fri03a, Fri03b, Luc05, Rel02]. Discover [Goo09b]. Discoveries [Dea05, Joh03a, Key02]. Discovering [Bro05b]. Discovery [Ben06, Dol04b, Goo06c, Haf08, Hal06, Hut04, Pad01, Ste05a, Ste04b, Ale02, BC01, Cad00, Her01, Hoo02, Hue03, Swe04, Wea03]. Discussing [Bro04a, LTGS02]. Disease [Biv06, Doo05, Har02c, Har04b, Kha06b, Kip04, Pal00, Tim04, EG04, Har00c, Hay01, Kip03, Llo03, Tha99, Wei02d, Wor00]. disks [Noa09]. Dispersion [Kra08a, Kai05b]. Display [Wil09a, Ade05, McT05]. Displaying [Bek01]. Dissection [Bro01b, Fre99]. Dissent [Dix09]. Distilling [Hen07a, Mor05a]. Distinguished [Joh06a, Wad03]. Distributed [Boo07]. disunities [Pic09]. Divall [Ver04]. Diversis [Hig05]. Diversity [Del09, Smi09b, Llo07]. Divide [Mar05, GHL03]. Divina [Kim05, vLP01]. division [Bor07]. divulgazione [Cor04b, Gov02]. Dixon [Far09a, Smi06g, Tur09b]. Djerassi [Bla03]. DM [Eag05a, Har01b, Har01c, Har02d, Hop01, Joh02b, Kra02, Maz02a, Nes01, Smi02b, Sta03, Wei02b]. DM198 [Hug00]. DMV [Set04]. DNA [Sen06a, Lee02]. Dobzhansky [Phi02, Cec01]. doctor [Sha00]. Documentales [Med00].

Documentary [Sle09b]. Documentos [Fox05]. Documents [Aga03b, Aga04, Kra07, Chi01b, Hul00, SS01]. Dodgen [Fan02]. Doel [dC08]. Dogs [Rit09]. Doing [Gou03c, RT00, Ben03]. Dolan [Kno01]. Dolling [Bru06]. Domenico [Dea08]. domestic [Aga04]. Dominance [Edg08]. Domingo [Han07b, Wal04]. Dominique [Tur06a]. Don [Sta03]. Donahue [Smi02a]. Donald [Col02, Ghi01, Emm01]. Donna [Ree07]. donné [Rud04b]. don’t [May05a]. Doors [WM08, Elz06]. Dordrecht [Chi06, D’A03, DM05b, Ghi01, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Joh02d, Joh06a, Kim05, Kle04, Loc02, Lüt04a, Man04, Maz07a, Mos08b, Plo05, Spa05, Tso05]. Dorinda [Coo07a]. Dorman [Jam01b, Jam05]. Dorn [Ril01a]. Dorofeeva [Elm09]. Dorofeeva-Lichtmann [Elm09]. Doron [Cro04]. Dörries [Kra07]. Dosi [Qui08]. Douglas [Aga04, Hal04a, Har02c]. Downes [Hig06]. Downlink [Aga04, Mud01]. Dr [Edd02, Pem07]. Dr. [CP05]. Dragon [Fan02, Dod01]. Drake [Han07b]. Drama [Sch03a, Hil00]. Drawing [Kai05b, Kra08a]. Drawings [Raj09, Sch02a, HW01a, Nol07]. Drayton [Sch02b]. Dream [Smi06f, Sha03a]. Driver [Bee01, Siv08]. Drouin [Coo05a]. Druckwerke [Smi06c]. Drug [Tim03a, Tim03d]. Drummond [Sco04]. Dublin [Ade06, Bri04, Lea04, Ade05]. Dubow [Gre02, Poo09]. Duke [GG03a, Ren07, Smi08f, Ste09a, Wor07]. Dukhavina [Sir09]. Duncan [Hod01, Hod06a]. Duncanville [Hig04b]. Dünkel [Bru09]. Dunlap [Bar01b]. Dunn [Eag06a]. Durbach [Wor07]. Durham [GG03a, Ren07, Smi08f, Ste09a, Wor07]. during [Ben04a, Ben06, Daw06, Hue03, Whi05, Zim00b]. Dutch [Gin04, Kno08, Coo07b, Ver02]. Duxford [CK02]. DVDs [Noa09]. Dyck [Ash01].

Dyestuffs [Kle04, NG01c]. Dying [Lev02, Whi04]. Dynamics [Cha03, Fri01]. Dyson [Sta03].

e-book [Mos08b]. E. [Far00c, GG03a, Hay03, Sch01]. each [BV00]. Eagleton [Hig06]. Earle [Haf08]. Earliest [Goo09b]. Earlswood [Tho03a, Wri01]. Early [Alb08, App05a, App05b, Ber08, De 07, Dea07, Dea09, Del06a, Dod02, Edd09, Fie05, Gou09a, Gou09b, Han07b, Har05a, Hen07a, Hob06, Hoc08, Hut07, IH05, Kus07, Loc02, Luc05, Mac09a, Man04, Maz07a, Mcc05, Mor03b, Mor06e, Mor08, MW07, Pag03, Qui04, Sha02, Sle09a, Smi08c, Smi09a, Ste05b, Wee07, Wee08, Wil09a, Aab01, Arn06b, Bow06b, BM05, Cla07b, Coo07c, Cro01a, Dav09b, Del06b, Dry09, Edd02, Fin04, Gau01, Gra07a, Has09, Hel05, Hof07, Hos01, Hun07b, Ili00, Joh00, KM06, LRS03, MS05a, Mar09, Mar04, McL09, McT05, Mor09a, Mor98, Mus06, NG01a, NG04, PS05, Pri07, Rat07, Rel02, Ren05, RSD07, Sch06a, SS05, Sil01, Smi06b, SM05b, Ste00a, Wal04, Wea00, Win09b, Zui05, Zui06]. Early-Modern [Sha02, Gau01]. Early-Nineteenth-Century [Har05a, Mor98]. Earth [Cla04b, CKM99, Eds04, Fin02, Mor06c, Rup08, Shi03b, Smi09a, Loc07, Roo07, Rud04a, Rud05b, Smi01a, Loz00, O’C07, Ore01]. Earthquake [Mor07c, KMO05]. East [Chi02b, Hal04a, Sto00, Win09a]. Easy [Sta07a]. echo [TDNS08]. echo-sounding [TDNS08]. Eckert [Sta04]. Eclipse [Big06, Daw06, Loc02, Ben04a, Pan02, Spo02, Ste00a]. Ecological [Web07]. Ecology [Hal02a, Hal03, Hal04a, Web07, Ank01, Elt01, Kin05, Wei00]. Economic [Coo03, Coo06b, Coo05b]. Economics [GG03a, Mir07, Por03, Smi07b, Maa05, MS02, Sch05a, Tei00, Wei02c]. economies [CC07]. Ecstatic [Gor02, Row00]. ECU [Big01]. ed [Aga04, Ash04b, Ber06a, Big07, Biv06, Boh01, Bow06a, Bro05a, Bro04c, Can01b, Cha01, Chi01b, Chi04b, Cla04a, Cro05b, D’A03, Dea06a, DiM07, Div01, Dix09, Dou06, Eag05b, Edd06, Elm02, Fis04, For05, Fox05, Fra05, Ful09a, Fyf02b, Fyf03a, Fyf05, Fyf07, Gar03b, Gis09, Goo01, GG06b, Gre02, Han07a, Har04c, Hen01, Hen04c, Hen07a, Hen03b, Hen04e, Hes01, Hod06b, Hun05b, Hun07a, Jen02, Joh07a, Jor01, Jor02, Jor06, Kha06b, Kno01, Kra07, Mac01a, Man04, Man01b, Maz02a, Maz07a, Mil02, Mor06a, Mor01a, Mor09b, Mos08b, MW07, Mül01, Nea00a, Nes01, Nes06a, Nye09, Par01b, Pet07a, Plo05, Pye06, Qui02, Rad03b, Rei07b, Ric01b, Rob03, Rob08, Rud04b, Set04, Sha09, Shi03b, Sho03, Shr08, Shr09, Siv06, Smi00a, Smi05b, Smi09a, Sta03, Sta06c, Ste09b, Ste04b]. ed [Tho03c, Twe09, Wil06, Zig04]. Eddington [Bro09, MCS09, Sta07e, Sta07f]. Eddy [Bar07]. Edgar [Hen03a]. Edgerton [Far07b, Mac09b]. Edinburgh [Edd06, Fin04, Hul07, Joh04b, Raj09, Law05]. Edited [Bar01a, Can02b, Hen02a, Kim05, Man01a, Man06, Dea02, Man00a]. Editeur [Coo05a, Rud04b]. Editing [Lev00]. Edition [Ber06a, Can01b, Can02b, Eag05b, Elm02, Fyf02b, Mel01, Ric08, oAEN04, Ber06b, Bra00, Coo07a, Hop04, ML00, Mor02a, Mor03b, Mor01c, Ric06, Smi03b, Yam04b]. Editions [Chi03b, Edd07a, Fox04, Hen05b, Hoc09, Rud04b, Rud04b, Spr06]. editore [Cor04b]. Editorial [ACC09, Ano00f, Ano09n, Goo02d, Sta03, Ste04a, Liv03]. Edmund [Lam02]. Edney [Cor00]. eds [Aga03a, Aga05, Ash01, Bar07, Bre05, Bru06, Bru09, Can08a, Cha05, Chi06, Chi03b, Cla04b, Coo05a, Coo06c, Coo02b, Das07, De 07, Dea09, Del06a, Dix03a, Dix05, Eag08, Elm09, End04b, Far07a, Far06, Fie05, Ful09b, Fyf02a, Ghi01, Goo02b, Goo03, Goo05, Gou09a, Gou03c, GG09, Haf04, Haf08, Hag04, Har02d, Hem02b, Hen03a, Her05b, Hig06, Hig08, Hoc09, Hod01, Hod06a, Hug07b, Hun05a, Hur07, Jam01b, Jam05, Joh02b, Joh05b, Jor02, Jor08, Kem09, Kha06b, Kir05, Kle09, Kra02, Kus03, Kus07, Lau06, Lea04, Lea05, Led08, Lev07, Lip08, Mac09a, Mac01a, Mac06b, Mac08c, Mag06a, Mar05, Maz04a, Maz07a, McG08, Med00, Mor07a, Mor07b, Mor02a, Mor06d, Mül01, Mun07, Mus03, Nay09, Nea00b, Ner05, Olb02, Pag03, Pan06, Par01a, Pau09, Pic07, Por01, Por03]. eds [Pye09, Qui08, Ram02, Ras06, Ren06, Ric04, Ric09b, Ric01a, Ric08, Ril02, Rob09b, Rol06, Rud04b, Siv08, Sle02, Sle05, Smi05a, Smi08a, Smi08f, Smi08e, Smi08c, Sta03, Sta04, Sta07d, Ste09a, Ter08, Tos09, Tso05, Tur06a, Van02b, Vid05, Web07, Wee07, Wee08, Wee09a, Wer09, Win08, Yam04b, dC08]. Education [Aug09, Coo03, Nut05, Cro01b, Gay07]. Edward [Dea02, Seg05, Wil06, Gau08, Ric08, Tho03b]. Edwards [Hen04e]. Effect [Edg04, Ham07b, Har03b, RH05]. Efimovich [Yam00]. Efron [Can08b]. Eighteenth [Can04a, Del08, Edd08, Goo06c, Hoc08, Liv05a, Maz07a, Spa02, Spa03, Yeo04, Die06, ED06, Goo02e, Gra07a, Hay02a, Hee08, Sim02, Zui05, Zui06, tH02, Bro02a, Rob02, Swe04, Wal06a]. eighteenth- [Gra07a]. Eighteenth-Century [Can04a, Edd08, Goo06c, Hoc08, Liv05a, Maz07a, Spa03, Yeo04, Die06, ED06, Hay02a, Hee08, Sim02, Zui05, Zui06, tH02, Bro02a, Rob02, Swe04]. Einblicke [Man01b]. Einstein [MW09a, Sle03, Sta03, Sta09b, WM08, War08, Bar00c, Cre06, EC00, Elz06, Hol00, HS00, Mil01, MCS09, Sch08b, Sta08a, Whi01]. Eisner [Eag05b]. Ekmeleddin [Bek06, Brö02b]. elasmobranchs [Bla07]. Elder [Doo04, Hea99]. Electric [BP01, Goo02a, Goo05, Mor00]. Electrical [But06, Can01b, Emm01, Hun02, Joh02c, Ber06d, Goo04b]. Electricity [Ber08, Goo05, Har05a, Hen05b, Hoc03, Mor01c, BP01, Del06b, Far00b, Mor98]. electromagnetic [Bek01]. Electron [Hun05a, BW01]. electronic [Sec05, SD00]. Electronics [Mül01]. Electrons [Goo08a, Ara06]. Elektrizität [Hoc03, Nes06c]. Element [Kle09]. Elementa [Ste04a]. Elements [Cle00, Kle05, Lüt04a, Ste04a, Sie02, AH02, EoToBB01]. Elephant [Rob02, Spa03]. Elias [Hig04c]. Elie [Ghi01, GWT00]. Elizabeth [Ste09a, Coe09, Mor08]. Elizabethan [Hen01, Hig01, Tra07, Tre06, Wee09b, Ash04a, FH00, Har07a, Tur00]. Elliott [Ric01a]. Elliston [End04a]. Elman [Bra07]. Elmer [Kha06b]. Elsevier [Maz04a, Pet07a]. Elsevier/Musée [Maz04a]. Elton [Hal02a]. Elusive [Web06, Fic04]. Elwick [Cla09]. Elzinga [WM08]. embodiment [Smi00b]. Embryo [Bow07, Amu05]. Embryologe [Hop01, Men00]. Embryological [Alb03a]. Embryologist [Men00]. Embryos [Alb03a, Hop02, Mai09]. Emergence [Bro08b, Hig04b, Gau06]. Emerging [Pac03]. Emilie [Ter08, ZH06]. Emily [Pem04]. Eminent [Dou06, Bai03]. Emma [Har01a]. Emotions [Smi06g, Dix03b]. Empire [Bee01, Bek06, Big06, Bre06, Cor00, Del09, Dod06, Gas98a, Hal03, Mac00b, Maz07a, Mil05, Pan02, Rob06, Rob09b, Roo08, Siv06, Siv08, Sta07b, Ste06, DD08, Dri01, DM05c, Edn97, Fan03, tF04, JS04, RH04, Sch04a, Siv05, Ank01, Cro01c, Num07, çIh04]. Empires [Rei07a, Rei07b, MS05b]. Empiricism [De 07, PS05]. Empiricists [Out04, Ris02]. enactment [Hee08]. Enchantment [Hay06]. enclosed [End00]. Encounter [Dod06, tF04]. Encounters [Ful01c, Mus03, LM02]. Encyclopaedia [Sho03]. Encyclopaedic [Joh03a, Yeo01a]. Encyclopedia [Coo07a, Mur99, Yeo04, Bur03, tH02]. End [Sta03, Hol00]. Ende [Man01b]. Endersby [McO09]. Endocrinology [Nor07]. Energy [Her06, Yos00]. Engel [Dea09]. Engine [Can04c]. Engineer [Dou06, Bai03, MT07]. Engineering [Aga03a, MS05b, Mor09b, Rei07a, SD00, HH00, Pap06]. Engineers [Can01b, Emm01, Fan02, Hun02, Joh02c, Dod01]. England [Can04a, Coo03, Dig08, Dix00, Dol02a, Goo06a, Gou06, Hob06, Ren07, Ril01b, Smi09a, Sum05, Tho03a, Tre06, Whi04, Wor07, Alb02, ABB01, Ash04a, Ash03, Bic03, Cro01b, Dur05, ED06, Gol04, Gou99, Has09, Hay02a, Joh00, KG04, Lev02, Mac00c, MF06, Roo07, Sch06b, Suz06, Wri01]. English [Alb01, Alb08, Bar01b, Ber06b, Ber06a, Bro02e, For05, Hei08, Kas02, Lin05, Mel01, Mos08b, Noa09, Rut08, Sir09, Smi02a, oAEN04, Arn06b, BI06, Clu06, Dun99, Jan00, Kne00, ML00, Wea00]. English-Speaking [Lin05]. Englishmen [Rit09]. engraving [Hig04c]. Enigma [Cop04]. Enlightened [Ril02, CGS99, Hee08]. Enlightenment [Bax09, Ber08, Coo07a, Del09, Eag08, Fri03a, Fri03b, Gol09, Jac09, Joh03a, Jor01, Liv05a, Nay09, Out01, Out04, Smi06e, Spa05, Ste07, Ter08, Wer09, Del06b, Gol07, Isr01, Jon08, Yeo01a, BVB08, Bro02a, Bur03, EM06, Gue00, Rei05b, Ris02, Sha08a, She00, Wil03, Wit07, ZH06]. Enriching [Fin02, Smi01a]. Enrique [Goo02d]. Enrollment [Joh04c, Mor02b]. enseignement [Chi01b, Chi03c, Hul00, HB02]. enter [Can01a]. Enterprise [Dod02, Hos01]. Entities [Goo08a, Ara06]. Entomological [Cla04a]. entre [Maz04a]. Entropic [Coe09, Kra08b]. Entropie [Coe09]. Environment [Bar01b, Cla07a, Cla04b, Haf08, Hal04b, Lam02, Mor01a, CKM99, Dun99, McC02, Tay04, RV04]. Environmental [Hal03, Ank01]. Episodes [Aab01, Goo05, GG09, GP07, Mcc05, BP01].

Epistemological [Olb02, BFR00, Sch06a]. Epistemology [Whi04, Lev02]. Epitome [Mur02, Car01b]. equation [Hof07]. Equations [Coh07, Gra08a]. Era [Mag06a]. Eradication [Web09, Cue07]. Erasmus [Far07a, SA05]. Erfindung [Smi06c]. Erhard [Set04]. Eric [Nea00b, Goo02b, Nea00b, Tre06]. Ernan [Han07a]. Ernest [Ste04b, Val05]. Ernst [Web08, Wei02b, Zig04]. Ernst-Haeckel-Haus [Zig04]. Ernst-Haeckel-Haus-Studien [Wei02b]. Erudition [De 07, PS05]. erudizione [Pug09]. Erwin [Sta03]. Escola [Fox05]. Escuela [Fox05]. Essay [Gou06, Mer08, Twe09, Had06, KG04]. Essays [Bar01a, Dix03a, Fis04, Goo03, Han07b, Kra02, Kra07, Mor07a, Pic07, Por03, Sle02, Sle05, Smi08e, Ste05b, Ste09b, Tim08, Wee07, BBW05, BLR01, Deb06, GL00, HS98, HHAH01, KCH00, MS02, Pur01, Ris07, SM05b, Wal04, WPM05]. Essential [Cop04, Hod06b]. Establishment [Her06]. Estado [Fox05]. Estate [Cro05a, Pad01, Ske04]. Esther [Por03]. Esther-Mirjam [Por03]. Estudios [Med00]. Ethnography [Dod02, Hos01]. Ethnology [Bro03a]. Ethology [Dew07, Bur05]. Etudes [Chi01b, Cor05, Hul00]. Euclid [Ste04a, EoToBB01]. Euclides [Ste04a]. Eugenics [Wal06b, Wal06c, Ric03]. Eugenios [Pat07]. Euklid [Høy05]. Euler [Jon04, Jam02]. Eumo [Goo02d]. EUR [Eag05a]. EUR144.00 [Mos08b]. EUR149.00 [Dea09]. EUR17.50 [Ful09a]. EUR20.00 [Bur07a, Hoc09]. EUR22.00 [Pug09]. EUR22.50 [Web08]. EUR25.50 [Edg08, Mer08]. EUR28.00 [Fin09]. EUR29.80 [Bra07, Bru09]. EUR32.30 [Fox07]. EUR48.00 [Coe09]. EUR51.00 [Mac08c]. EUR75.00 [Edd07a]. EUR89.00 [Mac09a]. EUR99.00 [Ter08]. Euro [Chi01b]. EUR100.00 [Ste04a]. EUR105.00 [Nut05]. EUR112.00 [Gui04].

EUR113.00 [Gol05]. EUR118.00 [Fie05]. EUR138.00 [Tso05]. EUR14.00 [Fox04]. EUR17.50 [Chi03b]. EUR23.00 [Chi04b]. EUR23.50 [Chi03c, Cor04b]. EUR232.00 [Fin03a]. EUR24.00 [Ade04, Zig04, Spr06]. EUR248.00 [Smi06c]. EUR25.90 [Vid05]. EUR33.00 [Coo05a]. EUR33.50 [Smi04]. EUR35.00 [Iro04, Nes07]. EUR37.00 [App05b]. EUR45.00 [Ade06]. EUR46.00 [Cor05, GG04]. EUR49.00 [Gin04, Kus03, Pan06]. EUR49.80 [Nes06a]. EUR59.00 [Maz04a]. EUR59.80 [Hop04]. EUR68.00 [Sta04]. EUR73.83 [Ric04]. EUR75.00 [Hig05]. EUR82.24 [Høy05]. EUR83.00 [Sta03]. EUR90.00 [DM05b]. EUR93.50 [Man04]. EUR95.00 [Hen03b]. EUR97.00 [Kle04]. EUR98.00 [Sta03]. Europe [Das09, De 07, Dew01, Edd09, Gou09a, Hen04c, Hug08, Hut09, Kha06b, Kle04, Luc05, Mac09a, Man04, Mor07b, Pag03, Ril02, Smi08c, Van02b, Wil09b, CGS99, Coo07c, Cro01a, EG04, GS99, Kri06, KM06, Mil00, NG01a, NG01c, Ols08, Ost02a, Ost02b, PS05, Rel02, TR06]. European [Ack08, Aga04, Bro01b, Coo07a, Edd06, Edd07a, Gou09b, Hal04a, Hen04b, Hoc08, Joh07a, Jor01, Kus03, Nes02, Pye06, Rei07b, Siv06, Wei02a, Wer09, BVB08, Cla07b, CSML04, Dea01, Fre99, Joh02a, Loc07, NG04]. Eutociu [Ber06b, Ber06a]. Eutocius [oAEN04]. Evangelical [Ste06, Tur06b, Fyf04, Siv05]. Evangelicalism [Fyf03b, Ast01]. Evangelicals [Fyf02a, LHN99]. Evans [Eag08, Goo09a]. Evelyn [Pal02]. Everett [Chi06, AM01]. Every [Fit03, Har06]. Everything [Hun06, Mah03]. Evidence [Sen06a, Elz06, Lee02]. Evo [Bow07, Amu05]. Evo-Devo [Bow07, Amu05]. Evolution [Ade05, Ade08, Alb03b, BS01, Ben04a, Bow06a, Daw06, Dig09, Olb02, Qui08, Rad03b, Rob05, Ske04, Web06, Web07, Asm01, BFR00, Bru01, Cla08, Fic04, Key01, Kin05, Moo02, NS00]. Evolutionary [Bow07, Ghi01, Goo01, GWT00, Her05b, Amu05, Rad00b, Zim00a, BHR00]. Evolutionary-biology [GWT00]. Evolutionsbiologie [Har02d, BHR00]. Evolutionsmodelle [Web08]. Evolving [Wee09a, BVN07]. Ewin [Smi03a]. Exact [Har00a, Maz00]. example [Rup00a, Tay06]. Examples [Pen06]. excavation [Fre01]. Exceeding [Sta06a, Vic07]. Exchange [tF00, Kno08, Urr07, CC07, Coo07b, Hen05a, Hun07b, Wat97]. Excise [Sum05, Ash03]. Excised [Can02b]. Exclusion [Sta07a]. exegesis [Sco04]. Exemplary [Ste09a, CLW07]. Exercises [Twe09]. Exhibition [Fis04, For04, For05, Fri04, Har05a, Lar01, LFT02, Mai09, Mor98, Pur01]. Exit [Can01a]. Exotic [Spa03, Rob02]. Expanded [Sta03, EC00]. expectations [Nor07]. expedition [LS05, MCS09, Dri05]. Expeditions [Big06, Pan02]. Experience [Hen06b, Pen06]. Experiences [Ric09b, Gol01]. Experiment [Goo06b, Har05a, Kra02, Mei00, Par01b, Tim05, Tim08, Deb06, Fra02, How00, Mor98, HS98]. Experimental [Dol01, Far00b, HS98, Kra02, Mor01c, BC01, Can01a, Ful00b, UU07]. Experimentation [Fra05, Ram02, HL00, Rad03a]. Experiments [Car06, Edd07a, Hen05b, KLG03, Pal09, Spo02, Sta01b]. Expert [Goo06a, Sch03a, Gol04, Hil00]. Expertise [Goo09a, Tre06, Ash04a, CE07, Gra07a, Mae06]. Experts [Aga03a, HH00, Bow06b]. Explain [Pal00, Tha99]. Explaining [Wil09d, Tau08]. Explanation [Cha01, Fet00, Hoc03]. explanatory [Kim05]. Exploration [Bee01, Haf08, Mor07b, Rob08, Siv06, Bui07, Dri01]. Explorations [Van05]. Exploring [Aga04, Hen04d, Sta03, RP02a]. Exposed [Pro08, Tuc05]. Exposing [Mül01]. Exquisite [Gol01]. extension [Fre01]. External [Aga04]. extract [Sum01].

Extraordinary [Top02, Sec00]. Eye [Can09, Gou09b, Jar08a, Mac08b, Mor05b, Cla07b, Ott08, Sch07, Sch04b]. Eye-Witness [Mor05b]. Eyewitness [Pro08, Tuc05]. Ezio [Har00b].

F [Ade08, Aga08a, Biv06, Biv08, Bru06, Coo06a, Das07, Doo04, Emm01, Gor02, Kre06, Sil03, Smi00a, Smi08f]. Fa [Dod06]. Fa-Ti [Dod06]. Fabulous [Dol04b]. face [BVB02, Chi03b, Jam08]. Facing [Fre04d, Hen02a, Man06, Wei01]. facing-page [Hen02a, Man06]. Facsimile [Can02b]. Fact [Dol04b, Sle09b]. Factory [Har06, Fit03]. faculty [Zim00b]. Failed [Aga01a, Lig02, Sco98]. Faith [Gra08a, Jam03, Coh07, Rus00]. Fakultät [Wei02b, Zim00b]. Falcon [Dep07, Mac06a, Mun08]. Falk [Olb02, Nes07]. Fall [Gau09, Har07b]. Fallible [Edd04a]. Fame [Hay02b, Mac00d]. Family [Dig08, LS05, Suz06]. Fan [Dod06]. Fantasies [Sle09a, McL09]. Fara [Gil09, Hig03a]. Faraday [Can01b, Jam03, Mor01c, Mor09b, Twe09, FJ08, Ful00a, Jam08, MT07, Rus00]. Farber [Sle01]. Farm [Har06, Fit03]. Fascist [Wei02a, Gas98b]. Fasti [Har01a, Gee00]. Faszinationsgeschichte [Coe09]. Fate [Edd05b, Van04a, Coh02, NP02]. fated [Spa05, Wil03]. Fathoming [Haf08]. Fauque [Cro05b]. favourable [Spo02]. Fe [Mor01c, Smi02a]. Feenberg [Aga05]. Feferman [GG06a]. Feingold [Maz07a]. Felicity [Ben02]. Felix [Bee01, Siv08]. Fellows [Mor05b, Mar04]. Female [And03a, NS01]. Femmes [Chi03c, HB02]. Ferdinand [End04b, Har01c]. Fermentations [Wag05]. Fernand [Spr06]. Fernando [Far06]. Ferry [Cro05a]. Fetzer [Cha01]. Feudtner [Stu06]. Fevers [Web09, Cue07]. Feynman [Kra08a, Kai05b]. Feza [Brö02b]. FF [Big01, Chi01a, Chi01b, NG01b]. Fichman [Web06]. Fiction [Dol04b, Sle09a, McL09]. Field [Gar03b, Jor02, Ken08, LS05, WC02]. Fielder [Pen04]. Fields [Ham01a, Jor02]. Fight [Mor05b, Pen04, TF02]. Fighting [Lam02, Swe01, Wal03, Rus01]. Figuring [Jor08, SL06]. Filippo [Han07b]. Films [Sle09b]. Filosofia [Goo05]. Filosofici [Man00b]. Filth [Ham07a, Bar06]. Finches [Dea05, Key02]. Finding [Far00c, Sle01]. Fine [Ewi01, Smi03a]. Finger [Van05]. Fingerprinting [Sen02b, Col01a]. Finish [Gan07, Rea05]. finite [Val05]. Finn [Mül01, Pal06]. Finocchiaro [Han07a]. Finucane [Lea04]. Fire [Aga00, Cla04b, Edd05b, Roo08, CKM99, NP02, New03, RH97]. Firenze [Pan06]. Firework [Wer04]. First [Can02b, Cro04, Cro05a, Goo09b, Gui05, Hei08, Nea00b, BI06, Fer03, Mor01c, Nea00a, Wil97, Sel08]. Fischer [Her05b]. Fischler [Art04]. Fish [Smi03a, Ewi01]. Fisher [Mor01c, Rob09a]. fishery [Dry09]. Fitness [Wal06c]. Fitzgerald [Har06]. Fitzroy [Fyf03a]. Five [HW01a, Sch02a]. Flamsteed [Nea00b, Nea00a, Wil97]. Fled [Tur02a]. Fleming [Nay09, Ris01, Bur07c]. Flesh [Smi06e]. Flight [Hem02b, Rol06, Bed03, GR00]. Florian [Hoc09]. fluens [Mcg06]. fluens/fluxus [Mcg06]. fluorine [GO09]. fluxions [Ram01]. fluxus [Mcg06]. Folkerts [Ste05b]. Folletos [Med00]. Food [Fin02, Wag05, Smi01a]. Footprints [Fri04]. Forbes [Nea00b]. Force [Bou02, Gui04, Maz07a]. Forces [Yam04a, Row02]. Ford [Hen05b]. Foreskin [Hou07, Dar05]. Forest [Hag04, CW02]. Foreword [Gar03b, Haf08, Rud04b, Tot04, Wil06, Hor03, Sta03, Van04a]. forgeries [Hig03b]. Forgotten [Cha05, Eag05a, AO03, Kin01]. förlag [Hes01]. forma [Mcg06]. formae [Mcg06]. Formal [Sta09a, FK06]. Forman [Pag01, Tra07, Tra01]. Format [Fox04]. formation [Ver03]. Formative [Ful01a, Sta03, Hac00, HS00]. formazione [Cor04b, Gov02]. Forschungen [App05b]. Forsythe [Dig08]. Fortun [Wil09c]. Forward [Edd05b]. Fossil [Fin03a, SFS^{+}00]. Fossils [Dea05, Key02, Rup08, Van04a, Coh02, O’C07]. Foundation [EB07, Ras06, Spa02, Ter08, Hör78]. Foundations [CW02, Gau09, Hag04, Har07b]. founders [Tho07]. Founding [Dew07, Bur05]. fountain [Wer01]. Four [Ade06, Gil09, Far09b]. Fourth [Cro05a, Fyf02b, Pad01, Ske04]. Fowler [GS02]. Fox [Har00a, Hen01, Pal02, Pye06, Sta07d]. frameworks [Ili00]. francaise [Chi04b, Pye06]. France [Bur07a, Chi03c, Cor06, Fox04, Gil08, Hoc08, Liv05a, Qui04, Ren07, Rob07, Wil09a, Bro02a, Cro01c, Cro07, Die06, Gil04, McT05, Pap06, Sch06b, Sil01, Sim05, Win09b]. Frances [Nea00b, Win08, Nea00a]. Francesc [Luc05]. Francesca [Elm09]. Francis [Dea09, Gou06, Hen02a, Man00a, Man06, Ren09, Sha02, Wal03, Wee07, Bru08, Bur01, Gau01, Hen02a, KG04, Man00a, Man01a, Man03, Man06, SM05b, Swe01, Zag01]. Francisco [Sle02]. Frank [Can01b, Har08, Jor02, Mor09b, Sen06a, Web09, Mor09b]. Frankenstein [Har05a, Hig04a, LFT02, Mor98]. Frankfurt [Far07c, Joh02b]. Franklin [Goo06b, Hen05b, Sta07a]. Franz [Nes01, Ste04a, Van02a]. FranzSteiner [Eag05a]. Frasca [Por01]. Frasca-Spada [Por01]. Fraunhofer [Mor02c, Jac00a]. Frayling [Boo07]. Frayn [Kra07]. Fred [Smi07a, Gre05, Mit05, Mor03b]. Frèdéric [Jor02, Ram02]. Frederick [Hod01, Hod06a]. Free [NG04]. Freedberg [Fin03a]. Freeman [Mor07c, Sta03]. Freemasonry [ED06]. Freiburg [Coe09]. French [Bax09, Bro01b, Car01a, Hul00, Out04, Ree08, Sch01, Spa05, Sta09a, BV00, BVB02, BVB03, Car07, Cha02b, CGGT06, Fau03, Hal04c, HP00, HB02, Ris02, Roc01, Rou07, Sha08a, Spa00, Wil03]. Frenkel [Hug00, Fre96]. Fresh [Pye09]. Fresnel [Lev00]. Freudenthal [Sam07]. Fried [Cuo02]. Friedman [Ben02, Can08a, Cha03, Ful01b]. Friedrich [Alb03a, Kni04, Kra02, Tso05, Cha02a, Hof07]. Friendship [Sta07b, Mil05]. Fritz [Fin02, Smi06c, Smi01a]. frog [Can01a]. Front [Ano09b, Ano09d, Ano09f, Ano09h]. Frosh [Can07]. Frost [Rei05a]. Frühe [Web08]. Fry [HW01a, Sch02a]. Fuchs [Hen05b]. Fuegians [Dea05, Key02]. Fugitive [Bro02e, Car99]. Fuller [Dix09, Gat05, Hay03]. Fullick [Jam03]. Fullmer [Dol01]. Function [Van05]. fungi [Jar08b]. fungifunguli [Jar08b]. Furqan [EB07]. Fyfe [Ric09b, Tur06b].

G [Big01, Cha01, Chi02b, De 07, Deb06, Dix03a, Fet00, Ful09b, Fyf02a, Goo08b, Hen07a, McL09, Nea00b, Nut05, Rad00a, Ric01a, Sle09a, Smi06c, Tim08, Wil09b]. G. [Cam08, Dix03a, Doo05, Hes03, Hsu06, Nut04, Smi09b]. GA [Jen01, Kno01]. Gabler [Har08]. Gabriel [Bur07a]. Gabrielle [Goo03]. Gad [Sam07]. Gaetana [Maz07b, Pet09]. Gage [Hay02b, Mac00d]. Galan [Edg05, Kle04]. Galdeano [Goo02d]. Gale [Sha09]. Gale/Cengage [Sha09]. Galilei [Han07b, Hen05b]. Galileo [Han07a, Han07b, Hei08, Hen05b, She07, Bia06, BI06, Fin05b, McM05, Nay03, Pal09, SA06, Wal04, Wil08]. Galison [Hem02b, Ren00]. Gall [Van02a]. Gallery [Bro03c, Jor02, NFJ97]. Galluzzi [Pan06]. Galton [Ren06, Wal03, Wal06c, Bur01, Swe01]. Galway [Nes06d]. Gamper [Smi04]. Garden [Haf08, Sch01, End00, RV04, Spa00].

Gardens [Raj09, Whi05]. Garland [Chi06]. Garner [Rad00b]. Gascoigne [Mac00b]. Gasentladungsforschung [Nes07]. Gasman [Wei02a]. Gaston [Til04]. Gates [Bro05a]. Gaukroger [Bro08b, Sha02]. Gave [Hod06b]. Gay [Aug09]. Gayon [Bur07a]. Gaze [Siv08, Arn06a]. Gee [Har01a]. Gegenbaur [Her05b]. Gegenwart [Hör78]. Gehennical [Edd05b, New03]. Gender [Hob06, Jor08, Maz04b, SL06]. Gene [Olb02, Pal02, BFR00, Kel00]. General [Set04, Mor06a, WS^{+}01b]. Generation [Brö04b, MW07, Rad03b, Sta09b, BS01, Har02b, Smi06b, Sta08a, Str00]. Generations [Fre04c]. Genesis [Ris07, Rob05, Ste09b, Moo02]. Genetic [Har03c, Mar05, WO01, GHL03]. Genetics [For08, Rob05, Wil09c, Moo02]. Geneva [Coo05a, Rud04b]. Geneva/Cours [Rud04b]. Genève [Rud04b]. Genius [Far07a, Hes04, Hig03a, MW09a, Bek01, Far02, HD02, Sch08b, SA05]. Genomics [Gan07, Wil09c, For08, Rea05]. gentility [Maz04b]. Geoffrey [Eag05b, Ner05, Tot04, Lip08, Smo08, Sta07c]. Geographical [Cor00, Eds04, Edn97, Liv05b, LS05, Loz00]. Geographically [Gol09, Wit07]. géographie [Rud04b]. geographies [Nay05, WHF08]. Geography [Bee01, Dri01, Edd03a, LW05, Qui02, Rud04b, Sha03b, Smi08a, Tyb04, Dri05, Liv03, May05b, Rup00b, Wit01, BJ00]. Geohistory [Coh09, Mor06c, Rud05a, Rud08]. Geological [Fin03a, Mor05b, Mor07a, Mor07b, Bur07c, Fre01, SFS^{+}00, Mor06c]. Geologist [Mor06c, Her05a]. Geologists [Mor06c, Rud05b]. Geology [Alb01, Edd05a, Mor06c, Mor06b, Rob03, Her04, Kne00, Lyn02, Rud04a, Tor02, O’C05]. Geometer [Rav01, vD05]. Geometrical [Ale02, Ste05a]. Geometrico [Han07b].

Geometrie [Cor04a]. Geometry [EB07, GG02, Mag01, Val05, AH02]. Geophysical [Mor07c]. Georg [Coo05a, Rud04b]. George [Edd05b, Fra09, Fri03a, Fri03b, Rob09a, Ack08, Gue00, New03, Tso07, Whi04]. Georges [Elm09, Sir06]. Georgina [Cro05a]. Gerald [Sta03, Wal03]. Gerard [Bro04a, McG08, Hig01, Hig07]. Gerhard [Wer04]. German [Gor09, Har01c, Har04c, Hör78, Kre03, Men00, Nes02, Rei07b, Smi08f, Zic00, BSD08, BHR00, Gos00, Hen07b, Hoc03, Mei00, Sch00, SEM04, Sta01a, Tei00, Zim00b]. Germany [Chi02b, Dea07, Kim07, Nes06b, Pan08, Tei00, Tur02a, Wer04, Har05b, Hau03, Hel05, Jac06, Sch06a, Sto00]. Germs [Ham07a, Kip04, Bar06, Wor00]. Gerovitch [Cla06]. Geschichte [Har02d, Hör78, Par01b, Sch00, Smi02b, Sta04, Wei02b, Zig04]. gesellschaftlicher [Nes06a]. Ghiselin [Sle02]. GI [Mor07c]. Gianelli [Bru06]. Gianna [De 07]. Giant [Ken08]. Giants [Ben06, Hue03]. Gibbons [Goo09b]. Gibson [Smi08f]. Gidlunds [Hes01]. Gift [Tur02a, Win08]. Gillian [Joh05b, Sle05]. Gillies [Aga01b]. Gillispie [Cor06, Har00a]. Gimelli [Wee07]. Gingerich [Gol05, Hen05b, Hen06a]. Giovanni [Qui08]. Gisela [Dea09]. Gispert [Chi04b]. Gitelman [Ner05]. Giuliano [Goo05]. Giuseppe [Mel01, ML00]. given [Aga04, Bro03b, Chu02, Cro05b, Edg05, Fox05, Goo05, Her05b, Hes01, Hig04b, Hig06, Hig08, Hug00, Jam01b, Kha06b, Kle05, Man01b, Mor07c, Pye06, Rud04b, Set04, Sha03b, Sir09, Whi05, tF00, Goo02c, Goo02d, Mag06a, Med00, Ris01]. Givens [Eag06b, Kem09]. glacial [Fin04]. Glands [Log08, Sen06b]. Glass [Ben00]. Glen [LS05]. Glenn [Bru06]. Glick [Ade08]. Glimpse [WM08, Elz06]. Global [Far07b, Hal06, Nay09, FJC06, Wea03]. GNT [Nes07, Par01b]. GNT-Verlag [Nes07, Par01b]. God [Dig09, Eag07, Fyf03b, Gra08a, Pet09, Ast01, Cla08, Coh07, Maz07b, Nor05]. Goddard [Col00, Zen98]. Godly [Ste06, Siv05]. goes [Fin03b]. Goethe [Kni05, Ham01b, Ric02b]. Gohau [Bur07a]. Golan [Goo06a]. Golden [Coo07b, Kno08]. Goldman [Mac05]. Goldstein [Goo02c]. Golem [Pem07, CP05]. Golinski [Bro04a, Edd04a, Nay09, Ril02]. Good [Goo02b, Sha07, Wal03, HLS00, Jon06b, Swe01]. Gooday [But06, Sta07d]. Goodman [Mar05, Tim03a, Tim03d]. goods [BC01]. Gorilla [Dig09, Cla08]. Goschler [Nes01]. Gothic [Eag06b, Giv05]. Göttingen [Nes06c]. Gouk [Dix00]. Gould [Bow04, Gar03b, Van04a]. Gourko [Ghi01]. Governance [Gan07, Rea05]. Government [Sch02b, Dra00]. governmental [Tur03a]. Govindoo [Raj09, Nol07]. Govoni [Cor04b]. Gowan [Tur08]. Gower [Fyf02b]. Gracewing [Hur07]. Graeme [Sta07d, But06]. Grafton [Pag03, Van02b]. Graham [Hen02a, Joh09, Man00a, Man06, Smi00a, Smi03b]. Granada [Man00b]. Grand [Shi03b]. Grandin [Ras06]. Grant [Gau08]. Graphics [BDLM07, Elm09]. Grasp [Vic07, Sta06a]. Grasseni [Shr09]. Grattan [Har00a, Pet07a]. Grattan-Guinness [Har00a, Pet07a]. Gravitation [Hen04e]. Gravitational [Aga06, Col04]. Gravity [Aga06, Col04, Hen04e]. Gray [GG09, Maz02b]. Great [Art04, Fis04, For04, Ham07a, Sho03, Bar06, Mur99, UU07, GS02, Pur01]. greater [Cai00]. Greece [Nut04, Tot04, Wil09d, Sas01, Tau08]. Greek [Cam08, Doo05, Hsu06, Rob09c, Cuo07, Llo03, Llo04, Llo06]. Green [Mor01c, Mor08, Smi02a]. Greene [Bro03b]. Greenwich [Eag06a, Hig07, NHA^{+}02, RN09, vCA^{+}05]. Greenwood [Can08b]. Gregg [Mun07, Stu09]. Gregory [Mor09b, Smi07a, Dol04a, Smi08d]. Greta [Ade04, Ren06]. Grew [Gar03a]. Grip [Doo05, Llo03]. Grob [Hig08]. Gross [Ful09b]. Grossbritannien [Har01c]. ground [GS02]. Groundbreakers [Jam03]. Group [Hig06]. Growth [Mok06, Mea04]. Grubb [Ben00, Gla97]. Grundlagen [Cor04a]. Guagnini [Mac06b]. Guerrini [Fri03a, Fri03b]. guidance [Ste09c]. Guide [Big07, Can01b, Cla04a, Ful03, Fyf03a, Lin05, Mor06d, Bro01a, Hei05, Hes00, RP02b]. Guided [Ful09b, HG07]. Guides [Can08b]. Guildhall [Can01b]. Guillem [Pye09]. Guillem-Llobat [Pye09]. Guillermo [Fox05, Fox05]. Guinness [Har00a, Pet07a]. Gummett [Coo02b]. Günergun [Brö02b]. Gunpowder [Wer04]. Gustav [Hes01]. Gustave [Pap06]. Gustave-Adolphe [Pap06]. guten [Man01b]. Guy [Edg04].

H [Aga02, Bre05, Bro04a, Cha01, Cor00, Dol02b, End04b, GS02, Gou09a, GG04, GG09, Hen06b, Jam01b, Kir05, Kni02, Mag06a, Mar05, Mor09b, MW07, Nes09, Ram02, Rud04b, Rut08, Sle03, Sle09a, Smi02a, Smi02b, Ste09a, Tre06]. H. [McL09, Smi00a, Ste04a, Wag05]. H.-B [Rud04b]. H.M.S. [Key00]. Haber [Fin02, Smi01a]. habits [Liv05b]. Hackett [Ric08]. Hacking [Rad02]. Hacohen [Ful01a]. Hadot [Mer08]. Haeckel [Web08, Wei02a, Wei02b, Zig04, Gas98b]. Hahn [Fox07, Smi08f]. Haim [Ful01a]. Haines [Gou03b]. Hakfoort [Can00]. Haley [Lev07]. Hall [Goo08b, Lüt04b]. Haller [Edd07a]. Hallyn [Spr06]. Ham [Ken07]. Hambledon [Eag07, Fra09, Goo06c, Har03a]. Hamblin [Tur09a]. Hamilton [Bro03c, Jor02, Woo06]. Hammond [Dea06a]. Hand [Mac09a, RSD07]. Handbook [Ric08, Hac08]. Hannah [Aug09]. Hanne [Sta08b]. Hanns [Ste07]. Hans [Hig08, Olb02, Smi02b, Sta03, Fra05, GG06b].

Hans-Jörg [Olb02]. Hans-Werner [Smi02b]. Hansen [Ham07b]. happened [Aga08b]. Haraway [Ree07, Har04a]. hardback [Abr08, Ack08, Ade06, Ade08, Aga01a, Aga02, Aga03a, Aga03b, Aga04, Aga05, Aga08a, Alb01, Alb03b, Alb08, And09, Ash04b, Bar01a, Bar07, Bar08, Bax09, Bek06, Ben02, Ben06, Ber06a, Ber08, Big07, Bla03, Boo07, Bow06a, Bow07, Bra07, Bre04, Bre05, Bre06, Bro04a, Bro05b, Bro06a, Brö02b, Brö04b, Bro08b, Bro09, Bro03b, Bro03c, Bro01b, Bro02e, Bro04c, Bud09, Bur08, But06, Cam08, Can08a, Can04a, Can07, Can08b, Car01a, Car06, Cha01, Chi08, Chi02b, Chi04a, Chi06, Chu02, Cla04a, Cla06, Cla07a, Cla09, Coe09, Coh09, Col02, Coo07a, Coo03, Coo04, Coo06a, Coo06b, Coo02b, Cor06, Cro04, Cro05a, Cuo02, D’A03, Das07, Das09, Daw06, Daw07a, De 07, Dea05, Dea07, Dea08, Dea09, Del06a, Del08, Del09, DM05b, Dep07, Dew07, DiM07, Dig08, Dig09, Dix05]. hardback [Dix08a, Dix09, Dod02, Dod06, Dol01, Dol02a, Dol02b, Dol02c, Dol04b, Dol04c, Doo04, Doo05, Dou06, Dri03, Duf09, Dun08, ETV04, Eag05b, Eag06a, Eag06b, Eag07, Eag08, Edd03a, Edd03b, Edd05a, Edd05b, Edd06, Edd07a, Edd07b, Edd08, Edd09, Edg08, Eds04, EB07, Elm09, Elm02, End04a, End04b, Far07a, Far07b, Far07c, Fin03a, Fin02, For03, For04, Fox05, Fox07, Fre04d, Ful01a, Ful02, Ful03, Ful09a, Ful09b, Fyf02b, Fyf02a, Fyf03a, Fyf03b, Fyf05, Fyf07, Gar03b, Gau08, Gau09, Ghi01, Gil08, Gil09, Gin04, Gli06, Gol05, Gol09, Goo04a, Goo06a, Goo08a, Goo09a, Goo06b, Goo06c, Gor09, Gou03a, Gou06, Gou09b, Gou03b, Gra08a, GG02, GG04, GG06a, GG06b, GG09, Gra06, Gra08b, Gre02, Gui04, Gui05, Haf04, Haf08, Hal03, Hal06, Ham07a, Ham08b, Han07a]. hardback [Han07b, Har05a, Har03a, Har02c, Har01b, Har03c, Har06, Hay02b, Hay03, Hay06, Hei07a, Hei08, Hem02a, Hem02b, Hem03, Hen01, Hen02a, Hen03a, Hen04a, Hen06a, Hen06b, Hen07a, Hen03b, Hen04d, Hes03, Hes04, Hig03a, Hig01, Hig05, Hig07, Hob06, Hod06a, Hod06b, Hop01, Hor03, Hou07, Høy05, Hsu06, Hug07a, Hug08, Hug09, Hul07, Hun02, Hun05a, Hun05b, Hun06, Hun07a, Hur07, Hut07, Hut09, Iro04, Jac09, Jam01b, Jam03, Jam05, Jar08a, Jen01, Jen02, Joh03a, Joh07a, Joh07b, Joh09, Joh02c, Joh02d, Joh03b, Joh04b, Joh04c, Joh05c, Joh06a, Joh06c, Jon02a, Jon06a, Jor01, Jor02, Jor06, Jor08, Kai05a, Kem09, Kha06a, Kim05, Kip04, Kir05, Kle04, Kni03, Kni04, Kni05, Kni06a, Kni08, Kno08, Kre06, Kus03, Kus07, Lea04, Lea05, Led08, Leh08, Leh09, Lev07]. hardback [Lin05, Lin07, Lip08, Liv05a, Loc02, Log08, Luc05, Lüt04a, Lüt04b, Lyn09, Mac08b, Mac08a, Mac09a, Mac07, Mac08c, Mag06b, Man04, Man01a, Man06, MW09a, Maz02b, Maz02a, Maz07a, Maz09, McG08, McO09, Mel01, Mer08, Mil02, Mil04, Mir07, Mok06, Mor03a, Mor06a, Mor06c, Mor06b, Mor07a, Mor07b, Mor07c, Mor01a, Mor06d, Mor06e, Mor01c, Mor08, Mor09b, Mos08a, Mos08b, MW07, MW09b, Mun07, Mun09, Nay09, Ner05, Nes06a, Nes09, Noa04, Nut05, Nye09, Olb02, Out01, Pag01, Pal00, Pal02, Pal06, Pan08, Pau09, Pem04, Pem07, Pen04, Pet07a, Pet07b, Pet09, Pet08, Pic07, Plo05, Poo09, Pre03, Pro08, Pye09, Qui02, Qui08, Rad00a, Rad03b, Ras06, Ree08, Rei07a, Rei07b, Ren06, Ren09, Ric09a, Ric09b, Ric01a, Ric08, Ril01a, Ril02, Rit02, Rit09]. hardback [Rob03, Rob05, Rob07, Rob06, Rob08, Roo08, Rud04b, Rud07, Rup08, Rut08, Sam07, Sch08a, Sch02a, Sch02b, Sec08, Seg05, Sen02b, Sen06a, Sha07, Sha09, She07, Shi03b, Sho03, Shr09, Sir06, Siv06, Siv08, Ske04, Sle02, Sle03, Sle05, Sle09a, Smi00a, Smi01b, Smi02b, Smi03a, Smi04, Smi06d, Smi06g, Smi06c, Smi07a, Smi07b, Smi08b, Smi08a, Smi08f, Smi08e, Smi08c, Smi09b, Smi09a, Smo08, Spa03, Spa05, Sta03, Sta04, Sta07a, Sta09a, Sta08b, Sta06c, Sta07b, Sta07c, Sta07d, Ste05a, Ste05b, Ste07, Ste04b, Ste06, Stu06, Stu09, Sum05, Sum06, Sut03, Suz05, Tho03a, Tho03b, Til04, Tim03a, Tim03b, Tim03c, Tim03d, Tim04, Tim08, Tot04, Tra07, Tre06, Tso05, Tur09a, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van02b, Ver04, Vic07, Wag05]. hardback [Wal06b, WM08, War08, Web06, Web07, Web08, Web09, Wee07, Wee08, Wee09a, Wee09b, Wei02a, Wer09, Whi06, Whi04, Wil06, Wil09a, Wil09d, Win08, Wor07, Yam04a, Yos00]. hardcover [Eag05a, Gis09, Pag03, Top02]. Hardwicke [Hur07]. Hardy [Biv08]. Hargreaves [Bro03c]. Harkness [Wee09b]. Harloe [Hig06]. Harman [Hun05b, Yam04b]. Harmattan [Ful09a, Vid05]. Harmon [Ful09b]. Harmonics [Tyb02, Bar00a]. Harmonious [Jac06, Pan08]. Harold [Ril01a, Kno08]. HarperCollins [Dea05]. Harri [Joh02b]. Harriot [Hen01, FH00]. Harris [Brö04b]. Harrison [Biv06, Gau09]. Harro [Mir07]. Harry [Ful03, Aga06, Goo09a, Pem07, Tim05]. Hart [Fyf02a]. Hartcup [Edg04]. Hartz [Sir06]. Harvard [Ber08, Bra07, Dod06, Edg08, Fox07, Fre04d, Ful03, Goo06a, Haf08, Hal03, Hal06, Hen07a, Log08, Mac07, Mac08c, MW09a, Maz07a, Mer08, MW09b, Pal02, Pet08, Rad02, Rad03b, Rob06, Sen02b, Sta03, Tho02b, CC07]. harvest [Sec05]. Harvey [Fin03a, Hod01, Sch02a]. Harwood [Coo02b, Kim07, Mül01]. Hasok [Kle09, Kni06a]. Hate [Can07, Fro05]. Hau [Nes06b]. Haugeland [Bar01a]. Hauksbee [Bru08]. Haus [Wei02b, Zig04]. Haven [Aga01a, Dew01, Dix00, For03, Har06, Hut09, Kno08, Mor07c, Sch02b, Web09, Wee09b]. Havoc [Sta03, Mil01]. Hawaii [Fan02]. Hay [Sch02a]. Haycock [Can04a]. Hayes [Ter08]. Haynes [Har02c]. Heads [Alb03b, Asm01]. Healing [Dol02b, Kha06b, Suf07, Web09, Bau01, Ber06d]. Health [EG04, Kha06a, Kha06b, Nes06b, Pic07, Tim03b, Bas04a, Hau03, Jef01, WPM05]. Healy [Doo04]. Hearing [Tho03c, Pol02]. Heart [Tim04, Wei02d, BZ04]. Hearts [Tim03b, Jef01]. heat [Tay06]. Heath [Mar05]. Heavens [Smi08b, BZ04, Mos07, Win09b]. Hebrew [Sam07, Fre05]. Hecht [Goo03]. Hedemora [Hes01]. Hedin [Mac01a]. Heesen [Yeo04]. Hegemony [Hug08, Kri06]. Heidegger [Ful01b, Fri00]. Heidelberg [Sha03b]. Heidelberger [Kra02, Tso05]. Heike [Hop01]. Heilbron [Big07, Fyf05]. Heinemann [Chi08, Hen06a, Jam03]. Heinrich [Hop01, Men00]. Heisenberg [Kra07]. Helaine [D’A03, Plo05]. held [Hig02]. Helden [Hen05b]. Helen [Gou03b, Haf08, Mun08, Haf08, Hen04d, Sle05]. Helena [Ghi01]. Hélène [Chi04b]. Helge [Coe09, Fre04c]. Hellyer [Dea07]. Helm [Kus03]. Helmontian [Edd05b, NP02]. Helmuth [Mül01, Coo06c]. Hempel [Cha01, Fet00]. Henare [Urr07]. Henrika [Shr08]. Henry [Aga02, Bro05b, Cha05, Col00, Hes04, Jam01a, Jam01b, Jam05, Joh05b, Lüt04b, Mac08b, Maz07a, Sta03, AO03, Brö04b, DeV00, Dol02b, Fox04, Hal02b, HRR08, Moy97, Raj09, Sch04b, Sco04, Sen06a, Zen98]. Henslow [Mor03a, WS01a]. Hentschel [Gor09, Joh03b, Sta03, Yeo04]. Her [And09, Rei08]. Herbert [Col00, Hut04, Mor06c, Ren06, Ren09, Zen98]. here [Far00a]. Hereditarian [Wal03, Swe01]. heredity [Bur01]. Heresies [Ren06, LRR04]. Heretic [Web06, Slo04]. Heringman [Mor06b]. Heritage [EB07, Kni03, Mor03b, Far00a]. Herman [Iro04, Kno02]. Hermann [Set04, WS^{+}01b]. hermaphrodite [Mae06]. Hermeneutics [Sha03b, Liv02]. Hermes [Smi05b, Lin03]. Hermione [For04]. Herran [Pye09]. Herschel [Ash01, Ash04b, Bre06, CDK98, HH03, Hos03, RH04]. Herz [Art04]. Herz-Fischler [Art04]. Herzig [Lin07]. Hessen [Chi03a]. Hessenbruch [Fyf03a]. Hester [Eag06a]. Heterodoxies [Dol02b, Bau01]. Heterodoxy [BM05, Kus07]. Hettner [Sha03b, Liv02]. Hettner-Lecture [Liv02]. Hettner-Lectures [Sha03b]. heuristics [Tho07]. heute [Har02d, BHR00]. HFL75 [Rav00]. hidden [Ram01]. Hiebert [Sta03]. hierarchy [Sim02]. Higgitt [Maz07a, Maz09]. Higgs [Goo02b]. Higher [Aug09, Gay07]. Higton [Eag06a]. Hilaire [NG01b]. Hilaire-Pérez [NG01b]. Hilbert [GG04, Gra06, Maz02b, Cor04a, Gra00, Yan02]. Hilgartner [Sch03a]. Hill [Rol06, Stu06]. Hills [Coo04, Mil04]. Himrod [Lin05]. Hippocrates [Bro04c, Can02a]. Hirn [Pap06]. Hispanica [Med00]. histoire [Chi03c]. Historia [De 07, Med00, PS05, Hen05b]. Historian [Dol04c, Wil02]. Historica [Joh02b]. Historical [Aga03b, Biv06, Can08b, Dix03a, Edd03a, Edd08, Fle98, Ful02, Fyf02a, GG02, Haf08, Har01c, Har04c, Hei07b, Hun07a, Kir05, Kra07, Led08, Maz04a, Mor06d, Olb02, Qui02, Rei07b, Ric04, Ric09a, Ric01a, Ris01, Smi03b, Smi08f, Sta03, Sta09a, Tho03a, Tra07, Whi05, Win09a, WHF08, BFR00, BLR01, Cav08, CJ02, DS02, Efr07, FK06, Ham01b, Kai05c, Kip03, LHN99, MPB00, Mae06, Mag01, Mei00, Nay05, Ric02a, RV04, Rup00b]. Históricos [Med00]. Histories [BW01, Hal06, Hun05a, Ste06, Ver04, DS01]. Historiography [dC08, DS06]. historiques [Maz04a]. historische [Mei00, Par01b]. History [Abr08, Ade04, Aga03a, Aga03b, Aga04, Aga08a, Alb03a, Alb03b, And09, Ano00e, App05a, Ash04b, Aug09, Bar01b, Bea01, Ben02, Big07, BBW05, Bow04, Bow06a, Bra07, Brö02b, Bro02e, Bro04c, Bud09, Bur08, Byn08, Can09, Chi06, Chr05, Cla01, Cla05, Cla06, Col01b, Coo07a, Coo03, Coo06b, Coo06c, Coo02b, D’A03, DM05b, Dol04b, Dri03, Duf09, Edd04a, Edd07a, Edd09, Emm01, End04a, tF00, Far07b, Far09a, Fin03a, For04, For05, Fre04c, Fyf02b, Fyf03a, Fyf05, Gar03b, Gat05, Gau08, Gil09, Gin04, Goo02a, Goo05, Goo06a, GG06b, GG09, Gra06, Haf04, Haf08, Han07b, HS02, Hei08, Hem02b, Hem03, Hen07a, Hes01, Hes03, Hig05, Hig06, Hoc08, Hug07b, Hul07, Hun02, Hun05a, Iro04, Jam05, Jar08a, Jen01, Joh07a, Joh02b, Joh02c, Joh06a, Jor01, Jor02]. History [Kem09, Kha06a, Kip04, Kir05, Kle04, Kle05, Kle09, Kni02, Kni03, Kno01, Kra07, Kre06, Lam02, Loc02, Lüt04a, Mac01a, Mac06b, Man04, Man01a, Maz07a, Maz09, Mcc05, Mer08, Mor05b, Mor07a, Mor01a, Mor02a, Mor03b, Mor06e, Mor02c, Mos08b, Mül01, Nay09, Nes06b, Nye09, Pag03, Par01a, Plo05, Pre03, Pye09, Qui02, Ram02, Ras06, Rei07a, Ric04, Ril01a, Rob06, Roo08, Sch01, Sec00, Sec08, Sen02b, Shi03b, Shr08, Siv06, Sle01, Sle02, Sle09b, Smi00a, Smi08f, Smi08e, Smi09a, Söd07, Spa02, Spa03, Spa05, Spa06, Sta03, Sta06c, Ste09b, Stu09, Sum06, Suz05, Tim05, Tim08, Top02, Tso05, Van04a, Van05, Van02b, Vic07, WM08, Wee08, Wei02a, Whi06, Win08, Yam04a, Yos00, Aab01, Ade05, Aft01, AAH00a, Alb02, AA01, AM01, AAH00b, Asm01, Bai05, Bas04a]. history [BW06, BP01, Biv07, Bod06, Bow06b, BS02, BSD08, Bro02c, Coh02, Col01a, Coo07c, DS02, Dun99, Dun09, Elm06, Far09b, Fin05a, FJC06, FSJ00, Gas09, Gay07, Ger02, GL00, GRT06, Gol04, Gol98, Goo02e, GWW04, Gra07b, GP07, Had06, Hau03, HBB^{+}03, Hei05, Hes00, Hof07, Hol99, HL00, Hol00, Hör78, Hug04a, Jah03, Jan00, Joh01, Joh06b, KNRW03, Kni06b, Kuk08, Lev01, Lig02, Lov01, MS05b, Maz00, MD99, Mud01, NFJ97, NG01c, Ore01, Osl02, Ott08, Pic01, Por02, Por09, Pri07, Ras01, Rei08, Ris07, Row02, Rup00a, Sch07, Sch00, SS00, Sel00, Sha08b, Sie02, SS01, Sle07, Spa00, Sta06a, SBF00, Tei00, Top00, TDNS08, Wal04, Wat97, Wei02d, Por01, Mor09a]. Hitler [Tur02a]. HMS [Ric01b]. Hoax [Ful09a, Sok08]. Hobhouse [For04]. Hochadel [Nes06c]. Hoddeson [Aga00, Hes04]. Hofmeier [Smi04]. Hogan [Mor02a]. Hogarth [Jor02]. Hokusen [tF00]. Hokusen-Sha [tF00]. Holes [Sta07b, Mil05]. Hollar [Hig04c]. Hollinger [Hug07b]. Holmberg [Hes01]. Holmes [Ram02]. Holographic [Aga08a, Joh06b]. Holton [Sta03]. Holy [Roo08, Num07]. Home [Dig08, End04b, Suz06]. Homer [Shi03b]. homme [Cha02b, Cor05]. Hong [Hem03]. Honolulu [Fan02]. Honor [Goo03, HHAH01]. Honors [GG04, Yan02]. Honour [Pan06, Sle05, Smi08e]. Hooijmaijers [Hig08]. Hook [Ste04b]. Hooke [Hen05b]. Hooker [McO09]. Hopkins [Ade08, Aga04, Bur08, Chi02b, Dea08, Dig09, Dix09, ETV04, Gar03b, Gra08a, Gra08b, Ham07a, Hem02a, Kni02, Kni04, Lyn09, Pet09, Pro08, Rad00a, Ril01a, Sle01, Spa03, Tim03b, Tre06, Web07, Web09, Whi06]. Hopwood [Alb03a, Lau06, Mai09]. Horace [Rud04b]. Horace-Bénédict [Rud04b]. Horary [Eag06a, NHA^{+}02]. Hormones [Log08, Sen06b]. Horses [Whi06, Der03]. Horst [Bru09, Web08]. Hoskin [Ash04b]. Hospital [For03, Ste00b]. Hossfeld [Har02d, Her05b]. Hostetler [Dod02]. Hot [Coo02b, Hou06, Log08, BG02]. Houck [Log08]. Houghton [Sta07b]. Houndmills [Das09]. House [Bro04a, Wee09b, Har07a, LTGS02]. Hove [Smi03b]. Howard [Ben00, Can04c, Gla97, Mor01c, Pag01, Sta03]. Howse [Wat00]. Hoyle [Smi07a, Gre05, Mit05]. Huang [Wag05]. Huberman [Sir06]. Hubert [Edd07a]. Huddersfield [Hen07a]. Huerta [Ben06]. Hugh [Smi03a, Edd05a, Ewi01, Hem02a, Hut09].

Hughes [Aga03a, Goo03, Joh06c, HHAH01]. Hulin [Chi01b, Chi03c]. Human [Aga01a, Goo09b, Hug04b, Joh06c, Kir05, Lyn09, Ric09a, Ske04, Smi09b, Can01a, CJ02, GO09, Ken07, Key01, Liv08, Llo07, Nor07, Sco98, Van02a]. Human-Built [Joh06c, Hug04b]. humanism [May05a]. Humanist [Wil06, WR04]. Humanistic [Cla04b, CKM99]. humanités [Chi01b, Hul00]. Humanities [Nes02]. Humanity [Ham07b, Kuk06, Hul00]. Humankind [Goo08b]. Humans [Lam02, Cha02b, Rus01]. Humboldt [Far07c]. Humphries [Bro03c]. Humphry [Dol01, Ful00b]. Hundred [Sch02a, HW01a]. Hunger [Gra08b]. Hunter [Dea02, Elm02, Fyf02b, Tho09]. Hunters [Pad01, Cad00]. Hutcheson [Del09]. Hutton [Maz07a, Noa09]. Huxley [Dix06, Lig02, Whi03]. Hydrographer [Hom09]. Hygiene [Kha06a, Bas04a]. Hysteria [Sir06, DHC03].

I. [Ben00]. Ian [Kus07, Mac06b, Smi08c, Yam04b, Ham08b, Mor01b, Mur02, Rad02]. Iarosh [Sir09]. Iberian [Goo02c, CG00a]. ice [Fin04]. Iceland [Wil09c, For08, Osl02]. Icon [CK02, Dix09, Ful01c]. Iconography [Kni03, KMO05, Maz07a, Mor07c, Sir06, Ber01, DHC03, Key05]. Icy [Tso07, Rei05c]. Ida [Mag06a]. Idea [Mer08, Smi00a, Had06]. Ideal [Har06, Fit03]. Ideas [DM05b, Hes01, Hod06b, Lüt04a, Man04, Maz07a, Mos08b, Nut04, Sec08, Spa05, Kni06b, Nor07]. idées [Kim05]. Idéhistoriska [Whi05]. Identification [Sen02b, Col01a]. Identities [Sen02b, Col01a]. Identity [Bro03c, Edd03a, FK06, Gan07, Jon02a, Sta09a, HH01, Mur00, Rea05, Wit01]. Ideology [Wei02a, Gas98b]. idols [Zag01]. IEE [Emm01, Hun02]. Ignorance [Tos09, BV00, Chi01a, PS08]. Igo [Pet08]. Ihsanoglu [Bek06, Brö02b]. II [Aga03a, End04b, Hun05b, Jon02a, Man06, Mor01c, Mur00, Ste04a, Ade04, HH00]. III [Hun05b, Mor01c, Ril01a, Ste04a, EoToBB01]. Ileana [Big01]. Ilich [Fre96, Hug00]. Iliffe [Maz07a]. Ill [Spa05, Wil03]. Ill-fated [Spa05, Wil03]. Illinois [DiM07, Dol02b, Wil09b]. Illness [Hay03, Stu06, TM01]. Illumination [And03a, NS01]. Illusion [Hen04d, RP02a, Sch06a]. Illustrate [Jor02, NFJ97]. Illustrated [Dol02a, Sch02a, ABB01, HW01a]. Illustration [Bro05a, Gat02]. im [Hop04, Kre03, Nes06a, Nes07]. Image [Eag06b, Emm01, Gal97, Jor01, Rad00a, Ren00, Alt99, Giv05, McL00, She00]. Image-Making [Eag06b, Giv05]. Imagery [Smi02b]. Images [Ben02, Dig09, Fie05, She07, Smi08c, Bia06, Cla08, KM06, LRS03, SBF00]. Imagination [Bre06, Doo05, Ham08b, Til04, Kuk06, Llo03, RH04]. imagined [May05b]. Imaging [Ber01, Kni03]. Imagining [App05b, Osl02]. imitation [HP02]. Imitations [Hig02]. immunologie [Maz04a]. Immunology [Maz04a]. Immunology/Dialogues [Maz04a]. Imperatives [Edg08]. Imperial [Ank01, Aug09, Bas04a, Chu02, Fan02, Hal03, Har02c, Kha06a, McO09, Mil02, Sch02b, Urr07, Wri00, Bra08, Dod01, Dra00, Hen05a, Gay07, Hay01]. Imperialism [Gre02]. Implantable [Tim03b, Jef01]. Implications [Ras06, GWW04]. Improve [Aga01a, Sco98]. Improvement [Sch02b, Nes06d, Dra00]. improving [Kum07]. incarnations [Sco04]. Inchofer [Hei08, BI06]. includes [Hig08]. Including [Hei08, BI06]. incorporación [Fox05]. Index [Ano09o, Mac05]. India [Cor00, Siv08, Win09a, Arn06a, Bro03a, Edn97, Kum07]. Indian [Nai06]. Indigenous [Edd09, Coo07c, Kuk06]. indigo [Kum07]. individual [Can01b, Noa09]. Industrial [Coo03, Div01, Dun08, Fox05, Har06, Jac09, Jon08, Kle04, Mor08, Cos00, Cro01b, Fit03, Hor00, ML07, Mus06, NG01c]. Industriales [Fox05]. Industrialisation [Mac09a, RSD07]. Industrialisierungsgeschichte [Har01b, Tei00]. industrialization [Tei00]. industrie [Fox04]. Industrieforschung [Har01c]. Industry [Dol02a, For04, Jon09, Maz07a, Mor01a, Mor03b, Mor06d, Qui08, Ras06, Aft01, ABB01, GWW04, JS04]. inédite [Big01]. Inequality [Ree08, Car07]. inerrant [Edd04a]. Infinite [Gar03b, WB02]. Influence [Bek06, Jor02, Ham01a, Ken07, çIh04]. informal [Fan03]. Information [Aga00, Gov02, RH97]. Ingenieros [Fox05]. ingeniou [Edd02, Edd04b]. Ingrid [Gor02]. Inkster [Mac06b]. Inner [Bor07]. Innovation [Goo01, Zim00a]. Innovators [Lea04, MDF02]. Inquietudes [Fox05]. Inquiry [App05a, Mac09a, RSD07]. Insanity [Dig08, MF06]. Inscribed [Win09b]. Insects [Lam02, Rus01]. Insider [Shi03b, Ore01]. insignes [Hen05b]. instant [Mcg06]. Instauratio [Hen02a, Man06, BR00, BRW04]. Institut [Goo02d, Ren05]. Institute [Ade04, Aga01c, Aga03a, Alb03a, Ben00, Bro02e, Har01c, Hun05a, Jen01, Jor02, Kno01, Kre06, Lea05, Mac08a, Mor07c, Mor02a, Nea00b, Ram02, Rei07b, Ren06, Smi08f, Tso05, Van02b, Wal06c, Nes06d]. Instituting [CC07]. Institution [Can01b, Emm01, Hun02, Jam01a, Jam01b, Jam05, Joh02c, Mor09b, J^{+}07, UU07]. Institution/Science [Jam05]. institutional [Noa09]. Institutions [Bek06, Brö02b, Gil08, Sle02, Cro07, GL00, çG00b, çIh04]. Instituto [Med00]. Instrument [Hig01, Mei00, Pan08, Par01b, Hig04c, Jac06, Tur00, Mei00]. instrument-maker [Hig04c]. Instrumental [Che00, Joh02d, Win09b]. instrumentation [Zui05, Zui06].

Instruments [Ben02, Dun08, Goo06b, Hig02, Hig05, Hig06, Hig08, HL00, Joh02d, Pan06, Ram02, She07, Smi08c, Win08, Bai04, Ben03, Bia06, Che00, KM06, ML07, SBF00, vC02]. Insulin [Stu06]. integrating [Cav08]. integration [Cai00]. Intellect [Hug09, Tho06]. Intellectual [Han07b, Jon06a, Mac08b, Ren06, CC07, Pet04, Sch04b, Wal04]. Intelligence [Col00, Cop04, Dol02c, Ree08, Sum06, Car07, Ric00, RS02, Zen98]. Intelligent [Dix09, Com07]. Intelligibility [Hei07a, Dea06b]. inter [Hor00, May05a, Sch05b]. inter-war [Hor00, May05a, Sch05b]. Interdisciplinary [Fis04, Gou03c, Kra02, Mos08b, Clu06, Pur01, Sta03]. interests [Atm04, Mae06]. Intermolecular [Yam04a, Row02]. International [Big01, Brö02b, DM05b, Gou03b, Hai01, Hig05, Lüt04a, Man04, Maz07a, Mor03b, Mos08b, Smi08f, Spa05, Aft01, BW06, çG00b, How00]. Internationale [App05b]. internationales [Kim05]. Internationalization [Aga03b, Hor07, SS01]. internationally [Aga04]. interpretation [Goo02e]. Interpretations [Win08]. Intersections [Dea09]. Interview [Bar01a, KCH00]. Interwar [Ful01a, Hac00]. Intimate [FJC06, Nay09]. Introduction [Ahm84, Can08b, Far09a, Hen04a, McG08, Nay05, Ril01a, Set04, Smi07a, Smi08f, Ste04a, Top00, BZ04, Byn08, Dix08c, Dol02a, Efr07, Goo02d, Gor02, Han07b, Hen02a, Man06, MD99, Mor01c, Ric06, WS^{+}01b]. Introductory [Hal02a, Kni02, Sle01, Can02b]. Intuitionist [Rav01, vD05]. invalidate [Bul04]. Inventing [Cha04, Coo07c, Edd09, Gra07a, Kni06a, SD00]. Invention [Chi08, DHC03, Eag07, Mac09a, NG01b, Noa04, Ren09, Sir06, Tur09b, Dix08b, HP00, HP02, Jea07, Luc02, MT07, Nor05, RSD07]. Inventions [Coo04, Jon06a, Pet04]. inventor [HP02, Rat07]. inventores [Edg05, NG01d]. inventors [NG01d]. investigation [Zui05, Zui06]. Invisible [Bro02c, Hay03, Hes03, TM01, Ver05]. Ioan [Jon04]. Iqbal [Can08b]. IRCICA [Brö02b]. Ireland [Ade04, Lea05, Hol08, MW03, Nes06d, WHF08]. Irish [Lea04, Ade05, MDF02]. Irony [But06, Goo04b]. Irritability [Edd07a]. Irritating [Edd07a]. Irving [Del09]. Isaac [Edg05, Fri04, Gou06, Key05, Maz07a, Smi05b, Bra00, Far00a, Fei04, IH05, KG04, Lin03, Liv03, NG01d, Pat07, Rei05a]. Isambard [Har03a, Buc02, Mil08]. ISBN [Abr08, Ack08, Ade04, Ade06, Ade08, Aga00, Aga01a, Aga01b, Aga01c, Aga02, Aga03a, Aga03b, Aga04, Aga05, Aga06, Aga08a, Alb01, Alb03a, Alb03b, Alb08, And03a, And09, App05a, App05b, Art04, Ash01, Ash04b, Aug09, Bar01a, Bar01b, Bar07, Bar08, Bax09, Bee01, Bek06, Ben00, Ben02, Ben04b, Ben06, Ber06a, Ber08, Big01, Big06, Big07, Biv06, Biv08, Bla03, Boh01, Boo07, Bor03, Bow04, Bow06a, Bow07, Bra00, Bra07, Bre04, Bre05, Bre06, Bro04a, Bro05a, Bro05b, Bro06a, Brö02b, Brö04b, Bro08b, Bro09, Bro03b, Bro03c, Bro01b, Bro02e, Bro04c, Bru06, Bru09, Bud09, Bur07a, Bur08, But06, CK02, Cam08, Can09, Can08a, Can01b, Can02b, Can04a, Can07, Can08b, Can04c, Car01a, Car06, Cha05, Cha01, Cha03, Chi08, Chi02b, Chi04a, Chi06, Chi01a, Chi01b]. ISBN [Chi03b, Chi03c, Chi04b, Chr05, Chu02, Cla01, Cla04a, Cla05, Cla06, Cla07a, Cla09, Cla04b, Coe09, Coh09, Col00, Col01b, Col02, Coo05a, Coo07a, Coo03, Coo04, Coo06a, Coo06b, Coo06c, Coo02b, Cor00, Cor04b, Cor05, Cor06, Cro04, Cro05a, Cro05b, Cuo02, D’A03, Dan01, Das07, Das09, Daw06, Daw07a, De 07, Dea05, Dea02, Dea06a, Dea07, Dea08, Dea09, Del06a, Del08, Del09, DM05b, Dep07, Dew01, Dew07, DiM07, Dig08, Dig09, Div01, Dix00, Dix03a, Dix05, Dix06, Dix08a, Dix09, Dod02, Dod06, Dol01, Dol02a, Dol02b, Dol02c, Dol04a, Dol04b, Dol04c, Doo04, Doo05, Dou06, Dri03, Duf09, Dun08, ETV04, Eag05a, Eag05b, Eag06a, Eag06b, Eag07, Eag08, Edd03a, Edd03b, Edd04a, Edd05a, Edd05b, Edd06, Edd07a, Edd07b, Edd08, Edd09, Edg04, Edg05, Edg08, Eds04, EB07, Elm09]. ISBN [Elm02, Emm01, End04a, End04b, tF00, Fan02, Far07a, Far07b, Far07c, Far09a, Far06, Fie05, Fin01, Fin03a, Fin09, Fin02, Fis04, For03, For04, For05, Fox04, Fox07, Fra09, Fra05, Fre04c, Fre04d, Fri03a, Fri03b, Fri04, Ful01a, Ful01b, Ful01c, Ful02, Ful03, Ful09a, Ful09b, Fyf02b, Fyf02a, Fyf03a, Fyf03b, Fyf05, Fyf07, Gan07, Gar03b, Gat05, Gau08, Gau09, Ghi01, Gil08, Gil09, Gin04, Gis09, Gli06, Gol05, Gol09, Goo01, Goo02a, Goo02b, Goo03, Goo04a, Goo05, Goo06a, Goo08a, Goo09a, Goo06b, Goo02c, Goo02d, Goo06c, Goo08b, Goo09b, Gor09, Gor02, Gou03a, Gou06, Gou09a, Gou09b, Gou03b, Gou03c, Gra08a, GG02, GG03b, GG03a, GG04, GG06a, GG06b, GG09, Gra05, Gra06, Gra08b, Gre02, Gui04, Gui05, Haf04, Haf08, Hag04, Hal02a]. ISBN [Hal03, Hal04a, Hal04b, Hal06, Ham07a, Ham08b, Ham07b, Han07a, Han07b, Har01a, Har00a, Har00b, Har05a, Har03a, Har02c, Har01b, Har01c, Har02d, Har03c, Har04c, Har06, Har08, Hay02b, Hay03, Hei07b, Hei07a, Hei08, Hem02a, Hem02b, Hem03, Hen01, Hen02a, Hen03a, Hen04c, Hen04b, Hen04a, Hen05b, Hen06a, Hen06b, Hen07a, Hen03b, Hen04d, Hen04e, Hes01, Hes03, Hes04, Hig04a, Hig03a, Hig01, Hig02, Hig04b, Hig05, Hig06, Hig07, Hig08, Hob06, Hoc08, Hoc09, Hod01, Hod06a, Hod06b, Hop01, Hop04, Hor03, Hou07, Høy01, Høy05, Hsu06, Hug00, Hug01, Hug07a, Hug07b, Hug08, Hug09, Hul07, Hun02, Hun05a, Hun05b, Hun06, Hun07a, Hur07, Hut04, Hut07, Hut09, Iro04, Jac09, Jam01a, Jam01b, Jam03, Jam05, Jar08a, Jen01, Jen02, Joh03a, Joh07a, Joh07b, Joh09, Joh02b, Joh02c]. ISBN [Joh02d, Joh03b, Joh04b, Joh04c, Joh05b, Joh05c, Joh06a, Joh06c, Jon02a, Jon04, Jon06a, Jon09, Jor01, Jor02, Jor06, Jor08, Kai05a, Kas02, Kem09, Kha06a, Kha06b, Kim05, Kim07, Kip04, Kir05, Kle04, Kle05, Kle09, Kni02, Kni03, Kni04, Kni05, Kni06a, Kni08, Kno08, Kno01, Kra02, Kra07, Kra08a, Kre06, Kus00, Kus03, Kus07, Lam02, Lau06, Lea04, Lea05, Led08, Leh08, Leh09, Lev07, Lin05, Lin07, Lip08, Liv05a, Loc02, Log08, Luc05, Lüt04a, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac08a, Mac09b, Mac09a, Mac01a, Mac05, Mac06b, Mac07, Mac08c, Mag06a, Mag06b, Man04, Man00a, Man01a, Man01b, Man06, Mar05, MW09a, Maz04a, Maz02b, Maz02a, Maz07a, Maz09, McG08, McO09, Mcc05, Med00, Mel01, Mer08, Mil02, Mil04, Mir07, Mok06, Mor03a]. ISBN [Mor05b, Mor06a, Mor06c, Mor06b, Mor07a, Mor07b, Mor07c, Mor01a, Mor02a, Mor03b, Mor06d, Mor06e, Mor01b, Mor01c, Mor02c, Mor08, Mor09b, Mos08a, Mos08b, MW07, MW09b, Mül01, Mun07, Mun08, Mun09, Mur02, Mus03, Nay09, Nea00b, Nea00a, Ner05, Nes01, Nes02, Nes06a, Nes06b, Nes06c, Nes07, Nes09, NG01b, Noa04, Noa09, Nut04, Nut05, Nye09, Olb02, Out01, Out04, Pad01, Pag01, Pag03, Pal00, Pal02, Pal06, Pan06, Pan08, Par01a, Par01b, Pau09, Pem04, Pem07, Pen04, Pet07a, Pet07b, Pet09, Pet08, Phi02, Pic07, Plo05, Poo09, Por01, Por03, Pre03, Pro08, Pug09, Pye09, Qui02, Qui04, Qui08, Rad00a, Rad02, Rad03b, Raj09, Ram02, Ras06, Rav00, Rav01, Ree08, Ree07, Rei07a, Rei07b, Ren00, Ren06, Ren07, Ren09, Ric04, Ric06, Ric09a, Ric09b]. ISBN [Ric01a, Ric01b, Ric08, Ril00, Ril01a, Ril01b, Ril02, Ris01, Rit02, Rit09, Rob03, Rob05, Rob07, Rob06, Rob08, Rob09b, Rob09c, Rol06, Roo08, Rud04b, Rud07, Rup08, Rut08, Sam07, Sch08a, Sch03a, Sch02a, Sch01, Sch02b, Sec08, Seg05, Sen02b, Sen06a, Set04, Sha02, Sha03b, Sha07, Sha09, She07, Shi03b, Sho03, Shr08, Shr09, Sir06, Sir09, Siv06, Siv08, Ske04, Sle01, Sle02, Sle03, Sle05, Sle09a, Sle09b, Smi00a, Smi01b, Smi02a, Smi02b, Smi03b, Smi03a, Smi04, Smi05a, Smi05b, Smi06d, Smi06e, Smi06f, Smi06g, Smi06c, Smi07a, Smi07b, Smi08b, Smi08a, Smi08d, Smi08f, Smi08e, Smi08c, Smi09b, Smi09a, Smo08, Spa02, Spa03, Spa05, Spa06, Spr06, Sta03, Sta04, Sta09a, Sta08b, Sta06b, Sta06c, Sta07b, Sta07c, Sta07d, Sta09b, Ste04a, Ste05a, Ste05b, Ste09a, Ste09b]. ISBN [Ste07, Ste04b, Ste06, Stu06, Stu09, Sum05, Sum06, Sut03, Suz05, Ter08, Tho03a, Tho02b, Tho03b, Tho03c, Til04, Tim02, Tim03b, Tim03c, Tim04, Tim05, Tim08, Top02, Top08, Tos09, Tot04, Tra07, Tre06, Tso05, Tso07, Tur02a, Tur06a, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Van02b, Ver04, Vic07, Vid05, Wag05, Wal03, Wal06b, Wal06c, WM08, War08, Web06, Web07, Web08, Web09, Wee07, Wee08, Wee09a, Wee09b, Wei02a, Wei02b, Wer09, Whi05, Whi06, Whi04, Wil06, Wil09a, Wil09d, Wil09c, Wil09b, Win08, Wor07, Yam00, Yam04a, Yam04b, Yeo04, Yos00, Zig04, dC08]. Isis [Mer08, Had06]. iskusstve [Sir09]. Islam [Can08b, EB07, Iqb07]. Islamic [Ack08, Brö02b, EB07, çG00b]. Isotope [Her06]. Israel [Out01, Mor09a]. ISSN [Fox05, Her05b, Pye06, Wer04]. issue [AAH00a, Ano00a, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, AAH00b, Her05b]. Issues [GG02, Maz04a, Smi00a, Mag01]. Istanbul [Brö02b]. Istituto [Man00b]. Italian [Nut05, Gov02, Sir01]. Italiano [Man00b]. Italy [Web09, Gov02, Sno06]. Ithaca [Mor06c, Mor06b, Sut03, Tho03c]. IV [Jor01, CK02, Can08a, Hen04e, Hig06]. Ivor [Har00a, Pet07a]. Iwan [Gor07, Hei07b, Har05a, Sta06b]. ix [Aga05, Alb03a, App05a, App05b, Bor03, Bro06a, Brö04b, Bro08b, Bro01b, Cla04b, Cor06, Dan01, Edd03b, Eds04, Fan02, Fie05, Fyf03b, Goo06b, Goo09b, GG04, Haf04, Hen04d, Hig04a, Hig07, Hoc08, Høy01, Hsu06, Hul07, Joh07a, Kha06a, Kle09, Maz09, Mor06d, Mus03, Noa04, Nut05, Pem04, Pet07b, Por03, Pro08, Ren07, Sle09a, Spa02, Tre06, Yam00].

J

[Aga04, Aga05, Alb01, Ash01, But06, Can01b, Can08b, Car01a, Chi04a, Coh09, Coo06b, Dew07, Dix08a, Eag08, Edd03a, End04b, Fyf03b, Gol09, Gra08a, GG09, Hal02a, Har03c, Har08, Hei07b, Hei08, Hen05b, Hug07a, Joh05c, Joh06a, Jor02, Kim05, Kni05, Kno08, Mac07, Maz02b, Maz07a, Mor06c, Mor07a, Mor06e, Mor09b, Olb02, Raj09, Rav01, Ren06, Ric04, Ric08, Ril02, Smi05b, Smi08a, Sum05, Tho03c, vD05]. J. [Bax09, Bro05c, Gui04, Pol02, Tyb04, Yam04a]. Jack [Hod06b, Col01b, Hug01, Mok06, Ril00, Rud07]. Jackson [Kle09, Mor02c, Pan08, Stu09]. Jacob [Maz07a, Tur09a]. Jacqueline [Hut07]. Jacyna [Biv08, Dan01]. Jagdish [Kai05a]. Jahnke [GG06b]. Jahr [Smi06c]. Jahre [Hör78]. Jahrhundert [Hop04, Nes07]. Jahrhunderts [Hop01, Men00]. James [Ash01, Bar01a, Bro05b, Can01b, Ful03, Gra08b, Hun05b, Hun06, Jon04, Jor02, Mil04, Mor09b, Nay09, Rad03b, Ric06, Aga01a, Aga01b, Ber08, Bor07, Cha01, Har02a, Hil05, Mah03, Par06, Rad03b, Ril01a, Ris01, Rob09b, Top02]. Jamie [Cla09, Gou03a, Gou06]. Jan [Bro04a, Edd04a, Mor07c, Nay09]. Jane [Gar03b, Bah01, Smi07a]. Janet [Rol06, Gli06, Hur07]. Jankovic [Cla01, Nay09]. Janssen [Hig04b]. Janssens [McG08]. January [Jam01b, Jam05]. Janus [Jam08]. Janze [Har04c]. Japan [tF00, Wat97]. Jardine [Hig06, Man01a, Por01]. Java [Spa06, KNRW03]. Jay [Bow04, Gar03b, Har08, Van04a, Ful03]. Jazz [Dig09, Cla08]. Jean [Bur07a, Coo05a, Coo05a, Cor05, Eag06b, Kem09, Maz02a, Tur02a]. Jean-Daniel [Coo05a]. Jean-Luc [Cor05]. Jean-Marc [Coo05a]. Jean-Paul [Maz02a]. Jeannette [Smi03a]. Jeans [Sta07f]. Jeater [Chi08]. Jed [Hun05a]. Jeff [Spa02]. Jeffrey [Tim03b, War08]. Jeffries [Jam01b, Jam05]. Jena [Her05b, Wei02b, Zim00b, Zim00b]. Jene [Pau09]. Jenkin [Bul04]. Jenkins [Hor03, Twe09]. Jennifer [Hig06, Pro08]. Jenny [Gan07]. Jeremy [Hen07a, GG09, Maz02b]. Jerry [Wil06, WR04]. Jersey [Har00a]. Jessica [Out04, Ste09b]. Jesuit [Dea07, Hel05]. Jevons [Mir07, Maa05]. Jewel [Wee09b, Har07a]. Jewish [Can07, Gra08b, Lip08, CS06, Fro05, Par06]. Jews [Sta07c, Can05]. Jill [Cla04b]. Jim [McO09]. Jo [Sta06c]. Johannes [Ben06, Ste04a, EoToBB01, Hue03, Smi02a]. Johansen [Leh08]. John [Aga04, Ash01, Bar01a, Bre06, CDK98, Daw00, Edd02, Edd04b, Hes04, Hor03, Hun06, Hut07, Jam01b, Joh02c, Joh04b, Maz07a, Mor03a, Mor07b, Mos08b, Nea00b, Rud04b, Rud07, Sch02a, Smi05a, Smi08e, Sta03, Yam00, Yam04b, AH02, Ber06d, Big07, Bro03b, Bur00, Bur07c, Chr05, Cla05, Clu06, Dol04b, Doo04, Eag07, Fyf05, Goo01, HD02, Hor03, Hug08, Joh05b, KB02, Kus07, Loc02, Mac00b, MS05a, Mor05b, Mor05c, Ree08, Rob03, Rob05, RH04, Smi06d, Sum01, Tho02b, Tur06a, WS01a]. Johns [Ade08, Aga04, Bur08, Chi02b, Dea08, Dig09, Dix09, ETV04, Gar03b, Gra08a, Gra08b, Ham07a, Hem02a, Kni02, Kni04, Lyn09, Pet09, Pro08, Rad00a, Ril01a, Sle01, Spa03, Tim03b, Tre06, Web07, Web09, Whi06]. Johnson [Aga04, Leh08]. Johnston [Aga08a, Kre06]. Joined [Joh05c, Dic03].

Jon [CK02]. Jonathan [Gli06, Hod01, Daw07a, Kim07, Mun08, Out01, Rob07]. Jones [Ade04, Jac09, Ren06, Rit02, Sha07, Tyb04]. Jordan [Kir05, Tim03a, Tim03d]. Jörg [Olb02]. José [Goo02c, Med00]. Josep [Pye09]. Joseph [Dig08, Gra08b, Hes04, Jam01a, Jam01b, Jam05, Mac00b, McO09, Mil02, Mor02c, Par06, Sta03, Van02a, Wag05, Ful09b, Gas98a, HRR08, Jac00a, Moy97, Pac03, Ric04]. Josephine [Mus03]. Journal [Wer04]. journals [Joh00]. Journey [Gor02, TS01, Tho03b, Row00]. Joy [Hod01]. Judaism [Can08b, Kus03, Efr07]. judgement [Dry09, Sta01b]. Judith [Log08, Tho03b, DiM07, Ter08]. Judy [Hay03, Ric08]. Jules [McG08]. Julian [Kai05a, MM00]. Julie [Led08, Mor06d, Ter08, Wee07]. June [Dol01]. Jung [Smi06f, Sha03a]. Jürgen [Fie05, Høy05, Kus03]. Jury [War08, Cre06]. Just [Can02b]. Justin [MW07]. Jutta [Jar08a].

Kaiser [Hun07a, Kra08a]. Kalb [DiM07]. Kamm [Joh04b]. Kampfplatz [Coe09]. Kant [Cha03, Fri01]. Kantian [Can08a, FN06]. Kapil [Das09]. Karafyllis [Dea09]. Karch [Spa06]. Karen [GG09, Hen05b, Kem09, Gra08b]. Karl [Ful01a, Mag06b, Sta04, Hac00, Por04, Ras06]. Kassell [Tra07]. Kassler [Gou03a, Gou06]. Katharine [Fin01, Edd07b, Wee08]. Katherine [Hig06]. Kathleen [Jam01b, Jam05]. Kathryn [And03a]. Keller [Pal02]. Ken [Alb08, Cha05, Dol02c, Goo04a, Hun02]. Kenelm [Hed08]. Keniston [Cla04b]. Kennedy [Wil06]. Kenneth [Biv06, Cla04b, Coo06a]. Kent [Daw06]. Kenyon [Rit02]. Kenyon-Jones [Rit02]. Kepler [Kus00, Smi02a, Spr06, Hal04c, Met98]. Ker [Cla04b]. Keuren [Haf08]. Kevin [Ash01, Ade06]. Kew [Spa06, KNRW03]. Key [Smi00a]. Keynes [Dea05, Maz07a, Ric01b, Ske04]. Keyworth [Mor07b]. Kim [Big06, Web07]. Kimball [Kai05a]. Kind [Hay02b, Mac00d]. Kindred [KJ01, Rit02]. King [Eag05a, Sta06b, Mor05d, Maz00, Gor07]. Kingdom [Buc02, Har03a, Mil08]. Kingsland [Web07]. Kiple [Biv06]. Kipnis [Yam00]. Kircher [Gor02, Row00]. Kirk [Elm02, Tim03b]. Kirsch [Goo02a]. Kiyoshi [Eag06a]. Klaas [Rav00]. Klaus [Sta03, Gor09, Joh03b]. Klaushentschel [Joh02b]. Klein [Car06, Edd08, Hen03b]. Klep [Mag06a]. Kluwer [Chi06, D’A03, DM05b, Ghi01, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Joh02d, Joh06a, Kim05, Kle04, Loc02, Lüt04a, Man04, Maz07a, Plo05, Spa05, Tso05]. Knell [Alb01]. Knight [Bar07, Bow06a, Sec08]. Knights [Can01b]. Knoeff [Iro04]. Knopf [Coo06a, Goo09b]. knot [Sta01b]. Know [Whi04, Lev02]. Knowing [Ben03, Chr05, Ken08, Pic01]. Knowledge [Bek06, Ben06, Bre04, Chi04a, Das09, ETV04, Edd04a, Edd09, Eds04, Elm09, Fyf07, Gis09, Goo06b, Gou09a, Hen04b, Hen07a, Joh04c, Jon09, Kas02, Kre03, Maz08, Mus09, Poo09, Qui08, Ric09a, Smi08c, Sta07a, Tre06, Wea00, Whi05, Ash04a, Bai04, Bir01, BDLM07, Bra08, Coo07c, Dau05, Dea01, Dub06, Fre01, Gol98, Hue03, KM06, Lon01, Loz00, May05a, Mor05a, Mor02b, Wig08, Woo06, Yeo01b, çIh04]. Knox [Win09a]. Koenraad [Hig05]. Koenraadvan [Hig07]. Koertge [Sha09]. Kohler [Joh07b]. Köln [Gol05]. Kompendium [BSD08, Bru09]. Koninklijke [Gin04, Iro04, Mac09a]. Konrad [Dew07, Bur05]. Konzepte [Kre03, Nes06a]. Korallen [Web08]. KozÁK [Mor07c]. Kragh [Coe09, Fre04c]. Kramer [Wer04]. Krause [Sta09a]. kreativitet [Lar01]. Krementsov [Tim03c]. Kretschmann [Nes06a]. Krige [Hug08, Tur06a]. Krohn [Hen03a]. Kuhn [Bar01a, Ful01c, Gat05]. Kuklick [Shr08]. Kultur [Hop04]. kultureller [Coe09]. Kumar [Mac00a]. Kusukawa [Smi08c]. Kyle [Vic07].

L [Big07, Can01b, Coo04, Dix05, Edd03b, Fox04, Fyf05, Hag04, Med00, Mil04, Mor09b, Ram02, Rav01, Sha07, Shi03b, Spa05, Ste04a, Goo02d, Jam05]. L. [Dan01, Liv05a, vD05]. Labinger [Ful03]. Laboratories [Coo02b, Sta07d, BG02, FG05]. Laboratory [Elm02, Hen03b, Hul07, Kni04, Cha02a, HK01, Kle01, Law05]. ladies [Maz04b]. Ladina [App05b]. Lady [MW09b, Som04, Som07]. Lagrange [Gui04, Bou02]. Laird [Mel01]. Lajos [Mac08c]. Lake [Dri05]. Lamarck [Bur07a, CGGT06]. Lambert [App05b]. Lampreys [Bla07]. Lancaster [Jon09]. Land [Web09, Loc07, Suf07]. land-based [Loc07]. Landed [Bow04]. Landmark [Coo04, GGC^{+}05, Mil04, Pet07a]. Lands [Siv06]. Landscape [Cla07a, Siv08, Arn06a, Tay04]. Landscapes [Ste05a, Stu09, Ale02, Mit07]. Lang [End04b, Far07c, Kim07, Wei02a]. Language [Chi08, Dan01, Gil08, Hen04a, Rad00a, Smi08d, Alt99, Jac00b, Jea07, Rad00b, Rad07]. languages [Has09]. Lanham [Lin05]. Lankester [Mor09b]. lantern [Ver05]. Lanuza [Pye09]. Lanuza-Navarro [Pye09]. Laplace [Fox07, Har00a, Hah05]. Larry [Bro04a, Maz07a]. Larsson [For05, Mac01a]. Last [Col01b, Hen02a, Pre98, BR00, GB02]. Late [Bud09, But06, Chu02, Coo05b, Elm02, Fan02, Høy01, Luc05, Mac09a, Smi08b, Wal06b, Bos02, Cuo00, Dod01, Edd04b, Fin05a, Goo04b, Gra07a, HK01, Mos07, Mus09, Rel02, Ric03, RSD07, Sha08b, Wri00]. Latin [Eds04, Hig06, Ste05b, Loz00].

Latinus [McG08]. Launius [Rol06]. Laura [Dix08a, Dod02]. Lauren [Tra07]. Laurent [Ber01, Kni03]. Laurier [Art04]. Lavoisier [Cro05b, Kni03, Ber01, Cro03, Fau03]. Law [Chi08, Jea07]. Lawrence [Biv08, Edd05b, Hul07, Sle01, Mac05, Smi09a]. Laws [Gol04, Goo06a, Ste09a, CLW07]. Leah [Phi02]. Lean [Can04c]. Learning [Bek06, Daw00, Dew01, Sha09, Sta07d, Wee07, FG05, Mil00, çIh04, SM05b]. Lebanon [Jor08]. Lecture [Liv02]. Lectures [Cha03, Fri01, Mor09b, Rud04b, Sha03b, J^{+}07, Sim04]. Lederer [Hig04a]. Lee [Sen06a]. Lefèvre [Edd08, Fie05]. Legacy [Can08a, Ren06, Wal03, FN06, Swe01]. Legenda [Nes02]. Legenda/Studies [Nes02]. Legs [Bur08, Sle07]. Leibniz [Har00b, Sha07, Jon06b]. Leibold [Hal02a]. Leicester [Dix03a, Ver04]. Leiden [Chu02, Cuo02, Dea09, Elm09, Gol05, Hig02, Hig08, Kus03, Nut05, Smi09a]. Leidensia [Chu02]. Leila [Col00]. Leipzig [Sta01b]. Lennart [Her05b, Tyb04]. LEO [Cro05a, Cla04b, Pan06, Fer03, Gra06]. Leominster [Hur07]. Leonhart [Hen05b]. Leshem [DM05b]. Lesley [Nea00b, Mor06d]. Lessons [Jor06, Das04]. Letté [Fox04]. Letters [Ben00, Hug00, Hun05b, Liv05a, Mil02, BC00, Bro02a, Fre96, Gla97, Har02a, Mar09, May05b]. Leuven [McG08]. Levere [Bro04a, Kni02, Ram02]. Leviathan [Joh09, Bur07b]. Levine [Whi04]. Leviton [Sle02]. Lewisburg [Ben06]. Leys [Dri03]. Lezioni [Man00b]. Lianne [Wil09a]. Liba [Wil09d, Win08]. Libavius [Roo08]. Libby [Ren07]. Liber [McG08]. Liberal [Pau09, PP07]. Liberalism [Mun09, Coe07]. Libraries [Fyf02b, TBbTc90]. Library [Can01b, Coo04, Fri04, Gor02, Jam03, LFT02, Lin05, Maz07a, Mil04, Mok06, Sch02a, HW01a]. libri [Hen05b]. Libros [Med00]. Lichtmann [Elm09]. lies [Tur03a]. Life [Brö04b, Bud09, Cla09, Coo04, Coo06a, Cop04, Fri03a, Fri03b, Goo02b, GG06a, Har00a, Har03a, Hes04, Hug00, Hun06, Hur07, Hut09, Joh04b, Kle09, Kni05, Log08, Mag06b, Mun09, Nea00a, Pal06, Rad03b, Rav01, Ren09, Sha07, Smi03a, Smi07a, Sta07d, Ste09b, Web06, Bow03, Buc02, Coe07, Ewi01, FG05, Fre96, Gue00, Har02b, HLS00, HD02, Jon06b, KB02, Mah03, Mit05, Por04, Ric02b, Ris07, Sel08, Sen06b, Sha08b, Sil03, Slo04, Str00, TR06, Wil97, vD05, Wil08]. Light [Can09, Goo02b, Joh02d, Kre06, Che00, Joh01, Lev00, Ott08]. Lightman [Bar08, Jor08, Mor06a, Ric09b]. Lightning [Coo06a, Sil03]. Lights [HP00]. Like [Aga01a, Sch05b, Sco98]. Liliane [NG01b]. Lilley [Ene09]. Lillian [Aga00, Hes04]. Limits [Col01b, Mor06c, Bah01, Pre98, Rud05a]. Lindberg [Dix05]. Lindee [Mar05]. Linden [Smi05b]. Lindqvist [Mac01a]. Linguistic [Rad00a, Alt99]. Linnaeus [Sle01, Far00c]. Linné [Kim05]. L’invention [HP00]. Lion [Fri04, Mor01c, Smi02a]. Lipperhey [Hig04b]. Lisa [Man01a, Bru06, Ner05, Qui04]. Lisbon [Mor07c, KMO05]. Lissa [Mac09a]. Listening [Pem04, Sta03, Bar00c, Tho02a]. Lists [Ric01b, Key00]. Literary [Spa02, Wee07, Lov01]. Literatur [Smi06c]. Literature [Edd03b, Ful09b, Lin05, Nes02, Sle03, Sle05, Tur08, Cow01, Daw07b, HG07, Whi01]. Literaturnoe [Sir09]. Literaturwissenschaft [App05b]. Little [Cro04, Goo04a]. Live [Mac08a, Nye06]. lived [Far00a]. Liverpool [Twe09]. Lives [Ben00, Edd05b, Stu09, Gla97, Mit07, New03]. Livingstone [Fyf02a, Lyn09, Sha03b, Smi08a]. lix [Hen03a]. Llobat [Pye09]. Lloyd [Cam08, Doo05, Hsu06, Nut04, Smi09b, Tot04]. Local [Bek06, Edd09, Nay09, Coo07c, Fin05a, FJC06, çIh04]. Locating [LS05]. location [Ren05]. Locke [Del09]. Logic [Cop04, Gal97, GG06a, Hen04a, Ren00, Tso07, Rei05c]. Logie [Bur00, Joh02c, Joh04b, KB02]. Logsdon [Aga04]. Lombroso [Smi08f]. Londa [Tos09, Del06a, Rob06]. London [Ack08, Aga00, Aga01a, Aga03a, Aga04, Aga05, Aga06, And09, App05a, Aug09, Bar01a, Bar08, Bax09, Ben06, Ber08, Boh01, Boo07, Bor03, Bow04, Bow06a, Bra07, Bre05, Bro05a, Bro09, Bro02e, Bud09, Can09, Can08a, Can02b, Can08b, Car01a, Cha05, Chi06, Chr05, Cla05, Cla06, Cla09, Coh09, Col02, Coo06c, Coo02b, Cro04, Cro05a, Daw06, De 07, Dea05, Dea02, Del08, Del09, DM05b, Dew01, Dew07, Dig08, Div01, Dix00, Dix05, Dix08a, Dod02, Dod06, Dol02a, Dol04c, Dri03, ETV04, Eag07, Edd05b, Edd07b, Edd08, Edg08, EB07, Emm01, Far07b, Far06, Fin02, For03, For04, Fox07, Fra09, Fre04d, Ful03, Ful09b, Fyf03a, Gar03b, Gat05, Ghi01, Gli06, Gol09, Goo02a, Goo02b, Goo03, Goo04a, Goo06a, Goo08a, Goo09a, Goo06b, Goo06c, Gou09a, Gou03c, GG02, GG03a]. London [GG06b, GG09, Gra05, Gra06, Gui04, Haf08, Hag04, Hal02a, Hal03, Hal06, Han07a, Har05a, Har03a, Har01c, Har06, Hay02b, Hay03, Hay06, Hei07b, Hei07a, Hem02a, Hem02b, Hem03, Hen03a, Hen04a, Hen06a, Hen06b, Hen07a, Hen03b, Hig04a, Hig03a, Hig01, Hor03, Hou07, Hug09, Hun02, Hun05a, Hun07a, Hut04, Hut09, Jam01a, Jam01b, Jar08a, Jen01, Jen02, Joh02c, Joh02d, Joh04c, Joh06a, Joh06c, Jon02a, Kim05, Kle04, Kle09, Kni02, Kni04, Kni05, Kni08, Kno08, Kra08a, Leh09, Lin05, Lin07, Lip08, Log08, Lüt04a, Mac08a, Mac06b, Mac07, Mac08c, Mag06b, Man04, Mar05, MW09a, Maz07a, Maz09, McO09, Mel01, Mer08, Mil02, Mok06, Mor05b, Mor06c, Mor06b, Mor07a, Mor07c, Mor02c, Mor09b, MW09b, Mun08, Mun09, Nay09, Ner05, Nes06b, Out04, Pad01, Pag01]. London [Pag03, Pal02, Pal06, Pau09, Pem04, Pem07, Pet07a, Pet08, Phi02, Plo05, Por03, Qui02, Qui04, Rad02, Rad03b, Ram02, Ree07, Rei07b, Ren06, Ren07, Ric06, Ric09b, Ric08, Ril02, Rob05, Rol06, Roo08, Rup08, Sch01, Sch02b, Seg05, Sen02b, Sha07, She07, Sir06, Siv08, Ske04, Sle01, Sle09b, Smi00a, Smi05a, Smi06e, Smi07a, Smi07b, Smi08a, Smi08d, Smi08f, Smo08, Spa03, Spa05, Spa06, Sta03, Sta06b, Sta09b, Ste09a, Ste09b, Ste07, Ste04b, Stu06, Stu09, Sum06, Sut03, Tho02b, Tho03c, Til04, Tim03b, Tim03c, Tim04, Top02, Tot04, Tra07, Tre06, Tso05, Tur02a, Tur06a, Tur09a, Tur06b, Twe09, Van04a, Van02b, Ver04, Wal06c, Web06, Web07, Web09, Wee09b, Wer04, Whi05, Whi06, Whi04, Wil06, Wor07, Yeo04, Yos00, dC08, Gay07, Har07a]. London [Mor98, Rat07, Tra01, Tur00, Wig08]. Long [ETV04, Smi08d, Rad07]. Look [Gou03c]. Looking [Alb08, And03b, Arn06b]. López [Med00]. L’opinion [BV00]. Lorenz [Dew07, Bur05]. Lorenzo [GG02]. Lorraine [Wee08, Boh01, Far06, Fin01, Jor06, Mun07]. Lost [Dan01, Jac00b, Mor07c, Web07, Fre04a]. Louis [Nor07]. Louise [Fis04, Rut08, Spa03]. Louvain [Hig05, vC02]. Love [Ric03, Wal06b]. Loveland [Spa02]. Lovie [Dix03a]. Low [Dun08, Edd06]. Lozovsky [Eds04]. Ltd [Goo06c]. Luc [Cor05]. Lucas [End04b]. Luciana [Siv08]. Luckhurst [Mus03, Noa04]. Lucy [Chi02c, GG03b]. Lumières [HP00, NG01b]. lunar [Har00c]. Lunbeck [Ste09a]. Lund [Bro03c, Hes01]. Lundgren [Par01a]. lungfish [Bla07]. Lusa [Fox05]. Lustig [Ren06]. Lutterworth [Gil08, GG03b]. lvi [Hal02a]. lviii [Mor09b]. lxvi [Ric06]. lxxii [Maz07a]. lxxxix [McG08]. Lyell [Mor06c, Rud05b]. Lynch [Ric08, Rob03]. Lyndy [Smi02b]. Lynn [Hoc08, Hoc09, Bre05]. Lyons [Cro05a, Fer03]. Lysenko [Ham07b, RH05].

M [Aga04, App05a, Bar07, Biv08, Bru06, Chi06, Cla07a, Edd05b, End04b, Far07a, For05, Gou03b, Haf08, Har02c, Hod01, Hod06a, Hun05b, Jac09, Jon02a, Jon09, Kem09, Kim05, Lin07, Loc02, Mag06a, Mag06b, Mar05, Maz04a, Mor03a, Mor09b, Pau09, Ril01b, Rob03, Sir09, Smi09a, Sta03, Ste09a, Tim05, Web09, Wil06, Yeo04]. M. [Fin03b, GG03b, Hig04b]. M.D [Deb84]. MA [Ack08, Aga03a, Aga05, Ber08, Bra07, Can08a, Car01a, Cla06, Col02, De 07, Dod06, Dol04c, Edd08, Edg08, Elm09, Fin02, For05, Fox07, Fre04d, Ful03, Gol05, Goo03, Goo06a, GG04, Haf04, Haf08, Hal03, Hal06, Hay02b, Hem03, Hen07a, Hug08, Hun05a, Hun07a, Jam05, Joh07a, Joh04c, Jon02a, Kni03, Log08, Mac08a, Mac07, Mac08c, Maz07a, Mer08, Mor02c, MW09b, Ner05, Pag03, Pal02, Pan08, Par01a, Pem04, Pet08, Rad02, Rad03b, Ras06, Ric08, Rob06, Roo08, Sen02b, Sir06, Smi09a, Smo08, Sum06, Tho02b, Tim08, Van02b, WM08, Wee09a, Wil06, Med00]. Maas [Mir07]. Macdonald [Ben04b, Ewi01, Mun08, Smi03a]. Machamer [Ful02]. Machine [Abr08, CK02, Haf08, Joh05b, Sum06, Sut03, Bod06, Des01, NG01d, RS02, RV04]. machinery [CC08, Mar04, Var00]. Machines [Jef01, Joh04c, Jor02, Tim03b, Mor02b, Rat07]. Mackenzie [Col02]. Maclean [Kus07, Smi08c]. MacLeod [Ril01b, Chi06]. Macmillan [Can07, Das09, Edg04, Hay02b, Hig03a, Kha06a, Led08, Rei07a, Rit09, Sle09a, Smi08e]. Mad [Boo07, Rit09, Seg05, Sch03b]. Madness [Dig08, Por02, Smi08e, Suz05, Suz06, MF06]. Madrid [Edg05]. Maggie [Joh04c]. Magic [Dix00, Elm02, Tra07, Gou99, HK01, Ver05]. Magna [Hen02a, Man06, BR00, BRW04]. Magnani [GG02]. Magnetic [Dol02b, Bau01]. Magnetism [Rob09a]. Magnificence [For03, Ste00b]. Mahon [Hun06]. Maidenhead [Top08]. Main [Far07c, Joh02b]. Maison [Pye06]. Majesty [And09, Rei08]. maker [Hig04c]. Makers [Ben00, Hig01, Hig07, Pan08, Gla97, Jac06, Tur00, Tur03b]. Makes [Hei07a, Dea06b]. Making [Ack08, CK02, DS01, Dix06, Dol01, Dol02c, Dun08, Eag06b, Edd04a, Gol98, Gou09a, Hei07b, Hig03a, Hig01, Mai09, Maz07a, Maz09, Mir07, ML07, Out01, Pet08, Smi06f, Sta01a, Tos09, Tur09b, Ver04, Woo06, Dix08b, Far02, Fei04, Ful00b, GB02, Giv05, Hun07b, Igo07, Isr01, Ken08, Maa05, PS08, Ric00, Sha03a, Tur00, Whi03]. Malachi [Ful01a]. Malady [Stu09]. Malaria [Web09, Cue07, Sno06, Suf07]. Malcolm [Hen04c, Hut07, Joh04b, Mac08c, Hay02b, Hen04c]. Malthus [Kno01, Dol00]. Malthusianism [Dol00, Kno01]. Mammoth [Van04a, Coh02]. Man [Bla03, Cha05, Coo06a, Dix06, Gau09, Hen01, Hig04b, Hun06, Ric06, Smi08f, Tot04, AO03, Dje01, FH00, Har07b, LGR06, Mah03, Sas01, Sil03, Whi03, Ken07]. Management [Aga03a, Aga04, HH00, Joh02a, Sel08]. Manchester [Chr05, Cla01, Fis04, Gre02, Ham08b, Hoc08, Jac09, Jor02, Kha06b, Mur02, Mus03, Ric09a, Rut08]. Mandelbrote [Fri04, Pic07]. Människor [Lar01]. Manson [Har02c, Hay01]. Manual [BSD08]. Manuscript [Goo02c, CG00a]. Manuscripts [Can01b, Maz07a, Ste04a]. Mapping [Cor00, Edn97, Hen02b, Joh03b, May05b]. Maps [Dol02c]. máquina [Edg05, NG01d]. Mara [Pan06]. Marburg [Hop01]. Marc [Coo05a, Jam01b, Lip08, Jam01b, Jam05]. Marcello [Ful02]. Marcet [Bah01]. March [Deb84, Fri04]. Marcia [Sta03]. Marco [Joh07a, Kni03, Pan06]. Marconi [Hem03, Hon01]. Marcos [Web09]. Marcus [Dea07]. Margaret [Maz07a, Yam04b, Abr08, Jon02a, Pic07, Smi07b, Whi06]. Margareta [Sir09]. Margit [Har04c]. Maria [Man06, Pet09, Maz07b, Tot04]. Mariana [Qui08]. Marianne [MW09b]. Mariano [Hei08, Ade08]. Marie [Smi09a, Lüt04b]. Marika [Mac01a]. Marilyn [Dol02a]. Marina [Por01]. Marine [Haf08, RV04]. Mario [She07]. Maritime [Dun09, Eag06a, Hig07, Nea00a, vCA^{+}05]. Mark [Fyf02a, Biv06, Lea05, Mac08c, Ren09, Stu09]. Märker [Sta04]. Marketplace [Ric09b]. Marks [Tim05]. Maroske [End04b]. Marr [Eag08]. Marsch [Har01c]. Marsden [Rei07a]. Marshall [Coo05b, Coo06b]. Martin [Hop01, Mor06c, Sch08a, Web06, Wee07, Coh09, Fyf07, Men00]. Martínez [Ade08]. Martins [Siv08]. Marvin [Ben02]. Marx [Cla04b]. Mary [And03a, Bro06a, Smi08f, Bro06b, Lev07, Mac07, NS01, Sta06c]. Masao [tF00]. masculine [Ili00]. Mass [Pet08, Igo07]. Massachusetts [Cla04b, Mac08a]. Massimi [Sta07a]. Massimo [Gis09, Pet09]. Masters [Gra05, War03]. matches [Sch05b]. Material [Can02b, Dun09, Gal97, Hal02a, Ren00]. Materiality [Nes02]. Materializing [Ben04b]. Materials [Edd08]. Materie [Set04, WS^{+}01b]. Mathematica [Høy05]. Mathematical [Bra00, GG03a, GG06b, GG09, Gra05, Jon04, Sta06c, Ste05a, Ale02, NCW99, Ver03, War03, Wei02c]. Mathematician [Gra08b, Hut07, Pet09, MS05a, Maz07b, Par06]. Mathematicians [Jon02a, Jon04, Jam02, Mur00]. Mathematics [Aga03b, DM05b, Gra08a, GG06b, Høy01, Kus00, Maz02a, Pet07a, Plo05, Ric04, Sel00, Ste05b, Coh07, Cuo00, DS02, GGC^{+}05, Les03, Met98, Pie00, SS01]. Mathew [Hal02a]. Matiushin [Sir09]. Matter [Chi02a, Coo02a, Cou04, Fre04b, Kni02, Maz07a, Sen02a, Shi03a, Tho03c, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Bra08, JS04, Lev01, Pol02]. Matters [Coo07b, Kno08, Mac08a, Wor07, Dur05, Nye06]. Matthew [Bar07, Bro09, Cor00, Hen04e, Mos08a, Sha07]. Matthias [Kra07]. Maudslay [Joh05b]. Maunder [Ogi00]. Maurice [Gil08, Han07a, Gil08]. Maxime [Seg05]. Maxwell [Hun05b, Hal04c, Har02a, Hun06, Mah03, Spr06]. Mayerne [Hut09, TR06]. Mazzotti [Gis09, Pet09]. Mazzucatto [Qui08]. McCartney [Lea05]. McClean [Sle09a]. McCook [Hal04b]. McDonagh [Mus03]. McLean [Emm01, Mos08a]. McLellan [Ril01a]. Mcmullin [Han07a]. McTavish [Wil09a]. MD [Lin05]. Me [Jen02, NM00]. Meadows [Mok06]. Meaning [MW09a, Che06, Goo02e, Lev00, Mco01, Sch08b]. Meanings [Tur09b, Dix08b]. Measure [Goo04a, Ree08, Car07]. Measurement [But06, Kni06a, Kre06, Cha04, Goo04b, Joh01]. Measures [Edd06, CSML04]. Measuring [Col00, Zen98]. Mechanic [Nes06d, Noa09]. Mechanical [Bou02, Gui04, Smi08g, Smi08h]. Mechanics [Bou02, Dea08, Gui04, Mel01, Ber06c, Mar04, ML00, ML00]. mechanism [Mar04]. Mechanizing [Col02, Mac01b]. Medal [Can04b]. Medawar [Tur02a]. Media [Ner05, GP03]. Medica [Ade04, Jen01, Kno01, Med00]. Medical [Alb03a, Biv08, Bro02e, Fyf02b, Goo05, Hul07, Kip04, Led08, Qui02, Rup00b, Sch02a, BP01, Har00c, Wor00, Zim00b, Zui05, Zui06]. Medicana [Edd07a]. Medicina [Med00]. Medicine [Ade04, AO03, Aug09, Bro02e, Bro04c, Cha05, Chr05, Col01b, Coo02b, Duf09, Edd07a, Far09a, For03, Gou03c, Hal06, Har02c, Hul07, Jen01, Kas02, Kem09, Kip04, Kle09, Kno08, Kno01, Log08, Nut05, Pem07, Pic07, Qui02, Rut08, Sch02a, Sir01, Smi05a, Smi08e, Smi09a, Spa06, Ste00b, Stu06, Tim05, Tra07, dC08, Biv07, Bru08, Byn08, CP05, Coo07b, DS06, Dol00, Gay07, GRT06, HW01a, Har00c, Hay01, Hou06, Law05, Pen06, Pic01, Pre98, Roo07, Wea00, WPM05, LFT02]. Medieval [Kem09, Luc05, Sam07, Ste05b, Wer04, Fre05, GRT06, Mcg06, Rel02].

Medizin [Wei02b, Zig04]. Medizinische [Wei02b, Zim00b]. Meet [Dix05, LN03]. Meeting [Kra07]. Meetings [Mor05b]. Meets [Dol02b, She01]. Mehra [Kai05a]. Meinel [Par01b]. melancholy [Has09]. Melbourne [Pre03]. Melchior [Hei08, BI06]. Meli [Dea08]. Melling [Dig08]. Mellon [For03]. Mémoires [Coo05a]. Memoirs [Dol01, Sle02, TS01, Wil06, WR04, Tho03b]. Memories [Wil06, WR04]. Memory [Bow05, Hen04a, Ric01a, Smo08, Cav08, Far00a]. Men [Ade04, DiM07, Goo06a, Zin05, Chi03a, Gol04]. Mendel [Har03c, WO01]. Mendelsohn [Chi06, AM01]. Menopause [Log08, Hou06]. Mensing [Hig02]. Menso [Ste05b]. Mental [Gor09, Hay03, Hut07, Tho03a, Twe09, Wri01, Hen07b, MS05a, TM01]. Mentality [Gor09, Hen07b]. mention [Jar08b]. Mentor [Mor03a, WS01a]. Menz [Hop01]. meraviglie [Ber07, Fin09]. Merit [Ree08, Car07]. Metabiography [Far07c]. Métailié [Elm09]. Metaphor [Nes02, Sle03, Whi01]. Metaphysics [Bou02, Gui04]. Metchnikoff [Ghi01, GWT00]. Météores [Wer01]. meteorites [Nai06]. meteorological [Zui05, Zui06, Dry09]. Meteorology [Edd07b, And03b, Nay06]. Method [Tyb02, Bar00a, GO09, Ram01, Tho07]. methods [Joh00]. Methuen [Kus00]. Mexico [Web09, Cue07]. MI [Sha09]. Michael [Ash04b, Can01b, Can02b, Dea02, Dol02b, Jam03, Kra07, Man01a, Mer08, Mor09b, Pen04, Ren06, Ric08, Rit09, Sch08a, Twe09, Aga00, And09, Ash01, Ash04b, Can08a, Cha03, Cuo02, Dou06, Edg08, Elm02, FJ08, Ful01b, Ful00a, Goo03, Haf04, Hoc08, Jam08, Kip04, Kra02, Mor07b, Mor07c, Mor01c, Nes06b, Rus00, SM05a, Sle02, Sle03, Sta04, Tso05]. Michel [Fox04, NG01b, Kni04]. Michela [Sta07a, Tot04]. Michigan [Eds04]. Micoud [Hoc09]. microfilm [Can01b]. Microform [Can01b]. Micrographia [Hen05b]. Microphysics [Gal97, Hun05a, Ren00, BW01]. Microscope [Jar08a, Cav08, Hee08, Sch07]. microscopes [Hee08]. Mid [Fra09, Fre01, Hor07, Pap06]. mid-1970s [Hor07]. mid-nineteenth-century [Fre01, Pap06]. Mid-Victorian [Fra09]. Middle [Eag05a, Kin01, Nut05]. Midlands [Jac09, Jon08]. Midwiving [Bic03, Hob06]. Mifflin [Sta07b]. Miguel [Man00b]. Mike [Wil09c]. Mikhail [Sir09]. Mikulás [Har01b]. Milestones [Mor02a]. Militant [Bee01, Dri01]. Militare [Han07b]. miljöer [Lar01]. Millar [Yam04b]. Millard [Jen02]. Miller [Bro05b, Dew01, Fin03a, Hay03, Sta03, Sta07b]. Millikan [Jam01b, Jam05]. Mills [Hen07a]. Milo [Maz07a]. Milton [Kai05a]. Mind [Abr08, And03a, Bod06, Chi08, Dol02c, Fra09, Mag06a, Smi09b, CC07, Jea07, Llo07, NS01, Zag01]. Mindful [Mac09a, RSD07]. Minds [Col00, Yam04b, HM02, Zen98]. mineralogical [Edd02, Edd04b]. mineralogist [Pac03]. Miniati [Pan06]. Minutes [Bro04a, LTGS02]. Mirjam [Por03]. Mirowski [Por03]. Misa [Aga05]. Miscellaneous [Joh00]. Misperception [Sch06a]. Mission [Edg08, Ste06, Siv05]. Mitman [Mun07, Stu09]. Mitton [Smi07a, Yam04b]. Mnemosyne [Cuo02]. moa [Tho09]. moa-hunter [Tho09]. Mobile [Die06]. Model [Ste09a, CLW07]. Models [Alb03a, Car06, Gou03a, Hal04a, Lau06, Wei00, dCH04, Hop02, Kas01, KLG03]. Modern [App05a, App05b, Biv06, Bra07, Bud09, CK02, Coo03, Coo06b, Das09, De 07, Dea07, Dea09, Del06a, DiM07, Dod02, Edd09, tF00, Fie05, Fyf05, Gou06, Gou09a, Gou09b, GG09, Hay06, Hei07b, Hen03a, Hen07a, Hob06, Hoc08, Hsu06, Hut07, Kus07, Lin07, Log08, Luc05, Mac01a, Man04, Maz07a, Mir07, MW07, Pag03, Qui04, Ren09, Sha02, Shi03b, Smi06f, Smi08c, Smi09a, Sta06c, Stu09, Wee07, Wee08, Wei02a, Wil09a, Zil03, Bor07, BM05, Cla07b, Coo07c, Cro01a, Elm06, Fau03, Fei04, Gau01, GP07, Har04b, Has09, HBB^{+}03, Hel05, Hos01, Hou06, Ili00, Joh00, KG04, KM06, LRS03, LHL00, Llo04, Maa05, Mar09, McT05, NG01a, Ore01, PS05, Pri07, Rel02, SS05, Set07, Sha03a, Sha08b, Sil01, Smi06b, SM05b, Wat97, Wea00, Win09b, Zin05]. moderne [Cro05b, Fau03]. Modernist [Sle03, Val05, Whi01]. Modernity [Aga05, Bro08b, MBF03, Mur02, Out01, Pem04, Sta07c, Wee07, Can05, Car01b, Gau06, Isr01, Jam08, Tho02a]. Modes [Hen03b, Kle01]. Molecular [Yam00]. Moleski [Sch08a]. Moletti [Mel01, ML00]. Mollan [Lea04]. Moment [Maz07a, Fei04]. Money [Hul07, Law05]. Monforte [Fox05]. Monism [Wei02a, Gas98b]. Monismus [Zig04, Zic00]. monkey [Mae06]. Monks [Eag05a, Kin01]. Monograph [Kle09]. Monographien [Wei02b, Zig04]. Monographs [Jor02, Mor05b, Tho03a, Tra07]. Mont [Rud04b]. Monte [Leh08]. Montreal [Hen04e]. Monturiol [Edg05, NG01d]. Moon [Hen04d, Wil09d, RP02a, Tau08]. Moore [Aga01c, Ric06, Rob05]. Moral [Bud09, Far06, Tur09b, DV04, Dix08b, Sha08b]. Morality [Kno01, Dol00]. Morals [But06, Goo04b]. Moran [Hen07a, Roo08]. Mordechai [Maz07a]. Moreira [Mor07c]. Morgan [Rad00b]. Morland [Rat07]. Morphology [Her05b]. Morrell [Hug01, Ril00, Rud07]. Morrison [Dun08, Edd06]. Morrison-Low [Dun08, Edd06]. Morse [Coo06a, Sil03]. Mort [Joh04c]. Morus [Har05a, Hei07b, Sta06b, Gor07]. Moscow [Joh06a, Sir09]. Mosley [Hig06, Smi08b]. Most [Ber08, Log08, Del06b, Sen06b]. motion [Mcg06, Nay03]. motu [McG08].

Moulin [Maz04a]. Mountain

[Kai05a, MM00]. Mountains [Mor07b]. Movement [Wor07, Dur05]. movements [Mor06c]. Moyer [Jam01a]. Mr [Can04c]. MR2093668 [Ber06b]. Mudgway [Aga04]. Mueller [End04b]. Müller [Nes07]. multinationals [Hor07]. München [Smi02b, Smi06c, Sta04]. Munich [Hop04, Web08]. Münster [Mos08a, McL07]. Murchison [Mor07b]. Murdin [Nea00b]. Murray [Hor03, Jon02a, Sho03]. Musa [Pan06]. Musaei [Pan06]. Musée [Maz04a]. Museum [Alb03a, Ben02, Ben04b, Bow06a, Cha05, Cla05, Coo06c, Coo02b, Div01, Dol02a, Eag06a, Fin03a, For05, Hig02, Hig06, Hig07, Hig08, Nea00a, Pre03, Sta04, Win08, vCA^{+}05, AO03, Mae06, Pac03, Ras01, Ade05, Dun09, Mac02, Mco01, Ras01]. Museums [Alb03b, Alb08, Hen05a, Joh07a, Joh04b, Mac01a, Urr07, Ver04, Arn06b, Asm01, Ber05, DS01, LHL00]. Music [Dix00, Gou99, Gou03a, Gou06, Kas01, KG04]. Musicians [Pan08, Jac06]. musick [KG04]. Musselman [Mor08]. Muzaffar [Can08b]. Myles [Mor02c, Pan08]. Myrmecology [Bur08, Sle07]. Mystery [Hen04d, Hig06, Spa06, KNRW03, RP02a]. Mystic [Bro09, Rav01, vD05, Sta07e]. Mysticism [Tim08, Deb06]. Myth [Van04a, Coh02]. Mythos [Man01b].

N [Bru06, But06, Cuo02, Dew01, Dol04a, Fyf02a, Kir05, Lyn09, Sha03b, Smi08a, Ste09a, Tim02, Tos09]. N. [GS02]. nach [Sch00, Smi02b]. Nadja [Wor07]. naissance [Cro05b, Fau03]. Name [Bro05a, Gat02]. Nancy [De 07, Van02b, Dri03, Nut05]. Nanni [Dea09]. Naomi [Shi03b]. Napoleonic [Cor06, Gil04]. Napoli [Man00b, Pug09]. Narrating [BS02, Haf04]. Narrative [Whi04, Lev02]. Narratives [Dan01, Daw06, Ste09a, Ben04a, CLW07, Jac00b, Osl02]. NASA [Aga04, Mud01]. Natalia [Eds04]. Nathaniel [Dix09].

Natick [GG04]. Nation [Web09, Suf07]. National [Aga04, Bro03c, Eag06a, Edd03a, Hig07, Joh04b, Jor02, Nea00a, Sta03, Wal06c, Hen96, Wit01, Dun09, LFT02, Zim00b, vCA^{+}05]. Nationalism [Kha06a, Bas04a]. Nationalizing [Car01a, Nay06, Roc01]. Nationalsozialismus [Wei02b, Zim00b]. Native [Chi08, Bro03a, Jea07]. Naturae [Pug09]. Natural [Alb03b, Bar07, Ben06, Bos02, Bow04, Bow06a, Bra00, Bre04, Cla05, Dea07, Del09, Dix00, Edd04a, Edd09, End04a, Fin03a, Fin05a, Gar03b, Gau08, Joh07a, Kem09, Kle04, Leh08, Rad00a, Sch01, Sec00, Sle02, Smi07b, Smi09a, Spa02, Top02, Van02b, Wee09a, Ade05, Alb02, AA01, Alt99, Asm01, BC01, BVN07, Ber05, Bru08, Bul04, Coo07c, CGGT06, ED06, Gar03a, Gas09, GL00, GRT06, Gol98, Goo02e, Gou99, Gra07b, Ham01b, Hue03, Ili00, Joh04a, Jon02b, Ken07, KE05, Kum07, Lov01, Mae06, NCW99, NG01c, Roo07, Sch05a, Spa00, Yeo01b, GS99, Hel05]. natural-historical [Ham01b]. Naturalism [Smi06d, Van04b]. Naturalist [Sle01, Far00c, Mco01]. naturalistes [Hoc09]. naturalistica [Pug09]. Naturalists [AA01, Dod06, End04a, Joh07b, Fan03, tF04]. naturalium [McG08]. Nature [Bar01b, Bar08, Bro05a, Cla05, Cla07a, Dep07, Dra00, Dri03, Dun99, Far06, Fin01, Goo06a, Goo08b, Hal04a, Hal04b, Hei07a, Hig04a, Joh09, Jor02, Lam02, Man04, Mar05, McO09, Mer08, Pag03, Pro08, Ric09a, Sch02b, Siv05, Sle01, Ste07, Ste06, Tim03a, Tim03d, Tuc05, Van02b, Web08, Wil09d, Alb02, Bur07a, Bur07b, CGGT06, Cro01a, DP98, DV04, Dea06b, Fal05, Far00c, Gat02, Gol04, GHL03, GS99, Had06, Ken07, LFT02, Lig07, McC02, NG01a, Osl02, Rei05b, Rus01, Sim02, Ste01, Tau08, Tay04, Van02a, Wei00]. Nature/Culture [Mar05, GHL03]. Naturgeschichte [Web08].

Naturvetenskapen [Hes01].

Naturwissenschaften [Par01b, Sta04]. naturwissenschaftliche [Hop04]. Naval [Coo06b, Dic03, Joh05c, Bas04b]. Navarro [Goo02d, Pye09]. Navy [And09, Rei08]. Nazi [Tim02, Tur02a]. Nazism [Can07, Fro05]. NC [GG03a, Ren07, Ste09a, Wor07]. Nebular [Hes01, Hol99]. Nederlandse [Gin04, Iro04, Mac09a]. Needham [Wag05]. Needs [Hig08, Gra07a]. Neeley [And03a]. Neglect [Ste04b, Hoo02]. Negotiating [Ade08]. Nehemiah [Gar03a]. Neil [Div01, Mil02, Rit09]. nel [Ber07, Fin09]. nell’Italia [Cor04b, Fin09, Pug09, Ber07, Gov02]. Nemesis [Kim05, vLP01]. Neptune [Haf08, RV04]. Nervous [Mor08, Mus06, Mor00]. Néstor [Pye09]. Neswald [Coe09]. Netherlands [Gin04, Iro04, Mag06a, Mar09, Zui05, Zui06]. Network [Aga04, Joh04c, Mor02b, Mud01]. Networks [Aga03b, Ric04, Sta03, Dry09, RN09, Tho09]. Netz [Ber06b, Ber06a]. Neumann [Jon04, Jam02]. Neuroscience [Van05]. never [Wer01]. Newcastle [Pye09]. Newly [Bra00]. Newman [Edd05b, Pag03, Rud04b, Wee09a]. newspapers [Wig08]. Newspeak [Cla06, Ger02]. Newton [Bax09, Bra00, DM05b, Fri04, Gou06, Gui04, Gui05, Hig03a, Maz07a, Maz09, Smi05b, BC01, Bou02, Dav09a, Far00a, Far02, Fei04, FH04, Hig03b, IH05, JS04, KG04, Key05, Les03, Lin03, Liv03, Maz04b, Pat07, Ram01, Sha08a]. Newtonian [Maz07a, Maz09, Fei04, Ver03]. Newtonianism [Maz07a, FH04]. Newtons [Maz07a]. Nexus [Ras06, GWW04]. NH [Chi08, Jor08]. Nicholas [Rob09b, Dea06a, Joh06a]. Nick [Hig06, Lau06, Mai09, Por01, Alb03a]. Nicolaas [Har02d, Far07c, Qui02]. Nicolau [Goo02d]. Nicolaus [Hen05b, Hen06a].

Nicole [Dea09, Smi08f, Chi01b, Chi03c]. Niels [GG06b, Kra07]. Nieto [Edg05, Kle04]. Nieto-Galan [Edg05, Kle04]. Niko [Dew07, Bur05]. Nikolai [Tim03c]. Nils [Ham07b]. Nina [Ras06]. nine [Bow06a]. Nineteenth [Ade04, Bro05b, Bro03c, Can08a, Car06, Cla07a, Gau08, Ham07a, Har05a, Joh09, Maz07a, Maz09, Men00, Mor06a, Mor01b, Mus03, Pan08, Rad00a, Rei07a, Ric09b, Rit02, Wal06b, Whi05, Wil09b, Alb02, Bar06, Bur07b, Can01a, Che06, Fre01, FN06, Gra07a, HH01, Hof07, How00, Maz00, Mus09, Nay06, Pap06, Ric03, Sch06a, Sta01a, Alt99, Bar00b, Bro03a, Gra07b, Jac06, KLG03, LM02, MS05b, Mor98, Ols08, Tay04]. Nineteenth-Century [Bro05b, Bro03c, Can08a, Cla07a, Ham07a, Joh09, Maz07a, Maz09, Men00, Mor06a, Mor01b, Pan08, Rei07a, Ric09b, Wil09b, Alb02, Bar06, Bur07b, Can01a, FN06, Gra07a, Hof07, How00, Maz00, Mus09, Nay06, Sch06a, Sta01a, Bar00b, Bro03a, Jac06, MS05b, Ols08, Tay04]. NITA [Fri03b]. Nivola [Edg05]. NJ [Edd07a, Goo08b, Har05a, Hig04a, Lin07, Mag06b, Tur09a]. NMS [Edd06]. No [Aga04, Bro03b, Chu02, Cro05b, Edg05, tF00, Fox05, Goo05, Goo02c, Goo02d, Her05b, Hes01, Hig04b, Hig06, Hig08, Jam01b, Kha06b, Kle05, Mag06a, Man01b, Med00, Mor07b, Mor07c, Pye06, Ris01, Rud04b, Set04, Sha03b, Sir09, Sta07a, Whi05, Hug00]. Noah [Can08b, Mor06b]. Nobel [For05, GWW04, Hes04, LHL00, Mac01a, Ras06, WM08, Elz06, HD02, Lar01]. Nobody [Hen06a]. Nocturnals [Eag06a, NHA^{+}02]. Noel [Hut07]. Noga [Bro02e]. Noll [Fyf02a]. Noltie [Raj09]. Nomos [Kra02]. Non [D’A03, Plo05, SS00, Sel00]. Non-Western [D’A03, Plo05, SS00, Sel00]. Nonsense [Dol02b, She01]. Nordic [Haf04, BS02]. Nordmann [Can08a]. Noretta [Sha09].

Norman [Eag05b, Fri03a, Fri03b]. Norris [Aga02, DeV00]. North [Boo07, Eag07, Gou06, Jen02, Rol06, Stu06, Hol08, KG04, NM00, Osl02]. Northern [DiM07]. Norton [Aga00, Ste09a]. Notation [Eag05a, Kin01]. note [Ano00f]. Notes [Jen02, NM00, Ric01b, Goo02d, Hen02a, Key00, Kim05, Man06]. Notorious [Pag01, Tra01]. Notre [Dea07, Han07a, Hei08]. Novatores [Edg05]. Novel [Jon06a, Pet04]. November [Deb84]. Novoe [Sir09]. Novum [Man06, BRW04]. Nuclear [Tur09a, Ham08a, Hug04a, Tur03a]. nucleonics [Her06]. nueva [Fox05]. Number [Eag05a, Kin01]. Number-Notation [Eag05a, Kin01]. Numbers [Dix05]. Nummedal [Roo08]. Nuncius [Hen05b, Pan06]. nuova [Pug09]. Nuremberg [Gol05, Gin02]. NY [Hul07, Mor08]. Nyassa [Dri05]. Nye [Mac08a, Mac07, Sta06c]. Nyhart [Bre05, Hoc09].

O [ETV04, Far00c, Mac07, Sle01]. Oakland [Han07b, Hen05b]. ob [Sir09]. Obesity [Fri03a, Fri03b, Gue00]. Obituary [Bro03d, Can00, Deb84, Wat00]. Object [Jor06, Das04]. objectivity [Zag01]. Objects [Boh01, Dea08, Gou09a, Ber06c, Das00, Die06]. Obligatory [Ogi00]. Oblivion [Joh06a, Wad03]. Obozrenie [Sir09]. observateurs [Cha02b, Cor05]. Observation [Eag06b, Giv05]. Observations [Hen05b, Loc02, Ste00a]. Observatoire [Big01]. Observatory [Joh05c, RN09, Loc07, Dic03]. Observed [Hei08, SA06]. observer [Hof07]. observers [Cha02b]. Observing [Ast01, Fyf03b]. Obsession [Sta07b, Mil05]. Occasional [Mor07b]. Occult [Elm02, BVB02, HK01]. occulte [BVB02, Chi03b]. Occultism [Hay06]. occurrence [Che06]. Ocean [And09, Haf08, Joh05c, Dic03, Rei08]. Oceans [Tur09a, Ham08a]. Ochre [MW09b, Som07]. O’Connor [Rup08]. Octavo [Han07b, Hen05b]. October [Fri04, Hig02, Hig08]. octubre [Fox05]. Odd [Hay02b, Mac00d]. Odyssey [Goo04a]. Öffentliche [Nes06c, Hoc03]. Öffentlichkeit [Gos00, Hop04, Nes01]. Office [Aga04, Cro05a, Hug07b, Kra07, Fer03]. Ogée [Jor02]. OH [Daw06]. oil [Ewi01]. Oils [Smi03a]. Oklahoma [Eag05b, Fri03a, Fri03b]. Olaf [Her05b, Web08]. Old [Far07b, Hul07, Sch01, Law05, Spa00]. Oldenbourg [Hop04]. Oldenburg [Lüt04b, Hal02b]. Oldroyd [Mor07c]. OLEpetergrell [Kha06b]. Olga [Ric08]. Oliver [Nes06c]. Olschki [Pan06]. Olsen [Cha05]. Olson [For05, Wil09b]. Olsson [Her05b]. One [Eag05b, Ful03, BSD08, LC01]. online [Mai09]. Only [Hes04, HD02]. Ontario [GG02]. Ontology [Edd08]. Open [Ful01b, Kha06b, Top08]. Opening [Dix09, Com07]. Openness [ETV04, Lon01]. operations [Tho07]. Operazioni [Han07b]. Opham [Hod01]. Opinion [Chi01a, Hoc08, BV00]. Opinionated [Ful03, Bro01a]. Oppenheimer [Hug07b, Hug09, MW09a, Tho03c, CH05, Pol02, Sch08b, Tho06]. opportunities [TDNS08]. Optical [Joh02d, Smi02a, Smi08g, Smi08h, Che00, KD00]. Optics [KD00, Mor02c, Smi02a, Jac00a]. optimism [Sch06a]. oral [Wei02d]. orbes [Goo02d]. orbium [Hen05b]. Order [Fin01, Gis09, Hal03, Joh09, Sle01, Ank01, Bur07b, DP98, Far00c, Maz08]. Ordinary [Hor03, Pey01]. Oregon [Wil09d]. Orel [Har03c]. Oreskes [Shi03b]. Oreste [Pug09]. Organic [Car06, KLG03, NG04]. Organisation [Fyf07, Dau05]. Organism [Sut03, Che06, Des01]. organization [Zui05, Zui06]. Organon [Man01a, BJS00]. Organum [Man06, BRW04]. Origin [Sta07a, Sta09b]. Originals [Hig02]. Origins [Col00, Del09, Hen03a, Hig01, Kni08, Lyn09, Smi00a, Smi06d, Smi07b, Van05, Zil03, Liv08, Nay03, Rad00b, Sch05a, Sta08a, Tur00, Van04b, Zen98]. Ort [Hör78]. Osiris [Bre05, Ric01a]. Osservatorio [Big01]. Oster [Hen04c]. Other [Dol02b, Elm02, Fin03a, Sta03, Bau01, Hol00, SFS^{+}00]. Ottaviani [Pug09]. Otter [Can09]. Otto [Mor03b]. Ottoman [Bek06, BZ04, çIh04]. OUPusa [Har00b]. Our [Eds04, Gat05, Goo09b, Smi06e, Stu09, Vic07, Jef01, Loz00, Mit07, Sta06a, Tim03b]. Outram [Coo07a]. outsourcing [Dry09]. overseas [Hor07]. Overview [Smi03b, Ric02a]. Ovid [Har01a, Gee00]. Owen [Hay06, Hen05b, Gol05, Hen06a, Jon02b]. Own [Bra07, BV00, Elm05]. owners [Alb02]. Oxford [Abr08, Aga01b, Aga02, Aga08a, Alb03b, Bee01, Ben04b, Big07, Bla03, Bro04a, Brö04b, Bro08b, Bru06, Cha01, Dan01, Das07, Dol02b, Dol04a, Dol04b, Doo04, Doo05, Duf09, Eag06a, Far09a, Ful02, Ful09a, Fyf02a, Fyf05, Fyf07, Gan07, Gil09, Gor09, Gou09b, Gou03b, Gui05, Har00b, Har03c, Har04c, Hen02a, Hen04d, Hig03a, Hig01, Hig07, Hod06b, Hsu06, Hug01, Hug07a, Hut07, Jam03, Joh07b, Joh03b, Jon06a, Kai05a, Kni06a, Kus07, Leh08, Liv05a, Lüt04b, Mac00a, Man00a, Man06, Maz02b, Maz07a, Nes02, Noa04, Out01, Pen04, Pet07a, Pet07b, Poo09, Pye06, Rav01, Rei07b, Ris01, Rob08, Sch08a, Sen06a, Shi03b, Shr08, Shr09, Sle03, Sle05, Smi01b, Smi03a, Smi07a, Smi08c, Smi09b, Spa02, Sta03, Sta09a, Sta07c, Sta07d, Sum05, Suz05, Ter08, Tho03a, Tho03b, Tim02, Tra07]. Oxford [Tur09b, Tyb04, Van05, Vic07, Wal06b, War08, Yam00, Bui07, BP05, FG05, HBB^{+}03, Hei05, Mor97a].

P [Aga03a, Bor03, DiM07, Joh06c, Mor09b, Ter08]. P. [Hun05b, Joh02b, Vic07]. Pacemaker [Tim03b, Jef01]. Pacific [Siv06, Ste06, Siv05]. Paderborn [Har01c]. paese [Ber07, Fin09]. page [Hen02a, Man06]. Pagel [Deb84]. Pain [Qui04, Sil01]. Palace [For04]. palaeoanthropology [GO09]. Palaeobotany [Mor07a, BBW05]. Paleoanthropology [Goo08b]. Palermo [Big01]. Palestine [Web09, Suf07]. Palgrave [Can07, Das09, Edg04, Hen04c, Hen04b, Kha06a, Led08, Rei07a, Rit09, Sle09a, Smi08e]. Palmer [Dol02a]. Pamela [ETV04, Gou09a, Hen06b]. Pampered [Spa03, Rob02]. Pancaldi [Goo05]. Panda [Dix09, Com07]. Pang [Big06]. Paola [Cor04b, Fin09, Goo05]. Paolo [Pan06, Hen04a]. Paper [Bro02e, Car99, Car06, Fin03a, Goo02c, KLG03, Sta06b]. paperback [Ade04, Aga01a, Aga01b, Aga06, Alb03a, And03a, App05a, App05b, Art04, Bar01b, Bee01, Ben04b, Big06, Biv06, Biv08, Boo07, Bor03, Bow04, Bra00, Bro05a, Bru06, Bru09, Bur07a, CK02, Can09, Can01b, Can02b, Can04c, Cha05, Cha03, Chi01b, Chi03b, Chi03c, Chi04b, Chr05, Cla01, Cla05, Cla04b, Coe09, Coo05a, Coo07a, Coo06c, Coo02b, Cor04b, Cor05, Cro05b, Dea06a, Dea08, Dew07, Div01, Dix03a, Dix06, Dol04a, Edd04a, Edd05b, Edg04, Edg05, Fan02, Far09a, Far06, Fie05, Fin09, Fis04, For05, Fox04, Fox05, Fra05, Fre04c, Fri03a, Fri03b, Fri04, Ful01b, Ful03, Ful09a, Ful09b, Gan07, Gar03b, Gat05, Gau08, Gil08, Goo02b, Goo03, Goo05, Goo08a, Goo08b, Goo09b, Gor02, Gou09a, Gou03c, GG03b, GG03a, Gra05, Gui05, Hag04, Hal02a, Hal04a, Hal04b, Ham07b, Han07a]. paperback [Har02d, Har04c, Har08, Hei07b, Hen04c, Hen04b, Hen07a, Hen04e, Her05b, Hes01, Hig04a, Hig02, Hig04b, Hig06, Hig08, Hod06b, Hop04, Hug07b, Hut04, Jen01, Joh02b, Joh05b, Jon04, Jon09, Jor02, Kas02, Kha06b, Kim07, Kle05, Kle09, Kni02, Kno01, Kra02, Kra07, Kra08a, Lam02, Lau06, Lip08, Mac09b, Man01a, Man01b, Mar05, Maz04a, Mcc05, Mor05b, Mor02a, Mor03b, Mor09b, Mül01, Mun08, Mur02, Mus03, Nes01, Nes02, Nes06b, Nes06c, Nes07, Nut04, Out04, Pan06, Phi02, Por03, Pug09, Pye06, Qui04, Rad02, Raj09, Ree07, Ren06, Ren07, Ric04, Ric06, Ric01a, Ril01a, Rob09b, Rob09c, Rol06, Rud04b, Sch03a, Sch01, Set04, Sha02, Sha03b, Sha07, Shr08, Sir09, Siv08, Sle01, Sle09b, Smi00a, Smi02b, Smi03b, Smi05a, Smi05b, Smi06e, Smi06f, Smi08f, Smi09b]. paperback [Spa02, Spa06, Sta03, Sta09b, Ste04a, Ste09a, Ste09b, Stu09, Ter08, Tho02b, Tho03c, Tim02, Tim05, Top08, Tos09, Tso07, Tur02a, Tur06a, Tur06b, Van05, Vid05, Wal03, Wal06c, Wei02b, Wer04, Whi05, Wil09c, Wil09b, Wor07, Yam04b, Yeo04, Yos00, Zig04, dC08]. Papers [Ghi01, GWT00, Hag04, Hen07a, Hug07b, Hun05b, Jam01b, Jam05, Kra07, CW02, Har02a, HRR08]. Pappus [Høy01, Cuo00]. Paradigms [Web07]. Paralipomena [Smi02a, KD00]. Parallax [Kni04]. Parallel [Ben06, Hue03]. Paris [Big01, Bur07a, Chi01a, Chi03b, Chi03c, Cor05, Cro05b, Fox04, Ful09a, Ham07a, NG01b, Rud04b, Spa03, Spr06, Vid05, Bar06, Fin03b, Gra07a, Rob02]. Park [Fin01, Wee08, Whi05]. Parke [Goo03]. Parrots [Spa03, Rob02]. Parshall [GG09, Gra08b]. Part [Eag05b, Fin03a, Goo02c, Man06, Smi02a, Wag05, Wal03, Zui05, Zui06, KD00, BRW04, Smi08g, Smi08h]. Particulars [Van02b, GS99]. Parting [Ful01b, Fri00]. Partnership [Ash04b, Hos03]. Pascal [Dea06a, Sha07, Ham03, Hig03b, Jon06b]. passages [Bra08]. Passion [Lea05, MW03]. Passions [Smi06g, Sta03, Dix03b, Hol00]. Past [Goo06c, Jor00, Swe04]. Patent [Jon06a, Pet04, Mil08]. Patents [Mac06b]. Patient [Dig08, Suz06]. Patricia [Gil09, Hig03a]. Patrick [Har02c, Hig06, Aga01c, Hay01]. patrie [Chi04b]. Patterns [Bur05, Dew07]. Paul [For03, Jam01b, Smi07a, Tot04, Can01b, Dix06, Mag06a, Maz02a, Pal00, Sle01, Zig04]. Pauli [Sta07a]. Paviland [MW09b, Som04, Som07]. Pearson [Goo08b, Mag06b, Por04]. Pedagogy [Hun07a, Kai05c]. Peder [Hal03]. Pedersen [Cla04a]. Pedro [Pye09]. Peel [Ren06]. Peeters [McG08]. Peiresc [Dew01, Mil00]. Pell [Hut07, MS05a]. Pelling [Pic07]. Pemberton [Rit09]. pendulums [Pal09]. Penelope [Dix00]. Penetrating [Hig04a, LFT02]. Penguin [Ric06, Smi06e]. Peninsula [Goo02c, CG00a]. Pennsylvania [Del06a, Har02c]. People [Joh04c, Pre03, Mor02b, Ras01]. Peoples [Siv06, Gas09]. Pepin [Hig04b]. Pera [Ful02]. Peral [Edg05, NG01d]. Perception [Hen04d, RP02a]. Pérez [NG01b]. Perfection [Whi06, Der03]. Perga [Cuo02, FU01]. Period [Rit02, KJ01, Sch05b]. periodical [Mus09]. Periphery [Pat07, Pye06, NG04]. Pérouse [Spa05, Wil03]. Perseus [Sen06a, Tho03b]. persona [Mil08]. Personal [Dix03a, BLR01, Cav08, Hof07]. Perspective [Dol02a, Edd06, Fyf02a, Led08, Qui02, Smi08f, ABB01, CSML04, LHN99, Rup00b, BW06]. Perspectives [Alb08, Cos00, Div01, Ful02, Haf08, Har04c, Hen04e, Hsu06, Hun07a, Jor02, Kir05, Mun07, Nea00a, Olb02, Pye09, Ric01a, Ris01, Tso05, BFR00, CJ02, DM05a, Fle98, HS02, Hug04a, Kai05c, Llo04, MPB00, RV04, Wil97]. Perversion [Chi04a, Bir01]. Pestre [Tur06a]. Peter [Dol02a, End04b, Far07c, Jor02, Kim07, Mac09a, Pau09, Sta08b, Wei02a, Bro03c, Dew01, Ful02, Gau09, Hei07b, Hei07a, Hem02b, Hen04b, Hig02, Jac09, Jen02, Kha06b, Man01b, Mor06e, Nes02, Olb02, Ren00, Smi08f, Ste07, Top08, Yam04b]. PETERelmer [Kha06b]. Peters [GG04].

Petry [Kim05]. Pettitt [Jon06a]. Peyton [Hor03]. Pharma [Qui08]. Pharma-Biotech [Qui08]. Pharmaceutical [Jon09, Mor06d]. Pharmacopoeias [Cow01, Edd03b]. Pharmacy [Rob07, Sim05]. Phenomena [Dol02b, Bau01]. phenomenology [Pal09, Sta07f]. Philadelphia [Aga01c, Ben00, Del06a, Dol01, Goo02c, Har02c, Kle05, Kre06, Lea05, Mor03b, Wal03, Dri05]. Philip [Aga05, Bro05b, Coo02b, Sum06, Jen01, Por03]. Phillips [GB02, Pau09, Rud07, Mor05c]. philosophe [Bur07a, CGGT06]. Philosopher [Sch08a, BC01, CGGT06, SM05a]. Philosophers [Ben06, Hue03]. Philosophica [Man01b]. Philosophical [BR96, Bar01a, Bro04a, Dol01, Ful02, Gat05, Goo02c, Gou06, Hsu06, Kle05, Man00a, Sta09a, Wal03, FK06, KG04, KCH00, LTGS02, Llo04, MPB00, Tay06, KG04]. Philosophies [Dea09]. philosophizing [Bos02]. Philosophy [Bar07, Bow07, Bra00, Cha01, Chi06, Cop04, Dea07, Dix08a, Fra05, Ful01a, Ful09a, Gau08, Ghi01, Goo06b, Gou03a, Gou06, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Joh02d, Joh06a, Kle04, Kni05, Kni06a, Leh08, Loc02, Mag01, Man01a, MW07, Out01, Pet07b, Sha02, Sle02, Smi09a, Ste09b, Ter08, Til04, Tim08, Tso05, Tso07, Bai04, Bru08, Deb06, ED06, Fet00, Gar03a, Gau01, GL00, Gra07b, Hac00, HS02, Hel05, Ili00, Isr01, Joh04a, Kas01, KG04, KE05, NCW99, Rad03a, Rei05c, Ric02b, Ris07, Roo07, Ryc05, Smi06b, Sny06, Sok08, Ver03, ZH06]. Phineas [Hay02b, Mac00d]. Phlegraei [Woo06]. phlogiston [Tay06]. phonograph [Rad00b]. Photographic [Sir06, DHC03]. photographs [Mus09]. Photography [Bro03c, Pro08, HH01, Tuc05]. Phrenology [Smi06d, Van04b]. Physical [Bru06, Rud04b, Sta06c, DGS03]. Physician [Hut09, Iro04, Pag01, Tra07, Kno02, Tra01, TR06]. physicist [GS02]. Physicists [Gor09, Lea05, MW03, Pan08, Hen07b, Jac06]. Physics [Aga01c, Ben00, Big07, Cor04a, Dea07, FG05, Fre04c, Gor07, Gra05, Gra06, Hen96, Hes04, Hul00, Jam03, Kra08a, Kre06, Lea05, Mac07, Nea00b, Pet07b, Sta03, Sta09a, Sta06b, Sta07d, Yos00, FK06, Hei05, Hel05, HD02, Kai05b, Mor05d, Nye04, Rus00, Ryc05, Set07, War03]. Physik [Cor04a]. physiology [Can01a]. Physique [Chi01b, Hul00, Rud04b]. Piatkus [Tur02a]. Picasso [Sta03, Mil01]. Pickering [Cla09, Dea02, Del09, Maz07a, Maz09]. Pickled [Alb03b, Asm01]. Pickstone [Chr05, Smi05a, Smi08e]. Picture [Yeo04, BSD08, tH02]. pictures [BSD08]. Picturing [Dri03, Ste01]. Pier [Maz02a]. Pierre [Fox07, Har00a, Hah05, Mer08]. Pierre-Simon [Har00a]. Pietro [Bur07a]. Pill [Bla03, Dje01]. Pinch [Pem07]. Piñero [Med00]. Pingree [Ner05]. Pioneer [Joh02c, Bur00, Hor00]. Pioneers [Joh05b]. Pitt [Hal04a]. Pittsburgh [Fra05, Goo06b, Hal04a, Sta07a]. Place [Gli06, Goo08b, Hay06, Smi03b, Ric02a]. places [Nay05]. Placing [Alb02, Gol09, Wit07]. Plague [Hay03, TM01]. Planetarium [Ben02]. planetary [Har00c]. Plant [Spa06, KNRW03]. Plantation [Kum07]. Plants [Rob06, Sch04a]. Plate [Ore01, Shi03b]. Plato [Leh08, Joh04a]. Playfair [AH02]. Plaza [Edg05]. Pleasure [Whi05]. plebeian [Rei05a]. plena [Fox05]. Pliny [Doo04, Hea99]. Plug [Hen04d]. Plurality [Can02b, WR01]. Plus [Can02b, Cop04]. Pnina [Ric01a]. Poetics [App05b, Kni04, Nye09, Rup08, Söd07, Cha02a, O’C07]. Poetry [Kni04, Cha02a, Win09b]. Poison [Ham08a, Ham08b, Tur09a]. Polanyi [Sch08a, SM05a]. Polarity [Wee09a, BVN07]. Polemical [Roo08, Spa02, Lov01]. Polenberg [Tho03c].

Policy [Ras06, GWW04]. Politècnica [Fox05]. Politic [Mor08, Mus06]. Political [Can09, Ott08]. Politics [Del06a, Dig08, Ful01a, Ham07b, Hei08, Lyn09, Mac05, Mac07, Ric01a, Ril00, TS01, Tim03a, Tim03c, Tim03d, Bla07, Bor07, Bru01, Hac00, Kre02, Liv08, MF06, Mor97b, Nai06, Nye04, RH05, SS05, SA06, Tur03a, Wig08, Tho03b]. politique [Chi04b]. Polity [Biv06, Cor06, Gil04]. Pomata [De 07]. Pontecorvo [Tur03a]. Pope [Maz00]. Pope-king [Maz00]. Popper [Ful01a, Hac00]. Popular [Cro01c, Fyf03b, Hoc08, Mur02, Rup08, Rut08, Top08, Tur06b, Ast01, Bow06b, Bro06c, Fyf04, O’C07, Sim04]. popularization [Kre03]. Popularizers [Bar08, Lig07]. Porter [Bro03d, Das07, Hod01, Hod06a, Mag06b, Pau09, Smi06e, Smi08e, Suz05]. Pòrtic [Goo02d]. Portrait [Bro03c, Hig04c]. Portraits [Jam03]. Portsmouth [Chi08]. Portugal [MCS09]. Portuguese [MCS09]. Post [Jon02a, BZ04, Mur00]. post-Copernican [BZ04]. Post-World [Jon02a, Mur00]. Postmodern [Ful01c]. postpaid [Aga04]. Postwar [Hug08, Kra08a, Kai05b, Kri06]. pounds [Sum01]. pounds-per-barrel [Sum01]. pour [Chi04b]. powder [Hed08]. Power [Ash04a, Coo06b, Fie05, Gli06, Goo03, Tre06, Bas04b, HHAH01, LRS03, Mco01]. Pozzo [Fin03a]. Pp [Abr08, Ack08, Ade04, Ade06, Ade08, Aga00, Aga01a, Aga01b, Aga01c, Aga02, Aga03a, Aga03b, Aga04, Aga05, Aga06, Aga08a, Alb01, Alb03a, Alb03b, Alb08, And03a, And09, App05a, App05b, Art04, Ash01, Ash04b, Bar01a, Bar01b, Bar07, Bar08, Bax09, Bee01, Bek06, Ben00, Ben02, Ben04b, Ben06, Ber06a, Ber08, Big01, Big06, Big07, Biv06, Biv08, Bla03, Boh01, Boo07, Bor03, Bow04, Bow06a, Bow07, Bra00, Bra07, Bre04, Bre05, Bre06, Bro04a, Bro05a, Bro05b, Bro06a, Brö02b, Brö04b, Bro08b, Bro09, Bro03b, Bro03c, Bro01b, Bro02e, Bro04c, Bru06, Bru09, Bud09, Bur07a, Bur08, But06, CK02, Cam08, Can09, Can08a, Can01b, Can02b, Can04a, Can07, Can08b, Can04c, Car01a, Car06, Cha05, Cha01, Cha03, Chi08, Chi02b, Chi04a, Chi06, Chi01a, Chi01b, Chi03b]. Pp [Chi03c, Chi04b, Chr05, Chu02, Cla01, Cla04a, Cla05, Cla07a, Cla09, Cla04b, Coe09, Coh09, Col00, Col01b, Col02, Coo05a, Coo07a, Coo03, Coo04, Coo06a, Coo06b, Coo06c, Coo02b, Cor00, Cor04b, Cor05, Cor06, Cro04, Cro05a, Cro05b, Cuo02, D’A03, Dan01, Das07, Das09, Daw06, Daw07a, De 07, Dea05, Dea06a, Dea07, Dea08, Dea09, Del06a, Del08, Del09, DM05b, Dep07, Dew01, Dew07, DiM07, Dig08, Div01, Dix00, Dix03a, Dix05, Dix06, Dix08a, Dix09, Dod02, Dod06, Dol01, Dol02a, Dol02b, Dol02c, Dol04a, Dol04b, Dol04c, Doo04, Doo05, Dou06, Dri03, Duf09, Dun08, ETV04, Eag05a, Eag05b, Eag06a, Eag06b, Eag07, Eag08, Edd03a, Edd03b, Edd04a, Edd05a, Edd05b, Edd06, Edd07a, Edd07b, Edd08, Edd09, Edg04, Edg05, Edg08, Eds04, EB07, Elm09, Elm02, Emm01, End04a]. Pp [End04b, tF00, Fan02, Far07a, Far07b, Far07c, Far09a, Far06, Fie05, Fin01, Fin03a, Fin09, Fin02, Fis04, For03, For04, For05, Fox04, Fox05, Fox07, Fra09, Fra05, Fre04c, Fre04d, Fri03a, Fri03b, Fri04, Ful01a, Ful01b, Ful01c, Ful02, Ful03, Ful09a, Ful09b, Fyf02b, Fyf02a, Fyf03a, Fyf03b, Fyf05, Fyf07, Gan07, Gar03b, Gat05, Gau08, Gau09, Ghi01, Gil08, Gil09, Gin04, Gis09, Gli06, Gol05, Gol09, Goo01, Goo02a, Goo02b, Goo03, Goo04a, Goo05, Goo06a, Goo08a, Goo09a, Goo06b, Goo02c, Goo02d, Goo06c, Goo08b, Goo09b, Gor09, Gor02, Gou03a, Gou06, Gou09a, Gou09b, Gou03b, Gou03c, Gra08a, GG02, GG03b, GG03a, GG04, GG06a, GG06b, GG09, Gra05, Gra06, Gra08b, Gre02, Gui04, Gui05, Haf04, Haf08, Hag04, Hal02a, Hal03, Hal04a, Hal04b]. Pp [Hal06, Ham07a, Ham08b, Ham07b, Han07a, Han07b, Har01a, Har00a, Har00b, Har05a, Har03a, Har02c, Har01b, Har01c, Har02d, Har03c, Har04c, Har06, Har08, Hay02b, Hay03, Hay06, Hei07b, Hei07a, Hei08, Hem02a, Hem02b, Hem03, Hen01, Hen02a, Hen03a, Hen04c, Hen04b, Hen04a, Hen06a, Hen06b, Hen07a, Hen03b, Hen04d, Hen04e, Her05b, Hes01, Hes03, Hes04, Hig04a, Hig03a, Hig01, Hig02, Hig04b, Hig05, Hig06, Hig07, Hig08, Hob06, Hoc08, Hoc09, Hod01, Hod06a, Hod06b, Hop01, Hop04, Hor03, Hou07, Høy01, Høy05, Hsu06, Hug00, Hug01, Hug07a, Hug07b, Hug08, Hug09, Hul07, Hun02, Hun05a, Hun05b, Hun06, Hun07a, Hur07, Hut04, Hut07, Hut09, Iro04, Jac09, Jam01a, Jam01b, Jam03, Jam05, Jar08a, Jen01, Jen02, Joh03a, Joh07a, Joh07b, Joh09, Joh02b, Joh02c, Joh02d]. Pp [Joh03b, Joh04b, Joh04c, Joh05b, Joh05c, Joh06a, Joh06c, Jon02a, Jon04, Jon06a, Jon09, Jor01, Jor02, Jor06, Jor08, Kai05a, Kas02, Kem09, Kha06a, Kha06b, Kim05, Kim07, Kip04, Kir05, Kle04, Kle05, Kle09, Kni02, Kni03, Kni04, Kni05, Kni06a, Kni08, Kno08, Kno01, Kra02, Kra07, Kra08a, Kre06, Kus00, Kus03, Kus07, Lam02, Lau06, Lea04, Lea05, Led08, Leh08, Leh09, Lev07, Lin05, Lin07, Lip08, Liv05a, Loc02, Log08, Luc05, Lüt04a, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac08a, Mac09b, Mac09a, Mac01a, Mac05, Mac06b, Mac07, Mac08c, Mag06a, Mag06b, Man04, Man00a, Man00b, Man01a, Man01b, Man06, Mar05, MW09a, Maz04a, Maz02b, Maz02a, Maz07a, Maz09, McG08, McO09, Mcc05, Med00, Mel01, Mer08, Mil02, Mil04, Mir07, Mok06, Mor03a]. Pp [Mor05b, Mor06c, Mor06b, Mor07a, Mor07b, Mor07c, Mor01a, Mor02a, Mor03b, Mor06d, Mor06e, Mor01b, Mor01c, Mor02c, Mor08, Mor09b, Mos08a, Mos08b, MW07, MW09b, Mül01, Mun07, Mun08, Mun09, Mur02, Mus03, Nay09, Nea00b, Nea00a, Ner05, Nes01, Nes02, Nes06a, Nes06b, Nes06c, Nes07, Nes09, NG01b, Noa04, Nut04, Nut05, Nye09, Olb02, Out01, Out04, Pad01, Pag01, Pag03, Pal00, Pal02, Pal06, Pan06, Pan08, Par01a, Par01b, Pau09, Pem04, Pem07, Pen04, Pet07a, Pet07b, Pet09, Pet08, Phi02, Pic07, Plo05, Poo09, Por01, Por03, Pre03, Pro08, Pug09, Pye06, Pye09, Qui02, Qui04, Qui08, Rad00a, Rad02, Rad03b, Raj09, Ram02, Ras06, Rav00, Rav01, Ree08, Ree07, Rei07a, Rei07b, Ren00, Ren06, Ren07, Ren09, Ric04, Ric06, Ric09a, Ric09b, Ric01a]. Pp [Ric01b, Ric08, Ril00, Ril01a, Ril01b, Ril02, Ris01, Rit02, Rit09, Rob03, Rob05, Rob07, Rob06, Rob08, Rob09b, Rob09c, Rol06, Roo08, Rud04b, Rud07, Rup08, Rut08, Sam07, Sch08a, Sch03a, Sch02a, Sch01, Sch02b, Sec08, Seg05, Sen02b, Sen06a, Set04, Sha02, Sha03b, Sha07, She07, Shi03b, Sho03, Shr08, Shr09, Sir06, Siv06, Siv08, Ske04, Sle01, Sle02, Sle03, Sle05, Sle09a, Sle09b, Smi00a, Smi01b, Smi02a, Smi02b, Smi03b, Smi03a, Smi04, Smi05a, Smi05b, Smi06d, Smi06e, Smi06f, Smi06g, Smi06c, Smi07a, Smi07b, Smi08b, Smi08a, Smi08d, Smi08f, Smi08e, Smi08c, Smi09b, Smi09a, Smo08, Spa02, Spa03, Spa05, Spa06, Spr06, Sta03, Sta04, Sta07a, Sta09a, Sta08b, Sta06b, Sta06c, Sta07b, Sta07c, Sta07d, Sta09b, Ste04a, Ste05a, Ste05b, Ste09b, Ste07, Ste04b, Ste06]. Pp [Stu06, Stu09, Sum05, Sum06, Sut03, Suz05, Ter08, Tho03a, Tho02b, Tho03b, Tho03c, Til04, Tim02, Tim03a, Tim03b, Tim03c, Tim03d, Tim04, Tim05, Tim08, Top02, Top08, Tos09, Tot04, Tra07, Tre06, Tso05, Tso07, Tur02a, Tur06a, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Van02b, Ver04, Vic07, Vid05, Wag05, Wal03, Wal06b, Wal06c, WM08, War08, Web06, Web07, Web08, Web09, Wee07, Wee08, Wee09a, Wee09b, Wei02a, Wei02b, Wer04, Wer09, Whi05, Whi06, Whi04, Wil06, Wil09a, Wil09d, Wil09c, Wil09b, Win08, Wor07, Yam00, Yam04a, Yam04b, Yeo04, Yos00, Zig04, dC08, Aug09, Cla06, Sir09]. Practical [Bro09, JS04, Maz07a, Sta07e, Hof07, Var00, Wig08]. Practice [Edd05a, Gil08, Hun07a, Joh02b, Kas02, Kim07, Kip04, Pic07, Ste05a, Ale02, Cro07, Goo04b, Har05b, Kai05c, Mil08, Sco04, Tor02, Wea00, WPM05, Wor00]. Practices [Bor03, But06, Cam08, Gou09a, Haf04, McO09, Ric01a, Smo08, Bow05, BS02, Llo06, Tur02b]. Prague [Mor07c]. Pre [Mag06a]. Pre-statistical [Mag06a]. Precision [Hig01, Lea05, Mor02c, Jac00a, MW03, Tur00]. Predecessors [Pre03, Ras01]. Predicting [Edd07b]. Predictions [Loc02, Ste00a]. Preface [Edg05, Fox04, Gar03b]. Prehistoric [Mor07c, Pad01, Cad00, Fre04a, Goo02e]. Prehistory [Har03c, GO09, WO01]. Prelude [Har03c, WO01]. Prematurity [Hoo02, Ste04b]. Prentice [Goo08b]. Preparing [GS02]. Presbyterians [Hol08]. Present [Biv06, Hay03, Har04b, Hör78, TM01]. Presidential [Bow06b, Bro05d, Jor00]. Press [Abr08, Ack08, Ade06, Ade08, Aga01a, Aga01b, Aga02, Aga03a, Aga04, Aga05, Aga06, Aga08a, Alb03b, And03a, And09, App05a, Art04, Ash01, Aug09, Bar01a, Bar01b, Bar08, Bax09, Ben02, Ben06, Ber06b, Ber06a, Ber08, Big06, Big07, Biv06, Biv08, Bla03, Boh01, Boo07, Bor03, Bow07, Bra00, Bra07, Bre05, Bro04a, Bro05a, Brö04b, Bro08b, Bro09, Bru06, Bud09, Bur08, But06, Can09, Can08a, Can02b, Can04a, Can08b, Car01a, Car06, Cha05, Cha01, Chi02b, Chi04a, Chr05, Cla01, Cla05, Cla06, Coh09, Col00, Col01b, Col02, Coo07a, Cor00, Cor06, Dan01, Das07, Daw00, Daw06, Daw07a, De 07, Dea06a, Dea07, Dea08, Del06a, Del08, Del09, Dep07, Dew01, Dew07, DiM07, Dig08, Dig09, Dix00, Dix05, Dix06, Dix08a, Dix09, Dod02, Dod06, Dol02b, Dol02c, Dol04a, Dol04b]. Press [Dol04c, Doo04, Doo05, Duf09, ETV04, Eag06a, Eag06b, Edd03a, Edd04a, Edd05b, Edd07b, Edd08, Edd09, Edg08, Eds04, Elm02, Fan02, Far09a, Far06, Fin02, Fis04, For03, Fox07, Fra05, Fre04c, Fre04d, Fri03a, Fri03b, Ful01a, Ful01b, Ful02, Ful03, Ful09a, Ful09b, Fyf02a, Fyf05, Fyf07, Gan07, Gar03b, Gat05, Gau08, Gau09, Gil08, Gil09, Gol09, Goo01, Goo02a, Goo02b, Goo03, Goo06a, Goo08a, Goo09a, Goo06b, Gor09, Gou09a, Gou09b, Gra08a, GG03b, GG03a, GG06a, Gra05, Gra08b, Gre02, Gui05, Haf08, Hag04, Hal02a, Hal03, Hal04a, Hal04b, Hal06, Ham07a, Ham08b, Han07a, Har01a, Har00a, Har00b, Har05a, Har02c, Har03c, Har06, Har08, Hay02b, Hay03, Hay06, Hei07b, Hei07a, Hei08, Hem02a, Hem03, Hen02a, Hen04a, Hen06b, Hen07a, Hen04d, Hes04, Hig04a, Hig01]. Press [Hig07, Hoc08, Hod01, Hod06a, Hod06b, Hou07, Høy01, Hsu06, Hug01, Hug07a, Hug08, Hug09, Hul07, Hun05a, Hun05b, Hun07a, Hut04, Hut07, Hut09, Jac09, Jam01a, Jam01b, Jam03, Jar08a, Jen02, Joh03a, Joh07b, Joh09, Joh03b, Joh04c, Joh05c, Joh06c, Jon02a, Jon04, Jon06a, Jor02, Jor06, Jor08, Kai05a, Kas02, Kha06b, Kip04, Kir05, Kni02, Kni04, Kni05, Kni06a, Kni08, Kno08, Kra08a, Kus07, Lam02, Lau06, Leh08, Leh09, Lev07, Lin05, Lin07, Lip08, Liv05a, Log08, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac08a, Mac09b, Mac09a, Mac05, Mac07, Mac08c, Mag06b, Man00a, Man01a, Man06, Mar05, MW09a, Maz02b, Maz07a, McO09, Mel01, Mer08, Mil02, Mir07, Mor03a, Mor06c, Mor06b, Mor07c, Mor02a, Mor03b, Mor01b, Mor01c, Mor02c, Mor08, MW07, MW09b, Mun07, Mun09]. Press [Mur02, Mus03, Nay09, Nea00a, Ner05, Nes06b, Noa04, Nut04, Olb02, Out01, Out04, Pag01, Pag03, Pal00, Pal02, Pan08, Pau09, Pem04, Pem07, Pen04, Pet07b, Pet09, Pet08, Phi02, Poo09, Por01, Por03, Pro08, Qui04, Qui08, Rad00a, Rad02, Rad03b, Ram02, Rav01, Ree08, Rei07b, Ren00, Ren06, Ren07, Ric09a, Ric09b, Ric01a, Ric01b, Ric08, Ril01a, Ril02, Ris01, Rob03, Rob05, Rob06, Rob08, Rob09c, Rol06, Roo08, Rup08, Rut08, Sch08a, Sch03a, Sch01, Sch02b, Seg05, Sen02b, Sha02, Sha07, She07, Shi03b, Sir06, Siv08, Sle01, Sle03, Sle05, Sle09b, Smi00a, Smi01b, Smi02a, Smi02b, Smi03b, Smi03a, Smi05b, Smi06f, Smi06g, Smi07a, Smi07b, Smi08b, Smi08a, Smi08d, Smi08f, Smi08c, Smi09b, Smo08, Spa03, Spa06, Sta03, Sta07a, Sta09a, Sta08b, Sta06b]. Press [Sta06c, Sta07c, Sta07d, Sta09b, Ste05a, Ste09a, Ste09b, Ste07, Ste04b, Ste06, Stu06, Stu09, Sum05, Sum06, Sut03, Suz05, Tho03a, Tho02b, Tho03b, Tho03c, Tim02, Tim03a, Tim03b, Tim03c, Tim03d, Tim04, Tim05, Top02, Top08, Tos09, Tot04, Tra07, Tre06, Tso07, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Van02b, Ver04, Vic07, Wag05, Wal06b, War08, Web06, Web07, Web09, Wee08, Wee09a, Wee09b, Whi06, Whi04, Wil06, Wil09d, Wil09c, Wil09b, Wor07, Yam00, Yam04a, Yam04b, Yeo04, Yos00, Mus09]. Press/German [Rei07b]. Press/OUPusa [Har00b]. Presse [Hop01]. Presses [Ben06, Bur07a, Chi01b, Chi03c, Chi04b]. Pressman [Col01b]. Prestel [Web08]. Preston [Mun08]. Previously [Can02b]. price [Aga04, Bro03b, Chu02, Cro05b, Edg05, tF00, Fox05, Goo05, Goo02c, Goo02d, Her05b, Hes01, Hig04b, Hig06, Hig08, Hug00, Jam01b, Kha06b, Kle05, Mag06a, Man01b, Med00, Mor07c, Pye06, Ris01, Rud04b, Set04, Sha03b, Sir09, Whi05]. Primary [Hen04c, Sta03, Hen96, Ost02a]. primo [Pug09]. primus [McG08]. Prince [Jar08b]. Princeton [Aga02, Bru06, Cor06, Dan01, Dol04a, Fre04c, Gan07, Har00a, Har05a, Joh07b, Joh09, Mag06b, Pal00, Ree08, Sta03, Tim02, Tyb04, War08]. Principe [Edd05b, Smi09a, MCS09]. Principia [Bra00, Gui05, NCW99, Dav09a]. Principle [Mil08, Sta07a]. Principles [Bra00, Cam08, Llo06, Lüt04a, Cle00, NCW99]. Print [Goo02c, CG00a]. Prints [Sch02a, HW01a]. Private [Bre06, Joh07a, Maz07a, Mun09, Ber05, Coe07, Ham01b, RH04]. Prize [For05, WM08, Lar01, Elz06]. Prizes [Hes04, HD02]. Problem [Daw06, MW07, Vic07, Ben04a, Smi06b, Sta06a]. Problems [Goo06b, GG04, Fra02, Yan02]. Proceedings [Brö02b, Hig02, çG00b]. Process [Goo01, Zim00a]. Proctor [Tim02, Tos09]. Producing [Win09a]. Production [Elm09, Fin02, Sum05, Ash03, BDLM07, Smi01a]. Productions [Hig04b]. Productive [For04]. Profession [Mok06, Hor00, Mea04]. Professional [Jon02a, Mur00, Ogi00]. Professor [Edd04b]. Profile [Far07b]. Profit [Jon09]. progeny [Mar04]. Program [Aga04]. Programs [Aga04, Joh02a]. Progress [Kni06a, Tim05, Cha04]. Project [Fri03a, Fri03b, Web09, Suf07]. projection [Sta01a]. projections [Cav08, Hee08]. projet [Chi04b]. Promise [Tim08, Deb06]. Promising [For08, Wil09c, Hor00]. promote [Spo02]. promoting [Ade05, NS00]. Proof [Col02, Mac01b]. Property [Jon06a, Pet04]. prophecy [Liv03, Sch00]. proposals [Var00]. Prose [Del08, Eag05b, Wal06a]. prospect [Ken08]. protesting [NS00]. Providence [GG06b, GG09]. provincial [Alb02, Nay06, WHF08]. proyecto [Fox05]. Pseudoscience [Dol02b, Bau01]. Psychiatrist [Dig08, Suz06]. Psychic [Dol02b, Bau01]. Psychoanalysis [Can07, Sle05, Fro05]. Psychological [Dix03a, Smi06g, Dix03b]. Psychology [BLR01, Dix03a, Joh06a, Smi00a, Smi01b, Smi03b, Smi06f, Wal03, Can01a, Ric02a, Ryl00, Sha03a, Swe01]. Psychosurgery [Col01b, Pre98]. Ptolemy [Tyb02, Tyb04, Bar00a, BJ00]. pubblico [Cor04b, Gov02]. Public [Bre04, Bre06, Hoc08, Joh07a, Kha06a, Kni06b, Maz07a, Pau09, Pet08, Sch03a, Sec08, Bah01, Bas04a, Ber05, Bro06b, Gos00, Gov02, Ham01b, Hil00, Igo07, RN09, RH04, Wig08, Yeo01b, BV00, Hoc03, PP07]. Publication [Mor07a, Mor07b, Top02, Ken07, Sec00]. Publications [Ash04b, Cha03, For05, Haf04, Haf08, Har01c, Hei08, Jam05, Joh07a, Jor01, Kni03, Mac01a, Nay09, Par01a, Pre03, Ras06, Roo08, Smi09a, Tim08, WM08, Mar09]. Publications/USA [Haf04, Hei08, Nay09, Roo08, Smi09a, Tim08, WM08]. publique [BV00, Chi01a]. published [Gui05]. Publisher [Mor07c]. Publishers [Can01b, Chi06, D’A03, DM05b, Fin03a, Ghi01, GG02, Gra06, Gui04, Hem02b, Hen03a, Hen03b, Joh02d, Joh06a, Kim05, Kle04, Loc02, Lüt04a, Man04, Mül01, Plo05, Spa05, Tso05, Bow06b]. Publishing [Aga01c, Ben00, Bro06a, Coo04, Edd06, Joh04b, Lea05, Mil04, Pye09, Ren09, Sen06a, Tur06b, Fyf04]. pump [Bru08]. Purbrick [Fis04]. Pure [Gra08a, Coh07]. Purity [DM05b, Les03]. Pursuit [Sta07a, Tim03a, Tim03d, Fin05a]. Pushing [Hen04e]. Put [Joh09, Bur07b]. Putting [Ric02a, Smi03b]. Pyke [Tur02a]. Pyramid [Art04]. Pyramus [Coo05a].

Qing [Dod02, Dod06, tF04, Hos01].

Quadrants [Eag06a, NHA^{+}02]. Quakers [Sta07c, Can05]. Quammen [Gar03b]. quantification [Zui05, Zui06]. quantity [Sum01]. Quantum [Fre04c, GS02, Sim04]. Queen [Fra09]. Quest [Cro04, Hen04a, Sta07b, Sum06, GO09, Mil05, RS02]. Question [Hoc09, Bro03a]. Questions [Mac08a, Maz04a, Nye06]. Quintessence [Log08, Sen06b]. Quotable [Sta03, EC00]. Quotations [Das07, BP05]. Quranic [Ahm84].

R [ABB01, Ash01, Bar01b, Cam08, Coe09, Dew07, Dol02a, Doo05, Dou06, Edd05b, Gar03b, Goo02c, Hal04a, Hal06, Hei08, Hem02a, Hen04e, Hoc08, Hod01, Hsu06, Jam01b, Jam05, Jor01, Mel01, Mor07a, Mor07c, Mor01b, Mor09b, Nay09, Nut04, Pag03, Rol06, Smi09b, Ste05a, Wee09a, Win08]. R. [ABB01, Dol02a, Eag08, Edd06, End04b, GS02, Har03a]. Rabi [Tho02b, Rig00]. Rabies [Rit09]. Race [Gan07, Goo09b, Lyn09, Rad00a, Rea05, Smi00a, War08, Alt99, Cre06, Liv08, Sch05b]. Racism [Smi00a]. Radder [Fra05]. Radical [Hut04, Isr01, Out01, Her01, Maz04b]. radicals [NG04]. Radick [Smi08d]. Radio [Hem02a, Slo00, TDNS08]. Radioactive [Tur09a, Ham08a]. Radioactivity [Hes03, Bro02c]. Radiumforschung [Ren05]. Rafael [Ade08]. Rafter [Smi08f]. Railway [Atm04, Fre01]. Railways [Car01b, Mur02]. rainbow [Wer01]. Rainer [Har02d]. Raising [Dav09b]. Raisonné [Hig05, vC02]. Raj [Das09, Mac00a]. Ralph [Rup08]. Ramberg [Mor06e]. Randal [Ske04]. Randall [Fan02]. Ransome [Leh08]. Ranyard [Mus09]. Raphael [Olb02]. Rashed [EB07]. Rasmussen [Pre03]. Rat [Mun08]. Rathke [Hop01, Men00, Men00]. ratio [Spa02]. Rational [Wal06b, Ric03]. rationalism [Sha00]. Rationality [Cha01, CC08, Fet00]. Raum [Set04, WS^{+}01b]. Raum-Zeit-Materie [Set04]. Raven [Hen03a]. Raymond [Chi02b]. Rays [Hes03, Bro02c]. re [Hee08]. re-enactment [Hee08]. Reaching [Hes01, Hol99]. reactions [MCS09]. Read [Hen06a]. Reader [Fyf03a, Gar03b, Gui05, Hen04c, Hes00, Ree07, Smi05b, Har04a, Lin03, Ost02a, Ost02b, WC02]. Reading [Cla01, Jan00, Liv03]. Readings [Bru06, DGS03]. Reaktion [Boo07, Dri03, Mun08, Stu09]. Reappraising [CH05, Hug07b]. Reardon [Gan07]. Reason [Cha03, Gol09, Lin07, Man04, Smi06e, Cro01a, Fri01, Rad00b, Wit07]. Reasoning [Cla09, Coo05b]. reassessed [Pat07]. Rebecca [Lin07]. Rebekah [Maz07a, Maz09]. Rebellion [Sta03, Hol00]. rebound [Smi08g, Smi08h]. Rebutting [Can02b]. received [Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09i, Ano09j, Ano09k, Ano09l]. Recentiores [Eds04]. Reception [Gin04, Top02, Sec00, Spo02, Ver02, WHF08]. reciprocal [Hun07b]. Reconsidering [Bed03, Rol06]. Reconstructing [Har08]. Reconstruction [Coh09, Hug08, Mor06c, Kri06, Rud05a, Rud08]. Records [Mor06d]. Recreating [Maz09]. Red [MW09b, Som04, Som07]. Redaction [Ste04a, EoToBB01]. Redruth [Can04c]. Redux [Ste09b, Ris07]. Reeds [Hen05b, Kem09]. reels [Can01b]. Rees [Hen02a, Man00a, Man06]. Reference [Smi05a, Tur06a]. Refiguring [Wee07, SM05b]. reflection [Smi08g, Smi08h]. Reflections [Bla03, Bow04, Dix03a, Hsu06, Hug07b, Jar08a, Smi09b, Wal03, BLR01, CH05, Dav09a, Dje01, Liv05b, Llo04, Llo07, Sch07, Swe01]. Reflexive [Smi00a]. Reform [Coh09, Mac05, Mor06c, Pic07, Tim05, Hul00, Rud05b, Rud08, WPM05]. reformas [Fox05]. Reformation [Mor09a, Mos08a, McL07]. réforme [Chi01b, Hul00]. Reforming [Dix08a, Sny06, Var00]. Regard [Rud04b]. Regardfully [End04b]. Regency [Twe09]. Regent [Whi05]. Regime [Sch01, Spa00]. Regina [Bro03b]. Regulation [Hem02a, RN09, Slo00]. Reich [Har04c].

Reid [Gou03c]. Reidy [And09]. Reign [Pet07b, Ryc05]. Reill [Ste07]. Reinhard [Aga03b]. Reinventing [Bro04c, MT07, Can02a]. Reisch [Tso07]. Relaño [Luc05]. Related [Edd03b, Cow01]. Relation [Fyf02b, Ric06, Win09a, TBbTc90]. Relations [Aga04]. Relativism [Bor03, Tur02b]. Relativity [Hut04, Pet07b, Sle03, Sta09b, War08, Cre06, Her01, Sta08a, Whi01, Ryc05]. Religion [Bro09, Can04a, Can08b, Far09a, Fyf02a, Kus07, Lin05, Lyn09, Man04, BM05, Dix08c, Hay02a, Liv03, Liv08, Sta07e, Cro01a]. Religious [Coe09, HW01b, Kus03, Mor09a, Sta07c, Ber06d, Can05, Kra08b]. Relocating [Das09]. Remaking [Coo06b, Bas04b]. Remarkable [Jam02, Jon04]. Remarks [Ste04a]. remembrance [Jor00]. Renaissance [Ack08, Bro01b, Daw00, ETV04, Eag08, GS99, Hig07, Mac09a, Mos08b, Van02b, Clu06, EM06, Fre99, Lon01, Nut05, RSD07, Tur03b]. René [Rud04b]. Renn [Fie05]. Rennes [Chi04b]. Report [Ano00e]. Reporters [Can04c]. Representation [Hen03b, Joh03b, Hen02b, Kle01, Pic09]. Representations [Joh02b]. Representing [Ara06, Goo08a, Jor02]. Reprint [Alb03a]. Reprinted [Can02b, Mor01c]. Reproduction [Wal06b, Mus09, Ric03]. Republic [Gin04, Liv05a, Mor07c, Bro02a, Mar09, May05b, Ver02]. Republics [Ree08, Car07]. reputation [HP02]. Rescuers [Spa05, Wil03]. Research [Aug09, Brö02b, Coo02b, Cro03, Her06, Joh02b, Joh03b, Kni08, Nes02, BG02, Gay07, Hen02b, Hor07, Rob09a, Tho07]. Researches [Mor01c, Far00b]. Resistance [Ste04b, Hoo02, Rei05a]. Resort [Col01b, Pre98]. Respectability [Tur08, Daw07b]. response [Hig03b]. Responses [Sta07c, Can05]. Restoration [Mor09a, Rat07]. Restoring [Emm01, McL00]. Rethinking [CE07, Gis09, Goo09a, Maz08]. Retooling [Dol04c, Wil02]. Retours [Ful09a, Rou07]. Retrying [Fin05b, Han07a]. return [Wil08, Rou07]. Rev. [Edd02]. Revealed [Alb01, Kne00]. Revealing [Ber06d]. Reverend [Rob09a]. Reviel [Ber06b, Ber06a]. Review [Abr08, Ack08, Ade04, Ade06, Ade08, Aga00, Aga01a, Aga01b, Aga01c, Aga02, Aga03a, Aga05, Aga06, Aga08a, Alb01, Alb03a, Alb03b, Alb08, And03a, And09, App05a, App05b, Art04, Ash01, Aug09, Bar01a, Bar01b, Bar07, Bar08, Bax09, Bee01, Bek06, Ben00, Ben02, Ben04b, Ben06, Ber06a, Ber08, Big01, Big06, Big07, Biv08, Boh01, Bor03, Bow04, Bow06a, Bow07, Bra00, Bre04, Bre05, Bre06, Bro04a, Bro05a, Bro05b, Bro06a, Brö02b, Brö04b, Bro08b, Bro09, Bro03b, Bro03c, Bro01b, Bro02e, Bro04c, Bru06, Bru09, Bud09, Bur07a, Bur08, But06, CK02, Cam08, Can09, Can08a, Can01b, Can02b, Can04a, Can07, Can04c, Car01a, Car06, Cha05, Cha01, Cha03, Chi08, Chi02b, Chi04a, Chi06, Chi01a, Chi01b, Chi03b, Chi03c, Chi04b, Chr05, Chu02, Cla01, Cla04a]. Review [Cla05, Cla06, Cla07a, Cla09, Cla04b, Coh09, Col00, Col01b, Col02, Coo05a, Coo03, Coo04, Coo06a, Coo06b, Coo06c, Cor00, Cor04b, Cor05, Cor06, Cro04, Cro05a, Cro05b, Cuo02, D’A03, Dan01, Das07, Das09, Daw00, Daw06, Daw07a, De 07, Dea05, Dea02, Dea06a, Dea07, Dea09, Del06a, Del08, DM05b, Dep07, Dew01, DiM07, Dig09, Div01, Dix00, Dix03a, Dix05, Dix06, Dix08a, Dod02, Dod06, Dol01, Dol02a, Dol02c, Dol04a, Dol04b, Dol04c, Doo04, Doo05, Dou06, Dri03, Duf09, Dun08, ETV04, Eag05a, Eag05b, Eag06a, Eag06b, Eag07, Eag08, Edd03a, Edd03b, Edd04a, Edd05a, Edd06, Edd07a, Edd07b, Edd08, Edd09, Edg04, Edg05, Edg08, Eds04, EB07, Elm09, Elm02, Emm01, End04a, End04b, tF00, Fan02, Far07a, Far07b, Far07c, Far06, Fie05, Fin01, Fin03a, Fin09, Fin02, Fis04]. Review [For03, For04, For05, Fox04, Fox05, Fox07, Fra09, Fra05, Fre04c, Fre04d, Fri03a, Fri04, Ful01a, Ful01b, Ful01c, Ful02, Fyf02a, Fyf03a, Fyf03b, Fyf05, Fyf07, Gan07, Gat05, Gau09, Ghi01, Gil09, Gin04, Gis09, Gli06, Gol05, Gol09, Goo01, Goo02a, Goo02b, Goo03, Goo04a, Goo05, Goo06a, Goo09a, Goo02c, Goo02d, Goo06c, Goo08b, Goo09b, Gor07, Gor09, Gor02, Gou03a, Gou06, Gou09a, Gou09b, Gou03b, Gou03c, Gra08a, GG02, GG03b, GG03a, GG04, GG06a, GG06b, GG09, Gra05, Gra06, Gra08b, Gre02, Gui04, Gui05, Haf04, Hag04, Hal02a, Hal03, Hal04a, Hal04b, Hal06, Ham07a, Ham08b, Ham07b, Har01a, Har00a, Har00b, Har05a, Har03a, Har02c, Har01b, Har01c, Har02d, Har03c, Har04c, Har06, Har08, Hay02b, Hay03, Hay06, Hei07b, Hei07a, Hem02a]. Review [Hem02b, Hem03, Hen01, Hen02a, Hen03a, Hen04b, Hen04a, Hen06a, Hen06b, Hen03b, Hen04d, Hen04e, Her05b, Hes01, Hes03, Hes04, Hig04a, Hig03a, Hig01, Hig02, Hig04b, Hig05, Hig08, Hob06, Hoc08, Hoc09, Hod01, Hop01, Hop04, Hor03, Hou07, Høy01, Høy05, Hsu06, Hug01, Hug07a, Hug07b, Hug08, Hug09, Hul07, Hun02, Hun05a, Hun06, Hun07a, Hur07, Hut04, Hut07, Hut09, Iro04, Jac09, Jam01a, Jam05, Jar08a, Jen02, Joh03a, Joh07a, Joh07b, Joh09, Joh02b, Joh02c, Joh02d, Joh03b, Joh04b, Joh04c, Joh05b, Joh05c, Joh06a, Joh06c, Jon02a, Jon04, Jon06a, Jon09, Jor01, Jor06, Jor08, Kai05a, Kas02, Kem09, Kha06a, Kim05, Kim07, Kip04, Kir05, Kle04, Kle05, Kle09, Kni02, Kni03, Kni04, Kni05, Kni06a, Kni08, Kno08, Kno01, Kra02, Kra07, Kra08a, Kre06, Kus00]. Review [Kus03, Kus07, Lam02, Lau06, Lea04, Lea05, Led08, Leh09, Lev07, Lin05, Lin07, Liv05a, Loc02, Luc05, Lüt04a, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac08a, Mac09b, Mac09a, Mac01a, Mac05, Mac06b, Mac07, Mac08c, Mag06a, Mag06b, Mai09, Man04, Man00a, Man00b, Man01a, Man01b, Man06, Mar05, MW09a, Maz04a, Maz02b, Maz02a, Maz09, McG08, McO09, Mcc05, Med00, Mel01, Mer08, Mil02, Mil04, Mir07, Mok06, Mor03a, Mor05b, Mor06a, Mor06b, Mor07a, Mor07b, Mor01a, Mor02a, Mor03b, Mor06d, Mor06e, Mor01b, Mor02c, Mor08, Mos08a, MW07, MW09b, Mül01, Mun07, Mun09, Mur02, Mus03, Nea00a, Ner05, Nes01, Nes02, Nes06a, Nes06b, Nes06c, Nes07, Nes09, NG01b, Noa04, Noa09, Nut04, Nut05, Nye09, Olb02, Out01, Out04, Pad01, Pag01, Pag03, Pal00]. Review [Pal02, Pal06, Pan06, Pan08, Par01a, Par01b, Pau09, Pem04, Pem07, Pen04, Pet07a, Pet07b, Pet09, Pet08, Phi02, Pic07, Plo05, Poo09, Por01, Por03, Pre03, Pro08, Pug09, Pye06, Pye09, Qui02, Qui04, Qui08, Rad00a, Rad02, Raj09, Ram02, Ras06, Rav00, Rav01, Ree08, Ree07, Rei07a, Rei07b, Ren00, Ren07, Ren09, Ric04, Ric06, Ric09a, Ric09b, Ric01a, Ric01b, Ric08, Ril00, Ril01a, Ril01b, Ril02, Ris01, Rit02, Rit09, Rob03, Rob05, Rob07, Rob08, Rob09b, Rob09c, Rol06, Rud07, Rup08, Rut08, Sam07, Sch84, Sch08a, Sch03a, Sch02a, Sch01, Sch02b, Sec08, Seg05, Sen02b, Sen06a, Set04, Sha02, Sha03b, Sha09, She07, Shi03b, Sho03, Shr08, Shr09, Sir06, Sir09, Siv06, Ske04, Sle01, Sle02, Sle03, Sle05, Sle09a, Sle09b, Smi01b]. Review [Smi02a, Smi02b, Smi03b, Smi03a, Smi04, Smi05a, Smi05b, Smi06d, Smi06e, Smi06f, Smi06g, Smi06c, Smi07b, Smi08b, Smi08a, Smi08d, Smi08e, Smi08c, Smo08, Spa02, Spa03, Spa05, Spa06, Spr06, Sta04, Sta07a, Sta09a, Sta08b, Sta06b, Sta06c, Sta07b, Sta07c, Sta07d, Sta09b, Ste04a, Ste05a, Ste05b, Ste09a, Ste09b, Ste07, Ste04b, Ste06, Stu06, Sum05, Sum06, Sut03, Suz05, Ter08, Tho03a, Tho02b, Tho03b, Tho03c, Til04, Tim02, Tim03a, Tim03b, Tim03c, Tim04, Tim05, Tim08, Top02, Top08, Tos09, Tot04, Tra07, Tre06, Tso05, Tso07, Tur02a, Tur06a, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Van02b, Ver04, Vic07, Vid05, Wag05, Wal03, Wal06b, Wal06c, WM08, War08, Web08, Wee07, Wee08, Wee09a, Wee09b, Wei02a, Wei02b, Wer04]. Review [Wer09, Whi05, Whi06, Whi04, Wil06, Wil09a, Wil09d, Wil09c, Wil09b, Win08, Yam00, Yam04a, Yeo04, Zig04, dC08, Ber06b]. reviewed [Ano01e, Ano04e, Ano09m]. Reviews [Aga03b, Aga04, Ash04b, Biv06, Bla03, Boo07, Bra07, Can08b, Coe09, Coo07a, Coo02b, Dea08, Del09, Dew07, Dig08, Dix09, Dol02b, Edd05b, Far09a, Fri03b, Ful03, Ful09a, Ful09b, Fyf02b, Gar03b, Gau08, Gil08, Goo08a, Goo06b, Haf08, Han07a, Han07b, Hei08, Hen04c, Hen05b, Hen07a, Hig06, Hig07, Hod06a, Hod06b, Hug00, Hun05b, Jam01b, Jam03, Jen01, Jor02, Kha06b, Leh08, Lip08, Log08, Maz07a, Mor06c, Mor07c, Mor01c, Mor09b, Mos08b, Mun08, Nay09, Nea00b, Rad03b, Ren06, Rob06, Roo08, Rud04b, Sha07, Siv08, Smi00a, Smi07a, Smi08f, Smi09b, Smi09a, Sta03, Stu09, Web06, Web07, Web09, Wor07, Yam04b, Yos00]. Revised [Mor02a, Mor03b]. Revisited [Brö04b, Pye06, Ene09, Har02b]. Revolucions [Goo02d]. Revolution [Dun08, Edd05b, Gas98a, Hal04a, Hen06b, Hen07a, Joh02d, Mac00b, Mor06c, Rav01, Rob07, Sch01, Sha07, Smi08a, Sta09b, Wee09b, Che00, Har07a, Jon06b, LW05, Mor05a, ML07, New03, Rud04a, Rud05a, Sim05, Spo02, Wei00, Spa00, Sta08a]. Revolutionary [Cor06, Gil04]. revolutionibus [Gol05, Hen05b, Gin02]. Revolutionising [Dea01]. Revolutionizing [Hen04b, Sen06a, Lee02]. révolutionnaire [Cor05]. Revolutions [CK02, Hen06a, Sta08b, ABC06]. Revue [Pye06]. Rewriting [Ter08, ZH06]. Reymond [Fin03b]. Rheinberger [Olb02]. Rhetoric [Lov01, Phi02, Spa02, Cec01]. rhetorical [Hal04c, Lig02]. rhétoriques [Hal04c, Spr06]. Rhodesia [Chi08, Jea07].

Rhys [Gor07, Har05a, Hei07b, Sta06b]. RI [GG06b, GG09]. ricerca [Pug09]. Richard [Coo04, Eag06a, Eag07, Hei08, Smi08f, Bre04, Coo04, Dea05, Dew07, Goo08b, Joh03a, Lin05, Mil04, Nor05, Ric01b, Sch02b, Sta09b, Tho03c, Wil09b]. Richards [Dix03a, Kni05, Ren06, Smi00a, Smi03b]. Richardson [Dol02c, Wal06b, Sum01]. Richmond [Mor06d]. Rick [Smi01b]. Riemannian [Val05]. Rienk [Gin04]. Riet [McG08]. Rigden [Tho02b]. rights [HP02]. Rina [Iro04]. Riordan [Aga00]. Rise [Gra05, Hay03, Hou07, Jam01a, Dar05, Moy97, TM01, War03]. Risk [Col02, Mac01b]. Riskin [Out04, Ste09b]. Ritual [Daw00]. Rivalry [Pad01, Cad00]. River [Fan02, Goo08b, Dod01]. Road [Bar01a, KCH00]. Rob [Maz07a]. Robbins [Spa03]. Robert [Aga01b, Dea02, Dou06, Elm02, Far07a, Ful03, Goo09a, Har00a, Hen03a, Hen05b, Mül01, Pol02, Raj09, Ren06, Tho03c, Bai03, Ben06, Coo02b, Hen01, Hou07, HD00, Hun07b, Joh07b, Jon02b, Kle05, Kni05, Man04, Mar04, Nol07, Pye06, Sta07d, Tim02, Tos09, Win09a]. Roberta [Duf09, Smi08e]. Roberts [Coo04, Mac09a]. robespierristes [Cor05]. Robin [Wee07, Hag04]. Robinson [Hur07]. Rochester [Hul07, Kir05]. Rocke [Car01a]. Rockefeller [Aga03b, Hul07, Law05, SS01]. Rocks [Mor06b, Her04]. Rodarmor [Van04a]. Roddey [Gou03c]. Rodger [Nay09, Ris01]. Rodopi [Ade04, App05b, Edd07a, Jen01, Kno01]. Rodríguez [Goo02d]. Roger [Bro03c, Gou06, Jen02, Art04, Bro01b, Fox07, Har03c, KG04, Mus03, Noa04, NM00, Ric09a, Rol06, Smi05a, Spa05]. Roland [Hem02b, Sum06]. Role [Bow07, Joh02b, Amu05]. Roll [Ham07b]. Roll-Hansen [Ham07b]. ROM [Han07b, Hen05b, Hig08]. Roman [Num07, Rob09c, Roo08, Cuo07]. Romancing [Hed08]. Romano [Dea09].

Romantic [Her04, Kni05, Mor06b, Rit02, Ric02b, KJ01]. Romantic-Period [Rit02]. Rombach [Coe09]. Rome [Cor04b, Gor02, Wil09d, Maz00, Row00, Tau08]. Ron [Bow07]. Ronald [Dix05, dC08]. Roos [Smi09a]. Roots [Bow07, Amu05]. rope [Atm04]. Roper [Hut09]. Rosalind [Dol04c]. Rose [Rud04b]. Rosemary [Goo06c]. Rosenblith [Wil06]. Roshdi [EB07]. Röslin [Man00b]. Ross [Hen04d, Web06]. Rossi [Hen04a]. Rothenberg [Jam01b, Jam05]. Rothmann [Man00b]. Rousseau [Vid05, BVB03]. Routes [Web07]. Routledge [Bow06a, Dig08, Gou03c, Ree07, Rob09b, Sec08, Smi00a, Smi03b, Smi05a, Til04, Tur06a, dC08]. Roux [Ful09a]. Rowland [Gor02]. Rowlinson [Yam00, Yam04a]. Roy [Chi06, Das07, GG03a, Hor03, LS05, Smi08e, Bro03d, Jon09, Ril01b, Smi06e, Suz05]. Royal [Ade06, Can01b, Cla04a, For04, Lea04, Lüt04b, Mor09a, Mor01a, Mor09b, Nea00b, Nea00a, Raj09, Spa06, Bri04, Can04b, Gas09, Hal02b, Hom09, Hun07b, J^{+}07, UU07, Wil97, Win09a]. Rózsa [Mac08c]. Rozwadowski [Haf08]. Rudolph [Smi04]. Rudwick [Coh09, Mor06c]. Ruiz [Pye09]. Rules [Ful03, Bro01a]. Rumm [Jam01b]. Rungiah [Raj09, Nol07]. Ruosch [Smi04]. Rupke [Far07c, Har02d, Qui02]. Ruse [Can02b, Ren06]. Ruskin [Bre06]. Russel [Fic04, Gar03b, Jon02b, Slo04, Web06]. Russell [Aga02, Jam03, Lam02, Mor01a, DeV00]. Russellburns [Joh02c]. Russia [Hal04a, Mor07b, Wei00]. Russian [Hal04a]. Rutgers [Goo02a, Hay03, Hig04a, Lin07, Tim04, Tur09a]. Ryckman [Pet07b]. Rylance [Smi01b].

S [And09, Bar01a, Ben00, Big01, Bro09, Coe09, Coh09, Dan01, Ham07a, Hen03a, Lin05, MW09a, Maz07a, Mor06c, Mor07c, Mor09b, Noa09, Pan06, Ril02, Smi00b, Sta07e, Tho02b, Win08, Yam04a, Gas09]. S. [Gui05, Mor03a]. Sabetai [Cuo02]. saccharometry [Sum01]. Sachiko [Smi08c]. Sacrifice [Lin07]. Saddle [Goo08b]. Safety [Dry09]. Sagamore [Haf08, Hei08, Joh07a, Nay09, Ras06, Roo08, Smi09a, Tim08, WM08]. Sage [Hen04e, Hug07a, Bro05c]. said [Col02]. Saliba [Ack08]. Salpêtrière [Sir06, DHC03]. Salt [Smi09a, Roo07]. Salusbury [Wil08]. Salvation [Tur06b, Fyf04]. salve [Hed08]. samhället [Hes01]. Samuel [Coo06a, Rat07, Sil03]. San [Pet07a, Sle02]. Sandra [Mor06c, Web09]. Sanofi [Chi01a]. Sanofi-Synthélabo [Chi01a]. Santa [Coo07a, Gou03b, Mor01c, Sho03]. Sante [Mor01c, Smi02a]. Santponc [Edg05, NG01d]. Sara [End04b]. Sarah [Del09, Pet08]. Sardar [Ful01c]. Sassi [Tot04]. Saturnino [Edg05]. Saul [Gre02, Poo09]. Saur [Smi06c]. Saussure [Rud04b]. savante [Chi04b]. Savants [Chi03b, BVB02]. Scarecrow [Lin05]. Scene [Ber08, Del06b]. Scenes [Ben04b, Hei08, BI06, Mac02]. Schabas [Smi07b]. Schädellehre [Van02a]. Schaffer [Mac09a, Ril02]. Schemes [Aga01a, Sco98]. Schickore [Jar08a]. Schiebinger [Del06a, Rob06, Tos09]. Schlegel [Kni04, Cha02a]. Schmidt [Gou09a]. Schneider [Bru09, Seg05]. Schoeningh [Har01c]. Schoepflin [Fie05]. Scholar [Bro03b, Gre99]. Scholars [Pye09]. Scholarship [Gin04, Iro04]. Scholz [Set04]. Schönbeck [Høy05]. School [Hig05, vC02, ED06, Her06]. schools [Cro03]. Schrödinger [Phi02, Cec01]. Schultz [Mac08b]. Schultze [Aga03b]. Schütt [Smi02b]. Schwabe [Smi04]. Schwartz [Seg05]. Schweber [MW09a, Ren07]. Schwinger [Kai05a, MM00]. Science [Abr08, Ack08, Ade05, Aga03a, Aga03b, Aga08a, Alb01, Alb03a, AM01, And03a, And09, Ano00e, Ash04b, Aug09, Bar07, Bar08, Bau01, Bek06, Ben04b, BVB08, Big06, Bor03, Bow03, Bra07, Bra08, Bre04, Bre05, Bre06, Bro05b, Bro06a, Brö02b, Bro08b, Bro09, Bur03, CK02, Cam08, Can08a, Can04a, Can07, Can08b, Car01a, Car06, Cha01, Chi08, Chi04a, Chi06, Chi03c, Chi04b, Chr05, Chu02, Coo07a, Coo06c, Coo02b, Cor06, CLW07, D’A03, Das09, Del06a, DD08, Del09, Dep07, DiM07, Div01, Dix00, Dix05, Dix06, Dix08a, Dix08c, Dod06, Dol02a, Dol02b, Dol04a, Dol04b, Doo04, Edd03a, Edd04a, Edd07b, Edd08, Edg04, Elm02, tF00, Far09a, Far09b, Fet00, Fie05, For04, For05, Fox04, Fre04d, Fre05, Ful01c, Ful03, Ful09a, Fyf02b, Fyf02a, Fyf03a, Fyf03b, Fyf04, Fyf05, Gas98a]. Science [Gau09, Ghi01, Gil04, Gil08, Gil09, Gin04, Goo05, Gou03a, Gou03b, Gou03c, GG02, GG03a, Gra06, Gre02, Gui04, Haf04, Haf08, Hal04b, Hal06, Ham07b, Har00a, Har04c, Hei07b, Hei07a, Hei08, Hem02b, Hem03, Hen01, Hen03a, Hen04c, Hen03b, Hes01, Hes04, Hig05, Hig06, Hil00, Hoc08, Hsu06, Hug01, Hug07a, Hug07b, Hug08, Hul07, Hun05a, Hun07a, Iqb07, Iro04, Jac09, Jam03, Jam05, Joh07a, Joh02d, Jor01, Jor02, Jor06, Jor08, Kem09, Kim07, Kle04, Kle09, Kni02, Kni03, Kni05, KE05, Kni06a, Kno08, Kra07, Kre06, Kus07, Lau06, Lea04, Lin05, Lin07, Liv02, Liv03, Loc02, Luc05, Mac00a, Mac00b, Mac01a, Mac05, Mar05, Maz07a, Maz09, MD99, McO09, MS02, Mor97b, Mor97a, Mor05b, Mor06c, Mor02c, Mor08, Mül01, Mun09]. Science [Mus03, NFJ97, Nai06, Nay09, NW08, Nes02, Nes06d, Nes09, Noa09, NB02, Ols08, Ost02a, Ost02b, Out04, Pag03, Pal06, Par01a, Pau09, Phi02, Pic07, Plo05, Poo09, Por03, Pro08, Pye06, Pye09, Ram02, Ras06, Rei07b, Ric04, Ric09b, Ric01a, Ric08, Ril00, Ril01a, Ril01b, Ris02, Rob07, Rob09b, Roo08, Rud07, Rup08, Sam07, Sch03a, Sch02b, Sec08, Sha03b, She00, Siv06, Siv08, Sle01, Sle02, Sle05, Sle09a, Sle09b, Smi06d, Smi06f, Smi07a, Smi09a, Sta03, Sta07a, Sta06c, Sta07c, Ste05a, Ste09a, Ste06, TS01, Ter08, Tho03b, Til04, Tim05, Tim08, Top08, Tot04, Tso05, Tso07, Tur06a, Tur06b, Van02b, Vic07, Wag05, WM08, Wat97, Web08, Wee08, Wer09, Wil09b, Win08, Yam00, Yeo01b, Yos00, Zil03, çG00b, çIh04, dCH04, dC08, AAH00a]. science [Arn06a, AAH00b, Ast01, Bah01, BV00, BVB03, Ber01, Bir01, Bod06, Bow06b, Bro08a, Bro06c, BM05, Bro01a, Bro05c, BG02, Can05, Cec01, Chi01a, Coe07, CC07, Coo05b, Coo07b, Cro01c, Das04, Dea06b, DS06, Dra00, Dub06, Dun09, Efr07, ED06, Elm05, Elm06, Ene09, Fal05, Fan03, tF04, Fin05a, FH00, FN06, Fro05, Gau06, Gay07, GL00, Gol98, GHL03, Gou99, Gov02, GWW04, Hac08, Hai01, Hal04c, Har07b, Har03b, Har05b, Hay02a, Hea99, HS02, HBB^{+}03, Hes00, Hoc03, HD02, Hol08, Hol00, Hor07, How00, HB02, HK01, Hun07b, JS04, Jea07, Joh01, Joh06b, Jon08, Jor00, Kai05c, Kas01, Kne00, Kni06b, Kri06, Kum07, LC01, Law05, LRS03, Lig07, LN03, LHL00, LHN99, Llo04, Llo06, LS05, LM02, Mac00c, May05b, Maz00, McC02, McL09]. science [Mit05, MDF02, Mor09a, Mor05c, Mus06, Nay03, Nay05, NS01, O’C07, PP07, Pic01, Por09, RT00, Rei08, Rei05c, Rel02, Ric02b, Roc01, Rud04a, Rup00a, RH04, Sas01, SS05, Sha03a, SA06, She01, SL06, Sim04, Siv05, Smi00b, Sny06, Sok08, Spo02, Spr06, Sta06a, Sta07e, Tei00, Tuc05, Tur03a, Tur02b, WPM05, Wei01, Wei02c, Whi03, Wig08, Wit01, WHF08, Wri00, Zin05, ZH06, Gos00, Mac02, çG00b, RH05]. science-industry [GWW04]. Sciences [Chi03b, Cla09, Fyf02b, Hen04b, Hen03b, Hoc09, Kus03, Mor06c, Mor07c, Por01, Ril02, Sle02, Smo08, Sta06c, Sta07c, BVB03, Bow05, Can05, Chi04b, CGS99, Coo05a, Dea01, FSJ00, HW01b, Kle01, Loc07, Mar04, Maz00, Rud04b, Rud04a, Rud05b, TBbTc90, Top00, UU07, Vid05]. Scientific [Bar00a, Boh01, Bro08b, Bud09, Cro07, Das07, Dol04b, Dun08, Edd05b, Fra05, Ful02, Ful09b, Fyf02b, Gil08, Goo06a, Goo06b, Haf04, Hen06b, Hen07a, Hig02, Hig05, Hig08, How00, Hun05b, Hut04, Joh03a, Joh06a, Kai05a, Kni03, Kni06a, MPB00, Mag06b, Mor09b, Nye09, Pad01, Pan06, Sec08, Set04, Sha07, Sha09, Smi06d, Söd07, Sta07a, Sta08b, Ste04b, Tyb02, Wee07, Wee09b, Whi04, Yam04a, Yam04b, ABC06, Bai04, BS02, BSD08, BP05, Cad00, Cha04, Das00, ED06, Far00a, Fin04, Gau06, GB02, Gol04, Gov02, Har07a, Har02a, HM02, HG07, Her01, Hoo02, Hul00, Jon06b, Kni06b, Lev02, Lig02, Liv05b, MM00, Mor05a, ML07, New03, Por04, Rad03a, Ren05, Row02, Sel08, Sha08b, Ste09c, Tho07, Van04b, Var00, Wad03, WS^{+}01b, Yeo01a, Zic00, vC02, SEM04]. scientifica [Cor04b, Gov02]. scientifique [Chi03c, HB02]. scientifiques [Chi01b, Hul00]. Scientism [Wil09b, Ols08]. Scientist [Boo07, Fox07, Hor03, Jam01a, Mok06, Sch08a, Tho02b, Wil06, Hah05, MT07, Mea04, Moy97, Pey01, Rig00, SM05a, WR04]. Scientists [Jor02, Mor06a, Pal00, Smi08f, Tur02a, Yam04b, BW06, Ham01a, MMMM02, Sch05b, Tha99, Var00, BVB02]. Scienza [Fin09, Ber07, Cor04b, Gov02, Pug09]. Scotland [Edd03a, Edd06, Elm02, Joh04b, Mil04, Wit01, CSML04, Fin05a, HK01, LS05]. Scott [Aga01a, Fin03a, Pic07, Sch08a, Ver04, Fri04]. Scottish [Edd02, Edd04b]. Scriba [Ric04]. Scribe [Pre03]. Scribner [Sha09]. script [Ahm84]. Scriptural [Rob03, Lyn02, Rei05a]. Scuola [Man00b]. Sea [Haf08]. Sean [Aga08a, Kre06]. Search [Aga06, Ben06, Col04, Hue03, Ken08, Sch00]. Seattle [Siv08]. Sebastian [Mos08a, McL07]. Second [Can02b, Coo07a, Edg04, Elm02, Hop04, Mor03b, Smi03b, Yam04b, HK01, Cro01c, Har03b]. Secondary [Hen04c, Lin05, Ost02b]. Secord [Top02]. Secrecy [ETV04, She07, Bia06, Lon01]. Secret [Aga04, Elm02, Log08, Top02, Joh02a, Sec00, Sen06b]. Secrets [Cop04, Hig04a, NG01a, Pag03, LFT02, Tur03a]. section [Top00]. Secular [Smi06g, Dix03b]. secundus [McG08]. Security [Tho03c, Pol02, Tur03a]. seditious [Chi03a]. Sedley [Leh09]. seducción [Edg05, NG01d]. seduction [NG01d]. See [Mor05b, Smi06e]. Seeds [Pal06, Bow03]. Seeing [Aga01a, Sco98]. Seicento [Pug09]. Selected [Aga04, End04b, Mor07a, Tim08, BBW05, Deb06]. Selection [Gar03b, Mil02, Ric06, BC00, Bla07, Bul04, Jon02b, Ken07, WC02]. Selective [Har03c, WO01]. Selectivity [Fra02, Goo06b]. self [Bor07]. Selin [D’A03, Plo05]. Selling [Bar00b, Mor01b]. Selves [Maz04a]. Seminal [Cop04]. Seminar [Set04]. Semitism [Can07]. Sengoopta [Log08]. Sensation [Top02, Sec00]. Sense [Hei07a, Dea06b, She01]. Sensibility [Edd07a, Out04, Poo09, Dub06, Ris02]. Sent [Por03]. Sentimental [Out04, Ris02]. Separating [Roo08]. September [Ano00a]. Septentrion [Chi01b]. Serafina [Høy01, Rob09c]. Series [Ade04, Aga04, Bek06, Bre04, Cam08, Coo03, Coo06c, Edd03b, Edd05a, Edd07a, Emm01, End04a, Fin03a, Gol05, Gou03a, GG02, Hal04a, Han07b, Hun02, Jen01, Joh02c, Jon04, Kle09, Kno01, Lin05, Mun08, Nut04, Ril00, Ril01b, Sam07, Ste05b, Dea09, How00]. sermons [Sim04]. Service [Gas98a, Mac00b, Maz07a, JS04]. set [Bow06a, Noa09, Raj09]. Settecento [Fin09, Ber07]. setting [WHF08]. Seuil [Spr06]. Seven [Goo04a]. Seven-Year [Goo04a]. Seventeenth [Dea08, Dew01, Dix00, Elm02, Lüt04a, Ber06c, Bos02, Che06, Gou99, Ken08, Smi08g, Smi08h, Cle00, HK01, Man03, Mil00]. Seventeenth-Century [Dix00, Elm02, Bos02, Gou99, HK01]. Sex [Daw06, Log08, Ric06, Ben04a, Sen06b]. SFR [Sta03, Høy05, Ric04, Smi04]. Sha [tF00]. Shadow [Aga06, Col04]. Shadows [Kre06, Joh01]. Shaftesbury [Del09]. Shakespeare [Hob06, Bic03]. Shamdasani [Smi06f]. Shank [Bax09]. Shape [Art04, Stu09, Mit07]. Shapin [Bud09]. Shaping [Bro08b, Cec01, Luc05, Lüt04b, Phi02, Gau06, Hal02b, Rel02]. Shared [Cla09]. Sharon [Gou03c, Ben04b, Web07]. Shea [Hei08, Jor01]. Sheets [Bro02e, Car99]. shells [Die06]. Shermer [Dol02b]. Sherrington [Smi00b]. Sherwood [Gor02]. shift [Smi08g, Smi08h]. Shifts [Har08]. Shiman [Sum06]. ship [Loc07]. Shipley [Yam00]. Shock [Far07b]. Shoolery [Tho03b]. Short [Far09a, Byn08, Dix08c]. Shorthorn [Whi06, Der03]. Should [GB02]. Show [Rup08, O’C07]. Shteir [Jor08]. Siderius [Hen05b]. Sidgwick [Mac08b, GS02, Sch04b]. siècle [HP00, NG01b]. Siegfried [Kle05]. Siegmund [Aga03b]. Siegmund-Schultze [Aga03b]. Sight [Elm02, HK01]. Sigla [Ste04a]. siglo [Fox05]. Siglorum [Ste04a]. Sigmund [Eag05b]. signalling [Hom09]. Sigrist [Rud04b]. Silent [Lam02, Rus01]. Silke [Eag06a]. Silvan [MW09a]. Silvanus [GB02]. Silver [GG03b, Chi02c]. Silverman [Coo06a, Qui04]. Silverthorne [Man01a]. Simian [Smi08d, Rad07]. Simon [Fox07, Hah05, Har00a, Mac09a, Pag01, Pye09, Rob07, Smi07a, Tra07, Yam04b, Alb01, Sen02b, Tra01]. Simone [McG08]. simple [Sta07f]. Simpson [Edd06]. since [Bar01a, Der03, Edd03a, Far07b, KCH00, Mac08c, Whi06, Wit01]. Sinclair [Smi03a, Ewi01]. Singapore [Mor09b]. Singer [May05a]. Single [Noa09]. Singular [Maz04a]. Sinica [Chu02]. Sinkora [End04b]. Sir [Ash01, Coo06b, Hut07, Hut09, Maz07a, Mil02, Tot04, BC00, Bas04b, CDK98, Hed08, Key05, Liv03, MS05a, TR06, Woo06]. Siraisi [De 07, Nut05, Van02b]. site [Ken08]. Sites [Far00a, Ric09b]. Sivasundaram [Ste06]. Six [Bur08, Sle07]. Sixteenth [Kus03, Smi08b, Ben03, HW01b, Mos07]. sixties [Aga08b]. Size [Hen04d, RP02a]. Skies [Cla01, Jan00]. skill [Sta01b]. Skilled [Shr09]. Skin [Jen01]. skrifter [Whi05]. Skull [Fra09]. Sky [Dic03, Joh05c, And03b, BZ04]. Slatkine [Rud04b]. Slava [Cla06]. Slaves [Spa03, Rob02]. Sleigh [Bur08]. Slopes [Tso07, Rei05c]. Slotten [Hem02a, Web06]. Small [Sle05]. Smil [Fin02]. Smith [Daw07a, Far07a, Gou09a, Hen06b, MW07, Nes02, Rei07a, Ric09a, Yos00]. Smithsonian [Jam01a, Jam01b, Jam05]. Snowden [Web09]. Snyder [Dix08a]. So-called [Ste04a, EoToBB01]. sobre [Med00]. sociability [Nes06d]. Social [Aga03b, Bor03, Chi06, Coo03, Coo06b, Gis09, Hen03a, Mac05, Nes06b, Rad02, Smi08e, Ste06, Stu06, Zil03, AM01, Bur01, Hau03, Maz08, SS01, Tur02b]. Socialism [Chi02b, Sta03, Hen96, Sto00, Zim00b]. Société [Cor05, Cha02b, Chi03b, Chi03c, Chi04b]. societies [Fin05a, Joh00]. Society [Ade06, Ano00e, Bre05, Bro04a, Can04b, Can04c, Cla04a, Dig08, Dix03a, Dix08a, Dol01, End04a, Goo02c, GG06b, GG09, Gre02, Hen07a, Jor02, Kha06b, Kle05, Kle09, Lüt04b, Mor05b, Mor07a, Mor01a, Nut05, Spa06, Wal03, Wer04, Whi05, AA01, Cha02b, EG04, Ene09, LTGS02, MF06, NB02, Sny06, Zui05, Zui06, Bri04, Cai00, Gas09, Hal02b, Hun07b, RP02b, Win09a, Mor09a]. Sociology [Gis09, Maz08]. Söderqvist [Nye09, dC08]. Sofia [Whi05]. softcover [Aug09]. soi [Maz04a]. Sokal [Ful09a, Rou07]. sol [Har00c]. sol-lunar [Har00c]. Solar [Big06, Cav08, Hee08, Pan02, Spo02]. Sold [Por03, MS02]. Sole [Pug09]. Solly [Hor03, Pey01]. Solomon [GG06a, Wee07]. Solve [Sen06a, Lee02]. Solvers [GG04, Yan02]. Some [Dav09a]. Somerville [And03a, Bro06a, Bro06b, NS01]. Sommer [MW09b]. Sommerfeld [Sta04]. son [BV00, Chi01a]. Sons [Sha09]. Sonu [Smi06f]. Soojung [Big06]. Soojung-Kim [Big06]. Sophie [Ful09a]. Soraya [Lau06]. Sörlin [Haf04]. Soto [Han07b, Wal04]. Souls [Smi06e]. sounding [TDNS08]. Sounds [Sta03, Bar00c]. Soundscape [Pem04, Tho02a]. Source [Kha06b, EG04]. Sources [Brö02b, Hen04c, Sta03, Hen96, Mcg06, Ost02a, Ost02b]. South [Das09, Poo09, Dub06]. Southern [Chi08, Gre02, Jea07]. souvenirs [Coo05a]. Soviet [Chi04a, Cla06, Hal04a, Bir01, Ger02, Wei00]. Space [Aga04, Mud01, Sha03b, Sta03, Stu09, Bar00c, Die06, Joh02a, Liv02, Mil01, Mit07, Mor00]. space-time [Bar00c]. Spada [Por01]. Spanish [Hal04b, McC02, NG01d]. Sparks [Rad03b, Str00, Ber06d]. Spary [Sch01]. Spatial [Mor06e]. Speaking [Lin05]. Special [AAH00b, Mor07a, AAH00a, Her05b, Top00]. Speciation [Cai00]. species [Mco01]. Specimen [Ric01b, Key00]. Specimens [Fin03a, SFS^{+}00]. Spectacle [Wer09, BVB08]. Spectrum [Jac00a, Joh03b, Jon04, Mor02c, Hen02b]. Speculation [Wil09c, For08]. Spencer [Ren06, Ren09, Hal06]. Sphere [Ber06a, Bre04, Yeo01b]. spiritistic [Sta01b]. Spirits [Des01, Sut03]. Spiritual [DM05b, Les03]. Splashes [Bow04]. Spontaneous [Brö04b, Rad03b, BS01, Har02b, Str00]. Sprat [Mor09a]. Spreading [Her06, Kip04, Wor00, Gov02, Kre03].

Spring [Lam02, Rus01]. Springer

[Mcc05, Mos08b]. Springer-Verlag [Mcc05]. Stachel [Sta03]. Stack [Fra09]. Stadler [Tso05]. Stage [Sch03a, Hil00, Kuk06]. staging [Liv03]. Staley [Sta09b]. Stamhuis [Mag06a]. standard [RN09]. Standardizing [Sel08]. Standards [Shr09]. Stanford [Big06, Car06, Cha03, Fri01, Har08, Jor02, Lau06, Mor01b, Sch03a, Ste05a, Tos09, Vic07]. Stanley [Bro09, Maa05, Mir07, Van05]. Stanton [Smi05b]. star [Sta07f]. Stargazer [Hur07]. Starkey [Edd05b, NP02, New03]. Stars [Hes01, Nea00a, Sta07b, Hol99, Mil05, Sta01a, Wil97]. State [Aga01a, Coo06b, Daw06, Dig08, Gas98a, Mac00b, Mac09b, Mor08, Wil09d, Bas04b, Edg06, Mae06, MF06, Sco98]. States [Bar01b, Hal04b, Hem02a, McC02, Tim05, Dun99, Slo00]. Statesman [Wil06, WR04]. Statile [Bru06]. Statistical [Mag06a, Mag06b, Por04]. Statistics [Ren07, Sch06b]. Steam [Bro04a, NG01d, LTGS02]. Stedall [Hut07]. Steele [Loc02]. Stefan [Coo06c]. Stein [Sch00, Smi02b]. Steiner [Nes01, Ste04a]. Steinke [Edd07a]. Steinle [Kra02]. Stellar [Hes01, Hol99]. Stepan [Dri03]. Stéphane [Bur07a]. Stephen [Biv08, Gar03b, Hen04a, Van04a, Aga04, Alb03b, Bor03, Bow04, Bro08b, Can07, Mos08b, Rad00a, Sch03a, Sha02]. Stephenson [Ben02, Dou06, Bai03]. sterbende [Man01b]. Stereochemistry [Mor06e]. sterling [Hug00]. Steve [Dix09, Gat05]. Steven [Bre06, Bud09, Fre04d, Joh05c, Sle09a, Spa06, Sta09a]. Stevens [Gou03b, Mor03a, WS01a]. Stevenson [For03, Mor06d]. Stewart [Bro04a, Maz07a, Mor09b]. stick [GB02]. Stift [Hör78]. Stilgeschichte [BSD08, Bru09]. Stillman [Han07b]. Stimulus [Kus00, Met98]. Stink [Ham07a, Bar06]. stirpium [Hen05b].

Stocksfield [Ren09]. Stokes [Chi02b]. stone [Goo02e, Sch00]. storici [Cor04b]. Stories [Hay02b, FS03, Mac00d]. Story [Chi04a, GG03b, Pad01, Tim03a, Tim03c, Tim03d, Yeo04, Bir01, Cad00, Chi02c, GC00, Kre02, tH02]. Stow [Mor03a]. strange [Lig02]. Strategic [RS02, Sum06]. strategy [Lig02]. Strick [Rad03b]. Striving [Wal06c]. Strong [Goo02b]. Stroud [Hes03, Joh05b]. Structure [Bar01a, Mor06e, Sta08b, ABC06, KCH00]. Structures [Pye06, Spr06, Hal04c]. Struggle [Ham07a, Bar06]. Stuart [Gou09b, Hal04b]. Stuchtey [Rei07b]. Studi [Cor04b, Man00b, Pan06]. Studia [Gol05]. Studien [Par01b, Wei02b, Zig04, Mei00]. Studies [Aga03a, Aga03b, Bar01b, Bek06, Bow07, Bre04, Brö02b, Cam08, Chi06, Cla04b, Col01b, Coo06c, Coo02b, Dea09, Doo05, Edd03a, Edd03b, Edd05a, End04a, For03, Ghi01, Goo05, Gou03a, Gou03c, Gra06, Gre02, Gui04, Hal04a, Han07b, Har01a, Har06, Hem02b, Hem03, Hen03a, Hen03b, Hes01, Hig05, Hoc08, Høy01, Hug07b, Hul07, Hun05a, Joh07a, Jor01, Kem09, Kir05, Kle04, Kni02, Kni03, Kni06a, Kra02, Kus03, Lam02, Loc02, Man00a, Maz07a, Mor06c, Mor02c, Mos08b, MW07, Mül01, Mur02, Nes02, Pag03, Pan06, Ram02, Ric04, Ric08, Ril00, Ril01b, Sam07, Sle01, Smi09a, Spa02, Sta03, Ste05b, Stu06, Til04, Van02b, Wei02a, Whi05, BR96, CH05, Clu06, CKM99, Dun09, FH04, Hac08, Hol99, Hul00, Llo03, Pri07, Rud04a, Rud05b, Rup00a, Sch06a, SS01, CC07]. Studio [Alb03a, Hop02]. Study [Cai00, Fyf02b, Har00b, Joh04c, Leh08, Lüt04a, Cle00, Joh04a, Mei00, Mor02b, TBbTc90]. Stuffed [Alb03b, Asm01]. Stukeley [Can04a, Hay02a]. Stuttgart [Eag05a, Nes01, Ste04a]. style [BSD08]. Styles [Cla09]. Subjects [Hob06, Qui04, Bic03, Sil01]. submarines [NG01d]. submarinos [Edg05, NG01d]. subscription [UU07]. Subtext [Cuo02, FU01]. Such [Chi02c, GG03b]. Suche [Sch00, Smi02b]. Suffering [Lin07]. Suffolk [Hul07, Nea00a]. Sufian [Web09]. Sujit [Ste06]. Sun [Big06, Pan02]. Sundberg [Del08]. Sundials [Eag06a, NHA^{+}02]. Sungook [Hem03]. Superior [Fox05]. Superiori [Man00b]. Supplement [Bro02e, Hod06a, Qui02, Bra00, Cuo02]. supply [Gra07a]. support [Big01, Mai09]. Surgery [Jen01]. Surgical [Hou07, Dar05]. Survey [Hei07b, Mor07b]. Surveys [Pet08, Igo07]. Susan [Mar05, Hig04a, Sch02a]. Susanne [Wei02b]. Suspect [Col01a, Sen02b]. Sutton [Hes03]. Suzuki [Dig08]. Svante [Mac01a]. Sven [Ras06, Hes01]. svenska [Hes01]. Sverker [Haf04]. Swade [Cro04]. swamping [Bul04]. Swan [Del06a]. Swedish [Dol02a, Hes01, ABB01, Hol99]. Sweeney [Wal03]. Sweet [Goo06c, Stu06]. Swetlitz [Lip08]. Sydney [Bro03d, End00]. Syllepticus [Hei08, BI06]. Sylvester [Can04b, Gra08b, Par06]. Sylvia [Haf08]. sympathy [Hed08]. Symphony [Sta03, Bar00c]. Symposia [Brö02b, çG00b]. Symposium [GWW04, Hig02, LHL00, Mac01a, Ras06]. Synthélabo [Chi01a]. synthesis [AH02]. Syphilis [Jen01]. system [Mor00]. Systems [Aga03a, Aga04, HH00, Rei05a, Ste09a, CLW07, Joh02a]. Szöllösi [Har04c]. Szöllösi-Janze [Har04c].

T [Alb03b, Bro05a, Hag04, Hen07a, Hig04b, Hod01, Mor07c, Roo08, Sch08a, Sle02, Wag05]. T. [Leh08]. Tackling [Coo06c]. Takada [Eag06a]. Tal [Goo06a]. tale [Mae06]. Talent [Ree08]. talents [Car07]. Talk [Jor06, Das04, Liv05b, May05a]. Tansey [Biv08, Jon09]. Tara [Roo08]. Tarski [GG06a]. Tate [Sle05]. Tatjana [Mai09]. Taub [Wil09d, Win08]. Tauber [Ghi01]. Tauri [Hig04b]. Taxol [Tim03a, Tim03d]. Taylor [Cla07a]. taylorisme [Fox04]. Tayra [Pye09]. Teaching [Joh03b, Hen02b, Hul00, HB02]. Teashops [Cro05a, Fer03]. Technical [Coo03, ETV04, Elm09, Gil08, BDLM07, BSD08, Cro01b, HP00, Lon01]. Technik [Par01b, Sta04]. technique [Bra08, HP00, NG01b]. Techniques [Joh03b, Hen02b]. Technische [BSD08, Bru09]. Technological [Dol04c, Edg08, Goo01, Zim00a, Wil02]. Technologies [Goo03, HHAH01, Led08]. Technology [Aga03a, Aga05, Aug09, Bar07, Bek06, Bre06, Bro05b, Brö02b, Chi02b, Chr05, Col02, Coo03, Coo06c, Coo02b, Cro01b, Cuo07, Dea09, Doo04, Emm01, Far07b, Goo02b, Gra06, Haf08, Hal06, Har05b, Hem02a, Hem02b, Hem03, HLS00, Hug07b, Hun02, Hun05a, Jac09, Joh02c, Joh04c, Joh06c, Jor02, Kem09, Kim07, Kle09, Kra07, Lea04, Loc02, Mac08a, Mac06b, Mor02c, Mor09b, Mül01, Nye06, Pag03, Ram02, Rei07a, Ric08, Ril01a, Rob07, Rob09c, Smi06d, Van02b, Wag05, dC08, Bra08, DS06, Gay07, Hac08, Hea99, Hug04b, Jon08, MS05b, MD99, MBF03, MDF02, NFJ97, Pic01, RV04, Slo00, Sto00, çG00b, çIh04]. Tècnica [Fox05]. Tectonics [Shi03b, Ore01]. Tedeschi [Daw00]. Teich [Har01b]. telegraph [Bek01, Mor00]. Telegraphy [Hun02, Bea01]. Teleology [Leh08, Gar03a, Joh05a]. Telepathy [Noa04, Luc02]. Telescope [Ben00, Hig04b, Gla97]. Telescopes [She07, Bia06]. Television [Hem02a, Joh02c, Sle09b, Bur00, Slo00]. Teller [Tho03b]. Telling [FS03]. Temperature [Kni06a, Cha04]. temporal [Ili00]. temps [Cha02b, Cor05]. Temptation [Hou07, Dar05]. Tempus [Joh05b]. Terms [Bra07, Elm05]. terre [Rud04b]. Test [War08, Cre06]. Testi [Pan06]. Testimony [Goo06a, Gol04, Liv05b]. Testing [Col00, Zen98]. Tests [Bru06, DGS03]. Texas [Hal04b]. Text [Cuo02, Elm09, Liv05b, BDLM07, FU01]. Textbooks [Par01a, Sim04, LBV00]. Textiles [Kle04, NG01c]. Texts [Eds04, Elm02, Fis04, Gou09a, Man01a, Man06, Maz07a, BRW04]. Thackray [Cla05, Mor05b]. Thad [Goo03]. Thagard [Pal00]. Theatrum [Pug09]. Theerman [Jam01b]. Their [Bra07, GG04, Har00b, Smi08f, Alb02, BW06, Elm05, Hig07, Hig08, LBV00, Par01a, Tur03b, Yan02]. Themes [Nay09, FJC06, WHF08]. Theodor [Mor03b]. Theodore [Hut09, TR06, Goo08a, Mag06b]. Theological [Kus00, Lin05, Met98]. Theology [Rad00a, Alt99, Val05]. Theoretical [Goo08a, GG02, Tyb04, Ara06, BJ00, Mag01, Set07]. Theorie [Har02d]. Theories [Joh02d, Kip04, Kra08a, Che00, Gol01, Har00c, Kai05b, Wor00]. Theory [Bor03, Bru06, Fri03a, Fri03b, GG03a, Gra05, Hen04e, Her05b, Shi03b, Web07, Bru01, Bul04, DGS03, Fin04, Jon02b, Ore01, Tay06, Tur02b, War03]. Therapeutic [Tim05]. therapies [Sha00]. Therapy [Hig06]. Thermodynamics [Coe09, Kra08b, Pap06]. Thermodynamik [Coe09]. these [Ade04, Mar04]. Thing [Bai04, Goo06b, Sta07f]. Things [Brö04b, Das04, Del08, Goo04a, Har02b, Jor06, May05a, Wal06a]. Think [Joh06c, Pem07, CP05, Hug04b]. Thinking [Ber06c, Dea08, Gol09, Mun07, Wit07, DM05a]. Third [Fyf02b, Har04c, Lau06, Ric08, dCH04]. Thoemmes [Maz07a, Mor06a, Rad03b, Rob03, Smi00a]. Thomas [Ade08, Aga03a, Ben00, Bre05, Far09a, Ful01c, Fyf03b, Gat05, Goo03, Hen01, Hur07, Smi04, dC08, Aga05, Ast01, Bar01a, Bar01b, FH00, Gla97, HHAH01, Joh06c, Mor09a, Nye09, Pet07b, Smi06g, Tur09b, Whi03, Wil08]. Thompson [Mor09b, Pem04, GB02]. Thornton [Fyf02b]. Thorpe [Hug09]. Thought [Bow07, Gou03a, Hut04, Mos08b, Wee07, Amu05, Clu06, Her01, Kas01, SM05b]. Thousand [Gil09, Far09b]. Three [Fin03a, Mor01c]. Thun [Joh02b]. Ti [Dod06]. Tides [And09, Rei08]. Tillberg [Sir09]. Tim [Sho03]. Timaeus [Leh08, Joh04a]. Timaeus-Critas [Joh04a]. Timaeus-Critias [Leh08]. Time [Bru06, Eag07, Mil04, Mor06c, Mor01b, Sta03, AAH00a, AAH00b, Bar00b, Bar00c, Cha02b, DGS03, Hom09, Hör78, Ili00, Mil01, Mor00, Nor05, RN09, Rud05a, SD00, Zim00b, RN09]. Timekeeping [Mor01b, Bar00b]. Times [Fri03a, Fri03b, Gat05, Har03a, Loc02, Buc02, Gue00, Ste00a]. Timmerman [Led08]. Timothy [Hal02a, Sle09b]. Tinbergen [Dew07, Bur05]. Tinemue [Ste04a, EoToBB01]. Tirard [Bur07a]. Tirnady [Sen06a]. today [BHR00]. together [oAEN04, Ber06b, Ber06a]. Tokyo [tF00]. Tongue [Smi08d, Rad07]. Tony [Siv06]. Tools [Car06, Hen03b, Kle01, KLG03]. Toronto [Jen02, Mel01, Pau09]. Torrens [Edd05a]. Torrey [Hay03]. Torsten [Dol02a]. Tortoise [Col02]. Tortured [Qui04, Sil01]. Tour [Ful09b, Hoc09, HG07]. Touwaide [Kem09]. Tr [Mer08, Ber06a]. tra [Pug09]. Trabucco [Pug09]. Tracks [Fre01, Mor07c, Fre04a]. Tractatus [Hei08, McG08, BI06]. Trade [Fyf02b, Hig01, Sum05, Ash03, How00, TBbTc90, Tur00]. Tradition [Biv08, Lip08, Sle01, Ste05b, Web08, CS06, Far00c]. Traditions [Joh02d, Sam07, Che00, Fre05]. Tragic [Hug09, Tho06]. Traister [Pag01]. Transactions [Goo02c, Gou06, Kle05, LM02, Mus03, Wal03, KG04]. Transfer [Bek06, çIh04]. Transformation [Dea08, Fin02, Jon09, Lev07, Roo08, Sha02, Ste05a, Stu06, Ale02, AH05, Ber06c, Gau01, Smi01a]. Transformations [Del08, Hem03, Mor02c, Pag03, Wal06a]. Transformed [Goo04a, Smi06e, Tso07, Rei05c]. Transforming [Hug01, Kni02, Lev01, Mor97a]. Transition [Goo02c, Ner05, CG00a]. Transl [Sir09]. Translated [Ber06a, Daw00, Dol02a, Goo02d, Hen04a, Mor03b, Seg05, Sir06, Smi08f, Tot04, Van04a, Yeo04, oAEN04, Ber06b, Bra00, Kim05]. Translating [Chu02, Wri00, NG04]. Translation [Hei08, Hig06, Mel01, Rup00a, Tyb04, BJ00, BI06, Bro02e, For05, Han07b, ML00, Smi02a]. translations [Hen02a, Man06]. Translator [Sta03]. Transmission [Chu02, Wri00]. Transmitting [KM06, Smi08c]. Transport [Coo06c, Ver04, DS01]. travaux [Spr06]. Travel [Dol02a, ABB01]. Traveling [Siv08, Arn06a]. Traweek [Gou03c]. Treasurehouse [Cla05]. Treatise [Eag05b, CE02]. Treatises [Eag05b]. Trends [Tso05, HS02]. Trevithick [Can04c]. Trevor [Hut09, Pem07, Ram02, Bro04a, Kni02]. Trevor-Roper [Hut09]. Triads [Pan08, Jac06]. Trial [Hei08, Joh09, BI06, Bur07b]. Triarco [Pan06]. Tribute [Chi06, AM01]. Trident [Joh04c, Mor02b]. Tried [Edd05b, NP02]. Trischler [Coo06c, Mül01]. Trismegistus [Smi05b, Lin03]. Tropical [Dri03, DM05c, Hag04, Har02c, Siv08, CW02, Hay01, Ste01]. Tropics [Gar03b, Siv08, Arn06a, WB02]. Trudi [Sle05]. True [Chi04a, Hes04, HD02, Mor01b, Pad01, Bar00b, Bir01, ED06, Ken08]. Trust [But06, Col02, Jon09, Mai09, Qui02, Goo04b, Mac01b]. Truth [Qui04, Ram01, Sil01]. Trying [Bur07b, Joh09]. Tsvetnaia [Sir09]. Tuberculosis [Ade04]. Tübingen [Kus00, Met98]. Tucker [Pro08, Tot04].

Tully [Fyf02b]. Turin [Fin09, Vid05]. Turing [CK02, Hod06b, Cop04]. Turkish [Brö02b, çG00b]. Turned [Seg05, Sch03b]. Turner [Bor03, Bro04a, Eag06a, Hig01, Hig07, Jen01]. Turnhout [Fin03a, Hig05]. Turton [Mor06d]. Tuscan [Bos02]. Twentieth [Ade04, Fre04c, Har08, Hem02b, Mac07, Smi05a, Tho03b, Til04, Tim04, Tur06a, Bow06b, Dav09b, Ene09, Jam08, NG04, Ren05, Sta03, GR00, Hol00, Nye04, TS01, Wei02d]. Twentieth-Century [Tho03b, Til04, Tim04, Bow06b, Dav09b, Ren05, Sta03, TS01]. twins [Bur01]. Two [Aga03b, Ber06a, Goo06b, Hes04, Mar05, Fra02, GHL03, HD02, Mar04, Maz07a, SS01, Wig08, Rei05a]. Two-Culture [Mar05, GHL03]. TX [Hig04b]. Tycho [Smi08b, Mos07]. Tying [Sta01b]. Tyndall [Mor09b]. typological [Sco04]. Tyrrell [Pen04].

U [Far07a]. U.S [Aga04, Joh05c]. U.S. [Dic03]. Ubiratan [Plo05]. UCL [Qui02]. Ugglan [Hes01]. Uitgeverij [Rav00]. UK [Mac09a]. Ulf [Mac01a, For05]. Ulrich [Smi04, Har01c]. Umeå [Whi05]. uncertain [Bor07]. Uncertainty [Mun09, Coe07, Pen06]. Unconceived [Vic07, Sta06a]. Understanding [Bro06c, Nut04, Sec08, Top08, Kni06b]. Unfinished [Mel01, Sta03, Bar00c, ML00]. Unguru [Cuo02]. Unification [Tay06]. Uniformity [Dep07, Fal05]. Unimaginable [App05b]. Union [Big01, Edd07a]. United [Bar01b, Hem02a, Tim05, Dun99, Slo00]. Unity [Dep07, Smi09b, Fal05, Ken07, Llo07]. Univ [Ber06b]. Universal [CK02, Hen04a, ED06]. Universality [Nay09, FJC06]. Universe [Ben02, Gou03a, Mac08b, Smi07a, Gre05, Kas01, Sch04b, SBF00, Val05]. Università [Goo05]. Universitaires [Bur07a, Chi04b, Chi01b, Chi03c].

Universitat [Fox05, Med00, Man01b, Wei02b, Zim00b]. Universities [Nut05, Sir01]. University [Abr08, Ade08, Aga01a, Aga01b, Aga02, Aga04, Aga06, Aga08a, Alb03b, And03a, And09, App05a, Art04, Ash01, Bar01a, Bar01b, Bar08, Bax09, Ben02, Ben06, Ber06a, Ber08, Big06, Big07, Biv06, Biv08, Bla03, Boh01, Boo07, Bor03, Bow07, Bra00, Bra07, Bre05, Bro04a, Bro05a, Brö04b, Bro09, Bru06, Bud09, Bur08, But06, Can09, Can02b, Car06, Cha01, Cha03, Chi02b, Chr05, Cla01, Cla04b, Coh09, Col00, Col01b, Coo07a, Cor00, Cor06, Dan01, Das07, Daw00, Daw06, Daw07a, Dea06a, Dea07, Dea08, Del06a, Del08, Del09, Dep07, Dew01, Dew07, DiM07, Dig08, Dig09, Dix00, Dix05, Dix06, Dix08a, Dix09, Dod02, Dod06, Dol02b, Dol02c, Dol04a, Dol04b, Doo04, Doo05, Duf09, ETV04, Eag05b, Eag06a, Eag06b, Edd03a, Edd04a, Edd05b, Edd07b, Edd09, Edg08, Eds04, Fan02, Far09a]. University [Far06, Fis04, For03, Fox07, Fra05, Fre04c, Fre04d, Fri03a, Fri03b, Fri04, Ful01a, Ful02, Ful03, Ful09a, Ful09b, Fyf02a, Fyf05, Fyf07, Gan07, Gar03b, Gat05, Gau08, Gau09, Gil09, Gol09, Goo01, Goo02a, Goo02b, Goo06a, Goo08a, Goo09a, Goo06b, Gor09, Gor02, Gou09a, Gou09b, Gra08a, GG03a, GG06a, Gra05, Gra08b, Gre02, Haf08, Hag04, Hal02a, Hal03, Hal04a, Hal04b, Hal06, Ham07a, Ham08b, Han07a, Har01a, Har00a, Har00b, Har05a, Har02c, Har03c, Har06, Har08, Hay03, Hay06, Hei07b, Hei07a, Hei08, Hem02a, Hen06b, Hen07a, Hen04d, Hig04a, Hig01, Hig07, Hoc08, Hod01, Hod06a, Hou07, Høy01, Hug01, Hug07a, Hug07b, Hug09, Hul07, Hun05b, Hut04, Hut07, Hut09, Jac09, Jam03, Jar08a, Jen02, Joh03a, Joh07b, Joh09, Joh03b, Joh05c, Joh06c, Jon04, Jon06a]. University [Jor02, Kai05a, Kas02, Kha06b, Kip04, Kir05, Kni02, Kni04, Kni05, Kni06a, Kni08, Kno08, Kra08a, Kus07, Lam02, Lau06, Leh08, Leh09, Lev07, Lin07, Lip08, Liv05a, Log08, Lüt04b, Lyn09, Mac00a, Mac00b, Mac08b, Mac09b, Mac09a, Mac05, Mac07, Mac08c, Mag06b, Mai09, Man01a, Mar05, MW09a, Maz02b, Maz07a, McO09, Mel01, Mer08, Mir07, Mor03a, Mor06c, Mor06b, Mor07c, Mor01b, Mor08, MW07, MW09b, Mun07, Mun09, Mur02, Mus03, Nay09, Nes02, Nes06b, Noa04, Nut04, Olb02, Out01, Out04, Pag01, Pal00, Pal02, Pau09, Pem07, Pen04, Pet07b, Pet09, Pet08, Phi02, Poo09, Por01, Por03, Pro08, Qui04, Qui08, Rad00a, Rad02, Rad03b, Ree08, Rei07b, Ren00, Ren06, Ren07, Ric09a, Ric09b, Ric01a, Ric01b, Ril01a, Ril02, Ris01, Rob05, Rob06]. University [Rob08, Rob09c, Rol06, Roo08, Rup08, Rut08, Sch08a, Sch03a, Sch02a, Sch01, Sch02b, Seg05, Sen02b, Sha02, Sha03b, Sha07, She07, Siv08, Sle01, Sle03, Sle05, Smi01b, Smi02b, Smi03a, Smi05b, Smi06f, Smi06g, Smi07a, Smi07b, Smi08b, Smi08a, Smi08d, Smi08f, Smi08c, Spa03, Sta03, Sta07a, Sta09a, Sta08b, Sta06b, Sta06c, Sta07c, Sta07d, Sta09b, Ste05a, Ste09a, Ste09b, Ste07, Ste04b, Ste06, Stu06, Stu09, Sum05, Sut03, Suz05, Tho02b, Tho03c, Tim02, Tim03a, Tim03b, Tim03c, Tim03d, Tim04, Tim05, Top02, Top08, Tos09, Tot04, Tre06, Tso07, Tur09a, Tur06b, Tur08, Tur09b, Twe09, Tyb02, Tyb04, Urr07, Van04a, Van05, Ver04, Vic07, Wag05, Wal06b, War08, Web06, Web07, Web09, Wee08, Wee09b, Whi06, Whi04, Wil09d, Wil09c, Wil09b, Win08, Wor07, Yam04a]. University [Yam04b, Yeo04, Yos00, Mor97a, Fri01, HW01a, Zim00b]. Unknown [Aga04, Hig04b]. Unmaking [Tos09, PS08]. Unpacking [Ham01b]. Unpublished [Can02b, Maz07a]. unseen [Sco04]. untitled [Sch84]. Unveiled [Ben02, SBF00]. Uplink [Aga04, Mud01]. Uplink-Downlink [Aga04, Mud01]. Upper [Goo08b]. Uppsala [Joh07a, Kni03]. Ural [Mor07b]. Urban [Her05b]. Urbana [Dol02b, Wil09b]. Urs [Fie05]. Ursula [Car06, Edd08, Hen03b]. Ursus [Man00b]. US$19.95 [Fri03b]. US$22.50 [Bla03]. US$27.95 [Tim03d]. US$94.95 [Aga03b]. USA [Haf04, Hei08, Nay09, Roo08, Smi09a, Tim08, WM08]. useful [ED06]. Users [Hig08]. Uses [Gas98a, Joh02d, Mac00b, Che00]. using [Win09a]. utility [Ber06d]. Utopia [Sch01, Spa00]. Uwe [Har02d, Her05b].

V [Aga04, Chr05, Edd07a, GS02, Jor02, Med00, Mor07a, Pet07a, Rud04b, Smi08e, Wag05, Can02b, Dix09, Kno01]. V. [Hug00]. vaccination [Dur05, Wor07]. Vaclav [Fin02]. Vadim [Chi04a]. Vaiana [Big01]. Vailati [Har00b]. Valais [Rud04b]. Vale [Jor02]. Valencia [Med00, Med00]. València-C.S.I.C [Med00]. Valencianos [Med00]. Validation [Sta07a]. value [Smi00b]. Vanities [Cla07b, Gou09b]. vapores [Edg05, NG01d]. Variations [Smi09b, Llo07]. Variorum [Bek06, Bre04, Cam08, Eag05b, Edd03b, Edd05a, End04a, Gou03a, Han07b, Ril00, Ril01b, Sam07, Ste05b]. Various [Hut09, TR06]. Vatican [Ade08]. Vehicle [Goo02a]. Veil [Mer08, Had06]. Vera [Bru09, Elm09]. Verbeke [McG08]. vergleichenden [App05b]. Verhandlungen [Har02d]. Verlag [Aga03b, Coe09, Eag05a, Fie05, Joh02b, Mcc05, Nes01, Nes06a, Nes06c, Nes07, Par01b, Sta04, Ste04a, Web08, Zig04]. Verlag2008 [Bru09]. Verlagsgesellschaft [Kra02]. Vermeer [Ben06, Hue03]. Vermij [Gin04]. Vernacular [Gil08]. Version [Ste04a, EoToBB01]. Verso [Gar03b, Pal06]. Versuch [Kra02]. Versuche [HS98]. versus [BVB02]. Very [Far09a, Byn08, Dix08c]. Vestiges [Top02, Rup00a, Sec00]. Vetenskapsbärarna [Hes01]. VI [Eag05b, Hen05b, Man00a, Biv06, Brö02b, Bro03b, Bro03c, Can07, Coo06a, Coo06c, Del06a, Fyf02a, Gor09, GG06a, Hun07a, Leh08, Mok06, Rob07]. Viaggio [Ber07, Fin09]. Vicki [Hes04]. Victor [Goo02d, Mor07c, Goo02d]. Victoria [Fra09, Pre03, Ras01]. Victorian [Bar08, Ben00, Big06, But06, Can09, Daw07a, Dix08a, Fra09, Fyf03b, Fyf07, Gra08a, Gra08b, Ham08b, Hut04, Jon06a, Mac05, McO09, Mok06, Pro08, Rad03b, Ril01b, Rud07, Smi01b, Smi06d, Tho03a, Top02, Tur06b, Tur08, Tur09b, Web06, Whi04, Yos00, And03b, Ast01, Coh07, Coo05b, Dau05, Daw07b, Dix08b, Fan03, Fic04, Fin04, Fin05a, Fyf04, Gla97, Goo04b, Her01, Lev02, Lig07, Mac00c, Mar04, Mea04, Mor05c, Mor00, Ott08, Pan02, Par06, Pet04, RN09, Ryl00, Sec00, Smi06a, Sny06, Str00, Tuc05, Van04b, Wri01]. Victorians [Edd07b, Mor07c, Fre04a]. Vienna [Ful01a, Har01b, Mun09, Ren05, Tso05, Coe07, Hac00]. View [Nes02, Zic00]. Viewed [Ash04b, Hos03]. views [HP02]. vii [Col00, Elm02, Far06, Goo08b, GG09, Hal03, Har04c, Mun07, Par01a, Pye09, Ren06, Rob08, Set04, Sle05, Sta07d, Wei02a]. viii [Ash04b, Bar01a, Bee01, Bro04a, Bur08, De 07, DiM07, Dol02b, Dol02c, For03, Goo06a, Gou03c, Ham08b, Hei07b, Hen04b, Hes03, Hod06b, Hug00, Jor01, Kas02, Kra07, Mac00b, Mag06b, Maz07a, Mor01c, Mos08a, Pan06, Ree07, Rei07b, Ric09a, Ril01a, Rit02, Roo08, Sec08, Seg05, Shr09, Smi08a, Smi09b, Tos09, Tyb02, Web09, Wee09a, Yam04a, Yam04b]. Vilich [Hör78, Hör78]. Villeneuve [Chi01b]. Vincent [Chi01a, Chi03b, Par01a, Vid05, Wee09a, Wer09]. Vintage [Bow04]. Virgin [Win09b]. Virtue [Dew01, Sha07, Jon06b, Mil00]. Visible [Mai09, Sco04]. Vision [Can09, Gou09b, Joh06a, Cla07b, Ott08, Sch06a, Wad03]. Visions [Aga08a, Joh03a, Shr09, Siv08, Web08, DM05c, Joh06b, Yeo01a]. Visual [Daw07a, Joh02b, Joh03b, Jor01, Jor08, And03b, Coo05b, Hen02b, She00, SL06, Smi06a]. visualization [Ver05]. Visualizing [GRT06, Kem09]. Vita [Høy05]. Vital [Cla07a, Far06, Ren07, Sch06b, Tay04]. Vitalism [Gar03a]. Vitalizing [Rei05b, Ste07]. Vitezslav [Har03c]. Vivien [Tim03a, Tim03d]. Vivisection [Bro01b, Fre99]. Vladimir [Nay09, Cla01]. Vocation [Bud09, BC01, Sha08b]. void [Man03]. Voigt [End04b]. Vol [Ber06b, Ber06a, Hul07, Joh07a, Kle09, Kra07, Mac06b, Maz07a, Nea00b, Rob08, Siv06, Dea09]. Volker [Smi06c]. Vols [Noa09, Abr08, Dea02, Dea09, Maz07a, Mor06a, Rad03b, Raj09, Sha09]. Voltaire [Spa02, Ter08]. Volume [Aga04, Eag05b, End04b, Gli06, Har01c, Hod01, Hod06a, Hug07b, Hun05b, Jam01b, Jam05, Mil04, Mor01c, Mor09b, Rav01, Sta06c, Ste04a, Wag05, Wee08, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Hun05b]. Volumen [Med00]. volumes [Bow06a, Mor01c]. Vorkin [Aga02]. Voulgaris [Pat07]. Voyage [Bre06, RH04]. Voyages [Rud04b, Spa05, Ste05a, Ale02, Wil03]. vselennaia [Sir09]. Vuibert [Cro05b]. VWB [Har02d, Wei02b]. vxiii [Jon02a].

W [Aga00, Dew07, Eag08, Edd03a, End04b, Gol09, Hop04, Jam01b, Jam05, Liv05a, Mor02c, Mor09b, Pan08, Pau09, Ric04, Smi08a, Wer04]. W. [Biv08, Das07, Mel01, Ril02]. Waals [Yam00]. Wade [Joh06a]. Wagenbach [Web08]. Wagner [Jor02]. während [Wei02b, Zim00b]. Wajcman [Ric08]. Wake [Sle03, Whi01]. Wakefield [Can01b]. Wakely [Man06]. Wales [Dol02a, ABB01, Som04]. Walker [Edd02, Edd04b]. Wall [Del08]. Wallace [Gar03b, Han07b, Web06, Fic04, Jon02b, Slo04, WC02, WB02]. Waller [Dol04b]. Wallflower [Sle09b]. Wallingford [Eag07, Nor05]. Wallstein [Nes06c]. Walsh [Tim03a, Tim03d]. Walter [Deb84, Wil06]. Walters [Mor03a]. Waltham [Noa09]. Wandel [Nes06a, Kre03]. Wandering [Aga01c, Moo99]. Wanderings [Mor07b]. War [Aga03a, Coo02b, Edg04, HH00, Jon02a, Kle09, Lam02, Mac07, Mur00, Rus01, Tim02, Tim03c, Tso07, Cue07, Hor00, May05a, Nye04, Sch05b, Tho07, BG02, Har03b, Kre02, Rei05c, Sel08]. Warfare [Edg06, Mac09b]. Warming [Hal06, Wea03]. warp [BDLM07]. Wars [Aga03b, Bax09, Ful01c, Ful03, Pen04, Web09, Bro01a, Sha08a, TF02, SS01]. Warwick [Gra05, Hun05a]. Was [Bou02, Aga01b, GC00, Gui04]. Washington [Aga04, Hes04, Jam01a, Jam01b, Siv08, Smi08f, Sta03]. Waste [Tur09a, Ham08a]. Watanabe [tF00]. Water [Bro05b, Cla04b, CKM99, Gra07a]. Waterloo [Art04]. Watt [Bro05b, Mil04, Hil05, MT07]. Waves [Aga06, Col04]. Wax [Alb03a, Hop02]. Way [Gou03c, Sen06a, Smi06e, Bru08, Lee02]. Ways [Chr05, Ful01b, Pic01, Fri00]. Weapons [Mac08c]. Wear [Kas02]. Weart [Hal06]. Weather [Cla01, Edd07b, Nay09, FJC06, Gol01, Gol07, Jan00, Nay06]. Web [Aga01b, Liv05a, Bro02a, GC00]. Webb [Hur07]. Webster [Pic07, WPM05]. weft [BDLM07]. Weights [CSML04, Edd06]. Weinberg [Fre04d]. Weiner [Hal04a]. Weintraub [GG03a]. Weisen [Sch00, Smi02b]. Weisse [Tim04]. Well [Tur09a, Ham08a]. Wellcome [Ade04, Bro02e, Cha05, Edd07a, Jen01, Jon09, Kno01, Mai09, Qui02, AO03]. Wells [Joh06a, Sle09a, McL09, Wad03]. Weltanschauung [Zig04, Zic00]. Wenceslaus [Hig04c]. Went [Can02b]. were [Ben03]. Werner [Kra07, Smi02b]. Wesley [Ber06d]. West [Eds04, tF00, Jac09, Jon08, Loz00, Wat97]. Western [Bek06, Biv08, Chu02, D’A03, GG02, Pet07a, Plo05, çIh04, GGC^{+}05, SS00, Sel00, Wri00]. Westport [Can08b]. Westview [Chi04a, Shi03b]. Wetenschappen [Gin04, Iro04, Mac09a]. Wetzell [Smi08f]. Weyl [Set04, WS^{+}01b]. Whale [Joh09, Bur07b]. Wheeler [Sch02a]. Wheelis [Mac08c]. Where [Dol02b, She01]. Whewell [Can02b, Mar04]. Whipple [Alb03a, Hig06, Win08]. Whitaker [Lea05]. White [Dix06, Poo09, Dub06]. Whitla [Liv03]. Whitman [Bra00]. Whitmore [Hag04]. Whitney [Sch02a]. Whitworth [Sle03]. Who [Bro01a, Ful03, Hig08, Hun06, Mah03, Tur02a]. Wide [GC00]. wider [Bek01]. Widmalm [Hes01, Ras06]. Wiesner [Wil06, WR04]. Wight [Raj09, Nol07]. Wilding [Mor07a]. Wiley [Hun06]. Wilfrid [Art04]. Wilhelm [Hof07]. William [Bas04b, Can04a, Coo07a, Coo06b, Edd05b, GG03b, Hen06a, Hur07, Joh06a, Kir05, Lea04, Liv03, Mir07, Nes09, Smi02a, Val05, Van04a, Wee09a, Woo06, Bor07, Bro08a, Can02b, Chi02c, Cla07a, Edd05b, Far09a, Fyf03b, Han07b, Hay02a, Hei08, Jor01, Kni08, Maa05, Mar04, Nes09, Pag03, Sch08a, Som04, Sum05, Tay06, Wad03]. Williams [Dol04c, Spa05]. Williamson [Ghi01, Mor02a]. Willis [Mar04]. Willmoth [Nea00b, Nea00a, Win08]. Wilson [Jen01, Phi02, Sle01, Cec01, Far00c]. Wines [Smi03a, Ewi01]. Winkelmann [Kus03]. Winner [Hes04, HD02]. Wireless [Hem03, Hon01]. Wirtschaft [Har01b, Har01c, Tei00]. Wise [Ste09a]. Wissenschaft [Har01b, Har01c, Hoc03, Nes01, Nes06c, Tei00, Zig04, Gos00]. Wissenschaftlicher [Sta04, BSD08, Bru09, SEM04]. Wissenschaftskultur [Zic00, Zig04]. Wissenschaftspopularisierung [Hop04]. Wissenskultur [Nes06a]. Wissenspopularisierung [Nes06a, Kre03]. Wissensverbreitung [Kre03, Nes06a]. Witelo [Smi02a, KD00]. Withers [Edd03a, Gol09, Smi08a]. without [CLW07, Dep07, Fal05, Ste09a]. Witness [Mor05b, Cav08]. Wittmann [Joh02b]. Wolfgang [Edd08, Hen03a, Fie05]. Woman [Wal06b, Ric03]. Women [Bro05a, DiM07, Gou03b, Hob06, HB02, Jon02a, Log08, Mur00, Gat02, Hai01, Hou06, Ogi00, Zin05, Hor00]. Wonder [Eag08, EM06]. Wonders [Ber08, DP98, Fin01, Wer01, Del06b]. Wood [Har03c]. Woodbridge [Can04a, Elm02, Hul07, Kir05, Maz07a, Nea00a]. Woods [Fin03a, SFS^{+}00]. Woodwardian [O’C05]. Wootton [Hal02a]. Worboys [Kip04, Rit09]. word [Che06]. Words [Dan01, Smi08c, Jac00b, KM06, Tho09]. Work [Fri04, Hug00, Hur07, Kni04, Nea00a, Set04, Cha02a, Fin04, Fre96, GS02, WS^{+}01b, Wil97]. Works [Ber06a, Dea02, Dol04a, Eag05b, Jor02, Pag01, oAEN04, Ber06b, HD00, NFJ97, Tra01]. World [Aga03a, Aga03b, Biv06, Bro06a, Brö02b, Cro05a, Del06a, Edg04, Fin02, Gas09, Gau08, Goo04a, Gra08b, Har03b, Hei07a, Hob06, Hut07, Joh06c, Jon02a, Lam02, Lin05, Mor07c, Mor09b, Mos08a, Nes02, Noa09, Nut04, Pad01, Pet09, Ril01a, Rob06, Rob08, Rob09b, Sch02b, Sel08, Siv06, Smi05a, Tur06a, Wil09c, Yeo04, Bek01, Bui07, Cad00, Dea06b, DD08, Fer03, For08, Fre01, Fre04a, Gol01, Gra07b, Har04b, Hug04b, MS05a, Maz07b, MD99, McL07, Par06, Sch04a, SS05, Smi01a, Zic00, çG00b, tH02, Dra00, GC00, HH00, Mur00, Rus01, SS01]. World-View [Nes02]. Worlds [Can02b, Coh09, Hsu06, Rud08, Llo04, WR01]. Worldwide [Har08]. Wormbs [Ras06]. worth [Bro06b]. wounds [Sel08]. Wright [Chu02, Tho03a]. wristwatch [SD00]. Writing [Big06, Bro05a, Car06, DS02, Ric04, Rit02, Sch03a, Ste05a, dC08, Bow06b, DS06, Gat02, KJ01, Maz00]. Writings [Cop04, Gar03b, Hen02a, Pet07a, BR00, GGC^{+}05, Liv03, WC02]. Wundarztes [Smi04]. Wurtz [Car01a, Cro03, Roc01]. Wussing [Sta03]. Wyhe [Smi06d]. Wyndham [Wal06c].

X [Aga04, Alb08, Ben02, Bow06a, Bro05b, Coo03, Fin02, Fis04, Fre04d, Fyf03a, Goo03, Goo02d, Goo09b, Gor02, Gra06, Gui05, Haf04, Har03a, Hem02a, Hen04d, Hor03, Hun06, Jam01b, Jon06a, Kha06a, Kni04, Liv05a, Man00a, Mor01c, Mun08, Pen04, Rad03b, Rob08, Sch08a, Siv08, Smi04, Smi08c, Sta07b, Win08, Yos00, Aga00, Ber06a, Boh01, Bre05, Bro09, Bro04c, Cam08, Can09, Chi06, Cla04a, Cla09, Cro04, Del09, DM05b, Dix08a, Eag06a, Edd07b, Edd08, Ful02, Ful09a, Ghi01, Gin04, Gra08a, GG03b, Gre02, Hen06b, Høy05, Joh04c, Kle05, Kni02, Luc05, Lyn09, Mac07, Maz02a, McO09, Mor07c, Mor02c, Pad01, Pan08, Rad02, Ric09b, Rit09, Rud04b, Sle01, Sle09b, Smi01b, Smi06g, Smi07a, Smi08e, Sta03, Sta09b, Ste07, Tim02, Tso05, Tur09a, Ver04]. x [Wal03, Web06, Web07, Wil09c]. xcvi [Hen02a]. xcviii [Coo07a]. XI [Man06, Ack08, Ber08, Bla03, Car06, Chi02b, Chi03c, Col02, Cro05a, Dod06, Dol04a, Dol04b, tF00, Fre04d, Ful03, Gau09, Gis09, Gou09a, GG06b, Hal06, Ham07a, Har01a, Har06, Hei08, Hes04, Hun05a, Joh06a, Kre06, Kus00, Leh08, Loc02, Lüt04a, Mac08c, Mel01, Mor08, Mun09, Phi02, Rad03b, Rei07a, Rob09c, Rut08, She07, Sle03, Smi07a, Smi07b, Smo08, Spa06, Sta03, Ter08, Til04, Top08, Van02b, Wer09, Whi04, Yos00]. Xiang [Sta08b, Joh02d]. xii [Aga01b, Alb08, Art04, Big06, Bur07a, Chr05, Coo06b, Cuo02, Dea07, Dea08, Dew07, Dig08, Dix00, Dix05, Dol02a, EB07, Fis04, Fox07, Fra05, Fyf02b, Goo02b, Goo02c, Gou09b, Haf08, Hal04a, Han07a, Har00a, Har00b, Hei07a, Hen01, Hen04c, Hen06a, Hig05, Hig07, Hob06, Hop04, Hug07b, Hut09, Jac09, Jam01a, Joh04b, Joh06c, Kai05a, Lip08, Log08, Lüt04b, Mac01a, Maz02b, Mer08, Mor02a, MW09b, Nes02, Pau09, Qui02, Ric01a, Ril00, Rob06, Rol06, Sha02, Sir06, Siv08, Smi06c, Sta03, Sta07a, Sta06b, Sta07c, Ste06, Suz05, Tho03a, Tho03b, Tim05, Tur08, Twe09, WM08, Yam04b, Yeo04]. XIII [Hen02a, Aga03b, Ben00, Ben04b, Biv06, Biv08, Bow07, Bro05b, Can04a, Car01a, Coo07a, Del08, Del09, Dol02b, ETV04, Edd09, Elm09, Ful01a, Ful03, Gol09, Goo02a, Gor02, GG03a, Gra08b, Hay02b, Hei08, Joh07b, Joh03b, Joh05c, Jor02, Kra02, MW07, Out04, Pal06, Rad00a, Ris01, Rup08, Shr08, Smi08f, Smi09a, Sum05, Sut03, Tim03a, Tim03b, Tim03d, Tyb04, Wee07, Wor07]. Ximo [Pye09]. xiv [Aga01a, And09, Bar07, Bek06, Cha03, Cla01, Cla06, Coo03, Das09, Dix06, Edd04a, End04a, Fre04c, Goo04a, Goo08a, Gou06, Gra05, Hal04b, Han07a, Han07b, Har05a, Hay06, Hig01, Hug07a, Joh09, Jon04, Jon06a, Kni04, Kno08, Kus03, Lea05, Led08, Mac08a, MW09a, Mül01, Mur02, Nea00a, Pen04, Por01, Ren09, Ric08, Ril02, Rob09b, Siv08, Ske04, Smi03b, Smi08b, Smi08d, Spa03, Sta07a, Tso07, Tur06b, Tur09b, Vic07, Wil09a, Wil09d]. xix [Aga02, Coo02b, Dea05, Eag07, Gar03b, Gou03b, GG02, Hen04c, Hur07, Jen02, Kni05, Man04, Out01, Rud07, Smi05a, Top02, Urr07]. xl [Hod01, Hod06a]. xliii [Bru06, Maz07a, Mil02, Sta03]. XLIX [Pan06, Nea00b]. xlvii [Jam01b]. xlviii [Abr08, Aga04, Jam01b]. xv [Alb03b, Bar01b, Bra07, Col01b, Coo05a, Dix09, Dol04c, Doo04, Eag08, Edd05b, Ful01b, Gil09, Hem02a, Hem03, Hen07a, Hen03b, Hut04, Jen01, Lea04, Mac00a, Mac09b, Maz07a, Mcc05, Nye09, Pic07, Qui04, Ren06, Sch03a, Sch01, Smi02a, Sta09a, Stu09, Web09, dC08]. xvi [And03a, Bar08, Bax09, Bro02e, Daw06, Dea06a, Dew01, Dix03a, Dol01, Dun08, Edd06, Gau08, Gou03a, Gui04, Hay03, Hem02b, Hig03a, Iro04, Jor02, Kip04, Lau06, Mor07b, Mor01a, Mor01b, Nut04, Olb02, Pre03, Rob05, Smi06f, Smi08c, Smi09a, Tim03c, Whi06]. xvii [Aga04, Bud09, Cla07a, Das07, Dep07, Duf09, Edd03a, Edg04, Far07a, Fin02, Gar03b, Gat05, Goo01, Gra06, Hag04, Har03c, Har08, Kni03, Lam02, Leh09, Mar05, Maz07a, Mos08b, Pet07a, Qui08, Ras06, Ree08, Roo08, Sha07, Smi00a, Smi06d, Smi08f, Sta08b, Ste05a, Ste09b, Tim04]. xviii [Big01, Dig08, Dig09, Far07b, Fra09, Hor03, Kim05, Kir05, Kni06a, Mor05b, Mor06c, Pag01, Pal00, Rud04b, Smi06e, Tra07, Van05, Wag05, Web09, Wee09b]. xx [Can08b, Chi04a, Cor00, Dod02, Emm01, For04, Fri03a, Fri03b, Goo03, Hug01, Hug09, Hun06, Kem09, Kha06b, Kra08a, Liv05a, Mac08b, Mor03a, Mor06c, Nay09, Pet09, Plo05, Sam07, Sch08a, Smi03a, Sta07b, Ste04b, Tur02a, Win08]. xxi [Aga08a, Alb01, Big07, Doo05, Ful09a, Fyf02b, Goo06b, Goo06c, Joh03a, Kus07, Lev07, Mac06b, Nes09, Ram02, Sch02b]. xxii [Chi08, Coh09, Mir07, Mor06b, Mor03b, Rud04b, Sho03, Smi02b]. xxiii [Abr08, Aga06, D’A03, Daw07a, Gui05, Hun02, Joh02d]. xxiiii [Stu06]. xxiv [Eag05b, Eag06b, Edd05b, Ful09b, Har03a, Mor06c, Mor06e, Rob03, Shi03b, Web07, Wil06]. xxix [Bre06, Cha01, Dea09, Fyf03a, Kha06b, Sta06c, Tho02b, War08]. xxv [Joh02c, Kle04, Ren00, Smi05b, Tim08]. xxvi [Bro05a, But06, Chu02, Sum06]. xxvii [Aug09, Dou06, Hun05b, Jon09, Mac09a, Ric04, Wee08]. xxviii [Aga04, Fyf05, Hen04a, Hun05b, Rob08, Sch02a]. xxx [Hun05b, Jor08, Sen06a, Tot04]. xxxi [Fox04, Gol05]. XXXII [Man01b, Tho03c]. xxxiii [Mor09b, Smi00a]. xxxiv [Lin05, Ner05, Ric01b, Van04a]. xxxv [Man01a, Siv06, Tur06a]. xxxvii [Nea00b]. xxxviii [Bra07].

Yakov [Hug00, Fre96]. Yakovlevich [Yam00]. Yale [Aga01a, Dew01, Dix00, For03, Har06, Hut09, Kno08, Mor07c, Sch02a, Sch02b, Stu09, Web09, Wee09b, HW01a]. Yandell [GG04]. Yavelov [Yam00]. Year [Gil09, Goo04a, Ano00e, Far09b]. Yearbook [Dea09, Tso05]. Years [Cor06, Ful01a, Jam01b, Jam05, Sch02a, Sta03, Gil04, HW01a, Hac00, Hör78, HS00]. Yellow [Fan02, Dod01]. Yeo [Bre04, Joh03a]. York [Ade04, Aga00, App05b, Ben04b, Big07, Bow06a, But06, Can07, Coo06a, Das09, Dol02c, Duf09, Eag06a, Eag07, Fin01, Fis04, For04, Ful02, Ful09a, Fyf02a, Fyf05, Gar03b, Goo06c, Goo09b, Gor09, Gou03c, Gre02, Ham07b, Har00b, Har03a, Har03c, Har04c, Jac09, Jam03, Joh09, Joh06a, Jor02, Jor06, Kha06b, Led08, Mac06b, Maz07a, Mcc05, Mor08, Mun07, Mur02, Mus03, Pal06, Poo09, Ree07, Rei07a, Rei07b, Ric09a, Ris01, Rit09, Rob09b, Sec08, Shr09, Smi03b, Smi05a, Smi08e, Sta03, Sta09a, Sta07b, Ste06, Til04, Tur06a, Ver04, Wal06b, Web06, Wei02a, dC08, Bur07b]. Young [Dol01, Ful00b]. Yours [End04b].

Z [Dol01, Hun05a]. Zacut [Goo02c, CG00a]. Zambesi [Dri05]. Zealand [Bar01b, Mor02a, Dun99, NS00, Tho09]. Zeilinger [Coo06c]. Zeit [Set04, Wei02b, WS^{+}01b, Zim00b]. Zeiten [Man01b]. Zenderland [Col00]. Ziauddin [Ful01c]. Ziche [Zig04]. Ziegler [Alb03a, Hop02]. Zigman [Man01b]. Zilsel [Hen03a]. Ziman [Goo01]. Zimmermann [Wei02b]. Zinsser [DiM07, Ter08]. Zionist [Web09, Suf07]. Zittel [Dea09]. Zoé [Rud04b]. Zone [Fin01, Jor06]. Zoological [Whi05]. Zoology [Ric01b, Bla07, Key00]. zrenii [Sir09]. Zuckerman [Hor03, Pey01]. zum [HS98, Kra02, Smi06c]. Zur [Coe09, App05b, Har01b, Har02d, Hör78, Tei00, Wei02b, Zig04]. Zurich [NG04]. Zwischen [Har01c].

References

Allen:2001:NSC

[AA01] David Elliston Allen and David Elliston Allen, editors. Naturalists and society: the culture of natural history in Britain, 1700–1900, volume CS724 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, England, 2001. ISBN 0-86078-863-6. xiv + 298 pp. LCCN QH137 .N28

                                                   2001. URL http://www.loc.
                                                   gov/catdir/toc/fy02/
                                                   2001022813.html. Reprints of

articles originally published in various journals from 1856–2000.

Aaboe:2001:EEH

[Aab01] Asger Aaboe. Episodes from the early history of astronomy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95136-9 (softcover). xv + 172 pp. LCCN QB15 .A15

                                                   2001. URL http://www.loc.
                                                   gov/catdir/enhancements/
                                                   fy0816/00061919-d.html;
                                                   http://www.loc.gov/catdir/
                                                   toc/fy02/00061919.html.

Agar:2000:BSI

[AAH00a] Jon Agar, William J. Ashworth, and Jeff Hughes. BJHS special issue: On time: history, science and commemoration. British Journal for the History of Science, 33(4):385, December 2000.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028026.

Ashworth:2000:THS

[AAH00b] William Ashworth, Jon Agar, and Jeff Hughes. On time: history, science and commemoration. special issue. British Journal for the History of Science, 33(4):385–514, 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Angerstein:2001:RRA

[ABB01] R. R. (Reinhold Rücker) Angerstein, Torsten Berg, and Peter Bergf. R. R. Angerstein’s illustrated travel diary, 1753– 1755: industry in England and Wales from a Swedish perspective. Science Museum, London, UK, 2001. ISBN 1-900747-24-3. xxii + 378 pp. LCCN T26.E7 .A54 2001.

Andersen:2006:CSS

[ABC06] Hanne Andersen, Peter Barker, and Xiang Chen. The cognitive structure of scientific revolutions. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85575-6. xvii + 199 pp. LCCN Q174.8 .A53

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0633/2005021713-d.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2005021713-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0516/2005021713.html.

Abraham:2008:BRM

[Abr08] Tara H. Abraham. Book review: Margaret A. Boden, Mind as Machine: A History of Cognitive Science. 2 vols. Oxford: Oxford University Press, 2006. Pp. xlviii+xxiii + 1631. ISBN 0-19-924144-9. £125.00 (hardback). British Journal for the History of Science, 41(4):623–624, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agar:2009:E

[ACC09] Jon Agar, Joe Cain, and Hasok Chang. Editorial. British Journal for the History of Science, 42(3):319–320, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ackermann:2008:BRG

[Ack08] Silke Ackermann. Book review: George Saliba, Islamic Science and the Making of the European Renaissance. Cambridge, MA and London: MIT Press, 2007. Pp. xi + 315. ISBN 978-0-262-19557-7. £24.95 (hardback). British Journal for the History of Science, 41(4):602–603, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Adelman:2004:BRG

[Ade04] Juliana Adelman. Book review: Greta Jones, ‘Captain of All these Men of Death’: The History of Tuberculosis in Nineteenth and Twentieth Century Ireland. Wellcome Institute Series in the History of Medicine: Clio Medica, 62. Amsterdam and New York: Rodopi, 2001. Pp. ii + 263. ISBN 90-420-1031-2. $22.50, EUR24.00 (paperback). British Journal for the History of Science, 37(4): 481–482, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028658.

Adelman:2005:EDP

[Ade05] Juliana Adelman. Evolution on display: promoting Irish natural history and Darwinism at the Dublin Science and Art Museum. British Journal for the History of Science, 38(4):411–436, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Adelman:2006:BRK

[Ade06] Juliana Adelman. Book review:

Kevin Bright, The Royal Dublin Society, 1815–1845. Dublin: Four Courts Press, 2004. Pp. 288. ISBN 1-85182-813-3. EUR45.00, £45.00, $45.00 (hardback). British Journal for the History of Science, 39(4):611–612, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028533.

Adelman:2008:BRM

[Ade08] Juliana Adelman. Book review: Mariano Artigas, Thomas F. Glick and Rafael A. Martínez, Negotiating Darwin: The Vatican Confronts Evolution, 1877– 1902. Baltimore: Johns Hopkins University Press, 2006. Pp. 326. ISBN 0-8018-8389-7. £33.50, $50.00 (hardback). British Journal for the History of Science, 41(3):471–472, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Aftalion:2001:HIC

[Aft01] Fred Aftalion. A history of the international chemical industry. Chemical Heritage Press, Philadelphia, PA, USA, second edition, 2001. ISBN 0-941901-29-7. xxii + 442 pp. LCCN HD9650.5 .A3813

                                        2001. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0630/2001047224-d.html.

Agar:2000:BRM

[Aga00] Jon Agar. Book review: Michael Riordan and Lillian Hoddeson, Crystal Fire: The Birth of the Information Age. New York and London: W. W. Norton & Company, 1997. Pp. x + 352. ISBN 0-393-04124-7. £27.50, $56.95. British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4027926.

Agar:2001:BRJa

[Aga01a] Jon Agar. Book review: James C. Scott, Seeing Like a State:

How Certain Schemes to Improve the Human Condition Have Failed. New Haven and London: Yale University Press, 1998. Pp. xiv + 445. ISBN 0-300-07016-0. £27.50, $37.50 (hardback); 0-300-07815-3. £11.95, $16.95 (paperback). British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028041.

Agar:2001:BRJb

[Aga01b] Jon Agar. Book review: James Gillies and Robert Cailliau, How the Web was Born. Oxford: Oxford University Press, 2000. Pp. xii + 372. ISBN 0-19-286207-3. £8.99, $15.95 (paperback). British Journal for the History of Science, 34(3):341– 373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028102.

Agar:2001:BRP

[Aga01c] Jon Agar. Book review: Patrick Moore, The Wandering Astronomer. Bristol and Philadelphia: Institute of Physics Publishing, 1999. Pp. 208. ISBN 0-7503-0693-9. £19.95. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028042.

Agar:2002:BRD

[Aga02] Jon Agar. Book review: David H. De Vorkin, Henry Norris Russell: Dean of American Astronomers. Princeton and Oxford: Princeton University Press, 2000. Pp. xix + 499. ISBN 0-691-04918-1. £30.00, $49.50 (hardback). British Journal for the History of Science, 35(4):475– 485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agar:2003:BRA

[Aga03a] Jon Agar. Book review: Agatha C. Hughes and Thomas P. Hughes (eds.), Systems, Experts, and Computers: The Systems Approach in Management and Engineering, World War II and After. Dibner Institute Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2000. Pp. ii + 513. ISBN 0-262-08285-3. £34.50 (hardback). British Journal for the History of Science, 36(4):485–487, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028224;
                                        http://www.jstor.org/
                                        stable/4028284.

Agar:2003:BRR

[Aga03b] Jon Agar. Book reviews: Reinhard Siegmund-Schultze, Rockefeller and the Internationalization of Mathematics between the Two World Wars: Documents and Studies for the Social History of Mathematics in the 20th Century. Science Networks – Historical Studies, 25. Basel, Boston and Berlin: Birkhäuser Verlag, 2001. Pp. xiii + 341. ISBN 3-7643-6468-8. US$94.95 (hardback). British Journal for the History of Science, 36 (1):87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028295.

Agar:2004:BRS

[Aga04]  Jon Agar.  Book reviews:

Stephen B. Johnson, The Secret of Apollo: Systems Management in American and European Space Programs. New Series in NASA History. Baltimore and London: Johns Hopkins University Press, 2002. Pp. xvii + 290. ISBN 0-8018-6898-X. £30.50 (hardback). John M. Logsdon (ed.), Exploring the Unknown: Selected Documents in the History of the U.S. Civil Space Program. Volume V: Exploring the Cosmos. NASA History Series. Washington: NASA, 2001. Pp. xxviii+796. ISBN 0-16-061774-X. No price given (hardback). Douglas J. Mudgway, Uplink-Downlink: A History of the Deep Space Network 1957–1997. NASA History Series. Washington, DC: National Aeronautics and Space Administration Office of External Relations, 2001. Pp. xlviii+674. ISBN 0-16-066599-X. $82.00 (US domestic postpaid), $102.50 (internationally) (hardback). British Journal for the History of Science, 37(2):231–233, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028359.

Agar:2005:BRT

[Aga05] Jon Agar. Book review: Thomas J. Misa, Philip Brey and Andrew Feenberg (eds.), Modernity and Technology. Cambridge, MA and London: MIT Press, 2003. Pp. ix + 421. ISBN 0-262-13421-7. £26.50 (hardback). British Journal for the History of Science, 38(4):471–473, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028458.

Agar:2006:BRH

[Aga06] Jon Agar. Book review: Harry Collins, Gravity’s Shadow: The Search for Gravitational Waves. Chicago and London: University of Chicago Press, 2004. Pp. xxiii + 870. ISBN 0-226-11378-7. $39.00 (paperback). British Journal for the History of Science, 39(4):624–625, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028542.

Agar:2008:BRS

[Aga08a] Jon Agar. Book review: Sean F. Johnston, Holographic Visions: A History of New Science. Oxford: Oxford University Press, 2006. Pp. xxi + 518. ISBN 0-19-857122-4. £75.00 (hardback). British Journal for the History of Science, 41(1): 126–127, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Agar:2008:WHS

[Aga08b] Jon Agar. What happened in the sixties? British Journal for the History of Science, 41(4): 567–600, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ackerberg-Hastings:2002:ASJ

[AH02] Amy Ackerberg-Hastings. Analysis and synthesis in John Playfair’s elements of geometry. British Journal for the History of Science, 35(1):43–72, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028402.

Archer:2005:CCC

[AH05] Mary D. Archer and Christopher D. Haley, editors. The 1702 Chair of Chemistry at Cambridge: transformation and change. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-82873-2 (hardcover). xxi + 318 pp. LCCN QD18.G7

C48 2002. URL http://www.
loc.gov/catdir/description/
cam041/2004051857.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2004051857-b.html;

                                        http://www.loc.gov/catdir/
                                        toc/cam051/2004051857.html.

Ahmad:1984:IQS

[Ahm84] Syed Barakat Ahmad. Introduction to Quranic script. Macmillan India, New Delhi, India, 1984. ISBN 0-333-90735-3. xi + 168 pp. LCCN PJ6696.Z5 A4 1984. Rs150.00.

Alberti:2001:BRS

[Alb01] Samuel J. M. M. Alberti. Book review: Simon J. Knell, The Culture of English Geology, 1815–1851: A Science Revealed Through its Collecting. Aldershot: Ashgate, 2000. Pp. xxi + 377. ISBN 1-84014-625-7. £59.50 (hardback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028374.

Alberti:2002:PNN

[Alb02] Samuel J. M. M. Alberti. Placing nature: natural history collections and their owners in nineteenth-century provincial England. British Journal for the History of Science, 35(3): 291–311, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028125.

Alberti:2003:BRN

[Alb03a] Samuel J. M. M. Alberti. Book review: Nick Hopwood, Embryos in Wax: Models from the Ziegler Studio. With a Reprint of Embryological Wax Models by Friedrich Ziegler. Cambridge: Whipple Museum of the History of Science and Bern: Institute of Medical History, 2002. Pp. ix + 206. ISBN 0-906271-18-5. £13.50 (paperback). British Journal for the History of Science, 36(3): 372–373, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028165;
http://www.jstor.org/
stable/4028450.

Alberti:2003:BRS

[Alb03b] Samuel J. M. M. Alberti. Book review: Stephen T. Asma, Stuffed Animals & Pickled Heads: The Culture and Evolution of Natural History Museums. Oxford: Oxford University Press, 2001. Pp. xv + 302. ISBN 0-19-513050-2. £22.99, $30.00 (hardback). British Journal for the History of Science, 36(2):236–237, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028238.

Alberti:2008:BRK

[Alb08] Samuel J. M. M. Alberti. Book review: Ken Arnold, Cabinets for the Curious: Looking back at Early English Museums. Perspectives on Collecting. Aldershot: Ashgate, 2006. Pp. xii + 297. ISBN 0-7546-0506-X. £47.50 (hardback). British Journal for the History of Science, 41 (1):139–140, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Alexander:2002:GLV

[Ale02] Amir R. Alexander. Geometrical landscapes: the voyages of discovery and the transformation of mathematical practice, volume 1 of Writing science. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-3260-4 (cloth). xv + 293 pp. LCCN QA24 .A54

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0710/2002009239-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0710/
                                        2002009239-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy032/2002009239.html.

Alter:1999:DLI

[Alt99] Stephen G. Alter. Darwinism and the linguistic image: language, race, and natural theology in the Nineteenth Century. New studies in American intellectual and cultural history. The Johns Hopkins University Press, Baltimore, MD, USA, 1999. ISBN 0-8018-5882-8. xiii + 193 pp. LCCN P75 .A47

                                        1999. URL http://www.h-net.
                                        org/review/hrev-a0b8y2-aa;
                                         http://www.loc.gov/catdir/
                                        bios/jhu051/98014270.html;
                                         http://www.loc.gov/catdir/
                                        description/jhu052/
                                        98014270.html.

Allen:2001:SHS

[AM01] Garland E. Allen and Roy M.

MacLeod, editors. Science, history, and social activism: a tribute to Everett Mendelsohn, volume 228 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0495-8 (hardcover), 94-017-2956-5. ISSN 0068-0346. x + 388 pp. LCCN Q174 .B67 vol. 228; Q175.55. URL http://

www.loc.gov/catdir/
enhancements/fy0821/
2002023597-d.html;
http://www.loc.gov/catdir/
toc/fy022/2002023597.html;
 https://link.springer.com/
book/10.1007/978-94-017-
2956-7.

Amundson:2005:CRE

[Amu05] Ronald Amundson. The changing role of the embryo in evolutionary thought: roots of evodevo. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-80699-2 (hardcover). xiii + 280 pp. LCCN QH360.5 .A48

2005. URL http://www.loc.
gov/catdir/description/
cam051/2004054409.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2004054409-b.html;
http://www.loc.gov/catdir/
toc/cam051/2004054409.html.

Anderson:2003:BRK

[And03a] Katharine Anderson. Book review: Kathryn A. Neeley, Mary Somerville: Science, Illumination, and the Female Mind. Cambridge Science Biographies. Cambridge: Cambridge University Press, 2001. Pp. xvi + 263. ISBN 0-521-62672-2. £14.95, $23.00 (paperback). British Journal for the History of Science, 36(2):237–238, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028239.

Anderson:2003:LSV

[And03b] Katharine Anderson. Looking at the sky: the visual context of Victorian meteorology. British Journal for the History of Science, 36(3):301–332, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028157.

Anderson:2009:BRM

[And09] Katharine Anderson. Book review: Michael S. Reidy, Tides of History: Ocean Science and Her Majesty’s Navy. Chicago and London: University of Chicago Press, 2008. Pp. xiv + 389. ISBN 978-0-226-70932-1. £55.00 (hardback). British Journal for the History of Science, 42(3): 464–465, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anker:2001:IEE

[Ank01] Peder Anker. Imperial ecology: environmental order in the British Empire, 1895–1945. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00595-3 (cloth). vii + 343 pp. LCCN GF551 .A65 2001.

Anonymous:2000:BSI

[Ano00a] Anonymous. BJHS September 1999 issue. British Journal for the History of Science, 33(1):127, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028077.

Anonymous:2000:BRa

[Ano00b] Anonymous. Books received.

British Journal for the History of Science, 33(1):125, March 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2000:BRb

[Ano00c] Anonymous. Books received.

British Journal for the History of Science, 33(2):255–256, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027942.

Anonymous:2000:BRc

[Ano00d] Anonymous. Books received.

British Journal for the History of Science, 33(3):380–382, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027965.

Anonymous:2000:BSH

[Ano00e] Anonymous. British Society for the History of Science Report of Council for the year 1999. British Journal for the History of Science, 33(4):501–514, December 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028034.

Anonymous:2000:EN

[Ano00f] Anonymous. Editorial note. British Journal for the History of Science, 33(1):1–2, March 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2001:BRa

[Ano01a] Anonymous. Books received. British Journal for the History of Science, 34(1):125–128, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028060.

Anonymous:2001:BRb

[Ano01b] Anonymous. Books received. British Journal for the History of Science, 34(2):251–253, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028023.

Anonymous:2001:BRc

[Ano01c] Anonymous. Books received. British Journal for the History of Science, 34(3):375–378, September 2001. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028121.

Anonymous:2001:BRd

[Ano01d] Anonymous. Books received.

British Journal for the History of Science, 34(4):483–485, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028398.

Anonymous:2001:BRe

[Ano01e] Anonymous. Books reviewed.

British Journal for the History of Science, 34(4):486–488, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2002:BRa

[Ano02a] Anonymous. Books received.

British Journal for the History of Science, 35(1):125–127, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2002:BRb

[Ano02b] Anonymous. Books received.

British Journal for the History of Science, 35(2):251–253, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028208.

Anonymous:2002:BRc

[Ano02c] Anonymous. Books received.

British Journal for the History of Science, 35(3):380–382, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028154.

Anonymous:2003:BRa

[Ano03a] Anonymous. Books received. British Journal for the History of Science, 36(2):255–259, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028252.

Anonymous:2003:BRb

[Ano03b] Anonymous. Books received. British Journal for the History of Science, 36(3):385–387, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028176.

Anonymous:2003:BRc

[Ano03c] Anonymous. Books received. British Journal for the History of Science, 36(4):489–491, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2004:BRa

[Ano04a] Anonymous. Books received. British Journal for the History of Science, 37(1):115–118, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028271.

Anonymous:2004:BRb

[Ano04b] Anonymous. Books received.

British Journal for the History of Science, 37(2):241–243, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2004:BRc

[Ano04c] Anonymous. Books received.

British Journal for the History of Science, 37(3):369–371, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028147.

Anonymous:2004:BRd

[Ano04d] Anonymous. Books received.

British Journal for the History of Science, 37(4):493–494, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2004:BRe

[Ano04e] Anonymous. Books reviewed.

British Journal for the History of Science, 37(4):496–499, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2005:BRa

[Ano05a] Anonymous. Books received.

British Journal for the History of Science, 38(1):123–125, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2005:BRb

[Ano05b] Anonymous. Books received. British Journal for the History of Science, 38(2):245–249, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2005:BRc

[Ano05c] Anonymous. Books received. British Journal for the History of Science, 38(3):374–378, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2005:BRd

[Ano05d] Anonymous. Books received. British Journal for the History of Science, 38(4):497–498, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2006:BRa

[Ano06a] Anonymous. Books received. British Journal for the History of Science, 39(1):155–157, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2006:BRb

[Ano06b] Anonymous. Books received. British Journal for the History of Science, 39(2):315–318, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2006:BRc

[Ano06c] Anonymous. Books received.

British Journal for the History of Science, 39(3):473–475, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2006:BRd

[Ano06d] Anonymous. Books received.

British Journal for the History of Science, 39(4):626–628, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2007:BRa

[Ano07a] Anonymous. Books received.

British Journal for the History of Science, 40(1):156–158, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500712.

Anonymous:2007:BRb

[Ano07b] Anonymous. Books received.

British Journal for the History of Science, 40(2):311–313, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500743.

Anonymous:2007:BRc

[Ano07c] Anonymous. Books received.

British Journal for the History of Science, 40(3):467–469, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2007:BRd

[Ano07d] Anonymous. Books received. British Journal for the History of Science, 40(4):623–625, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2008:BRa

[Ano08a] Anonymous. Books received. British Journal for the History of Science, 41(1):157–159, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2008:BRb

[Ano08b] Anonymous. Books received. British Journal for the History of Science, 41(2):315–317, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2008:BRc

[Ano08c] Anonymous. Books received. British Journal for the History of Science, 41(3):473–476, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2008:BRd

[Ano08d] Anonymous. Books received. British Journal for the History of Science, 41(4):631–633, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIa

[Ano09a] Anonymous. BJH volume 42 issue 1 cover and back matter. British Journal for the History of Science, 42(1):b1–b3, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIb

[Ano09b] Anonymous. BJH volume 42 issue 1 cover and front matter. British Journal for the History of Science, 42(1):f1–f2, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIc

[Ano09c] Anonymous. BJH volume 42 issue 2 cover and back matter. British Journal for the History of Science, 42(2):b1–b4, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVId

[Ano09d] Anonymous. BJH volume 42 issue 2 cover and front matter. British Journal for the History of Science, 42(2):f1–f2, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIe

[Ano09e] Anonymous. BJH volume 42 issue 3 cover and back matter. British Journal for the History of Science, 42(3):b1–b3, September 2009. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIf

[Ano09f] Anonymous. BJH volume 42 issue 3 cover and front matter. British Journal for the History of Science, 42(3):f1–f2, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIg

[Ano09g] Anonymous. BJH volume 42 issue 4 cover and back matter. British Journal for the History of Science, 42(4):b1–b4, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BVIh

[Ano09h] Anonymous. BJH volume 42 issue 4 cover and front matter. British Journal for the History of Science, 42(4):f1–f2, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BRa

[Ano09i] Anonymous. Books received. British Journal for the History of Science, 42(1):157–159, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BRb

[Ano09j] Anonymous. Books received. British Journal for the History of Science, 42(2):315–318, June 2009. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BRc

[Ano09k] Anonymous. Books received.

British Journal for the History of Science, 42(3):474–477, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BRd

[Ano09l] Anonymous. Books received.

British Journal for the History of Science, 42(4):635–637, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:BRe

[Ano09m] Anonymous. Books reviewed.

British Journal for the History of Science, 42(4):641–644, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:E

[Ano09n] Anonymous. Editorial. British Journal for the History of Science, 42(1):1–4, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Anonymous:2009:IA

[Ano09o] Anonymous. Index of authors.

British Journal for the History of Science, 42(4):639, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Arnold:2003:MMF

[AO03] Ken Arnold and Danielle Olsen, editors. Medicine man: the forgotten museum of Henry Wellcome. British Museum Press, London, UK, 2003. ISBN 0-7141-2794-9 (paperback). 397 pp. LCCN R891.L86 M436 2003. US$19.99.

Applebaum:2005:BRB

[App05a] Wilbur Applebaum. Book review: Barbara M. Benedict, Curiosity: A Cultural History of Early Modern Inquiry. Chicago and London: University of Chicago Press, 2003. Pp. ix + 321. ISBN 0-226-04264-2. £17.50, $25.00 (paperback). British Journal for the History of Science, 38(2):226–227, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028703.

Applebaum:2005:BRL

[App05b] Wilbur Applebaum. Book review: Ladina Bezzola Lambert, Imagining the Unimaginable: The Poetics of Early Modern Astronomy. Internationale Forschungen zur allgemeinen und vergleichenden Literaturwissenschaft, 58. Amsterdam and New York: Rodopi, 2002. Pp. ix + 182. ISBN 90-420-1578-0. $34.50, EUR37.00 (paperback). British Journal for the History of Science, 38(1): 105–106, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028588.

Arabatzis:2006:REB

[Ara06] Theodore Arabatzis. Representing electrons: a biographical approach to theoretical entities. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-02420-2, 0-226-02421-0 (paperback). xiv + 295 pp. LCCN QC793.5.E62 A73 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0621/
2005004218-b.html;
http://www.loc.gov/catdir/
enhancements/fy0621/
2005004218-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip058/2005004218.html.

Arnold:2006:TTG

[Arn06a]  David Arnold.  The tropics

and the traveling gaze: India, landscape, and science, 1800– 1856. Culture, place, and nature. University of Washington Press, Seattle, WA. USA, 2006. ISBN 0-295-98581-X (hardcover). xiv + 298 pp. LCCN GF661 .A76

2006. URL http://www.loc.
gov/catdir/toc/ecip0517/
2005023751.html.

Arnold:2006:CCL

[Arn06b] Ken Arnold. Cabinets for the curious: looking back at early English museums. Perspectives on collecting. Ashgate, Aldershot, Hampshire, England, 2006. ISBN 0-7546-0506-X. xii + 297 pp. LCCN AM42.E54 .A76

2006. URL http://www.loc.

                                        gov/catdir/toc/ecip0422/
                                        2004020631.html.

Artmann:2004:BRR

[Art04] Benno Artmann. Book review: Roger Herz-Fischler, The Shape of the Great Pyramid. Waterloo: Wilfrid Laurier University Press, 2000. Pp. xii + 293. ISBN 0-88920-324-5. $29.95 (paperback). British Journal for the History of Science, 37(3):355–356, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4028137;   http://www.
                                        jstor.org/stable/4028439.

Ashworth:2001:BRMa

[Ash01] William J. Ashworth. Book review: Michael J. Crowe, David R. Dyck and James R. Kevin (eds.), A Calendar of the Correspondence of Sir John Herschel. Cambridge: Cambridge University Press, 1998. Pp. 828. ISBN 0-521-63149-1. £100.00. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028043.

Ashworth:2003:CET

[Ash03] William Ashworth. Customs and excise: trade, production and consumption in England 1640–1845. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-925921-6. xii + 396 pp. LCCN HJ5707.G7 A84

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0615/2003270849-d.html;
http://www.loc.gov/catdir/
enhancements/fy0615/
2003270849-t.html.

Ash:2004:PKE

[Ash04a] Eric H. Ash. Power, knowledge, and expertise in Elizabethan England. The Johns Hopkins University Press, Baltimore, MD, USA, 2004. ISBN 0-8018-7992-2 (hardcover). viii + 265 pp. LCCN DA356 .A69

2004. URL http://www.h-net.
org/review/hrev-a0e7b8-aa;
 http://www.loc.gov/catdir/
bios/jhu051/2004003188.
html;   http://www.loc.gov/
catdir/description/jhu051/
2004003188.html;   http://
www.loc.gov/catdir/toc/
ecip0414/2004003188.html.

Ashworth:2004:BRMb

[Ash04b] William Ashworth. Book reviews: Michael Hoskin (ed.), Caroline Herschel’s Autobiographies. Cambridge: Science History Publications, 2003. Pp. viii +

147. ISBN 0-905193-06-7. £25.00 (hardback). Michael Hoskin, The Herschel Partnership as Viewed by Caroline. Cambridge: Science History Publications, 2003. Pp. viii+182. ISBN 0-905193-05-9. £25.00 (hardback). British Journal for the History of Science, 37(3):350–351, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4028131;   http://www.
                                        jstor.org/stable/4028433.

Asma:2001:SAP

[Asm01] Stephen T. Asma. Stuffed animals and pickled heads: the culture and evolution of natural history museums. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513050-2. xv + 302 pp. LCCN QH70.A1 .A75

                                        2001. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0638/00040674-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0725/
                                        00040674-b.html.

Astore:2001:OGT

[Ast01] William J. Astore. Observing God: Thomas Dick, evangelicalism, and popular science in Victorian Britain and America. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-7546-0202-8. ix + 303 pp. LCCN BL245 .A78 2001.

Atmore:2004:RIR

[Atm04] Henry Atmore. Railway interests and the ‘rope of air’, 1840–8. British Journal for the History of Science, 37(3): 245–279, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028424.

Auger:2009:BRH

[Aug09] Jean-François Auger. Book review: Hannah Gay, The History of Imperial College London, 1907–2007: Higher Education and Research in Science, Technology and Medicine. London: Imperial College Press, 2007. pp. xxvii + 825. ISBN 1-86094-709-3. £122.40, $152.00 (softcover). British Journal for the History of Science, 42(4):630–631, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bahar:2001:JML

[Bah01] Saba Bahar. Jane Marcet and the limits to public science. British Journal for the History of Science, 34(1):29–49, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028037.

Bailey:2003:RSE

[Bai03] Michael R. (Michael Reeves)

Bailey, editor. Robert Stephenson — the eminent engineer. Ashgate, Aldershot, Hampshire, England, 2003. ISBN 0-7546-3679-8. xxvii + 401 pp. LCCN TA140.S72 .R63 2003.

Baird:2004:TKP

[Bai04] Davis Baird. Thing knowledge:

a philosophy of scientific instruments. University of California Press, Berkeley, CA, USA, 2004. ISBN 0-520-23249-6 (cloth). xxi + 273 pp. LCCN Q185 .B24

2004. URL http://www.loc.
gov/catdir/bios/ucal052/
2003005683.html;
http://www.loc.gov/catdir/
description/ucal042/
2003005683.html;   http://
www.loc.gov/catdir/toc/
ecip041/2003005683.html.

Bailyn:2005:AHC

[Bai05] Bernard Bailyn. Atlantic history: concept and contours. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01688-2. 149 pp. LCCN D13.5.A75 B35 2005.

Barker:2000:SMP

[Bar00a]  Andrew Barker.  Scientific

method in Ptolemy’s Harmonics. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-55372-5 (hardcover). viii + 281 pp. LCCN MT5.5.P823 B37

2000. URL http://www.loc.
gov/catdir/description/
cam021/00028912.html;
http://www.loc.gov/catdir/
samples/cam031/00028912.
html;
http://www.loc.gov/catdir/
toc/cam023/00028912.html.

Bartky:2000:STT

[Bar00b] Ian R. Bartky. Selling the true time: Nineteenth-Century timekeeping in America. Stanford University Press, Stanford, CA, USA, 2000. ISBN 0-8047-3874-2. xvi + 310 pp. LCCN QB210.U5

B37 2000. URL http://www.
loc.gov/catdir/description/
cam0210/99086739.html;

http://www.loc.gov/catdir/
enhancements/fy0710/
99086739-b.html;
http://www.loc.gov/catdir/
toc/cam027/99086739.html.

Bartusiak:2000:EUS

[Bar00c] Marcia Bartusiak. Einstein’s unfinished symphony: listening to the sounds of space-time. Joseph Henry Press, Washington, DC, USA, 2000. ISBN 0-309-06987-4. xii + 249 pp. LCCN QC173.59.S65 B39 2000.

Barnes:2001:BRT

[Bar01a] Barry Barnes. Book review:

Thomas S. Kuhn, The Road since Structure : Philosophical Essays, 1970–1993, with an Autobiographical Interview. Edited by James Conant and John Haugeland. Chicago and London: University of Chicago Press, 2000. Pp. viii + 335. ISBN 0-226-45798-2. £16.00, $25.00 (hardback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028103.

Barrow:2001:BRT

[Bar01b] Mark V. Barrow, Jr. Book review: Thomas R. Dunlap, Nature and the English Diaspora: Environment and History in the United States, Canada, Australia, and New Zealand. Studies in Environment and History. Cambridge: Cambridge University Press, 1999. Pp. xv + 350.

ISBN 0-521-65700-8. £13.95, $19.95 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028375.

Barnes:2006:GSP

[Bar06] David S. Barnes. The great stink of Paris and the nineteenthcentury struggle against filth and germs. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8349-0 (hardcover). xi + 314 pp. LCCN RA418.3.E85 B37

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005023385-b.html;
http://www.loc.gov/catdir/
enhancements/fy0623/
2005023385-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0518/2005023385.html.

Barton:2007:BRD

[Bar07]  Ruth  Barton.

Book review: David M. Knight and Matthew D. Eddy (eds.), Science and Beliefs: From Natural Philosophy to Natural Science, 1700–1900. Science, Technology and Culture, 1700–1945. Aldershot: Ashgate, 2005. Pp. xiv + 272. ISBN 0-7546-3996-7. £47.50 (hardback). British Journal for the History of Science, 40 (1):121–123, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500690.

Barton:2008:BRB

[Bar08] Ruth Barton. Book review:

Bernard Lightman, Victorian Popularizers of Science: Designing Nature for New Audiences. Chicago and London: University of Chicago Press, 2007. Pp. xvi + 545. ISBN 978-0-226-48118-0. $37.50, £23.50 (hardback). British Journal for the History of Science, 41(4):616–617, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bashford:2004:IHC

[Bas04a] Alison Bashford. Imperial hygiene: a critical history of colonialism, nationalism and public health. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2004. ISBN 1-4039-0488-X (hardcover). xiv + 264 pp. LCCN RA553 .B37

2004. URL http://www.loc.
gov/catdir/bios/hol053/
2003048280.html;
http://www.loc.gov/catdir/
description/hol041/
2003048280.html;
http://www.loc.gov/catdir/
toc/hol041/2003048280.html.

Bastable:2004:ASS

[Bas04b] Marshall J. Bastable. Arms and the state: Sir William Armstrong and the remaking of British naval power, 1854–1914. Modern economic and social history series. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-7546-3404-3. xii + 300 pp. LCCN HD9743.G72 B37 2004.

Bauer:2001:SPM

[Bau01] Henry H. Bauer. Science or pseudoscience: magnetic healing, psychic phenomena, and other heterodoxies. University of Illinois Press, Urbana, IL, USA, 2001. ISBN 0-252-02601-2 (cloth). xi + 275 pp. LCCN Q172.5.P77 B38 2001.

Baxfield:2009:BRJ

[Bax09] Christopher Baxfield. Book review: J. B. Shank, The Newton Wars and the Beginning of the French Enlightenment. Chicago and London: University of Chicago Press, 2008. Pp. xvi + 571. ISBN 978-0-226-74945-7. $55.00 (hardback). British Journal for the History of Science, 42(2):302–304, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bowden:2005:HPS

[BBW05] A. J. Bowden, Cynthia V. Burek, and Richard Wilding, editors. History of palaeobotany: selected essays, volume 241 of Geological Society special publication. The Geological Society, London, UK, 2005. ISBN 1-86239-174-2. 304 pp. LCCN QE904.A1 H55 2005. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy0602/2005432693.html.

Banks:2000:LSJ

[BC00] Joseph Banks and Neil Chambers. The letters of Sir Joseph Banks: a selection, 1768–1820. Imperial College Press, London, UK, 2000. ISBN 1-86094-204-0. xliii + 420 pp. LCCN Q143.B17 A4 2000.

Ben-Chaim:2001:DNG

[BC01] Michael Ben-Chaim. The discovery of natural goods: Newton’s vocation as an ‘experimental philosopher’. British Journal for the History of Science, 34(4):395–416, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028371.

Bray:2007:GTP

[BDLM07] Francesca Bray, Vera Dorofeeva-Lichtmann, and Georges Métailie, editors. Graphics and text in the production of technical knowledge in China: the warp and the weft, volume 79 of Sinica Leidensia, 0169-9563. Brill, Leiden, The Netherlands, 2007. ISBN 90-04-16063-9 (hardcover). xiii + 772 pp. LCCN T10.5 .G685 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0727/2007038666.html.

Beauchamp:2001:HT

[Bea01] K. G. Beauchamp. History of telegraphy, volume 26 of IEE history of technology series. Institution of Electrical Engineers (IEE), London, UK, 2001. ISBN 0-85296-792-6. xxii + 413 pp. LCCN TK5115 .B43

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0631/2004533187-d.html;
http://www.loc.gov/catdir/
enhancements/fy0631/
2004533187-t.html.

Bednarek:2003:RCF

[Bed03] Janet R. Daly (Janet Rose Daly) Bednarek, editor. Reconsidering a century of flight. University of North Carolina Press, Chapel Hill, NC, USA, 2003. ISBN 0-8078-2815-7 (cloth), 0-8078-5488-3 (paperback). xii + 300 pp. LCCN TL521 .R335

2003. URL http://www.loc.
gov/catdir/description/
unc041/2003006420.html;
http://www.loc.gov/catdir/
toc/ecip041/2003006420.
html.

Beecroft:2001:BRF

[Bee01] Mark Beecroft. Book review: Felix Driver, Geography Militant: Cultures of Exploration and Empire. Oxford: Blackwell, 2001. Pp. viii + 258. ISBN 0-631-20112-2. £16.99, $29.95 (paperback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028104.

Bektas:2001:DAG

[Bek01] Yakup Bektas. Displaying the American genius: the electromagnetic telegraph in the wider world. British Journal for the History of Science, 34(2): 199–232, June 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028010.

Bektas:2006:BRE

[Bek06] Yakup Bektas. Book review:

Ekmeleddin Ihsanoglu, Science, Technology and Learning in the Ottoman Empire: Western Influence, Local Institutions and the Transfer of Knowledge. Variorum Collected Studies Series, CS773. Aldershot: Ashgate, 2004. Pp. xiv + 352. ISBN 0-86078-924-1. £59.50 (hardback). British Journal for the History of Science, 39(2):287–288, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028615.

Bennett:2000:BRG

[Ben00] Jim Bennett. Book review: I.

S. Glass, Victorian Telescope Makers: The Lives and Letters of Thomas and Howard Grubb. Bristol and Philadelphia: Institute of Physics Publishing, 1997. Pp. xiii + 279. ISBN 0-7503-0454-5. £30.00, $50.00. British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027927.

Bennett:2002:BRB

[Ben02] Jim Bennett. Book review:

Bruce Stephenson, Marvin Bolt and Anna Felicity Friedman, The Universe Unveiled: Instruments and Images through History. Chicago: Adler Planetarium and Astronomy Museum and Cambridge: Cambridge University Press, 2000. Pp. 152. ISBN 0-521-79143-X. £19.95, $29.95 (hardback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bennett:2003:KDS

[Ben03] Jim Bennett. Knowing and doing in the Sixteenth Century: what were instruments for? British Journal for the History of Science, 36(2):129–150, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028230.

Bender:2004:ESP

[Ben04a] Bert Bender. Evolution and “the sex problem”: American narratives during the eclipse of Darwinism. Kent State University Press, Kent, OH, USA, 2004. ISBN 0-87338-809-7. xvi + 389 pp. LCCN PS374.S33 B46

                                        2004. URL http://www.loc.
                                        gov/catdir/toc/ecip0417/
                                        2004008855.html.

Bennett:2004:BRS

[Ben04b] Jim Bennett. Book review: Sharon Macdonald, Behind the Scenes at the Science Museum. Materializing Culture. Oxford and New York: Berg, 2002. Pp. xiii + 293. ISBN 1-85973-571-1. £14.00 (paperback). British Journal for the History of Science, 37(1):99–100, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028259.

Bennett:2006:BRR

[Ben06] Jim Bennett. Book review:

Robert D. Huerta, Giants of Delft. Johannes Vermeer and the Natural Philosophers: The Parallel Search for Knowledge during the Age of Discovery. Lewisburg: Bucknell University Press and London: Associated University Presses, 2003. Pp. 156. ISBN 0-8387-5538-0. £39.50 (hardback). British Journal for the History of Science, 39(1):131, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028558.

Beretta:2001:ICS

[Ber01] Marco Beretta. Imaging a career in science: the iconography of Antoine Laurent Lavoisier, volume 1 (Bologna), 29 (Uppsala) of Bologna studies in scientific heritage; Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2001. ISBN 0-88135-294-2. xvii + 126 pp. LCCN QD22.L4

B29 2001. URL http://www.

                                        loc.gov/catdir/toc/fy031/
                                        2001049615.html.

Beretta:2005:PPN

[Ber05] Marco Beretta, editor. From private to public: natural collections and museums, volume 32; 5 of Uppsala studies in history of science; European studies in science history and the arts. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2005. ISBN 0-88135-360-4. ix + 252 pp. LCCN Q105.E8 F76 2005.

Bernard:2006:BRR

[Ber06a] Alain Bernard. Book review: Reviel Netz (ed. and tr.), The Works of Archimedes: Translated into English, together with Eutocius’ Commentaries, with Commentary, and Critical Edition of the Diagrams. Vol. 1: The Two Books on the Sphere and the Cylinder. Cambridge: Cambridge University Press, 2004. Pp. x + 375. ISBN 0-521-66160-9. £75.00 (hardback). British Journal for the History of Science, 39(4):604–605, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028527.

Bernard:2006:BRB

[Ber06b] Alain Bernard. Book review: The works of Archimedes. Vol. I, translated into English, together with Eutocius’ commentaries, with commentary, and critical edition of the diagrams by Reviel Netz [Cambridge Univ. Press, Cambridge, 2004; MR2093668]. British Journal for the History of Science, 39(4(143)):604– 605, 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

BertoloniMeli:2006:TOT

[Ber06c] Domenico Bertoloni Meli. Thinking with objects: the transformation of mechanics in the seventeenth century. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8426-8 (hardcover), 0-8018-8427-6 (paperback). xii + 389 pp. LCCN QC125.2 .M45

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2006002642-b.html;
http://www.loc.gov/catdir/
enhancements/fy0668/
2006002642-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip068/2006002642.html.

Bertucci:2006:RSJ

[Ber06d]  Paola Bertucci.  Revealing

sparks: John Wesley and the religious utility of electrical healing. British Journal for the History of Science, 39(3):341–362, September 2006. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bertucci:2007:VNP

[Ber07] Paola Bertucci. Viaggio nel paese delle meraviglie: scienza e curiosità nell’Italia del Settecento, volume 161 of Nuova cultura. Bollati Boringhieri, Torino, Italy, 1. edition, 2007. ISBN 88-339-1802-5. 294 + 4 pp. LCCN Q127.I8 B484 2007. 28.00

EUR. URL http://www.loc.
gov/catdir/toc/casalini07/
0758296X.pdf.

Bertucci:2008:BRJ

[Ber08] Paola Bertucci. Book review: James Delbourgo, A Most Amazing Scene of Wonders: Electricity and Enlightenment in Early America. Cambridge, MA and London: Harvard University Press, 2006. Pp. xi + 367. ISBN 0-674-02299-8. $29.95, £19.95 (hardback). British Journal for the History of Science, 41(2):297–298, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Beurton:2000:CGD

[BFR00] Peter J. Beurton, Raphael Falk, and Hans-Jörg Rheinberger, editors. The concept of the gene in development and evolution: historical and epistemological perspectives. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77187-0. xvi + 384 pp. LCCN QH447 .C66 2000. URL http:

//www.loc.gov/catdir/
description/cam0210/
99042106.html;   http://
www.loc.gov/catdir/samples/
cam032/99042106.html;
http://www.loc.gov/catdir/
toc/cam025/99042106.html.

Bud:2002:CWH

[BG02] Robert Bud and Philip Gummett, editors. Cold War, hot science: applied research in Britain’s defence laboratories, 1945–1990. NMSI Trading Ltd., London, UK, 2002. ISBN 1-900747-47-2. xix + 426 pp. LCCN U395.G7 C65 2002.

Bromer:2000:EDB

[BHR00] Rainer Brömer, Uwe Hossfeld, and Nicolaas A. Rupke, editors. Evolutionsbiologie von Darwin bis heute. (German) [Evolutionary biology from Darwin to today], volume 4 of Verhandlungen zur Geschichte und Theorie der Biologie, 1435-7852. VWB, Verlag für Wissenschaft und Bildung, Berlin, Germany, 2000. ISBN 3-86135-382-2. 425 pp. LCCN QH360 .E86 2000.

Blackwell:2006:BSG

[BI06] Richard J. Blackwell and Melchior Inchofer. Behind the scenes at Galileo’s trial: including the first English translation of Melchior Inchofer’s Tractatus syllepticus. University of Notre Dame Press, Notre Dame, IN, USA, 2006. ISBN 0-268-02201-1. xiii + 245 pp. LCCN QB36.G2 B573 2006. URL http://www.

loc.gov/catdir/toc/
ecip0614/2006016783.html.

Biagioli:2006:GIC

[Bia06] Mario Biagioli. Galileo’s instruments of credit: telescopes, images, secrecy. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-04561-7 (cloth). 302 pp. LCCN Q180.55.M67 B53

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0662/2005014562-b.html;
http://www.loc.gov/catdir/
enhancements/fy0662/
2005014562-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0514/2005014562.html.

Bicks:2003:MSS

[Bic03] Caroline Bicks. Midwiving subjects in Shakespeare’s England. Women and gender in the early modern world. Ashgate, Aldershot, Hampshire, England, 2003. ISBN 0-7546-0938-3. xi + 211 pp. LCCN PR2992.M53 B53

2003. URL http://www.loc.
gov/catdir/toc/fy042/
2002032614.html.

Bigg:2001:BRI

[Big01] Charlotte Bigg. Book review: Ileana Chinnici, La Carte du ciel: Correspondance inédite conservée dans les Archives de l’Observatoire de Paris. Paris: Observatoire de Paris; Palermo: Osservatorio Astronomico di Palermo G. S. Vaiana (with the support of the International Astronomical Union), 1999. Pp. xviii + 475. ISBN 2-901057-40-3. FF 150.00, ECU 22.87. British Journal for the History of Science, 34 (1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028044.

Bigg:2006:BRA

[Big06] Charlotte Bigg. Book review:

Alex Soojung-Kim Pang, Empire and the Sun: Victorian Solar Eclipse Expeditions. Writing Science. Stanford: Stanford University Press, 2002. Pp. xii + 196. ISBN 0-8047-3926-9. £16.95 (paperback). British Journal for the History of Science, 39 (1):134–135, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028561.

Bigg:2007:BRJ

[Big07] Charlotte Bigg. Book review:

John L. Heilbron (ed.), The Oxford Guide to the History of Physics and Astronomy. New York: Oxford University Press, 2005. Pp. xxi + 358. ISBN 0-19-517198-5. $35.00 (hardback). British Journal for the History of Science, 40(4):614–615, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Birstein:2001:PKT

[Bir01] Vadim J. Birstein. The perversion of knowledge: the true story of Soviet science. Westview Press, Boulder, CO, USA, 2001. ISBN 0-8133-3907-3. xx + 492 pp. LCCN Q127.S65 B57

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0831/2001057013-b.html;
http://www.loc.gov/catdir/
enhancements/fy0831/
2001057013-d.html.

Bivins:2006:BRM

[Biv06] Roberta Bivins. Book reviews: Mark Harrison, Disease and the Modern World: 1500 to the Present Day. Cambridge: Polity Press, 2004. Pp. vi + 270. ISBN 0-7456-2810-9. £17.99, $26.95 (paperback). Kenneth F. Kiple (ed.), The Cambridge Historical Dictionary of Disease. Cambridge: Cambridge University Press, 2003. Pp. xiii+412. ISBN 0-521-53026-1. £19.95, $27.00 (paperback). British Journal for the History of Science, 39 (2):282–283, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028611.

Bivins:2007:AMH

[Biv07] Roberta E. Bivins. Alternative medicine?: a history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-921887-0 (hardcover). xvii + 238 pp. LCCN R733 .B58 2007.

Bivins:2008:BRW

[Biv08]  Roberta Bivins.  Book re-

view: W. F. Bynum, Anne Hardy, Stephen Jacyna, Christopher Lawrence and E. M. Tansey, The Western Medical Tradition 1800 to 2000. Cambridge: Cambridge University Press, 2006. Pp. xiii + 614. ISBN 0-5214-7565-5. £19.99, $29.99 (paperback). British Journal for the History of Science, 41(2):279– 280, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Berggren:2000:PBA

[BJ00] J. L. Berggren and Alexander Jones. Ptolemy’s Geography: an annotated translation of the theoretical chapters. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-01042-0. xii + 192 pp. LCCN GA881.1 .B47 2000. URL http://www.

loc.gov/catdir/bios/
prin051/00060625.html;
http://www.loc.gov/catdir/
description/prin021/
00060625.html;
http://www.loc.gov/catdir/
toc/prin031/00060625.html.

Bacon:2000:NO

[BJS00] Francis Bacon, Lisa Jardine, and Michael Silverthorne. The new organon. Cambridge texts in the history of philosophy. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-56399-2 hb, 0-521-56483-2 pb. xxxv + 252 pp. LCCN B1168.E5 J3713 2000. URL http://www.

loc.gov/catdir/description/
cam029/99023266.html;
http://www.loc.gov/catdir/
samples/cam032/99023266.
html;
http://www.loc.gov/catdir/
toc/cam024/99023266.html.

Blackman:2003:BRC

[Bla03] Helen Blackman. Book reviews:

Carl Djerassi, This Man’s Pill: Reflections on the 50th Birthday of the Pill. Oxford: Oxford University Press, 2001. Pp. xi + 308. ISBN 0-19-850872-7. £12.99, US$22.50 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028296.

Blackman:2007:LLE

[Bla07] Helen Blackman. Lampreys, lungfish and elasmobranchs: Cambridge zoology and the politics of animal selection. British Journal for the History of Science, 40(3):413–437, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500750.

Bunn:2001:PBH

[BLR01] G. C. (Geoffrey C.) Bunn, A. D. (Alexander D.) Lovie, and Graham Richards, editors. Psychology in Britain: historical essays and personal reflections. BPS Books, Leicester, UK, 2001. ISBN 1-85433-332-1 (paperback). xiv + 495 pp. LCCN BF108.G7 P78 2001.

Brooke:2005:HEM

[BM05]  John  Hed-

ley Brooke and Ian Maclean, editors. Heterodoxy in early modern science and religion. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-926897-5. xxi + 373 pp. LCCN BL240.3 .H48

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0637/2005019524-d.html;
http://www.loc.gov/catdir/
enhancements/fy0637/
2005019524-t.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2005019524-b.html.

Boden:2006:MMH

[Bod06] Margaret A. Boden. Mind as machine: a history of cognitive science. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-924144-9 (set), 0-19-929237-X (vol. 1), 0-19-929238-8 (vol. 2). xlviii + 700 (vol. 1), xxiii + 702–1631 (vol.

2) pp. LCCN BF311 .B576

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0640/2006011795-t.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
2006011795-d.html.

Bohlin:2001:BRL

[Boh01]  Ingemar Bohlin.  Book re-

view: Lorraine Daston (ed.), Biographies of Scientific Objects. Chicago and London: University of Chicago Press, 2000. Pp. x + 308. ISBN 0-226-13672-8. £13.50, $19.00. British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028105.

Boon:2007:BRC

[Boo07] Timothy Boon. Book reviews:

Christopher Frayling, Mad, Bad and Dangerous? The Scientist and the Cinema. London: Reaktion Books, 2005. Distributed in North America by University of Chicago Press. Pp. 239. ISBN 1-86189-255-1. £19.95, $35.00 (hardback). ISBN 1-8618-9285-0. £12.95, $19.95 (paperback). British Journal for the History of Science, 40(1):114–115, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500685.

Borck:2003:BRS

[Bor03] Cornelius Borck. Book review: Stephen P. Turner, Brains/ Practices/ Relativism: Social Theory after Cognitive Science. Chicago and London: University of Chicago Press, 2002. Pp. ix + 214. ISBN 0-226-81740-7. £12.00, $19.00 (paperback). British Journal for the History of Science, 36(3): 383–384, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028174.

Bordogna:2007:IDU

[Bor07] Francesca Bordogna. Inner division and uncertain contours: William James and the politics of the modern self. British Journal for the History of Science, 40(4):505–536, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Boschiero:2002:NPI

[Bos02] Luciano Boschiero. Natural philosophizing inside the late seventeenth-century Tuscan court. British Journal for the History of Science, 35(4):383–410, December 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028274.

Boudri:2002:WWM

[Bou02] Johan Christiaan Boudri. What Was Mechanical About Mechanics: the Concept of Force Between Metaphysics and Mechanics from Newton to Lagrange, volume 224 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0233-5, 94-017-3672-3. ISSN 0068-0346. xvi + 276 pp. LCCN Q174 .B67 vol. 224;

QC73. URL http://www.loc.
gov/catdir/enhancements/
fy0814/2001050645-b.html;
http://www.loc.gov/catdir/
enhancements/fy0814/
2001050645-t.html;
http://www.loc.gov/catdir/
enhancements/fy0823/
2001050645-d.html;   https:
//link.springer.com/book/
10.1007/978-94-017-3672-5.

Bowring:2003:SSC

[Bow03] Finn Bowring. Science, seeds, and cyborgs: biotechnology and the appropriation of life. Verso, London, UK, 2003. ISBN 1-85984-687-4. xiii + 338 pp. LCCN TP248.6 .B695 2003.

Bowler:2004:BRS

[Bow04] Peter J. Bowler. Book review: Stephen Jay Gould, I Have Landed: Splashes and Reflections in Natural History. London: Vintage, 2002. Pp. 418. ISBN 0-09-974971-8. £7.99 (paperback). British Journal for the History of Science, 37(2):223–224, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028351.

Bowker:2005:MPS

[Bow05] Geoffrey C. Bowker. Memory practices in the sciences. Inside technology. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-02589-2. xi + 261 pp. LCCN Q387 .B69 2005. URL http://

                                        www.loc.gov/catdir/toc/
                                        fy0613/2005050492.html.

Bowler:2006:BRD

[Bow06a] Peter J. Bowler. Book review: David Knight (ed.), The Evolution Debate 1813–1870. London and New York: Routledge in association with the Natural History Museum, 2003. Pp. 3748. ISBN 0-415-28922-X (set of nine volumes). £895.00 (hardback). British Journal for the History of Science, 39(1):140–141, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028565.

Bowler:2006:PAE

[Bow06b] Peter J. Bowler. Presidential address: Experts and publishers: writing popular science in early twentieth-century Britain, writing popular history of science now. British Journal for the History of Science, 39(2):159–187, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bowler:2007:BRR

[Bow07] Peter J. Bowler. Book review:

Ron Amundson, The Changing Role of the Embryo in Evolutionary Thought: The Roots of Evo-Devo. Cambridge Studies in Philosophy and Biology. Cambridge: Cambridge University Press, 2005. Pp. xiii + 280. ISBN 0-521-80699-2. $75.00 (hardback). British Journal for the History of Science, 40(3): 460–462, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500764.

Bertucci:2001:EBE

[BP01] Paola Bertucci and Giuliano Pancaldi, editors. Electric bodies: episodes in the history of medical electricity. Università di Bologna, Bologna, Italy, 2001. ISBN 88-900162-2-1. 298 pp. LCCN RM871 .E325 2001. URL

http://www.loc.gov/catdir/
toc/fy0604/2002476564.html.

Bynum:2005:ODS

[BP05] W. F. (William F.) Bynum and Roy Porter, editors. Oxford dictionary of scientific quotations. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-858409-1. xvi + 712 pp. LCCN Q173 .O915 2005.

URL http://www.e-streams.
com/es0910/es0910_4499.
html;   http://www.loc.gov/
catdir/enhancements/fy0640/
2005277260-d.html;
http://www.loc.gov/catdir/
enhancements/fy0640/
2005277260-t.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
2005277260-b.html.

Bacon:1996:PSC

[BR96] Francis Bacon and Graham Rees. Philosophical studies, c. 1611–c. 1619, volume 6 of The Oxford Francis Bacon. Clarendon Press, New York, NY, USA, 1996. ISBN 0-19-812290-X (cloth). cxvi + 503 pp. LCCN B1153

1996. URL http://www.loc.
gov/catdir/enhancements/
fy0605/95038755-d.html.

Bacon:2000:IML

[BR00] Francis Bacon and Graham Rees. The Instauratio magna: last writings, volume 13 of The Oxford Francis Bacon. Clarendon Press, New York, NY, USA, 2000. ISBN 0-19-818470-0. xcvi + 363 pp. LCCN B1180.I572

E5 2000. URL http://www.loc.
gov/catdir/enhancements/
fy0610/00033977-d.html;

http://www.loc.gov/catdir/
enhancements/fy0610/
00033977-t.html.

Brackenridge:2000:BRI

[Bra00] J. Bruce Brackenridge. Book review: Isaac Newton. The Principia: Mathematical Principles of Natural Philosophy, 3rd edition (1726). Newly translated by I. Bernard Cohen and Anne Whitman. With a supplement by I. Bernard Cohen. Berkeley: University of California Press, 1999. Pp. 1025. ISBN 0-520-08816-6. £60.00, $75.00 (cloth); 0-520-08817-4, £24.95, $35.00 (paperback). British Journal for the History of Science, 33 (2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027928.

Bray:2007:BRB

[Bra07]  Francesca Bray.  Book re-

views: Benjamin A. Elman, On Their Own Terms: Science in China, 1550–1900. Cambridge, MA and London: Harvard University Press, 2005. Pp. xxxviii + 567. ISBN 0-674-01685-8. £35.95 (hardback). Benjamin A. Elman, A Cultural History of Modern Science in China. Cambridge, MA and London: Harvard University Press, 2006. Pp. xv+308. ISBN 0-674-02306-4. $35.00, £22.95, EUR29.80 (hardback). British Journal for the History of Science, 40(3): 439–441, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500751.

Bray:2008:STT

[Bra08] Francesca Bray. Science, technique, technology: passages between matter and knowledge in imperial Chinese agriculture. British Journal for the History of Science, 41(3):319–344, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bregman:2004:BRR

[Bre04] Leigh D. Bregman. Book review: Richard Yeo, Science in the Public Sphere: Natural Knowledge in British Culture 1800–1860. Variorum Collected Studies Series, CS726. Aldershot: Ashgate, 2001. Pp.

324. ISBN 0-86078-865-2. £57.50 (hardback). British Journal for the History of Science, 37(4): 477–478, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028654.

Bregman:2005:BRL

[Bre05] Leigh D. Bregman. Book review: Lynn K. Nyhart and Thomas H. Broman (eds.), Science and Civil Society. Osiris,

17. Chicago and London: University of Chicago Press, 2002. Pp. x + 373. ISBN 0-226-07371-8. $50.50 (hardback). British Journal for the History of Science, 38(2):220–221, June 2005.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028698.

Bregman:2006:BRS

[Bre06] Leigh D. Bregman. Book review:

Steven Ruskin, John Herschel’s Cape Voyage: Private Science, Public Imagination and the Ambitions of Empire. Science, Technology and Culture, 1700–1945. Aldershot: Ashgate, 2004. Pp. xxix + 229. ISBN 0-7546-3558-9. £45.00 (hardback). British Journal for the History of Science, 39(2):298–299, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028624.

Bright:2004:RDS

[Bri04] Kevin Bright. The Royal Dublin Society, 1815–45. Maynooth historical studies. Four Courts Press, Dublin, Ireland, 2004. ISBN 1-85182-813-3 (hard). 282 + 32 pp. LCCN Q41 .B658 2004.

Brown:2001:WRS

[Bro01a] James Robert Brown. Who rules in science?: an opinionated guide to the wars. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00652-6. xi + 236 pp. LCCN Q175.5 .B758 2001.

Brownstein:2001:BRR

[Bro01b] Daniel Brownstein. Book review: Roger French, Dissection and Vivisection in the European Renaissance. Aldershot: Ashgate, 1999. Pp. ix + 289. ISBN 1-85928-361-6. £55.00, $99.95 (hardback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028376.

Brockliss:2002:CWE

[Bro02a] L. W. B. Brockliss. Calvet’s web: Enlightenment and the republic of letters in Eighteenth-Century France. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-924748-X. xiv + 471 + 12 pp. LCCN R507.C285 B76

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0614/2002512708-d.html;
http://www.loc.gov/catdir/
enhancements/fy0614/
2002512708-t.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2002512708-b.html.

Bromer:2002:BRE

[Brö02b] Rainer Brömer. Book review: Ekmeleddin Ihsanoglu and Feza Günergun, Science in Islamic Civilisation: Proceedings of the International Symposia ‘Science Institutions in Islamic Civilisation’ and ‘Science and Technology in the Turkish and Islamic World’. Studies and Sources on the History of Science, 9. Istanbul: Research Centre for Islamic History, Art and Culture (IRCICA) 2000. Pp. vi + 289. ISBN 92-9063-095-7. $40.00 (hardback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brown:2002:IRH

[Bro02c] George Ingham Brown. Invisible rays: the history of radioactivity. Sutton, Stroud, UK, 2002. ISBN 0-7509-2667-8. 248 pp. LCCN QC794.98 .B76 2002. US$19.99.

URL http://www.gbv.de/dms/
bowker/toc/9780750926676.

Browne:1995:CDB

[Bro02d] E. J. (E. Janet) Browne. Charles Darwin: a biography. Knopf, New York, NY, USA, 1995–2002. ISBN 0-394-57942-9 (vol. 1), 0-679-42932-8 (vol. 2 : jkt.). xiii + 605 + 32 (vol. 1), 591 + 24 (vol. 2) pp. LCCN QH31.D2

B84 1995. URL http://www.
loc.gov/catdir/bios/
random057/94006598.html;
http://www.loc.gov/catdir/
description/random041/
94006598.html;   http://
www.loc.gov/catdir/samples/
random041/94006598.html.

Brownstein:2002:BRA

[Bro02e] Daniel Brownstein. Book review:

Andrea Carlino, Paper Bodies: A Catalogue of Anatomical Fugitive Sheets 1538–1687. English translation by Noga Arikha. Medical History, Supplement 19. London: Wellcome Institute for the History of Medicine, 1999. Pp. xvi + 352. ISBN 0-85484-069-9. $50.00 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028404.

Brown:2003:ECA

[Bro03a] Mark Brown. Ethnology and colonial administration in Nineteenth-Century British India: the question of native crime and criminality. British Journal for the History of Science, 36(2):201–219, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028233.

Browne:2003:BRJ

[Bro03b] Janet Browne. Book review: John C. Greene, Debating Darwin: Adventures of a Scholar. Claremont: Regina Books, 1999. Pp. vi + 288. ISBN 0-941690-85-7. No price given (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028297.

Browne:2003:BRP

[Bro03c] Janet Browne. Book review: Peter Hamilton and Roger Hargreaves, The Beautiful and the Damned: The Creation of Identity in Nineteenth-Century Photography. Aldershot: Lund Humphries in association with the National Portrait Gallery, 2001. Pp. vi + 122. ISBN 0-85331-821-2. £25.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028298.

Browne:2003:ORS

[Bro03d] Janet Browne. Obituary: Roy Sydney Porter (1946–2002). British Journal for the History of Science, 36(1):83–86, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028294.

Brock:2004:BRT

[Bro04a] William H. Brock. Book review:

Trevor H. Levere and Gerard L’E. Turner, Discussing Chemistry and Steam: The Minutes of a Coffee House Philosophical Society 1780–1787. With contributions from Jan Golinski and Larry Stewart. Oxford: Oxford University Press, 2002. Pp. viii + 284. ISBN 0-19-851530-8. £65.00 (hardback). British Journal for the History of Science, 37(1): 106–107, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028264.

Bromer:2004:BRH

[Brö04b] Rainer Brömer. Book review:

Henry Harris, Things Come to Life: Spontaneous Generation Revisited. Oxford: Oxford University Press, 2002. Pp. ix + 168. ISBN 0-19-851538-3. £20.00 (hardback). British Journal for the History of Science, 37 (2):224–225, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028352.

Brownstein:2004:BRD

[Bro04c] Daniel Brownstein. Book review: David Cantor (ed.), Reinventing Hippocrates. The History of Medicine in Context. Aldershot: Ashgate, 2002. Pp. x + 341. ISBN 0-7546-0528-0. £55.00 (hardback). British Journal for the History of Science, 37(2):197–199, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028332.

Brock:2005:BRB

[Bro05a] Claire Brock. Book review: Barbara T. Gates (ed.), In Nature’s Name: An Anthology of Women’s Writing and Illustration, 1780–1930. Chicago and London: University of Chicago Press, 2002. Pp. xxvi + 673. ISBN 0-226-28446-8. £17.50, $27.50 (paperback). British Journal for the History of Science, 38(1):110–111, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028592.

Brock:2005:BRD

[Bro05b] W. H. Brock. Book review:

David Philip Miller, Discovering Water: James Watt, Henry Cavendish and the Nineteenth-Century ‘Water Controversy’. Science, Technology and Culture, 1700–1945. Aldershot: Ashgate, 2004. Pp. xiii + 316. ISBN 0-7546-3177-X. £55.00 (hardback). British Journal for the History of Science, 38(2):232–234, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028708.

Brown:2005:JDB

[Bro05c] Andrew Brown. J. D. Bernal:

the sage of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-851544-8. xiv + 562 pp. LCCN QC16.B457 B76

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0635/2005023505-d.html;
http://www.loc.gov/catdir/
enhancements/fy0723/
2005023505-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0517/2005023505.html.

Browne:2005:PAC

[Bro05d] Janet Browne. Presidential address: Commemorating Darwin. British Journal for the History of Science, 38(3):251–274, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brock:2006:BRA

[Bro06a] Claire Brock. Book review: Allan Chapman, Mary Somerville and the World of Science. Bristol: Canopus Publishing, 2004. Pp. ix + 157. ISBN 0-9537868-4-6. £12.95 (hardback). British Journal for the History of Science, 39(2):297–298, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028623.

Brock:2006:PWM

[Bro06b] Claire Brock. The public worth of Mary Somerville. British Journal for the History of Science, 39(2):255–272, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Broks:2006:UPS

[Bro06c] Peter Broks. Understanding popular science. Issues in cultural and media studies. Open University Press, Maidenhead, England, 2006. ISBN 0-335-21548-3 (paper), 0-335-21549-1 (cloth), 0-335-22437-7 (e-book). x + 183 pp. LCCN Q175.5.

                                        URL http://www.myilibrary.
                                        com?id=95087.

Brock:2008:WCC

[Bro08a] W. H. (William Hodson) Brock. William Crookes (1832–1919) and the commercialization of science. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2008. ISBN 0-7546-6322-1. xxvii + 556 pp. LCCN Q143.C7 B76 2008.

Brooke:2008:BRS

[Bro08b] John Hedley Brooke. Book review: Stephen Gaukroger, The Emergence of a Scientific Culture: Science and the Shaping of Modernity, 1210–1685. Oxford: Clarendon Press, 2006. Pp. ix + 563. ISBN 0-19-929644-8. £35.00 (hardback). British Journal for the History of Science, 41(2):276–277, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brooke:2009:BRM

[Bro09] John Hedley Brooke. Book review: Matthew Stanley, Practical Mystic: Religion, Science and A. S. Eddington. Chicago and London: University of Chicago Press, 2007. Pp. x + 313. ISBN 978-0-226-77097-0. $37.50, £22.00 (hardback). British Journal for the History of Science, 42 (1):125–127, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brundell:2001:CCP

[Bru01] Barry Brundell. Catholic Church politics and evolution theory, 1894–1902. British Journal for the History of Science, 34 (1):81–95, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028040.

Brush:2006:BRL

[Bru06] Stephen G. Brush. Book review: Lisa M. Dolling, Arthur F. Gianelli and Glenn N. Statile (eds.), The Tests of Time: Readings in the Development of Physical Theory. Princeton and Oxford: Princeton University Press, 2003. Pp. xliii + 716. ISBN 0-691-09085-8. $25.95 (paperback). British Journal for the History of Science, 39(1):125–126, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028553.

Brundtland:2008:MNP

[Bru08]  Terje  Brundtland.

From medicine to natural philosophy: Francis Hauksbee’s way to the air-pump. British Journal for the History of Science, 41(2):209–240, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Brusius:2009:BRH

[Bru09] Mirjam Brusius. Book review: Horst Bredekamp, Birgit Schneider and Vera Dünkel (eds.), Das technische Bild. Kompendium zu einer Stilgeschichte wissenschaftlicher Bilder. Berlin: Akademie Verlag2008. Pp. 231. ISBN 978-3-05-004496-5. EUR29.80 (paperback). British Journal for the History of Science, 42(4):611– 612, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Bacon:2004:IMP

[BRW04] Francis Bacon, Graham Rees, and Maria Wakely. The Instauratio magna. Part 2, Novum organum and associated texts, volume 11 of The Oxford Francis Bacon. Clarendon Press, New York, NY, USA, 2004. ISBN 0-19-924792-7. cxxviii + 634 pp. LCCN B1168.E5 R44

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0620/2004302456-d.html;
http://www.loc.gov/catdir/
enhancements/fy0620/
2004302456-t.html.

Bastian:2001:ESG

[BS01] Henry Charlton Bastian and James Edgar Strick. Evolution and the spontaneous generation debate. Thoemmes, Bristol, UK, 2001. ISBN 1-85506-872-9. 8007 pp. LCCN ???? £395.00. Six volumes.

Bravo:2002:NAC

[BS02] Michael Bravo and Sverker Sörlin, editors. Narrating the Arctic: a cultural history of Nordic scientific practices. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2002. ISBN 0-88135-385-X. ix + 373 pp. LCCN G630.S33 N37 2002.

Bredekamp:2008:TBK

[BSD08] Horst Bredekamp, Birgit Schneider, and Vera Dünkel, editors. Das Technische Bild: Kompendium zu einer Stilgeschichte wissenschaftlicher Bilder. (German) The technical picture: Manual for one style of the history of scientific pictures. Akademie-Verlag, Berlin, Germany, 2008. ISBN 3-05-004496-9. 231 pp. LCCN ????

Buchanan:2002:BLT

[Buc02] R. A. (Robert Angus) Buchanan. Brunel: the life and times of Isambard Kingdom Brunel. Hambledon and London, London, UK, 2002. ISBN 1-85285-331-X. xxiv + 294 + 16 pp. LCCN TA140.B75 B83

2002. URL http://www.h-net.
org/review/hrev-a0c8t4-aa;
 http://www.loc.gov/catdir/
bios/hol052/2002392805.
html;   http://www.loc.gov/
catdir/description/hol032/
2002392805.html;
http://www.loc.gov/catdir/
toc/fy0709/2002392805.html.

Bud:2009:BRS

[Bud09]  Robert Bud.  Book review:

Steven Shapin, The Scientific Life: A Moral History of a Late Modern Vocation. Chicago and London: University of Chicago Press, 2008. Pp. xvii + 468. ISBN 978-0-226-75024-8. $29.00 (hardback). British Journal for the History of Science, 42(4): 632–634, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Buisseret:2007:OCW

[Bui07] David Buisseret, editor. The Oxford companion to world exploration. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-514922-X (hardcover). xxviii + 478 (vol. 1), vii + 501 (vol. 2) pp. LCCN G80 .O95

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0726/2006027968-b.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
2006027968-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0619/2006027968.html.

Bulmer:2004:DJS

[Bul04] Michael Bulmer. Did Jenkin’s swamping argument invalidate Darwin’s theory of natural selection? British Journal for the History of Science, 37(3): 281–297, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028425.

Burns:2000:JLB

[Bur00] R. W. Burns. John Logie Baird:

television pioneer, volume 28 of IEE history of technology series. Institution of Electrical Engineers (IEE), London, UK, 2000. ISBN 0-85296-797-7. xxv + 417 pp. LCCN TK6635.B3 B87 2000.

Burbridge:2001:FGT

[Bur01] David Burbridge. Francis Galton on twins, heredity and social class. British Journal for the History of Science, 34(3): 323–340, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028101.

Burns:2003:SEE

[Bur03] William E. Burns. Science in the Enlightenment: an encyclopedia. ABC-CLIO’s history of science series. ABC-CLIO, Santa Barbara, CA, USA, 2003. ISBN 1-57607-886-8, 1-57607-887-6 (ebook). xxvii + 355 pp. LCCN Q121 .B87 2003. URL http:/

                                        /www.e-streams.com/es0708/
                                        es0708_3497.html;
                                        http://www.loc.gov/catdir/
                                        description/abcclio041/
                                        2003011342.html;   http://
                                        www.loc.gov/catdir/toc/
                                        ecip044/2003011342.html.

Burkhardt:2005:PBK

[Bur05] Richard W. (Richard Wellington) Burkhardt. Patterns of behavior: Konrad Lorenz, Niko Tinbergen, and the founding of ethology. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-08089-7 (cloth), 0-226-08090-0 (paperback). xii + 636 pp. LCCN QL750.5 .B87

                                        2005. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/2004004083.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0616/
                                        2004004083-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0415/2004004083.html.

Burkhardt:2007:BRP

[Bur07a] Richard W. Burkhardt. Book review: Pietro Corsi, Jean Gayon, Gabriel Gohau and Stéphane Tirard, Lamarck, philosophe de la nature. Paris: Presses Universitaires de France, 2006. Pp. xii + 167 pp. ISBN 2-13-051976-8. EUR20.00 (paperback). British Journal for the History of Science, 40(4):611–613, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Burnett:2007:TLN

[Bur07b]  D. Graham Burnett.  Try-

ing Leviathan: the nineteenthcentury New York court case that put the whale on trial and challenged the order of nature. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12950-9 (hardcover). xiv + 266 + 16 pp. LCCN KF228.M38

B87 2007. URL http://www.
loc.gov/catdir/toc/
ecip0712/2007008378.html.

Burns:2007:JFG

[Bur07c] James Burns. John Fleming and the geological deluge. British Journal for the History of Science, 40(2):205–225, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500717.

Burkhardt:2008:BRC

[Bur08] Richard W. Burkhardt. Book review: Charlotte Sleigh, Six Legs Better: A Cultural History of Myrmecology. Baltimore: Johns Hopkins University Press, 2007. Pp. viii + 302. ISBN 0-8018-8445-4. £36.50 (hardback). British Journal for the History of Science, 41(3):447–449, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Butrica:2006:BRG

[But06] Andrew J. Butrica. Book review: Graeme J. N. Gooday, The Morals of Measurement: Accuracy, Irony, and Trust in Late Victorian Electrical Practices. Cambridge and New York: Cambridge University Press, 2004. Pp. xxvi + 285. ISBN 0-521-43098-4. £55.00, $85.00 (hardback). British Journal for the History of Science, 39(4):621– 622, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028539.

Bensaude-Vincent:2000:LPS

[BV00] Bernadette Bensaude-Vincent. L’opinion publique et la science: à chacun son ignorance. (French) [Public opinion and science: to each its own ignorance]. Collection Les empêcheurs de penser en rond. Institut d’édition Sanofi-Synthélabo, Paris, France, 2000. ISBN 2-84324-141-3. 238 pp. LCCN Q172.5.P82 B46 2000.

Bensaude-Vincent:2002:SFO

[BVB02] Bernadette Bensaude-Vincent and Christine Blondel, editors. Des savants face à l’occulte, 1870–1940. (French) [Scientists versus the occult, 1870–1940]. Sciences et société. Éditions La Découverte, Paris, France, 2002. ISBN 2-7071-3616-6. 231 pp. LCCN BF1409.5 .D47 2002.

Bensaude-Vincent:2003:RSF

[BVB03]  Bernadette  Bensaude-

Vincent and Bruno Bernardi, editors. Rousseau et les sciences. (French) [Rousseau and science]. Collection Epistémologie et philosophie des sciences. L’Harmattan, Paris, France, 2003. ISBN 2-7475-5100-8. 316 pp. LCCN Q174.8 .R68 2003.

Bensaude-Vincent:2008:SSE

[BVB08]  Bernadette  Bensaude-

Vincent and Christine Blondel, editors. Science and spectacle in the European Enlightenment. Science, technology and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2008. ISBN 0-7546-6370-1. xi + 164 pp. LCCN Q175.52.E85 S354

2008. URL http://www.loc.
gov/catdir/toc/ecip0727/
2007038995.html.

Bensaude-Vincent:2007:ANE

[BVN07] Bernadette Bensaude-Vincent and William R. Newman, editors. The artificial and the natural: an evolving polarity. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-02620-1 (hardcover). vi + 331 pp. LCCN Q127.E8 A78 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip078/2007001896.html.

Buchwald:2001:HEB

[BW01] Jed Z. Buchwald and Andrew Warwick, editors. Histories of the electron: the birth of microphysics. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02494-2 (hardcover). xi + 514 pp. LCCN QC793.5.E62

                                        H57 2001. URL http://www.e-
                                        streams.com/es0509/es0509_
                                        2104.htm;
                                        http://www.loc.gov/catdir/
                                        toc/fy033/00046580.html.

Becker:2006:CTS

[BW06] Peter Becker and Richard F. Wetzell, editors. Criminals and their scientists: the history of criminology in international perspective, Publications of the German Historical Institute. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-81012-4. LCCN HV6021 .C75

                                        2006. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/2003055193.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam031/2003055193.html.

Bynum:2008:HMV

[Byn08] W. F. (William F.) Bynum. History of medicine: a very short introduction. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-921543-X. 169 pp. LCCN R131 .B974 2008. URL http://www.

loc.gov/catdir/toc/fy0805/
2008016934.html.

Ben-Zaken:2004:HSH

[BZ04] Avner Ben-Zaken. The heavens of the sky and the heavens of the heart: the Ottoman cultural context for the introduction of post-Copernican astronomy. British Journal for the History of Science, 37(1):1–28, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028254.

Cadbury:2000:DHS

[Cad00] Deborah Cadbury. The dinosaur hunters: a story of scientific rivalry and the discovery of the prehistoric world. Fourth Estate, London, UK, 2000. ISBN 1-85702-959-3. 384 pp. LCCN QE705.G7 C33 2000.

Cain:2000:TGD

[Cai00] Joe Cain. Towards a ‘greater degree of integration’: the Society for the Study of Speciation, 1939–1941. British Journal for the History of Science, 33(1):85–108, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028067.

Campbell:2008:BRG

[Cam08] Gordon Campbell. Book review: G. E. R. Lloyd, Principles and Practices in Ancient Greek and Chinese Science. Variorum Collected Studies Series.

Aldershot: Ashgate, 2006. Pp. x + 320. ISBN 0-86078-993-4. £60.00 (hardback). British Journal for the History of Science, 41(2):290–291, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cantor:2000:OCH

[Can00]  Geoffrey Cantor.  Obituary:

Casper Hakfoort (1955–1999). British Journal for the History of Science, 33(2):227–229, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027925.

Canales:2001:EFE

[Can01a] Jimena Canales. Exit the frog, enter the human: physiology and experimental psychology in nineteenth-century astronomy. British Journal for the History of Science, 34(2): 173–197, June 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028009.

Cantor:2001:BRP

[Can01b] Geoffrey Cantor. Book review: Paul Knights (ed.), The Manuscripts of Michael Faraday (1791–1867) from the Collections of the Royal Institution – the Institution of Electrical Engineers – the Guildhall Library. Wakefield: Microform Academic Publishers, 2001. 23 microfilm reels. £1058.00 (complete), £46.00 (individual reels).

Frank A. J. L. James, Guide to the Microfilm Edition of the Manuscripts of Michael Faraday (1791–1867) from the Collections of the Royal Institution – the Institution of Electrical Engineers – the Guildhall Library. Wakefield: Microform Academic Publishers, 2000. Pp.

94. ISBN 1-851170-31-6. £10.00 (paperback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028106.

Cantor:2002:RH

[Can02a] David Cantor, editor. Reinventing Hippocrates, The history of medicine in context. Ashgate, Aldershot, Hampshire, England, 2002. ISBN 0-7546-0528-0. LCCN R126.H8 R35

2002. URL http://www.loc.
gov/catdir/toc/fy031/
2001022834.html.

Cantor:2002:BRW

[Can02b] Geoffrey Cantor. Book review:

William Whewell, Of the Plurality of Worlds. A Facsimile of the First Edition of 1853: Plus Previously Unpublished Material Excised by the Author Just Before the Book Went to Press; and Whewell’s Dialogue Rebutting His Critics, Reprinted from the Second Edition. Edited and with new introductory material by Michael Ruse. Chicago and London: University of Chicago Press, 2001. Pp. v + 510. ISBN 0-226-89436-3. £13.00, $20.00 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cantor:2004:BRD

[Can04a] Geoffrey Cantor. Book review: David Boyd Haycock, William Stukeley: Science, Religion and Archaeology in Eighteenth-Century England. Woodbridge: Boydell Press, 2002. Pp. xiii + 290. ISBN 0-85115-864-1. £40.00, $70.00 (hardback). British Journal for the History of Science, 37 (2):209–210, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028339.

Cantor:2004:CRS

[Can04b] Geoffrey Cantor. Creating the Royal Society’s Sylvester Medal. British Journal for the History of Science, 37(1):75–92, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028257.

Cantrell:2004:BRB

[Can04c] John Cantrell. Book review: Bridget Howard, Mr Lean and the Engine Reporters. Redruth: Trevithick Society, 2002. Pp. 99. ISBN 0-904040-56-9. £9.99 (paperback). British Journal for the History of Science, 37(3): 353, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028134;
http://www.jstor.org/
stable/4028436.

Cantor:2005:QJS

[Can05] G. N. Cantor. Quakers, Jews, and science: religious responses to modernity and the sciences in Britain, 1650–1900. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-927668-4. xi + 420 pp. LCCN BL240.3 .C36

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0638/2005019301-d.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2005019301-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0515/2005019301.html.

Cantor:2007:BRS

[Can07] Geoffrey Cantor. Book review:

Stephen Frosh, Hate and the ‘Jewish Science’: Anti-Semitism, Nazism and Psychoanalysis. Basingstoke and New York: Palgrave Macmillan, 2005. Pp. vi + 228. ISBN 1-4039-2170-9. £25.00 (hardback). British Journal for the History of Science, 40 (2):307–309, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500741.

Caneva:2008:BRM

[Can08a] Kenneth L. Caneva. Book review: Michael Friedman and Alfred Nordmann (eds.), The Kantian Legacy in Nineteenth-Century Science. Cambridge, MA and London: MIT Press, 2006. Pp. iv + 370. ISBN 0-262-06254-2. £29.95(hardback). British Journal for the History of Science, 41(2):307–308, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cantor:2008:BRN

[Can08b] Geoffrey Cantor. Book reviews: Noah J. Efron, Judaism and Science: A Historical Introduction. Greenwood Guides to Science and Religion. Westport, CT and London: Greenwood Press, 2006. Pp. xx + 348. ISBN 978-0-313-33053-7. $65.00, £37.95 (hardback). Muzaffar Iqbal, Science and Islam. Greenwood Guides to Science and Religion. Westport, CT and London: Greenwood Press, 2007. Pp. xx+348. ISBN 978-0-313-33576-1. $65.00, £37.95 (hardback). British Journal for the History of Science, 41(2):302–304, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Canales:2009:BRC

[Can09] Jimena Canales. Book review: Chris Otter, The Victorian Eye: A Political History of Light and Vision in Britain, 1800–1910. Chicago and London: University of Chicago Press, 2008. Pp. x + 382. ISBN 978-0-226-64077-8. £13.00 (paperback). British Journal for the History of Science, 42(2):305–306, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Carlino:1999:PBC

[Car99] Andrea Carlino. Paper bodies: a catalogue of anatomical fugitive sheets, 1538–1687, volume 19 of Medical History. Supplement. Wellcome Institute for the History of Medicine, London, UK, 1999. ISBN 0-85484-069-9. xvi + 352 pp. LCCN QM11 .C37 1999.

Carneiro:2001:BRA

[Car01a] Ana Carneiro. Book review:

Alan J. Rocke, Nationalizing Science: Adolphe Wurtz and the Battle for French Chemistry. Cambridge, MA and London: MIT Press, 2001. Pp. xiii + 443. ISBN 0-262-18204-1. £29.50 (hardback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028377.

Carter:2001:RCB

[Car01b] Ian Carter. Railways and culture in Britain: the epitome of modernity. Studies in popular culture. Manchester University Press, Manchester, UK, 2001. ISBN 0-7190-5965-8 (hardcover), 0-7190-5966-6 (paperback). xi + 338 pp. LCCN PR468.R35 C37

2001. URL http://www.loc.
gov/catdir/bios/hol058/

2001044117.html;
http://www.loc.gov/catdir/
description/hol052/
2001044117.html;
http://www.loc.gov/catdir/
toc/hol051/2001044117.html.

Carneiro:2006:BRU

[Car06] Ana Carneiro. Book review: Ursula Klein, Experiments, Models, Paper Tools: Cultures of Organic Chemistry in the Nineteenth Century. Writing Science. Stanford: Stanford University Press, 2003. Pp. xi + 305. ISBN 0-8047-4359-2. £49.50 (hardback). British Journal for the History of Science, 39(1): 135–136, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028562.

Carson:2007:MMT

[Car07] John Carson. The measure of merit: talents, intelligence, and inequality in the French and American republics, 1750– 1940. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-01715-8. xvii + 401 pp. LCCN BF431.5.F8 C37

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0654/2006006604-d.html;
http://www.loc.gov/catdir/
enhancements/fy0734/
2006006604-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip069/2006006604.html.

Cavicchi:2008:WAS

[Cav08] Elizabeth Cavicchi. A witness account of solar microscope projections: collective acts integrating across personal and historical memory. British Journal for the History of Science, 41(3): 369–383, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cohen-Cole:2007:ISM

[CC07] Jamie Cohen-Cole. Instituting the science of mind: intellectual economies and disciplinary exchange at Harvard’s Center for Cognitive Studies. British Journal for the History of Science, 40(4):567–597, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cohen-Cole:2008:CMR

[CC08] Jamie Cohen-Cole. Cybernetics and the machinery of rationality. British Journal for the History of Science, 41(1):109–114, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Crowe:1998:CCS

[CDK98] Michael J. Crowe, David R.

Dyck, and James J. Kevin, editors. A calendar of the correspondence of Sir John Herschel. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63149-1. 828 pp. LCCN QB36.H59

C35 1998. URL http://www.

loc.gov/catdir/description/
cam029/98022928.html;
http://www.loc.gov/catdir/
toc/cam026/98022928.html.

Chaucer:2002:TA

[CE02] Geoffrey Chaucer and Sigmund Eisner. A treatise on the astrolabe, volume 6 of A variorum edition of the works of Geoffrey Chaucer; The prose treatises; pt.

1. University of Oklahoma Press, Norman, OK, USA, 2002. ISBN 0-8061-3413-5 (hardcover). xxiv + 358 pp. LCCN PR1891 .E38 2002.

Collins:2007:RE

[CE07] H. M. (Harry M.) Collins and Robert Evans. Rethinking expertise. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-11360-4 (cloth). 159 pp. LCCN HM651 .C64 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0732/
2007022671-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007022671-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0720/2007022671.html.

Ceccarelli:2001:SSR

[Cec01] Leah Ceccarelli. Shaping science with rhetoric: the cases of Dobzhansky, Schrödinger, and Wilson. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-09906-7 (cloth), 0-226-09907-5 (paperback). xi + 204 pp. LCCN QH303.6 .C433

2001. URL http://www.loc.
gov/catdir/bios/uchi051/

00012179.html;
http://www.loc.gov/catdir/
description/uchi051/
00012179.html;
http://www.loc.gov/catdir/
toc/uchi051/00012179.html.

Chabas:2000:AIP

[CG00a] José Chabás and Bernard R.

Goldstein. Astronomy in the Iberian Peninsula: Abraham Zacut and the transition from manuscript to print, volume 90, pt. 2 of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 2000. ISBN 0-87169-902-8 (paperback). ISSN 0065-9746. xii + 196 pp. LCCN QB36.Z3 C48 2000.

cIhsanoglu:2000:SIC

[çG00b] Ekmeleddin çIhsanoğlu and Feza Günergun, editors. Science in Islamic civilisation: proceedings of the international symposia “Science institutions in Islamic civilisation” and “Science and technology in the Turkish and Islamic world”, volume 9 of Studies and sources on the history of science series. Organisation of the Islamic Conference, Istanbul, Turkey, 2000. ISBN 92-9063-095-7. LCCN BP190.5.S3 S35 2000.

Corsi:2006:LPN

[CGGT06] Pietro Corsi, Jean Gayon, Gabriel Gohau, and Stéphane Tirard, editors. Lamarck, philosophe de la nature. (French) [Lamarck, natural philosopher]. Science, histoire et société. Presses universitaires de France, Paris, France, 2006. ISBN 2-13-051976-8. xii + 167 pp. LCCN QH31.L2 L353 2006.

Clark:1999:SEE

[CGS99] William Clark, Jan Golinski, and Simon Schaffer, editors. The sciences in enlightened Europe. University of Chicago Press, Chicago, IL, USA, 1999. ISBN 0-226-10939-9 (cloth), 0-226-10940-2 (paperback). xi + 566 pp. LCCN Q127.E8 S356

                                        1999. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/98051992.html;
                                        http://www.loc.gov/catdir/
                                        toc/uchi052/98051992.html.

Carson:2005:ROC

[CH05] Cathryn Carson and David A. Hollinger, editors. Reappraising Oppenheimer: centennial studies and reflections, volume 21 of Berkeley papers in history of science. Office for History of Science and Technology, Berkeley, CA, USA, 2005. ISBN 0-9672617-3-2. xii + 413 pp. LCCN QC16.O62 R47 2005.

Chart:2001:BRJ

[Cha01] David Chart. Book review: James H. Fetzer (ed.), Science, Explanation, and Rationality: Aspects of the Philosophy of Carl G. Hempel. Oxford: Oxford University Press, 2000. Pp. xxix + 342. ISBN 0-19-512137-6. £37.50 (hardback). British Journal for the History of Science, 34(4):453–481, December 2001.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028378.

Chaouli:2002:LPC

[Cha02a] Michel Chaouli. The laboratory of poetry: chemistry and poetics in the work of Friedrich Schlegel. Parallax. The Johns Hopkins University Press, Baltimore, MD, USA, 2002. ISBN 0-8018-6884-X (hardcover). xiv + 281 pp. LCCN PT2503.S7

C43 2002. URL http://www.
loc.gov/catdir/bios/jhu051/
2001006620.html;
http://www.loc.gov/catdir/
description/jhu051/
2001006620.html;
http://www.loc.gov/catdir/
toc/jhu051/2001006620.html.

Chappey:2002:SOH

[Cha02b] Jean-Luc Chappey. La société des observateurs de l’homme, 1799–1804: des anthropologues au temps de Bonaparte. (French) [The society of observers of humans, 1799–1804: anthropologists in Bonaparte’s time]. Bibliothèque d’histoire révolutionnaire. Société des études robespierristes, Paris, France, 2002. ISBN 2-908327-45-7. 573 pp. LCCN GN2 .C53 2002.

Chart:2003:BRM

[Cha03] David Chart. Book review:

Michael Friedman, Dynamics of Reason: The 1999 Kant Lectures at Stanford University. Stanford, CA: CSLI Publications, 2001. Pp. xiv + 141. ISBN 1-57586-292-1. £12.50, $19.50 (paperback). British Journal for the History of Science, 36(1): 87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028299.

Chang:2004:ITM

[Cha04] Hasok Chang. Inventing temperature: measurement and scientific progress. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-517127-6. xviii + 286 pp. LCCN QC271.6 .C46

                                        2004. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0614/2003058489-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0726/
                                        2003058489-b.html.

Chaplin:2005:BRK

[Cha05] Simon Chaplin. Book review: Ken Arnold and Danielle Olsen (eds.), Medicine Man: The Forgotten Museum of Henry Wellcome. London: British Museum Press, 2003. Pp. 397. ISBN 0-7141-2794-9. £19.99 (paperback). British Journal for the History of Science, 38(2): 243–244, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028717.

Chen:2000:ITT

[Che00] Xiang Chen. Instrumental traditions and theories of light: the uses of instruments in the optical revolution, volume 9 of Science and philosophy. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6349-3. xxiii + 211 pp. LCCN QC371 .C385

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0821/00030182-d.html;
http://www.loc.gov/catdir/
enhancements/fy0821/
00030182-t.html.

Cheung:2006:OBB

[Che06] Tobias Cheung. From the organism of a body to the body of an organism: occurrence and meaning of the word ‘organism’ from the seventeenth to the nineteenth centuries. British Journal for the History of Science, 39(3): 319–339, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Chimisso:2001:BRB

[Chi01a] Cristina Chimisso. Book review: Bernadette Bensaude-Vincent, L’Opinion publique et la science: à chacun son ignorance. Paris: Sanofi-Synthélabo, 2000. Pp. 239. ISBN 2-84324-141-3. FF 94.00. British Journal for the History of Science, 34(1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028045.

Chimisso:2001:BRN

[Chi01b] Cristina Chimisso. Book review: Nicole Hulin (ed.), Physique et humanités scientifiques: autour de la réforme de l’enseignement de 1902. Etudes et documents. Villeneuve d’Ascq: Presses universitaires du Septentrion, 2000. Pp. 339. ISBN 2-85939-624-1. FF 170.00, Euro 24.92 (paperback). British Journal for the History of Science, 34(3):341– 373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028107.

Chilvers:2002:BM

[Chi02a] C. A. J. Chilvers. Back matter. British Journal for the History of Science, 35(2):254, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028210.

Chilvers:2002:BRR

[Chi02b] C. A. J. Chilvers. Book review: Raymond G. Stokes, Constructing Socialism: Technology and Change in East Germany, 1945– 1990. Baltimore: Johns Hopkins University Press, 2000. Pp. xi + 260. ISBN 0-8018-6391-0. £33.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028182.

Chisholm:2002:SSC

[Chi02c] Monty Chisholm. Such silver currents: the story of William and Lucy Clifford, 1845–1929. Lutterworth Press, Cambridge, UK, 2002. ISBN 0-7188-3017-2 (paperback). ix + 198 pp. LCCN QA29.C56 C45 2002.

Chilvers:2003:DSM

[Chi03a] C. A. J. Chilvers. The dilemmas of seditious men: the Crowther– Hessen correspondence in the 1930s. British Journal for the History of Science, 36(4):417– 435, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028214.

Chimisso:2003:BRB

[Chi03b] Cristina Chimisso. Book review: Bernadette Bensaude-Vincent and Christine Blondel (eds.), Des Savants face à l’occulte 1870–1940. Sciences et société. Paris: Editions la découverte, 2002. Pp. 233. ISBN 2-7071-3616-6. EUR17.50 (paperback). British Journal for the History of Science, 36(1): 87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028300.

Chimisso:2003:BRN

[Chi03c] Cristina Chimisso. Book review:

Nicole Hulin, Les Femmes et l’enseignement scientifique. Science, histoire, société. Paris:

Presses universitaires de France, 2002. Pp. xi + 227. ISBN 2-13-052659-4. EUR23.50 (paperback). British Journal for the History of Science, 36(3):373– 374, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028166.

Chilvers:2004:BRV

[Chi04a] C. A. J. Chilvers. Book review: Vadim J. Birstein, The Perversion of Knowledge: The True Story of Soviet Science. Boulder: Westview Press, 2001. Pp. xx + 492. ISBN 0-8133-3907-3. $32.50 (hardback). British Journal for the History of Science, 37(2):229–230, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028356.

Chimisso:2004:BRH

[Chi04b] Cristina Chimisso. Book review: Hélène Gispert (ed.), ‘Par la Science, pour la patrie’: L’Association française pour l’avancement des sciences (1872–1914), un projet politique pour une société savante. Collection Carnot. Rennes: Presses Universitaires de Rennes, 2002. Pp. 372. ISBN 2-86847-681-3. EUR23.00 (paperback). British Journal for the History of Science, 37(2):225–227, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028353.

Chilvers:2006:BRG

[Chi06] C. A. J. Chilvers. Book review:

Garland E. Allen and Roy M. Macleod (eds.), Science, History and Social Activism: A Tribute to Everett Mendelsohn. Boston Studies in the Philosophy of Science, 228. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. x + 388. ISBN 1-4020-0495-8. £85.00, $124.00 (hardback). British Journal for the History of Science, 39(1): 148–149, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028572.

Chiang:2008:BRD

[Chi08] Howard Hsueh-Hao Chiang.

Book review: Diana Jeater, Law, Language, and Science: The Invention of the ‘Native Mind’ in Southern Rhodesia, 1890–1930. Portsmouth, NH: Heinemann, 2007. ISBN 978-0-325-07108-4. Pp. xxii + 274. £54.95, $94.95 (hardback). British Journal for the History of Science, 41(3): 453–455, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Christie:2005:BRJ

[Chr05] J. R. R. Christie. Book review:

John V. Pickstone, Ways of Knowing: A New History of Science, Technology and Medicine. Manchester: Manchester University Press, 2000; Chicago and London: University of Chicago Press, 2001. Pp. xii + 273. ISBN 0-226-66795-2. £14.00, $27.50 (paperback). British Journal for the History of Science, 38(3): 350–351, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028674.

Chung:2002:BRD

[Chu02] Juliette Chung. Book review: David Wright, Translating Science: The Transmission of Western Chemistry into Late Imperial China, 1840–1900. Sinica Leidensia, 48. Leiden: Brill, 2000. Pp. xxvi + 558. ISBN 90-04-11776-8. No price given (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

cIhsanoglu:2004:STL

[çIh04] Ekmeleddin çIhsanoğlu. Science, technology, and learning in the Ottoman Empire: Western influence, local institutions, and the transfer of knowledge. Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, England, 2004. ISBN 0-86078-924-1. xiv + 338 pp. LCCN Q127.I742 I47 2004.

Creager:2002:AHB

[CJ02] Angela N. H. Creager and William C. Jordan, editors. The animal-human boundary: historical perspectives. Studies in comparative history, 1539-4905. University of Rochester Press, Rochester, NY, USA, 2002. ISBN 1-58046-120-4. xviii + 342 pp. LCCN QL85 .A5143

2002. URL http://www.loc.
gov/catdir/toc/fy035/
2002072458.html.

Campbell-Kelly:2002:BRJ

[CK02] Martin Campbell-Kelly. Book review: Jon Agar, Turing and the Universal Machine: The Making of the Modern Computer. Revolutions in Science. Duxford: Icon Books, 2001. Pp. iv + 153. ISBN 1-84046-250-7. £5.99, $9.95 (paperback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Conway:1999:EAF

[CKM99] Jill K. Conway, Kenneth Keniston, and Leo Marx, editors. Earth, air, fire, water: humanistic studies of the environment. University of Massachusetts Press, Amherst, MA, USA, 1999. ISBN 1-55849-220-8 (cloth), 1-55849-221-6 (paperback). ix + 349 pp. LCCN GE140 .E18 1999.

Clark:2001:BRV

[Cla01] John F. M. Clark. Book review: Vladimir Jankovic, Reading the Skies: A Cultural History of the Weather, 1650–1820. Manchester: Manchester University Press, 2000. Pp. xiv + 272.

ISBN 0-7190-5989-5. £16.99 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028379.

Clark:2004:BRB

[Cla04a] J. F. M. Clark. Book review: Berit Pedersen (ed.), A Guide to the Archives of the Royal Entomological Society. Ashgate: Aldershot, 2002. Pp. x + 198. ISBN 0-7546-0106-4. £42.50 (hardback). British Journal for the History of Science, 37(4):476–477, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028653.

Clayton:2004:BRJ

[Cla04b] Leigh Clayton. Book review: Jill Ker Conway, Kenneth Keniston and Leo Marx (eds.), Earth, Air, Fire, Water: Humanistic Studies of the Environment. Amherst: University of Massachusetts, 1999. Pp. ix + 349. ISBN 1-55849-221-6. £17.00 (paperback). British Journal for the History of Science, 37(4): 490–491, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028665.

Clark:2005:BRJ

[Cla05] J. F. M. Clark. Book review: John Thackray and Bob Press, The Natural History Museum: Nature’s Treasurehouse. London: Natural History Museum, 2001. Pp. 144. ISBN 0-565-09164-6. £11.00 (paperback). British Journal for the History of Science, 38(1): 114–115, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028595.

Clark:2006:BRS

[Cla06] D. J. Clark. Book review:

Slava Gerovitch, From Newspeak to Cyberspeak: A History of Soviet Cybernetics. Cambridge, MA and London: MIT Press, 2002. pp. xiv + 369. ISBN 0-262-07232-7. £25.95 (hardback). British Journal for the History of Science, 39(1):146–148, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028571.

Clark:2007:BRW

[Cla07a]  J.  F.  M.  Clark.

Book review: William M. Taylor, The Vital Landscape: Nature and the Built Environment in Nineteenth-Century Britain. Aldershot: Ashgate, 2004. Pp. xvii + 252. ISBN 0-7546-3069-2. £45.00 (hardback). British Journal for the History of Science, 40(3):446–447, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500755.

Clark:2007:VEV

[Cla07b] Stuart Clark. Vanities of the eye: vision in early modern European culture. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-925013-8 (hardcover). xi + 415 pp. LCCN B846 .C53 2007.

Clark:2008:GGI

[Cla08] Constance Areson Clark. God – or gorilla: images of evolution in the jazz age. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8825-5 (hardcover). xvi + 289 pp. LCCN GN281 .C513

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0805/2007041258-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0805/
                                        2007041258-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip082/2007041258.html.

Clarke:2009:BRJ

[Cla09] Ellen Clarke. Book review: Jamie Elwick, Styles of Reasoning in the British Life Sciences: Shared Assumptions, 1820–1858. London: Pickering & Chatto, 2007. Pp. x + 233. ISBN 978-1-85196-920-3. £60.00, $99.00 (hardback). British Journal for the History of Science, 42(1): 143–145, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Clericuzio:2000:EPC

[Cle00] Antonio Clericuzio. Elements, principles, and corpuscles: a study of atomism and chemistry in the Seventeenth Century, volume 171 of Archives internationales d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6782-0. xi + 223 pp. LCCN QD18.E85 C57

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0821/00067519-d.html;
http://www.loc.gov/catdir/
enhancements/fy0821/
00067519-t.html.

Clucas:2006:JDI

[Clu06] Stephen Clucas, editor. John Dee: interdisciplinary studies in English Renaissance thought, volume 193 of Archives internationales d’ histoire des idées = International archives of the history of ideas. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-4245-0. xvii + 366 pp. LCCN BF1598.D5 J64

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0904/2007451201-d.html;
http://www.loc.gov/catdir/
toc/fy0715/2007451201.html.

Creager:2007:SLM

[CLW07] Angela N. H. Creager, Elizabeth Lunbeck, and M. Norton Wise, editors. Science without laws: model systems, cases, exemplary narratives. Science and cultural theory. Duke University Press, Durham, NC, USA, 2007. ISBN 0-8223-4046-1 (cloth) 0-8223-4068-2 (paperback). 287 pp. LCCN QH324.8 .S25 2007. URL http://www.

loc.gov/catdir/toc/
ecip0710/2007006303.html.

Coen:2007:VAU

[Coe07] Deborah R. Coen. Vienna in the age of uncertainty: science, liberalism, and private life. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-11172-5 (cloth). xi + 380 pp. LCCN DB844.E98 C64

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2007003723-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2007003723-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0710/2007003723.html.

Coen:2009:BRE

[Coe09] Deborah R. Coen. Book reviews: Elizabeth R. Neswald, Thermodynamik als kultureller Kampfplatz: Zur Faszinationsgeschichte der Entropie, 1850– 1915. Freiburg: Rombach Verlag, 2006. Pp. 475. ISBN 978-3-7930-9448-7. EUR48.00 (paperback). Helge S. Kragh, Entropic Creation: Religious Contexts of Thermodynamics and Cosmology. Aldershot: Ashgate, 2008. pp. 272. ISBN 978-0-7546-6414-7. £60.00 (hardback). British Journal for the History of Science, 42(3):467–468, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cohen:2002:FMF

[Coh02] Claudine Cohen. The fate of the mammoth: fossils, myth, and history. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-11292-6 (cloth). xxxiv + 297 pp. LCCN QE882.P8 C6413 2002. URL

http://www.loc.gov/catdir/
bios/uchi051/2001048011.
html;   http://www.loc.gov/
catdir/description/uchi051/
2001048011.html;   http://
www.loc.gov/catdir/toc/
uchi051/2001048011.html.

Cohen:2007:EGP

[Coh07] Daniel J. (Daniel Jared) Cohen.

Equations from God: pure mathematics and Victorian faith. Johns Hopkins studies in the history of mathematics. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8553-1 (hardcover). x + 242 pp. LCCN BL265.M3 C64

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006019751-b.html;
http://www.loc.gov/catdir/
enhancements/fy0703/
2006019751-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0616/2006019751.html.

Cohen:2009:BRM

[Coh09]  Claudine Cohen.  Book re-

view: Martin J. S. Rudwick, Worlds before Adam: The Reconstruction of Geohistory in the Age of Reform. Chicago and London: University of Chicago Press, 2008. Pp. xxii + 614. ISBN 978-0-226-73128-5. £23.50 (hardback). British Journal for the History of Science, 42(3): 434–435, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Collins:2000:BRL

[Col00] Alan Collins. Book review: Leila Zenderland, Measuring Minds: Henry Herbert Goddard and the Origins of American Intelligence Testing. Cambridge: Cambridge University Press, 1998. Pp. vii + 466. ISBN 0-521-44373-3. £45.00, $64.95. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027929.

Cole:2001:SIH

[Col01a] Simon A. Cole. Suspect identities: a history of fingerprinting and criminal identification. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00455-8. 369 pp. LCCN HV6074 .C557 2001.

Collins:2001:BRJ

[Col01b] Alan Collins. Book review: Jack D. Pressman, Last Resort: Psychosurgery and the Limits of Medicine. Studies in the History of Medicine. Cambridge: Cambridge University Press, 1998. Pp. xv + 555. ISBN 0-521-35371-8. £40.00, $49.95. British Journal for the History of Science, 34(1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028046.

Collins:2002:BRW

[Col02]  Harry  Collins.

Book review: What the Tortoise said to Achilles Donald Mackenzie, Mechanizing Proof: Computing, Risk, and Trust. Inside Technology. Cambridge, MA and London: MIT Press, 2001. Pp. xi + 427. ISBN 0-262-13393-8. £30.95 (hardback). British Journal for the History of Science, 35(4):469–474, December 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028277.

Collins:2004:GSS

[Col04] H. M. (Harry M.) Collins. Gravity’s shadow: the search for gravitational waves. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-11377-9 (hardcover), 0-226-11378-7 (paperback). xxiii + 870 pp. LCCN QC179 .C65

2004. URL http://www.loc.
gov/catdir/description/
uchi052/2003023823.html;
http://www.loc.gov/catdir/
enhancements/fy0614/
2003023823-b.html;   http:/

                                        /www.loc.gov/catdir/toc/
                                        ecip0410/2003023823.html.

Comfort:2007:PBB

[Com07] Nathaniel C. Comfort, editor. The panda’s black box: opening up the intelligent design controversy. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8599-X (hardcover). xv + 165 pp. LCCN BS652 .P35

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0703/2006034764-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0703/
                                        2006034764-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip073/2006034764.html.

Coopey:2002:BM

[Coo02a] Richard Coopey. Back matter. British Journal for the History of Science, 35(4):495– 508, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028288.

Coopey:2002:BRR

[Coo02b] Richard Coopey. Book reviews: Robert Bud and Philip Gummett (eds.), Cold War, Hot Science: Applied Research in Britain’s Defence Laboratories 1945–1990. Studies in the History of Science, Technology and Medicine, 7. Amsterdam: Harwood Academic/Science Museum, 1999. Pp. xix + 426. ISBN 90-5702-481-0. £42.00, $62.00 (hardback); London: Science Museum, 2002. Pp. xix+426. ISBN 1-900747-47-2. £34.95 (paperback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cookson:2003:BRB

[Coo03] Gillian Cookson. Book review:

Bernard Cronin, Technology, Industrial Conflict and the Development of Technical Education in 19th-Century England. Modern Economic and Social History Series, 11. Aldershot: Ashgate, 2001. Pp. xiv + 301. ISBN 0-7546-0313-X. £55.00 (hardback). British Journal for the History of Science, 36(2):252–253, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028249.

Cookson:2004:BRR

[Coo04] Gillian Cookson. Book review:

Richard L. Hills, Life and Inventions of Richard Roberts, 1789– 1864. Landmark Collector’s Library. Ashbourne: Landmark Publishing, 2002. Pp. 255. ISBN 1-84306-027-2. £29.95 (hardback). British Journal for the History of Science, 37(3):352–353, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028133;   http://www.
jstor.org/stable/4028435.

Cook:2005:BRJ

[Coo05a] Alexandra Cook. Book review: Jean-Daniel Candaux and Jean-Marc Drouin (eds.), Augustin-Pyramus de Candolle: Mémoires et souvenirs (1778–1841). Bibliothèque d’histoire des sciences,

5. Geneva: Georg Editeur, 2003. Pp. xv + 591. ISBN 2-8257-0832-1. EUR33.00 (paperback). British Journal for the History of Science, 38(4):482– 483, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028466.

Cook:2005:LVV

[Coo05b] Simon Cook. Late Victorian visual reasoning and Alfred Marshall’s economic science. British Journal for the History of Science, 38(2):179–195, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cookson:2006:BRK

[Coo06a] Gillian Cookson. Book review: Kenneth Silverman, Lightning Man: The Accursed Life of Samuel F. B. Morse. New York: Alfred A. Knopf, 2003. Pp. vi + 503. ISBN 0-375-40128-8. $35.00 (hardback). British Journal for the History of Science, 39(2):295–296, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028621.

Cookson:2006:BRM

[Coo06b] Gillian Cookson. Book review:

Marshall J. Bastable, Arms and the State: Sir William Armstrong and the Remaking of British Naval Power, 1854–1914. Modern Economic and Social History. Aldershot: Ashgate, 2004. Pp. xii + 300. ISBN 0-7546-3404-3. £49.99 (hardback). British Journal for the History of Science, 39(4):622, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028540.

Cooper:2006:BRH

[Coo06c] Martin Cooper. Book review:

Helmuth Trischler and Stefan Zeilinger (eds.), Tackling Transport. Artefects Series: Studies in the History of Science and Technology. London: Science Museum, 2003. Pp. vi + 186. ISBN 1-900747-53-7. £21.95 (paperback). British Journal for the History of Science, 39 (2):308–309, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028632.

Cook:2007:BRD

[Coo07a] Alexandra Cook. Book reviews: Dorinda Outram, The Enlightenment. Second edition. New Approaches to European History. Cambridge: Cambridge University Press, 2005. Pp. xiii + 165. ISBN 0-521-546681-8. £14.99, $24.99 (paperback).

William E. Burns, Science in the Enlightenment: An Encyclopedia. Santa Barbara: ABC-CLIO, 2003. Pp. xcviii+165. ISBN 1-57607-887-6. $65.00 (hardback). British Journal for the History of Science, 40(1):137–139, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500700.

Cook:2007:MEC

[Coo07b] Harold John Cook. Matters of exchange: commerce, medicine, and science in the Dutch Golden Age. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-14321-4. xiv + 562 pp. LCCN Q127.N2 C66

2007. URL http://www.loc.
gov/catdir/toc/ecip0619/
2006026973.html.

Cooper:2007:IIL

[Coo07c] Alix Cooper. Inventing the indigenous: local knowledge and natural history in early modern Europe. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87087-9 (hardcover). xiii + 218 pp. LCCN GR880 .C67

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006037259-d.html;
http://www.loc.gov/catdir/
enhancements/fy0703/
2006037259-t.html;
http://www.loc.gov/catdir/
enhancements/fy0729/
2006037259-b.html.

Copeland:2004:ETS

[Cop04] B. Jack Copeland, editor. The Essential Turing: Seminal Writings in Computing, Logic, Philosophy, Artificial Intelligence, and Artificial Life, plus The Secrets of Enigma. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-825079-7 (hardcover), 0-19-825080-0 (paperback). viii + 613 pp. LCCN QA29.T8 E77

2004. URL ftp://uiarchive.
cso.uiuc.edu/pub/etext/
gutenberg/;
http://www.loc.gov/catdir/
toc/fy053/2004275594.html.

Cormack:2000:BRM

[Cor00] Lesley B. Cormack. Book review: Matthew H. Edney, Mapping an Empire: The Geographical Construction of British India, 1765–1843. Chicago: University of Chicago Press, 1997. Pp. xx + 480. ISBN 0-226-18487-0. £27.95, $35.00. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027958.

Corry:2004:DHA

[Cor04a] Leo Corry. David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2004. ISBN 1-4020-2777-X (hardcover), 1-4020-2778-8 (ebook). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 513 pp. LCCN QA29.H5 C67

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0823/2006280837-d.html;
http://www.loc.gov/catdir/
enhancements/fy0823/
2006280837-t.html.

Corsi:2004:BRP

[Cor04b] Pietro Corsi. Book review: Paola Govoni, Un pubblico per la scienza: La divulgazione scientifica nell’Italia in formazione. Studi storici Carocci, 27. Rome: Carocci editore, 2002. Pp. 351. ISBN 88-430-2321-7. EUR23.50 (paperback). British Journal for the History of Science, 37 (2):216–217, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028345.

Corsi:2005:BRJ

[Cor05] Pietro Corsi. Book review: Jean-Luc Chappey, La Société des observateurs de l’homme (1799–1804): Des Anthropologues au temps de Bonaparte. Bibliothèque d’histoire révolutionnaire, 5. Paris: Société des études robespierristes, 2002. Pp. 573. ISBN 2-908327-45-7. EUR46.00 (paperback). British Journal for the History of Science, 38(2):228–230, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028705.

Corsi:2006:BRC

[Cor06] Pietro Corsi. Book review:

Charles Coulston Gillispie, Science and Polity in France: The Revolutionary and Napoleonic Years. Princeton: Princeton University Press, 2004. Pp. ix + 751. ISBN 0-691-11541-9. £52.95 (hardback). British Journal for the History of Science, 39(3): 451–452, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028496.

Cossons:2000:PIA

[Cos00] Neil Cossons, editor. Perspectives on industrial archaeology. Science Museum, London, UK, 2000. ISBN 1-900747-31-6 (paperback). 176 pp. LCCN T26.G7 P48 2000.

Coulls:2004:BM

[Cou04] Anthony Coulls. Back matter.

British Journal for the History of Science, 37(2):244, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028367.

Cowen:2001:PRL

[Cow01]  David L. Cowen.  Pharma-

copoeias and related literature in Britain and America, 1618– 1847, volume CS700 of Variorum collected studies series. Ashgate/ Variorum, Aldershot, Hampshire, England, 2001. ISBN 0-86078-842-3. viii + 296 pp. LCCN RS141.3 .C675 2001.

Collins:2005:DGH

[CP05] H. M. (Harry M.) Collins and T. J. (Trevor J.) Pinch. Dr. Golem: how to think about medicine. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-11366-3 (cloth). xii + 246 pp. LCCN RC81 .C695

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0621/2005001621-b.html;
http://www.loc.gov/catdir/
enhancements/fy0621/
2005001621-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip056/2005001621.html.

Crelinsten:2006:EJR

[Cre06] Jeffrey Crelinsten. Einstein’s jury: the race to test relativity. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12310-1. xxix + 397 pp. LCCN QC173.585 .C74

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0654/2005032681-d.html;
http://www.loc.gov/catdir/
enhancements/fy0734/
2005032681-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip063/2005032681.html.

Crocker:2001:RRN

[Cro01a] Robert Crocker, editor. Religion, reason, and nature in early modern Europe, volume 180 of Archives internationales d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0047-2. LCCN BL182 .R45

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0823/2001045958-d.html;
http://www.loc.gov/catdir/
enhancements/fy0823/
2001045958-t.html.

Cronin:2001:TIC

[Cro01b] B. P. (Bernard P.) Cronin. Technology, industrial conflict, and the development of technical education in 19th-Century England. Modern economic and social history series. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-7546-0313-X. xiii + 301 pp. LCCN T173.8 .C75 2001.

Crosland:2001:PSA

[Cro01c] Maurice Crosland. Popular science and the arts: challenges to cultural authority in France under the Second Empire. British Journal for the History of Science, 34(3):301–322, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028100.

Crosland:2003:RSC

[Cro03]  Maurice Crosland.  Re-

search schools of chemistry from Lavoisier to Wurtz. British Journal for the History of Science, 36(3):333–361, September 2003.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028158.

Croarken:2004:BRD

[Cro04] Mary Croarken. Book review: Doron Swade, The Cogwheel Brain: Charles Babbage and the Quest to Build the First Computer. London: Little, Brown and Company, 2000. Pp. x + 342. ISBN 0-316-64847-7. £14.99 (hardback). British Journal for the History of Science, 37(3): 351–352, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028132;
                                        http://www.jstor.org/
                                        stable/4028434.

Croarken:2005:BRG

[Cro05a] Mary Croarken. Book review: Georgina Ferry, A Computer Called LEO: Lyons Teashops and the World’s First Office Computer. London: Fourth Estate, 2003. Pp. xi + 221. ISBN 1-84115-185-8. £15.99 (hardback). British Journal for the History of Science, 38(4):491–492, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028473.

Crosland:2005:BRD

[Cro05b] Maurice Crosland. Book review: Danielle Fauque (ed.), Lavoisier et la naissance de la chimie moderne. Paris: Vuibert, 2003. Pp. 233. ISBN 2-7117-5353-0. No price given (paperback). British Journal for the History of Science, 38(3):371, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028689.

Crosland:2007:SIP

[Cro07]  Maurice P. Crosland.  Sci-

entific institutions and practice in France and Britain, c.1700– c.1870, volume 870 of Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2007. ISBN 0-7546-5913-5. 286 pp. LCCN Q180.F7

C76 2007. URL http://www.
loc.gov/catdir/toc/ecip072/
2006034000.html.

Cantor:2006:JTC

[CS06] G. N. Cantor and Marc Swetlitz, editors. Jewish tradition and the challenge of Darwinism. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-09276-3 (cloth) 0-226-09277-1 (paperback). xii + 260 pp. LCCN BM538.E8 J49

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0666/2006017533-b.html;
http://www.loc.gov/catdir/
enhancements/fy0666/
2006017533-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0614/2006017533.html.

Connor:2004:WMS

[CSML04] R. D. Connor, A. D. C. (Allen David Cumming) Simpson, and A. D. Morrison-Low. Weights and measures in Scotland: a European perspective. National Museums of Scotland, Edinburgh, Scotland, 2004. ISBN 1-901663-88-4. xvi + 842 pp. LCCN QC89.G8 C66 2004.

Cueto:2007:CWD

[Cue07] Marcos Cueto. Cold war, deadly fevers: malaria eradication in Mexico, 1955–1975. Woodrow Wilson Center Press, Washington, DC, USA, 2007. ISBN 0-8018-8645-7 (cloth). xvi + 264 pp. LCCN RC162.M6 C84

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0731/2007004959-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0731/
                                        2007004959-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0710/2007004959.html.

Cuomo:2000:PAM

[Cuo00] S. (Serafina) Cuomo. Pappus of Alexandria and the mathematics of late antiquity. Cambridge classical studies. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-64211-6. ix + 234 pp. LCCN QA22 .C86 2000.

                                        URL  http://www.loc.gov/
                                        catdir/description/cam029/
                                        99012317.html;   http://
                                        www.loc.gov/catdir/samples/
                                        cam032/99012317.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam023/99012317.html.

Cuomo:2002:BRM

[Cuo02] Serafina Cuomo. Book review: Michael N. Fried and Sabetai Unguru, Apollonius of Perga’s Conica : Text, Context, Subtext. Mnemosyne, supplement 222. Leiden: Brill, 2001. Pp. xii + 499. ISBN 90-04-11977-9. ???105.00, $122.00 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Cuomo:2007:TCG

[Cuo07] S. (Serafina) Cuomo. Technology and culture in Greek and Roman antiquity. Key themes in ancient history. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-81073-6 (hardcover), 0-521-00903-0 (paperback), (paperback). xi + 212 pp. LCCN T16 .C78

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0805/2007008985-b.html;
http://www.loc.gov/catdir/
enhancements/fy0805/
2007008985-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0712/2007008985.html.

Chazdon:2002:FTF

[CW02] Robin Lee Chazdon and T. C.

(Timothy Charles) Whitmore, editors. Foundations of tropical forest biology: classic papers with commentaries. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-10224-6 (cloth), 0-226-10225-4 (paper). xv + 862 pp. LCCN QH84.5 .F68 2002. URL http://www.

loc.gov/catdir/description/

                                        uchi051/2001005769.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0612/
                                        2001005769-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy031/2001005769.html.

DAmbrosio:2003:BRH

[D’A03] Ubiratan D’Ambrosio. Book review: Helaine Selin (ed.), Astronomy Across Cultures: The History of Non-Western Astronomy. Science Across Cultures: The History of Non-Western Science, 1. Dordrecht: Kluwer Academic Publishers, 2000. Pp. xxiii + 665. ISBN 0-7923-6363-9. £215.00, $345.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028301.

Danziger:2001:BRJ

[Dan01] Kurt Danziger. Book review: L. S. Jacyna, Lost Words: Narratives of Language and the Brain, 1825–1926. Princeton and Oxford: Princeton University Press, 2000. Pp. ix + 241. ISBN 0-691-00413-7. £28.50, $45.00. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028011.

Darby:2005:STD

[Dar05] Robert (Robert J. L.) Darby. A surgical temptation: the demonization of the foreskin and the rise of circumcision in Britain. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-13645-0. xi + 374 pp. LCCN GN484 .D37 2005. URL http:/

/www.loc.gov/catdir/bios/
uchi051/2004027996.html;
http://www.loc.gov/catdir/
enhancements/fy0617/
2004027996-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip054/2004027996.html.

Daston:2000:BSO

[Das00] Lorraine Daston, editor. Biographies of scientific objects. University of Chicago Press, Chicago, IL, USA, 2000. ISBN 0-226-13670-1 (cloth), 0-226-13672-8 (paperback). ix + 307 pp. LCCN Q175.32.K45 B56

2000. URL http://www.loc.
gov/catdir/description/
uchi051/99043745.html;
http://www.loc.gov/catdir/
toc/uchi052/99043745.html.

Daston:2004:TTO

[Das04] Lorraine Daston, editor. Things that talk: object lessons from art and science. Zone Books, New York, NY, USA, 2004. ISBN 1-890951-43-9 (cloth). 447 + 8 pp. LCCN N72.S3 T55

2004. URL http://www.h-net.
org/review/hrev-a0e3o8-aa;
http://www.h-net.org/
review/hrev-a0e3v1-aa.

Daston:2007:BRW

[Das07] Lorraine Daston. Book review:

W. F. Bynum and Roy Porter (eds.), Oxford Dictionary of Scientific Quotations. Oxford: Oxford University Press, 2005. Pp. xvii + 712. ISBN 0-19-858409-1. £30.00, $50.00 (hardback). British Journal for the History of Science, 40(2):276–277, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500720.

Daston:2009:BRK

[Das09] Lorraine Daston. Book review: Kapil Raj, Relocating Modern Science: Circulation and the Construction of Knowledge in South Asia and Europe, 1650–1900. Houndmills and New York: Palgrave Macmillan, 2007. Pp. xiv + 285. ISBN 978-0-230-50708. £53.00 (hardback). British Journal for the History of Science, 42(2):298–299, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Daunton:2005:OKV

[Dau05] M. J. (Martin J.) Daunton, editor. The organisation of knowledge in Victorian Britain. [British Academy centenary monographs]. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-726326-7 (hardcover). vii + 424 pp. LCCN DA533 .O65

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0638/2005296111-d.html;
http://www.loc.gov/catdir/

enhancements/fy0638/
2005296111-t.html.

Davies:2009:SRN

[Dav09a] E. B. Davies. Some reflections on Newton’s principia. British Journal for the History of Science, 42(2):211–224, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Davis:2009:RAE

[Dav09b] Sophia Davis. Raising the aerocompass in early twentiethcentury Britain. British Journal for the History of Science, 42(1):73–94, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawbarn:2000:BRA

[Daw00] Frances Dawbarn. Book review: Andrea Carlino (Translated by John Tedeschi and Anne C. Tedeschi). Books of the Body: Anatomical Ritual and Renaissance Learning. Chicago: University of Chicago Press, 1999. British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027930.

Dawson:2006:BRB

[Daw06] Gowan Dawson. Book review:

Bert Bender, Evolution and ‘the Sex Problem’: American Narratives during the Eclipse of Darwinism. Kent, OH and London: Kent State University Press, 2004. Pp. xvi + 389. ISBN 0-87338-809-7. $59.95 (hardback). British Journal for the History of Science, 39(3): 463–464, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028503.

Dawson:2007:BRJ

[Daw07a] Gowan Dawson. Book review: Jonathan Smith, Charles Darwin and Victorian Visual Culture. Cambridge: Cambridge University Press, 2006. Pp. xxiii + 349. ISBN 0-521-85690-6. £60.00 (hardback). British Journal for the History of Science, 40(4): 618–619, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dawson:2007:DLV

[Daw07b] Gowan Dawson. Darwin, literature and Victorian respectability, volume 57 of Cambridge studies in nineteenth-century literature and culture. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87249-9. xii + 282 pp. LCCN PR468.S34 D39

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0729/2007277752-b.html;
http://www.loc.gov/catdir/
enhancements/fy0729/
2007277752-d.html;
http://www.loc.gov/catdir/
enhancements/fy0729/
2007277752-t.html.

deChadarevian:2008:BRR

[dC08] Soraya de Chadarevian. Book review: Ronald E. Doel and Thomas Söderqvist (eds.) The Historiography of Contemporary Science, Technology, and Medicine: Writing Recent Science. London and New York: Routledge, 2006. Pp. xv + 313. ISBN 0-415-39142-3. £80.00 (paperback). British Journal for the History of Science, 41(1):127– 129, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

deChadarevian:2004:MTD

[dCH04] Soraya de Chadarevian and Nick Hopwood, editors. Models: the Third Dimension of Science. Writing science. Stanford University Press, Stanford, CA, USA, 2004. ISBN 0-8047-3971-4 (hardcover), 0-8047-3972-2 (paperback). xvi + 464 pp. LCCN Q223 .M58 2004. URL http:/

/www.e-streams.com/es0803/
es0803_4012.html;
http://www.loc.gov/catdir/
enhancements/fy0710/
2003025168-b.html;
http://www.loc.gov/catdir/
enhancements/fy0710/
2003025168-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0411/2003025168.html.

Delbourgo:2008:SEA

[DD08] James Delbourgo and Nicholas Dew, editors. Science and empire in the Atlantic world. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2008. ISBN 0-415-96126-2 (hardcover), 0-415-96127-0 (paperback), 0-203-93384-2 (e-book). xiv + 365 pp. LCCN Q175.52.A43 S32

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip0716/
                                        2007016373.html.

DeRenzi:2007:BRG

[De 07] Silvia De Renzi. Book review: Gianna Pomata and Nancy G. Siraisi (eds.), Historia: Empiricism and Erudition in Early Modern Europe. Cambridge, MA and London: The MIT Press, 2005. Pp. viii + 490. ISBN 0-262-16229-6. £32.95 (hardback). British Journal for the History of Science, 40 (2):282–284, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500724.

Dear:2001:RSE

[Dea01] Peter Dear. Revolutionising the sciences: European knowledge and its ambitions, 1500– 1700. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2001. ISBN 0-333-71573-X, 0-333-71574-8 (paperback). viii + 208 pp. LCCN Q125.2 .D339 2001.

Dear:2002:BRB

[Dea02] Peter Dear. Book review: Boyle in the bag! Robert Boyle, The Works of Robert Boyle, edited by Michael Hunter and Edward B. Davis, 14 vols. London: Pickering & Chatto, 1999–2000. ISBN 1-85196-109-7. £995.00, $1,495.00. British Journal for the History of Science, 35(3): 335–340, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028127.

Dean:2005:BRR

[Dea05] Sheila Ann Dean. Book review:

Richard Keynes, Fossils, Finches and Fuegians: Charles Darwin’s Adventures and Discoveries on the Beagle, 1832–1836. London: HarperCollins, 2002. Pp. xix + 428. ISBN 0-00-710189-9. £25.00 (hardback). British Journal for the History of Science, 38(1): 112–113, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028593.

Dear:2006:BRN

[Dea06a]  Peter Dear.  Book review:

Nicholas Hammond (ed.), The Cambridge Companion to Pascal. Cambridge: Cambridge University Press, 2003. Pp. xvi + 287. ISBN 0-521-00611-2. £15.95, $23.00 (paperback). British Journal for the History of Science, 39(1): 131–133, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028559.

Dear:2006:INH

[Dea06b] Peter Dear. The intelligibility of nature: how science makes sense of the world. Science.culture.

University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-13948-4. xii + 242 pp. LCCN Q175.32.R45 D43

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0702/2005026032-b.html;
http://www.loc.gov/catdir/
enhancements/fy0702/
2005026032-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0518/2005026032.html.

Dear:2007:BRM

[Dea07] Peter Dear. Book review: Marcus Hellyer, Catholic Physics: Jesuit Natural Philosophy in Early Modern Germany. Notre Dame, IN: University of Notre Dame Press, 2005. Pp. xii + 337. ISBN 0-268-03071-5. $50.00 (hardback). British Journal for the History of Science, 40(1): 135–137, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500699.

Dear:2008:BRD

[Dea08]  Peter Dear.  Book re-

views: Domenico Bertoloni Meli, Thinking with Objects: The Transformation of Mechanics in the Seventeenth Century. Baltimore: Johns Hopkins University Press, 2006. Pp. xii + 389. ISBN 0-8018-8426-8. £46.50 (hardback). ISBN 0-8018-8427-6. £20.00 (paperback). British Journal for the History of Science, 41(2):296–297, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dear:2009:BRC

[Dea09] Peter Dear. Book review: Claus Zittel, Gisela Engel, Romano Nanni and Nicole C. Karafyllis eds., Philosophies of Technology: Francis Bacon and His Contemporaries. Intersections: Yearbook for Early Modern Studies series vol. 11. Leiden: Brill, 2008. 2 vols. Pp. xxix + 577. ISBN 978-9-0041-70506. EUR149.00 (hardback). British Journal for the History of Science, 42(4):616– 617, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Debus:1984:OWP

[Deb84] Allen G. Debus. Obituary: Walter Pagel, M.D.: 12 November 1898–25 March 1983. British Journal for the History of Science, 17(1):83–84, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4026020.

Debus:2006:CPE

[Deb06] Allen G. Debus. The chemical promise: experiment and mysticism in the chemical philosophy, 1550–1800: selected essays of Allen G. Debus. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2006. ISBN 0-88135-296-9. xxv + 548 pp. LCCN QD14 .D44 2006.

Delbourgo:2006:BRL

[Del06a] James Delbourgo. Book review: Londa Schiebinger and Claudia Swan (eds.), Colonial Botany: Science, Commerce, and Politics in the Early Modern World. Philadelphia: University of Pennsylvania Press, 2004. Pp. vi + 346. ISBN 0-8122-3827-3. £36.00, $55.00 (hardback). British Journal for the History of Science, 39(4):590– 591, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028518.

Delbourgo:2006:MAS

[Del06b] James Delbourgo. A most amazing scene of wonders: electricity and enlightenment in early America. Harvard University Press, Cambridge, MA, USA, 2006. ISBN 0-674-02299-8. 367 pp. LCCN E162 .D395 2006.

Delbourgo:2008:BRC

[Del08] James Delbourgo. Book review: Cynthia Sundberg Wall, The Prose of Things: Transformations of Description in the Eighteenth Century. Chicago and London: The University of Chicago Press, 2006. Pp. xiii + 316. ISBN 0-226-87158-4. $35.00 (hardback). British Journal for the History of Science, 41(1): 140–141, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Delbourgo:2009:BRD

[Del09] James Delbourgo. Book reviews:

Daniel Carey, Locke, Shaftesbury, and Hutcheson: Contesting Diversity in the Enlightenment and Beyond. Cambridge: Cambridge University Press, 2006. Pp. x + 260. ISBN 978-0-5214-4502-1. £51.00 (hardback). Sarah Irving, Natural Science and the Origins of the British Empire. London: Pickering and Chatto, 2008. Pp. xiii+183. ISBN 978-1-85196-889-3. £60.00 (hardback). British Journal for the History of Science, 42(3):459–461, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Depew:2007:BRA

[Dep07] David Depew. Book review: Andrea Falcon, Aristotle and the Science of Nature: Unity without Uniformity. Cambridge: Cambridge University Press, 2005. Pp. xvii + 139. ISBN 0-521-85439-3. £45.00 (hardback). British Journal for the History of Science, 40 (1):123–124, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500691.

Derry:2003:BPS

[Der03] Margaret Elsinor Derry. Bred for perfection: shorthorn cattle, collies, and arabian horses since 1800. Animals, history, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2003. ISBN 0-8018-7344-4 (hardcover). xvi + 198 pp. LCCN SF199.S56 D47 2003. URL

                                        http://www.e-streams.com/
                                        es0708/es0708_3426.html;
                                        http://www.loc.gov/catdir/
                                        bios/jhu051/2002015859.
                                        html;   http://www.loc.gov/
                                        catdir/description/jhu051/
                                        2002015859.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy043/2002015859.html.

DesChene:2001:SCM

[Des01] Dennis Des Chene. Spirits and clocks: machine and organism in Descartes. Cornell University Press, Ithaca, NY, USA, 2001. ISBN 0-8014-3764-4 (cloth). xiii + 181 pp. LCCN QP21 .D378 2001.

DeVorkin:2000:HNR

[DeV00] David H. DeVorkin. Henry Norris Russell: dean of American astronomers. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-04918-1 (cloth). xix + 499 pp. LCCN QB36.R78 D49 2000. URL http:

                                        //www.loc.gov/catdir/
                                        description/prin021/
                                        99059472.html;   http://
                                        www.loc.gov/catdir/samples/
                                        prin031/99059472.html;
                                        http://www.loc.gov/catdir/
                                        toc/prin032/99059472.html.

Dew:2001:BRP

[Dew01] Nicholas Dew. Book review: Peter N. Miller, Peiresc’s Europe: Learning and Virtue in the Seventeenth Century. New Haven and London: Yale University Press, 2000. Pp. xvi + 234. ISBN 0-300-08252-5. £30.00. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028012.

Dewsbury:2007:BRR

[Dew07] Donald A. Dewsbury. Book reviews: Richard W. Burkhardt, J R, Patterns of Behavior: Konrad Lorenz, Niko Tinbergen, and the Founding of Ethology. Chicago and London: University of Chicago Press, 2005. Pp. xii + 636. ISBN 0-226-08089-7. $80.00 (hardback). ISBN 0-226-08090-0. $29.00 (paperback). British Journal for the History of Science, 40(1):147–149, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500707.

Dolling:2003:TTR

[DGS03] Lisa M. Dolling, Arthur F. Gianelli, and Glenn N. Statile, editors. The tests of time: readings in the development of physical theory. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-09084-X, 0-691-09085-8 (paperback). xlii + 716 pp. LCCN QC125 .T47

2003. URL http://www.loc.
gov/catdir/bios/prin051/
2002074976.html;
http://www.loc.gov/catdir/

                                        description/prin031/
                                        2002074976.html.

Didi-Huberman:2003:IHC

[DHC03] Georges Didi-Huberman and J. M. (Jean Martin) Charcot. Invention of hysteria: Charcot and the photographic iconography of the Salpêtrière. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-04215-0 (hardcover). xii + 373 pp. LCCN RC532 .D5313

                                        2003. URL http://www.loc.
                                        gov/catdir/toc/fy038/
                                        2002029382.html.

Dick:2003:SOJ

[Dic03] Steven J. Dick. Sky and ocean joined: the U.S. Naval Observatory, 1830–2000. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81599-1. xiii + 609 pp. LCCN QB82.U62 U553 2003. URL http://www.

                                        loc.gov/catdir/description/
                                        cam022/2002017386.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0731/
                                        2002017386-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam021/2002017386.html.

Dietz:2006:MOS

[Die06] Bettina Dietz. Mobile objects: the space of shells in eighteenthcentury France. British Journal for the History of Science, 39(3): 363–382, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Digby:2008:BRA

[Dig08] Anne Digby. Book reviews: Akihito Suzuki, Madness at Home:

The Psychiatrist, the Patient and the Family in England, 1820–1860. Berkeley and London: University of California Press, 2006. Pp. xii + 260. ISBN 0-520-24580-6. $49.95, £32.50 (hardback). Joseph Melling and Bill Forsythe, The Politics of Madness: The State, Insanity and Society in England, 1845– 1914. Abingdon: Routledge, 2006. Pp. xviii +278. ISBN 0-415-30174-2. £75.00 (hardback). British Journal for the History of Science, 41(2):283–285, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Digrius:2009:BRC

[Dig09] Dawn M. Digrius. Book review: Constance Areson Clark, God or Gorilla: Images of Evolution in the Jazz Age. Baltimore: Johns Hopkins University Press, 2008. xviii + 289. ISBN 978-0-8018-8825-0. £23.50 (hardback). British Journal for the History of Science, 42(4):608–609, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

DiMeo:2007:BRJ

[DiM07]  Michelle DiMeo.  Book re-

view: Judith P. Zinsser (ed.), Men, Women, and the Birthing of Modern Science. De Kalb, IL: Northern Illinois University Press, 2005. Pp. viii + 215. ISBN 0-87580-340-7. $38.00 (hardback). British Journal for the History of Science, 40(4):

607–608, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Divall:2001:BRN

[Div01] Colin Divall. Book review: Neil Cossons (ed.), Perspectives on Industrial Archaeology. London: Science Museum, 2000. Pp. 176. ISBN 1-900747-31-6. £19.95 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028380.

Dixon:2000:BRP

[Dix00] Tom Dixon. Book review: Penelope Gouk, Music, Science and Natural Magic in Seventeenth-Century England. New Haven and London: Yale University Press, 1999. Pp. xii + 308. ISBN 0-300-07383-6. £30.00, $35.00. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4027959.

Dixon:2003:BRG

[Dix03a] Thomas Dixon. Book review: G. C. Bunn, A. D. Lovie and G. D. Richards (eds.), Psychology in Britain: Historical Essays and Personal Reflections. Leicester: British Psychological Society, 2001. Pp. xvi + 495. ISBN 1-85433-332-1. £26.95 (paperback). British Journal for the History of Science, 36(3): 375–377, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028168.

Dixon:2003:PEC

[Dix03b] Thomas (Thomas M.) Dixon.

From passions to emotions: the creation of a secular psychological category. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-82729-9 (hardcover). x + 287 pp. LCCN B105.E46

D59 2003. URL http://www.
loc.gov/catdir/description/
cam032/2003273091.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2003273091-b.html;
http://www.loc.gov/catdir/
samples/cam034/2003273091.
html;
http://www.loc.gov/catdir/
toc/cam032/2003273091.html.

Dixon:2005:BRD

[Dix05]  Thomas Dixon.  Book re-

view: David C. Lindberg and Ronald L. Numbers (eds.), When Science and Christianity Meet. Chicago and London: University of Chicago Press, 2003. Pp. xii + 357. ISBN 0-226-48214-6. £20.50, $29.00 (hardback). British Journal for the History of Science, 38(4):469– 471, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028457.

Dixon:2006:BRP

[Dix06]  Thomas Dixon.  Book re-

view: Paul White, Thomas Huxley: Making the ‘Man of Science’. Cambridge Science Biographies. Cambridge: Cambridge University Press, 2003. Pp. xiv + 205. ISBN 0-521-64967-6. £16.99, $22.00 (paperback). British Journal for the History of Science, 39(1): 138–140, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028564.

Dixon:2008:BRL

[Dix08a] Thomas Dixon. Book review: Laura J. Snyder, Reforming Philosophy: A Victorian Debate on Science and Society. Chicago and London: University of Chicago Press, 2006. Pp. x + 386. ISBN 0-226-76733-7. $45.00, £23.50 (hardback). British Journal for the History of Science, 41(3): 462–464, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dixon:2008:IAM

[Dix08b] Thomas (Thomas M.) Dixon. The invention of altruism: making moral meanings in Victorian Britain. A British Academy postdoctoral fellowship monograph. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-726426-3 (hardcover). xiv + 420 pp. LCCN BJ1474 .D59 2008.

Dixon:2008:SRV

[Dix08c] Thomas (Thomas M.) Dixon.

Science and religion: a very short introduction. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-929551-4. 150 pp. LCCN BL240.3 .D57 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0820/2008023565.html.

Dixon:2009:BRS

[Dix09] Thomas Dixon. Book reviews:

Steve Fuller, Dissent over Descent: Intelligent Design’s Challenge to Darwinism. Cambridge: Icon Books, 2008. Pp. v + 272. ISBN 978-184046804-5. £12.99 (hardback). Nathaniel C. Comfort (ed.), The Panda’s Black Box: Opening up the Intelligent Design Controversy. Baltimore: Johns Hopkins University Press, 2007. Pp. xv+165. ISBN 978-0-8018-8599-0. £13.50 (hardback). British Journal for the History of Science, 42(3):440–442, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Djerassi:2001:MPR

[Dje01] Carl Djerassi. This man’s pill:

reflections on the 50th birthday of the pill. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850872-7 (hardcover). viii + 308 pp. LCCN RG137.5 .D57

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0638/2001021982-d.html;

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0725/
                                        2001021982-b.html.

Daston:2005:TAN

[DM05a] Lorraine Daston and Gregg Mitman, editors. Thinking with animals: new perspectives on anthropomorphism. Columbia University Press, New York, NY, USA, 2005. ISBN 0-231-13038-4 (cloth). LCCN BL215 .T48

                                        2004.  URL http://www.e-
                                        streams.com/es0809/es0809_
                                        4341.html.

Delgado-Moreira:2005:BRA

[DM05b] Raquel Delgado-Moreira. Book review: Ayval Leshem, Newton on Mathematics and Spiritual Purity. International Archives of the History of Ideas, 183. Dordrecht, Boston and London: Kluwer Academic Publishers, 2003. Pp. x + 230. ISBN 1-4020-1151-2. £58.00, $86.00, EUR90.00 (hardback). British Journal for the History of Science, 38(4):479–480, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028464.

Driver:2005:TVA

[DM05c] Felix Driver and Luciana Martins, editors. Tropical visions in an age of empire. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-16471-3 (cloth), 0-226-16472-1 (paperback). xii + 279 + 8 pp. LCCN G905 .T75

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0621/2005001424-b.html;
http://www.loc.gov/catdir/
enhancements/fy0621/
2005001424-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip056/2005001424.html.

Dodgen:2001:CDC

[Dod01] Randall A. Dodgen. Controlling the dragon: Confucian engineers and the Yellow River in the late imperial China. University of Hawai’i Press, Honolulu, HI, USA, 2001. ISBN 0-8248-2191-2, 0-8248-2366-4 (paperback). ix + 243 pp. LCCN DS793.Y45 D63 2001.

Dodgen:2002:BRL

[Dod02] Randall Dodgen. Book review:

Laura Hostetler, Qing Colonial Enterprise: Ethnography and Cartography in Early Modern China. Chicago and London: University of Chicago Press, 2001. Pp. xx + 257. ISBN 0-226-35420-2. £22.50, $35.00 (hardback). British Journal for the History of Science, 35(3):347– 379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dodgen:2006:BRF

[Dod06] Randall Dodgen. Book review:

Fa-Ti Fan, British Naturalists in Qing China: Science, Empire and Cultural Encounter. Cambridge, MA and London: Harvard University Press, 2004. Pp. xi + 238. ISBN 0-674-01143-0. £32.95 (hardback). British Journal for the History of Science, 39(2):292–294, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028619.

Dolan:2000:MMM

[Dol00] Brian Dolan, editor. Malthus, medicine and morality: Malthusianism after 1798, volume 59 of Wellcome Institute series in the history of medicine; Clio medica, 0045-7183. Rodopi, Amsterdam, The Netherlands, 2000. ISBN 90-420-0841-5 (paper), 90-420-0851-2 (bound). LCCN HB863 .M237

                                        2000. URL http://www.loc.
                                        gov/catdir/toc/fy046/
                                        2004301817.html.

Dolan:2001:BRJ

[Dol01] Brian Dolan. Book review: June Z. Fullmer, Young Humphry Davy: The Making of an Experimental Chemist. Memoirs of the American Philosophical Society, 237. Philadelphia: American Philosophical Society, 2000. Pp. xvi + 385. ISBN 0-87169-237-6. $30.00 (hardback). British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028108.

Dolan:2002:BRR

[Dol02a] Brian Dolan. Book review: R. R. Angerstein, R. R. Angerstein’s Illustrated Travel Diary, 1753– 1755: Industry in England and Wales from a Swedish Perspective. Translated by Torsten and Peter Berg. With an introduction by Marilyn Palmer. London: Science Museum, 2001. Pp. xii + 378. ISBN 1-900747-24-3. £34.95 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028405.

Dolby:2002:BRH

[Dol02b] Alex Dolby. Book reviews:

Henry H. Bauer, Science or Pseudoscience: Magnetic Healing, Psychic Phenomena, and Other Heterodoxies. Urbana and Chicago: University of Illinois Press, 2001. Pp. xiii + 275. ISBN 0-252-02601-2. $29.95 (cloth). Michael Shermer, The Borderlands of Science: Where Science Meets Nonsense. Oxford: Oxford University Press, 2001. Pp. viii+360. ISBN 0-19-514326-4. £17.95 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028406.

Dolby:2002:BRK

[Dol02c] R. G. A. Dolby. Book review:

Ken Richardson, The Making of Intelligence. Maps of the Mind. New York: Columbia University Press, 2000. Pp. viii + 210. ISBN 0-231-12004-4. $24.95 (hardback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028183.

Dolby:2004:BRG

[Dol04a] Alex Dolby. Book review: Gregory N. Derry, What Science Is and How It Works. Princeton and Oxford: Princeton University Press, 2002. Pp. xi + 311. ISBN 0-691-09550-7. £13.95 (paperback). British Journal for the History of Science, 37(3): 365–366, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028145;
                                        http://www.jstor.org/
                                        stable/4028447.

Dolby:2004:BRJ

[Dol04b] Alex Dolby. Book review: John Waller, Fabulous Science: Fact and Fiction in the History of Scientific Discovery. Oxford: Oxford University Press, 2002. Pp. xi + 308. ISBN 0-19-280404-9. £18.99 (hardback). British Journal for the History of Science, 37(2):236–237, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028363.

Dolby:2004:BRR

[Dol04c] Alex Dolby. Book review: Rosalind Williams, Retooling: A Historian Confronts Technological Change. Cambridge, MA and London: MIT Press, 2002. Pp. xv + 252. ISBN 0-262-23223-5. £18.50 (hardback). British Journal for the History of Science, 37(2):237–238, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028364.

Doody:2004:BRJ

[Doo04] Aude Doody. Book review: John F. Healy, Pliny the Elder on Science and Technology. Oxford: Oxford University Press, 1999. Pp. xv + 467. ISBN 0-19-814687-6. £65.00 (hardback). British Journal for the History of Science, 37(3):345–346, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028428.

Doody:2005:BRG

[Doo05] Aude Doody. Book review: G.

E. R. Lloyd, In the Grip of Disease: Studies in the Greek Imagination. Oxford: Oxford University Press, 2003. Pp. xxi + 258. ISBN 0-19-925323-4. £25.00 (hardback). British Journal for the History of Science, 38(4): 476–477, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028461.

Dougherty:2006:BRM

[Dou06] Carolyn Dougherty. Book review: Michael R. Bailey (ed.), Robert Stephenson: The Eminent Engineer. Aldershot: Ashgate, 2003. Pp. xxvii + 401. ISBN 0-7546-3679-8. £55.00 (hardback). British Journal for the History of Science, 39(2):296–297, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028622.

Daston:1998:WON

[DP98] Lorraine Daston and Katharine Park. Wonders and the order of nature, 1150–1750. Zone Books, New York, NY, USA, 1998. ISBN 0-942299-90-6. 511 pp. LCCN Q124.97 .D38 1998.

Drayton:2000:NGS

[Dra00] Richard Harry Drayton. Nature’s government: science, imperial Britain, and the ‘Improvement’ of the World. Yale University Press, New Haven, CT, USA, 2000. ISBN 0-300-05976-0. xxi + 346 pp. LCCN DA470 .D73 2000.

Driver:2001:GMC

[Dri01] Felix Driver. Geography militant: cultures of exploration and empire. Blackwell Publishers, Oxford, UK, 2001. ISBN 0-631-20111-4 (hardcover), 0-631-20112-2 (paperback). viii + 258 pp. LCCN G70 .D75 2001.

Driver:2003:BRN

[Dri03]  Felix Driver.  Book review:

Nancy Leys Stepan, Picturing Tropical Nature. Picturing History. London: Reaktion Books, 2001. Pp. 283. ISBN 1-86189-084-2. £25.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028302.

Dritsas:2005:LNP

[Dri05] Lawrence Dritsas. From Lake Nyassa to Philadelphia: a geography of the Zambesi Expedition, 1858–1864. British Journal for the History of Science, 38(1):35–52, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dry:2009:SNF

[Dry09] Sarah Dry. Safety networks: fishery barometers and the outsourcing of judgement at the early Meteorological Department. British Journal for the History of Science, 42(1):35–56, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Divall:2001:MHT

[DS01] Colin Divall and Andrew Scott.

Making histories in transport museums. Making histories in museums. Leicester University Press, London, UK, 2001. ISBN 0-7185-0106-3. viii + 221 pp. LCCN TA1006.A1 D58 2001.

Dauben:2002:WHM

[DS02] Joseph Warren Dauben and Christoph J. Scriba, editors.

Writing the history of mathematics: its historical development, volume 27 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 2002. ISBN 0-8176-6166-2, 3-7643-6166-2 (hardcover), 3-7643-6167-0 (softcover). xxxvii + 689 pp. LCCN QA21 .W75

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0829/2002027806-d.html;
http://www.loc.gov/catdir/
enhancements/fy0829/
2002027806-t.html.

Doel:2006:HCS

[DS06]  Ronald  Edmund  Doel

and Thomas Söderqvist, editors. The historiography of contemporary science, technology, and medicine: writing recent science, volume 23 of Routledge studies in the history of science, technology, and medicine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2006. ISBN 0-415-27294-7 (hardcover), 0-415-39142-3 (paperback), 0-203-32388-2 (e-book). xv + 313 pp. LCCN Q175 .H496 2006.

URL  http://www.loc.gov/
catdir/enhancements/fy0654/
2005031144-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip062/2005031144.html.

Dubow:2006:CKS

[Dub06] Saul Dubow. A commonwealth of knowledge: science, sensibility, and white South Africa, 1820– 2000. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-929663-4. viii + 296 pp. LCCN DT1756 .D83 2006. URL http://www.

loc.gov/catdir/
enhancements/fy0725/
2006021303-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0616/2006021303.html.

Duffin:2009:BRR

[Duf09]  Jacalyn Duffin.  Book re-

view: Roberta Bivins, Alternative Medicine? A History. New York: Oxford University Press, 2007. Pp. xvii + 238. ISBN 978-0-19-921887-5. £14.99 (hardback). British Journal for the History of Science, 42(2):284– 285, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunlap:1999:NED

[Dun99] Thomas R. Dunlap. Nature and the English diaspora: environment and history in the United States, Canada, Australia, and New Zealand. Studies in environment and history. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-65173-5 (hardcover), 0-521-65700-8 (paperback). xv + 350 pp. LCCN QH15 .D85 1999.

URL  http://www.loc.gov/
catdir/description/cam029/
98043736.html;   http://
www.loc.gov/catdir/samples/
cam032/98043736.html;
http://www.loc.gov/catdir/
toc/cam026/98043736.html.

Dunn:2008:BRD

[Dun08] Richard Dunn. Book review: A.

D. Morrison-Low, Making Scientific Instruments in the Industrial Revolution. Aldershot: Ashgate, 2007. Pp. xvi + 408. ISBN 978-0-7546-5758-3. £55.00 (hardback). British Journal for the History of Science, 41(3): 459–460, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Dunn:2009:MCH

[Dun09] Richard Dunn. Material culture in the history of science: case studies from the National Maritime Museum. British Journal for the History of Science, 42(1):31–33, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Durbach:2005:BMA

[Dur05] Nadja Durbach. Bodily matters: the anti-vaccination movement in England, 1853–1907. Radical perspectives. Duke University Press, Durham, NC, USA, 2005. ISBN 0-8223-3412-7 (cloth), 0-8223-3423-2 (paperback). xiii + 276 pp. LCCN RA644.S6 D874

                                        2005. URL http://www.loc.
                                        gov/catdir/toc/ecip0418/
                                        2004013140.html.

Daston:2004:MAN

[DV04] Lorraine Daston and Fernando Vidal, editors. The moral authority of nature. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-13680-9, 0-226-13681-7 (paperback). vii + 519 pp. LCCN BD581 .M78

                                        2004. URL http://www.loc.

gov/catdir/bios/uchi051/
2003008351.html;
http://www.loc.gov/catdir/
description/uchi052/
2003008351.html;   http://
www.loc.gov/catdir/toc/
ecip042/2003008351.html.

Eagleton:2005:BRD

[Eag05a] Catherine Eagleton. Book review: David A. King, The Ciphers of the Monks: A Forgotten Number-Notation of the Middle Ages. Boethius, 44. Stuttgart: FranzSteiner Verlag, 2001. Pp. 506. ISBN 3-515-07640-9. DM 199.49, EUR 102.00 (hardcover). British Journal for the History of Science, 38(3): 359–360, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028680.

Eagleton:2005:BRS

[Eag05b] Catherine Eagleton. Book review: Sigmund Eisner (ed.), A Variorum Edition of the Works of Geoffrey Chaucer. Volume VI: The Prose Treatises. Part One: A Treatise on the Astrolabe. Norman: University of Oklahoma, 2002. Pp. xxiv + 358. ISBN 0-8061-3413-5. $75.00 (hardback). British Journal for the History of Science, 38 (4):478, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028463.

Eagleton:2006:BRH

[Eag06a] Catherine Eagleton. Book review: Hester Higton, with contributions from Silke Ackermann, Richard Dunn, Kiyoshi Takada and Anthony Turner, Sundials at Greenwich: A Catalogue of the Sundials, Horary Quadrants and Nocturnals in the National Maritime Museum, Greenwich. Oxford and New York: Oxford University Press and National Maritime Museum, 2002. Pp. x + 463. ISBN 0-19-850877-8. £99.50 (hardback). British Journal for the History of Science, 39 (1):129–130, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028556.

Eagleton:2006:BRJ

[Eag06b] Catherine Eagleton. Book review: Jean A. Givens, Observation and Image-Making in Gothic Art. Cambridge: Cambridge University Press, 2004. Pp. xxiv + 231. ISBN 0-521-83031-1. £45.00, $80.00 (hardback). British Journal for the History of Science, 39(3):444– 445, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028491.

Eagleton:2007:BRJ

[Eag07] Catherine Eagleton. Book review: John North, God’s Clockmaker: Richard of Wallingford and the Invention of Time. London and New York: Hambledon and London, 2005. Pp. xix + 441. ISBN 1-85285-451-0. £25.00 (hardback). British Journal for the History of Science, 40(1): 127–128, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500694.

Eagleton:2008:BRR

[Eag08] Catherine Eagleton. Book review: R. J. W. Evans and Alexander Marr (eds.) Curiosity and Wonder from the Renaissance to the Enlightenment. Aldershot: Ashgate, 2006. Pp. xv + 265. ISBN 0-7546-4102-3. £55.00, $99.95 (hardback). British Journal for the History of Science, 41(2):293–294, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

El-Bizri:2007:BRR

[EB07] Nader El-Bizri. Book review:

Roshdi Rashed, Geometry and Dioptrics in Classical Islam. London: al-Furqan Islamic Heritage Foundation, 2005. Pp. xii + 1178. ISBN 1873992998. £50.00 (hardback). British Journal for the History of Science, 40(1): 124–126, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500692.

Einstein:2000:EQE

[EC00]  Albert  Einstein  and

Alice Calaprice. The expanded quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-07021-0 (cloth). xliii + 407 pp. LCCN QC16.E5 A25 2000. URL http:

                                        //www.loc.gov/catdir/bios/
                                        prin051/00026873.html;
                                        http://www.loc.gov/catdir/
                                        description/prin022/
                                        00026873.html;   http://
                                        www.loc.gov/catdir/samples/
                                        prin031/00026873.html;
                                        http://www.loc.gov/catdir/
                                        toc/prin031/00026873.html.

Elliott:2006:STU

[ED06] Paul Elliott and Stephen Daniels. The ‘school of true, useful and universal science’? Freemasonry, natural philosophy and scientific culture in eighteenth-century England. British Journal for the History of Science, 39(2):207–229, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eddy:2002:SCC

[Edd02] M. D. Eddy. Scottish chemistry, classification and the early mineralogical career of the ‘ingenious’ Rev. Dr. John Walker (1746 to 1779). British Journal for the History of Science, 35(4):411–438, December 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028275.

Eddy:2003:BRC

[Edd03a] M. D. Eddy. Book review:

Charles W. J. Withers, Geography, Science and National Identity: Scotland since 1520. Cambridge Studies in Historical Geography, 33. Cambridge: Cambridge University Press, 2001. Pp. xvii + 310. ISBN 0-521-64202-7. £45.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028303.

Eddy:2003:BRD

[Edd03b] M. D. Eddy. Book review: David L. Cowen, Pharmacopoeias and Related Literature in Britain and America, 1618–1847. Variorum Collected Studies Series, CS700. Aldershot: Ashgate, 2001. Pp. ix + 296. ISBN 0-86078-842-3. £55.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028304.

Eddy:2004:BRF

[Edd04a] M. D. Eddy. Book review:

Fallible or inerrant? A belated review of the ‘constructivist’s bible’ Jan Golinski, Making Natural Knowledge: Constructivism and the History of Science. Cambridge History of Science. Cambridge: Cambridge University Press, 1999. Pp. xiv + 236. ISBN 0-521-44913-8. £15.95 (paperback). British Journal for the History of Science, 37 (1):93–98, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028258.

Eddy:2004:SCC

[Edd04b] M. D. Eddy. Scottish chemistry, classification and the late mineralogical career of the ‘ingenious’ Professor John Walker (1779–1803). British Journal for the History of Science, 37(4): 373–399, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028148;
                                        http://www.jstor.org/
                                        stable/4028640.

Eddy:2005:BRH

[Edd05a] M. D. Eddy. Book review: Hugh Torrens, The Practice of British Geology, 1750–1850. Variorum Collected Studies Series, CS736. Aldershot: Ashgate, 2002. Pp.

372. ISBN 0-86078-876-8. £59.50 (hardback). British Journal for the History of Science, 38 (2):231–232, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028707.

Eddy:2005:BRW

[Edd05b] M.  D.  Eddy.

Book reviews: William R. Newman, Gehennical Fire: The Lives of George Starkey, an American Alchemist in the Scientific Revolution. With a New Forward. Chicago and London: University of Chicago Press, 2003. Pp. xxiv + 348. ISBN 0-226-57714-7. £19.50, $27.50 (paperback). William R. Newman and Lawrence M. Principe, Alchemy Tried in the Fire: Starkey, Boyle, and the Fate of Helmontian Chymistry. Chicago and London: University of Chicago Press, 2002. Pp. xv+344. ISBN 0-226-57711-2. £28.00, $40.00 (hardback). British Journal for the History of Science, 38(3): 364–366, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028685.

Eddy:2006:BRR

[Edd06] Matthew D. Eddy. Book review: R. D. Connor and A. D. Simpson, with A. D. Morrison-Low (ed.), Weights and Measures in Scotland: A European Perspective. Edinburgh: NMS Publishing, 2004. Pp. xvi + 842. ISBN 1-901663-88-4. £50.00 (hardback). British Journal for the History of Science, 39(4): 601–602, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028525.

Eddy:2007:BRH

[Edd07a] Matthew D. Eddy. Book review:

Hubert Steinke, Irritating Experiments: Haller’s Concept and the European Controversy on Irritability and Sensibility, 1750– 1790. The Wellcome Series in the History of Medicine. Clio Medicana 76. Union, NJ and Amsterdam: Editions Rodopi B. V., 2005. Pp 354. ISBN 90-420-1852-6. EUR75.00, $94.00 (hardback). British Journal for the History of Science, 40(4):610–611, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eddy:2007:BRK

[Edd07b] Matthew D. Eddy. Book review: Katharine Anderson, Predicting the Weather: Victorians and the Science of Meteorology. Chicago and London: University of Chicago Press, 2005. Pp. x + 331. ISBN 0-226-019680-3. £31.50, $45.00 (hardback). British Journal for the History of Science, 40(2):295–297, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500733.

Eddy:2008:BRU

[Edd08] Matthew D. Eddy. Book review: Ursula Klein and Wolfgang Lefèvre, Materials in Eighteenth-Century Science: A Historical Ontology. Cambridge, MA and London: MIT Press, 2007. Pp. x + 345. ISBN 978-0-262-11306-6. £24.95 (hardback). British Journal for the History of Science, 41(4):610–611, December 2008.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Eddy:2009:BRA

[Edd09] Matthew D. Eddy. Book review:

Alix Cooper, Inventing the Indigenous: Local Knowledge and Natural History in Early Modern Europe. Cambridge: Cambridge University Press, 2007. Pp. xiii + 218. ISBN 978-0-521-87087-0. £45.00, $75.00 (hardback). British Journal for the History of Science, 42(4):617–619, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Edgerton:2004:BRG

[Edg04]  David Edgerton.  Book re-

view: Guy Hartcup, The Effect of Science on the Second World War. Basingstoke: Palgrave Macmillan, 2003. Pp. xvii + 214. ISBN 1-4039-0643-2. £17.99 (paperback). British Journal for the History of Science, 37(2):230, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028357.

Edgerton:2005:BRA

[Edg05] David Edgerton. Book review:

Augustí Nieto-Galan, Santponç – Monturiol – Isaac Peral: la seducción de la máquina vapores, submarinos e inventores. With a Preface by Saturnino de la Plaza. Novatores, 2. Madrid: Nivola, 2001. Pp. 129. ISBN 84-95599-10-4. No price given (paperback). British Journal for the History of Science, 38(2):234–235, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028709.

Edgerton:2006:WSB

[Edg06] David Edgerton. Warfare state: Britain, 1920–1970. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85636-1, 0-521-67231-7 (paperback). xv + 364 pp. LCCN DA566.7 .E26

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0642/2006296114-d.html;
http://www.loc.gov/catdir/
enhancements/fy0642/
2006296114-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2006296114-b.html.

Edgerton:2008:BRM

[Edg08] David Edgerton. Book review: Michael Adas, Dominance by Design: Technological Imperatives and America’s Civilizing Mission. Cambridge, MA and London: The Belknap Press of Harvard University Press, 2006. Pp. 542. ISBN 0-674-01876-2. £18.95, $29.95, EUR25.50 (hardback). British Journal for the History of Science, 41(1):123– 124, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Edney:1997:MEG

[Edn97] Matthew H. Edney. Mapping an empire: the geographical construction of British India, 1765– 1843. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-18487-0 (cloth), 0-226-18488-9 (paperback). xv + 458 pp. LCCN GA1131 .E36

1997. URL http://www.loc.
gov/catdir/description/
uchi052/96039703.html;
http://www.loc.gov/catdir/
toc/uchi051/96039703.html.

Edson:2004:BRN

[Eds04] Evelyn Edson. Book review: Natalia Lozovsky, ‘The Earth is Our Book’: Geographical Knowledge in the Latin West ca. 400–1000. Recentiores: Latin Texts and Contexts. Ann Arbor: University of Michigan Press, 2000. Pp. ix + 182. ISBN 0-472-11132-9. £28.00, $44.50 (hardback). British Journal for the History of Science, 37(2):196–197, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028331.

Efron:2007:JSH

[Efr07] Noah J. Efron. Judaism and science: a historical introduction. Greenwood guides to science and religion. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2007. ISBN 0-313-33053-0. xx + 348 pp. LCCN BM538.S3 E34 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006014784.html.

Elmer:2004:HDS

[EG04] Peter Elmer and Ole Peter Grell, editors. Health, disease, and society in Europe, 1500–1800: a source book. Manchester University Press, Manchester, UK, 2004. ISBN 0-7190-6736-7 (hardcover), 0-7190-6737-5 (paperback). xx + 380 pp. LCCN R131 .H36 2004. URL http:/

/www.loc.gov/catdir/bios/
hol055/2004351456.html;
http://www.loc.gov/catdir/
description/hol054/
2004351456.html;
http://www.loc.gov/catdir/
toc/hol052/2004351456.html.

Elmer:2002:BRM

[Elm02]  Peter Elmer.  Book

review: Michael Hunter (ed.), The Occult Laboratory: Magic, Science and Second Sight in Late Seventeenth-Century Scotland. A New Edition of Robert Kirk’s The Secret Commonwealth and Other Texts. Woodbridge: Boydell Press, 2001. Pp. vii + 247. ISBN 0-85115-801-3. £50.00, $90.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028184.

Elman:2005:TOT

[Elm05] Benjamin A. Elman. On their own terms: science in China, 1550–1900. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01685-8. xxxviii + 567 pp. LCCN Q127.C5 E48 2005. URL http:

//www.loc.gov/catdir/toc/
fy053/2004059654.html.

Elman:2006:CHM

[Elm06] Benjamin A. Elman. A cultural history of modern science in China. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2006. ISBN 0-674-02306-4. xii + 308 pp. LCCN Q127.C5 E46 2006. URL http:

//www.loc.gov/catdir/toc/
fy0706/2006042706.html.

Elman:2009:BRF

[Elm09] Benjamin A. Elman. Book review: Francesca Bray, Vera Dorofeeva-Lichtmann and Georges Métailié (eds.), Graphics and Text in the Production of Technical Knowledge in China. Leiden and Boston, MA: Brill, 2007. Pp. xiii + 772. ISBN 978-90-04-16063-7. $149.00, $199.00 (hardback). British Journal for the History of Science, 42(3): 450–451, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Elton:2001:AE

[Elt01] Charles S. (Charles Sutherland)

Elton. Animal ecology. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-20639-4 (paperback). lvi + 209 pp. LCCN QH541 .E398 2001.

URL http://www.e-streams.
com/es0411/es0411_1566.htm;
http://www.loc.gov/catdir/
bios/uchi051/00069087.html;
http://www.loc.gov/catdir/

description/uchi051/
00069087.html;
http://www.loc.gov/catdir/
toc/uchi051/00069087.html.

Elzinga:2006:ENP

[Elz06] Aant Elzinga. Einstein’s Nobel Prize: a glimpse behind closed doors: the archival evidence, volume 6 of Archives of the Nobel Museum. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2006. ISBN 0-88135-283-7. xii + 228 pp. LCCN QC16.E5 E68 2006.

Evans:2006:CWR

[EM06] Robert John Weston Evans and Alexander Marr, editors. Curiosity and wonder from the Renaissance to the Enlightenment. Ashgate, Aldershot, Hampshire, England, 2006. ISBN 0-7546-4102-3. xvi + 265 pp. LCCN CB361 .C87 2006. URL http:

//www.loc.gov/catdir/toc/
ecip066/2005037430.html.

Emmerson:2001:BRD

[Emm01] Andrew Emmerson. Book review: Donald F. McLean, Restoring Baird’s Image. IEE History of Technology Series, 27. London: Institution of Electrical Engineers, 2000. Pp. xx + 295. ISBN 0-85296-795-0. £29.00, $55.00. British Journal for the History of Science, 34(2):233–250, June 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028013.

Endersby:2000:GEB

[End00] Jim Endersby. A garden enclosed: botanical barter in Sydney, 1818–1839. British Journal for the History of Science, 33(3):313–334, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027955.

Endersby:2004:BRD

[End04a] Jim Endersby. Book review:

David Elliston Allen, Naturalists and Society: The Culture of Natural History in Britain, 1700–1900. Variorum Collected Studies Series: CS724. Aldershot: Ashgate, 2001. Pp. xiv + 298. ISBN 0-86078-863-6. £55.00 (hardback). British Journal for the History of Science, 37(1): 108–110, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028266.

Endersby:2004:BRR

[End04b] Jim Endersby. Book review: R.

W. Home, A. M. Lucas, Sara Maroske, D. M. Sinkora and J. H. Voigt (eds.), Regardfully Yours: Selected Correspondence of Ferdinand von Mueller. Volume II: 1860–1875. Bern: Peter Lang, 2002. Pp. 865. ISBN 3-906757-09-9. £36.00 (hardback). British Journal for the History of Science, 37(2):217–219, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028346.

Enebakk:2009:LRS

[Ene09] Vidar Enebakk. Lilley revisited: or science and society in the twentieth century. British Journal for the History of Science, 42(4):563–593, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Euclid:2001:JTR

[EoToBB01] Euclid, John of Tynemouth d 1290?, Adelard of Bath, ca 1116– 1142, and H. L. L. (Hubertus Lambertus Ludovicus) Busard. Johannes de Tinemue’s redaction of Euclid’s Elements, the socalled Adelard III version, volume 45 of Boethius. F. Steiner, Stuttgart, Germany, 2001. ISBN 3-515-07975-0. 632 pp. LCCN QA31 .E83165 2001. Two volumes.

Eagleton:2004:BRP

[ETV04] Catherine Eagleton, Karin Tybjerg, and Koen Vermeir. Book review: Pamela O. Long, Openness, Secrecy, Authorship: Technical Arts and the Culture of Knowledge from Antiquity to the Renaissance. Baltimore and London: Johns Hopkins University Press, 2001. Pp. xiii + 364. ISBN 0-8018-6606-5. £38.00 (hardback). British Journal for the History of Science, 37(1): 101–103, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028261.

Ewin:2001:FWF

[Ewi01] Jeannette Ewin. Fine wines and fish oil: the life of Hugh Macdonald Sinclair. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-262927-1 (hardcover). xx + 338 pp. LCCN QP143 .E955

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0612/2001034060-d.html;
http://www.loc.gov/catdir/
enhancements/fy0612/
2001034060-t.html.

Falcon:2005:ASN

[Fal05] Andrea Falcon. Aristotle and the science of nature: unity without uniformity. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-85439-3 (hardcover). xvii + 139 pp. LCCN B491.N3

F35 2005. URL http://www.
loc.gov/catdir/
enhancements/fy0632/
2006295711-d.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2006295711-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2006295711-b.html.

Fan:2002:BRR

[Fan02] Fa-Ti Fan. Book review: Randall A. Dodgen, Controlling the Dragon: Confucian Engineers and the Yellow River in Late Imperial China. Honolulu: University of Hawaii Press, 2001.

Pp. ix + 243. ISBN 0-8248-2366-4. $29.95 (paperback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028185.

Fan:2003:VNC

[Fan03] Fa-Ti Fan. Victorian naturalists in China: science and informal empire. British Journal for the History of Science, 36(1):1–26, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028290.

Fara:2000:INL

[Far00a] Patricia Fara. Isaac Newton lived here: Sites of memory and scientific heritage. British Journal for the History of Science, 33(4):407–426, December 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028028.

Faraday:2000:ERE

[Far00b] Michael Faraday. Experimental researches in electricity. Green Lion Press, Santa Fe, NM, USA, 2000. ISBN 1-888009-09-8 (vol. 1), 1-888009-10-1 (vol. 2), 1-888009-11-X (vol. 3). ???? pp. LCCN QC503 .F21 2000.

Farber:2000:FON

[Far00c] Paul Lawrence Farber. Finding order in nature: the naturalist tradition from Linnaeus to E. O. Wilson. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2000. ISBN 0-8018-6389-9, 0-8018-6390-2 (paperback). x + 136 pp. LCCN QH15 .F27

2000. URL http://www.loc.
gov/catdir/bios/jhu052/
99089621.html;
http://www.loc.gov/catdir/
description/jhu052/
99089621.html.

Fara:2002:NMG

[Far02] Patricia Fara. Newton: the making of a genius. MacMillan Publishing Company, New York, NY, USA, 2002. ISBN 0-333-90735-3. xvi + 347 + 16 pp. LCCN QC16.N7 .F37 2002.

Farber:2006:BRL

[Far06]  Paul Farber.  Book review:

Lorraine Daston and Fernando Vital (eds.), The Moral Authority of Nature. Chicago and London: University of Chicago Press, 2004. Pp. vii + 519. ISBN 0-226-13681-7. $26.00 (paperback). British Journal for the History of Science, 39(4): 595–596, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028521.

Fara:2007:BRC

[Far07a] Patricia Fara. Book review: C.

U. M. Smith and Robert Arnott (eds.), The Genius of Erasmus Darwin. Aldershot: Ashgate, 2005. Pp. xvii + 416. ISBN 0-7546-3671-2. £60.00 (hardback). British Journal for the History of Science, 40(1):139–140, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500701.

Fara:2007:BRD

[Far07b] Patricia Fara. Book review: David Edgerton, The Shock of the Old: Technology and Global History since 1900. London: Profile Books, 2007. Pp. xviii + 270. ISBN 1-86197-296-2. £18.99 (hardback). British Journal for the History of Science, 40(4): 621–622, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2007:BRN

[Far07c] Patricia Fara. Book review: Nicolaas A. Rupke, Alexander von Humboldt: A Metabiography. Frankfurt am Main: Peter Lang, 2005. Pp. 320. ISBN 3-631-53932-0. £22.80 (hardback). British Journal for the History of Science, 40(2):293–294, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500731.

Fara:2009:BRW

[Far09a] Patricia Fara. Book reviews: William Bynum, The History of Medicine: A Very Short Introduction. Oxford: Oxford University Press, 2008. Pp. 169. ISBN 978-0-19-921543-0. £7.99 (paperback). Thomas Dixon, Science and Religion: A Very Short Introduction. Oxford: Oxford University Press, 2008. Pp. 150. ISBN 978-0-19-929551-7. £7.99 (paperback). British Journal for the History of Science, 42 (2):293–294, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fara:2009:SFT

[Far09b] Patricia Fara. Science: a four thousand year history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-922689-X. xv + 408 pp. LCCN Q125 .F27

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0909/2008050975-b.html;
http://www.loc.gov/catdir/
enhancements/fy0909/
2008050975-d.html;
http://www.loc.gov/catdir/
toc/fy0904/2008050975.html.

Fauque:2003:LNC

[Fau03] Danielle Fauque. Lavoisier et la naissance de la chimie moderne. (French) [Lavoisier and the birth of modern chemistry]. Vuibert, Paris, France, 2003. ISBN 2-7117-5353-0. 230 pp. LCCN QD14 .F38 2003.

Feingold:2004:NMI

[Fei04] Mordechai Feingold. The Newtonian moment: Isaac Newton and the making of modern culture. New York Public Library, New York, NY, USA, 2004.

ISBN 0-19-517735-5 (cloth), 0-19-517734-7 (paperback). xv + 218 pp. LCCN QC16.N54 F45

                                        2004. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0616/2004018090-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0724/
                                        2004018090-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0421/2004018090.html.

Ferry:2003:CCL

[Fer03] Georgina Ferry. A computer called LEO: Lyons teashops and the world’s first office computer. Fourth Estate, London, UK, 2003. ISBN 1-84115-185-8. xi + 220 pp. LCCN QA76.8.L46 F47 2003.

Fetzer:2000:SER

[Fet00] James H. Fetzer, editor. Science, explanation, and rationality: aspects of the philosophy of Carl G. Hempel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-512137-6. xxix + 342 pp. LCCN B945.H454 S35

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0603/99087642-d.html.

Fox:2005:POL

[FG05] Robert Fox and Graeme Gooday, editors. Physics in Oxford, 1839–1939: laboratories, learning, and college life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856792-8. xix + 363 pp. LCCN QC47.G73 O947

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0639/2005296514-d.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2005296514-b.html;
http://www.loc.gov/catdir/
toc/fy0601/2005296514.html.

Fox:2000:THE

[FH00] Robert Fox and Thomas Harriot, editors. Thomas Harriot: an Elizabethan man of science. Ashgate, Aldershot, Hampshire, England, 2000. ISBN 0-7546-0078-5. x + 317 pp. LCCN Q143.H36 .T45 2000.

Force:2004:NNN

[FH04] James E. Force and Sarah Hutton, editors. Newton and Newtonianism: new studies, volume 188 of Archives internationale d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-1969-6, 1-4020-2238-7 (ebook). LCCN B1299.N34 N48

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0663/2004044084-d.html;
http://www.loc.gov/catdir/
enhancements/fy0814/
2004044084-t.html. Proceed-

ings of a conference held in November 2000 at the William Andrews Clark Memorial Library.

Fichman:2004:EVE

[Fic04] Martin Fichman. An elusive Victorian: the evolution of Alfred Russel Wallace. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-24613-2. x + 382 pp. LCCN QH31.W2 F524 2004. URL http://www.

                                        loc.gov/catdir/bios/
                                        uchi051/2003010949.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi052/
                                        2003010949.html;   http://
                                        www.loc.gov/catdir/toc/
                                        ecip044/2003010949.html.

Field:2005:BRW

[Fie05] J. V. Field. Book review: Wolfgang Lefèvre, Jürgen Renn and Urs Schoepflin (eds.), The Power of Images in Early Modern Science. Basel, Boston and Berlin: Birkhäuser Verlag, 2003. Pp. ix + 308. ISBN 3-7643-2434-1. EUR118.00 (paperback). British Journal for the History of Science, 38(3):361, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028682.

Findlen:2001:BRL

[Fin01] Paula Findlen. Book review: Lorraine Daston and Katharine Park, Wonders and the Order of Nature, 1150–1750. New York: Zone Books, 1998. Pp. 511. ISBN 0-942299-90-6. £24.95, $36.50. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028014.

Finlay:2002:BRV

[Fin02] Mark R. Finlay. Book review: Vaclav Smil, Enriching the Earth: Fritz Haber, Carl Bosch, and the Transformation of World Food Production. Cambridge, MA and London: MIT Press, 2001. Pp. xvii + 338. ISBN 0-262-19449-X. £23.95 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028407.

Findlen:2003:BRA

[Fin03a] Paula Findlen. Book review: Andrew C. Scott and David Freedberg, Fossil Woods and Other Geological Specimens. The Paper Museum of Cassiano dal Pozzo, Series B: Natural History, Part Three. Turnhout: Harvey Miller Publishers, 2000. Pp. 424. ISBN 1-872501-91-5. EUR232.00 (hardback). British Journal for the History of Science, 36(4): 478–480, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028219;
http://www.jstor.org/
stable/4028279.

Finkelstein:2003:MDB

[Fin03b] Gabriel Finkelstein. M. du Bois-Reymond goes to Paris. British Journal for the History of Science, 36(3):261–300, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028156.

Finnegan:2004:WIG

[Fin04] Diarmid A. Finnegan. The work of ice: glacial theory and scientific culture in early Victorian edinburgh. British Journal for the History of Science, 37(1):29–52, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028255.

Finnegan:2005:NHS

[Fin05a] Diarmid A. Finnegan. Natural history societies in late Victorian Scotland and the pursuit of local civic science. British Journal for the History of Science, 38(1):53– 72, March 2005. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Finocchiaro:2005:RG

[Fin05b] Maurice A. Finocchiaro. Retrying Galileo, 1633–1992. University of California Press, Berkeley, CA, USA, 2005. ISBN 0-520-24261-0 (cloth). xii + 485 pp. LCCN QB36.G2 F56

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0641/2004001861-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0735/
                                        2004001861-b.html;   http:/

/www.loc.gov/catdir/toc/
ecip0414/2004001861.html.

Findlen:2009:BRP

[Fin09] Paula Findlen. Book review:

Paola Bertucci, Viaggio nel paese delle meraviglie. Scienza e curiosità nell’Italia del Settecento. Turin: Bollati Boringhieri, 2007. Pp. 294. ISBN 978-88-339-1802-0. EUR28.00 (paperback). British Journal for the History of Science, 42(1):135– 136, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fisher:2004:BRL

[Fis04]  Nick Fisher.  Book review:

Louise Purbrick (ed.), The Great Exhibition of 1851: New Interdisciplinary Essays. Texts in Culture. Manchester and New York: Manchester University Press, 2001. Pp. xii + 217. ISBN 0-7190-5592-X. £15.99 (paperback). British Journal for the History of Science, 37(4):478– 479, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028655.

Fitzgerald:2003:EFF

[Fit03] Deborah Kay Fitzgerald. Every farm a factory: the industrial ideal in American agriculture. Yale agrarian studies series. Yale University Press, New Haven, CT, USA, 2003. ISBN 0-300-08813-2 (cloth). xi + 242 pp. LCCN HD1761 .F564

                                        2003. URL http://www.loc.
                                        gov/catdir/description/
                                        yale051/2002012135.html.

Faraday:1991:CMF

[FJ08] Michael Faraday and Frank A. J. L. James. The correspondence of Michael Faraday. [History of technology series]. Institution of Electrical Engineers (IEE), London, UK, 1991–2008. ISBN 0-86341-248-3 (vol. 1), 0-86341-249-1 (vol. 2), 0-86341-250-5 (vol. 3), 0-86341-251-3 (vol. 4), 0-86341-823-6 (vol. 5). xlix + 673 (vol. 1), lv + 807 (vol. 2), lvii + 835 (vol. 3), lxv + 1003 (vol. 4), lviii + 835 (vol. 5) pp. LCCN QC16.F2 A4 1991.

Fleming:2006:IUL

[FJC06]  James  Rodger

Fleming, Vladimir Janković, and Deborah R. Coen, editors. Intimate universality: local and global themes in the history of weather and climate, volume 1 of Science-history studies on atmospheres. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2006. ISBN 0-88135-367-1. xx + 264 pp. LCCN QC855 .I58 2006.

French:2006:IPH

[FK06] Steven French and Decio Krause. Identity in physics: a historical, philosophical, and formal analysis. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-927824-5 (hardcover). xv + 422 pp. LCCN QC6.4.I34 F74

                                        2006. URL http://www.loc.

gov/catdir/enhancements/
fy0640/2006008189-d.html;
http://www.loc.gov/catdir/
enhancements/fy0640/
2006008189-t.html.

Fleming:1998:HPC

[Fle98] James Rodger Fleming. Historical perspectives on climate change. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-507870-5. xi + 194 pp. LCCN QC981.8.C5 F45

1998. URL http://www.loc.
gov/catdir/enhancements/
fy0604/97022545-d.html;
http://www.loc.gov/catdir/
enhancements/fy0604/
97022545-t.html.

Friedman:2006:KLN

[FN06] Michael Friedman and Alfred Nordmann, editors. The Kantian legacy in nineteenth-century science. Dibner Institute for the history of science and technology. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-06254-2. vi + 370 pp. LCCN Q174.8 .K36

2006. URL http://www.h-net.
org/review/hrev-a0f6w2-aa;
http://www.loc.gov/catdir/
toc/fy0703/2005058035.html.

Forgan:2003:BRC

[For03] Sophie Forgan. Book review:

Christine Stevenson, Medicine and Magnificence: British Hospital and Asylum Architecture, 1660–1815. New Haven and London: Yale University Press for the Paul Mellon Centre for Studies in British Art, 2000. Pp. viii + 312. ISBN 0-300-08536-2. £30.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028305.

Forgan:2004:BRH

[For04] Sophie Forgan. Book review: Hermione Hobhouse, The Crystal Palace and the Great Exhibition: Art, Science and Productive Industry. A History of the Royal Commission for the Exhibition of 1851. London and New York: Athlone, 2002. Pp. xx + 451. ISBN 0-485-11575-1. £40.00 (hardback). British Journal for the History of Science, 37(4):479–480, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028656.

Forgan:2005:BRU

[For05] Sophie Forgan. Book review: Ulf Larsson (ed.), Cultures of Creativity: The Centennial Exhibition of the Nobel Prize. English translation by Daniel M. Olson. Nobel Museum Archives, 2. Canton, MA: Science History Publications, 2001. Pp. 228. ISBN 0-88135-288-8. $40.00 (paperback). British Journal for the History of Science, 38(2):242–243, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028716.

Fortun:2008:PGI

[For08] Michael Fortun. Promising genomics: Iceland and deCODE Genetics in a world of speculation. University of California Press, Berkeley, CA, USA, 2008. ISBN 0-520-24750-7 (cloth), 0-520-24751-5 (paperback). x + 330 pp. LCCN QH438.7 .F67

2008. URL http://www.loc.
gov/catdir/toc/ecip083/
2007044597.html.

Fox:2004:BRH

[Fox04] Robert Fox. Book review: Henry L E Chatelier, Science et industrie: Les Débuts du taylorisme en France. With a preface by Michel Letté. Format, 46. Paris: Editions du CTHS, 2001. Pp. xxxi + 283. ISBN 2-7355-0487-5. EUR14.00 (paperback). British Journal for the History of Science, 37(2):227, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028354.

Fox:2005:BRG

[Fox05]  Robert Fox.  Book review:

Guillermo Lusa Monforte (ed.), La creación de la Escuela Industrial Barcelonesa (1851): 1 de octubre de 1851–1 de octubre de 2001. Documentos de la Escuela de Ingenieros Industriales de Barcelona, 11. Barcelona: Escola Tècnica Superior d’Enginyeria Industrial de Barcelona, Universitat Politècnica de Catalunya, 2001. Pp. 168. ISSN 1137-0238. No price given (hardback). Guillermo Lusa Monforte, Inquietudes y reformas de cambio de siglo: el proyecto de nueva Escuela Industrial (1899– 1910). Documentos de la Escuela de Ingenieros Industriales de Barcelona, 12. Barcelona: Escola Tècnica Superior d’Enginyeria Industrial de Barcelona, Universitat Politècnica de Catalunya, 2002. Pp. 191. ISSN 1137-0238. No price given (paperback). Guillermo Lusa Monforte, El conflicto con la Diputación (1915). La plena incorporación de la Escuela al Estado (1917). Documentos de la Escuela de Ingenieros Industriales de Barcelona, 13. Barcelona: Escola Tècnica Superior d’Enginyeria Industrial de Barcelona, Universitat Politècnica de Catalunya, 2003. Pp. 191. ISSN 1137-0238. No price given (paperback). British Journal for the History of Science, 38(1):116–117, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028597.

Fox:2007:BRR

[Fox07] Robert Fox. Book review: Roger Hahn, Pierre Simon Laplace 1749–1827:A Determined Scientist. Cambridge, MA and London: Harvard University Press, 2005. Pp. xii + 310. ISBN 0-674-01892-3. £22.95, $35.00, EUR32.30 (hardback). British Journal for the History of Science, 40(4):613–614, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Franklin:2002:SDT

[Fra02] Allan Franklin. Selectivity and discord: two problems of experiment. University of Pittsburgh Press, Pittsburgh, PA, USA, 2002. ISBN 0-8229-4191-0. 290 pp. LCCN Q175 .F795

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0708/2002011054-d.html.

Franklin:2005:BRH

[Fra05] Allan Franklin. Book review:

Hans Radder (ed.), The Philosophy of Scientific Experimentation. Pittsburgh: University of Pittsburgh Press, 2003. Pp. xii + 311. ISBN 0-8229-5795-7. $29.95 (paperback). British Journal for the History of Science, 38(2):241–242, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028715.

Francis:2009:BRD

[Fra09] Mark Francis. Book review:

David Stack, Queen Victoria’s Skull: George Combe and the Mid-Victorian Mind. London: Hambledon Continuum, 2008. Pp. xviii + 350. ISBN 978 1 84725 233 3. British Journal for the History of Science, 42(4): 620–621, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Frenkel:1996:YIF

[Fre96] Viktor Iakovlevich Frenkel. Yakov Ilich Frenkel: his work, life, and letters. Birkhäuser Verlag, Basel, Switzerland, 1996. ISBN 3-7643-2741-3 (Basel), 0-8176-2741-3 (Boston). viii + 323 pp. LCCN QC16.F715 A3 1996.

French:1999:DVE

[Fre99] R. K. (Roger Kenneth) French. Dissection and vivisection in the European Renaissance. The history of medicine in context. Ashgate, Aldershot, Hampshire, England, 1999. ISBN 1-85928-361-6 (hardcover). viii + 289 pp. LCCN QM33.4 .F74 1999.

Freeman:2001:TNW

[Fre01] Michael Freeman. Tracks to a new world: railway excavation and the extension of geological knowledge in mid-nineteenth-century Britain. British Journal for the History of Science, 34(1):51–65, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028038.

Freeman:2004:VPT

[Fre04a] Michael J. Freeman. Victorians and the prehistoric: tracks to a lost world. Yale University Press, New Haven, CT, USA, 2004. ISBN 0-300-10334-4. 310 pp. LCCN QE705.G7

                                        F74 2004. URL http://www.
                                        loc.gov/catdir/bios/
                                        yale051/2004007555.html;

http://www.loc.gov/catdir/
description/yale051/
2004007555.html;   http://
www.loc.gov/catdir/toc/
ecip0416/2004007555.html.

French:2004:BM

[Fre04b] Steven French. Back matter.

British Journal for the History of Science, 37(4):500, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028668.

French:2004:BRH

[Fre04c] Steven French. Book review:

Helge Kragh, Quantum Generations: A History of Physics in the Twentieth Century. Princeton: Princeton University Press, 2002. Pp. xiv + 494. ISBN 0-691-09552-3. £13.95 (paperback). British Journal for the History of Science, 37(2): 220–221, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028348.

French:2004:BRS

[Fre04d] Steven French. Book review:

Steven Weinberg, Facing Up: Science and Its Cultural Adversaries. Cambridge, MA and London: Harvard University Press, 2001. Pp. xi + 283. ISBN 0-674-00647-X. £17.95, $26.00 (hardback). British Journal for the History of Science, 37(4):491– 492, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028666.

Freudenthal:2005:SMH

[Fre05] Gad Freudenthal. Science in the medieval Hebrew and Arabic traditions, volume CS803 of Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-86078-952-7. xix + 370 pp. LCCN Q124.97 .F74 2005.

Friedman:2000:PWC

[Fri00] Michael Friedman. A parting of the ways: Carnap, Cassirer, and Heidegger. Open Court, Chicago, IL, USA, 2000. ISBN 0-8126-9425-2 (paperback), 0-8126-9424-4 (cloth). xv + 175 pp. LCCN B808.5 .F76 2000.

Friedman:2001:DRK

[Fri01] Michael Friedman. Dynamics of reason: the 1999 Kant Lectures at Stanford University. CSLI Publications, Stanford, CA, USA, 2001. ISBN 1-57586-291-3, 1-57586-292-1 (paperback). xiv + 141 pp. LCCN Q175 .F892

                                        2001. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/2001017171.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy037/2001017171.html.

Friesen:2003:BRA

[Fri03a] John Friesen. Book review: Anita Guerrini, Obesity and Depression in the Enlightenment: The Life and Times of George Cheyne. Oklahoma Project for Discourse and Theory. Norman: University of Oklahoma Press, 2000. Pp. xx + 283. ISBN 0-8061-3201-9. $19.95 (paperback). British Journal for the History of Science, 36(4):481–482, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028221;   http://www.
jstor.org/stable/4028281.

Friesen:2003:BRN

[Fri03b] John Friesen. Book reviews: A NITA Guerrini, Obesity and Depression in the Enlightenment: The Life and Times of George Cheyne. Oklahoma Project for Discourse and Theory. Norman: University of Oklahoma Press, 2000. Pp. xx + 283. ISBN 0-8061-3201-9. US$19.95 (paperback). British Journal for the History of Science, 36(4):481– 482, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028221.

Friesen:2004:BRS

[Fri04] John Friesen. Book review: Scott Mandelbrote, Footprints of the Lion: Isaac Newton at Work. Exhibition at Cambridge University Library 9 October 2001–23 March 2002. Cambridge: Cambridge University Library, 2001. Pp. 142. ISBN 0-902205-58-7. £7.50 (paperback). British Journal for the History of Science, 37(3):349–350, September 2004.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028130;
                                        http://www.jstor.org/
                                        stable/4028432.

Frosh:2005:HJS

[Fro05]  Stephen  Frosh.

Hate and the ‘Jewish science’: antisemitism, Nazism And psychoanalysis. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2005. ISBN 1-4039-2170-9 (cloth). vi + 228 pp. LCCN BM538.P68 F76

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0619/2004062501-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0619/
                                        2004062501-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0619/
                                        2004062501-t.html.

Fyfe:2003:TS

[FS03] Aileen Fyfe and Paul Smith. Telling stories. British Journal for the History of Science, 36(4):471–476, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028217.

Frasca-Spada:2000:BSH

[FSJ00]  Marina  Frasca-

Spada and Nicholas Jardine, editors. Books and the sciences in history. Cambridge University Press, Cambridge, UK, 2000.

ISBN 0-521-65063-1 (cloth), 0-521-65939-6 (paperback). xiv + 438 pp. LCCN Q125 .B623

2000. URL http://www.loc.
gov/catdir/description/
cam0210/99087281.html;
http://www.loc.gov/catdir/
toc/cam021/99087281.html.

Fried:2001:APB

[FU01] Michael N. Fried and Sabetai Unguru. Apollonius of Perga’s Conica: text, context, subtext, volume 222 of Mnemosyne, bibliotheca classica Batava. Supplementum, 0169-8958. Brill, Leiden, The Netherlands, 2001. ISBN 90-04-11977-9. xii + 499 pp. LCCN QA31 .F75 2001.

Fullick:2000:MF

[Ful00a] Ann Fullick. Michael Faraday. Heinemann Library, Oxford, UK, 2000. ISBN 0-431-10443-3, 0-431-10455-7 (paperback). 48 pp. LCCN QC16.F2 F85 2001.

Fullmer:2000:YHD

[Ful00b]  June  Z.

Fullmer. Young Humphry Davy: the making of an experimental chemist, volume 237 of Memoirs of the American Philosophical Society, 0065-9738. American Philosophical Society, Philadelphia, PA, USA, 2000. ISBN 0-87169-237-6 (cloth). xvi + 385 pp. LCCN Q11 .P612 vol. 237; QD22.D3F85.

Fuller:2001:BRMa

[Ful01a]  Steve Fuller.  Book review:

Malachi Haim Hacohen, Karl Popper – The Formative Years, 1902–1945: Politics and Philosophy in Interwar Vienna. Cambridge: Cambridge University Press, 2001. Pp. xiii + 610. ISBN 0-521-47053-6. £35.00, $54.95 (hardback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028109.

Fuller:2001:BRMb

[Ful01b]  Steve Fuller.  Book review:

Michael Friedman, A Parting of the Ways: Carnap, Cassirer, and Heidegger. Chicago: Open Court Press, 2000. Pp. xv + 175. ISBN 0-8126-9425-2. $24.95 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028381.

Fuller:2001:BRZ

[Ful01c] Steve Fuller. Book review: Ziauddin Sardar, Thomas Kuhn and the Science Wars. Postmodern Encounters. Cambridge: Icon Books, 2000. Pp. 76. ISBN 1-84046-136-5. £2.99. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028047.

Fuller:2002:BRP

[Ful02] Steve Fuller. Book review: Peter Machamer, Marcello Pera and Aristides Baltas, Scientific Controversies: Philosophical and Historical Perspectives. New York and Oxford: Oxford University Press, 2000. Pp. x + 278. ISBN 0-19-511987-8. £32.50 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028408.

Fuller:2003:BRJ

[Ful03] Steve Fuller. Book reviews: Jay A. Labinger and Harry Collins, The One Culture? A Conversation about Science. Chicago and London: University of Chicago Press, 2001. Pp. xi + 329. ISBN 0-226-46723-6. £11.00, $17.00 (paperback). James Robert Brown, Who Rules in Science? An Opinionated Guide to the Wars. Cambridge, MA and London: Harvard University Press, 2001. Pp. xiii+236. ISBN 0-674-00652-6. £17.95, $26.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028306.

Fuller:2009:BRA

[Ful09a] Steve Fuller. Book reviews: Alan Sokal, Beyond the Hoax: Science, Philosophy and Culture. Oxford and New York: Oxford University Press, 2008. Pp. xxi + 465. ISBN 978-0-19-923920-7. £20.00 (hardback). Sophie Roux (ed.), Retours sur l’affaire Sokal. Paris: L’Harmattan, 2007. Pp. x+190. ISBN 978-2-296-02389-5. EUR17.50 (paperback). British Journal for the History of Science, 42(3):442–444, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fuller:2009:BRJ

[Ful09b] Steve Fuller. Book reviews: Joseph E. Harmon and Alan G. Gross (eds.), The Scientific Literature: A Guided Tour. Chicago and London: University of Chicago Press, 2007. Pp. xxiv + 327. ISBN 978-0-22631656-7. $72.50, £46.00 (hardback). ISBN 978-0-226-31656-7. $29.00, £18.50 (paperback). British Journal for the History of Science, 42(4):612–613, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fyfe:2002:BRD

[Fyf02a] Aileen Fyfe. Book review: David N. Livingstone, D. G. Hart and Mark A. Noll (eds.), Evangelicals and Science in Historical Perspective. Religion in America. Oxford and New York: Oxford University Press, 1999. Pp. vi + 351. ISBN 0-19-511557-0. £39.99 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028409.

Fyfe:2002:BRA

[Fyf02b] Aileen Fyfe. Book reviews: Andrew Hunter (ed.), Thornton and Tully’s Scientific Books, Libraries and Collectors: A Study of Bibliography and the Book Trade in Relation to the History of Science. Fourth Edition. Aldershot: Ashgate, 2000. Pp. xii + 405. ISBN 0-566-05481-7. £80.00 (hardback). Alain Besson (ed.), Thornton’s Medical Books, Libraries and Collectors: A Study of Bibliography and the Book Trade in Relation to the Medical Sciences. Third Edition. Aldershot: Gower, 1990. Pp. xxi+417. ISBN 0-566-05481-7. £65.00 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fyfe:2003:BRA

[Fyf03a] Aileen Fyfe. Book review: Arne Hessenbruch (ed.), Reader’s Guide to the History of Science. London and Chicago: Fitzroy Dearborn, 2000. Pp. xxix + 934. ISBN 1-884964-29-X. £95.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028307.

Fyfe:2003:BRW

[Fyf03b]  Aileen Fyfe.  Book review:

William J. Astore, Observing God: Thomas Dick, Evangelicalism, and Popular Science in Victorian Britain and America. Aldershot: Ashgate, 2001. Pp. ix + 303. ISBN 0-7546-0202-8. £45.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028308.

Fyfe:2004:SSE

[Fyf04] Aileen Fyfe. Science and salvation: evangelical popular science publishing in Victorian Britain. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-27647-3, 0-226-27648-1 (paperback). xiv + 325 pp. LCCN BL245 .F94 2004. URL http://

                                        www.loc.gov/catdir/bios/
                                        uchi051/2003017513.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi052/
                                        2003017513.html;   http://
                                        www.loc.gov/catdir/toc/
                                        ecip047/2003017513.html.

Fyfe:2005:BRJ

[Fyf05]  Aileen Fyfe.  Book review:

John L. Heilbron (ed.), The Oxford Companion to the History of Modern Science. Oxford and New York: Oxford University Press, 2003. Pp. xxviii + 941. ISBN 0-19-511229-6. £80.00 (hardback). British Journal for the History of Science, 38(3): 351–353, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028675.

Fyfe:2007:BRM

[Fyf07] Aileen Fyfe. Book review: Martin Daunton (ed.), The Organisation of Knowledge in Victorian Britain. Oxford: Oxford University Press, 2005. Pp.

424. ISBN 0-19-726326-7. £55.00 (hardback). British Journal for the History of Science, 40(3): 443–444, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500753.

Galison:1997:ILM

[Gal97] Peter Galison. Image and Logic:

a Material Culture of Microphysics. University of Chicago Press, Chicago, IL, USA, 1997. ISBN 0-226-27916-2 (cloth), 0-226-27917-0 (paperback). xxv + 955 pp. LCCN QC173.4.M5 G35

1997. URL http://www.loc.
gov/catdir/description/
uchi051/96052177.html;
http://www.loc.gov/catdir/
enhancements/fy0609/
96052177-b.html;
http://www.loc.gov/catdir/
toc/uchi051/96052177.html.

Gannett:2007:BRJ

[Gan07] Lisa Gannett. Book review:

Jenny Reardon, Race to the Finish: Identity and Governance in an Age of Genomics. Princeton and Oxford: Princeton University Press, 2005. Pp. 256. ISBN 0-691-11857-4. £11.95 (paperback). British Journal for the History of Science, 40(3): 462–464, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500765.

Garrett:2003:VTN

[Gar03a] Brian Garrett. Vitalism and teleology in the natural philosophy of Nehemiah Grew (1641–1712). British Journal for the History of Science, 36(1):63–81, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028293.

Garwood:2003:BRA

[Gar03b] Christine Garwood. Book reviews: Andrew Berry (ed.), Infinite Tropics: An Alfred Russel Wallace Anthology. With a preface by Stephen Jay Gould. London and New York: Verso, 2002. Pp. xvii + 430. ISBN 1-85984-652-1. £19.00, $27.00 (hardback). Jane R. Camerini (ed.), The Alfred Russel Wallace Reader: A Selection of Writings from the Field. Foreword by David Quammen. Center Books in Natural History. Baltimore and London: Johns Hopkins University Press, 2002. Pp. xix+221. ISBN 0-8018-6789-4. £13.00 (paperback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028309.

Gascoigne:1998:SSE

[Gas98a] John Gascoigne. Science in the Service of Empire: Joseph Banks, the British State and the Uses of Science in the Age of Revolution. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-55069-6. vii + 247 pp. LCCN Q127.G7 G38

1998. URL http://www.loc.
gov/catdir/description/
cam029/98192120.html;
http://www.loc.gov/catdir/
toc/cam027/98192120.html.

Gasman:1998:HMB

[Gas98b]  Daniel Gasman.  Haeckel’s

monism and the birth of fascist ideology, volume 33 of Studies in modern European history, 0893-6897. P. Lang, New York, NY, USA, 1998. ISBN 0-8204-4108-2. viii + 482 pp. LCCN JC481 .G3 1998.

Gascoigne:2009:RSN

[Gas09] John Gascoigne. The Royal Society, natural history and the peoples of the ‘New World(s)’, 1660–1800. British Journal for the History of Science, 42(4): 539–562, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gates:2002:NNA

[Gat02] Barbara T. Gates, editor. In nature’s name: an anthology of women’s writing and illustration, 1780–1930. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-28444-1 (hardcover), 0-226-28446-8 (paperback). xxvi + 673 pp. LCCN PR1111.N3 I5

2002. URL http://www.h-net.
org/review/hrev-a0c4u3-aa;
 http://www.h-net.org/
review/hrev-a0c6k2-aa;
http://www.loc.gov/catdir/
bios/uchi051/2001005012.
html;   http://www.loc.gov/
catdir/description/uchi051/
2001005012.html;   http://
www.loc.gov/catdir/toc/
uchi051/2001005012.html.

Gattei:2005:BRS

[Gat05] Stefano Gattei. Book review: Steve Fuller, Thomas Kuhn: A Philosophical History for Our Times. Chicago and London: University of Chicago Press, 2002. Pp. xvii + 472. ISBN 0-226-26896-9. £14.50, $22.50 (paperback). British Journal for the History of Science, 38(1): 121–122, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028601.

Gaukroger:2001:FBT

[Gau01] Stephen Gaukroger. Francis Bacon and the transformation of early-modern philosophy. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80154-0 (hardcover), 0-521-80536-8 (paperback). xii + 249 pp. LCCN B1198 .G38 2001.

URL  http://www.loc.gov/
catdir/description/cam021/
00063097.html;   http://
www.loc.gov/catdir/samples/
cam031/00063097.html;
http://www.loc.gov/catdir/
toc/cam024/00063097.html.

Gaukroger:2006:ESC

[Gau06] Stephen Gaukroger. The emergence of a scientific culture: science and the shaping of modernity, 1210– 1685. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-929644-8 (hardcover). ix + 563 pp. LCCN Q124.97 .G38

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0725/2007296427-d.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2007296427-t.html.

Gaukroger:2008:BRE

[Gau08] Stephen Gaukroger. Book reviews: Edward Grant, A History of Natural Philosophy: From the Ancient World to the Nineteenth Century. Cambridge: Cambridge University Press, 2007. Pp. xvi + 361. ISBN 978-0-521-86931-7. £40.00, $70.00 (hardback). ISBN 978-0-521-68957-1. £14.99, $24.99 (paperback). British Journal for the History of Science, 41(2):291–292, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gaukroger:2009:BRP

[Gau09] Stephen Gaukroger. Book review: Peter Harrison, The Fall of Man and the Foundations of Science. Cambridge: Cambridge University Press, 2007. Pp. xi + 300. ISBN 978-0-521-87559-2. £50.00 (hardback). British Journal for the History of Science, 42(2):294–296, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gay:2007:HIC

[Gay07] Hannah Gay. The history of Imperial College London, 1907– 2007: higher education and research in science, technology and medicine. Imperial College Press, New York, NY, USA, 2007. ISBN 1-86094-708-5, 1-86094-709-3 (paperback). xxvii + 825 pp. LCCN T173.L9345

                                        G39 2007. URL http://www.
                                        loc.gov/catdir/toc/fy0803/
                                        2006049718.html.

Gay:2002:SCS

[GB02] Hannah Gay and Anne Barrett. Should the cobbler stick to his last? Silvanus Phillips Thompson and the making of a scientific career. British Journal for the History of Science, 35 (2):151–186, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028179.

Gillies:2000:HWW

[GC00] James Gillies and R. Cailliau. How the Web was born: the story of the World Wide Web. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-286207-3. xii + 372 + 8 pp. LCCN TK5105.888 .G544

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0637/2001274436-d.html;
http://www.loc.gov/catdir/
enhancements/fy0637/
2001274436-t.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2001274436-b.html.

Gee:2000:OAA

[Gee00] Emma Gee. Ovid, Aratus, and Augustus: astronomy in Ovid’s Fasti. Cambridge classical studies. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-65187-5. xi + 226 pp. LCCN PA6519.F9 G44 2000. URL

http://www.loc.gov/catdir/
description/cam0210/
99030735.html;   http://
www.loc.gov/catdir/samples/
cam032/99030735.html;
http://www.loc.gov/catdir/
toc/cam023/99030735.html.

Gerovitch:2002:NCH

[Ger02] Slava Gerovitch. From newspeak to cyberspeak: a history of Soviet cybernetics. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-07232-7 (hardcover). xiv + 369 pp. LCCN Q305 .G47

2002. URL http://www.loc.
gov/catdir/toc/fy031/
2001056241.html.

Grattan-Guinness:2002:BRL

[GG02] I. Grattan-Guinness. Book review: Lorenzo Magnani, Philosophy and Geometry: Theoretical and Historical Issues. Western Ontario Series in Philosophy of Science. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. xix + 249. ISBN 0-7923-6933-5. £59.00, $88.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028190.

Grattan-Guinness:2003:BRR

[GG03a] I. Grattan-Guinness. Book review: E. Roy Weintraub, How Economics Became a Mathematical Science. Science and Cultural Theory. Durham, NC and London: Duke University Press, 2002. Pp. xiii + 313. ISBN 0-8223-2871-2. £14.50 (paperback). British Journal for the History of Science, 36(3):380– 381, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028172.

Grattan-Guinness:2003:BRM

[GG03b] Ivor Grattan-Guinness. Book review: M. Chisholm, Such Silver Currents: The Story of William and Lucy Clifford 1845– 1929. Cambridge: Lutterworth Press, 2002. Pp. x + 198. ISBN 0-7188-3017-2. £17.50 (paperback). British Journal for the History of Science, 36(1):

87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028313.

Grattan-Guinness:2004:BRB

[GG04] I. Grattan-Guinness. Book review: Benjamin H. Yandell, The Honors Class: Hilbert’s Problems and Their Solvers. Natick, MA: A. K. Peters, 2002. Pp. ix + 486. ISBN 1-56881-141-1. £28.00, EUR46.00, $39.00 (hardback). British Journal for the History of Science, 37(1): 112–113, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028268.

Grattan-Guinness:2006:BRA

[GG06a] Ivor Grattan-Guinness. Book review: Anita Burdman Feferman and Solomon Feferman, Alfred Tarski: Life and Logic. Cambridge: Cambridge University Press, 2004. Pp vi + 425. ISBN 0-521-8049-7. £23.00, $35.00 (hardback). British Journal for the History of Science, 39(3):469–470, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028507.

Grattan-Guinness:2006:BRH

[GG06b] Ivor Grattan-Guinness. Book review: Hans Niels Jahnke (ed.), A History of Analysis. History of Mathematics, 24. Providence, RI: American Mathematical Society and London Mathematical Society, 2003. Pp. xi + 422. ISBN 0-8218-2623-9. $89.00 (hardback). British Journal for the History of Science, 39(1): 126–127, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028554.

Grattan-Guinness:2009:BRJ

[GG09] Ivor Grattan-Guinness. Book review: Jeremy J. Gray and Karen H. Parshall (eds.), Episodes in the History of Modern Algebra (1800–1950). Providence, RI: American Mathematical Society and London: London Mathematical Society, 2007. Pp. vii + 336 pp. ISBN 978-0-8218-4343-7. $69.00 (hardback). British Journal for the History of Science, 42(2):304–305, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Grattan-Guinness:2005:LWW

[GGC^{+}05] I. Grattan-Guinness, Roger Cooke, et al., editors. Landmark writings in Western mathematics 1640–1940. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-444-50871-6 (cloth), 0-08-045744-4 (e-book). xvii + 1022 pp. LCCN QA21.

Ghiselin:2001:BRH

[Ghi01] Michael T. Ghiselin. Book review: Helena Gourko, Donald I. Williamson and Alfred I. Tauber (eds.), The Evolutionary Biology Papers of Elie Metchnikoff. Boston Studies in the Philosophy of Science, 212. Dordrecht, Boston and London: Kluwer Academic Publishers, 2000. Pp. x + 221. ISBN 0-7923-6067-2. £89.00 (hardback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028382.

Goodman:2003:GNC

[GHL03] Alan H. Goodman, Deborah Heath, and M. Susan Lindee, editors. Genetic nature/culture: anthropology and science beyond the two-culture divide. University of California Press, Berkeley, CA, USA, 2003. ISBN 0-520-23792-7, 0-520-23793-5 (paperback). LCCN GN289 .G455

2003. URL http://www.loc.
gov/catdir/bios/ucal052/
2002152222.html;
http://www.loc.gov/catdir/
description/ucal042/
2002152222.html;
http://www.loc.gov/catdir/
toc/fy042/2002152222.html.

Gillispie:2004:SPF

[Gil04] Charles Coulston Gillispie. Science and polity in France: the revolutionary and Napoleonic years. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-11541-9. viii + 751 pp. LCCN Q127.F8 G535 2004. URL http://www.

loc.gov/catdir/description/

                                        prin051/2003055451.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0737/
                                        2003055451-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        prin051/2003055451.html.

Gillispie:2008:BRM

[Gil08] Charles C. Gillispie. Book reviews: Maurice Crosland, The Language of Science: From the Vernacular to the Technical. Cambridge: Lutterworth Press, 2006. Pp. 128. ISBN 978-0-7188-3060-1. £12.50, $27.50 (paperback). Maurice Crosland, Scientific Institutions and Practice in France and Britain, c. 1700–1870. Aldershot: Ashgate, 2007. Pp. 228. ISBN 978-0-7546-5913-6. £60.00 (hardback). British Journal for the History of Science, 41(4):611–613, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gillispie:2009:BRP

[Gil09] Charles C. Gillispie. Book review: Patricia Fara, Science: A Four Thousand Year History. Oxford: Oxford University Press, 2009. Pp. xv + 408. ISBN 978-0-19-922689-4. £20.00 (hardback). British Journal for the History of Science, 42(4): 613–615, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gingerich:2002:ACC

[Gin02] Owen Gingerich. An annotated census of Copernicus’ De revolutionibus (Nuremberg, 1543 and Basel, 1566), volume 2 of Studia Copernicana. Brill, Leiden, The Netherlands, 2002. ISBN 90-04-11466-1. ISSN 0925-6806. xxxi + 402 pp. LCCN Z8192.5 .G55 2002; QB41.

Gingerich:2004:BRR

[Gin04]  Owen Gingerich.  Book re-

view: Rienk Vermij, The Calvinist Copernicans: The Reception of the New Astronomy in the Dutch Republic, 1575–1750. History of Science and Scholarship in the Netherlands, 1. Amsterdam: Koninklijke Nederlandse Akademie van Wetenschappen, 2002. Pp. x + 433. ISBN 90-6984-340-4. EUR49.00 (hardback). British Journal for the History of Science, 37(4): 471–472, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028649.

Gissis:2009:BRM

[Gis09] Snait B. Gissis. Book review:

Massimo Mazzotti (ed.), Knowledge as Social Order: Rethinking the Sociology of Barry Barnes. Aldershot: Ashgate, 2008. Pp. xi + 184. ISBN 978-0-7546-4863-5. £50.00 (hardcover). British Journal for the History of Science, 42(3):445–447, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Givens:2005:OIM

[Giv05] Jean A. (Jean Ann) Givens. Observation and image-making in Gothic art. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83031-1 (hardcover). xiv + 231 + 8 pp. LCCN N6310 .G55 2005. URL http://www.

                                        loc.gov/catdir/description/
                                        cam041/2004043582.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0732/
                                        2004043582-b.html;
                                        http://www.loc.gov/catdir/
                                        samples/cam051/2004043582.
                                        html;
                                        http://www.loc.gov/catdir/
                                        toc/cam041/2004043582.html.

Ghiselin:2000:CIN

[GL00] Michael T. Ghiselin and Alan E. Leviton, editors. Cultures and institutions of natural history: essays in the history and philosophy of science, volume 25 of Memoir / California Academy of Sciences. California Academy of Sciences, Los Angeles, CA, USA, 2000. ISBN 0-940228-48-3. 363 pp. LCCN Q11 .C17 vol. 25; QH15.

Glass:1997:VTM

[Gla97] I. S. (Ian S.) Glass. Victorian telescope makers: the lives and letters of Thomas and Howard Grubb. Institute of Physics Pub., Bristol, UK, 1997. ISBN 0-7503-0454-5. xiii + 279 pp. LCCN QB84.75.G78 G56 1997.

Glick:2006:BRJ

[Gli06] Thomas F. Glick. Book review:

Janet Browne, Charles Darwin: The Power of Place. Volume 2 of a Biography. London: Jonathan Cape, 2002. Pp. 591. ISBN 0-679-42932-8. £25.00 (hardback). British Journal for the History of Science, 39(2):299–301, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028625.

Goodrum:2009:QAC

[GO09] Matthew R. Goodrum and Cora Olson. The quest for an absolute chronology in human prehistory: anthropologists, chemists and the fluorine dating method in palaeoanthropology. British Journal for the History of Science, 42(1):95–114, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Golinski:1998:MNK

[Gol98] Jan Golinski. Making natural knowledge: constructivism and the history of science. Cambridge history of science. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-44471-3, 0-521-44913-8 (paperback). xiv + 236 pp. LCCN Q125 .G63 1998. URL http://www.

loc.gov/catdir/description/
cam028/97024028.html;
http://www.loc.gov/catdir/
toc/cam026/97024028.html.

Golinski:2001:EAT

[Gol01] Jan Golinski. ‘exquisite atmography’: theories of the world and experiences of the weather in a diary of 1703. British Journal for the History of Science, 34(2):149–171, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028008.

Golan:2004:LML

[Gol04] Tal Golan. Laws of men and laws of nature: the history of scientific expert testimony in England and America. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01286-0. viii + 325 pp. LCCN KD7521 .G65 2004.

Goldstein:2005:BRO

[Gol05] Bernard R. Goldstein. Book review: Owen Gingerich, An Annotated Census of Copernicus’ De revolutionibus (Nuremberg, 1543 and Basel, 1566). Studia Copernicana – Brill’s Series, 2. Leiden, Boston, MA and Köln: Brill, 2002. Pp. xxxi + 402. ISBN 90-04-11466-1. $132.00, EUR113.00 (hardback). British Journal for the History of Science, 38(1):104–105, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028587.

Golinski:2007:BWC

[Gol07] Jan Golinski. British weather and the climate of enlightenment. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-30205-9 (cloth). xv + 284 pp. LCCN QC989.G69

G66 2007. URL http://www.h-
net.org/review/hrev-a0f8a2-
aa;   http://www.loc.gov/
catdir/enhancements/fy0702/
2006031722-t.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006031722-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006031722-d.html.

Golinski:2009:BRC

[Gol09] Jan Golinski. Book review:

Charles W. J. Withers, Placing the Enlightenment: Thinking Geographically about the Age of Reason. Chicago and London: University of Chicago Press, 2007. Pp. xiii + 330. ISBN 978-0-226-90405-4. $45.00, £28.50 (hardback). British Journal for the History of Science, 42 (1):137–138, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gooday:2001:BRJ

[Goo01] Graeme Gooday. Book review:

John Ziman (ed.), Technological Innovation as an Evolutionary Process. Cambridge: Cambridge University Press, 2000. Pp. xvii + 379. ISBN 0-521-62361-8. £40.00, $64.95. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028015.

Gooday:2002:BRD

[Goo02a] Graeme Gooday. Book review: David A. Kirsch, The Electric Vehicle and the Burden of History. New Brunswick and London: Rutgers University Press, 2000. Pp. xiii + 291. ISBN 0-8135-2809. $20.00. British Journal for the History of Science, 35(2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028186.

Gooday:2002:BRE

[Goo02b] Graeme Gooday. Book review: Eric Higgs, Andrew Light and David Strong (eds.), Technology and the Good Life? Chicago and London: University of Chicago Press, 2000. Pp. xii + 392. ISBN 0-226-33387-6. £16.00, $25.00 (paperback). British Journal for the History of Science, 35 (1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028410.

Goodman:2002:BRJ

[Goo02c] David Goodman. Book review: José Chabás and Bernard R. Goldstein, Astronomy in the Iberian Peninsula: Abraham Zacut and the Transition from Manuscript to Print. Transactions of the American Philosophical Society, 90, Part 2. Philadelphia: American Philosophical Society, 2000. Pp. xii + 196. ISBN 0-87169-902-8 (paper). No price given. British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028187.

Goodman:2002:BRN

[Goo02d] David Goodman. Book review: Nicolau Copèrnic, De les Revolucions dels orbes celestes. Translated into Catalan by Enrique Rodríguez Galdeano and Víctor Navarro Brotóns with introduction and notes by Victor Navarro Brotóns. Clàssics de la Ciència, 3. Barcelona: Editorial Pòrtic, l’Institut d’Estudis Catalans and Eumo Editorial, 2000. Pp. l + 61. ISBN 84-7306-626-X. No price given. British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028188.

Goodrum:2002:MCA

[Goo02e] Matthew R. Goodrum. The meaning of ceraunia: archaeology, natural history and the interpretation of prehistoric stone artefacts in the eighteenth century. British Journal for the History of Science, 35(3):255– 269, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028123.

Gooday:2003:BRM

[Goo03]  Graeme Gooday.  Book re-

view: Michael Thad Allen and Gabrielle Hecht (eds.), Technologies of Power. Essays in Honor of Thomas Parke Hughes and Agatha Chipley Hughes. Cambridge, MA and London: MIT Press, 2001. Pp. xx + 339. ISBN 0-262-51124-X. £16.95 (paperback). British Journal for the History of Science, 36(1): 87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028310.

Gooday:2004:BRK

[Goo04a] Graeme Gooday. Book review: Ken Alder, The Measure of All Things: The Seven-Year Odyssey that Transformed the World. London: Little, Brown, 2002. Pp. xiv + 466. ISBN 0-316-85989-3. £15.99 (hardback). British Journal for the History of Science, 37(4):474–475, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028652.

Gooday:2004:MMA

[Goo04b] Graeme Gooday. The morals of measurement: accuracy, irony, and trust in late Victorian electrical practice. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-43098-4. xxv + 285 pp. LCCN TK275 .G66

2004. URL http://www.loc.
gov/catdir/description/
cam032/2003053291.html;
http://www.loc.gov/catdir/
samples/cam041/2003053291.
html;
http://www.loc.gov/catdir/
toc/cam032/2003053291.html.

Gooday:2005:BRP

[Goo05] Graeme Gooday. Book review:

Paola Bertucci and Giuliano Pancaldi (eds.), Electric Bodies: Episodes in the History of Medical Electricity. Bologna Studies in History of Science 9. Bologna: Università di Bologna, Dipartimento di Filosofia, 2001. Pp. 298. ISBN 88-900162-2-1. No price given (paperback). British Journal for the History of Science, 38(4):474–475, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028460.

Gooday:2006:BRT

[Goo06a] Graeme J. N. Gooday. Book review: Tal Golan, Laws of Men and Laws of Nature: The History of Scientific Expert Testimony in England and America. Cambridge, MA and London: Harvard University Press, 2004. Pp. viii + 336. ISBN 0-674-01286-0. £33.95, $52.50 (hardback). British Journal for the History of Science, 39(3):452– 454, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028497.

Gooding:2006:BRD

[Goo06b] David C. Gooding. Book reviews: Davis Baird, Thing Knowledge: A Philosophy of Scientific Instruments. Berkeley and London: University of California Press, 2004. Pp. xxi + 273. ISBN 0-520-23249-6. £42.95, $65.00 (hardback). Allan Franklin, Selectivity and Discord: Two Problems of Experiment. Pittsburgh: University of Pittsburgh Press, 2002. Pp. ix+290. ISBN 0-8229-4191-0. £31.50 (hardback). British Journal for the History of Science, 39(4):598–599, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028523.

Goodrum:2006:BRR

[Goo06c] Matthew R. Goodrum. Book review: Rosemary Sweet, Antiquaries: The Discovery of the Past in Eighteenth-Century Britain. London and New York: Hambledon and London Ltd., 2004. Pp. xxi + 473. ISBN 1-85285-309-3. £25.00 (hardback). British Journal for the History of Science, 39(3):448–449, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028494.

Gooday:2008:BRT

[Goo08a] Graeme Gooday. Book reviews: Theodore Arabatzis, Representing Electrons: A Biographical Approach to Theoretical Entities. Chicago and London: The University of Chicago Press, 2006. Pp. xiv + 295. ISBN 0-226-02420-2. $70.00, £44.50 (hardback). ISBN 0-226-02421-0. $28.00, £18.00 (paperback). British Journal for the History of Science, 41(1):149–151, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goodrum:2008:BRR

[Goo08b] Matthew R. Goodrum. Book review: Richard G. Delisle, Debating Humankind’s Place in Nature 1860–2000: The Nature of Paleoanthropology. Upper Saddle River, NJ: Pearson Prentice Hall, 2007. Pp. vii + 447. ISBN 0-13-1777390-9. $53.00 (paperback). British Journal for the History of Science, 41(3):451– 452, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gooday:2009:BRH

[Goo09a] Graeme J. N. Gooday. Book review: Harry Collins and Robert Evans, Rethinking Expertise. Chicago and London: University of Chicago Press, 2007. Pp. 160. ISBN 978-0-226-11360-9. $37.50 (hardback). British Journal for the History of Science, 42(3): 444–445, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goodrum:2009:BRA

[Goo09b] Matthew R. Goodrum. Book review: Ann Gibbons, The First Human: The Race to Discover Our Earliest Ancestors. New York: Knopf, 2007. Pp. ix + 303. ISBN 1-4000-7696-X. $14.95 (paperback). British Journal for the History of Science, 42(1):154– 155, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gorman:2002:BRI

[Gor02] Michael John Gorman. Book review: Ingrid D. Rowland, The Ecstatic Journey: Athanasius Kircher in Baroque Rome. With an introduction by F. Sherwood Rowland. Chicago: University of Chicago Library, 2000. Pp. xiii + 109. ISBN 0-94-305625-X. £9.00, $15.00 (paperback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028189.

Gordin:2007:BRB

[Gor07] Michael D. Gordin. Book review: When Physics Became King by Iwan Rhys Morus. Journal of Interdisciplinary History, 38(1): 95–96, ???? 2007. ISSN 0022-1953 (print), 1530-9169 (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4139676.

Gordin:2009:BRK

[Gor09] Michael D. Gordin. Book review: Klaus Hentschel, The Mental Aftermath: The Mentality of German Physicists, 1945–1949. Oxford and New York: Oxford University Press, 2007. Pp. vi + 205. ISBN 978-0-19-920566-0. £25.00 (hardback). British Journal for the History of Science, 42 (1):147–148, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Goschler:2000:WOB

[Gos00] Constantin Goschler, editor. Wissenschaft und Öffentlichkeit in Berlin, 1870–1930. (German) [Science and the public in Berlin, 1870– 1930]. F. Steiner, Stuttgart, Germany, 2000. ISBN 3-515-07778-2. 232 pp. LCCN DD866.8 .W59

2000. URL http://www.h-net.
org/review/hrev-a0c1y1-aa.

Gouk:1999:MSN

[Gou99] Penelope Gouk. Music, science, and natural magic in seventeenth-century England. Yale University Press, New Haven, CT, USA, 1999. ISBN 0-300-07383-6 (cloth). xii + 308 pp. LCCN ML3800 .G66 1999.

Gouk:2003:BRJ

[Gou03a] Penelope Gouk. Book review:

Jamie C. Kassler, Music, Science, Philosophy: Models in the Universe of Thought. Variorum Collected Studies Series, CS713. Aldershot: Ashgate, 2001. Pp. xvi + 301. ISBN 0-86078-862-8. £55.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028311.

Gould:2003:BRC

[Gou03b] Paula Gould. Book review: Catharine M. C. Haines with Helen M. Stevens, International Women in Science: A Biographical Dictionary to 1950. Santa Barbara, Denver and Oxford: ABC-CLIO, 2001. Pp. xix + 383. ISBN 1-57607-090-5. £44.95 (hardback). British Journal for the History of Science, 36 (2):231–233, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028235.

Gould:2003:BRR

[Gou03c] Paula Gould. Book review: Roddey Reid and Sharon Traweek (eds.), Doing Science+Culture: How Cultural and Interdisciplinary Studies are Changing the Way We Look at Science and Medicine. New York and London: Routledge, 2000. Pp. viii + 339. ISBN 0-415-92112-0. £16.99 (paperback). British Journal for the History of Science, 36(1): 87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028312.

Gouk:2006:BRJ

[Gou06] Penelope Gouk. Book review: Jamie C. Kassler, The Beginnings of the Modern Philosophy of Music in England: Francis North’s A Philosophical Essay of Music (1677) with Comments of Isaac Newton, Roger North and in the Philosophical Transactions. Aldershot: Ashgate, 2004. Pp. xiv + 243. ISBN 0-7546-0139-0. £40.00 (hardback). British Journal for the History of Science, 39(3):446–448, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028493.

Gouk:2009:BRP

[Gou09a] Penelope Gouk. Book review:

Pamela H. Smith and Benjamin Schmidt (eds.), Making Knowledge in Early Modern Europe: Practices, Objects, and Texts, 1400–1800. Chicago and London: University of Chicago Press, 2008. Pp. xi + 360. ISBN 978-226-76329-3. $28.00 (paperback). British Journal for the History of Science, 42(1):132– 133, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gouk:2009:BRS

[Gou09b] Penelope Gouk. Book review:

Stuart Clark, Vanities of the Eye: Vision in Early Modern European Culture. Oxford: Oxford University Press, 2007. Pp. xii + 415. ISBN 978-0-19-925013-4. £35.00 (hardback). British Journal for the History of Science, 42(4):603–604, December 2009.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Govoni:2002:PPS

[Gov02] Paola Govoni. Un pubblico per la scienza: la divulgazione scientifica nell’Italia in formazione. (Italian) [A public for science: the spreading of scientific information in Italy], volume 27 of Studi storici Carocci. Carocci, Roma, Italy, 1a edition, 2002. ISBN 88-430-2321-7. 351 + 16 pp. LCCN Q127.I8 G68 2002.

Gitelman:2003:NM

[GP03] Lisa Gitelman and Geoffrey B. Pingree, editors. New media, 1740–1915. Media in transition. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-07245-9 (hardcover). xxxiii + 271 pp. LCCN P90 .N5 2003. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy046/2002029542.html.

Gray:2007:EHM

[GP07] Jeremy Gray and Karen Hunger Parshall, editors. Episodes in the history of modern algebra (1800– 1950), volume 32 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4343-5. viii + 336 pp. LCCN QA151 .E65 2007.

Galison:2000:AFT

[GR00] Peter Galison and Alex Roland, editors. Atmospheric flight in the Twentieth Century, volume 3 of Archimedes. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6037-0. xvi + 383 pp. LCCN TL559 .A86 1999.

Gray:2000:HC

[Gra00]  Jeremy Gray.  The Hilbert

challenge. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850651-1. xii + 315 pp. LCCN QA29.H5 G739

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0611/2001270619-d.html;
http://www.loc.gov/catdir/
enhancements/fy0611/
2001270619-t.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2001270619-b.html.

Gray:2005:BRA

[Gra05] Jeremy Gray. Book review:

Andrew Warwick, Masters of Theory: Cambridge and the Rise of Mathematical Physics. Chicago and London: University of Chicago Press, 2003. Pp. xiv + 572. ISBN 0-226-87375-7. £20.50, $29.00 (paperback). British Journal for the History of Science, 38(3):372– 373, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028690.

Gray:2006:BRL

[Gra06] Jeremy Gray. Book review: Leo Corry, David Hilbert and the Axiomatization of Physics (1898– 1918). Archimedes New Studies in the History and Philosophy of Science and Technology. Dordrecht, Boston and London: Kluwer Academic Publishers, 2004. Pp. xvii + 513. ISBN 1-4020-2777-X. $179.00 (hardback). British Journal for the History of Science, 39(3):467– 468, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028506.

Graber:2007:INE

[Gra07a]  Frédéric Graber.  Inventing

needs: expertise and water supply in late eighteenth- and early nineteenth-century Paris. British Journal for the History of Science, 40(3):315–332, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500746.

Grant:2007:HNP

[Gra07b] Edward Grant. A history of natural philosophy: from the ancient world to the Nineteenth Century. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-86931-5, (paperback), 0-521-68957-0 (paperback). xiv + 361 pp. LCCN QC7 .G73

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0642/2006012920-d.html;
http://www.loc.gov/catdir/
enhancements/fy0642/
2006012920-t.html;
http://www.loc.gov/catdir/

enhancements/fy0729/
2006012920-b.html.

Grabiner:2008:BRD

[Gra08a] Judith V. Grabiner. Book review: Daniel J. Cohen, Equations from God: Pure Mathematics and Victorian Faith. Baltimore: Johns Hopkins University Press, 2007. Pp. x + 242. ISBN 978-0-8018-8553-2. £33.50 (hardback). British Journal for the History of Science, 41 (2):298–300, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Gray:2008:BRK

[Gra08b] Jeremy Gray. Book review:

Karen Hunger Parshall, James Joseph Sylvester: Jewish Mathematician in a Victorian World. Baltimore: Johns Hopkins University Press, 2006. Pp. xiii + 461. ISBN 0-8018-8291-5. £46.50 (hardback). British Journal for the History of Science, 41 (2):300–302, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Greene:1999:DDA

[Gre99] John C. Greene. Debating Darwin: adventures of a scholar. Regina Books, Claremont, CA, USA, 1999. ISBN 0-941690-85-7 (cloth). 288 pp. LCCN QH366.2 .G74 1999.

Greenwood:2002:BRS

[Gre02] Anna Greenwood. Book review: Saul Dubow (ed.), Science and Society in Southern Africa.

Studies in Imperialism. Manchester and New York: Manchester University Press, 2000. Pp. x + 241. ISBN 0-7190-5812-0. £45.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028191.

Gregory:2005:FHU

[Gre05] Jane Gregory. Fred Hoyle’s universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-850791-7. x + 406 + 16 pp. LCCN QB36.H75 G74

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0636/2005279804-d.html;
http://www.loc.gov/catdir/
enhancements/fy0636/
2005279804-t.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2005279804-b.html.

Givens:2006:VMM

[GRT06]  Jean  A.  (Jean  Ann)

Givens, Karen Reeds, and Alain Touwaide, editors. Visualizing medieval medicine and natural history, 1200–1550, volume 5 of AVISTA studies in the history of medieval technology, science and art. Ashgate, Aldershot, Hampshire, England, 2006. ISBN 0-7546-5296-3. xx + 278 pp. LCCN R141 .V57 2006. URL http://

www.loc.gov/catdir/toc/
ecip063/2005033204.html.

Grafton:1999:NPN

[GS99] Anthony Grafton and Nancy G.

Siraisi, editors. Natural particulars: nature and the disciplines in Renaissance Europe, Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-07193-2 (hardcover). LCCN QH81 .N3 1999.

Gavroglu:2002:PGQ

[GS02]  Kostas  Gavroglu

and Ana Simões. Preparing the ground for quantum chemistry in Great Britain: the work of the physicist R. H. Fowler and the chemist N. V. Sidgwick. British Journal for the History of Science, 35(2):187–212, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028180.

Guerrini:2000:ODE

[Gue00] Anita Guerrini. Obesity and depression in the Enlightenment: the life and times of George Cheyne, volume 3 of Oklahoma project for discourse and theory. Series for science and culture. University of Oklahoma Press, Norman, OK, USA, 2000. ISBN 0-8061-3201-9. xx + 283 pp. LCCN R489.C45 G84 1999.

Guicciardini:2004:BRJ

[Gui04] Niccolò Guicciardini. Book review: J. Christiaan Boudri, What was Mechanical about Mechanics: The Concept of Force between Metaphysics and Mechanics from Newton to Lagrange. Boston Studies in the Philosophy of Science, 224. Dordrecht, Boston and London: Kluwer Academic Publishers, 2002. Pp. xvi + 276. ISBN 1-4020-0233-5. £75.00, $112.00, EUR112.00 (hardback). British Journal for the History of Science, 37(4):473–474, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028651.

Guicciardini:2005:BRC

[Gui05] Niccolò Guicciardini. Book review: S. Chandrasekhar, Newton’s Principia for the Common Reader. Oxford: Clarendon Press, 2003. Pp. xxiii + 595. ISBN 0-19-852675-X. £49.95 (paperback). (First published in hardback, 1995). British Journal for the History of Science, 38(3):366–368, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028686.

Gourko:2000:EBP

[GWT00] Helena Gourko, Donald I. Williamson, and Alfred I. Tauber. The Evolutionary-biology Papers of Elie Metchnikoff, volume 212 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6067-2, 94-015-9381-7. ISSN 0068-0346. x + 221 pp. LCCN Q174 .B67 vol. 212; QH366.2.

URL  http://www.loc.gov/
catdir/enhancements/fy0821/
99049680-d.html;
http://www.loc.gov/catdir/
enhancements/fy0821/
99049680-t.html;   https://
link.springer.com/book/10.
1007/978-94-015-9381-6.

Grandin:2004:SIN

[GWW04] Karl Grandin, Nina Wormbs, and Sven Widmalm, editors. The science-industry nexus: history, policy, implications: Nobel Symposium 123. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2004. ISBN 0-88135-365-5. LCCN Q180.A3 N63 2002. URL

http://www.loc.gov/catdir/
toc/fy054/2004059031.html.

Hacohen:2000:KPF

[Hac00] Malachi Haim Hacohen. Karl Popper, the formative years, 1902–1945: politics and philosophy in interwar Vienna. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-47053-6. xiii + 610 pp. LCCN B1649.P64

H33 2000. URL http://www.
loc.gov/catdir/description/
cam0210/99047921.html;
http://www.loc.gov/catdir/
toc/cam026/99047921.html.

Hackett:2008:HST

[Hac08] Edward J. Hackett, editor. The handbook of science and technology studies. MIT Press, Cambridge, MA, USA, third edition, 2008. ISBN 0-262-08364-7 (hardcover). xiv + 1065 pp. LCCN Q158.5 .H36 2008. URL http:

//www.loc.gov/catdir/toc/
ecip078/2007000959.html.

Hadot:2006:VIE

[Had06] Pierre Hadot. The veil of Isis: an essay on the history of the idea of nature. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2006. ISBN 0-674-02316-1. xii + 399 pp. LCCN BD581 .H2813 2006.

Hafsteinsson:2004:BRM

[Haf04] Sigurjon Baldur Hafsteinsson. Book review: Michael Bravo and Sverker Sörlin (eds.), Narrating the Arctic: A Cultural History of Nordic Scientific Practices. Canton, MA: Science History Publications/USA, 2002. Pp. ix + 373. ISBN 0-88135-385-X. $39.95 (hardback). British Journal for the History of Science, 37(3): 364–365, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028144;
http://www.jstor.org/
stable/4028446.

Hafsteinsson:2008:BRH

[Haf08]  Sigurjon  Bal-

dur Hafsteinsson. Book reviews: Helen M. Rozwadowski, Fathoming the Ocean: The Discovery and Exploration of the Deep Sea. With a Foreword by Sylvia A. Earle. Cambridge, MA and London: The Belknap Press of Harvard University Press, 2005. Pp xii + 276. ISBN 0-674-01691. £16.95 (hardback). Helen M. Rozwadowski and David K. van Keuren (eds.), The Machine in Neptune’s Garden: Historical Perspectives on Technology and the Marine Environment. Sagamore Beach, CA: Science History Publications, 2004. Pp. 399. ISBN 0-88135-372-8. £24.99 (hardback). British Journal for the History of Science, 41 (1):119–120, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hagen:2004:BRR

[Hag04] Joel B. Hagen. Book review:

Robin L. Chazdon and T. C. Whitmore (eds.), Foundations of Tropical Forest Biology: Classic Papers with Commentaries. Chicago and London: University of Chicago Press, 2002. Pp. xvii + 862. ISBN 0-226-10225-4. £24.50, $35.00 (paperback). British Journal for the History of Science, 37(2):223, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028350.

Hahn:2005:PSL

[Hah05] Roger Hahn. Pierre Simon Laplace, 1749–1827: a determined scientist. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01892-3. x + 310 pp. LCCN Q143.L36 H35313 2005.

Haines:2001:IWS

[Hai01] Catharine M. C. Haines. International women in science: a biographical dictionary to 1950. ABC-CLIO, Santa Barbara, CA, USA, 2001. ISBN 1-57607-090-5. xix + 383 pp. LCCN Q141 .H2167 2001.

Hale:2002:BRC

[Hal02a] Piers J. Hale. Book review: Charles Elton, Animal Ecology. With New Introductory Material by Mathew A. Leibold and J. Timothy Wootton. Chicago and London: Chicago University Press, 2001. Pp. lvi + 209. ISBN 0-226-20639-4. £13.00, $18.00 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hall:2002:HOS

[Hal02b] Marie Boas Hall. Henry Oldenburg: shaping the Royal Society. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-851053-5. xii + 369 pp. LCCN Q143.O4 H35

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0611/2001053128-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0611/
                                        2001053128-t.html.

Hale:2003:BRP

[Hal03] Piers J. Hale. Book review: Peder Anker, Imperial Ecology: Environmental Order in the British Empire, 1895–1945. Cambridge, MA and London: Harvard University Press, 2001. Pp. vii + 343. ISBN 0-674-00595-3. £41.50 (hardback). British Journal for the History of Science, 36(2):248–250, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028246.

Hale:2004:BRD

[Hal04a] Piers J. Hale. Book review: Douglas R. Weiner, Models of Nature: Ecology, Conservation and Cultural Revolution in Soviet Russia. With a New Afterword. Pitt Series in Russian and East European Studies. Pittsburgh: University of Pittsburgh Press, 2000. Pp. xii + 324. ISBN 0-8229-5733-7. $17.95 (paperback). British Journal for the History of Science, 37(2):219–220, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028347.

Hale:2004:BRS

[Hal04b]  Piers J. Hale.  Book re-

view: Stuart McCook, States of Nature: Science, Agriculture, and Environment in the Spanish Caribbean, 1760–1940. Austin: University of Texas Press, 2002. Pp. xiv + 201. ISBN 0-292-75257-1. £17.50 (paperback). British Journal for the History of Science, 37(2): 215–216, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028344.

Hallyn:2004:SRS

[Hal04c] Fernand Hallyn. Les structures rhétoriques de la science: de Kepler à Maxwell. (French). [The rhetorical structures of science: from Kepler to Maxwell]. Des travaux. Seuil, Paris, France, 2004. ISBN 2-02-063249-7. 321 pp. LCCN Q174.8 .H35 2004.

Hale:2006:BRS

[Hal06] Piers J. Hale. Book review: Spencer R. Weart, The Discovery of Global Warming. New Histories of Science, Technology and Medicine. Cambridge, MA and London: Harvard University Press, 2003. Pp. xi + 228. ISBN 0-674-01157-0. £19.95 (hardback). British Journal for the History of Science, 39(1): 149–150, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028573.

Hamilton:2001:FIC

[Ham01a] James Hamilton, editor. Fields of influence: conjunctions of artists and scientists, 1815– 1860. University of Birmingham Press, Birmingham, AL, USA, 2001. ISBN 0-902459-10-5. xiii + 174 + 24 pp. LCCN N6767 .F44 2001.

Hamm:2001:UGC

[Ham01b] E. P. Hamm. Unpacking Goethe’s collections: the public and the private in naturalhistorical collecting. British Journal for the History of Science, 34(3):275–300, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028099.

Hammond:2003:CCP

[Ham03] Nicholas Hammond, editor. The Cambridge companion to Pascal. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-80924-X, 0-521-00611-2 (paperback). xvi + 287 pp. LCCN B1903 .C327

2003. URL http://www.loc.
gov/catdir/description/
cam032/2003273207.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2003273207-b.html;
http://www.loc.gov/catdir/
samples/cam034/2003273207.
html;
http://www.loc.gov/catdir/
toc/cam032/2003273207.html.

Hamlin:2007:BRD

[Ham07a] Christopher Hamlin.

Book review: David S. Barnes, The Great Stink of Paris and the Nineteenth-Century Struggle against Filth and Germs. Baltimore: Johns Hopkins University Press, 2006. Pp xi + 314. ISBN 0-8018-8349-0. £23.50 (hardback). British Journal for the History of Science, 40(3):

449–450, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500757.

Hammond:2007:BRN

[Ham07b] Andy Hammond. Book review: Nils Roll-Hansen, The Lysenko Effect: The Politics of Science. New York: Humanity Books, 2005. Pp. 335. ISBN 1-59102-262-2. £17.99 (paperback). British Journal for the History of Science, 40(2): 309–310, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500742.

Hamblin:2008:PWR

[Ham08a] Jacob Darwin Hamblin. Poison in the well: radioactive waste in the oceans at the dawn of the nuclear age. Rutgers University Press, New Brunswick, NJ, USA, 2008. ISBN 0-8135-4220-0 (hardcover). x + 311 pp. LCCN TD898 .H35 2008. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0715/2007015498.html.

Hamlin:2008:BRI

[Ham08b] Christopher Hamlin. Book review: Ian Burney, Poison, Detection, and the Victorian Imagination. Manchester: Manchester University Press, 2006. Pp. viii + 169. ISBN 0-7190-7376-6. £35.00 (hardback). British Journal for the History of Science, 41(2):285–286, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hannam:2007:BRE

[Han07a] James Hannam. Book reviews: Ernan Mcmullin (ed.), The Church and Galileo. Notre Dame, IN: University of Notre Dame Press. Pp. xiv + 392. ISBN 0-268-03484-2. $30.00 (paperback). Maurice A. Finocchiaro, Retrying Galileo, 1633– 1992. Berkeley, Los Angeles and London: University of Califonia Press, 2005. Pp. xii+485. ISBN 0-520-24261-0. $50.00, £32.50 (hardback). British Journal for the History of Science, 40 (2):285–287, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500726.

Hannam:2007:BRW

[Han07b] James Hannam. Book reviews:

William A. Wallace, Domingo de Soto and the Early Galileo: Essays on Intellectual History. Variorum Collected Studies Series. Aldershot: Ashgate, 2004. Pp. xiv + 340. ISBN 0-86078-964-0, $105.95 (hardback). Galileo Galilei, Le Operazioni del Compasso Geometrico et Militare. With a translation by Stillman Drake and an introduction by Filippo Camerota. Oakland, CA: Octavo, 2005. ISBN 1-891788-97-3, $35.00 (CD-ROM). British Journal for the History of Science, 40(1):133–135, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500698.

Harman:2000:BRC

[Har00a]  P.  M.  Harman.

Book review: Charles Coulston Gillispie, with the collaboration of Robert Fox and Ivor Grattan-Guinness. Pierre-Simon Laplace, 1749–1827. A Life in Exact Science. Princeton University Press: Princeton, New Jersey, 1998. Pp. xii + 322. ISBN 0-691-10850-0. $49.50, £35.00. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027932.

Harman:2000:BRE

[Har00b] P. M. Harman. Book review: Ezio Vailati, Leibniz and Clarke. A Study of Their Correspondence. Oxford and New York: Oxford University Press/OUPusa, 1997. Pp. xii + 250. ISBN 0-19-5113993. £35.00. British Journal for the History of Science, 33 (1):115–124, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028069.

Harrison:2000:MAM

[Har00c] Mark Harrison. From medical astrology to medical astronomy: sol-lunar and planetary theories of disease in British medicine, c. 1700–1850. British Journal for the History of Science, 33 (1):25–48, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028064.

Hardie:2001:BRE

[Har01a] Philip Hardie. Book review:

Emma Gee, Ovid, Aratus and Augustus: Astronomy in Ovid’s Fasti. Cambridge Classical Studies. Cambridge: Cambridge University Press, 2000. Pp. xi + 226. ISBN 0-521-65187-5. $54.95. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028016.

Harwood:2001:BRM

[Har01b] Jonathan Harwood. Book review: Mikuláš Teich, Bier, Wissenschaft und Wirtschaft in Deutschland 1800–1914: Ein Beitrag zur deutschen Industrialisierungsgeschichte. Vienna: Böhlau, 2000. Pp. 353. ISBN 3-205-99239-3. DM 69.80 (hardback). British Journal for the History of Science, 34(4):453– 481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028383.

Harwood:2001:BRU

[Har01c] Jonathan Harwood. Book review: Ulrich Marsch, Zwischen Wissenschaft und Wirtschaft: Industrieforschung in Deutschland und Grossbritannien 1880– 1936. Publications of the German Historical Institute, London, Volume 47. Paderborn: Ferdinand Schoeningh, 2000. Pp.

559. ISBN 3-506-72045-7. DM 168.00. British Journal for the History of Science, 34(1): 97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028048.

Harman:1990:SLP

[Har02a] P. M. (Peter Michael) Harman, editor. The scientific letters and papers of James Clerk Maxwell. Cambridge University Press, Cambridge, UK, 1990– 2002. ISBN 0-521-25625-9 (vol. 1), 0-521-25626-7 (vol. 2), 0-521-25627-5 (vol. 3). xxviii + 748 (vol. 1), xxx + 999 + 15 (vol. 2), xxvii + 932 pp. LCCN QC670 .M385 1990. URL http://www.

loc.gov/catdir/description/
cam023/89000452.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
89000452-b.html;
http://www.loc.gov/catdir/
toc/cam021/89000452.html.

Harris:2002:TCL

[Har02b] Henry Harris. Things come to life: spontaneous generation revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-851538-3. viii + 168 pp. LCCN QH325 .H35

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0636/2001054856-d.html;
http://www.loc.gov/catdir/
enhancements/fy0723/
2001054856-b.html;
http://www.loc.gov/catdir/
toc/fy033/2001054856.html.

Harrison:2002:BRD

[Har02c]  Mark Harrison.  Book re-

view: Douglas M. Haynes, Imperial Medicine: Patrick Manson and the Conquest of Tropical Disease. Philadelphia: University of Pennsylvania Press, 2001. Pp. 229. ISBN 0-8122-3598-3. £26.50, $37.50 (hardback). British Journal for the History of Science, 35(3):347– 379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Harwood:2002:BRR

[Har02d] Jonathan Harwood. Book review: Rainer Brömer, Uwe Hossfeld and Nicolaas A. Rupke (eds.), Evolutionsbiologie von Darwin bis heute. Verhandlungen zur Geschichte und Theorie der Biologie, 4. Berlin: VWB, 2000. Pp. 425. ISBN 3-86135-382-2. DM 68.00 (paperback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Harrington:2003:BRR

[Har03a] Ralph Harrington. Book review: R. Angus Buchanan, Brunel: The Life and Times of Isambard Kingdom Brunel. London and New York: Hambledon and London, 2002. Pp. xxiv + 294. ISBN 1-85285-331-X. £25.00 (hardback). British Journal for the History of Science, 36(3):370–371, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028163.

Hartcup:2003:ESS

[Har03b] Guy Hartcup. The effect of science on the Second World War. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2003. ISBN 1-4039-0643-2 (paperback), 0-333-67061-2 (hardcover). xv + 214 + 8 pp. LCCN D810.S2 H37

2003. URL http://www.loc.
gov/catdir/bios/hol052/
2002042819.html;
http://www.loc.gov/catdir/
description/hol032/
2002042819.html;
http://www.loc.gov/catdir/
toc/hol031/2002042819.html.

Harwood:2003:BRR

[Har03c] Jonathan Harwood. Book review: Roger J. Wood and Vitezslav Orel, Genetic Prehistory in Selective Breeding: A Prelude to Mendel. Oxford and New York: Oxford University Press, 2001. Pp. xvii + 323. ISBN 0-19-850584-1. £49.50 (hardback). British Journal for the History of Science, 36 (2):239–241, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028241.

Haraway:2004:HR

[Har04a] Donna Jeanne Haraway. The Haraway reader. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2004. ISBN 0-415-96688-4 (hardcover), 0-415-96689-2 (paperback). viii + 352 pp. LCCN HQ1190 .H364

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0651/2003005861-d.html;
http://www.loc.gov/catdir/
toc/ecip041/2003005861.
html.

Harrison:2004:DMW

[Har04b] Mark Harrison. Disease and the modern world: 1500 to the present day. Themes in history (Polity Press); Themes in history series. Polity, Cambridge, UK, 2004. ISBN 0-7456-2809-5, 0-7456-2810-9 (paperback). 270 pp. LCCN R131 .H247

2004. URL http://www.loc.
gov/catdir/toc/ecip047/
2003017371.html.

Harwood:2004:BRM

[Har04c] Jonathan Harwood. Book review: Margit Szöllösi-Janze (ed.), Science in the Third Reich. German Historical Perspectives, 12. Oxford and New York: Berg, 2001. Pp. vii + 289. ISBN 1-85973-421-9. £14.99 (paperback). British Journal for the History of Science, 37(2):227–229, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028355.

Harman:2005:BRI

[Har05a] P. M. Harman. Book review: Iwan Rhys Morus, Frankenstein’s Children: Electricity, Exhibition, and Experiment in Early-Nineteenth-Century London. Princeton, NJ: Princeton University Press, 1998. Pp. xiv + 324. ISBN 0-691-05952-7. $45.00, £32.50 (hardback). British Journal for the History of Science, 38(4):484–485, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028468.

Harwood:2005:TDA

[Har05b] Jonathan Harwood. Technology’s dilemma: agricultural colleges between science and practice in Germany, 1860–1934. P. Lang, Oxford, UK, 2005. ISBN 3-03910-299-0 (paperback), 0-8204-6987-4 (US: paperback). 288 pp. LCCN S535.G3 H37 2005.

Harwood:2006:BRD

[Har06] Jonathan Harwood. Book review: Deborah Fitzgerald, Every Farm a Factory: The Industrial Ideal in American Agriculture. Yale Agrarian Studies. New Haven and London: Yale University Press, 2003. Pp. xi + 241. ISBN 0-300-08813-2. £35.00 (hardback). British Journal for the History of Science, 39(1): 142–143, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028567.

Harkness:2007:JHE

[Har07a] Deborah E. Harkness. The Jewel house: Elizabethan London and the scientific revolution. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-11196-7. xviii + 349 pp. LCCN Q127.G4 H37

2007. URL http://www.h-net.
org/review/hrev-a0f5v3-aa;
 http://www.loc.gov/catdir/
enhancements/fy0833/
2007002683-b.html;
http://www.loc.gov/catdir/
enhancements/fy0833/
2007002683-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip079/2007002683.html.

Harrison:2007:FMF

[Har07b] Peter Harrison. The fall of man and the foundations of science. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87559-5. xi + 300 pp. LCCN BL245 .H355

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2008272575-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2008272575-t.html.

Harwood:2008:BRD

[Har08] Jonathan Harwood. Book review: David J. Frank and Jay Gabler, Reconstructing the University: Worldwide Shifts in Academia in the Twentieth Century. Stanford: Stanford University Press, 2006. Pp. xvii + 248. ISBN 0-8047-5376-8. $19.95 (paperback). British Journal for the History of Science, 41(1):147– 148, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Haskell:2009:LME

[Has09] Yasmin Haskell. The languages of melancholy in early modern England. British Journal for the History of Science, 42(2):275– 280, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hau:2003:CHB

[Hau03] Michael Hau. The cult of health and beauty in Germany: a social history, 1890–1930. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-31974-1 (cloth), 0-226-31976-8 (paperback). x + 286 pp. LCCN RA418.3.G3 H375

                                        2003. URL http://www.h-net.
                                        org/review/hrev-a0d2e7-aa;
                                         http://www.loc.gov/catdir/
                                        bios/uchi051/2002014839.
                                        html;   http://www.loc.gov/
                                        catdir/description/uchi051/
                                        2002014839.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy041/2002014839.html.

Haynes:2001:IMP

[Hay01] Douglas Melvin Haynes. Imperial medicine: Patrick Manson and the conquest of tropical disease. University of Pennsylvania Press, Philadelphia, 2001. ISBN 0-8122-3598-3 (cloth). 229 pp. LCCN RC962.G7 H39 2001.

Haycock:2002:WSS

[Hay02a] David Boyd Haycock. William Stukeley: science, religion, and archaeology in eighteenth-century England. Boydell Press, Woodbridge, Suffolk, UK, 2002. ISBN 0-85115-864-1. xiii + 290 pp. LCCN DA93.S85 H39 2002. URL

http://www.loc.gov/catdir/
toc/fy034/2001056099.html.

Hayward:2002:BRM

[Hay02b] Rhodri Hayward. Book review:

Malcolm Macmillan, An Odd Kind of Fame: Stories of Phineas Gage. Cambridge, MA and London: MIT Press, 2000. Pp. xiii + 562. ISBN 0-262-13363-6. £26.50 (hardback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hayward:2003:BRF

[Hay03] Rhodri Hayward. Book review:

E. Fuller Torrey and Judy Miller, The Invisible Plague: The Rise of Mental Illness from 1750 to the Present. New Brunswick and London: Rutgers University Press, 2001. Pp. xvi + 418. ISBN 0-8135-3003-2. $28.00 (hardback). British Journal for the History of Science, 36 (2):244–246, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028243.

Hayward:2006:BRA

[Hay06] Rhodri Hayward. Book review: Alex Owen, The Place of Enchantment: British Occultism and the Culture of the Modern. Chicago and London: University of Chicago Press, 2004. Pp. xiv + 335. $30.00, £21.50 (hardback). British Journal for the History of Science, 39(3): 459–461, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028501.

Hulin:2002:FES

[HB02]  Nicole  Hulin

and Bénédicte Bilodeau. Les femmes et l’enseignement scientifique. (French) [Women and the teaching of science]. Sciences, histoire, société, 1242-5087. Presses universitaires de France, Paris, France, 1. édition, 2002. ISBN 2-13-052659-4. xi + 227 pp. LCCN Q130 .H85 2002.

Heilbron:2003:OCH

[HBB^{+}03] J.  L.  Heilbron,

James Bartholomew, Jim Bennett, Frederic L. Holmes, Rachel Laudan, and Giuliano Pancaldi, editors. The Oxford companion to the history of modern science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-511229-6. xxviii + 941 pp. LCCN Q125 .O86

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0613/2002153783-d.html;
http://www.oxfordreference.
com/views/BOOK_SEARCH.html?
book=t124.

Hunter:1999:WRB

[HD00] Michael Cyril William Hunter and Edward Bradford Davis, editors. The works of Robert Boyle. Pickering and Chatto, London, UK, 1999–2000. ISBN 1-85196-109-7 (set), 1-85196-522-X (vol. 1–7 of set), 1-85196-523-8 (vol. 8–14 of set). ???? pp. LCCN Q155 .B667 1999.

Hoddeson:2002:TGL

[HD02]  Lillian  Hodde-

son and Vicki Daitch. True genius: the life and science of John Bardeen: the only winner of two Nobel Prizes in physics. Joseph Henry Press, Washington, DC, USA, 2002. ISBN 0-309-08408-3. xi + 467 pp. LCCN QC16.B27

H63 2002. URL http://www.
loc.gov/catdir/toc/fy037/
2002007967.html.

Healy:1999:PES

[Hea99] John F. Healy. Pliny the Elder on science and technology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-814687-6. xv + 467 pp. LCCN Q143.P64 H4

1999. URL http://www.loc.
gov/catdir/enhancements/
fy0606/99024316-d.html.

Hedrick:2008:RSS

[Hed08] Elizabeth Hedrick. Romancing the salve: Sir Kenelm Digby and the powder of sympathy. British Journal for the History of Science, 41(2):161–185, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heering:2008:EMR

[Hee08] Peter Heering. The enlightened microscope: re-enactment and analysis of projections with eighteenth-century solar microscopes. British Journal for the History of Science, 41(3):345– 367, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heilbron:2005:OGH

[Hei05] J. L. Heilbron, editor. The Oxford guide to the history of physics and astronomy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517198-5. xxi + 358 pp. LCCN QC7 .O94

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0639/2004020388-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0725/
                                        2004020388-b.html.

Heilbron:2007:BRPb

[Hei07a] John L. Heilbron. Book review: Peter Dear, The Intelligibility of Nature: How Science Makes Sense of the World. Chicago and London: University of Chicago Press, 2006. Pp. xii + 242. ISBN 0-226-13948-4. $27.50, £17.50 (hardback). British Journal for the History of Science, 40(4):599–601, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Heilbron:2007:BRPa

[Hei07b] John L. Heilbron. Book review: Peter J. Bowler and Iwan Rhys Morus, Making Modern Science: A Historical Survey. Chicago and London: University of Chicago Press, 2005. Pp. viii + 464. ISBN 0-226-0681-7. £17.50, $25.00 (paperback). British Journal for the History of Science, 40(1):118–119, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500688.

Heilbron:2008:BRW

[Hei08] John L. Heilbron. Book reviews: William R. Shea and Mariano Artigas, Galileo Observed: Science and the Politics of Belief. Sagamore Beach: Science History Publications/USA, 2006. Pp. xi + 212. ISBN 0-88135-356-6. $30.00 (hardback). Richard J. Blackwell, Behind the Scenes at Galileo’s Trial, Including the First English Translation of Melchior Inchofer’s Tractatus Syllepticus. Notre Dame: Notre Dame University Press, 2006. Pp. xiii+245. ISBN 0-268-02201-1. $35.00 (hardback). British Journal for the History of Science, 41(2):294–296, June 2008.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hellyer:2005:CPJ

[Hel05]  Marcus  Hellyer.

Catholic physics: Jesuit natural philosophy in early modern Germany. University of Notre Dame Press, Notre Dame, IN, USA, 2005. ISBN 0-268-03071-5 (cloth). xii + 336 pp. LCCN QC47.G3 H45 2005. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip053/2004024688.html.

Hempstead:2002:BRH

[Hem02a] Colin A. Hempstead. Book review: Hugh R. Slotten, Radio and Television Regulation: Broadcast Technology in the United States, 1920–1960. Baltimore and London: Johns Hopkins University Press, 2000. Pp. xv + 308. ISBN 0-8018-6450-X. £35.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028192.

Hempstead:2002:BRP

[Hem02b] Colin A. Hempstead. Book review: Peter Galison and Alex Roland (eds.), Atmospheric Flight in the Twentieth Century. Archimedes: New Studies in the History and Philosophy of Science and Technology,

3. Dordrecht, Boston and London: Kluwer Academic Publishers, 2000. Pp. xvi + 383. ISBN 0-7923-6037-0. £112.00 (hardback). British Journal for the History of Science, 35(3):347– 379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hempstead:2003:BRS

[Hem03] Colin A. Hempstead. Book review: Sungook Hong, Wireless: From Marconi’s Black-Box to the Audion. Transformations: Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2001. Pp. xv + 248. ISBN 0-262-08298-5. £23.95 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028314.

Hentschel:1996:PNS

[Hen96] Klaus Hentschel, editor. Physics and national socialism: an anthology of primary sources, volume 18 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 1996. ISBN 3-7643-5312-0 (Basel), 0-8176-5312-0 (Boston). ci + 406 + civ pp. LCCN Q127.G3 P48 1996.

Henry:2001:BRR

[Hen01]  John Henry.  Book review:

Robert Fox (ed.), Thomas Harriot: An Elizabethan Man of Science. Aldershot: Ashgate, 2000. Pp. xii + 317. ISBN 0-7546-0078-5. £47.50 (hardback). British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028110.

Henry:2002:BRF

[Hen02a] John Henry. Book review: Francis Bacon, The Instauratio Magna : Last Writings. Edited with introduction, notes, commentaries, and facing-page translations by Graham Rees. The Oxford Francis Bacon, XIII. Oxford: Clarendon Press, 2000. Pp. xcvi + 363. ISBN 0-19-818470-0. £80.00 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028411.

Hentschel:2002:MST

[Hen02b] Klaus Hentschel. Mapping the spectrum: techniques of visual representation in research and teaching. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850953-7 (hardcover). xii + 362 pp. LCCN QD95 .H47 2002.

                                        URL http://www.e-streams.
                                        com/es0511/es0511_2165.
                                        html;   http://www.loc.gov/
                                        catdir/enhancements/fy0613/
                                        2002072578-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy032/2002072578.html.

Henry:2003:BRD

[Hen03a] John Henry. Book review: Diderick Raven, Wolfgang Krohn and Robert S. Cohen (eds.), Edgar Zilsel: The Social Origins of Modern Science. Boston Studies in the Philosophy of Science,

200. Dordrecht, Boston and London: Kluwer Academic Publishers, 2000. Pp. lix + 267. ISBN 0-7923-6457-0. £89.00, $143.00 (hardback). British Journal for the History of Science, 36(4): 477–478, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028218;
http://www.jstor.org/
stable/4028278.

Hentschel:2003:BRU

[Hen03b] Klaus Hentschel. Book review: Ursula Klein (ed.), Tools and Modes of Representation in the Laboratory Sciences. Boston Studies in the Philosophy of Science, 222. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. xv + 259. ISBN 1-4020-0100-2. £59.00, $89.00, EUR95.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028315.

Henry:2004:BRPb

[Hen04a] John Henry. Book review: Paolo Rossi, Logic and the Art of Memory: The Quest for a Universal Language. Translated with an Introduction by Stephen Clucas, London: The Athlone Press, 2000. Pp. xxviii + 333. ISBN 0-485-11468-2. £50.00 (hardback). British Journal for the History of Science, 37(2):200–201, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028334.

Henry:2004:BRPa

[Hen04b] John Henry. Book review: Peter Dear, Revolutionizing the Sciences: European Knowledge and Its Ambitions, 1500–1700. Basingstoke: Palgrave, 2001. Pp. viii + 208. ISBN 0-333-71574-8. £14.99 (paperback). British Journal for the History of Science, 37(2):199–200, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028333.

Henry:2004:BRM

[Hen04c] John Henry. Book reviews: Malcolm Oster (ed.), Science in Europe, 1500–1800: A Primary Sources Reader. Basingstoke: Palgrave, 2002. Pp. xix + 282. ISBN 0-333-97002-0. £14.99 (paperback). Malcolm Oster (ed.), Science in Europe, 1500–1800: A Secondary Sources Reader. Basingstoke: Palgrave, 2001. Pp. xii+307. ISBN 0-333-97006-3. £14.99 (paperback). British Journal for the History of Science, 37(3):346–347, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028429.

Hentschel:2004:BRH

[Hen04d] Klaus Hentschel. Book review: Helen Ross and Cornelis Plug, The Mystery of the Moon Illusion: Exploring Size Perception. Oxford: Oxford University Press, 2002. Pp. ix + 277. ISBN 0-19-850862-X. £29.95 (hardback). British Journal for the History of Science, 37(2):233, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028360.

Hentschel:2004:BRM

[Hen04e]  Klaus Hentschel.  Book re-

view: Matthew R. Edwards (ed.), Pushing Gravity: New Perspectives on Le Sage’s Theory of Gravitation. Montreal: Apeiron, 2002. Pp. iv + 316. ISBN 0-9683689-7-2. $25.00 (paperback). British Journal for the History of Science, 37(2):234–235, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028361.

Henare:2005:MAI

[Hen05a] Amiria J. M. Henare. Museums, anthropology and imperial exchange. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83591-7. xix + 323 pp. LCCN GN36.N45 H46

2005. URL http://www.loc.

gov/catdir/description/
cam051/2004063436.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2004063436-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2004063436-b.html.

Henry:2005:BRL

[Hen05b] John Henry. Book reviews: Leonhart Fuchs, De historia stirpium commentarii insignes. With a Commentary by Karen Reeds. Octavo Digital Editions. Oakland: Octavo, 2003. ISBN 1-59110-051-8. £29.00, $30.00 (CD-ROM). Nicolaus Copernicus, De revolutionibus orbium coelestium libri VI. With a Commentary by Owen Gingerich. Octavo Digital Editions. Oakland: Octavo, 2003. ISBN 1-891788-14-0. £24.00, $40.00 (CD-ROM). Galileo Galilei, Siderius Nuncius. With a Commentary by Albert van Helden. Octavo Digital Editions. Oakland: Octavo, 2003. ISBN 1-891788-12-4. £15.00, $25.00 (CD-ROM). Robert Hooke, Micrographia. With a Commentary by Brian J. Ford. Octavo Digital Editions. Oakland: Octavo, 2003. ISBN 1-891788-02-7. £29.00, $30.00 (CD-ROM). Benjamin Franklin, Experiments and Observations on Electricity. With a Commentary by I. Bernard Cohen. Octavo Digital Editions. Oakland: Octavo, 2003. ISBN 1-891788-13-2. £23.00, $25.00 (CD-ROM). British Journal for the History of Science, 38(3):

361–362, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028683.

Henry:2006:BRO

[Hen06a] John Henry. Book review: Owen Gingerich, The Book Nobody Read: Chasing the Revolutions of Nicolaus Copernicus. London: William Heinemann, 2004. Pp. xii + 306. ISBN 0-434-01315-3. £12.99 (hardback). British Journal for the History of Science, 39(3):445–446, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028492.

Henry:2006:BRP

[Hen06b] John Henry. Book review:

Pamela H. Smith, The Body of the Artisan: Art and Experience in the Scientific Revolution. Chicago and London: University of Chicago Press, 2004. Pp. x + 367. ISBN 0-26-76399-4. £24.50, $35.00 (hardback). British Journal for the History of Science, 39(4):607–609, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028530.

Henry:2007:BRB

[Hen07a] John Henry. Book reviews: Bruce T. Moran, Distilling Knowledge: Alchemy, Chemistry, and the Scientific Revolution. Cambridge, MA and London: Harvard University Press, 2005. Pp. 210. ISBN 0-674-01495-2. £16.95 (hardback). Allen G. Debus (ed.), Alchemy and Early Modern Chemistry: Papers from Ambix. Huddersfield: Jeremy Mills (for the Society for the History of Alchemy and Chemistry), 2004. Pp. xv + 543. ISBN 0-9546484-1-2. £33.00, $60.00 (paperback). British Journal for the History of Science, 40(1): 130–132, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500696.

Hentschel:2007:MAM

[Hen07b] Klaus Hentschel. The mental aftermath: the mentality of German physicists 1945–1949. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920566-3 (hardcover). 205 pp. LCCN QC9.G3

H46 2007. URL http://www.
loc.gov/catdir/toc/fy0805/
2007280005.html.

Herbert:2001:VRR

[Her01] Christopher Herbert. Victorian relativity: radical thought and scientific discovery. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-32732-9, 0-226-32733-7 (paperback). xv + 302 pp. LCCN BD221 .H47

2001. URL http://www.loc.
gov/catdir/bios/uchi051/
00012177.html;
http://www.loc.gov/catdir/
description/uchi051/

00012177.html;
http://www.loc.gov/catdir/
toc/uchi051/00012177.html.

Heringman:2004:RRA

[Her04] Noah Heringman. Romantic rocks, aesthetic geology. Cornell University Press, Ithaca, NY, USA, 2004. ISBN 0-8014-4127-7 (cloth). xix + 304 pp. LCCN PR575.G46 H47

2004. URL http://www.loc.
gov/catdir/toc/ecip0410/
2003023065.html.

Herbert:2005:CDG

[Her05a] Sandra Herbert. Charles Darwin, geologist. Cornell University Press, Ithaca, NY, USA, 2005. ISBN 0-8014-4348-2. xx + 485 pp. LCCN QE22.D27 H47

2005. URL http://www.loc.
gov/catdir/toc/ecip057/
2005002690.html.

Hertler:2005:BRU

[Her05b] Christine Hertler. Book review: Uwe Hoßfeld, Lennart Olsson and Olaf Breidbach (eds.), Carl Gegenbaur and Evolutionary Morphology. Theory in Biosciences, 122, 2–3 (special issue). Jena: Urban & Fischer, 2003. Pp. 198. ISSN 1431-7613. No price given (paperback). British Journal for the History of Science, 38(2):235–236, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028710.

Herran:2006:SNI

[Her06] Néstor Herran. Spreading nucleonics: the Isotope School at the Atomic Energy Research Establishment, 1951–1967. British Journal for the History of Science, 39(4):569–586, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hessenbruch:2000:RGH

[Hes00] Arne Hessenbruch, editor. Reader’s guide to the history of science. Fitzroy Dearborn, London, UK, 2000. ISBN 1-884964-29-X. xxix + 934 pp. LCCN Q125 .R335

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0653/2001270888-d.html.

Hessenbruch:2001:BRS

[Hes01] Arne Hessenbruch. Book review: Sven Widmalm (ed.), Vetenskapsbärarna – Naturvetenskapen i det svenska samhället, 1880–1950. Hedemora: Gidlunds förlag, 1999. Pp. 368. ISBN 91-7844300-8. No price given. Gustav Holmberg, Reaching for the Stars: Studies in the History of Swedish Stellar and Nebular Astronomy, 1860–1940. Lund Studies in the History of Science and Ideas, 13. Lund: Ugglan, 1999. Pp. 243. ISBN 91-628-3837-7. £10.00 (paperback). British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028111.

Hessenbruch:2003:BRG

[Hes03] Arne Hessenbruch. Book review: G. I. Brown, Invisible Rays: The History of Radioactivity. Stroud: Sutton, 2002. Pp. viii + 248. ISBN 0-7509-2667-8. £19.99 (hardback). British Journal for the History of Science, 36(3):379–380, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028171.

Hessenbruch:2004:BRL

[Hes04] Arne Hessenbruch. Book review:

Lillian Hoddeson and Vicki Daitch, True Genius: The Life and Science of John Bardeen, the Only Winner of Two Nobel Prizes in Physics. Washington, DC: Joseph Henry Press, 2002. Pp. xi + 467. ISBN 0-309-08408-3. £20.95, $27.95 (hardback). British Journal for the History of Science, 37(2):230–231, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028358.

Harmon:2007:SLG

[HG07] Joseph E. Harmon and Alan G.

Gross, editors. The scientific literature: a guided tour. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-31655-6 (cloth), 0-226-31656-4 (paperback). xxiv + 327 pp. LCCN Q225.5 .S35

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0701/2006016547-t.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006016547-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006016547-d.html.

Hughes:2000:SEC

[HH00]  Agatha  C.  Hughes

and Thomas Parke Hughes, editors. Systems, experts, and computers: the systems approach in management and engineering, World War II and after. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-08285-3 (jacket). vi + 513 pp. LCCN TA168 .S887 2000. URL http:

//www.loc.gov/catdir/toc/
fy02/00021959.html.

Hamilton:2001:BDC

[HH01] Peter Hamilton and Roger Hargreaves. The beautiful and the damned: the creation of identity in nineteenth century photography. Lund Humphries, London, UK, 2001. ISBN 0-85331-821-2. 121 pp. LCCN TR15 .H295 2001. URL http://www.

loc.gov/catdir/toc/fy031/
2001088090.html.

Herschel:2003:CHA

[HH03] Caroline Lucretia Herschel and Michael A. Hoskin. Caroline Herschel’s autobiographies. Science History Publications, Cambridge, UK, 2003. ISBN 0-905193-06-7. 147 pp. LCCN QB36.H6 H47 2003.

Hughes:2001:TPE

[HHAH01] Thomas  Parke  Hughes,

Agatha C. Hughes, Michael Thad Allen, and Gabrielle Hecht, editors. Technologies of power: essays in honor of Thomas Parke Hughes and Agatha Chipley Hughes. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-01184-0 (hardcover), 0-262-51124-X (paperback). xx + 339 pp. LCCN T19 .T36 2001. URL http://

www.loc.gov/catdir/toc/
fy031/00048041.html.

Higton:2001:BRG

[Hig01] Hester Higton. Book review:

Gerard L’E. Turner, Elizabethan Instrument Makers: The Origins of the London Trade in Precision Instrument Making. Oxford: Oxford University Press, 2000. Pp. xiv + 305. ISBN 0-19-856566-6. £79.50 (hardback). British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028112.

Higton:2002:BRP

[Hig02] Hester Higton. Book review:

Peter de Clercq, Scientific Instruments: Originals and Imitations. The Mensing Collection. Proceedings of a Symposium, held at the Museum Boerhaave, Leiden, 15–16 October 1999. Leiden: Museum Boerhaave, 2000. Pp. 144. ISBN 90-6292-129-9. Dfl 75.00, ???30.00 (paperback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028193.

Higgitt:2003:BRP

[Hig03a] Rebekah Higgitt. Book review: Patricia Fara, Newton: The Making of Genius. London, Basingstoke and Oxford: Macmillan, 2002. Pp. xvi + 347. ISBN 0-333-90735-3. £20.00 (hardback). British Journal for the History of Science, 36(3): 368–370, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028162.

Higgitt:2003:NDB

[Hig03b]  Rebekah Higgitt.  ‘Newton

dépossédé!’ The British response to the Pascal forgeries of 1867. British Journal for the History of Science, 36(4): 437–453, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028215.

Higginson:2004:BRS

[Hig04a] Ian Higginson. Book review: Susan E. Lederer, Frankenstein: Penetrating the Secrets of Nature. New Brunswick, NJ and London: Rutgers University Press, 2002. Pp. ix + 78. ISBN 0-8135-3200-0. $30.00 (paperback). British Journal for the History of Science, 37(3):354–355, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028136;   http://www.
jstor.org/stable/4028438.

Higton:2004:BRM

[Hig04b] Hester Higton. Book review:

M. Barlow Pepin, The Emergence of the Telescope: Janssen, Lipperhey, and the Unknown Man. Duncanville, TX: T. Tauri Productions, 2002. Pp. 41. No ISBN. No price given (paperback). British Journal for the History of Science, 37(3):347– 348, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028430.

Higton:2004:PIM

[Hig04c] Hester Higton. Portrait of an instrument-maker: Wenceslaus Hollar’s engraving of Elias Allen. British Journal for the History of Science, 37(2):147–166, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028328.

Higton:2005:BRK

[Hig05] Hester Higton. Book review:

Koenraad van Cleempoel, A Catalogue Raisonné of Scientific Instruments from the Louvain School, 1530 to 1600. De Diversis Artibus: Collection of Studies from the International Academy of the History of Science, 65. Turnhout: Brepols, 2002. Pp. xii + 284. ISBN 2-503-51218-6. EUR75.00 (hardback). British Journal for the History of Science, 38(2):225–226, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028702.

Higton:2006:BRC

[Hig06]  Hester Higton.  Book re-

views: Catherine Eagleton, Jennifer Downes, Katherine Harloe, Boris Jardine, Nick Jardine and Adam Mosley, Instruments of Translation. Cambridge: Cambridge Latin Therapy Group and the Whipple Museum of the History of Science, 2003. Pp. 54. ISBN 0-906271-21-5. No price given (paperback). Patrick Boner and Catherine Eagleton (eds.), Instruments of Mystery. Cambridge: Cambridge Latin Therapy Group and the Whipple Museum of the History of Science, 2004. Pp. iv + 65. ISBN 0-906271-22-3. No price given (paperback). British Journal for the History of Science, 39(2):286–287, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028614.

Higton:2007:BRG

[Hig07] Hester Higton. Book reviews:

Gerard L’E. Turner, Renaissance Astrolabes and their Makers. Aldershot: Ashgate, 2003. Pp. xii + 294. ISBN 0-86078-903-9. £67.50, $124.95 (hardback). Koenraadvan Cleempoel, Astrolabes at Greenwich: A Catalogue of the Astrolabes in the National Maritime Museum, Greenwich. Oxford and Greenwich: Oxford University Press and the National Maritime Museum, 2005. Pp. ix+339. ISBN 0-19-853069-2. £99.50 (hardback). British Journal for the History of Science, 40 (1):129–130, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500695.

Higton:2008:BRB

[Hig08] Hester Higton. Book review:

Bart Grob and Hans Hooijmaijers (eds.), Who Needs Scientific Instruments? Conference on Scientific Instruments and their Users, 20-22 October 2005. Leiden: Museum Boerhaave, 2006. Pp. 272. ISBN 906292-158-2. No price given (paperback, includes CD-ROM). British Journal for the History of Science, 41(3): 458–459, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hilgartner:2000:SSE

[Hil00] Stephen Hilgartner. Science on stage: expert advice as public drama. Writing science. Stanford University Press, Stanford, CA, USA, 2000. ISBN 0-8047-3645-6, 0-8047-3646-4 (paperback). xi + 213 pp. LCCN Q127.U5 H55

                                        2000. URL http://www.loc.
                                        gov/catdir/description/
                                        cam021/00055706.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0710/
                                        00055706-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam027/00055706.html.

Hills:2002:JW

[Hil05] Richard Leslie Hills. James Watt. Landmark collector’s library. Landmark Pub., Ashbourne, England, 2002–2005. ISBN 1-84306-045-0 (vol. 1), 1-84306-046-9 (vol. 2). 480 + 16 (vol. 1), 256 (vol. 2) pp. LCCN TA140.W3 H55 2002.

Hunter:2001:OLM

[HK01] Michael Cyril William Hunter and Robert Kirk. The occult laboratory: magic, science, and second sight in late Seventeenth-Century Scotland. Boydell Press, Woodbridge, Suffolk, UK, 2001. ISBN 0-85115-801-3. vii + 247 pp. LCCN BF1434.S3 H86

                                        2001. URL http://www.h-net.
                                        org/review/hrev-a0c0u8-aa.

Holmes:2000:IEH

[HL00] Frederic Lawrence Holmes and Trevor Harvey Levere, editors. Instruments and experimentation in the history of chemistry. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-08282-9 (hardcover). xxi + 415 pp. LCCN QD53 .I57 2000.

Higgs:2000:TGL

[HLS00] Eric Higgs, Andrew Light, and David Strong, editors. Technology and the good life? University of Chicago Press, Chicago, IL, USA, 2000. ISBN 0-226-33386-8 (cloth), 0-226-33387-6 (paper). xii + 392 pp. LCCN T14 .T386

2000. URL http://www.loc.
gov/catdir/bios/uchi051/
00008237.html;
http://www.loc.gov/catdir/
description/uchi051/
00008237.html;
http://www.loc.gov/catdir/
toc/fy02/00008237.html.

Harman:2002:CSM

[HM02]  P.  M.

(Peter Michael) Harman and Simon Mitton, editors. Cambridge scientific minds. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78100-0, 0-521-78612-6 (paperback). viii + 343 pp. LCCN Q141 .C1283

2002. URL http://www.loc.
gov/catdir/description/
cam021/2001025664.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2001025664-b.html;
http://www.loc.gov/catdir/
samples/cam033/2001025664.
html;
http://www.loc.gov/catdir/
toc/cam021/2001025664.html.

Hobby:2006:BRC

[Hob06] Elaine Hobby. Book review:

Caroline Bicks, Midwiving Subjects in Shakespeare’s England. Women and Gender in the Early Modern World. Aldershot: Ashgate, 2003. Pp. xii + 212. ISBN 0-7546-0938-3. £40.00 (hardback). British Journal for the History of Science, 39(1): 130–131, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028557.

Hochadel:2003:OWE

[Hoc03] Oliver Hochadel. Öffentliche Wissenschaft: Elektrizität in der deutschen Aufklärung. (German) [Public science: Electricity in the German explanation]. Wallstein, Göttingen, Germany, 2003. ISBN 3-89244-629-6. 364 pp. LCCN QC507 .H63 2003.

Hochadel:2008:BRM

[Hoc08]  Oliver Hochadel.  Book re-

view: Michael R. Lynn, Popular Science and Public Opinion in Eighteenth-Century France. Studies in Early Modern European History. Manchester: Manchester University Press, 2006. Pp. ix + 177. ISBN 0-7190-7373-1. £50.00. British Journal for the History of Science, 41(1):143– 144, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hochadel:2009:BRF

[Hoc09] Oliver Hochadel. Book review: Florian Charvolin, André Micoud and Lynn K. Nyhart (eds.), Des Sciences citoyennes? La Question de l’amateur dans les sciences naturalistes. La Tour d’Aigues: Editions de l’Aube, 2007. Pp. 254. ISBN 9-782752-602305. EUR20.00. British Journal for the History of Science, 42 (1):140–142, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hodge:2001:BRF

[Hod01]  Jon Hodge.  Book review:

Frederick Burkhardt, Duncan M. Porter, Joy Harvey and Jonathan R. T Opham (eds.), The Correspondence of Charles Darwin, Volume 10, 1862. Cambridge: Cambridge University Press, 1997. Pp. xl + 936. ISBN 0-521-59032-9. £50.00. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028049.

Hodge:2006:BRF

[Hod06a] Jonathan Hodge. Book reviews:

Frederick Burkhardt, Duncan M. Porter et al. (eds.), The Correspondence of Charles Darwin, Volume 12: 1864. Cambridge: Cambridge University Press, 2001. Pp. xl + 694. ISBN 0-521-59034-5. £55.00 (hardback). Volume 13: 1865. With Supplement to the Correspondence 1822–1864. Cambridge: Cambridge University Press, 2002. Pp. xl+695. ISBN 0-521-82413-3. £65.00 (hardback). Volume 14: 1866. Cambridge: Cambridge University Press, 2004. Pp. xl+655. ISBN 0-521-84459-2. £75.00 (hardback). British Journal for the History of Science, 39(2):301–302, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028626.

Hodges:2006:BRB

[Hod06b] Andrew Hodges. Book reviews: B. Jack Copeland (ed.), The Essential Turing: The Ideas that Gave Birth to the Computer Age. Oxford: Clarendon Press, 2004. Pp. viii + 613. ISBN 0-19-825079-7. £50.00 (hardback). ISBN 0-19-825080-0. £14.99 (paperback). British Journal for the History of Science, 39(3): 470–471, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028508.

Hoffmann:2007:CDF

[Hof07]  Christoph Hoffmann.  Con-

stant differences: Friedrich Wilhelm Bessel, the concept of the observer in early nineteenthcentury practical astronomy and the history of the personal equation. British Journal for the History of Science, 40(3):333– 365, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500747.

Holmberg:1999:RSS

[Hol99] Gustav Holmberg. Reaching for the stars: studies in the history of Swedish stellar and nebular astronomy, 1860–1940, volume 13 of Ugglan, 1102-4313. Lund University, Lund, Sweden, 1999. ISBN 91-628-3837-7. 243 pp. LCCN QB33.S8 H65 1999.

Holton:2000:EHO

[Hol00] Gerald Holton. Einstein, history, and other passions: the rebellion against science at the end of the Twentieth Century. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00433-7. xii + 240 + 4 pp. LCCN ????

Holmes:2008:PSN

[Hol08] Andrew R. Holmes. Presbyterians and science in the north of Ireland before 1874. British Journal for the History of Science, 41(4):541–565, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Homes:2009:ARH

[Hom09] Caitlin Homes. The Astronomer Royal, the Hydrographer and the time ball: collaborations in time signalling 1850–1910. British Journal for the History of Science, 42(3):381–406, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hong:2001:WMB

[Hon01] Sungook Hong. Wireless: from Marconi’s black-box to the audion. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-08298-5 (hardcover). xiv + 248 pp. LCCN TK6547 .H66

2001. URL http://www.loc.
gov/catdir/toc/fy031/
2001030636.html.

Hook:2002:PSD

[Hoo02] Ernest B. Hook, editor. Prematurity in scientific discovery: on resistance and neglect. University of California Press, Berkeley, CA, USA, 2002. ISBN 0-520-23106-6 (cloth), 0-520-23107-4 (paper). xx + 378 pp. LCCN Q180.55.D57 P74 2002. URL

http://www.loc.gov/catdir/
bios/ucal052/2001004245.
html;   http://www.loc.gov/
catdir/description/ucal042/
2001004245.html;   http://
www.loc.gov/catdir/toc/
ucal041/2001004245.html.

Hopwood:2001:BRH

[Hop01] Nick Hopwood. Book review: Heike Menz, Martin Heinrich Rathke (1793–1860): Ein Embryologe des 19. Jahrhunderts. Acta Biohistorica, 7. Marburg: Basilisken-Presse, 2000. Pp. 280. ISBN 3-925347-59-3. DM 78.00 (hardback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028384.

Hopwood:2002:EWM

[Hop02] Nick Hopwood, editor. Embryos in wax: models from the Ziegler studio. Whipple Museum of the History of Science, University of Cambridge, Cambridge, UK, 2002. ISBN 0-906271-18-5 (paperback). ix + 206 pp. LCCN ????

Hopwood:2004:BRA

[Hop04] Nick Hopwood. Book review: Andreas W. Daum, Wissenschaftspopularisierung im 19. Jahrhundert. Bürgerliche Kultur, naturwissenschaftliche Bildung und die deutsche Öffentlichkeit, 1848– 1914. Second edition. Munich: Oldenbourg, 2002. Pp. xii + 619. ISBN 3-486-56551-6. EUR59.80 (paperback). British Journal for the History of Science, 37(3): 357–358, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028138;
http://www.jstor.org/
stable/4028440.

Horoldt:1978:JSV

[Hör78] Dietrich Höroldt, editor. 1000 Jahre Stift Vilich: 978–1978: Beiträge zur Geschichte und Gegenwart von Stift und Ort Vilich. (German). [1000 years of the Vilich Foundation, 978–1978: Contributions to history and the present time ???]. Röhrscheid, Bonn, Germany, 1978. ISBN 3-7928-0412-3. 208 pp. LCCN BX4630.V54 S757.

Horrocks:2000:PPP

[Hor00] Sally M. Horrocks. A promising pioneer profession? Women in industrial chemistry in interwar Britain. British Journal for the History of Science, 33(3): 351–367, September 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027957.

Horrocks:2003:BRJ

[Hor03] Sally M. Horrocks. Book review: John Peyton, Solly Zuckerman: A Scientist out of the Ordinary. With a foreword by Roy Jenkins. London: John Murray, 2001. Pp. xviii + 252. ISBN 0-7195-6283-X. £22.50 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028316.

Horrocks:2007:ISC

[Hor07] Sally M. Horrocks. The internationalization of science in a commercial context: research and development by overseas multinationals in Britain before the mid-1970s. British Journal for the History of Science, 40 (2):227–250, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500718.

Hostetler:2001:QCE

[Hos01] Laura Hostetler. Qing colonial enterprise: ethnography and cartography in early modern China. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-35420-2. xx + 257 + 16 pp. LCCN GN635.C5 H67

2001. URL http://www.loc.
gov/catdir/bios/uchi051/
00010974.html;
http://www.loc.gov/catdir/
description/uchi052/
00010974.html;
http://www.loc.gov/catdir/
toc/uchi051/00010974.html.

Hoskin:2003:HPV

[Hos03] Michael A. Hoskin. The Herschel partnership: as viewed by Caroline. Science History Publications, Cambridge, UK, 2003. ISBN 0-905193-05-9. 182 pp. LCCN QB36.H6 H69 2003.

Houck:2006:HBW

[Hou06]  Judith  A.

(Judith Anne) Houck. Hot and bothered: women, medicine, and menopause in modern America. Harvard University Press, Cambridge, MA, USA, 2006. ISBN 0-674-01896-6. xii + 328 pp. LCCN RG186 .H7792 2006.

Houck:2007:BRR

[Hou07] Judith A. Houck. Book review: Robert Darby, A Surgical Temptation: The Demonization of the Foreskin and the Rise of Circumcision in Britain. Chicago and London: University of Chicago Press, 2005. Pp. 368. ISBN 978-0-226-13645-5. $35.00 (hardback). British Journal for the History of Science, 40(3): 450–452, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500758.

Howsam:2000:ESN

[How00] Leslie Howsam. An experiment with science for the nineteenthcentury book trade: the International Scientific Series. British Journal for the History of Science, 33(2):187–207, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027922.

Hoyrup:2001:BRS

[Høy01] Jens Høyrup. Book review: Serafina Cuomo, Pappus of Alexandria and the Mathematics of Late Antiquity. Cambridge Classical Studies. Cambridge: Cambridge University Press, 2000. Pp. ix + 234. ISBN 0-521-64211-6. $59.95. British Journal for the History of Science, 34(2):233–250, June 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028017.

Hoyrup:2005:BRJ

[Høy05] Jens Høyrup. Book review: Jürgen Schönbeck, Euklid. Vita Mathematica, 12. Basel, Boston and Berlin: Birkhäuser, 2003. Pp. x + 264. ISBN 3-7643-6584-6. EUR82.24, SFr 128.00 (hardback). British Journal for the History of Science, 38 (2):223–225, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028701.

Hilaire-Perez:2000:LTA

[HP00]  Liliane  Hilaire-

Pérez. L’invention technique au siècle des Lumières. (French) [The technical invention of the Century of Lights]. Evolution de l’humanité, 0755-1843. Albin Michel, Paris, France, 2000. ISBN 2-226-11537-4. 443 pp. LCCN T26.F8 H55 2000.

Hilaire-Perez:2002:DVA

[HP02] Liliane Hilaire-Pérez. Diderot’s views on artists’ and inventors’ rights: invention, imitation and reputation. British Journal for the History of Science, 35(2):129–150, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028178.

Henry:1972:PJH

[HRR08] Joseph Henry, Nathan Reingold, and Marc Rothenberg. The papers of Joseph Henry. Smithsonian Institution Press, Washington, DC, USA, 1972–2008. ISBN 0-87474-123-8 (vol. 1), 0-87474-164-5 (vol. 2), 0-87474-174-2 (vol. 3), 0-87474-792-9 (vol. 4), 0-87474-793-7 (vol. 5), 1-56098-112-1 (vol. 6), 1-56098-533-X (vol. 7), 1-56098-891-6 (vol. 8), 1-881353-63-9 (vol. 9), 1-881353-58-2 (vol. 10), 0-88135-390-6 (vol. 11), 0-88135-391-4 (vol. 12). xxxix + 496 (vol. 1), xxxix + 524 (vol 2.), xxxiv + 585 (vol. 3), xli + 662 (vol. 6) pp. LCCN Q113 .H43.

Heidelberger:1998:EEV

[HS98]  Michael  Hei-

delberger and Friedrich Steinle, editors. Experimental essays = Versuche zum Experiment, volume 3 of Interdisziplinäre Studien = Interdisciplinary studies / ZIF. Nomos, Baden-Baden, Germany, 1998. ISBN 3-7890-5367-8. xiii + 408 pp. LCCN Q125 .E87 1998.

Howard:2000:EFY

[HS00]  Don  Howard  and

John J. Stachel, editors. Einstein: the formative years, 1879– 1909, volume 8 of Einstein studies. Birkhäuser Verlag, Basel, Switzerland, 2000. ISBN 0-8176-4030-4, 3-7643-4030-4. viii + 258 pp. LCCN QC16.E5 E524

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0812/98006996-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0812/
                                        98006996-t.html.

Heidelberger:2002:HPS

[HS02]  Michael  Heidelberger

and Friedrich Stadler, editors. History of philosophy of science: new trends and perspectives, volume 9 of Vienna Circle Institute yearbook 0929-6328. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0509-1. x + 440 pp. LCCN

???? URL http://www.loc.
gov/catdir/enhancements/
fy0826/2002284609-d.html;
http://www.loc.gov/catdir/
enhancements/fy0829/
2002284609-t.html.

Hsu:2006:BRG

[Hsu06] Elisabeth Hsu. Book review:

G. E. R. Lloyd, Ancient Worlds, Modern Reflections: Philosophical Perspectives on Greek and Chinese Science and Culture. Oxford: Clarendon Press, 2004. Pp. ix + 222. ISBN 0-19-927016-3. £25.00 (hardback). British Journal for the History of Science, 39(4):602–603, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028526.

Huerta:2003:GDJ

[Hue03] Robert D. Huerta. Giants of Delft: Johannes Vermeer and the natural philosophers: the parallel search for knowledge during the age of discovery. Bucknell University Press, London, UK and Lewisburg, PA, USA, 2003. ISBN 0-8387-5538-0. 156 pp. LCCN N72.S3 H84 2003. URL

http://www.loc.gov/catdir/
toc/fy037/2002151834.html.

Hughes:2000:BRV

[Hug00] Jeff Hughes. Book reviews: V. Y. Frenkel, Yakov Ilich Frenkel. His Work, Life and Letters. Basel, Boston and Berlin: Birkhauser, 1996. Pp. viii + 323. ISBN 3-7643-2741-3 (Basel); ISBN 0-8176-2741-3 (Boston). DM198 (no sterling price given). British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027933.

Hughes:2001:BRJ

[Hug01]  Jeff Hughes.  Book review:

Jack Morrell, Science at Oxford, 1914–1939: Transforming an Arts University. Oxford: Clarendon Press, 1997. Pp. xx + 473. ISBN 0-19-820657-7. £55.00. British Journal for the History of Science, 34 (2):233–250, June 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028018.

Hughes:2004:DBR

[Hug04a] J. Hughes. Deconstructing the bomb: recent perspectives on nuclear history. British Journal for the History of Science, 37(4):455–464, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028151;
                                        http://www.jstor.org/
                                        stable/4028643.

Hughes:2004:HBW

[Hug04b] Thomas Parke Hughes. Humanbuilt world: how to think about technology and culture. Science.culture. University of Chicago Press, Chicago, IL, USA, 2004. ISBN 0-226-35933-6. xii + 223 pp. LCCN T14.5 .H84 2004. URL http://www.

loc.gov/catdir/bios/
uchi051/2003018426.html;
http://www.loc.gov/catdir/
description/uchi052/
2003018426.html;   http://
www.loc.gov/catdir/toc/
ecip048/2003018426.html.

Hughes:2007:BRA

[Hug07a] Jeff Hughes. Book review: Andrew Brown, J. D. Bernal: The Sage of Science. Oxford: Oxford University Press, 2005. Pp. xiv + 562. ISBN 0-19-851544-8. £25.00 (hardback). British Journal for the History of Science, 40 (1):149–150, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500708.

Hughes:2007:BRC

[Hug07b] Jeff Hughes. Book review:

Cathryn Carson and David A. Hollinger (eds.), Reappraising Oppenheimer: Centennial Studies and Reflections. Berkeley Papers in History of Science, Volume 21. Berkeley: Office for History of Science and Technology, University of California, 2005. Pp. xii + 413. ISBN 0-9672617-3-2. $14.00 (paperback). British Journal for the History of Science, 40(3):459–460, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500763.

Hughes:2008:BRJ

[Hug08] Jeff Hughes. Book review: John Krige, American Hegemony and the Postwar Reconstruction of Science in Europe. Cambridge, MA: MIT Press, 2006. Pp.

384. ISBN 0-262-11297-3. $40.00, £25.95 (hardback). British Journal for the History of Science, 41(4):624–626, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hughes:2009:BRC

[Hug09]  Jeff Hughes.  Book review:

Charles Thorpe, Oppenheimer: The Tragic Intellect. Chicago and London: University of Chicago Press, 2006. Pp. xx + 413. ISBN 0-226-79845-3. $37.50, £24.00 (hardback). British Journal for the History of Science, 42 (1):127–128, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hulin:2000:PHS

[Hul00] Nicole Hulin, editor. Physique et humanités scientifiques: autour de la réforme de l’enseignement de 1902: études et documents. (French) [Physics and scientific humanity: the teaching reform of 1902: studies and documents]. Histoire des sciences.

Presses universitaires de Septentrion, Villeneuve d’Ascq (Nord), France, 2000. ISBN 2-85939-624-1. 339 pp. LCCN QC47.F8 P49 2000.

Hull:2007:BRC

[Hul07] Andrew Hull. Book review:

Christopher Lawrence, Rockefeller Money, the Laboratory and Medicine in Edinburgh 1919– 1930: New Science in an Old Country. Rochester Studies in Medical History. Vol. 5. Rochester, NY and Woodbridge, Suffolk: University of Rochester Press, 2005. Pp. ix + 373. ISBN 1-58046-195-6. £60.00, $85.00 (hardback). British Journal for the History of Science, 40 (2):304–305, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500739.

Hunt:2002:BRK

[Hun02]  Bruce J. Hunt.  Book re-

view: Ken Beauchamp, History of Telegraphy: Its Technology and Application. IEE History of Technology Series,

26. London: Institution of Electrical Engineers, 2001. Pp. xxiii + 413. ISBN 0-85296-792-6. £55.00, $95.00 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hunt:2005:BRJ

[Hun05a] Bruce J. Hunt. Book review:

Jed Z. Buchwald and Andrew Warwick (eds.), Histories of the Electron: The Birth of Microphysics. Dibner Institute Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2001. Pp. xi + 514. ISBN 0-262-02494-2. £37.95 (hardback). British Journal for the History of Science, 38 (1):117–118, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028598.

Hunt:2005:BRP

[Hun05b] Bruce J. Hunt. Book reviews: P. M. Harman (ed.), The Scientific Letters and Papers of James Clerk Maxwell. Volume I: 1846– 1862. Cambridge: Cambridge University Press, 1990. Pp. xxviii + 748. ISBN 0-521-25625-9. $195.00 (hardback). Volume II: 1862–1873. Cambridge: Cambridge University Press, 1995. Pp. xxx + 999. ISBN 0-521-25626-7. $285.00 (hardback). Volume III: 1873–1879. Cambridge: Cambridge University Press, 2002. Pp. xxvii + 932. ISBN 0-521-25627-5. £210.00, $315.00 (hardback). British Journal for the History of Science, 38(4):485– 487, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028469.

Hunt:2006:BRB

[Hun06] Bruce J. Hunt. Book review:

Basil Mahon, The Man Who Changed Everything: The Life of James Clerk Maxwell. Chichester: John Wiley, 2003. Pp. xx + 226. ISBN 0-470-86088-X. £18.99 (hardback). British Journal for the History of Science, 39(2):303, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028627.

Hunt:2007:BRD

[Hun07a] Bruce J. Hunt. Book review:

David Kaiser (ed.), Pedagogy and the Practice of Science: Historical and Contemporary Perspectives. Cambridge, MA and London: MIT Press, 2005. Pp. vi + 426. ISBN 0-262-11288-4. £29.95 (hardback). British Journal for the History of Science, 40 (1):144–145, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500705.

Hunter:2007:RBE

[Hun07b] Michael Hunter. Robert Boyle and the early Royal Society: a reciprocal exchange in the making of Baconian science. British Journal for the History of Science, 40(1):1–23, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500680.

Hurn:2007:BRJ

[Hur07] Mark Hurn. Book review: Janet and Mark Robinson (eds.), The Stargazer of Hardwicke: The Life and Work of Thomas William Webb. Leominster: Gracewing, 2006. Pp. xix + 259. ISBN 0-85244-666-7. £14.99 (hardback). British Journal for the History of Science, 40(4):615–616, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hutchison:2004:BRC

[Hut04] Hazel Hutchison. Book review: Christopher Herbert, Victorian Relativity: Radical Thought and Scientific Discovery. Chicago and London: University of Chicago Press, 2001. Pp. xv + 302. ISBN 0-226-32733-7. £10.50, $16.00 (paperback). British Journal for the History of Science, 37(4): 482–483, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028659.

Hutton:2007:BRN

[Hut07] Sarah Hutton. Book review: Noel Malcolm and Jacqueline Stedall, John Pell (1611–1685) and his Correspondence with Sir Charles Cavendish: The Mental World of an Early Modern Mathematician. Oxford: Oxford University Press, 2005. Pp.

657. ISBN 0-19-856484-8. £90.00 (hardback). British Journal for the History of Science, 40 (2):287–288, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500727.

Hutton:2009:BRH

[Hut09] Sarah Hutton. Book review:

Hugh Trevor-Roper, Europe’s Physician: The Various Life of Sir Theodore de Mayerne. New Haven and London: Yale University Press, 2006. Pp. xii + 438. ISBN 0-300-11263-7. $35.00. £25.00 (hardback). British Journal for the History of Science, 42(3):456–457, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

HCJHWML:2001:FHY

[HW01a] Harvey Cushing and John Hay Whitney Medical Library and Susan Wheeler. Five hundred years of medicine in art: an illustrated catalogue of prints and drawings in the Clements C. Fry Collection in the Harvey Cushing/John Hay Whitney Medical Library at Yale University. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-85967-992-6. xxviii + 363 pp. LCCN N8223 .H355 2001.

Helm:2001:RCS

[HW01b] Jürgen Helm and Annette Winkelmann, editors. Religious confessions and the sciences in the Sixteenth Century, volume 1 of Studies in European Judaism, 1568-5004. Brill, Leiden, The Netherlands, 2001. ISBN 90-04-12045-9. xiv + 161 pp. LCCN BL240.3 .R45 2001.

Igo:2007:AAS

[Igo07] Sarah Elizabeth Igo. The averaged American: surveys, citizens, and the making of a mass public. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02321-8. 398 pp. LCCN HN29 .I44 2007.

Iliffe:2005:EBI

[IH05] Rob Iliffe and Rebekah Higgitt, editors. Early biographies of Isaac Newton: 1660–1885. Pickering, London, UK, 2005. ISBN 1-85196-778-8 (hardcover). xliii + 420 pp. LCCN QC16.N7 E18

                                        2005. URL http://www.loc.
                                        gov/catdir/toc/ecip058/
                                        2005004907.html. Preface by

Milo Keynes.

Iliffe:2000:MBT

[Ili00] Rob Iliffe. The masculine birth of time: temporal frameworks of early modern natural philosophy. British Journal for the History of Science, 33(4): 427–453, December 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028029.

Iqbal:2007:SI

[Iqb07] Muzaffar Iqbal. Science and Islam. Greenwood guides to science and religion. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2007. ISBN 0-313-33576-1. xxxv + 233 pp. LCCN BP190.5.S3 I672

2007. URL http://www.loc.
gov/catdir/toc/ecip078/
2007000423.html.

Ironside:2004:BRR

[Iro04] Georgette Ironside. Book review: Rina Knoeff, Herman Boerhaave (1668–1738): Calvinist Chemist and Physician. History of Science and Scholarship in the Netherlands, 3. Amsterdam: Koninklijke Nederlandse Akademie van Wetenschappen, 2002. Pp. xvi + 237. ISBN 90-6984-342-0. EUR35.00 (hardback). British Journal for the History of Science, 37(4):472– 473, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028650.

Israel:2001:REP

[Isr01] Jonathan I. (Jonathan Irvine)

Israel. Radical enlightenment: philosophy and the making of modernity, 1650–1750. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-820608-9. xvi + 810 pp. LCCN B802 .I87

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0639/00044086-d.html;
http://www.loc.gov/catdir/
enhancements/fy0639/
00044086-t.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
00044086-b.html.

James:2007:CRI

[J^{+}07] Frank A. J. L. James et al., editors. Christmas at the Royal Institution: an anthology of lectures. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-277-108-5, 981-277-109-3 (paperback). xxvii + 366 pp. LCCN Q180.55.D57 C47

                                        2007. URL http://www.loc.
                                        gov/catdir/toc/fy1001/
                                        2008273188.html.

Jackson:2000:SBJ

[Jac00a] Myles W. Jackson. Spectrum of belief: Joseph von Fraunhofer and the craft of precision optics. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-10084-3. x + 284 pp. LCCN QC450.5 .J33 2000.

Jacyna:2000:LWN

[Jac00b] L. S. Jacyna. Lost words: narratives of language and the brain, 1825–1926. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-00413-7 (hardcover). 241 pp. LCCN RC425 .J33 2000. URL http://www.

                                        loc.gov/catdir/description/
                                        prin021/99089724.html;
                                        http://www.loc.gov/catdir/
                                        toc/prin032/99089724.html.

Jackson:2006:HTP

[Jac06] Myles W. Jackson. Harmonious triads: physicists, musicians, and instrument makers in Nineteenth-Century Germany. Transformations. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-10116-5. x + 395 pp. LCCN ML3805 .J33

2006. URL http://www.loc.
gov/catdir/toc/fy0710/
2006044991.html.

Jacob:2009:BRP

[Jac09] Margaret C. Jacob. Book review: Peter M. Jones, Industrial Enlightenment: Science, Technology, and Culture in Birmingham and the West Midlands, 1760–1820. Manchester and New York: Manchester University Press, 2008. Pp. xii + 260. ISBN 978-0-7190-7770-8. £55.00 (hardback). British Journal for the History of Science, 42(3): 462–463, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jahnke:2003:HA

[Jah03] H. N. (Hans Niels) Jahnke, editor. A history of analysis, volume 24 of History of mathematics, 0899-2428. American Mathematical Society, Providence, RI, USA, 2003. ISBN 0-8218-2623-9. ix + 422 pp. LCCN QA300

.H55 2003. URL http://www.e-
streams.com/es0703/es0703_
3102.html.

James:2001:BRA

[Jam01a] Frank A. J. L. James. Book review: Albert E. Moyer, Joseph Henry: The Rise of an American Scientist. Washington and London: Smithsonian Institution Press, 1997. Pp. xii + 348. ISBN 1-56098-776-6. £34.95, $45.00. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028050.

James:2001:BRM

[Jam01b] Frank A. J. L. James. Book reviews: Marc Rothenberg, Paul H. Theerman, Kathleen W. Dorman, John C. Rumm and Deborah Y. Jeffries (eds.), The Papers of Joseph Henry. Volume 7. The Smithsonian Years, January 1847–December 1849. Washington and London: Smithsonian Institution Press, 1996. Pp. xlviii + 707. ISBN 1-56098-533-X. No price given (hardback). Marc Rothenberg, Kathleen W. Dorman, Deborah Y. Jeffries and Frank R. Millikan (eds.), The Papers of Joseph Henry. Volume 8. The Smithsonian Years, January 1850–December 1853. Washington and London: Smithsonian Institution Press, 1998. Pp. xlvii+548. ISBN 1-56098-891-6. No price given (hardback). British Journal for the History of Science, 34(4):453– 481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028385.

James:2002:RME

[Jam02] I. M. (Ioan Mackenzie) James. Remarkable mathematicians: from Euler to von Neumann. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2002.

ISBN 0-521-81777-3 (hardcover), 0-521-52094-0 (paperback). xiv + 433 pp. LCCN QA28 .J36

2002. URL http://www.loc.
gov/catdir/description/
cam022/2002022266.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2002022266-b.html;
http://www.loc.gov/catdir/
samples/cam033/2002022266.
html;
http://www.loc.gov/catdir/
toc/cam026/2002022266.html.

James:2003:BRC

[Jam03] Frank A. J. L. James. Book reviews: Colin A. Russell, Michael Faraday: Physics and Faith. Oxford Portraits in Science. New York: Oxford University Press, 2000. Pp. 124. ISBN 0-19-511763-8. £15.20, $24.00 (hardback). Ann Fullick, Michael Faraday. Groundbreakers. Oxford: Heinemann Library, 2000. Pp. 48. ISBN 0-431-10443-3. £11.25. (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028317.

James:2005:BRM

[Jam05] Frank A. J. L. James. Book review: Marc Rothenberg, Kathleen W. Dorman and Frank R. Millikan (eds.), The Papers of Joseph Henry. Volume 9. The Smithsonian Years, January 1854–December 1857. With the assistance of Deborah Y. Jeffries.

Canton, MA: Smithsonian Institution/Science History Publications, 2002. Pp. l + 516. ISBN 1-88135-363-9. $79.95 (hardback). British Journal for the History of Science, 38(1):115–116, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028596.

James:2008:JFM

[Jam08] Frank A. J. L. James. The Janus face of modernity: Michael Faraday in the twentieth century. British Journal for the History of Science, 41(4):477–516, December 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jankovic:2000:RSC

[Jan00] Vladimir Janković. Reading the skies: a cultural history of English weather, 1650– 1820. Manchester University Press, Manchester, UK, 2000. ISBN 0-7190-5988-7 (cloth), 0-7190-5989-5 (paperback). xiv + 272 pp. LCCN QC857.G77 J36

                                        2000. URL http://www.h-net.
                                        org/review/hrev-a0c2m3-aa;
                                         http://www.loc.gov/catdir/
                                        bios/uchi051/00053250.html;
                                         http://www.loc.gov/catdir/
                                        toc/uchi051/00053250.html.

Jardine:2008:BRJ

[Jar08a] Boris Jardine. Book review: Jutta Schickore, The Microscope and the Eye: A History of Reflections, 1740–1870. Chicago and London: University of Chicago Press, 2007. Pp. 320. ISBN 978-0-226-73784-3. $40.00 (hardback). British Journal for the History of Science, 41(3):460– 462, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jardine:2008:PCF

[Jar08b] Nicholas Jardine. Prince Cesi and fungi, not to mention fungifunguli. British Journal for the History of Science, 41(2):267– 273, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jeater:2007:LLS

[Jea07] Diana Jeater. Law, language, and science: the invention of the “native mind” in Southern Rhodesia, 1890–1930. Social history of Africa series, 1099-8098. Heinemann, Portsmouth, NH, 2007. ISBN 0-325-07108-X (hardcover), 0-325-07109-8 (paperback). xxi + 273 pp. LCCN P81.Z55 J43 2007. URL http:

//www.loc.gov/catdir/toc/
ecip071/2006031039.html.

Jeffrey:2001:MOH

[Jef01]  Kirk Jeffrey.  Machines in

our hearts: the cardiac pacemaker, the implantable defibrillator, and American health care. The Johns Hopkins University Press, Baltimore, MD, USA, 2001. ISBN 0-8018-6579-4. xiii + 370 pp. LCCN RC684.P3 J444

2001. URL http://www.loc.
gov/catdir/bios/jhu052/
00009627.html;

                                        http://www.loc.gov/catdir/
                                        description/jhu051/
                                        00009627.html;
                                        http://www.loc.gov/catdir/
                                        toc/jhu051/00009627.html.

Jenner:2001:BRP

[Jen01] Mark S. R. Jenner. Book reviews: Philip K. Wilson, Surgery, Skin and Syphilis: Daniel Turner’s London (1667– 1741). Wellcome Institute Series in the History of Medicine: Clio Medica 54. Amsterdam and Atlanta, GA: Rodopi, 1999. Pp. xv + 312. ISBN 90-420-0526-2. £48, $88.50 (hardback); ISBN 90-420-0516-5. £13.50, $22.50 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028386.

Jenner:2002:BRP

[Jen02] Mark Jenner. Book review: Peter Millard (ed.), Notes of Me: The Autobiography of Roger North. Toronto, Buffalo and London: University of Toronto Press, 2000. Pp. xix + 353. ISBN 0-8020-4471-9. £42.00, $65.00 (hardback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johns:2000:MMA

[Joh00] Adrian Johns. Miscellaneous methods: authors, societies and journals in early modern England. British Journal for the History of Science, 33(2): 159–186, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027921.

Johnston:2001:HLC

[Joh01] Sean Johnston. A history of light and colour measurement: science in the shadows. Institute of Physics Pub., Bristol, UK, 2001. ISBN 0-7503-0754-4 (hardcover). xi + 281 pp. LCCN QC391 .J64

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0807/2008295506-d.html.

Johnson:2002:SAS

[Joh02a] Stephen B. Johnson. The secret of Apollo: systems management in American and European space programs. New series in NASA history. The Johns Hopkins University Press, Baltimore, MD, USA, 2002. ISBN 0-8018-6898-X (hardcover). xvii + 290 pp. LCCN UG1523 .J645 2002. URL

http://www.loc.gov/catdir/
bios/jhu051/2001005688.
html;   http://www.loc.gov/
catdir/description/jhu051/
2001005688.html;
http://www.loc.gov/catdir/
toc/fy035/2001005688.html.

Johnston:2002:BRP

[Joh02b] Sean Johnston. Book review:

P. Klaushentschel and Axel D. Wittmann (eds.), The Role of Visual Representations in Astronomy: History and Research Practice. Acta Historica Astronomiae, 9. Thun and Frankfurt am Main: Verlag Harri Deutsch, 2000. Pp. 148. ISBN 3-8171-1630-6. DM 24.00 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Johnston:2002:BRR

[Joh02c]  Sean Johnston.  Book re-

view: Russellburns, John Logie Baird, Television Pioneer. History of Technology Series, 28. London: Institution of Electrical Engineers, 2000. Pp. xxv + 417. ISBN 0-85296-797-7. £55.00 (hardback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028194.

Johnston:2002:BRX

[Joh02d] Sean Johnston. Book review: Xiang Chen, Instrumental Traditions and Theories of Light: The Uses of Instruments in the Optical Revolution. Science and Philosophy, 9. Dordrecht, Boston and London: Kluwer Academic Publishers, 2000. Pp. xxiii + 211. ISBN 0-7923-6349-3. £60.00, $99.00 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028412.

Johns:2003:BRR

[Joh03a] Adrian Johns. Book review:

Richard Yeo, Encyclopaedic Visions: Scientific Discoveries and Enlightenment Culture. Cambridge: Cambridge University Press, 2001. Pp. xxi + 336. ISBN 0-521-65191-3. £40.00 (hardback). British Journal for the History of Science, 36(3): 365–366, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028160.

Johnston:2003:BRK

[Joh03b] Sean Johnston. Book review:

Klaus Hentschel, Mapping the Spectrum: Techniques of Visual Representation in Research and Teaching. Oxford: Oxford University Press, 2002. Pp. xiii + 562. ISBN 0-19-850953-7. £75.00 (hardback). British Journal for the History of Science, 36 (1):87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028318.

Johansen:2004:PNP

[Joh04a] T. K. Johansen. Plato’s natural philosophy: a study of the Timaeus-Critas. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-79067-0. vi + 218 pp. LCCN B387 .J64

Johnston:2004:BRA

[Joh04b] Sean Johnston. Book review: Antony Kamm and Malcolm Baird, John Logie Baird: A Life. Edinburgh: National Museums of Scotland Publishing, 2002. Pp. xii + 465. ISBN 1-901663-76-0. £25.00 (hardback). British Journal for the History of Science, 37(2):221–222, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028349.

Johnston:2004:BRM

[Joh04c] Sean Johnston. Book review: Maggie Mort, Building the Trident Network: A Study of the Enrollment of People, Knowledge, and Machines. Inside Technology. Cambridge, MA and London: MIT Press, 2002. Pp. x + 217. ISBN 0-262-13397-0. £22.50 (hardback). British Journal for the History of Science, 37(4):485–486, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028661.

Johnson:2005:AT

[Joh05a] Monte Ransome Johnson. Aristotle on teleology. Oxford Aristotle studies. Clarendon Press, New York, NY, USA, 2005. ISBN 0-19-928530-6. xi + 339 pp. LCCN B491.T4 J64

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0637/2005020165-d.html;

http://www.loc.gov/catdir/
enhancements/fy0724/
2005020165-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0516/2005020165.html.

Johnston:2005:BRJ

[Joh05b] Sean Johnston. Book review:

John Cantrell and Gillian Cookson (eds.), Henry Maudslay and the Pioneers of the Machine Age. Stroud and Charleston: Tempus, 2002. Pp. 192. ISBN 0-7524-2766-0. £16.99, $26.99 (paperback). British Journal for the History of Science, 38(4): 483–484, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028467.

Johnston:2005:BRS

[Joh05c]  Sean Johnston.  Book re-

view: Steven J. Dick, Sky and Ocean Joined: The U.S. Naval Observatory 1830–2000. Cambridge: Cambridge University Press, 2003. Pp. xiii + 609. ISBN 0-521-81599-1. £90.00, $130.00 (hardback). British Journal for the History of Science, 38 (2):236–237, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028711.

Johnston:2006:BRN

[Joh06a] Sean Johnston. Book review:

Nicholas J. Wade, Destined for Distinguished Oblivion: The Scientific Vision of William Charles Wells (1757–1817). History and Philosophy of Psychology. New York, Boston, Dordrecht, London and Moscow: Kluwer Academic Publishers, 2003. Pp. xi + 310. ISBN 0-306-47385-2. $95.00 (hardback). British Journal for the History of Science, 39(2):292, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028618.

Johnston:2006:HVH

[Joh06b] Sean Johnston. Holographic visions: a history of new science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-857122-4. xxi + 518 pp. LCCN QC449 .J66

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0637/2005036594-d.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2005036594-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip065/2005036594.html.

Johnston:2006:BRT

[Joh06c] Sean F. Johnston. Book review: Thomas P. Hughes, Human-Built World: How to Think about Technology and Culture. Chicago and London: University of Chicago Press, 2004. Pp. xii + 223. ISBN 0-226-35933-6. £16.00, $22.50 (hardback). British Journal for the History of Science, 39(3):441– 442, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028489.

Johnson:2007:BRM

[Joh07a] Kristin Johnson. Book review:

Marco Beretta (ed.), From Private to Public: Natural Collections and Museums. Uppsala Studies in History of Science. Vol. 31. European Studies in Science History and the Arts. Vol.

5. Sagamore Beach, MA: Science History Publications, 2005. Pp. ix + 252. ISBN 0-88-135360-4. $39.95 (hardback). British Journal for the History of Science, 40(2):279–280, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500722.

Johnson:2007:BRR

[Joh07b] Kristin Johnson. Book review:

Robert E. Kohler, All Creatures: Naturalists, Collectors, and Biodiversity, 1850–1950. Princeton and Oxford: Princeton University Press, 2006. Pp. xiii + 363. ISBN 978-0-691-12539-2. $35.00, £22.95 (hardback). British Journal for the History of Science, 40(3):452–453, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500759.

Johnson:2009:BRD

[Joh09]  Kristin Johnson.  Book re-

view: D. Graham Burnett, Trying Leviathan: The Nineteenth-Century New York Court Case that Put the Whale on Trial and Challenged the Order of Nature. Princeton: Princeton University Press, 2007. Pp. xiv + 266. ISBN 978-0-691-12950-1. £17.95 (hardback). British Journal for the History of Science, 42(1):142– 143, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jones:2002:BRM

[Jon02a] Claire Jones. Book review: Margaret A. M. Murray, Women Becoming Mathematicians: Creating a Professional Identity in Post-World War II America. Cambridge, MA and London: MIT Press, 2000. Pp. vxiii + 277. ISBN 0-262-13369-5. £10.50 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028195.

Jones:2002:ARW

[Jon02b] Greta Jones. Alfred Russel Wallace, Robert Owen and the theory of natural selection. British Journal for the History of Science, 35(1):73–96, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028403.

Jones:2004:BRI

[Jon04] Claire Jones. Book review: Ioan James, Remarkable Mathematicians: From Euler to von Neumann. Spectrum Series of the Mathematical Association of America. Cambridge: Cambridge University Press, 2002. Pp. xiv + 433. ISBN 0-521-52094-0. £19.95, $25.00 (paperback). British Journal for the History of Science, 37(4):487– 488, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028663.

Jones:2006:BRC

[Jon06a] Greta Jones. Book review:

Clare Pettitt, Patent Inventions: Intellectual Property and the Victorian Novel. Oxford: Oxford University Press, 2004. Pp. xiv + 341. ISBN 0-19-925320-X. £58.00 (hardback). British Journal for the History of Science, 39(3):437–438, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028486.

Jones:2006:GLS

[Jon06b] Matthew L. (Matthew Laurence)

Jones. The good life in the scientific revolution: Descartes, Pascal, Leibniz, and the cultivation of virtue. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-40954-6 (cloth), 0-226-40955-4 (paperback), (cloth,, ). xvi + 384 pp. LCCN Q125.2 .J66

2006. URL http://www.h-net.
org/review/hrev-a0f4r7-aa;

                                         http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2006040491-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2006040491-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0707/
                                        2006040491-t.html.

Jones:2008:IES

[Jon08] Peter Jones. Industrial enlightenment: science, technology and culture in Birmingham and the West Midlands, 1760– 1820. Manchester University Press, Manchester, UK, 2008. ISBN 0-7190-7770-2. xii + 260 pp. LCCN HC258.B6 J66

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0913/2009284137-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0913/
                                        2009284137-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0913/
                                        2009284137-t.html.

Jones:2009:BRR

[Jon09] Claire Jones. Book review: Roy Church and E. M. Tansey, Burroughs Wellcome & Co.: Knowledge, Trust and Profit, and the Transformation of the British Pharmaceutical Industry, 1880– 1940. Lancaster: Crucible, 2007. Pp. xxvii + 564. ISBN 978-1-905-472-07-9. £19.99 (paperback). British Journal for the History of Science, 42(2):288– 289, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jordanova:2000:PAR

[Jor00] Ludmilla Jordanova. Presidential address: remembrance of science past. British Journal for the History of Science, 33(4): 387–406, December 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028027.

Jordanova:2001:BRW

[Jor01] Ludmilla Jordanova. Book review: William R. Shea (ed.), Science and the Visual Image in the Enlightenment. European Studies in Science History and the Arts, IV. Canton: Science History Publications, 2000. Pp. viii + 232. ISBN 0-88135-285-3. $39.95 (hardback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028387.

Jordanova:2002:BRJ

[Jor02] Ludmilla Jordanova. Book reviews: And? J. V. Field and Frank James (eds.), Science in Art: Works in the National Gallery that Illustrate the History of Science and Technology. BSHS Monographs, 11. Stanford in the Vale: British Society for the History of Science, 1997. Pp. 110. ISBN 0-906450-13-6. £15.00, $26.00 (paperback). James Hamilton (ed.), Fields of Influence: Conjunctions of Artists and Scientists 1815– 1860. Birmingham: University of Birmingham Press, 2001. Pp. xiii + 174. ISBN 0-902459-10-5. £20.00, $35.00 (hardback). David Bindman, Frèdéric Ogée and Peter Wagner (eds.), Hogarth: Representing Nature’s Machines. Barber Institute’s Critical Perspectives in Art History. Manchester and New York: Manchester University Press, 2001. Pp. xvi+287. ISBN 0-7190-5919-4. £18.99 (paperback). British Journal for the History of Science, 35(3): 341–345, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028128.

Jordanova:2006:BRL

[Jor06] Ludmilla Jordanova. Book review: Lorraine Daston (ed.), Things that Talk: Object Lessons from Art and Science. New York: Zone Books/MIT Press, 2004. Pp. 447. ISBN 1-890951-43-9. $37.50 (hardback). British Journal for the History of Science, 39(3):436–437, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028485.

Jordanova:2008:BRA

[Jor08] Ludmilla Jordanova. Book review: Ann B. Shteir and Bernard Lightman (eds.) Figuring It Out: Science, Gender, and Visual Culture. Lebanon, NH: Dartmouth College Press, 2006. Pp. xxx + 386. ISBN 1-58465-602-6. $65.00 (hardback). British Journal for the History of Science, 41(3): 467–469, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Jacob:2004:PMN

[JS04] Margaret C. Jacob and Larry Stewart. Practical matter: Newton’s science in the service of industry and empire, 1687–1851. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01497-9. 201 pp. LCCN Q174 .8 .J33

2004. URL http://www.h-net.
org/review/hrev-a0e6q9-aa;
http://www.loc.gov/catdir/
toc/fy051/2004052293.html.

Kaiser:2005:BRJ

[Kai05a] David Kaiser. Book review:

Jagdish Mehra and Kimball A. Milton, Climbing the Mountain: The Scientific Biography of Julian Schwinger. Oxford: Oxford University Press, 2000. Pp. xii + 677. ISBN 0-19-850658-9. £45.00 (hardback). British Journal for the History of Science, 38(2):239–240, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028714.

Kaiser:2005:DTA

[Kai05b] David Kaiser. Drawing theories apart: the dispersion of Feynman diagrams in postwar physics. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4 K35 2005. URL

http://www.loc.gov/catdir/
bios/uchi051/2004023335.
html;   http://www.loc.gov/
catdir/enhancements/fy0617/
2004023335-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip051/2004023335.html.

Kaiser:2005:PPS

[Kai05c] David Kaiser, editor. Pedagogy and the practice of science: historical and contemporary perspectives. Inside technology. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-11288-4. vi + 426 pp. LCCN Q181 .P344 2005.

Kassler:2001:MSP

[Kas01] Jamie Croy Kassler. Music, science, philosophy: models in the universe of thought, volume CS713 of Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-86078-862-8. xvi + 301 pp. LCCN ML3800 .K187 2001.

Kassell:2002:BRA

[Kas02] Lauren Kassell. Book review: Andrew Wear, Knowledge and Practice in English Medicine, 1550–1680. Cambridge: Cambridge University Press, 2000. Pp. viii + 496. ISBN 0-521-55827-1. £16.95, $27.95 (paperback). British Journal for the History of Science, 35(3):347– 379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kamm:2002:JLB

[KB02] Antony Kamm and Malcolm H. I. Baird. John Logie Baird: a life. National Museum of Scotland Publ., Edinburgh, Scotland, 2002. ISBN 1-901663-76-0. xi + 465 + 33 pp. LCCN TK6635.B3 K36 2002.

Kuhn:2000:RSS

[KCH00] Thomas S. Kuhn, James Conant, and John Haugeland. The road since structure: philosophical essays, 1970–1993, with an autobiographical interview. University of Chicago Press, Chicago, IL, USA, 2000. ISBN 0-226-45798-2 (cloth). viii + 335 pp. LCCN Q175 .K94 2000. URL http:/

/www.loc.gov/catdir/bios/
uchi051/00131583.html;
http://www.loc.gov/catdir/
description/uchi051/
00131583.html;
http://www.loc.gov/catdir/
toc/uchi051/00131583.html.

Kepler:2000:OPW

[KD00] Johannes Kepler and William H.

Donahue. Optics: Paralipomena to Witelo and optical part of astronomy. Green Lion Press, Santa Fe, NM, USA, 2000. ISBN 1-888009-12-8. xv + 459 pp. LCCN QC380 .K4613 2000.

Knight:2005:SBN

[KE05] David M. Knight and Matthew Eddy, editors. Science and beliefs: from natural philosophy to natural science, 1700–1900. Science, technology, and culture, 1700–1945. Ashgate, 2005. ISBN 0-7546-3996-7. xiv + 272 pp. LCCN Q125 .S43177

2005. URL http://www.loc.
gov/catdir/toc/ecip0417/
2004008415.html.

Keller:2000:CG

[Kel00] Evelyn Fox Keller. The century of the gene. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00372-1 (hardcover). 186 pp. LCCN QH428 .K448 2000.

Kemp:2009:BRJ

[Kem09] Martin Kemp. Book review: Jean A. Givens, Karen M. Reeds and Alain Touwaide (eds.), Visualizing Medieval Medicine and Natural History, 1200–1550. AVISTA Studies in the History of Medieval Technology, Science and Art. Aldershot: Ashgate, 2006. Pp. xx + 278. ISBN 0-7546-5296-3. £55.00 (hardback). British Journal for the History of Science, 42(4): 602–603, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kenny:2007:CHC

[Ken07] Robert Kenny. From the Curse of Ham to the curse of nature: the influence of natural selection on the debate on human unity before the publication of the descent of man. British Journal for the History of Science, 40(3):367–388, September 2007.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500748.

Kennedy:2008:STP

[Ken08] Alasdair Kennedy. In search of the ‘true prospect’: making and knowing the Giant’s Causeway as a field site in the seventeenth century. British Journal for the History of Science, 41(1):19–41, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Keynes:2000:CDZ

[Key00] R. D. Keynes, editor. Charles Darwin’s zoology notes and specimen lists from H.M.S. Beagle. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-46569-9 (hardcover), 0-521-67350-X (paperback). xxxiv + 430 pp. LCCN QH365.Z9 K4

2000. URL http://www.loc.
gov/catdir/description/
cam022/2001269276.html;
http://www.loc.gov/catdir/
toc/cam026/2001269276.html.

Keynes:2001:ABC

[Key01] Randal Keynes. Annie’s box:

Charles Darwin, his daughter, and human evolution. Fourth Estate, London, UK, 2001. ISBN 1-84115-060-6. xiv + 331 pp. LCCN QH31.D2 K48 2001. URL

http://www.loc.gov/catdir/
toc/fy053/2004484653.html.

Keynes:2002:FFF

[Key02] R. D. Keynes. Fossils, finches and fuegians: Charles Darwin’s adventures and discoveries on the Beagle, 1832–1836. Harper-Collins, London, UK, 2002. ISBN 0-00-710189-9. xix + 428 pp. LCCN QH31.D2 K478 2002.

Keynes:2005:ISI

[Key05] Milo Keynes. The iconography of Sir Isaac Newton to 1800. Boydell Press, Woodbridge, Suffolk, UK, 2005. ISBN 1-84383-133-3 (hardcover). viii + 120 + xxiv pp. LCCN N7628.N48 A4

2005. URL http://www.loc.
gov/catdir/toc/ecip051/
2004023329.html.

Kassler:2004:BMP

[KG04] Jamie Croy Kassler and Baron Francis North Guilford. The beginnings of the modern philosophy of music in England: Francis North’s A philosophical essay of musick (1677) with comments of Isaac Newton, Roger North and in the Philosophical Transactions. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-7546-0139-0. xiv + 243 pp. LCCN ML3800 .K186 2004.

Khalid:2006:BRA

[Kha06a] Amna Khalid. Book review: Alison Bashford, Imperial Hygiene: A Critical History of Colonialism, Nationalism and Public Health. Basingstoke: Palgrave Macmillan, 2004. Pp. ix + 264. ISBN 1-4039-0488-X. £50.00 (hardback). British Journal for the History of Science, 39(2):306–307, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028630.

Khalid:2006:BRP

[Kha06b] Amna Khalid. Book reviews: Peter Elmer (ed.), The Healing Arts: Health, Disease and Society in Europe, 1500–1800. Manchester and New York: Manchester University Press in association with The Open University, 2004. Pp. xxix + 408. ISBN 0-7190-6734-0. No price given (paperback). PETERelmer and OLEpetergrell (eds.), Health, Disease and Society in Europe, 1500–1800: A Source Book. Manchester and New York: Manchester University Press in association with The Open University, 2004. Pp. xx+380. ISBN 0-7190-6737-5. £16.99 (paperback). British Journal for the History of Science, 39(4):605–606, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028528.

Kimler:2005:BRC

[Kim05] William C. Kimler. Book review: Carl von Linné, Nemesis Divina. Edited and translated with explanatory notes by M. J. Petry. Archives internationales d’histoire des idées,

177. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. xviii + 483. ISBN 0-7923-6820-7. £119.00, $169.00 (hardback). British Journal for the History of Science, 38 (2):227–228, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028704.

Kimmelman:2007:BRJ

[Kim07] Barbara Kimmelman. Book review: Jonathan Harwood, Technology’s Dilemma: Agricultural Colleges between Science and Practice in Germany, 1860–1934. Bern: Peter Lang AG, 2005. Pp.

288. ISBN 3-03910-299-0. £29.00 (paperback.). British Journal for the History of Science, 40 (1):146–147, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500706.

King:2001:CMF

[Kin01]  David A. King.  The ci-

phers of the monks: a forgotten number-notation of the Middle Ages, volume 44 of Boethius: Texte und Abhandlungen zur Geschichte der Mathematik der Naturwissenschaften. F. Steiner, Stuttgart, Germany, 2001. ISBN 3-515-07640-9. 506 pp. LCCN QA141.2 .K56 2001.

Kingsland:2005:EAE

[Kin05]  Sharon E. Kingsland.  The

evolution of American ecology, 1890–2000. The Johns Hopkins University Press, Baltimore, MD, USA, 2005. ISBN 0-8018-8171-4 (hardcover). x + 313 pp. LCCN QH540.83.U6 K56

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0617/2004026034-b.html;
http://www.loc.gov/catdir/
enhancements/fy0617/
2004026034-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip053/2004026034.html.

Kiple:2003:CHD

[Kip03] Kenneth F. Kiple, editor. The Cambridge historical dictionary of disease. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-80834-0 (hardcover), 0-521-53026-1 (paperback). xiii + 412 pp. LCCN RC41 .C365

2003.  URL http://www.e-
streams.com/es0704/es0704_
3180.html;   http://www.
loc.gov/catdir/description/
cam031/2002031368.html;
http://www.loc.gov/catdir/
samples/cam033/2002031368.
html.

Kiple:2004:BRM

[Kip04] Kenneth F. Kiple. Book review: Michael Worboys, Spreading Germs: Disease Theories and Medical Practice in Britain, 1865–1900. Cambridge History of Medicine. Cambridge: Cambridge University Press, 2000. Pp. xvi + 327. ISBN 0-521-77302-4. £37.50, $59.95 (hardback). British Journal for the History of Science, 37(4):480– 481, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028657.

Kirk:2005:BRA

[Kir05] Robert G. W. Kirk. Book review: Angela N. H. Creager and William Chester Jordan (eds.), The Animal/Human Boundary: Historical Perspectives. Studies in Comparative History, 2. Woodbridge: University of Rochester Press, 2002. Pp. xviii + 342. ISBN 1-58046-120-4. £50.00, $75.00 (hardback). British Journal for the History of Science, 38(3):353– 355, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028676.

Kenyon-Jones:2001:KBA

[KJ01] Christine Kenyon-Jones. Kindred brutes: animals in Romantic period writing. The nineteenth century. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-7546-0332-6. 229 pp. LCCN PR468.A56 K46 2001.

Klein:2001:TMR

[Kle01] Ursula Klein, editor. Tools and modes of representation in the laboratory sciences, volume 222 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0100-2, 94-015-9737-5. ISSN 0068-0346. xv + 259 pp. LCCN Q174 .B67 vol. 222;

                                        QD7. URL http://www.loc.
                                        gov/catdir/enhancements/

fy0814/2001050224-d.html;
http://www.loc.gov/catdir/
enhancements/fy0814/
2001050224-t.html;   https:
//link.springer.com/book/
10.1007/978-94-015-9737-1.

Klein:2004:BRA

[Kle04] Ursula Klein. Book review:

Agustí Nieto-Galan, Colouring Textiles: A History of Natural Dyestuffs in Industrial Europe. Boston Studies in the Philosophy of Science, 217. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. xxv + 246. ISBN 0-7923-7022-8. £59.00, $84.00, EUR97.00 (hardback). British Journal for the History of Science, 37 (2):214–215, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028343.

Klein:2005:BRR

[Kle05] Ursula Klein. Book review:

Robert Siegfried, From Elements to Atoms: A History of Chemical Composition. Transactions of the American Philosophical Society, 92 (4). Philadelphia: American Philosophical Society, 2002. Pp. x + 278. ISBN 0-87169-924-9. No price given (paperback). British Journal for the History of Science, 38(2):221–222, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028699.

Klein:2009:BRH

[Kle09] Ursula Klein. Book review: Hasok Chang and Catherine Jackson (eds.), An Element of Controversy: The Life of Chlorine in Science, Medicine, Technology and War. BSHS Monograph Series, Vol. 13. London: British Society for the History of Science, 2007. Pp. ix + 407. ISBN 978-0-906450-01-7. £15.00 (paperback). British Journal for the History of Science, 42(1):128– 130, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Klein:2003:EMP

[KLG03] Ursula Klein, Timothy Lenoir, and Hans Ulrich Gumbrecht. Experiments, models, paper tools: cultures of organic chemistry in the Nineteenth Century. Writing science. Stanford University Press, Stanford, CA, USA, 2003. ISBN 0-8047-4359-2 (cloth). xi + 305 pp. LCCN QD291 .K57 2003. URL http://www.

                                        e-streams.com/es0611/
                                        es0611_2833.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0710/
                                        2002003144-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0710/
                                        2002003144-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy035/2002003144.html.

Kusukawa:2006:TKW

[KM06] Sachiko Kusukawa and Ian Maclean, editors. Transmitting knowledge: words, images, and instruments in early modern Europe. Oxford-Warburg studies. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-928878-X. xv + 274 pp. LCCN Q127.E8 T73

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0637/2005027342-d.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2005027342-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0519/2005027342.html.

Kozak:2005:ILE

[KMO05] Jan Kozák, Victor S. Moreira, and D. R. (David Roger) Oldroyd. Iconography of the 1755 Lisbon earthquake. Geophysical Institute of the Academy of Sciences of the Czech Republic, Praha, Czech Republic, 2005. ISBN 80-200-1322-9 (Academia: hardback), 80-239-4390-1 (GI: hardback). 81 pp. LCCN DP762 .K69 2005.

Knell:2000:CEG

[Kne00]  Simon J. Knell.  The cul-

ture of English geology, 1815– 1851: a science revealed through its collecting. Ashgate, Aldershot, Hampshire, England, 2000. ISBN 1-84014-625-7 (hardcover). xxi + 377 pp. LCCN QE262.E3 K64 2000.

Knight:2002:BRT

[Kni02] David Knight. Book review:

Trevor H. Levere, Transforming Matter: A History of Chemistry from Alchemy to the Buckyball.

Introductory Studies in the History of Science. Baltimore and London: Johns Hopkins University Press, 2001. Pp. x + 215. ISBN 0-8018-6610-3. £12.50 (paperback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028196.

Knight:2003:BRM

[Kni03] David Knight. Book review: Marco Beretta, Imaging a Career in Science: The Iconography of Antoine Laurent Lavoisier. Bologna Studies in Scientific Heritage, 1. Uppsala Studies in History of Science, 29. Canton, MA: Science History Publications, 2001. Pp. xvii + 126. ISBN 0-88135-294-2. $29.95 (hardback). British Journal for the History of Science, 36 (1):87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028319.

Knight:2004:BRM

[Kni04] David Knight. Book review: Michel Chaouli, The Laboratory of Poetry: Chemistry and Poetics in the Work of Friedrich Schlegel. Parallax. Baltimore and London: Johns Hopkins University Press, 2002. Pp. xiv + 290. ISBN 0-8018-6884-X. £33.50, $45.00 (hardback). British Journal for the History of Science, 37(2):211–212, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028341.

Knight:2005:BRR

[Kni05] David Knight. Book review:

Robert J. Richards, The Romantic Conception of Life: Science and Philosophy in the Age of Goethe. Chicago and London: University of Chicago Press, 2002. Pp. xix + 587. ISBN 0-226-71210-9. £24.00, $35.00 (hardback). British Journal for the History of Science, 38 (2):230–231, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028706.

Knight:2006:BRH

[Kni06a] David Knight. Book review: Hasok Chang, Inventing Temperature: Measurement and Scientific Progress. Oxford Studies in Philosophy of Science. Oxford: Oxford University Press, 2004. Pp. xviii + 286. ISBN 0-19-517127-6. £32.50 (hardback). British Journal for the History of Science, 39(4):599–600, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028524.

Knight:2006:PUS

[Kni06b] David M. Knight. Public understanding of science: a history of communicating scientific ideas, volume 26 of Routledge studies in the history of science, technology, and medicine. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2006. ISBN 0-415-20638-3 (ebook), 0-203-96642-2 (e-book). 232 pp. LCCN Q223 .K55 2006.

URL  http://www.loc.gov/
catdir/enhancements/fy0668/
2006004556-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip068/2006004556.html.

Knight:2008:BRW

[Kni08]  David Knight.  Book re-

view: William Clark, Academic Charisma and the Origins of the Research University. Chicago and London: The University of Chicago Press, 2006. Pp.

662. ISBN 0-226-1091-6. $45.00, £28.50 (hardback). British Journal for the History of Science, 41 (1):145–147, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Knox:2001:BRB

[Kno01] Kevin C. Knox. Book review: Brian Dolan (ed.), Malthus, Medicine, & Morality: ‘Malthusianism’ after 1798. Wellcome Institute Series in the History of Medicine: Clio Medica, 59. Amsterdam and Atlanta, GA: Rodopi, 2000. Pp. v + 232. ISBN 90-420-0851-2. £42.00 (paperback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028113.

Knoeff:2002:HBC

[Kno02] Rina Knoeff. Herman Boerhaave (1668–1738): Calvinist chemist and physician, volume 3 of History of science and scholarship in the Netherlands. Koninklijke Nederlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2002. ISBN 90-6984-342-0 (hardcover). xvi + 237 pp. LCCN Q143.B53 K66

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0628/2005419548-d.html.

Knoeff:2008:BRH

[Kno08] Rina Knoeff. Book review:

Harold J. Cook, Matters of Exchange: Commerce, Medicine, and Science in the Dutch Golden Age. New Haven and London: Yale University Press, 2007. Pp. xiv + 562. ISBN 978-0-300-11796-7. £28.00 (hardback). British Journal for the History of Science, 41(2):278–279, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Karch:2003:HCM

[KNRW03] Steven B. Karch, Josef Nevinný, Emma Reens, and Theodor Walger, editors. A history of cocaine: the mystery of coca java and the Kew plant. Royal Society of Medicine, London, UK, 2003. ISBN 1-85315-547-0 (paperback). x + 224 pp. LCCN HD9019.C63 H57 2003.

Kragh:2002:BRM

[Kra02]  Helge Kragh.  Book re-

view: Michael Heidelberger and Friedrich Steinle (eds.), Experimental Essays – Versuch zum Experiment. Interdisciplinary Studies, 3. Baden-Baden: Nomos Verlagsgesellschaft, 1998. Pp. xiii + 408. ISBN 3-7890-5367-8. DM 88.00 (paperback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kragh:2007:BRM

[Kra07] Helge Kragh. Book review: Matthias Dörries (ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Berkeley Papers in History of Science Vol. 20. Berkeley, CA: Office for History of Science and Technology, 2005. Pp. viii + 195. ISBN 0-9672617-2-4. $12.00 (paperback). British Journal for the History of Science, 40(1):115–116, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500686.

Kragh:2008:BRD

[Kra08a] Helge Kragh. Book review: David Kaiser, Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. Chicago and London: The University of Chicago Press, 2005. Pp. xx + 469. ISBN 0-226-42267-4. £21.00, $30.00 (paperback). British Journal for the History of Science, 41(1):151– 152, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kragh:2008:ECR

[Kra08b] Helge Kragh. Entropic creation:

religious contexts of thermodynamics and cosmology. Science, technology and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2008. ISBN 0-7546-6414-7. 272 pp. LCCN QB981 .K732 2008. URL http://www.

loc.gov/catdir/toc/
ecip0727/2007038440.html.

Kremenetisov:2002:CSC

[Kre02] N. L. Kremenëtìsov. The cure:

a story of cancer and politics from the annals of the Cold War. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-45284-0. xvi + 261 pp. LCCN RC271.I45 K74

2002.  URL http://www.e-
streams.com/es0607/es0607_
2584.html;   http://www.
loc.gov/catdir/bios/
uchi051/2001003291.html;
http://www.loc.gov/catdir/
description/uchi051/
2001003291.html;
http://www.loc.gov/catdir/
toc/fy022/2001003291.html.

Kretschmann:2003:WKW

[Kre03] Carsten Kretschmann, editor. Wissenspopularisierung: Konzepte der Wissensverbreitung im Wandel. (German) [Knowledge popularization: Concepts and changes in knowledge spreading (???)], volume 4 of Wissenskultur und gesellschaftlicher Wandel. Akademie-Verlag, Berlin, Germany, 2003. ISBN 3-05-003770-9 (hardcover). 408 pp. LCCN Q225 .W575 2003.

Kremer:2006:BRS

[Kre06]  Richard Kremer.  Book re-

view: Sean F. Johnston, A History of Light and Colour Measurement: Science in the Shadows. Bristol and Philadelphia: Institute of Physics, 2001. Pp. xi + 281. ISBN 0-7503-0754-4. £45.00, $75.00 (hardback). British Journal for the History of Science, 39(3):442– 444, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028490.

Krige:2006:AHP

[Kri06] John Krige. American hegemony and the postwar reconstruction of science in Europe. Transformations. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-11297-3. viii + 376 pp. LCCN Q127.E8 K75

                                        2006. URL http://www.h-net.
                                        org/review/hrev-a0f3p6-aa.

Kuklick:2006:HCS

[Kuk06] Henrika Kuklick. ‘humanity in the chrysalis stage’: indigenous Australians in the anthropological imagination, 1899–1926. British Journal for the History of Science, 39(4):535–568, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kuklick:2008:NHA

[Kuk08] Henrika Kuklick, editor. A new history of anthropology. Blackwell Publishers, Oxford, UK, 2008. ISBN 0-631-22599-4 (hardcover), 0-631-22600-1 (paperback). xiii + 402 pp. LCCN GN17 .N49

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0802/2006034807-b.html;
http://www.loc.gov/catdir/
enhancements/fy0802/
2006034807-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip073/2006034807.html.

Kumar:2007:PSI

[Kum07] Prakash Kumar. Plantation science: improving natural indigo in colonial India, 1860–1913. British Journal for the History of Science, 40(4):537–565, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Kusukawa:2000:BRC

[Kus00] Sachiko Kusukawa. Book review: Charlotte Methuen, Kepler’s Tübingen: Stimulus to a Theological Mathematics. Aldershot: Ashgate, 1998. Pp. 288+xi. ISBN 1-85928-397-7. £45.00. British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027934.

Kusukawa:2003:BRJ

[Kus03] Sachiko Kusukawa. Book review: Jürgen Helm and Annette Winkelmann (eds.), Religious Confessions and the Sciences in the Sixteenth Century. Studies in European Judaism, 1. Leiden: Brill, 2001. Pp. xiv + 161. ISBN 90-04-12045-9. $58.00, EUR49.00 (hardback). British Journal for the History of Science, 36(3):363–365, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028159.

Kusukawa:2007:BRJ

[Kus07] Sachiko Kusukawa. Book review: John Brooke and Ian Maclean (eds.), Heterodoxy in Early Modern Science and Religion. Oxford: Oxford University Press, 2005. Pp. xxi + 373. ISBN 0-19-926897-5. $120.00, £65.00 (hardback). British Journal for the History of Science, 40(4): 603–605, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lambert:2002:BRE

[Lam02] Cornelia Lambert. Book review: Edmund Russell, War and Nature: Fighting Humans and Insects with Chemicals from World War I to Silent Spring. Studies in Environment and History. Cambridge: Cambridge University Press, 2001. Pp. xvii + 315. ISBN 0-521-79937-6. £12.95, $19.95 (paperback). British Journal for the History of Science, 35(3):347– 379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Larsson:2001:MMO

[Lar01] Ulf Larsson, editor. Människor, miljöer och kreativitet. Cultures of creativity: the centennial exhibition of the Nobel prize, volume 2 of Nobel Museum Archives. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2001. ISBN 0-88135-288-8. 228 pp. LCCN AS911.N9 M34 2001.

Laubichler:2006:BRS

[Lau06] Manfred D. Laubichler. Book review: Soraya de Chadarevian and Nick Hopwood (eds.), Models: The Third Dimension of Science. Stanford: Stanford University Press, 2004. Pp. xvi + 464. ISBN 0-8047-3972-2. £17.50, $24.95 (paperback). British Journal for the History of Science, 39(4):596– 597, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028522.

Lawrence:2005:RML

[Law05] Christopher Lawrence. Rockefeller money, the laboratory, and medicine in Edinburgh, 1919– 1930: new science in an old country. Rochester studies in medical history. University of Rochester Press, Rochester, NY, USA, 2005. ISBN 1-58046-195-6 (hardcover). ISSN 1526-2715. ix + 373 pp. LCCN R497.E34

                                        L39 2005. URL http://www.
                                        loc.gov/catdir/toc/
                                        ecip0511/2005010469.html.

Lundgren:2000:CCT

[LBV00]  Anders  Lund-

gren and Bernadette Bensaude-Vincent, editors. Communicating chemistry: textbooks and their audiences, 1789–1939, volume 3 of European studies in science history and the arts. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2000. ISBN 0-88135-274-8. vii + 465 pp. LCCN QD8.5 .C66 2000. URL http://

                                        www.loc.gov/catdir/toc/
                                        fy032/99052211.html.

Labinger:2001:OCC

[LC01] Jay A. Labinger and H. M. (Harry M.) Collins, editors. The one culture?: a conversation about science. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-46722-8 (cloth), 0-226-46723-6 (paper). xi + 329 pp. LCCN Q175.55 .O54 2001. URL http://www.

loc.gov/catdir/bios/
uchi051/00013097.html;
http://www.loc.gov/catdir/
description/uchi051/
00013097.html;
http://www.loc.gov/catdir/
toc/uchi051/00013097.html.

Leaney:2004:BRC

[Lea04] Enda Leaney. Book review:

Charles Mollan, William Davis and Brendan Finucane (eds.), Irish Innovators in Science and Technology. Dublin: Royal Irish Academy, 2002. Pp. xv + 256. ISBN 1-874045-88-7. £28.00 (hardback). British Journal for the History of Science, 37(1): 100–101, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028260.

Leaney:2005:BRM

[Lea05] Enda Leaney. Book review:

Mark McCartney and Andrew Whitaker (eds.), Physicists of Ireland: Passion and Precision. Bristol and Philadelphia: Institute of Physics Publishing, 2003. Pp. xiv + 298. ISBN 0-7503-0866-4. £40.00, $58.00 (hardback). British Journal for the History of Science, 38 (2):238–239, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028713.

Lederer:2008:BRC

[Led08] Susan E. Lederer. Book review: Carsten Timmerman and Julie Anderson (eds.) Devices and Designs: Medical Technologies in Historical Perspective. Basingstoke and New York: Palgrave Macmillan, 2006. Pp. xiv + 284. ISBN 1-4039-86444-4. £55.00 (hardback). British Journal for the History of Science, 41 (1):122–123, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lee:2002:BEH

[Lee02] Henry C. Lee. Blood evidence: how DNA is revolutionizing the way we solve crimes. Perseus Pub., Cambridge, MA, USA, 2002. ISBN 0-7382-0602-4. xxx + 418 pp. LCCN HV8073 .L32 2002; RA1057.55 .L44

                                        2003. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0832/2002105970-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0832/
                                        2002105970-d.html.

Lehoux:2008:BRK

[Leh08] Daryn Lehoux. Book reviews: T. K. Johansen, Plato’s Natural Philosophy: A Study of the Timaeus-Critias. Cambridge: Cambridge University Press, 2004. Pp. vi + 218. ISBN 0-521-79067-0. £45.00 (hardback). Monte Ransome Johnson, Aristotle on Teleology. Oxford: Oxford University Press, 2005. Pp. xi+339. ISBN 0-19-928530-6. £45.00 (hardback). British Journal for the History of Science, 41 (1):129–130, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Lehoux:2009:BRD

[Leh09] Daryn Lehoux. Book review:

David Sedley, Creationism and Its Critics in Antiquity. Berkeley and London: University of California Press, 2007. Pp. xvii + 269. ISBN 978-0-520-25364-3. £17.95 (hardback). British Journal for the History of Science, 42(3):453–454, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Leshem:2003:NMS

[Les03]  Ayval Leshem.  Newton on

mathematics and spiritual purity, volume 183 of International archives of the history of ideas = Archives internationales d’histoire des idées. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1151-2. x + 230 pp. LCCN QC6 .L4123 2003. URL http://www.

loc.gov/catdir/
enhancements/fy0813/
2003041597-d.html;
http://www.loc.gov/catdir/
toc/fy038/2003041597.html.

Levitt:2000:ECF

[Lev00] Theresa Levitt. Editing out caloric: Fresnel, Arago and the meaning of light. British Journal for the History of Science, 33(1):49–65, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028065.

Levere:2001:TMH

[Lev01] Trevor Harvey Levere. Transforming matter: a history of chemistry from alchemy to the buckyball. Johns Hopkins introductory studies in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2001. ISBN 0-8018-6609-X, 0-8018-6610-3 (paperback). x + 215 pp. LCCN QD11

.L45 2001. URL http://www.e-
streams.com/es0805/es0805_
2024.htm;
http://www.loc.gov/catdir/
bios/jhu051/00011487.html;
 http://www.loc.gov/catdir/
description/jhu051/
00011487.html;
http://www.loc.gov/catdir/
toc/jhu051/00011487.html.

Levine:2002:DKS

[Lev02] George Lewis Levine. Dying to know: scientific epistemology and narrative in Victorian England. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-47536-0. xi + 326 pp. LCCN PR788.S33 L48

2002. URL http://www.loc.
gov/catdir/bios/uchi051/
2001006417.html;
http://www.loc.gov/catdir/
description/uchi051/
2001006417.html;
http://www.loc.gov/catdir/

enhancements/fy0612/
2001006417-t.html.

Levere:2007:BRM

[Lev07] Trevor H. Levere. Book review:

Mary Archer and Christopher Haley (eds.), The 1702 Chair of Chemistry at Cambridge: Transformation and Change. Cambridge: Cambridge University Press, 2005. Pp. xxi + 318. ISBN 0-521-82873-2. £50.00, $90.00 (hardback). British Journal for the History of Science, 40 (2):289–290, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500728.

Lederer:2002:FPS

[LFT02] Susan E. Lederer, Elizabeth Fee, and Patricia Tuohy, editors. Frankenstein: penetrating the secrets of nature: an exhibition by the National Library of Medicine. Rutgers University Press, New Brunswick, NJ, USA, 2002. ISBN 0-8135-3200-0. ix + 78 pp. LCCN PR5397.F73 F74 2002.

Lombroso:2006:CM

[LGR06] Cesare Lombroso, Mary Gibson, and Nicole Hahn Rafter. Criminal man. Duke University Press, Durham, NC, USA, 2006. ISBN 0-8223-3711-8 (cloth), 0-8223-3723-1 (paperback). xviii + 424 pp. LCCN HV6038 .L5813

2006. URL http://www.h-net.
org/review/hrev-a0f5o2-aa;
 http://www.loc.gov/catdir/

                                        toc/ecip064/2005034058.
                                        html.

Lindqvist:2000:MMS

[LHL00] Svante Lindqvist, Marika Hedin, and Ulf Larsson, editors. Museums of modern science: Nobel Symposium 112. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2000. ISBN 0-88135-299-3. LCCN Q105.A1 N63 1999.

Livingstone:1999:ESH

[LHN99] David N. Livingstone, D. G. (Darryl G.) Hart, and Mark A. Noll, editors. Evangelicals and science in historical perspective. Religion in America series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-511557-0 (cloth). vi + 351 pp. LCCN BL245 .E93

                                        1999. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0602/97024739-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0602/
                                        97024739-t.html.

Lightman:2002:HSA

[Lig02]  Bernard Lightman.  Huxley

and scientific agnosticism: the strange history of a failed rhetorical strategy. British Journal for the History of Science, 35(3): 271–289, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028124.

Lightman:2007:VPS

[Lig07] Bernard V. Lightman. Victorian popularizers of science: designing nature for new audiences. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-48118-2 (cloth). xvi + 545 pp. LCCN Q127.G4 L54

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0714/2007015179-t.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2007015179-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007015179-b.html.

Linden:2003:ARH

[Lin03] Stanton J. Linden, editor. The alchemy reader: from Hermes Trismegistus to Isaac Newton. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-79234-7, 0-521-79662-8 (paperback). [xi], 260 pp. LCCN QD26 .A585 2003. URL http://www.

loc.gov/catdir/description/
cam031/2002041453.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2002041453-b.html;
http://www.loc.gov/catdir/
toc/cam031/2002041453.html.

Lindberg:2005:BRR

[Lin05] David C. Lindberg. Book review: Richard S. Brooks and David K. Himrod, Science and Religion in the English-Speaking World, 1600–1727: A Bibliographic Guide to the Secondary Literature. American Theological Library Association Bibliography Series, 46. Lanham, MD and London: Scarecrow Press, 2001. Pp. xxxiv + 620. ISBN 0-8018-4011-1. $85.00 (hardback). British Journal for the History of Science, 38(1):107, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028589.

Lindee:2007:BRR

[Lin07] Susan Lindee. Book review: Rebecca M. Herzig, Suffering for Science: Reason and Sacrifice in Modern America. New Brunswick, NJ and London: Rutgers University Press, 2005. Pp. 194. ISBN: 0-8135-3662-6. $39.95 (hardback). British Journal for the History of Science, 40(2):302–304, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500738.

Lipphardt:2008:BRG

[Lip08] Veronika Lipphardt. Book reviews: Geoffrey Cantor and Marc Swetlitz (eds.), Jewish Tradition and the Challenge of Darwinism. Chicago and London: The University of Chicago Press, 2006. Pp. xii + 260. ISBN 0-226-09276-3. $60.00, £38.00 (hardback). ISBN 0-226-09277-1. $24.00, £15.00 (paperback). British Journal for the History of Science, 41(2):304–305, June 2008. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Livingstone:2002:SSH

[Liv02] David N. Livingstone, editor.

Science, space and hermeneutics: Hettner-Lecture 2001, volume 5 of Hettner-Lectures. Department of Geography, University of Heidelberg, Heidelberg, Germany, 2002. ISBN 3-88570-505-2. 116 pp. LCCN GF21 .S37 2002.

Livingstone:2003:SRG

[Liv03] David N. Livingstone. Science, religion and the geography of reading: Sir William Whitla and the editorial staging of Isaac Newton’s writings on biblical prophecy. British Journal for the History of Science, 36(1):27–42, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028291.

Livesey:2005:BRW

[Liv05a] James Livesey. Book review:

L. W. B. Brockliss, Calvet’s Web: Enlightenment and the Republic of Letters in Eighteenth-Century France. Oxford: Oxford University Press, 2002. Pp. xx + 471. ISBN 0-19-924748-X. £55.00 (hardback). British Journal for the History of Science, 38 (1):109–110, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028591.

Livingstone:2005:TTT

[Liv05b] David N. Livingstone. Text, talk and testimony: geographical reflections on scientific habits. An afterword. British Journal for the History of Science, 38(1):93– 100, March 2005. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Livingstone:2008:AAR

[Liv08] David N. Livingstone. Adam’s ancestors: race, religion, and the politics of human origins. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8813-1 (hardcover). x + 301 pp. LCCN BL256 .L578

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0805/2007033706-b.html;
http://www.loc.gov/catdir/
enhancements/fy0805/
2007033706-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0725/2007033706.html.

Lloyd:2003:GDS

[Llo03]  G.  E.  R.

(Geoffrey Ernest Richard) Lloyd. In the grip of disease: studies in the Greek imagination. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-925323-4. xxi + 258 pp. LCCN B187.M4 L56

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0615/2003271017-d.html;
http://www.loc.gov/catdir/
enhancements/fy0615/

2003271017-t.html;
http://www.loc.gov/catdir/
enhancements/fy0724/
2003271017-b.html.

Lloyd:2004:AWM

[Llo04]  G.  E.

R. (Geoffrey Ernest Richard) Lloyd. Ancient worlds, modern reflections: philosophical perspectives on Greek and Chinese science and culture. Clarendon Press, New York, NY, USA, 2004. ISBN 0-19-927016-3. xi + 222 pp. LCCN Q124.95 .L582

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0620/2004298453-d.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
2004298453-b.html;
http://www.loc.gov/catdir/
toc/fy051/2004298453.html.

Lloyd:2006:PPA

[Llo06]  G.  E.

R. (Geoffrey Ernest Richard) Lloyd. Principles and practices in ancient Greek and Chinese science, volume CS849 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, England, 2006. ISBN 0-86078-993-4. xvi + 302 pp. LCCN Q127.G7 L594

2006. URL http://www.loc.
gov/catdir/toc/ecip0614/
2006016387.html.

Lloyd:2007:CVR

[Llo07]  G.  E.  R.  (Geoffrey

Ernest Richard) Lloyd. Cognitive variations: reflections on the unity and diversity of the human mind. Clarendon Press, New York, NY, USA, 2007. ISBN 0-19-921461-1. viii + 201 pp. LCCN BF311 .L595 2007.

Luckhurst:2002:TES

[LM02] Roger Luckhurst and Josephine McDonagh, editors. Transactions and encounters: science and culture in the Nineteenth Century. Manchester University Press, Manchester, UK, 2002. ISBN 0-7190-5910-0 (hardcover), 0-7190-5911-9 (paperback). viii + 239 pp. LCCN Q175.5 .T73

                                        2002. URL http://www.loc.
                                        gov/catdir/bios/hol051/
                                        2002026513.html;
                                        http://www.loc.gov/catdir/
                                        description/hol031/
                                        2002026513.html;
                                        http://www.loc.gov/catdir/
                                        toc/hol031/2002026513.html.

Lindberg:2003:WSC

[LN03]  David  C.

Lindberg and Ronald L. Numbers, editors. When science and Christianity meet. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-48214-6. xii + 357 pp. LCCN BL245 .W485 2003. URL http://www.

                                        loc.gov/catdir/bios/
                                        uchi051/2002155569.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi052/
                                        2002155569.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy043/2002155569.html.

Locher:2002:BRJ

[Loc02] Kurt Locher. Book review: John M. Steele, Observations and Predictions of Eclipse Times by Early Astronomers. Archimedes: New Studies in the History and Philosophy of Science and Technology, 4. Dordrecht: Kluwer Academic Publishers, 2000. Pp. xi + 321. ISBN 0-7923-6298-5. £84.00 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Locher:2007:OLB

[Loc07] Fabien Locher. The observatory, the land-based ship and the crusades: earth sciences in European context, 1830–1850. British Journal for the History of Science, 40(4):491–504, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Logan:2008:BRC

[Log08] Cheryl Logan. Book reviews:

Chandak Sengoopta, The Most Secret Quintessence of Life: Sex, Glands and Hormones, 1850– 1950. Chicago and London: The University of Chicago Press, 2006. Pp. xii + 354. ISBN 0-226-74863-4. $45.00, £28.50 (hardback). Judith A. Houck, Hot and Bothered: Women, Medicine, and Menopause in Modern America. Cambridge, MA and London: Harvard University Press, 2006. Pp. xii+328. ISBN 0-674-01896-6. £24.95, $39.95 (hardback). British Journal for the History of Science, 41(2):286– 288, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Long:2001:OSA

[Lon01] Pamela O. Long. Openness, secrecy, authorship: technical arts and the culture of knowledge from antiquity to the Renaissance. The Johns Hopkins University Press, Baltimore, MD, USA, 2001. ISBN 0-8018-6606-5. xii + 364 pp. LCCN CB478 .L65 2001. URL http://www.

                                        loc.gov/catdir/bios/jhu051/
                                        00010152.html;
                                        http://www.loc.gov/catdir/
                                        description/jhu051/
                                        00010152.html;
                                        http://www.loc.gov/catdir/
                                        toc/jhu051/00010152.html.

Loveland:2001:RNH

[Lov01] Jeff Loveland. Rhetoric and natural history: Buffon in polemical and literary context, volume 2001:03 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2001. ISBN 0-7294-0747-0. ix + 214 pp. LCCN QH45.2 .L68 2001.

Lozovsky:2000:EOB

[Loz00] Natalia Lozovsky. The Earth is our book: geographical knowledge in the Latin West ca. 400– 1000. Recentiores. University of Michigan Press, Ann Arbor, MI, USA, 2000. ISBN 0-472-11132-9 (cloth). viii + 182 pp. LCCN G89 .L68

                                        2000. URL http://www.loc.

gov/catdir/description/
umich051/99050981.html.

Lustig:2004:DH

[LRR04]  Abigail  Lustig,  Robert

J. (Robert John) Richards, and Michael Ruse, editors. Darwinian heresies. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-81516-9. viii + 200 pp. LCCN QH366.2 .D342

2004. URL http://www.loc.
gov/catdir/description/
cam041/2003069670.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2003069670-b.html;
http://www.loc.gov/catdir/
samples/cam051/2003069670.
html;
http://www.loc.gov/catdir/
toc/cam041/2003069670.html.

Lefevre:2003:PIE

[LRS03] Wolfgang Lefèvre, Jürgen Renn, and Urs Schoepflin, editors. The power of images in early modern science. Birkhäuser Verlag, Basel, Switzerland, 2003. ISBN 3-7643-2434-1, 0-8176-2434-1. ix + 308 pp. LCCN Q223 .P69

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0812/2003060089-d.html;
http://www.loc.gov/catdir/
enhancements/fy0812/
2003060089-t.html.

Lorimer:2005:LFS

[LS05] Hayden Lorimer and Nick Spedding. Locating field science: a geographical family expedition to Glen Roy, Scotland. British Journal for the History of Science, 38(1):13–33, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Levere:2002:DCS

[LTGS02] Trevor Harvey Levere, Gerard L’Estrange Turner, Jan Golinski, and Larry Stewart. Discussing chemistry and steam: the minutes of a coffee house philosophical society, 1780–1787. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-851530-8. viii + 284 pp. LCCN Q157 .L447 2002.

URL http://www.e-streams.
com/es0604/es0604_2436.
html;   http://www.loc.gov/
catdir/enhancements/fy0614/
2002510571-d.html;
http://www.loc.gov/catdir/
enhancements/fy0614/
2002510571-t.html.

Luckhurst:2002:IT

[Luc02] Roger Luckhurst. The invention of telepathy, 1870–1901. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-924962-8. viii + 324 pp. LCCN BF1171 .L77

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0614/2002512654-d.html;
http://www.loc.gov/catdir/
enhancements/fy0614/
2002512654-t.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2002512654-b.html.

Lucas:2005:BRF

[Luc05] Adam Lucas. Book review:

Francesc Relaño, The Shaping of Africa: Cosmographic Discourse and Cartographic Science in Late Medieval and Early Modern Europe. Aldershot: Ashgate, 2002. Pp. x + 271. ISBN 0-754-60239-7. £52.50 (hardback). British Journal for the History of Science, 38(4):477–478, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028462.

Luthy:2004:BRA

[Lüt04a] Christoph Lüthy. Book review:

Antonio Clericuzio, Elements, Principles and Corpuscles: A Study of Atomism and Chemistry in the Seventeenth Century. International Archives of the History of Ideas, 171. Dordrecht, Boston and London: Kluwer Academic Publishers, 2000. Pp. xi + 223. ISBN 0-7923-6782-0. £65.00, $103.00 (hardback). British Journal for the History of Science, 37(1):103–104, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028262.

Luthy:2004:BRM

[Lüt04b] Christoph Lüthy. Book review: Marie Boas Hall, Henry Oldenburg: Shaping the Royal Society. Oxford: Oxford University Press, 2002. Pp. xii + 369. ISBN 0-19-851053-5. £60.00 (hardback). British Journal for the History of Science, 37 (2):201–203, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028335.

Livingstone:2005:GR

[LW05]  David  N.

Livingstone and Charles W. J. Withers, editors. Geography and revolution. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-48733-4. viii + 433 pp. LCCN G70 .G4419

                                        2005. URL http://www.h-net.
                                        org/review/hrev-a0f4d6-aa;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0621/
                                        2005003443-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0621/
                                        2005003443-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip057/2005003443.html.

Lynch:2002:CSG

[Lyn02] John M. Lynch, editor. Creationism and scriptural geology. Thoemmes, Bristol, UK, 2002. ISBN 1-85506-928-8. 3645 pp. LCCN BS657 .C74 2002. Eight volumes.

Lynch:2009:BRD

[Lyn09] John M. Lynch. Book review: David N. Livingstone, Adam’s Ancestors: Race, Religion, and the Politics of Human Origins. Baltimore: Johns Hopkins University Press, 2008. Pp. x + 301. ISBN 978-0-8018-8813-7. £23.50 (hardback). British Journal for the History of Science, 42(3): 435–437, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Maas:2005:WSJ

[Maa05] Harro Maas. William Stanley Jevons and the making of modern economics. Historical perspectives on modern economics. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-82712-4. xxii + 330 pp. LCCN HB103.J5 M33

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2004052546-d.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2004052546-t.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2004052546-b.html.

MacKenzie:2000:BRD

[Mac00a] John M. MacKenzie. Book review: Deepak Kumar. Science and the Raj, 1857–1905. Delhi: Oxford University Press, 1997. Pp. xv + 273. ISBN 0-19-564194-9. £4.50, $10.95. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027935.

MacKenzie:2000:BRJ

[Mac00b] John M. MacKenzie. Book review: John Gascoigne, Science in the Service of Empire: Joseph Banks, the British State and the Uses of Science in the Age of Revolution. Cambridge: Cambridge University Press, 1998. Pp. viii + 247. ISBN 0-521-55069-6. £40.00, $64.95. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4027960.

MacLeod:2000:CSV

[Mac00c] Roy M. MacLeod. The “creed of science” in Victorian England, volume CS598 of Variorum collected studies series. Ashgate/ Variorum, Aldershot, Hampshire, England, 2000. ISBN 0-86078-669-2. xxiv + 318 pp. LCCN Q127.G7 M33 2000.

Macmillan:2000:OKF

[Mac00d] Malcolm Macmillan. An odd kind of fame: stories of Phineas Gage. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-13363-6. xiii + 562 pp. LCCN RC387.5.G34 M33 2000.

Macdonald:2001:BRS

[Mac01a] Sharon Macdonald. Book review: Svante Lindqvist (ed.) with Marika Hedin and Ulf Larsson (associate eds.), Museums of Modern Science. Nobel Symposium, 112. Canton: Science History Publications, 2000. Pp. xii + 200. ISBN 0-88135-299-3. $39.95. British Journal for the History of Science, 34(1): 97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028051.

MacKenzie:2001:MPC

[Mac01b] Donald A. MacKenzie. Mechanizing proof: computing, risk, and trust. Inside technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-13393-8 (hardcover). xi + 427 pp. LCCN QA76.76.R44 M36 2001. URL

http://www.loc.gov/catdir/
toc/fy022/2001018687.html.

Macdonald:2002:BSS

[Mac02] Sharon Macdonald. Behind the scenes at the Science Museum. Materializing culture. Berg, Oxford, UK, 2002. ISBN 1-85973-566-5 (cloth), 1-85973-571-1 (paper). xiii + 293 pp. LCCN Q105.G72 L6835 2002. URL

http://www.loc.gov/catdir/
bios/hol059/2001006856.
html;   http://www.loc.gov/
catdir/description/hol051/
2001006856.html;
http://www.loc.gov/catdir/
toc/fy034/2001006856.html.

Macleod:2005:BRL

[Mac05] Roy Macleod. Book review:

Lawrence Goldman, Science, Reform and Politics in Victorian Britain: The Social Science Association, 1857–1886. Cambridge: Cambridge University Press, 2002. Pp. 419+index. ISBN 0-521-33053. British Journal for the History of Science, 38(4):488–490, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028471.

Macdonald:2006:F

[Mac06a] Helen Macdonald. Falcon. Animal (Reaktion Books). Reaktion, London, UK, 2006. ISBN 1-86189-238-1 (paperback). 208 pp. LCCN QL696.F34 M33

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0701/2006491202-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0701/
                                        2006491202-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0701/
                                        2006491202-t.html.

Macleod:2006:BRA

[Mac06b] Christine Macleod. Book review: Anna Guagnini and Ian Inkster (eds.), Patents in History. History of Technology, Vol. 24. London and New York: Continuum, 2004. Pp. xxi + 242. ISBN 0-8264-7186-2. £75.00. British Journal for the History of Science, 39(3):438–439, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028487.

Macleod:2007:BRM

[Mac07]  Roy Macleod.  Book

review: Mary J O Nye, Blackett: Physics, War, and Politics in the Twentieth Century. Cambridge, MA and London: Harvard University Press, 2004. Pp. x + 255. ISBN 0-674-01548-7.

$25.95 (hardback). British Journal for the History of Science, 40 (1):150–152, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500709.

MacLeod:2008:BRD

[Mac08a] Christine MacLeod. Book review: David E. Nye, Technology Matters: Questions to Live With. Cambridge, MA and London: Massachusetts Institute of Technology Press, 2006. Pp. xiv + 282. ISBN: 0-262-14093-4. £18.95 (hardback). British Journal for the History of Science, 41 (1):120–122, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacLeod:2008:BRB

[Mac08b] Roy MacLeod. Book review:

Bart Schultz, Henry Sidgwick — Eye of the Universe: An Intellectual Biography. Cambridge: Cambridge University Press, 2004. Pp. xx + 858. ISBN 0-521-82967-4. £40.00 (hardback). British Journal for the History of Science, 41(3): 464–465, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Macleod:2008:BRM

[Mac08c] Roy Macleod. Book review:

Mark Wheelis, Lajos Rózsa and Malcolm Dando (eds.), Deadly Cultures: Biological Weapons since 1945. Cambridge, MA and London: Harvard University Press, 2006. Pp. xi + 479. ISBN 0-0674-01699-8. £37.95, EUR51.00, $59.95 (hardback). British Journal for the History of Science, 41(1):124–126, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacLeod:2009:BRL

[Mac09a] Christine MacLeod. Book review: Lissa Roberts, Simon Schaffer and Peter Dear (eds.), The Mindful Hand: Inquiry and Invention from the Late Renaissance to Early Industrialisation. Amsterdam: Koninklijke Nederlandse Akademie van Wetenschappen, 2007. Available outside the UK and Europe from University of Chicago Press. Pp. xxvii + 503. ISBN 978-90-6984-483-1. EUR89.00, $110.00 (hardback). British Journal for the History of Science, 42(2):296– 298, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

MacLeod:2009:BRD

[Mac09b] Roy MacLeod. Book review: David Edgerton, Warfare State: Britain, 1920–1970. Cambridge: Cambridge University Press, 2005. Pp. xv + 364. ISBN 0-521-67231-7. £45.00, $75.00 (paperback). British Journal for the History of Science, 42(4): 628–629, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Maerker:2006:THM

[Mae06] Anna Maerker. The tale of the hermaphrodite monkey: classification, state interests and natural historical expertise between museum and court, 1791– 1794. British Journal for the History of Science, 39(1):29–47, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Magnani:2001:PGT

[Mag01]  Lorenzo Magnani.  Philoso-

phy and geometry: theoretical and historical issues, volume 66 of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6933-5. xix + 249 pp. LCCN QA447 .M34

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0821/2001022473-d.html;
http://www.loc.gov/catdir/
enhancements/fy0821/
2001022473-t.html.

Magnello:2006:BRP

[Mag06a] M. Eileen Magnello. Book review: Paul M. M. Klep and Ida H. Stamhuis (eds.), The Statistical Mind in a Pre-statistical Era: The Netherlands 1750– 1850. Askant: Amsterdam, 2002. Pp. 374. ISBN 90-5742-0341. No price given. British Journal for the History of Science, 39(3): 449–451, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028495.

Magnello:2006:BRT

[Mag06b] M. Eileen Magnello. Book review: Theodore M. Porter, Karl Pearson: The Scientific Life in a Statistical Age. Princeton, NJ and London: Princeton University Press, 2004. Pp. viii + 342. ISBN 0-691-11445-5. £22.95, $35.00 (hardback). British Journal for the History of Science, 39(4):619–620, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028538.

Mahon:2003:MWC

[Mah03] Basil Mahon. The man who changed everything: the life of James Clerk Maxwell. John Wiley, New York, NY, USA, 2003. ISBN 0-470-86088-X, 0-470-86171-1 (paperback). xx + 226 + 16 pp. LCCN QC16.M4

                                        M34 2003. URL http://www.
                                        loc.gov/catdir/bios/
                                        wiley047/2004426225.html;
                                        http://www.loc.gov/catdir/
                                        description/wiley042/
                                        2004426225.html;   http://
                                        www.loc.gov/catdir/toc/
                                        wiley041/2004426225.html.

Maienschein:2009:BRT

[Mai09] Jane Maienschein. Book review: Tatjana Buklijas and Nick Hopwood, Making Visible Embryos. An online exhibition developed by the University of Cambridge with support from the Wellcome Trust, 2008. British Journal for the History of Science, 42(2):287–288, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). Available at

http://www.hps.cam.ac.uk/visibleembryos/.

Manzo:2000:BRF

[Man00a] Silvia Manzo. Book review:

Francis Bacon, Philosophical Studies, c. 1611 – c. 1619, edited by Graham Rees. The Oxford Francis Bacon, VI. Oxford: Clarendon Press, 1996. Pp. cxvi + 503. ISBN 0-19-812290-X. £80.00. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027936.

Manzo:2000:BRM

[Man00b] Silvia A. Manzo. Book review: Miguel Angel Granada, El debate cosmológico en 1588. Bruno, Brahe, Rothmann, Ursus, Röslin. Istituto Italiano per gli Studi Filosofici, Lezioni della Scuola di Studi Superiori in Napoli, Bibliopolis, Napoli, 1996. Pp. 166. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027961.

Manzo:2001:BRF

[Man01a] Silvia Manzo. Book review: Francis Bacon, The New Organon. Edited by Lisa Jardine and Michael Silverthorne. Cambridge Texts in the History of Philosophy. Cambridge: Cambridge University Press, 2000. Pp. xxxv + 252. ISBN 0-521-56399-2. $59.95 (hardback); 0–521–56483–2. $18.95 (paperback). British Journal for the History of Science, 34(1): 97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028052.

Manzo:2001:BRP

[Man01b] Silvia Manzo. Book review: Peter Zigman (ed.), Einblicke in eine sterbende Ära: Das Ende des Mythos der guten alten Zeiten. Philosophica XXXII. Bratislava: Comenius-Universität, 2000. Pp. 191. ISBN 80-223-1425-0. No price given (paperback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028114.

Manzo:2003:AVS

[Man03] Silvia Manzo. The arguments on void in the Seventeenth Century: the case of Francis Bacon. British Journal for the History of Science, 36(1):43–61, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028292.

Mandelbrote:2004:BRR

[Man04] Scott Mandelbrote. Book review: Robert Crocker (ed.), Religion, Reason and Nature in Early Modern Europe. International Archives of the History of Ideas, 180. Dordrecht, Boston and London: Kluwer Academic Publishers, 2001. Pp. xix + 228. ISBN 1-4020-0047-2. £65.00, $103.00, EUR93.50 (hardback). British Journal for the History of Science, 37(4): 470–471, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028648.

Manzo:2006:BRF

[Man06] Silvia Manzo. Book review:

Francis Bacon, The Instauratio Magna Part II: Novum Organum and Associated Texts. Edited with introduction, notes, commentaries and facing-page translations by Graham Rees with Maria Wakely. The Oxford Francis Bacon, XI. Oxford: Clarendon Press, 2004. Pp. cxxviii + 634. ISBN 0-19-924792-7. £120.00 (hardback). British Journal for the History of Science, 39(2):290–291, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028617.

Marsden:2004:PTF

[Mar04] Ben Marsden. ‘the progeny of these two “Fellows”’: Robert Willis, William Whewell and the sciences of mechanism, mechanics and machinery in early Victorian Britain. British Journal for the History of Science, 37(4):401–434, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028149;
                                        http://www.jstor.org/
                                        stable/4028641.

Marie:2005:BRA

[Mar05] Jenny Marie. Book review: Alan H. Goodman, Deborah Heath and M. Susan Lindee (eds.), Genetic Nature/Culture: Anthropology and Science beyond the Two-Culture Divide. Berkeley, Los Angeles and London: University of California Press, 2003. Pp. xvii + 311. ISBN 0-520-23793-5. £16.95, $24.95 (paperback). British Journal for the History of Science, 38(4): 495–496, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028476.

Margocsy:2009:ACR

[Mar09] Dániel Margócsy. Advertising cadavers in the republic of letters: anatomical publications in the early modern Netherlands. British Journal for the History of Science, 42(2):187– 210, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mayer:2005:WTD

[May05a] Anna-K. Mayer. When things don’t talk: knowledge and belief in the inter-war humanism of Charles Singer (1876–1960). British Journal for the History of Science, 38(3):325–347, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mayhew:2005:MSI

[May05b] Robert Mayhew. Mapping science’s imagined community: geography as a Republic of Letters, 1600–1800. British Journal for the History of Science, 38(1):73– 92, March 2005. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mazzotti:2000:SPK

[Maz00] Massimo Mazzotti. For science and for the Pope-king: writing the history of the exact sciences in nineteenth-century Rome. British Journal for the History of Science, 33(3):257– 282, September 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027953.

Mazzotti:2002:BRJb

[Maz02a] Massimo Mazzotti. Book review: Jean-Paul Pier (ed.), Development of Mathematics 1950– 2000. Basel, Boston and Berlin: Birkhäuser, 2000. Pp. x + 1372.

ISBN 3-7643-6280-4. DM 298.00 (hardback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028197.

Mazzotti:2002:BRJa

[Maz02b] Massimo Mazzotti. Book review: Jeremy J. Gray, The Hilbert Challenge. Oxford: Oxford University Press, 2000. Pp. xii + 315. ISBN 0-19-850651-1. £20.00 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028413.

Mazumdar:2004:BRM

[Maz04a] Pauline M. H. Mazumdar. Book review: A. M. Moulin and A. Cambrosio (eds.), Singular Selves: Historical Issues and Contemporary Debates in Immunology/Dialogues entre soi: Questions historiques et débats contemporains en immunologie. Amsterdam: Elsevier/Musée Claude Bernard, 2001. Pp. 303. ISBN 2-84-299-210-5. EUR59.00 (paperback). British Journal for the History of Science, 37(4): 486–487, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028662.

Mazzotti:2004:NLG

[Maz04b] Massimo Mazzotti. Newton for ladies: gentility, gender and radical culture. British Journal for the History of Science, 37 (2):119–146, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028327.

Mazzotti:2007:BRT

[Maz07a] Massimo Mazzotti. Book reviews: The two Newtons and beyond J. E. Force and S. Hutton (eds.), Newton and Newtonianism: New Studies. International Archives of the History of Ideas 188. Dordrecht, Boston and London: Kluwer, 2004. Pp. xvii + 246. ISBN 1-4020-1969-6. £67.00 (hardback). Rob Iliffe, Milo Keynes and Rebekah Higgitt (eds.), Early Biographies of Isaac Newton 1660–1885. Vol. 1: Eighteenth-Century Biography of Isaac Newton: The Unpublished Manuscripts and Early Texts. Vol. 2: Nineteenth-Century Biography of Isaac Newton: Private Debate and Public Controversy. London: Pickering and Chatto, 2006. Pp. lxxii+387 and xliii+420. ISBN 1-85-196778-8. £195.00 (hardback). Milo Keynes, The Iconography of Sir Isaac Newton to 1800. Woodbridge: The Boydell Press, 2005. Pp. viii+120. ISBN 1-84383-133-3. £40.00 (hardback). John Henry (ed.), Newtonianism in Eighteenth-Century Britain. 7 vols. Bristol: Thoemmes Continuum, 2004. ISBN 1-84371-113-3. £595.00 (hardback). Mordechai Feingold, The Newtonian Moment: Isaac Newton and the Making of Modern Culture. New York and Oxford: The New York Public Library and Oxford University Press, 2004. Pp. xv+218. ISBN 0-19-517735-5. £25.00 (hardback). Margaret C. Jacob and Larry Stewart, Practical Matter: Newton’s Science in the Service of Industry and Empire 1687–1851. Cambridge, MA and London: Harvard University Press, 2004. Pp. 201. ISBN 0-674-01497-9. £22.95 (hardback). British Journal for the History of Science, 40(1):105–111, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500684.

Mazzotti:2007:WMG

[Maz07b] Massimo Mazzotti. The world of Maria Gaetana Agnesi, mathematician of God. Johns Hopkins studies in the history of mathematics. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8709-7 (paperback). xx + 217 pp. LCCN QA29.A28 M39

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0713/2007006287-b.html;
http://www.loc.gov/catdir/
enhancements/fy0713/
2007006287-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0711/2007006287.html.

Mazzotti:2008:KSO

[Maz08] Massimo Mazzotti, editor. Knowledge as social order: rethinking the sociology of Barry Barnes. Ashgate, Aldershot, Hampshire, England, 2008. ISBN 0-7546-4863-X. xi + 184 pp. LCCN HM651 .K563

2008. URL http://www.loc.
gov/catdir/toc/ecip0725/
2007034124.html.

Mazzotti:2009:BRR

[Maz09] Massimo Mazzotti. Book review: Rebekah Higgitt, Recreating Newton: Newtonian Biography and the Making of Nineteenth-Century History of Science. London: Pickering & Chatto, 2007. Pp. ix + 286. ISBN 978-1-85196-906-7. £60.00, $99.00 (hardback). British Journal for the History of Science, 42 (1):121–122, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Misa:2003:MT

[MBF03] Thomas J. Misa, Philip Brey, and Andrew Feenberg, editors. Modernity and technology. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-13421-7. vi + 421 pp. LCCN T14.5 .M63

2003. URL http://www.loc.
gov/catdir/toc/fy038/
2002071754.html.

McCook:2002:SNS

[McC02] Stuart George McCook. States of nature: science, agriculture, and environment in the Spanish Caribbean, 1760–1940. University of Texas Press, Austin, TX, USA, 2002. ISBN 0-292-75256-3, 0-292-75257-1. xiv + 201 pp. LCCN S477.A1 M33

2002. URL http://www.h-net.
org/review/hrev-a0d4d9-aa;
 http://www.loc.gov/catdir/
description/texas041/
2001052228.html;
http://www.loc.gov/catdir/
toc/fy031/2001052228.html.

Mccluskey:2005:BRA

[Mcc05] Stephen C. Mccluskey. Book review: Asger Aaboe, Episodes from the Early History of Astronomy. New York: Springer-Verlag, 2001. Pp. xv + 172. ISBN 0-387-95136-9. £33.00, $49.95 (paperback). British Journal for the History of Science, 38(3): 356–358, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028678.

Mcginnis:2006:MAA

[Mcg06] Jon Mcginnis. A medieval Arabic analysis of motion at an instant: the Avicennan sources to the forma fluens/fluxus formae debate. British Journal for the History of Science, 39(2):189– 205, June 2006. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McGinnis:2008:BRS

[McG08] Jon McGinnis. Book review: Simone van Riet, Jules Janssens and André Allard (eds.) Avicenna Latinus, Liber primus naturalium, tractatus secundus: De motu et de consimilibus. Introduction by Gérard Verbeke. Leuven: Peeters, 2006. Pp. lxxxix + 373. ISBN 978-2-8031-0231-0. £173.70 (hardback). British Journal for the History of Science, 41(1):131– 132, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McLean:2000:RBI

[McL00] Donald F. McLean. Restoring Baird’s image, volume 27 of IEE history of technology series. Institution of Electrical Engineers (IEE), London, UK, 2000. ISBN 0-85296-795-0. xx + 295 pp. LCCN TK6635.B3 M36

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0631/2004533191-t.html.

McLean:2007:CSM

[McL07] Matthew McLean, (Matthew Adam). The cosmographia of Sebastian Münster: describing the world in the Reformation. St. Andrews studies in Reformation history. Ashgate, Aldershot, Hampshire, England, 2007. ISBN 0-7546-5843-0. viii + 378 pp. LCCN G116 .M39

2007. URL http://www.loc.
gov/catdir/toc/ecip072/
2006032293.html.

McLean:2009:EFH

[McL09]  Steven McLean.  The early

fiction of H. G. Wells: fantasies of science. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2009. ISBN 0-230-53562-3. ix + 242 pp. LCCN PR5778.S35 M35

                                        2009. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1008/2008050876-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1008/
                                        2008050876-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy1008/
                                        2008050876-t.html.

McMullin:2005:CG

[McM05] Ernan McMullin, editor. The Church and Galileo. Studies in science and the humanities from the Reilly Center for Science, Technology, and Values. University of Notre Dame Press, Notre Dame, IN, USA, 2005. ISBN 0-268-03483-4, 0-268-03484-2 (paperback). xii + 391 pp. LCCN QB36.G2 C48 2005. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip057/2005002602.html.

Mcouat:2001:CPD

[Mco01] Gordon Mcouat. Cataloguing power: delineating ‘competent naturalists’ and the meaning of species in the British Museum. British Journal for the History of Science, 34(1):1–28, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028036.

McOuat:2009:BRJ

[McO09] Gordon McOuat. Book review: Jim Endersby, Imperial Nature: Joseph Hooker and the Practices of Victorian Science. Chicago and London: University of Chicago Press, 2008. Pp. x + 429. ISBN 978-0-226-2-791-9. £20.50 (hardback). British Journal for the History of Science, 42 (1):122–124, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mota:2009:EPE

[MCS09] Elsa Mota, Paulo Crawford, and Ana Simões. Einstein in Portugal: Eddington’s expedition to Principe and the reactions of Portuguese astronomers (1917– 1925). British Journal for the History of Science, 42(2):245– 273, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McTavish:2005:CDA

[McT05] Lianne McTavish. Childbirth and the display of authority in early modern France. Women and gender in the early modern world. Ashgate, 2005. ISBN 0-7546-3619-4. xiv + 257 pp. LCCN RG652 .M385 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0417/2004008774.html.

McClellan:1999:STW

[MD99] James E. (James Edward) Mc-Clellan and Harold Dorn. Science and technology in world history: an introduction. The Johns Hopkins University Press, Baltimore, MD, USA, 1999. ISBN 0-8018-5868-2, 0-8018-5869-0 (paperback). viii + 404 pp. LCCN Q125 .M414

                                        1999. URL http://www.h-net.
                                        org/review/hrev-a0b4i1-aa;
                                         http://www.loc.gov/catdir/
                                        bios/jhu051/98028898.html;
                                         http://www.loc.gov/catdir/
                                        description/jhu052/
                                        98028898.html.

Mollan:2002:IIS

[MDF02] R. C. (R. Charles) Mollan, William (William J.) Davis, and Brendan Finucane, editors. Irish innovators in science and technology. Royal Irish Academy, Dublin, Ireland, 2002. ISBN 1-874045-88-7. xv + 256 pp. LCCN Q141 .I79 2002.

Meadows:2004:VSG

[Mea04] A. J. (Arthur Jack) Meadows. The Victorian scientist: the growth of a profession. British Library, London, UK, 2004. ISBN 0-7123-0894-6. 202 pp. LCCN Q141 .M383 2004.

Medina:2000:BRJ

[Med00] Rosa Medina. Book review: José Ma López Piñero et al. (eds.), Bibliographia Medica Hispanica, 1475–1950, Volumen V, Libros y Folletos, 1851–1900, Caudemos Valencianos de Historia de la Medicina y de la Ciencia, L. Valencia, Instituto de Estudios Documentales e Históricos sobre la Clencia Valencia: Universitat de València-C.S.I.C., 1996. Pp. 956. ISBN 84-370-2349-1. No price given. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027937.

Meinel:2000:IEH

[Mei00] Christoph Meinel, editor. Instrument — Experiment: historische Studien. (German) [Instrument — Experiment: historical study]. Verlag für Geschichte der Naturwissenschaften und der Technik, Berlin, Germany, 2000. ISBN 3-928186-51-5. LCCN Q180.A3 I57 2000.

Meli:2001:BRW

[Mel01] Domenico Bertoloni Meli. Book review: W. R. Laird, The Unfinished Mechanics of Giuseppe Moletti: An Edition and English Translation of His Dialogue on Mechanics, 1576. Toronto, Buffalo and London: University of Toronto Press, 2000. Pp. xi + 222. ISBN 0-8020-4699-1. £40.00, $55.00 (hardback). British Journal for the History of Science, 34(4):453– 481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028388.

Menz:2000:MHR

[Men00] Heike Menz. Martin Heinrich Rathke (1793–1860): ein Embryologe des 19. Jahrhunderts. (German) [Martin Heinrich Rathke (1793– 1860): A Nineteenth-Century Embryologist], volume 7 of Acta biohistorica. Basilisken-Presse, Marburg, Germany, 2000. ISBN 3-925347-59-3. 280 pp. LCCN QL954.2.R38 M45 2000.

Merchant:2008:BRP

[Mer08] Carolyn Merchant. Book review: Pierre Hadot, The Veil of Isis: An Essay on the History of the Idea of Nature. Tr. Michael Chase. Cambridge, MA and London: Belknap Press of Harvard University Press, 2006. Pp. xii + 399. ISBN 978-0-674-02316-1. $29.95, £19.95, EUR25.50 (hardback). British Journal for the History of Science, 41(2):288– 289, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Methuen:1998:KTS

[Met98] Charlotte Methuen. Kepler’s Tübingen: stimulus to a theological mathematics. St. Andrews studies in Reformation history. Ashgate, 1998. ISBN 1-85928-397-7. xi + 280 pp. LCCN Q127.G3 M475 1998.

Melling:2006:PMS

[MF06]  Joseph  Melling

and Bill Forsythe. The politics of madness: the state, insanity, and society in England, 1845– 1914. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2006. ISBN 0-415-30174-2 (hardcover). xvii + 278 pp. LCCN HV4995.G7 M45 2006.

                                        URL  http://www.loc.gov/
                                        catdir/enhancements/fy0654/
                                        2005018894-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0516/2005018894.html.

Miller:2000:PEL

[Mil00] Peter N. Miller. Peiresc’s Europe: learning and virtue in the Seventeenth Century. Yale University Press, New Haven, CT, USA, 2000. ISBN 0-300-08252-5. xv + 234 + 16 pp. LCCN DC121.8.P4 M55

2000. URL http://www.h-net.
org/review/hrev-a0c1i0-aa.

Miller:2001:EPS

[Mil01] Arthur I. Miller. Einstein, Picasso: space, time, and beauty that causes havoc. Basic Books, New York, NY, USA, 2001. ISBN 0-465-01859-9 (paperback). x + 357 pp. LCCN N72.S3 M55

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0831/00065130-d.html.

Miller:2002:BRN

[Mil02] David Philip Miller. Book review: Neil Chambers (ed.), The Letters of Sir Joseph Banks: A Selection, 1768–1820. London: Imperial College Press, 2001. Pp. xliii + 420. ISBN 1-86094-204-0. £33.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028198.

Miller:2004:BRR

[Mil04] David Philip Miller. Book review: Richard L. Hills, James Watt, Volume 1: His Time in Scotland, 1736–1774. Landmark Collector’s Library. Ashbourne: Landmark Publishing, 2002. Pp.

480. ISBN 1-84306-045-0. £29.95 (hardback). British Journal for the History of Science, 37 (2):203–206, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028336.

Miller:2005:ESO

[Mil05] Arthur I. Miller. Empire of the stars: obsession, friendship, and betrayal in the quest for black holes. Houghton-Mifflin, Boston, MA, USA, 2005. ISBN 0-618-34151-X. xx + 364 pp. LCCN QB35 .M55

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0618/2004060909-d.html;
http://www.loc.gov/catdir/
enhancements/fy0618/
2004060909-s.html;
http://www.loc.gov/catdir/
enhancements/fy0618/
2004060909-t.html.

Miller:2008:PPP

[Mil08] David Philip Miller. Principle, practice and persona in Isambard Kingdom Brunel’s patent abolitionism. British Journal for the History of Science, 41(1):43– 72, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mirowski:2007:BRH

[Mir07] Philip Mirowski. Book review: Harro Maas, William Stanley Jevons and the Making of Modern Economics. Cambridge: Cambridge University Press, 2005. Pp. xxii + 330. ISBN 0-521-82712-4. $75.00 (hardback). British Journal for the History of Science, 40 (2):297–298, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500734.

Mitton:2005:FHL

[Mit05] Simon Mitton. Fred Hoyle: a life in science. Aurum, London, UK, 2005. ISBN 1-85410-961-8 (hardcover). xi + 369 + 8 pp. LCCN QB36.H75 M584 2005.

Mitman:2007:BSH

[Mit07] Gregg Mitman. Breathing space:

how allergies shape our lives and landscapes. Yale University Press, New Haven, CT, USA, 2007. ISBN 0-300-11035-9. xv + 312 pp. LCCN RC589 .M58

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0704/2006052827-b.html;
http://www.loc.gov/catdir/
enhancements/fy0704/
2006052827-d.html.

Moleti:2000:UMG

[ML00] Giuseppe Moleti and Walter Roy Laird. The unfinished mechanics of Giuseppe Moletti: an edition and English translation of his Dialogue on Mechanics (1576). University of Toronto Press, Toronto, ON, Canada, 2000. ISBN 0-8020-4699-1. viii + 222 pp. LCCN QA804 .M6513 2000.

Morrison-Low:2007:MSI

[ML07] A. D. Morrison-Low. Making scientific instruments in the industrial revolution. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2007. ISBN 0-7546-5758-2. xvi + 408 pp. LCCN HD9706.6.E62

                                        M67 2007. URL http://www.
                                        loc.gov/catdir/toc/
                                        ecip0615/2006018488.html.

Mehra:2000:CMS

[MM00] Jagdish Mehra and Kimball A. Milton. Climbing the mountain: the scientific biography of Julian Schwinger. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850658-9 (hardcover), 0-19-852745-4 (paperback). xii + 677 pp. LCCN QC16.S29 M45

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0606/99087323-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0606/
                                        99087323-t.html.

Millar:2002:CDS

[MMMM02] David Millar, Ian Millar, John Millar, and Margaret Millar, editors. The Cambridge dictionary of scientists. Cambridge University Press, Cambridge, UK, second edition, 2002. ISBN 0-521-80602-X (hardcover), 0-521-00062-9 (paperback). xii + 428 pp. LCCN Q141 .C128 2002.

URL  http://www.loc.gov/
catdir/description/cam031/
2002512240.html;   http://
www.loc.gov/catdir/samples/
cam041/2002512240.html;
http://www.loc.gov/catdir/
toc/cam031/2002512240.html.

Moktefi:2006:BRJ

[Mok06] Amirouche Moktefi. Book review: Jack Meadows, The Victorian Scientist: The Growth of a Profession. London: The British Library, 2004. Pp. vi + 202. ISBN 0-7123-0894-6. £16.95 (hardback). British Journal for the History of Science, 39(4): 613–614, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028534.

Moore:1999:WA

[Moo99]  Patrick Moore.  The wan-

dering astronomer. Institute of Physics Pub., Bristol, UK, 1999. ISBN 0-7503-0693-9. 208 pp. LCCN QB44.2 .M694

1999a. URL http://www.loc.
gov/catdir/enhancements/
fy0653/00709282-d.html.

Moore:2002:GGC

[Moo02] John Alexander Moore. From Genesis to genetics: the case of evolution and creationism. University of California Press, Berkeley, CA, USA, 2002. ISBN 0-520-22441-8 (cloth). xvi + 223 pp. LCCN QH367 .M813

2002. URL http://www.loc.
gov/catdir/bios/ucal052/

2001044419.html;
http://www.loc.gov/catdir/
description/ucal042/
2001044419.html;
http://www.loc.gov/catdir/
toc/fy022/2001044419.html.

Morrell:1997:SOT

[Mor97a] Jack Morrell. Science at Oxford, 1914–1939: transforming an arts university. Clarendon Press, New York, NY, USA, 1997. ISBN 0-19-820657-7. xvii + 473 pp. LCCN Q183.4.G73 O946

1997. URL http://www.loc.
gov/catdir/enhancements/
fy0606/97007976-d.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
97007976-b.html.

Morrell:1997:SCP

[Mor97b] Jack Morrell. Science, culture, and politics in Britain, 1750– 1870, volume CS567 of Collected studies series. Variorum, Aldershot, Great Britain, 1997. ISBN 0-86078-633-1. xii + 336 pp. LCCN Q127.G4 M67 1997.

Morus:1998:FCE

[Mor98] Iwan Rhys Morus. Frankenstein’s children: electricity, exhibition, and experiment in early-Nineteenth-Century London. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-05952-7. xiv + 324 pp. LCCN QC527.5 .M67 1998. URL http://www.

loc.gov/catdir/description/
prin021/98009451.html;
http://www.loc.gov/catdir/
toc/prin032/98009451.html.

Morus:2000:NSB

[Mor00] Iwan Rhys Morus. ‘the nervous system of Britain’: space, time and the electric telegraph in the Victorian age. British Journal for the History of Science, 33(4):455–475, December 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028030.

Morris:2001:BRC

[Mor01a] Peter Morris. Book review:

Colin A. Russell (ed.), Chemistry, Society and Environment: A New History of the British Chemical Industry. Cambridge: The Royal Society of Chemistry, 2000. Pp. xvi + 360. ISBN 0-85404-599-6. £59.50 (hardback). British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028115.

Morus:2001:BRI

[Mor01b] Iwan Rhys Morus. Book review: Ian R. Bartky, Selling the True Time: Nineteenth-Century Timekeeping in America. Stanford: Stanford University Press, 2000. Pp. xvi + 310. ISBN 0-8047-3874-2. £27.50, $45.00. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028019.

Morus:2001:BRM

[Mor01c] Iwan Rhys Morus. Book reviews: Michael Faraday, Experimental Researches in Electricity. Volume I, reprinted from the first edition, 1839. With an introduction by Howard Fisher. Santa Fe: Green Lion Press, 2000. Pp. 10+viii + 574. ISBN 1-888009-09-8; Experimental Researches in Electricity. Volume II, reprinted from the first edition, 1844. Sante Fe: Green Lion Press, 2000. Pp. 2+viii+302. ISBN 1-888009-10-1; Experimental Researches in Electricity. Volume III, reprinted from the first edition, 1855. Sante Fe: Green Lion Press, 2000. Pp. viii+588. ISBN 1-888009-11-X. $120.00 (three hardback volumes). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028389.

Morris:2002:BRD

[Mor02a] Peter Morris. Book review: Denis Hogan and Bryce Williamson (eds.), New Zealand is Different: Chemical Milestones in New Zealand History. Revised edition. Christchurch: Clerestory Press for New Zealand Institute of Chemistry, 2000. Pp. xii + 308. ISBN 0-9583706-7-2. £20.00 (paperback). British Journal for the History of Science, 35(4): 475–485, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mort:2002:BTN

[Mor02b] Maggie Mort. Building the Trident network: a study of the enrollment of people, knowledge, and machines. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-13397-0 (hardcover). x + 217 pp. LCCN V993 .M67

2002. URL http://www.loc.
gov/catdir/toc/fy022/
2001034275.html.

Morus:2002:BRM

[Mor02c] Iwan Rhys Morus. Book review:

Myles W. Jackson, Spectrum of Belief: Joseph von Fraunhofer and the Craft of Precision Optics. Transformations: Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2000. Pp. x + 284. ISBN 0-262-10084-3. £23.95 (cloth). British Journal for the History of Science, 35 (1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028414.

Morrell:2003:BRM

[Mor03a] Jack Morrell. Book review: S.

M. Walters and E. A. Stow, Darwin’s Mentor: John Stevens Henslow 1796–1861. Cambridge: Cambridge University Press, 2001. Pp. xx + 338. ISBN 0-521-59146-5. £40.00 (hardback). British Journal for the History of Science, 36(4):482–483, December 2003. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4028222;   http://www.
                                        jstor.org/stable/4028282.

Morris:2003:BRF

[Mor03b] Peter J. T. Morris. Book review: Fred Aftalion, A History of the International Chemical Industry: From the ‘Early Days’ to 2000. Second revised edition. Translated by Otto Theodor Benfey. Philadelphia: Chemical Heritage Press, 2001. Pp. xxii + 442. ISBN 0-941901-29-7. $24.95 (paperback). British Journal for the History of Science, 36 (2):253–254, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028250.

Moran:2005:DKA

[Mor05a] Bruce T. Moran. Distilling knowledge: alchemy, chemistry, and the scientific revolution. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01495-2. 210 pp. LCCN QD15 .M67 2005.

Morrell:2005:BRJ

[Mor05b] Jack Morrell. Book review: John C. Thackray, To See the Fellows Fight: Eye-Witness Accounts of Meetings of the Geological Society of London and its Club, 1822–1868. BSHS Monographs, 12. London: British Society for the History of Science, 2003. Pp. xviii + 243. ISBN 0-906450-14-4. £15.00. $26.00 (paperback). British Journal for the History of Science, 38(4): 487–488, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028470.

Morrell:2005:JPB

[Mor05c] Jack Morrell. John Phillips and the business of Victorian science. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 1-84014-239-1. xix + 437 pp. LCCN QE22.P477 M67

2005. URL http://www.loc.
gov/catdir/toc/ecip0416/
2004006876.html.

Morus:2005:WPB

[Mor05d] Iwan Rhys Morus. When physics became king. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-54201-7, 0-226-54202-5 (paperback). xii + 303 pp. LCCN QC9.E89 M67

2005. URL http://www.loc.
gov/catdir/description/
uchi051/2004015207.html;
http://www.loc.gov/catdir/
enhancements/fy0616/
2004015207-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0419/2004015207.html.

Morrell:2006:BRB

[Mor06a] Jack Morrell. Book review:

Bernard Lightman (general ed.), Dictionary of Nineteenth-Century British Scientists. 4 vols. Bristol: Thoemmes Continuum, 2004. ISBN 1-85506-999-7. £650.00 (hardback). British Journal for the History of Science, 39(3):454–456, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028498.

Morrell:2006:BRN

[Mor06b] Jack Morrell. Book review: Noah Heringman, Romantic Rocks, Aesthetic Geology. Ithaca and London: Cornell University Press, 2004. Pp. xxii + 304. ISBN 0-8014-4127-7. £27.50, $47.50 (hardback). British Journal for the History of Science, 39(4): 614–615, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028535.

Morrell:2006:BRG

[Mor06c] Jack Morrell. Book reviews: Geological movements Sandra Herbert, Charles Darwin, Geologist. Ithaca and London: Cornell University Press, 2005. Pp. xx + 485. ISBN 0-8014-4348-2. £21.95, $39.95 (hardback). Martin J. S. Rudwick, The New Science of Geology: Studies in the Earth Sciences in the Age of Revolution. Aldershot: Ashgate, 2004. Pp. xviii+316. ISBN 0-86078-958-6. £60.00 (hardback). Martin J. S. Rudwick, Lyell and Darwin, Geologists: Studies in the Earth Sciences in the Age of Reform. Aldershot: Ashgate, 2005. Pp. xviii+316. ISBN 0-86078-959-4. £60.00 (hardback). Martin J. S. Rudwick, Bursting the Limits of Time: The Reconstruction of Geohistory in the Age of Revolution. Chicago and London: University of Chicago Press, 2005. Pp. xxiv+708. ISBN 0-226-73111-1. £28.50, $45.00 (hardback). British Journal for the History of Science, 39 (2):273–279, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morris:2006:BRL

[Mor06d] Peter Morris. Book review:

Lesley Richmond, Julie Stevenson and Alison Turton (eds.), The Pharmaceutical Industry: A Guide to Historical Records. Aldershot: Ashgate, 2003. Pp. ix + 561. ISBN 0-7546-3352-7. £55.00 (hardback). British Journal for the History of Science, 39 (1):144–145, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028569.

Morris:2006:BRP

[Mor06e] Peter Morris. Book review:

Peter J. Ramberg, Chemical Structure, Spatial Arrangement: The Early History of Stereochemistry, 1874–1914. Aldershot: Ashgate, 2003. Pp. xxiv + 399. ISBN 0-75546-0397-0. £57.50 (hardback). British Journal for the History of Science, 39(2):305–306, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028629.

Morrell:2007:BRJ

[Mor07a] Jack Morrell. Book review: A. J. Bowden, C. V. Burek and R. Wilding (eds.), History of Palaeobotany: Selected Essays. Geological Society Special Publication 241. London: The Geological Society, 2005. Pp.

304. ISBN 1-86239-174-2. £80.00 (hardback). British Journal for the History of Science, 40(3): 447–448, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500756.

Morrell:2007:BRMa

[Mor07b] Jack Morrell. Book review: Michael Collie and John Diemer (eds.), Murchison’s Wanderings in Russia: His Geological Exploration of Russia in Europe and the Ural Mountains, 1840 and 1841. British Geological Survey Occasional Publication No. 2. Keyworth: British Geological Survey, 2004. Pp. xvi + 474. ISBN: 0-85272-467-5. £40.00 (hardback). British Journal for the History of Science, 40 (1):140–141, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500702.

Morrell:2007:BRMb

[Mor07c] Jack Morrell. Book reviews:

Michael Freeman, Victorians and the Prehistoric: Tracks to a Lost World. New Haven and London: Yale University Press, 2004. Pp. x + 310. ISBN: 0-300-10334-4. £25.00 (hardback). Jan T. KozÁK, Victor S. Moreira and David R. Oldroyd, Iconography of the 1755 Lisbon Earthquake. Prague: Geophysical Institute of the Academy of Sciences of the Czech Republic and Academia, the Publisher of the Academy of Sciences of the Czech Republic, 2005. Pp. 84. ISBN: 80-239-4390-1 (GI), 80-200-1322-9 (Academia). No price given (hardback). British Journal for the History of Science, 40(2):295, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500732.

Morus:2008:BRE

[Mor08] Iwan Rhys Morus. Book review: Elizabeth Green Musselman, Nervous Conditions: Science and the Body Politic in Early Industrial Britain. Albany, NY: State University of New York Press, 2006. Pp. xi + 276. ISBN 0-7914-6679-5. $75.00 (hardback). British Journal for the History of Science, 41 (2):282–283, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morgan:2009:RCS

[Mor09a] John Morgan. Religious conventions and science in the early Restoration: Reformation and ‘Israel’ in Thomas Sprat’s history of the royal society (1667). British Journal for the History of Science, 42(3):321–344, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Morus:2009:BRF

[Mor09b] Iwan Rhys Morus. Book reviews: Frank A. J. L. James (ed.), The Correspondence of Michael Faraday: Volume 5, 1855–1860. London: Institution of Engineering and Technology, 2008. Pp. lviii + 835. ISBN 978-0-86341-823-5. £70.00 (hardback). Frank A. J. L. James (ed.), Christmas at the Royal Institution: An Anthology of Lectures by M. Faraday, J. Tyndall, R. S. Ball, S. P. Thompson, E. R. Lankester, W. H. Bragg, W. L. Bragg, R. L. Gregory, and I. Stewart. Singapore: World Scientific Books, 2007. Pp. xxxiii+366. ISBN 981-277-109-3. £39.00 (paperback). British Journal for the History of Science, 42(2):308– 310, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mosley:2007:BHT

[Mos07] Adam Mosley. Bearing the heavens: Tycho Brahe and the astronomical community of the late Sixteenth Century. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-83866-5. xiv + 354 pp. LCCN QB36.B8 M67

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0713/2007000713-d.html;
http://www.loc.gov/catdir/
enhancements/fy0729/
2007000713-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip078/2007000713.html.

Mosley:2008:BRM

[Mos08a] Adam Mosley. Book review:

Matthew McLean, The Cosmographia of Sebastian Münster: Describing the World in the Reformation. Aldershot: Ashgate, 2007. Pp. viii + 378. ISBN 978-0-7546-5843-6. £60.00 (hardback). British Journal for the History of Science, 41(4):603–605, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Mosley:2008:BRS

[Mos08b] Adam Mosley. Book reviews:

Stephen Clucas (ed.) John Dee: Interdisciplinary Studies in English Renaissance Thought. International Archives of the History of Ideas, 193. Dordrecht: Springer, 2006. Pp. xvii + 366. ISBN: 1-4020-4245-0. EUR144.00, $189.00, £111.00 (hardback). ISBN: 1-4020-4246-9 (e-book). British Journal for the History of Science, 41(1):132– 134, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Moyer:1997:JHR

[Moy97] Albert E. Moyer. Joseph Henry: the rise of an American scientist. Smithsonian Institution Press, Washington, DC, USA, 1997. ISBN 1-56098-776-6 (cloth). xii + 348 pp. LCCN QC16.H37 M69 1997.

Machamer:2000:SCP

[MPB00] Peter K. Machamer, Marcello Pera, and Aristeidåes Baltas, editors. Scientific controversies: philosophical and historical perspectives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-511987-8. x + 278 pp. LCCN Q175 .S42323

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0605/98038234-d.html.

Mirowski:2002:SBS

[MS02] Philip Mirowski and Esther-Mirjam Sent, editors. Science bought and sold: essays in the economics of science. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-53856-7, 0-226-53857-5 (paperback). ix + 573 pp. LCCN Q180.55.E25

                                        S33 2002. URL http://www.
                                        loc.gov/catdir/bios/
                                        uchi051/2001042486.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi051/
                                        2001042486.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy031/2001042486.html.

Malcolm:2005:JPH

[MS05a] Noel Malcolm and Jacqueline A.

Stedall. John Pell (1611–1685) and his correspondence with Sir Charles Cavendish: the mental world of an early mathematician. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856484-8 (hardcover). 657 pp. LCCN QA29.P6725 A4

2005. URL http://www.h-net.
org/review/hrev-a0e7r9-aa;
 http://www.loc.gov/catdir/
enhancements/fy0627/
2005297429-d.html;
http://www.loc.gov/catdir/
enhancements/fy0627/
2005297429-t.html.

Marsden:2005:EEC

[MS05b] Ben Marsden and Crosbie Smith.

Engineering empires: a cultural history of technology in Nineteenth-Century Britain. Palgrave Macmillan, Houndmills, Basingstoke, Hampshire England and New York, NY, USA, 2005. ISBN 0-333-77278-4. xi + 351 pp. LCCN T14.5 .M37 2005. URL http:/

/www.loc.gov/catdir/bios/
hol053/2004052595.html;
http://www.loc.gov/catdir/
description/hol053/
2004052595.html;
http://www.loc.gov/catdir/
enhancements/fy0618/
2004052595-t.html.

Macleod:2007:ESR

[MT07] Christine Macleod and Jennifer Tann. From engineer to scientist: reinventing invention in the Watt and Faraday centenaries, 1919–1931. British Journal for the History of Science, 40(3):389–411, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500749.

Mudgway:2001:UDH

[Mud01] Douglas J. Mudgway. Uplinkdownlink: a history of the NASA Deep Space Network, 1957–1997, volume 2001–4227 of NASA SP; The NASA history series. NASA, Washington, DC, USA, 2001. ISBN 0-16-066599-X. xlviii + 674 pp. LCCN TL3026 .M84 2001.

Muller:2001:BRB

[Mül01] Falk Müller. Book review: B. Finn (ed.), with Robert Bud and Helmuth Trischler (associate eds.), Exposing Electronics. Artefacts: Studies in the History of Science and Technology, 2. Amsteldijk: Harwood Academic Publishers, 2000. Pp. xiv + 199. ISBN 90-5823-057-0. £15.00, $23.00 (paperback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028390.

Munz:2007:BRL

[Mun07] Tania Munz. Book review: Lorraine Daston and Gregg Mitman (eds.), Thinking with Animals: New Perspectives on Anthropomorphism. New York: Columbia University Press, 2005. Pp. vii + 230. ISBN 0-231-13038-4. £32.00, $49.50 (hardback). British Journal for the History of Science, 40(2):277–279, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500721.

Munz:2008:BRJ

[Mun08] Tania Munz. Book reviews:

Jonathan Burt, Rat. Animal Series. London: Reaktion Books, 2006. Pp. 189. ISBN 1-86189-224-1. £12.95, $19.95 (paperback). Helen Macdonald, Falcon. Animal Series. London: Reaktion Books, 2006. Pp. 208. ISBN 1-86189-238-1. £12.95, $19.95 (paperback). Claire Preston, Bee. Animal Series. London: Reaktion Books, 2006. Pp. 206. ISBN 1-86189-256-X. £12.95, $19.95 (paperback). British Journal for the History of Science, 41(3): 445–447, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Munz:2009:BRD

[Mun09] Tania Munz. Book review:

Deborah Coen, Vienna in the Age of Uncertainty: Science, Liberalism, and Private Life. Chicago and London: University of Chicago Press, 2007. Pp. xi + 380. ISBN 978-0-226-11172-8. $45.00, £28.50 (hardback). British Journal for the History of Science, 42(4):624–626, December 2009. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Murray:1999:EAG

[Mur99] Tim Murray, editor. Encyclopedia of archaeology: The great archaeologists. ABC-CLIO, Santa Barbara, CA, USA, 1999. ISBN 1-57607-199-5 (set), 1-57607-577-X (e-book). xxii + 459 (vol. 1), 461–950 (vol. 2) pp. LCCN CC110 .E54 1999.

Murray:2000:WBM

[Mur00] Margaret Anne Marie Murray. Women becoming mathematicians: creating a professional identity in post-World War II America. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-13369-5 (hardcover). xviii + 277 pp. LCCN QA27.5 .M88 2000.

Murdoch:2002:BRI

[Mur02] Jill Murdoch. Book review: Ian Carter, Railways and Culture in Britain: The Epitome of Modernity. Studies in Popular Culture. Manchester and New York: Manchester University Press, 2001. Pp. xiv + 338. ISBN 0-7190-5966-6. £16.99 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Musselman:2003:BRR

[Mus03]  Elizabeth  Green  Mussel-

man. Book review: Roger Luckhurst and Josephine McDonagh (eds.), Transactions and Encounters: Science and Culture in the Nineteenth Century. Manchester and New York: Manchester University Press, 2002. Pp. ix + 239. ISBN 0-7190-5911-9. £15.99 (paperback). British Journal for the History of Science, 36(3): 371–372, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028164.

Musselman:2006:NCS

[Mus06] Elizabeth Green Musselman.

Nervous conditions: science and the body politic in early industrial Britain. SUNY series, studies in the long nineteenth century; SUNY series in science, technology, and society. State University of New York Press, Albany, NY, USA, 2006. ISBN 0-7914-6679-5 (hardcover). xi + 276 pp. LCCN RC464.A1 G74 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0512/2005014025.html.

Mussell:2009:ACR

[Mus09] James Mussell. Arthur Cowper Ranyard, knowledge and the reproduction of astronomical photographs in the late nineteenthcentury periodical press. British Journal for the History of Science, 42(3):345–380, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McCartney:2003:PIP

[MW03] Mark McCartney and Andrew Whitaker, editors. Physicists of Ireland: passion and precision. Institute of Physics Pub., Bristol, UK and Philadelphia, PA, USA, 2003. ISBN 0-7503-0866-4 (hardcover). xiv + 298 pp. LCCN QC15 .P47

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0806/2007298520-d.html;
http://www.loc.gov/catdir/
toc/fy0802/2007298520.html.

Muller-Wille:2007:BRJ

[MW07] Staffan Müller-Wille. Book review: Justin E. H. Smith (ed.), The Problem of Animal Generation in Early Modern Philosophy. Cambridge Studies in Philosophy and Biology. Cambridge: Cambridge University Press, 2006. Pp. xiii + 456. ISBN 978-0-521-84077-4. £45.00 (hardback). British Journal for the History of Science, 40(4): 605–606, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Marshall:2009:BRS

[MW09a] Trevor Marshall and Max Wallis. Book review: Silvan S. Schweber, Einstein and Oppenheimer: The Meaning of Genius. London: Harvard University Press, 2008. Pp. xiv + 412. ISBN 978-0-674-02828-9. £19.25 (hardback). British Journal for the History of Science, 42(3):471–473, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Muller-Wille:2009:BRM

[MW09b] Staffan Müller-Wille. Book review: Marianne Sommer, Bones and Ochre: The Curious Afterlife of the Red Lady of Paviland. Cambridge, MA and London: Harvard University Press, 2007. Pp. xii + 398. ISBN 978-0-674-02499-1. £25.95 (hardback). British Journal for the History of Science, 42(4):619–620, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nair:2006:SPC

[Nai06] Savithri Preetha Nair. Science and the politics of colonial collecting: the case of Indian meteorites, 1856–70. British Journal for the History of Science, 39(1):97–119, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Naylor:2003:GCO

[Nay03] Ron Naylor. Galileo, Copernicanism and the origins of the new science of motion. British Journal for the History of Science, 36(2):151–181, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028231.

Naylor:2005:IHG

[Nay05] Simon Naylor. Introduction: historical geographies of science — places, contexts, cartographies. British Journal for the History of Science, 38(1):1–12, March 2005.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Naylor:2006:NPW

[Nay06] Simon Naylor. Nationalizing provincial weather: meteorology in nineteenth-century Cornwall. British Journal for the History of Science, 39(3):407–433, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Naylor:2009:BRJ

[Nay09] Simon Naylor. Book reviews: Jan Golinski, British Weather and the Climate of Enlightenment. Chicago and London: University of Chicago Press, 2007. Pp. 272. ISBN 978-0-226-30205-8. $35.00 (hardback). James Rodger Fleming, Vladimir Jankovic and Deborah R. Coen (eds.), Intimate Universality: Local and Global Themes in the History of Weather and Climate. Sagamore Beach: Science History Publications/USA, 2006. Pp. xx + 264. ISBN 0-88135-367-1. $39.95 (hardback). British Journal for the History of Science, 42(1):138–140, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nyhart:2002:SCS

[NB02]  Lynn  K.

Nyhart and Thomas Hoyt Broman, editors. Science and civil society, volume 17 of Osiris: a research journal devoted to the history of science and its cultural influences. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-07371-8, 0-226-07372-6. x + 373 pp. LCCN Q125 .S43178 2002; Q1 .O7 ser.2, v.17

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0666/2006367560-d.html;
http://www.loc.gov/catdir/
enhancements/fy0666/
2006367560-t.html. Second se-

ries.

Newton:1999:PMP

[NCW99] Sir Isaac Newton, I. Bernard Cohen, and Anne Miller Whitman. The Principia: mathematical principles of natural philosophy. University of California Press, Berkeley, CA, USA, 1999. ISBN 0-520-08816-6, 0-520-08817-4. xvii + 966 pp. LCCN QA803 .N413 1999. URL

http://www.loc.gov/catdir/
bios/ucal052/99010278.html;
 http://www.loc.gov/catdir/
description/ucal041/
99010278.html.

Neal:2000:BRF

[Nea00a] Katherine Neal. Book review:

Frances Willmoth (ed.), Flamsteed’s Stars: New Perspectives on the Life and Work of the first Astronomer Royal, 1646–1719. Woodbridge, Suffolk: The Boydell Press in association with the National Maritime Museum, 1997. Pp. xiv + 271. ISBN 0-85115-706-8. £45.00. British Journal for the History of Science, 33(1):

115–124, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028071.

Neal:2000:BRE

[Nea00b] Katherine Neal. Book reviews: Eric G. Forbes, Lesley Murdin and Frances Willmoth (eds.), The Correspondence of John Flamsteed, the First Astronomer Royal. Vol. 1. 1666–1682. Bristol: Institute of Physics, 1995. Pp. xlix + 955. ISBN 0-7503-01473-3. £140.00. $280.00. Eric G. Forbes, Lesley Murdin and Frances Willmoth (eds.), The Correspondence of John Flamsteed, the First Astronomer Royal. Vol. 2. 1682–1703. Bristol: Institute of Physics, 1997. Pp. xxxvii+1095. ISBN 0-7503-0391-3. £140.00, $280.00. British Journal for the History of Science, 33(1): 115–124, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028070.

Nerone:2005:BRL

[Ner05] John Nerone. Book review: Lisa Gitelman and Geoffrey B. Pingree (eds.), New Media, 1740–1915. Media in Transition. Cambridge, MA and London: MIT Press, 2003. Pp. xxxiv + 271. ISBN 0-262-07245-9. £23.50 (hardback). British Journal for the History of Science, 38(4): 473–474, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028459.

Neswald:2001:BRC

[Nes01] Elizabeth Neswald. Book review: Constantin Goschler (ed.), Wissenschaft und Öffentlichkeit in Berlin, 1870–1930. Stuttgart: Franz Steiner Verlag, 2000. Pp.

232. ISBN 3-515-07778-2. DM 86.00 (paperback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028391.

Neswald:2002:BRP

[Nes02] Elizabeth Neswald. Book review:

Peter D. Smith, Metaphor and Materiality. German Literature and the World-View of Science 1780–1955. Legenda/Studies in Comparative Literature 4. Oxford: European Humanities Research Centre, University of Oxford, 2000. Pp. xii + 372. ISBN 1-900755-32-7. £27.50, $49.50 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Neswald:2006:BRC

[Nes06a] Elizabeth Neswald. Book review: Carsten Kretschmann (ed.), Wissenspopularisierung. Konzepte der Wissensverbreitung im Wandel. Wissenskultur und gesellschaftlicher Wandel 4. Berlin: Akademie Verlag, 2003. Pp. 409. ISBN 3-05-003770-9. EUR49.80 (hardback). British Journal for the History of Science, 39(2):284–285, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028612.

Neswald:2006:BRM

[Nes06b] Elizabeth Neswald. Book review: Michael Hau, The Cult of Health and Beauty in Germany: A Social History, 1890–1930. Chicago and London: University of Chicago Press, 2003. Pp. 272. ISBN 0-226-31976-8. £15.50, $22.00 (paperback). British Journal for the History of Science, 39(3): 466–467, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028505.

Neswald:2006:BRO

[Nes06c] Elizabeth Neswald. Book review: Oliver Hochadel, Öffentliche Wissenschaft. Elektrizität in der deutschen Aufklärung. Göttingen: Wallstein Verlag, 2003. Pp. 364. ISBN 3-89244-629-6. EUR 35.00 (paperback). British Journal for the History of Science, 39(1): 133–134, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028560.

Neswald:2006:SSI

[Nes06d] Elizabeth Neswald. Science, sociability and the improvement of Ireland: the Galway Mechanics’ Institute, 1826–1851. British Journal for the History of Science, 39(4):503–534, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Neswald:2007:BRF

[Nes07] Elizabeth Neswald. Book review:

Falk Müller, Gasentladungsforschung im 19. Jahrhundert. Diepholz: GNT-Verlag, 2004. Pp. 300. ISBN 3-928186-76-0. EUR35.00 (paperback). British Journal for the History of Science, 40(1): 143–144, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500704.

Neswald:2009:BRW

[Nes09] Elizabeth R. Neswald. Book review: William H. Brock, William Crookes (1832–1919) and the Commercialization of Science. Alderhot: Ashgate, 2008. Pp. xxi + 556. ISBN 978-0-7546-6322-5. £65.00 (hardback). British Journal for the History of Science, 42(3):465–466, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Newman:2003:GFL

[New03] William Royall Newman. Gehennical fire: the lives of George Starkey, an American alchemist in the scientific revolution. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-57714-7 (paperback). xxiv + 348 pp. LCCN QD24.S73 N49

2003. URL http://www.loc.
gov/catdir/bios/uchi051/
2002032350.html;
http://www.loc.gov/catdir/
description/uchi052/
2002032350.html;   http://
www.loc.gov/catdir/toc/
uchi051/2002032350.html.

NG:1997:SAW

[NFJ97] National Gallery (Great Britain), J. V. Field, and Frank A. J. L. James. Science in art: works in the National Gallery that illustrate the history of science and technology, volume 11 of BSHS monographs. British Society for the History of Science, Stanford in the Vale, England, 1997. ISBN 0-906450-13-6. 110 pp. LCCN ND1460.S34 N38 1997.

Newman:2001:SNA

[NG01a] William Royall Newman and Anthony Grafton, editors. Secrets of nature: astrology and alchemy in early modern Europe. Transformations. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-14075-6. 443 pp. LCCN BF1676 .S43 2001.

Nieto-Galan:2001:BRL

[NG01b] Agustí Nieto-Galan. Book review: Liliane Hilaire-Pérez, L’Invention technique au siècle des lumières. Paris: Albin Michel, 2000. Pp. 443. ISBN 2-226-11537-4. FF 140.00. British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028053.

Nieto-Galan:2001:CTH

[NG01c] Agustí Nieto-Galan. Colouring textiles: a history of natural dyestuffs in industrial Europe, volume 217 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7022-8, 94-017-1081-3. ISSN 0068-0346. xxv + 246 pp. LCCN TP919 .N54

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0813/2001033541-d.html;
http://www.loc.gov/catdir/
enhancements/fy0813/
2001033541-t.html;   https:
//link.springer.com/book/
10.1007/978-94-017-1081-7.

Nieto-Galan:2001:SMV

[NG01d] Agustí Nieto-Galan. La seducción de la máquina: vapores, submarinos e inventores: Santponç, Monturiol, Isaac Peral. (Spanish) [The seduction of the machine: Steam, submarines and inventors Santponç, Monturiol, Isaac Peral], volume 2 of Novatores. Nivola Libros y Ediciones, Madrid, Spain, 1. edition, 2001. ISBN 84-95599-10-4. 129 pp. LCCN TJ139 .N54 2001.

Nieto-Galan:2004:FRE

[NG04] Agustí Nieto-Galan. Free radicals in the European periphery: ‘translating’ organic chemistry from Zurich to Barcelona in the early twentieth century. British Journal for the History of Science, 37(2):167–191, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028329.

NMM:2002:SGC

[NHA^{+}02] National Maritime Museum (Great Britain), Hester Higton, Silke Ackermann, et al. Sundials at Greenwich: a catalogue of the sundials, nocturnals, and horary quadrants in the National Maritime Museum, Greenwich. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850877-8. x pp. LCCN QB215 .N37

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002066794-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0613/
                                        2002066794-t.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0725/
                                        2002066794-b.html.

North:2000:NMA

[NM00] Roger North and Peter Millard. Notes of me: the autobiography of Roger North. University of Toronto Press, Toronto, ON, Canada, 2000. ISBN 0-8020-4471-9 (bound). xix + 353 pp. LCCN DA437.N87 A3 2000.

Noakes:2004:BRR

[Noa04] Richard Noakes. Book review:

Roger Luckhurst, The Invention of Telepathy, 1870–1901. Oxford: Oxford University Press, 2002. Pp. ix + 324. ISBN 0-19-924962-8. £35.00 (hardback). British Journal for the History of Science, 37(1):110–112, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028267.

Noakes:2009:BRE

[Noa09] Richard Noakes. Book review:

The English Mechanic and World of Science. Vols. 1-124 on 14 DVDs. Waltham Abbey: E. S. Hutton, 2006. ISBN 1-905383-00-2. Single disks £22.00 (individual), £50.00 (institutional); complete set £308.00 (individual), £650.00 (institutional). British Journal for the History of Science, 42 (2):310–311, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Noltie:2007:RWB

[Nol07] Henry J. Noltie. Robert Wight and the botanical drawings of Rungiah and Govindoo. Royal Botanic Garden, Edinburgh, Scotland, 2007. ISBN 1-906129-02-9 (set), 1-906129-03-7 (Book 1), 1-906129-00-2 (Book 2), 1-906129-01-0 (Book 3). 215 (vol. 1), 208 (vol. 2), 88 (vol. 3) pp. LCCN QK31.W48 N66 2007.

North:2005:GCR

[Nor05] John David North. God’s clockmaker: Richard of Wallingford and the invention of time. Hambledon and London, London, UK, 2005. ISBN 1-85285-451-0. xvii + 441 pp. LCCN TS544.8.W35 N67

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0629/2006361208-b.html;
http://www.loc.gov/catdir/
enhancements/fy0629/
2006361208-d.html;
http://www.loc.gov/catdir/
enhancements/fy0629/
2006361208-t.html.

Nordlund:2007:EEA

[Nor07] Christer Nordlund. Endocrinology and expectations in 1930s America: Louis Berman’s ideas on new creations in human beings. British Journal for the History of Science, 40(1): 83–104, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500683.

Newman:2002:ATF

[NP02] William Royall Newman and Lawrence Principe. Alchemy tried in the fire: Starkey, Boyle, and the fate of Helmontian chymistry. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-57711-2 (cloth). xiv + 344 pp. LCCN QD18 .G7

N48 2002. URL http://www.
loc.gov/catdir/bios/
uchi051/2002012922.html;

http://www.loc.gov/catdir/
description/uchi052/
2002012922.html;
http://www.loc.gov/catdir/
toc/fy034/2002012922.html.

Numbers:2000:AAP

[NS00] Ronald L. Numbers and John Stenhouse. Antievolutionism in the Antipodes: from protesting evolution to promoting creationism in New Zealand. British Journal for the History of Science, 33(3):335–350, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027956.

Neeley:2001:MSS

[NS01] Kathryn A. (Kathryn Angelyn)

Neeley and Mary Somerville. Mary Somerville: science, illumination, and the female mind. Cambridge science biographies. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-62299-9, 0-521-62672-2 (paperback). xvi + 263 pp. LCCN Q143.S7 N44

2001. URL http://www.loc.
gov/catdir/description/
cam021/00049363.html;
http://www.loc.gov/catdir/
toc/cam025/00049363.html.

Nummedal:2007:AAH

[Num07] Tara E. Nummedal. Alchemy and authority in the Holy Roman Empire. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-60856-5 (cloth). xvii + 260 pp. LCCN QD13 .N86

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007002521-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007002521-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip079/2007002521.html.

Nutton:2004:BRG

[Nut04] Vivian Nutton. Book review: G. E. R. Lloyd, The Ambitions of Curiosity: Understanding the World in Ancient Greece and China. Ideas in Context Series, 64. Cambridge: Cambridge University Press, 2002. Pp. xvi + 175. ISBN 0-521-89461-1. £15.95, $22.00 (paperback). British Journal for the History of Science, 37(3): 344, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028427.

Nutton:2005:BRN

[Nut05] Vivian Nutton. Book review: Nancy G. Siraisi, Medicine and the Italian Universities, 1250– 1600. Education and Society in the Middle Ages and Renaissance, 12. Leiden: Brill, 2001. Pp. ix + 389. ISBN 90-04-11942-6. $122.00, EUR105.00 (hardback). British Journal for the History of Science, 38(3): 360, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028681.

Needham:1954:SCC

[NW08]  Joseph  Need-

ham and Ling Wang. Science and civilisation in China. University Press, Cambridge, UK, 1954–2008. ISBN 0-521-05799-X (vol. 1), 0-521-05800-7 (vol. 2), 0-521-05801-5 (vol. 3), 0-521-07060-0 (vol. 4, pt 3), 0-521-32727-X (vol. 5, pt 6), (vol. 5, pt. 11), 0-521-65270-7 (vol. 6, pt 5), 0-521-63262-5 (vol. 6, pt 6), 0-521-57143-X (vol. 7, pt 1). xxxviii + 319 (vol. 1), xxii + 697 (vol. 2), xlvii + 886 (vol. 3), lvii + 931 (vol. 4, pt 3), xxviii + 601 (vol. 5, pt 6), xxviii + 741 (vol. 6, pt 5), xviii + 261 (vol. 6, pt 6), xxiv + 479 (vol. 7, pt 1) pp. LCCN DS721 .N39. URL http:

//www.loc.gov/catdir/
description/cam022/
54004723.html;   http://
www.loc.gov/catdir/samples/
cam031/54004723.html;
http://www.loc.gov/catdir/
toc/cam022/54004723.html.

Nye:2004:BPW

[Nye04] Mary Jo Nye. Blackett: physics, war, and politics in the Twentieth Century. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01548-7 (hardcover). x + 255 pp. LCCN QC16.B59 N94 2004.

Nye:2006:TMQ

[Nye06] David E. Nye. Technology matters: questions to live with. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-14093-4. xiv + 282 pp. LCCN T14 .N88 2006.

Nye:2009:BRT

[Nye09] Mary Jo Nye. Book review: Thomas Söderqvist (ed.) The History and Poetics of Scientific Biography. Aldershot: Ashgate, 2007. Pp. xv + 270. ISBN 978-0-7546-5181-9. £55.00 (hardback). British Journal for the History of Science, 42(1):117– 119, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Archimedes:2004:WAT

[oAEN04] Archimedes of Ascalon Eutocius and Reviel Netz. The works of Archimedes: translated into English, together with Eutocius’ commentaries, with commentary, and critical edition of the diagrams. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-66160-9. x + 375 pp. LCCN QA31 .A692

                                        2004. URL http://www.loc.
                                        gov/catdir/description/
                                        cam051/2004301152.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0732/
                                        2004301152-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam051/2004301152.html.

OConnor:2005:CWC

[O’C05] Anne O’Connor. The competition for the Woodwardian Chair of Geology: Cambridge, 1873. British Journal for the History of Science, 38(4):437–461, December 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

OConnor:2007:ESF

[O’C07] Ralph O’Connor. The Earth on show: fossils and the poetics of popular science, 1802– 1856. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-61668-1. xiii + 541 + 6 pp. LCCN QE13.G7 O35

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0732/2007028061-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007028061-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0722/2007028061.html.

Ogilvie:2000:OAA

[Ogi00] Marilyn Bailey Ogilvie. Obligatory amateurs: Annie Maunder (1868–1947) and British women astronomers at the dawn of professional astronomy. British Journal for the History of Science, 33(1):67–84, March 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028066.

Olby:2002:BRP

[Olb02] Robert Olby. Book review: Peter J. Beurton, Raphael Falk and Hans-Jörg Rheinberger (eds.), The Concept of the Gene in Development and Evolution: Historical and Epistemological Perspectives. Cambridge: Cambridge University Press, 2000. Pp. xvi + 384. ISBN 0-521-77187-0. $59.95 (hardback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028199.

Olson:2008:SSN

[Ols08] Richard Olson. Science and scientism in Nineteenth-Century Europe. University of Illinois Press, Urbana, IL, USA, 2008. ISBN 0-252-03188-1 (cloth), 0-252-07433-5 (paperback). 349 pp. LCCN Q175.52.E85 O47

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0827/2007005146-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/ecip0710/2007005146.
                                        html.

Oreskes:2001:PTI

[Ore01] Noami Oreskes. Plate tectonics: an insider’s history of the modern theory of the Earth. Westview Press, Boulder, CO, USA, 2001. ISBN 0-8133-3981-2. xxiv + 424 pp. LCCN ????

Oslund:2002:IIN

[Osl02] Karen Oslund. Imagining Iceland: narratives of nature and history in the North Atlantic. British Journal for the History of Science, 35(3):313–334, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028126.

Oster:2002:SEP

[Ost02a] Malcolm Oster, editor. Science in Europe, 1500–1800: a primary sources reader. Palgrave, Houndmills, Basingstoke, Hampshire, 2002. ISBN 0-333-97001-2, 0-333-97002-0 (paperback). xiv + 282 pp. LCCN Q127.E8 S352

2002. URL http://www.loc.
gov/catdir/bios/hol058/
2001036833.html;
http://www.loc.gov/catdir/
description/hol052/
2001036833.html;
http://www.loc.gov/catdir/
toc/hol051/2001036833.html.

Oster:2002:SES

[Ost02b] Malcolm Oster, editor. Science in Europe, 1500–1800: a secondary sources reader. Palgrave, Houndmills, Basingstoke, Hampshire, 2002. ISBN 0-333-97005-5 (cloth), 0-333-97006-3 (paperback). xii + 307 pp. LCCN Q127.E8 S354 2002. URL http:

//www.loc.gov/catdir/bios/
hol058/2001045149.html;
http://www.loc.gov/catdir/
description/hol052/
2001045149.html;
http://www.loc.gov/catdir/
toc/fy02/2001045149.html.

Otter:2008:VEP

[Ott08] Chris Otter. The Victorian eye: a political history of light and vision in Britain, 1800– 1910. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-64076-0 (cloth), 0-226-64077-9 (paperback). x + 382 pp. LCCN TH7703 .O88

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0805/2007043184-d.html;

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0827/
                                        2007043184-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip082/2007043184.html.

Outram:2001:BRJ

[Out01] Dorinda Outram. Book review: Jonathan I. Israel, Radical Enlightenment: Philosophy and the Making of Modernity 1650–1750. Oxford: Oxford University Press, 2001. Pp. xix + 810. ISBN 0-19-820608-9. £30.00 (hardback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028392.

Outram:2004:BRJ

[Out04] Dorinda Outram. Book review: Jessica Riskin, Science in the Age of Sensibility: The Sentimental Empiricists of the French Enlightenment. Chicago and London: University of Chicago Press, 2002. Pp. xiii + 338. ISBN 0-226-72079-9. £17.70, $25.00 (paperback). British Journal for the History of Science, 37(2):210–211, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028340.

Pacey:2003:EMJ

[Pac03] A. J. Pacey. Emerging from the museum: Joseph Dawson, mineralogist, 1740–1813. British Journal for the History of Science, 36(4):455–469, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028216.

Padian:2001:BRD

[Pad01] Kevin Padian. Book review:

Deborah Cadbury, The Dinosaur Hunters: A True Story of Scientific Rivalry and the Discovery of the Prehistoric World. London: Fourth Estate, 2000. Pp. x + 374. ISBN 1-85702-959-3. £15.99. British Journal for the History of Science, 34 (1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028054.

Page:2001:BRB

[Pag01] Sophie Page. Book review:

Barbara Howard Traister, The Notorious Astrological Physician of London: Works and Days of Simon Forman. Chicago and London: University of Chicago Press, 2001. Pp. xviii + 250. ISBN 0-226-81140-9. £19.00, $30.00 (hardback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028393.

Page:2003:BRW

[Pag03] Sophie Page. Book review:

William R. Newman and Anthony Grafton (eds.), Secrets of Nature: Astrology and Alchemy in Early Modern Europe. Transformations: Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2001. Pp.

443. ISBN 0-262-14075-6. £34.50 (hardcover). British Journal for the History of Science, 36 (1):87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028320.

Palladino:2000:BRP

[Pal00]  Paolo Palladino.  Book re-

view: Paul Thagard. How Scientists Explain Disease. Princeton: Princeton University Press, 1999. Pp. xviii + 263. ISBN 0-691-00261-4, £18.95 (hardback). British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027938.

Palladino:2002:BRE

[Pal02] Paolo Palladino. Book review: Evelyn Fox Keller, The Century of the Gene. Cambridge, MA and London: Harvard University Press, 2000. Pp. 186. ISBN 0-674-00372-1. £15.95 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028200.

Palladino:2006:BRF

[Pal06] Paolo Palladino. Book review:

Finn Bowring, Science, Seeds and Cyborgs: Biotechnology and the Appropriation of Life. London and New York: Verso, 2003. Pp. xiii + 388. ISBN 1-85984-687-4. £19.00, $29.99, C$39.00 (hardback). British Journal for the History of Science, 39(1): 150–151, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028574.

Palmieri:2009:PGE

[Pal09] Paolo Palmieri. A phenomenology of Galileo’s experiments with pendulums. British Journal for the History of Science, 42(4): 479–513, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pang:2002:ESV

[Pan02] Alex Soojung-Kim Pang. Empire and the sun: Victorian solar eclipse expeditions. Writing science. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-3925-0, 0-8047-3926-9 (paperback). x + 196 + 12 pp. LCCN QB541 .P36

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0710/2001057661-b.html;
http://www.loc.gov/catdir/
enhancements/fy0710/
2001057661-d.html;
http://www.loc.gov/catdir/
toc/fy032/2001057661.html.

Pantalony:2006:BRM

[Pan06] David Pantalony. Book review: Marco Beretta, Paolo Galluzzi and Carlo Triarco (eds.), Musa Musaei: Studies on Scientific Instruments and Collections in Honour of Mara Miniati. Biblioteca di Nuncius Studi e Testi, XLIX. Firenze: Leo S. Olschki, 2003. Pp. viii + 486. ISBN 88-222-5238-1. EUR49.00 (paperback). British Journal for the History of Science, 39 (1):127–128, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028555.

Pantalony:2008:BRM

[Pan08] David Pantalony. Book review: Myles W. Jackson, Harmonious Triads: Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany. Cambridge, MA: MIT Press, 2006. Pp. x + 368. ISBN 0-262-10116-5. £25.95 (hardback). British Journal for the History of Science, 41(2):309–310, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Papanelopoulou:2006:GAH

[Pap06]  Faidra  Papanelopoulou.

Gustave-Adolphe Hirn (1815–1890): engineering thermodynamics in mid-nineteenth-century France. British Journal for the History of Science, 39(2):231–254, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Park:2001:BRA

[Par01a] Buhm Soon Park. Book review:

Anders Lundgren and Bernadette Bensaude-Vincent (eds.), Communicating Chemistry: Textbooks and their Audiences, 1789–1939. Canton, MA: Science History Publications, 2000. Pp. vii + 465. ISBN 0-88135-274-8. $56.00. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028020.

Parnes:2001:BRC

[Par01b]  Ohad Parnes.  Book

review: Christoph Meinel (ed.), Instrument–Experiment: historische Studien. Berlin: GNT-Verlag für Geschichte der Naturwissenschaften und der Technik, 2000. Pp. 423. ISBN 3-928186-51-5. £34.00. British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028116.

Parshall:2006:JJS

[Par06] Karen Hunger Parshall. James Joseph Sylvester: Jewish mathematician in a Victorian world. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8291-5. xiii + 461 pp. LCCN QA29.S96 P375

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0625/2005052035-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0625/
                                        2005052035-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy0701/2005052035.html.

Patiniotis:2007:PRE

[Pat07] Manolis Patiniotis. Periphery reassessed: Eugenios Voulgaris converses with Isaac Newton. British Journal for the History of Science, 40(4):471–490, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Paul:2009:BRJ

[Pau09] Diane B. Paul. Book review: Jene M. Porter and Peter W. B. Phillips (eds.), Public Science in Liberal Democracy. Toronto, Buffalo and London: University of Toronto Press, 2007. Pp. xii + 343. ISBN 978-0-8020-9359-2. £45.00 (hardback). British Journal for the History of Science, 42 (1):150–152, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pemberton:2004:BRE

[Pem04] Neil Pemberton. Book review: Emily Thompson, The Soundscape of Modernity: Architectural Acoustics and the Culture of Listening in America, 1900–1933. Cambridge, MA and London: MIT Press, 2002. Pp. ix + 500. ISBN 0-262-20138-0.

£30.95 (hardback). British Journal for the History of Science, 37 (1):113–114, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028269.

Pemberton:2007:BRH

[Pem07] Neil Pemberton. Book review:

Harry Collins and Trevor Pinch, Dr. Golem: How to Think about Medicine. Chicago and London: University of Chicago Press, 2005. Pp. 280. ISBN 0-226-11366-3. $25.00, £17.50 (hardback). British Journal for the History of Science, 40(3): 464–465, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500766.

Pennington:2004:BRD

[Pen04] Hugh Pennington. Book review: David Tyrrell and Michael Fielder, Cold Wars: The Fight Against the Common Cold. Oxford: Oxford University Press, 2002. Pp. xiv + 253. ISBN 0-19-263285-X. £17.99 (hardback). British Journal for the History of Science, 37(3):362–363, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028142;   http://www.
jstor.org/stable/4028444.

Pender:2006:EEU

[Pen06] Stephen Pender. Examples and experience: on the uncertainty of medicine. British Journal for the History of Science, 39(1):1– 28, March 2006. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pettitt:2004:PII

[Pet04] Clare Pettitt. Patent inventions — intellectual property and the Victorian novel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-925320-X. ix + 341 pp. LCCN PR878.I65 P47

                                        2004. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0618/2004049211-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0724/
                                        2004049211-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy045/2004049211.html.

Petrunic:2007:BRI

[Pet07a] Josipa G. Petrunić. Book review: Ivor Grattan-Guinness (ed.), Landmark Writings in Western Mathematics 1640–1940. Amsterdam, San Diego, Oxford and London: Elsevier B. V., 2005. Pp. xvii + 1022. ISBN 0-444-50871-6. £158.00 (hardback). British Journal for the History of Science, 40(4):608–610, December 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Petrunic:2007:BRT

[Pet07b] Josipa G. Petrunic. Book review: Thomas Ryckman, The Reign of Relativity: Philosophy in Physics, 1915–1925. Oxford: Oxford University Press, 2005. Pp. ix + 317. ISBN 0-19-517717-7. £36.50, $60.00 (hardback). British Journal for the History of Science, 40(2):306–307, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500740.

Pettit:2008:BRS

[Pet08] Michael Pettit. Book review:

Sarah E. Igo, The Averaged American: Surveys, Citizens, and the Making of a Mass Public. Cambridge, MA and London: Harvard University Press, 2007. Pp. 398. ISBN 978-0-674-02321-5. £22.95 (hardback). British Journal for the History of Science, 41(4):621–623, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Petrunic:2009:BRM

[Pet09] Josipa G. Petrunić. Book review: Massimo Mazzotti, The World of Maria Gaetana Agnesi, Mathematician of God. Baltimore: Johns Hopkins University Press, 2007. Pp. xx + 217. ISBN 978-0-8018-8709-3. £33.50 (hardback). British Journal for the History of Science, 42(1):119– 121, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Peyton:2001:SZS

[Pey01] John Peyton. Solly Zuckerman: a scientist out of the ordinary. John Murray, London, UK, 2001. ISBN 0-7195-6283-X. xviii + 252 pp. LCCN QL31.Z93 P48

                                        2001. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0739/2001430809-d.html.

Philippidis:2002:BRL

[Phi02] Chris Philippidis. Book review: Leah Ceccarelli, Shaping Science with Rhetoric: The Cases of Dobzhansky, Schrödinger, and Wilson. Chicago and London: University of Chicago Press, 2001. Pp. xi + 204. ISBN 0-226-09907-5. £13.00, $20.00 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pickstone:2001:WKN

[Pic01]  John  V.

Pickstone. Ways of knowing: a new history of science, technology, and medicine. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-66794-4 (cloth), 0-226-66795-2 (paperback). xii + 271 pp. LCCN Q125 .P595 2001. URL http://www.

                                        loc.gov/catdir/bios/
                                        uchi051/00053248.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi051/
                                        00053248.html;
                                        http://www.loc.gov/catdir/
                                        toc/uchi051/00053248.html.

Pickstone:2007:BRM

[Pic07] John Pickstone. Book review:

Margaret Pelling and Scott Mandelbrote (eds.), The Practice of Reform in Health, Medicine, and Science, 1500–2000: Essays for Charles Webster. Aldershot: Ashgate, 2005. Pp. xv + 376. ISBN 0-7546-3933-9. £60.00 (hardback). British Journal for the History of Science, 40(3): 465–466, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500767.

Pickstone:2009:DR

[Pic09] John V. Pickstone. The disunities of representation. British Journal for the History of Science, 42(4):595–600, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pier:2000:DM

[Pie00] Jean-Paul Pier, editor. Development of mathematics 1950– 2000. Birkhäuser Verlag, Basel, Switzerland, 2000. ISBN 3-7643-6280-4. ix + 1372 pp. LCCN QA26 .D52 2000.

Plofker:2005:BRH

[Plo05]  Kim  Plofker.

Book review: Helaine Selin (ed.) and Ubiratan D’Ambrosio (advisory ed.), Mathematics across Cultures: The History of Non-Western Mathematics. Science across Cultures: The History of Non-Western Science, 2. Dordrecht and London: Kluwer Academic Publishers, 2000. Pp. xx + 479. ISBN 0-7923-6481-3. £135.00, $217.00 (hardback). British Journal for the History of Science, 38(1):102–103, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028585.

Polenberg:2002:MJR

[Pol02] Richard Polenberg, editor. In the matter of J. Robert Oppenheimer: the security clearance hearing. Cornell paperbacks. Cornell University Press, Ithaca, NY, USA, 2002. ISBN 0-8014-8661-0 (paperback), 0-8014-3783-0 (cloth). xxxii + 409 pp. LCCN QC16.O62 I5

                                        2002. URL http://www.h-net.
                                        org/review/hrev-a0c5p6-aa.

Pooley:2009:BRS

[Poo09] Simon Pooley. Book review: Saul Dubow, A Commonwealth of Knowledge: Science, Sensibility and White South Africa 1820– 2000. Oxford and New York: Oxford University Press, 2006. Pp.

296. ISBN 978-0-199-29663-7. £60.00 (hardback). British Journal for the History of Science, 42 (1):152–154, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Porter:2001:BRM

[Por01] Roy Porter. Book review: Marina Frasca-Spada and Nick Jardine (eds.), Books and the Sciences in History. Cambridge: Cambridge University Press, 2000.

Pp. xiv + 438. ISBN 0-521-65939-6. £18.95, $29.95. British Journal for the History of Science, 34(2):233–250, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028021.

Porter:2002:MBH

[Por02] Roy Porter. Madness: a brief history. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-280266-6. xii + 241 pp. LCCN RC438 .P67

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0640/2001052329-d.html;
http://www.loc.gov/catdir/
enhancements/fy0726/
2001052329-b.html;
http://www.loc.gov/catdir/
toc/fy035/2001052329.html.

Porter:2003:BRP

[Por03] Theodore M. Porter. Book review: Philip Mirowski and Esther-Mirjam Sent (eds.), Science Bought and Sold: Essays in the Economics of Science. Chicago and London: University of Chicago Press, 2002. Pp. ix + 573. ISBN 0-226-53857-5. £21.00, $33.00 (paperback). British Journal for the History of Science, 36(3):381– 383, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028173.

Porter:2004:KPS

[Por04] Theodore M. Porter. Karl Pearson: the scientific life in a statistical age. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-11445-5. viii + 342 pp. LCCN QA276.P64 P67 2004.

URL  http://www.loc.gov/
catdir/description/prin051/
2003054820.html;   http://
www.loc.gov/catdir/toc/
prin051/2003054820.html.

Porter:2003:CHS

[Por09] Roy Porter, editor. The Cambridge history of science. Cambridge University Press, Cambridge, UK, 2003–2009. ISBN 0-521-57244-4 (vol. 3), 0-521-57243-6 (vol. 4), 0-521-57199-5 (vol. 5), 0-521-57201-0 (vol. 6), 0-521-59442-1 (vol. 7). xxvii + 865 (vol. 3), xxx + 912 (vol. 4), xxix + 678 (vol. 5), xxvi + 662 (vol. 6), xxvii + 762 (vol. 7) pp. LCCN Q125 .C32

2003. URL http://www.h-net.
org/review/hrev-a0d3l3-aa;
 http://www.h-net.org/
review/hrev-a0d7k4-aa;
http://www.h-net.org/
review/hrev-a0f4e9-aa;
http://www.loc.gov/catdir/
description/cam021/
2001025311.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2001025311-b.html;   http://
www.loc.gov/catdir/samples/cam033/
2001025311.html;   http://www.
loc.gov/catdir/toc/cam026/2001025311.
html.

Phillips:2007:PSL

[PP07] Peter W. B. Phillips and J. M.

(Jene M.) Porter, editors. Public science in liberal democracy. University of Toronto Press, Toronto, ON, Canada, 2007. ISBN 0-8020-9359-0 (hardcover). LCCN Q175.5 .P835 2007. URL

http://www.loc.gov/catdir/
toc/fy0803/2008371179.html.

Pressman:1998:LRP

[Pre98] Jack David Pressman. Last resort: psychosurgery and the limits of medicine. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-35371-8 (hardcover). xv + 555 pp. LCCN RD594 .P655 1998. URL http:

//www.loc.gov/catdir/
description/cam028/
97014074.html;   http://
www.loc.gov/catdir/samples/
cam031/97014074.html;
http://www.loc.gov/catdir/
toc/cam026/97014074.html.

Press:2003:BRC

[Pre03] Bob Press. Book review: Carolyn Rasmussen, A Museum for the People: A History of Museum Victoria and its Predecessors, 1854–2000. Melbourne: Scribe Publications, 2001. Pp. xvi + 423. ISBN 0-908011-69-5. AU$49.95 (hardback). British Journal for the History of Science, 36(2):235–236, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028237.

Principe:2007:CCS

[Pri07] Lawrence Principe, editor. Chymists and chymistry: studies in the history of alchemy and early modern chemistry. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2007. ISBN 0-88135-396-5. xiii + 274 pp. LCCN QD13 .I68 2007. URL http://www.

                                        loc.gov/catdir/toc/
                                        ecip0718/2007020807.html.

Prodger:2008:BRJ

[Pro08] Phillip Prodger. Book review: Jennifer Tucker, Nature Exposed: Photography as Eyewitness in Victorian Science. Baltimore: Johns Hopkins University Press, 2006. Pp. ix + 294. ISBN 0-8018-7991-4. £36.50 (hardback). British Journal for the History of Science, 41(3): 465–467, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Pomata:2005:HEE

[PS05] Gianna Pomata and Nancy G. Siraisi, editors. Historia: empiricism and erudition in early modern Europe. Transformations. MIT Press, Cambridge, MA, USA, 2005. ISBN 0-262-16229-6. viii + 490 pp. LCCN QH135 .H57

                                        2005. URL http://www.h-net.
                                        org/review/hrev-a0f4h0-aa;
                                         http://www.loc.gov/catdir/
                                        toc/fy0602/2005041560.html.

Proctor:2008:AMU

[PS08] Robert Proctor and Londa L.

Schiebinger, editors. Agnotology: the making and unmaking of ignorance. Stanford University Press, Stanford, CA, USA, 2008. ISBN 0-8047-5652-X (cloth), 0-8047-5901-4 (paperback). LCCN BD221 .A36 2008. URL http:

//www.loc.gov/catdir/toc/
ecip086/2007049811.html.

Pugliano:2009:BRA

[Pug09] Valentina Pugliano. Book review: Alessandro Ottaviani and Oreste Trabucco, Theatrum Naturae: La ricerca naturalistica tra erudizione e nuova scienza nell’Italia del primo Seicento. Napoli: La Città del Sole. 2007. Pp. 177. ISBN 978-88-8292-363-1. EUR22.00 (paperback). British Journal for the History of Science, 42(2):301– 302, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Purbrick:2001:GEN

[Pur01] Louise Purbrick, editor. The Great Exhibition of 1851: new interdisciplinary essays. Texts in culture. Manchester University Press, Manchester, UK, 2001. ISBN 0-7190-5591-1 (hardcover), 0-7190-5592-X (paperback). x + 217 pp. LCCN T690.B1 G75

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0612/2001044115-b.html;
http://www.loc.gov/catdir/
enhancements/fy0612/
2001044115-d.html;

                                        http://www.loc.gov/catdir/
                                        toc/fy034/2001044115.html.

Pyenson:2006:BRR

[Pye06]  Lewis Pyenson.  Book re-

view: Robert Fox (ed.), Centre and Periphery Revisited: The Structures of European Science, 1750–1914. Revue de la Maison française: Oxford, 2003. Pp.

191. ISSN 1477-3082. No price given (paperback). British Journal for the History of Science, 39 (1):122–123, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028551.

Pyenson:2009:BRJ

[Pye09] Lewis Pyenson. Book review: Josep Simon and Néstor Herran, with Tayra Lanuza-Navarro, Pedro Ruiz Castell and Ximo Guillem-Llobat (eds.), Beyond Borders: Fresh Perspectives in History of Science. Newcastle: Cambridge Scholars Publishing, 2008. Pp. vii + 377. ISBN 978-184-71843-2. £39.99 (hardback). British Journal for the History of Science, 42(2):299– 301, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Quinlan:2002:BRN

[Qui02] Sean M. Quinlan. Book review: Nicolaas A. Rupke (ed.), Medical Geography in Historical Perspective. Medical History, Supplement 20. London: Wellcome Trust Centre for the History of Medicine at UCL, 2000. Pp. xii + 227. ISBN 0-85484-072-9. £32.00, $50.00 (hardback). British Journal for the History of Science, 35(4):475–485, December 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Quinlan:2004:BRL

[Qui04] Sean M. Quinlan. Book review: Lisa Silverman, Tortured Subjects: Pain, Truth, and the Body in Early Modern France. Chicago and London: University of Chicago Press, 2001. Pp. xv + 264. ISBN 0-226-75754-4. $20.00 (paperback). British Journal for the History of Science, 37 (1):105–106, March 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028263.

Quirke:2008:BRM

[Qui08] Viviane Quirke. Book review:

Mariana Mazzucatto and Giovanni Dosi (eds.) Knowledge Accumulation and Industry Evolution: The Case of Pharma-Biotech. Cambridge: Cambridge University Press, 2006. Pp. xvii + 446. ISBN 978-0-521-85822-4. £48.00 (hardback). British Journal for the History of Science, 41(4):626–628, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Radick:2000:BRS

[Rad00a] Gregory Radick. Book review:

Stephen G. Alter, Darwinism and the Linguistic Image: Language, Race, and Natural Theology in the Nineteenth Century. Baltimore: Johns Hopkins University Press, 1999. Pp. xiii + 193. ISBN 0-8018-5882-8. £32.50 (hardback). British Journal for the History of Science, 33 (1):115–124, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028072.

Radick:2000:MCG

[Rad00b] Gregory Radick. Morgan’s canon, Garner’s phonograph, and the evolutionary origins of language and reason. British Journal for the History of Science, 33(1):3–23, March 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028063.

Radick:2002:BRI

[Rad02] Gregory Radick. Book review: Ian Hacking, The Social Construction of What? Cambridge, MA and London: Harvard University Press, 2000. Pp. x + 261. ISBN 0-674-00412-4. £11.50 (paperback). British Journal for the History of Science, 35 (1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028415.

Radder:2003:PSE

[Rad03a] Hans Radder, editor. The philosophy of scientific experimentation. University of Pittsburgh Press, Pittsburgh, PA, USA, 2003. ISBN 0-8229-5795-7 (paperback). xii + 311 pp. LCCN Q180.55.M4 P45

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0708/2002014515-d.html.

Radick:2003:BRJ

[Rad03b] Gregory Radick. Book reviews:

James E. Strick, Sparks of Life: Darwinism and the Victorian Debates over Spontaneous Generation. Cambridge, MA and London: Harvard University Press, 2000. Pp. xi + 283. ISBN 0-674-00292-X. £30.95 (hardback). James E. Strick (ed.), Evolution and the Spontaneous Generation Debate. Bristol: Thoemmes Press, 2001. Pp. 8007. 6 vols. ISBN 1-85506-872-9. £395.00, $630.00 (hardback). British Journal for the History of Science, 36 (2):241–244, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028242.

Radick:2007:STL

[Rad07] Gregory Radick. The simian tongue: the long debate about animal language. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-70224-3 (cloth). xiv + 577 pp. LCCN QL737.P9 R25

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0732/2007026793-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0827/
                                        2007026793-b.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0721/2007026793.html.

Raj:2009:BRH

[Raj09] Kapil Raj. Book review: Henry J. Noltie, Robert Wight and the Botanical Drawings of Rungiah and Govindoo. 3 vols. Edinburgh: Royal Botanic Gardens Edinburgh, 2008. Pp. 215, 208 and 88. ISBN 978-1-906129-02-6. £75.00 (paperback, boxed set). British Journal for the History of Science, 42(4):606–608, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ramati:2001:HTC

[Ram01] Ayval Ramati. The hidden truth of creation: Newton’s method of fluxions. British Journal for the History of Science, 34(4): 417–438, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028372.

Ramberg:2002:BRF

[Ram02] Peter J. Ramberg. Book review: Frederic L. Holmes and Trevor H. Levere (eds.), Instruments and Experimentation in the History of Chemistry. Dibner Institute Studies in the History of Science and Technology. Cambridge and London: MIT Press, 2000. Pp. xxi + 415. ISBN 0-262-08282-9. £34.50 (cloth). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028201.

Rasmussen:2001:MPH

[Ras01] Carolyn Rasmussen. A museum for the people: a history of Museum Victoria and its predecessors, 1854–2000. Scribe Publications, Melbourne, Australia, 2001. ISBN 0-908011-69-5. xiv + 423 pp. LCCN AM101.M52 R37 2001.

Rasmussen:2006:BRK

[Ras06] Nicolas Rasmussen. Book review: Karl Grandin, Nina Wormbs and Sven Widmalm (eds.), The Science– Industry Nexus: History, Policy, Implications. Nobel Symposium 123. Sagamore Beach, MA: Science History Publications and the Nobel Foundation, 2004. Pp. xvii + 457. ISBN 0-88135-365-5. $54.95 (hardback). British Journal for the History of Science, 39(3):440–441, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028488.

Ratcliff:2007:SMH

[Rat07] J. R. Ratcliff. Samuel Morland and his calculating machines c. 1666: the early career of a courtier-inventor in Restoration London. British Journal for the History of Science, 40 (2):159–179, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500715.

Raven:2000:BRK

[Rav00] Diederick Raven. Book review: Klaas van Berkel, Dijksterhuis: een biografie. Amsterdam: Uitgeverij Bert Bakker, 1996. Pp.

639. ISBN 90-351-16941. HFL75. British Journal for the History of Science, 33(2):231–254, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4027939.

Raven:2001:BRD

[Rav01] Diederick Raven. Book review: D. Van Dalen, Mystic, Geometer, and Intuitionist: The Life of L. E. J. Brouwer. Volume 1: The Dawning Revolution. Oxford: Clarendon Press, 1999. Pp. 456. ISBN 0-19-850297-4. £75.00. British Journal for the History of Science, 34(1): 97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028055.

Reardon:2005:RFI

[Rea05] Jenny Reardon. Race to the finish: identity and governance in an age of genomics. In-formation series. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11856-6 (cloth), 0-691-11857-4 (paperback). xii + 237 pp. LCCN QH431 .R248

2005. URL http://www.loc.
gov/catdir/description/
prin051/2004050569.html;
http://www.loc.gov/catdir/
enhancements/fy0734/
2004050569-b.html.

Rees:2007:BRD

[Ree07] Amanda Rees. Book review:

Donna Haraway, The Haraway Reader. New York and London: Routledge, 2004. Pp. viii + 352. ISBN 0-415-96688-4. £16.99 (paperback). British Journal for the History of Science, 40(1): 117–118, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500687.

Reed:2008:BRJ

[Ree08] James W. Reed. Book review: John Carson, The Measure of Merit: Talent, Intelligence, and Inequality in the French and American Republics, 1750–1940. Princeton: Princeton University Press, 2007. Pp. xvii + 401. ISBN 0-691-01715-8. £26.95 (hardback). British Journal for the History of Science, 41(4):613–614, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reid:2005:IFB

[Rei05a] Francis Reid. Isaac Frost’s two systems of astronomy (1846): plebeian resistance and scriptural astronomy. British Journal for the History of Science, 38(2):161–177, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Reill:2005:VNE

[Rei05b] Peter Hanns Reill. Vitalizing nature in the Enlightenment. University of California Press, Berkeley, CA, USA, 2005. ISBN 0-520-24135-5. x + 388 pp. LCCN Q175.32.V65 R45

2005. URL http://www.h-net.
org/review/hrev-a0e5a6-aa;
 http://www.loc.gov/catdir/
enhancements/fy0641/
2004023034-d.html;
http://www.loc.gov/catdir/
enhancements/fy0735/
2004023034-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip051/2004023034.html.

Reisch:2005:HCW

[Rei05c] George A. Reisch. How the Cold War transformed philosophy of science: to the icy slopes of logic. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83797-9 (hardcover), 0-521-54689-3 (paperback). xiv + 418 pp. LCCN Q174.8 .R45

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2004052545-d.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2004052545-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2004052545-b.html.

Reisz:2007:BRBa

[Rei07a] Emma Reisz. Book review:

Ben Marsden and Crosbie Smith, Engineering Empires: A Cultural History of Technology in Nineteenth-Century Britain. New York: Palgrave Macmillan, 2005. Pp. xi + 351. ISBN 0-333-77278-4. $65.00 (hardback). British Journal for the History of Science, 40(2):299–301, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500736.

Reisz:2007:BRBb

[Rei07b] Emma Reisz. Book review:

Benedikt Stuchtey (ed.), Science across the European Empires, 1800–1950. Oxford and New York: Oxford University Press/German Historical Institute London, 2005. Pp. viii + 376. ISBN 0-19-927629-2. £60.00 (hardback). British Journal for the History of Science, 40(3): 441–443, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500752.

Reidy:2008:THO

[Rei08] Michael S. Reidy. Tides of history: ocean science and Her Majesty’s Navy. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-70932-9 (cloth). xiv + 389 pp. LCCN VA454 .R25

2008. URL http://www.loc.
gov/catdir/enhancements/

                                        fy0801/2007028060-t.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0827/
                                        2007028060-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0827/
                                        2007028060-d.html.

Relaano:2002:SAC

[Rel02] Francesc Relaäno. The shaping of Africa: cosmographic discourse and cartographic science in late medieval and early modern Europe. Ashgate, 2002. ISBN 0-7546-0239-7. x + 271 pp. LCCN GA781 .R45 2002. URL

                                        http://www.loc.gov/catdir/
                                        toc/fy036/2001022170.html.

Rentetzi:2000:BRP

[Ren00] Maria Rentetzi. Book review: Peter Galison, Image and Logic: A Material Culture of Microphysics. Chicago: University of Chicago Press, 1997. Pp. xxv + 995. ISBN 0-226-27916-2. £63.00, $90.00. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027962.

Rentetzi:2005:DNS

[Ren05] Maria Rentetzi. Designing (for) a new scientific discipline: the location and architecture of the Institut für Radiumforschung in early twentieth-century vienna. British Journal for the History of Science, 38(3):275–306, September 2005. CODEN BJHSAT.

ISSN 0007-0874 (print), 1474-001X (electronic).

Renwick:2006:BRG

[Ren06] Chris Renwick. Book reviews:

Greta Jones and Robert A. Peel (eds.), Herbert Spencer: The Intellectual Legacy. London: The Galton Institute, 2004. Pp. xv + 154. ISBN 0-9504066-8-6. £5.00 (paperback). Abigail Lustig, Robert J. Richards and Michael Ruse (eds.), Darwinian Heresies. Cambridge: Cambridge University Press, 2004. Pp. vii+200. ISBN 0-521-81516-9. £40.00 (hardback). British Journal for the History of Science, 39(4):617–619, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028537.

Renwick:2007:BRL

[Ren07]  Chris Renwick.  Book re-

view: Libby Schweber, Disciplining Statistics: Demography and Vital Statistics in France and England, 1830–1885. Durham, NC and London: Duke University Press, 2006. Pp. ix + 277. ISBN 0-8223-3814-9. £14.99 (paperback). British Journal for the History of Science, 40(4): 616–618, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Renwick:2009:BRM

[Ren09] Chris Renwick. Book review:

Mark Francis, Herbert Spencer and the Invention of Modern Life. Stocksfield: Acumen Publishing, 2007. Pp. xiv + 434. ISBN 978-1-84465-086-6. £25.00 (hardback). British Journal for the History of Science, 42(1): 124–125, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Riordan:1997:CFB

[RH97] Michael Riordan and Lillian Hoddeson. Crystal fire: the birth of the information age. Norton, New York, NY, USA, 1997. ISBN 0-393-04124-7. x + 352 pp. LCCN TK7809 .R56 1997.

Ruskin:2004:JHC

[RH04] Steven Ruskin and Sir John Frederick Herschel. John Herschel’s Cape voyage: private science, public imagination, and the ambitions of empire. Science, technology, and culture, 1700–1945. Ashgate, 2004. ISBN 0-7546-3558-9. xxix + 229 pp. LCCN QB36.H59 R87 2004. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy045/2003056076.html.

Roll-Hansen:2005:LEP

[RH05] Nils Roll-Hansen, editor. The Lysenko effect: the politics of science, Control of nature. Humanity Books, Amherst, NY, USA, 2005. ISBN 1-59102-262-2 (paperback). LCCN QH31.L95

                                        R65 2005. URL http://www.
                                        loc.gov/catdir/toc/
                                        ecip0418/2004012970.html.

Richardson:2000:MI

[Ric00] Ken Richardson. The making of intelligence. Maps of the mind. Columbia University Press, New York, NY, USA, 2000. ISBN 0-231-12004-4 (hardcover). viii + 210 pp. LCCN BF431 .R413 2000.

Richmond:2001:BRP

[Ric01a] Marsha L. Richmond. Book review: Pnina G. Abir-Am and Clark A. Elliott (eds.), Commemorative Practices in Science: Historical Perspectives on the Politics of Collective Memory. Osiris, 14. Chicago: Chicago University Press, 2000. Pp. xii + 383. ISBN 0-226-00092-3. $39.00 (hardback); 0-226-00093-1. $25.00 (paperback). British Journal for the History of Science, 34(4):453–481, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028394.

Richmond:2001:BRR

[Ric01b] Marsha L. Richmond. Book review: Richard Keynes (ed.), Charles Darwin’s Zoology Notes and Specimen Lists from HMS Beagle. Cambridge: Cambridge University Press, 2000. Pp. xxxiv + 430. ISBN 0-521-46569-9. £95.00, $150. British Journal for the History of Science, 34 (1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028056.

Richards:2002:PPP

[Ric02a] Graham Richards. Putting psychology in its place: a critical historical overview. Routledge & Kegan Paul, London, UK and New York, NY, USA, second edition, 2002. ISBN 1-84169-233-6 (hardcover), 1-84169-234-4 (paperback). xiv + 368 pp. LCCN BF95 .R55

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0653/2002021328-d.html.

Richards:2002:RCL

[Ric02b] Robert J. (Robert John) Richards. The romantic conception of life: science and philosophy in the age of Goethe. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-71210-9. xix + 587 pp. LCCN PT361 .R53 2002. URL http://www.

                                        loc.gov/catdir/bios/
                                        uchi051/2002001769.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi051/
                                        2002001769.html;   http://
                                        www.loc.gov/catdir/toc/
                                        uchi051/2002001769.html.

Richardson:2003:LEL

[Ric03] Angelique Richardson. Love and eugenics in the late nineteenth century: rational reproduction and the new woman. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-818700-9. xv + 250 + 16 pp. LCCN PR868.L69 R53

                                        2003. URL http://www.h-net.

org/review/hrev-a0e2i5-aa;
 http://www.loc.gov/catdir/
enhancements/fy0615/
2003283013-d.html;
http://www.loc.gov/catdir/
enhancements/fy0615/
2003283013-t.html.

Rice:2004:BRJ

[Ric04]  Adrian  Rice.

Book review: Joseph W. Dauben and Christoph J. Scriba (eds.), Writing the History of Mathematics: Its Historical Development. Historical Studies: Science Networks, 27. Basel, Boston and Berlin: Birkhäuser, 2002. Pp. xxvii + 689. ISBN 3-7643-6167-0. EUR73.83, SFr 186.00 (paperback). British Journal for the History of Science, 37(4): 488–489, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028664.

Richards:2006:BRC

[Ric06] Robert J. Richards. Book review: Charles Darwin, The Descent of Man, and Selection in Relation to Sex. 2nd edition of 1874. With an introduction by James Moore and Adrian Desmond. Penguin Classics. London: Penguin, 2004. Pp. lxvi + 791. ISBN 0-140-43631-6. £9.99 (paperback). British Journal for the History of Science, 39(4):615–617, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028536.

Richmond:2008:BRE

[Ric08] Jesse Richmond. Book review: Edward J. Hackett, Olga Amsterdamska, Michael Lynch and Judy Wajcman (eds.) The Handbook of Science and Technology Studies, Third Edition. Cambridge, MA and London: MIT Press, 2007. Pp. xiv + 1065. ISBN 978-0-262-08364-5. £35.95 (hardback). British Journal for the History of Science, 41(4): 628–629, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Richards:2009:BRR

[Ric09a] Graham Richards. Book review: Roger Smith, Being Human: Historical Knowledge and the Creation of Human Nature. Manchester and New York: Manchester University Press, 2007. Pp. viii + 288. ISBN 978-0-7190-7498-1. £25.00 (hardback). British Journal for the History of Science, 42(3): 447–448, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Richardson:2009:BRA

[Ric09b] Angelique Richardson. Book review: Aileen Fyfe and Bernard Lightman (eds.), Science in the Marketplace: Nineteenth-Century Sites and Experiences. Chicago and London: University of Chicago Press, 2007. Pp. x + 410. ISBN 978 0 226 27650 2. $45.00, £23.50 (hardback). British Journal for the History of Science, 42(4):621–623, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rigden:2000:RSC

[Rig00] John S. Rigden. Rabi, scientist and citizen. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00435-3. xxix + 302 pp. LCCN QC16.R2 R54 2000.

Riley:2000:BRJ

[Ril00] David Riley. Book review: Jack Morrell, Science, Culture and Politics in Britain, 1750–1870. Variorum Collected Studies Series, CS567. Aldershot: Ashgate, 1997. Pp. xii + 336. ISBN 0-86078-633-1. £52.50. British Journal for the History of Science, 33(3):369–379, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027963.

Riley:2001:BRJ

[Ril01a] David Riley. Book review: James E. McLellan III and Harold Dorn, Science and Technology in World History: An Introduction. Johns Hopkins University Press, 1999. Pp. viii + 404. ISBN 0-8018-5868-2, £48 (hardback); 0-8018-5869-0, £13 (paperback). British Journal for the History of Science, 34(1):97–124, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028057.

Riley:2001:BRR

[Ril01b] David Riley. Book review: Roy M. MacLeod, The ‘Creed of Science’ in Victorian England. Variorum Collected Studies Series: CS598. Aldershot: Ashgate, 2000. Pp. 346. ISBN 0-86078-669-2. £57.50. British Journal for the History of Science, 34 (1):97–124, March 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028058.

Riley:2002:BRW

[Ril02]  Dave Riley.  Book review:

W. Clark, J. Golinski and S. Schaffer (eds.), The Sciences in Enlightened Europe. Chicago and London: University of Chicago Press, 1999. Pp. xiv + 566. ISBN 0-226-10939-9. £59.50 (hardback). British Journal for the History of Science, 35 (1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028416.

Risbey:2001:BRJ

[Ris01] James Risbey. Book review: James Rodger Fleming, Historical Perspectives on Climate Change. New York and Oxford: Oxford University Press, 1998. Pp. xiii + 194. ISBN 0-19-507870-5. No price given. British Journal for the History of Science, 34(3):341–373, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028117.

Riskin:2002:SAS

[Ris02] Jessica Riskin. Science in the age of sensibility: the sentimental empiricists of the French Enlightenment. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-72078-0, 0-226-72079-9 (paperback). xii + 338 pp. LCCN Q127.F8 R57

2002. URL http://www.loc.
gov/catdir/bios/uchi051/
2002005793.html;
http://www.loc.gov/catdir/
description/uchi051/
2002005793.html;
http://www.loc.gov/catdir/
toc/fy034/2002005793.html.

Riskin:2007:GRE

[Ris07] Jessica Riskin, editor. Genesis redux: essays in the history and philosophy of artificial life. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-72080-2 (cloth), 0-226-72081-0 (paperback). xv + 389 pp. LCCN BD418.8 .G46

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2007003437-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2007003437-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip079/2007003437.html.

Ritvo:2002:BRC

[Rit02]  Harriet  Ritvo.

Book review: Christine Kenyon-Jones, Kindred Brutes: Animals in Romantic-Period Writing. Nineteenth Century. Aldershot: Ashgate, 2001. Pp. viii + 229. ISBN 0-7546-0332-6. £42.50 (hardback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Ritvo:2009:BRN

[Rit09] Harriet Ritvo. Book review: Neil Pemberton and Michael Worboys, Mad Dogs and Englishmen: Rabies in Britain, 1830–2000. Basingstoke and New York: Palgrave Macmillan, 2007. Pp. x + 247. ISBN 978-0-230-54240-2. £45.00 (hardback). British Journal for the History of Science, 42(2):290–291, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rooney:2009:GOT

[RN09] David Rooney and James Nye. ‘Greenwich Observatory Time for the public benefit’: standard time and Victorian networks of regulation. British Journal for the History of Science, 42(1):5– 30, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Robbins:2002:ESP

[Rob02] Louise E. Robbins. Elephant slaves and pampered parrots: exotic animals in Eighteenth-Century Paris. Animals, history, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2002. ISBN 0-8018-6753-3 (hardcover). xiv + 349 pp. LCCN SF411.36.F8

R63 2002. URL http://www.e-
streams.com/es0601/es0601_
2241.html;   http://www.
loc.gov/catdir/bios/jhu051/
2001000243.html;
http://www.loc.gov/catdir/
description/jhu051/
2001000243.html;
http://www.loc.gov/catdir/
toc/fy032/2001000243.html.

Roberts:2003:BRJ

[Rob03] Michael B. Roberts. Book review: John M. Lynch (ed.), Creationism and Scriptural Geology, 1817–1857. Bristol: Thoemmes Press, 2002. Pp. xxiv + 2994. ISBN 1-85506-928-8. £595.00 (hardback). British Journal for the History of Science, 36(4): 483–485, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028223;
http://www.jstor.org/
stable/4028283.

Roberts:2005:BRJ

[Rob05] Michael Roberts. Book review:

John A. Moore, From Genesis to Genetics: The Case of Evolution and Creationism. Berkeley, Los Angeles and London: University of California Press, 2002. Pp. xvi + 223. ISBN 0-520-22441-8. £19.95, $27.50 (hardback). British Journal for the History of Science, 38(4):494– 495, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028475.

Robinson:2006:BRL

[Rob06] Michael Robinson. Book reviews: Londa Schiebinger, Plants and Empire: Colonial Bioprospecting in the Atlantic World. Cambridge, MA: Harvard University Press, 2004. Pp. xii + 306. ISBN 0-674-01487-1. £25.95 (hardback). Bernard Bailyn, Atlantic History: Concepts and Contours. Cambridge, MA: Harvard University Press, 2005. Pp.

149. ISBN 0-674-01688-2. £12.95 (hardback). British Journal for the History of Science, 39(4): 588–590, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028517.

Roberts:2007:BRJ

[Rob07]  Lissa Roberts.  Book

review: Jonathan Simon, Chemistry, Pharmacy and Revolution in France, 1777–1809. Science, Technology and Culture, 1700– 1945. Aldershot: Ashgate, 2005. Pp. vi + 189. ISBN 0-7546-5044-8. £45.00 (hardback). British Journal for the History of Science, 40(2):290–291, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500729.

Robinson:2008:BRD

[Rob08] Michael F. Robinson. Book review: David Buisseret (ed.) The Oxford Companion to World Exploration. Oxford: Oxford University Press, 2007. Vol. 1. Pp. xxviii + 478. Vol. 2. Pp. vii+501. ISBN 978-0-19-514922-7, 0-19-514922-X. £140.00 (hardback). British Journal for the History of Science, 41(3):457–458, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roberts:2009:MCR

[Rob09a] G. W. Roberts. Magnetism and chronometers: the research of the Reverend George Fisher. British Journal for the History of Science, 42(1):57–72, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Robinson:2009:BRJ

[Rob09b] Michael F. Robinson. Book review: James Delbourgo and Nicholas Dew (eds.), Science and Empire in the Atlantic World. New York: Routledge, 2008. Pp. xiv + 365. ISBN 978-0-415-96127-1. £18.99 (paperback). British Journal for the History of Science, 42(3): 461–462, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Robson:2009:BRS

[Rob09c]  Eleanor Robson.  Book re-

view: Serafina Cuomo, Technology and Culture in Greek and Roman Antiquity. Cambridge: Cambridge University Press, 2007. Pp. xi + 212. ISBN 978-0-521-00903-4. £15.99 (paperback). British Journal for the History of Science, 42(3): 451–453, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rocke:2001:NSA

[Roc01] Alan J. Rocke. Nationalizing science: Adolphe Wurtz and the battle for French chemistry. Transformations. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-18204-1 (hardcover). xi + 443 pp. LCCN QD22.W87 R63 2001.

Roland:2006:BRR

[Rol06] Alex Roland. Book review: Roger D. Launius and Janet R. Daley Bednarek (eds.), Reconsidering a Century of Flight. Chapel Hill and London: University of North Carolina Press, 2003. Pp. xii + 300. ISBN 0-8078-5488-3. £14.95, $19.95 (paperback). British Journal for the History of Science, 39(1): 145–146, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028570.

Roos:2007:SEN

[Roo07] Anna Marie Eleanor Roos. The salt of the earth: natural philosophy, medicine, and chymistry in England, 1650–1750, volume 3 of History of science and medicine library, 1872-0684. Brill, Leiden, The Netherlands, 2007. ISBN 90-04-16176-7. 293 pp. LCCN Q127.E8 R6712

2007. URL http://www.loc.
gov/catdir/toc/ecip0722/
2007027772.html.

Roos:2008:BRT

[Roo08] Anna Marie Roos. Book reviews:

Tara Nummedal, Alchemy and Authority in the Holy Roman Empire. Chicago and London: University of Chicago Press, 2007. Pp. xvii +260. ISBN 978-0-226-60856-3. $37.50, £22.00 (hardback). Bruce T. Moran, Andreas Libavius and the Transformation of Alchemy: Separating Chemical Cultures with Polemical Fire. Sagamore Beach, MA: Science History Publications/USA, 2007. Pp. viii + 344. ISBN 978-0-88135-395-7. $49.95 (hardback). British Journal for the History of Science, 41(4): 608–610, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Roux:2007:RAS

[Rou07] Sophie Roux, editor. Retours sur l’affaire Sokal. (French) [Return to the Sokal affair]. Collection Histoire des sciences. Série études. Harmattan, Paris, France, 2007. ISBN 2-296-02389-4. x + 190 pp. LCCN Q175.3 .R48 2007.

Rowland:2000:EJA

[Row00] Ingrid D. (Ingrid Drake) Rowland. The ecstatic journey: Athanasius Kircher in Baroque Rome. University of Chicago Library, Chicago, IL, USA, 2000.

ISBN 0-943056-25-X. xiii + 109 pp. LCCN Q127.I8 R68

                                        2000. URL http://www.loc.
                                        gov/catdir/description/
                                        uchi051/00027481.html;
                                        http://www.loc.gov/catdir/
                                        toc/uchi051/00027481.html.

Rowlinson:2002:CSH

[Row02] J. S. (John Shipley) Rowlinson. Cohesion: a scientific history of intermolecular forces. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-81008-6 (hardcover). viii + 333 pp. LCCN QC183 .R74

                                        2002. URL http://www.loc.
                                        gov/catdir/description/
                                        cam022/2002070878.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0731/
                                        2002070878-b.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam031/2002070878.html.

Ross:2002:MMI

[RP02a] Helen Elizabeth Ross and Cornelis Plug. The mystery of the moon illusion: exploring size perception. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-850862-X. viii + 277 pp. LCCN QP495 .R67

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002072659-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy033/2002072659.html.

RESL:2002:GAR

[RP02b] Royal Entomological Society of London and Berit Pedersen. A guide to the archives of the Royal Entomological Society. Ashgate, Aldershot, Hampshire, England, 2002. ISBN 0-7546-0106-4. x + 198 pp. LCCN QL461 .R6574 2002.

Roland:2002:SCD

[RS02] Alex Roland and Philip Shiman.

Strategic computing: DARPA and the quest for machine intelligence, 1983–1993. History of computing. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-18226-2 (hardcover). xxvi + 427 pp. LCCN QA76.88 .R65

2002. URL http://www.loc.
gov/catdir/toc/fy033/
2001056252.html.

Roberts:2007:MHI

[RSD07] Lissa Roberts, Simon Schaffer, and Peter Dear, editors. The mindful hand: inquiry and invention from the late Renaissance to early industrialisation, volume 9 of History of science and scholarship in the Netherlands. Koninkliijke Nederlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2007. ISBN 90-6984-483-4 (hardcover). xxvii + 503 pp. LCCN T173.8 .M565 2007.

Reid:2000:DSC

[RT00] Roddey Reid and Sharon Traweek, editors. Doing science + culture. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-92111-2, 0-415-92112-0. viii + 339 pp. LCCN Q175.5 .D65

2000. URL http://www.loc.

gov/catdir/enhancements/
fy0651/99056916-d.html;
http://www.loc.gov/catdir/
enhancements/fy1007/
99056916-b.html.

Rudwick:2004:NSG

[Rud04a] M. J. S. Rudwick. The new science of geology: studies in the earth sciences in the age of revolution, volume CS789 of Variorum collected studies. Ashgate, 2004. ISBN 0-86078-958-6. xviii + 316 pp. LCCN QE26.3 .R83

2004. URL http://www.loc.
gov/catdir/toc/ecip0411/
2003025799.html.

Rudwick:2004:BRR

[Rud04b] Martin Rudwick. Book reviews: René Sigrist (ed.), H.-B. de Saussure (1740–1799): Un Regard sur la terre. Bibliothèque d’histoire des sciences, 4. Geneva and Paris: Georg Editeur, 2001. Pp. x + 540. ISBN 2-8257-0740-6. No price given (paperback). Albert V. Carozzi and John K. Newman (eds.), Lectures on Physical Geography given in 1775 by Horace-Bénédict de Saussure at the Academy of Geneva/Cours de géographie physique donné en 1775 par Horace-Bénédict de Saussure à l’Académie de Genève. Geneva: Éditions Zoé, 2003. Pp. xxii+527. ISBN 2-88182-481-1. No price given (paperback). Horace-Bénédict de Saussure, Voyages dans les Alpes: Augmentés des Voyages en Valais, au Mont Cervin et autour du Mont Rose. With a Foreword by Albert V. Carozzi. Geneva: Editions Slatkine, 2002. Pp. xviii+300. ISBN 2-8321-0047-3. No price given (hardback). British Journal for the History of Science, 37(2):206–207, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028337.

Rudwick:2005:BLT

[Rud05a] M. J. S. Rudwick. Bursting the limits of time: the reconstruction of geohistory in the age of revolution. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-73111-1 (cloth). xxiv + 708 pp. LCCN Q127.E8 R83

2005. URL http://www.h-net.
org/review/hrev-a0e8e6-aa;
http://www.loc.gov/catdir/
enhancements/fy0617/
2004022007-b.html;
http://www.loc.gov/catdir/
enhancements/fy0617/
2004022007-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip051/2004022007.html.

Rudwick:2005:LDG

[Rud05b] M. J. S. Rudwick. Lyell and Darwin, geologists: studies in the earth sciences in the age of reform, volume CS818 of Variorum collected studies. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-86078-959-4. xvi + 185 + 6 pp. LCCN QE11 .R83 2005.

Rudwick:2007:BRJ

[Rud07] Martin Rudwick. Book review: Jack Morrell, John Phillips and the Business of Victorian Science. Aldershot: Ashgate, 2005. Pp. xix + 437. ISBN 1-84014-239-1. £57.50 (hardback). British Journal for the History of Science, 40(1):141–143, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500703.

Rudwick:2008:WBA

[Rud08] M. J. S. Rudwick. Worlds before Adam: the reconstruction of geohistory in the age of reform. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-73128-6 (cloth). xxii + 614 pp. LCCN QE13.E85 R83

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2007041676-b.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007041676-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip082/2007041676.html.

Rupke:2000:TSH

[Rup00a] Nicolaas Rupke. Translation studies in the history of science: the example of Vestiges. British Journal for the History of Science, 33(2):209–222, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027923.

Rupke:2000:MGH

[Rup00b] Nicolaas A. Rupke, editor. Medical geography in historical perspective, volume 20 (2000) of Medical history. Supplement. Wellcome Trust Centre for the History of Medicine at UCL, London, UK, 2000. ISBN 0-85484-072-9. 227 + 1 pp. LCCN RA792 .M425 2000.

Rupke:2008:BRR

[Rup08] Nicolaas A. Rupke. Book review:

Ralph O’Connor, The Earth on Show: Fossils and the Poetics of Popular Science, 1802–1856. Chicago and London: University of Chicago Press, 2007. Pp. xiii + 541. ISBN 978-0-226-61668-1. £23.50 (hardback). British Journal for the History of Science, 41(4):614–616, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Russell:2000:MFP

[Rus00] Colin Archibald Russell. Michael Faraday: physics and faith. Oxford portraits in science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-511763-8 (hardcover). 124 pp. LCCN QC16.F2 .R87

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0610/00027008-d.html.

Russell:2001:WNF

[Rus01] Edmund Russell. War and nature: fighting humans and insects with chemicals from World War I to Silent Spring. Studies in environment and history. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-79003-4 (hardcover), 0-521-79937-6 (paperback). xvii + 315 pp. LCCN QH545.C48 R87 2001.

                                        URL http://www.e-streams.
                                        com/es0805/es0805_2026.htm;
                                         http://www.loc.gov/catdir/
                                        description/cam021/
                                        00040323.html;   http://
                                        www.loc.gov/catdir/samples/
                                        cam031/00040323.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam026/00040323.html.

Rutkin:2008:BRL

[Rut08] H. Darrel Rutkin. Book review: Louise H. Curth, English Almanacs, Astrology and Popular Medicine: 1550–1700. Manchester: Manchester University Press, 2007. Pp. xi + 283. ISBN 978-0-7190-6928-4. £55.00 (hardback). British Journal for the History of Science, 41(4): 606–608, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Rozwadowski:2004:MNG

[RV04] Helen M. Rozwadowski and David K. Van Keuren, editors. The machine in Neptune’s garden: historical perspectives on technology and the marine environment. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2004. ISBN 0-88135-372-8. LCCN GC28 .M345 2001.

Ryckman:2005:RRP

[Ryc05] Thomas Ryckman. The reign of Relativity: philosophy in physics, 1915–1925. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517717-7. ix + 317 pp. LCCN QC173.52 .R93

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0617/2004041576-d.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2004041576-b.html.

Rylance:2000:VPB

[Ryl00] Rick Rylance. Victorian psychology and British culture, 1850– 1880. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-812283-7. x + 355 pp. LCCN BF108.G7

R95 2000. URL http://www.
loc.gov/catdir/
enhancements/fy0610/
00036748-d.html;
http://www.loc.gov/catdir/
enhancements/fy0610/
00036748-t.html.

Smith:2005:GED

[SA05] C. U. M. (Christopher Upham Murray) Smith and Robert Arnott, editors. The genius of Erasmus Darwin. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-7546-3671-2. xvii + 416 pp. LCCN QH31.D3 G46 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0415/2004001762.html.

Shea:2006:GOS

[SA06] William R. Shea and Mariano Artigas. Galileo observed: science and the politics of belief. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2006. ISBN 0-88135-356-6. xi + 212 pp. LCCN QB36.G2 S459 2006.

Samuelson:2007:BRG

[Sam07] Norbert Samuelson. Book review: Gad Freudenthal, Science in the Medieval Hebrew and Arabic Traditions. Variorum Collected Studies Series CS803. Aldershot: Ashgate, 2005. Pp. xx + 350. ISBN 0-86078-952-7. £62.50 (hardback). British Journal for the History of Science, 40 (1):126–127, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500693.

Sassi:2001:SMA

[Sas01] Maria Michela Sassi. The science of man in ancient Greece. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-73530-3. xxx + 224 pp. LCCN B187.M25 .S2713

                                        2001. URL http://www.loc.
                                        gov/catdir/bios/uchi051/
                                        00009177.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi052/
                                        00009177.html;
                                        http://www.loc.gov/catdir/
                                        toc/uchi051/00009177.html.

Stephenson:2000:UUI

[SBF00] Bruce Stephenson, Marvin Bolt, and Anna Felicity Friedman. The universe unveiled: instruments and images through history. Cambridge University Press, Cambridge, UK, 2000.

ISBN 0-521-79143-X (hardcover). 152 pp. LCCN QB86 .S74 2000. URL http://www.

loc.gov/catdir/description/
cam021/00059883.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
00059883-b.html;
http://www.loc.gov/catdir/
toc/cam021/00059883.html.

Schaffer:1984:RU

[Sch84] S. Schaffer. Review: [untitled].

British Journal for the History of Science, 17(1):86–87, March 1984. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4026021.

Schutt:2000:SNS

[Sch00] Hans-Werner Schütt. Auf der Suche nach dem Stein der Weisen: die Geschichte der Alchemie. (German) [On the search for the stone of prophecy: the history of alchemy]. C. H. Beck, München, Germany, 2000. ISBN 3-406-46638-9. 602 pp. LCCN QD13 .S38 2000.

Schwarz:2001:BRC

[Sch01]  Angela Schwarz.  Book re-

view: E. C. Spary, Utopia’s Garden: French Natural History from Old Regime to Revolution. Chicago and London: University of Chicago Press, 2000. Pp. xv + 321. ISBN 0-226-76863-5. £16.00, $25.00 (paperback). British Journal for the History of Science, 34(3):341– 373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028118.

Schupbach:2002:BRS

[Sch02a] William Schupbach. Book review: Susan Wheeler, Five Hundred Years of Medicine in Art: An Illustrated Catalogue of Prints and Drawings from the Clements C. Fry Collection in the Harvey Cushing/John Hay Whitney Medical Library at Yale University. Aldershot: Ashgate, 2001. Pp. xxviii + 363. ISBN 0-85967-992-6. £79.50 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schwarz:2002:BRR

[Sch02b]  Angela Schwarz.  Book re-

view: Richard Drayton, Nature’s Government: Science, Imperial Britain, and the ‘Improvement’ of the World. New Haven and London: Yale University Press, 2000. Pp. xxi + 346. ISBN 0-300-05976-0. £25.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028202.

Schmidgen:2003:BRS

[Sch03a] Henning Schmidgen. Book review: Stephen Hilgartner, Science on Stage: Expert Advice as Public Drama. Writing Science. Stanford: Stanford University Press, 2000. Pp. xv + 214. ISBN 0-8047-3646-4. £11.95, $18.95 (paperback). British Journal for the History of Science, 36 (1):87–127, March 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028321.

Schwartz:2003:HCT

[Sch03b] Maxime Schwartz. How the cows turned mad. University of California Press, Berkeley, CA, USA, 2003. ISBN 0-520-23531-2. viii + 238 pp. LCCN RA644.P93 S3913 2003. URL http://www.

loc.gov/catdir/bios/
ucal052/2002075514.html;
http://www.loc.gov/catdir/
description/ucal042/
2002075514.html;   http://
www.loc.gov/catdir/toc/
ucal041/2002075514.html.

Schiebinger:2004:PEC

[Sch04a] Londa L. Schiebinger. Plants and empire: colonial bioprospecting in the Atlantic world. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01487-1 (cloth). x + 306 pp. LCCN RG137.45 .S35

2004. URL http://www.h-net.
org/review/hrev-a0e5c8-aa.

Schultz:2004:HSE

[Sch04b] Bart Schultz. Henry Sidgwick, eye of the universe: an intellectual biography. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-82967-4. xx + 858 pp. LCCN BJ604.S5 S38

                                        2004. URL http://www.loc.
                                        gov/catdir/description/
                                        cam032/2003055354.html;
                                        http://www.loc.gov/catdir/
                                        toc/cam032/2003055354.html.

Schabas:2005:NOE

[Sch05a] Margaret Schabas. The natural origins of economics. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-73569-9 (cloth). xi + 231 pp. LCCN HB74.S35 S33

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0622/2005014061-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0622/
                                        2005014061-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0513/2005014061.html.

Schaffer:2005:LBB

[Sch05b] Gavin Schaffer. ‘like a baby with a box of matches’: British scientists and the concept of ‘race’ in the inter-war period. British Journal for the History of Science, 38(3):307–324, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schickore:2006:MIE

[Sch06a] Jutta Schickore. Misperception, illusion and epistemological optimism: vision studies in early nineteenth-century Britain and Germany. British Journal for the History of Science, 39(3): 383–405, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schweber:2006:DSD

[Sch06b] Libby Schweber. Disciplining statistics: demography and vital statistics in France and England, 1830/1885. Politics, history, and culture. Duke University Press, Durham, NC, USA, 2006. ISBN 0-8223-3825-4 (cloth), 0-8223-3814-9 (paperback). 277 pp. LCCN HB853.F8

S38 2006. URL http://www.
loc.gov/catdir/toc/
ecip0611/2006010426.html.

Schickore:2007:MEH

[Sch07] Jutta Schickore. The microscope and the eye: a history of reflections, 1740–1870. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-73784-5 (cloth). ix + 317 pp. LCCN QH204 .S34

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0732/2007006102-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007006102-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0710/2007006102.html.

Schickore:2008:BRW

[Sch08a] Jutta Schickore. Book review:

William T. Scott and Martin X. Moleski, Michael Polanyi, Scientist and Philosopher. Oxford: Oxford University Press, 2005. Pp. xx + 364. ISBN 0-19-517433-X. £26.99, $45.00 (hardback). British Journal for the History of Science, 41(1):152– 154, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Schweber:2008:EOM

[Sch08b] S. S. (Silvan S.) Schweber. Einstein and Oppenheimer: the meaning of genius. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02828-7. xiv + 412 pp. LCCN QC16.E5 S3285 2008. URL http://www.

                                        loc.gov/catdir/toc/ecip082/
                                        2007043108.html.

Scott:1998:SLS

[Sco98] James C. Scott. Seeing like a state: how certain schemes to improve the human condition have failed. Yale agrarian studies; The Yale ISPS series. Yale University Press, New Haven, CT, USA, 1998. ISBN 0-300-07016-0. xiv + 445 pp. LCCN HD87.5 .S365

                                        1998. URL http://www.h-net.
                                        org/review/hrev-a0a9o0-aa.

Scott:2004:VIU

[Sco04] Anne Scott. ‘visible incarnations of the unseen’: Henry Drummond and the practice of typological exegesis. British Journal for the History of Science, 37(4):435–454, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028150;
                                        http://www.jstor.org/
                                        stable/4028642.

Stephens:2000:ETI

[SD00] Carlene Stephens and Maggie Dennis. Engineering time: inventing the electronic wristwatch. British Journal for the History of Science, 33(4):477– 497, December 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028031.

Secord:2000:VSE

[Sec00] James A. Secord. Victorian sensation: the extraordinary publication, reception, and secret authorship of Vestiges of the Natural History of Creation. University of Chicago Press, Chicago, IL, USA, 2000. ISBN 0-226-74410-8 (cloth). xix + 624 pp. LCCN QH363 .S4 2000. URL

http://www.loc.gov/catdir/
bios/uchi051/00009124.html;
http://www.loc.gov/catdir/
description/uchi051/
00009124.html;
http://www.loc.gov/catdir/
toc/uchi051/00009124.html.

Secord:2005:EH

[Sec05] James Secord. The electronic harvest. British Journal for the History of Science, 38(4): 463–467, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Secord:2008:BRD

[Sec08] James Secord. Book review:

David Knight, Public Understanding of Science: A History of Communicating Scientific Ideas. Abingdon and New York: Routledge, 2006. Pp. viii + 232. ISBN 0-415-20638-3. £65.00 (hardback). British Journal for the History of Science, 41 (1):144–145, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sedley:2007:CCA

[Sed07] D. N. Sedley. Creationism and its critics in antiquity, volume 66 of Sather classical lectures; The Joan Palevsky imprint in classical literature. University of California Press, Berkeley, CA, USA, 2007. ISBN 0-520-25364-7 (cloth), (paperback), 0-520-26006-6 (paperback). xvii + 269 pp. LCCN BD581 .S377

2007. URL http://www.loc.
gov/catdir/toc/ecip0719/
2007021528.html.

Seguin:2005:BRM

[Seg05]  Eve  Seguin.

Book review: Maxime Schwartz, How the Cows Turned Mad. Translated by Edward Schneider. Berkeley and London: University of California Press, 2003. Pp. viii + 238. ISBN 0-520-23531-2. £17.95, $24.95 (hardback). British Journal for the History of Science, 38(4):492– 494, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028474.

Selin:2000:MAC

[Sel00] Helaine Selin, editor. Mathematics across cultures: the history of non-Western mathematics, volume 2 of Science across cultures. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6481-3 (hardbound). xx + 479 pp. LCCN QA21 . M3612

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0814/00056016-d.html;
http://www.loc.gov/catdir/
enhancements/fy0814/
00056016-t.html.

Selcer:2008:SWA

[Sel08] Perrin Selcer. Standardizing wounds: Alexis Carrel and the scientific management of life in the First World War. British Journal for the History of Science, 41(1):73–107, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sommerfeld:2000:WBG

[SEM04]  Arnold  Som-

merfeld, Michael Eckert, and Karl Märker. Wissenschaftlicher Briefwechsel (German) [Scientific Correspondence]. Deutsches Museum, Berlin, Germany, 2000–2004. ISBN 3-928186-44-2 (set), 3-928186-49-3 (Vol. 1), 3-928186-53-1 (Vol. 2). 694 (vol. 1), 735 (vol. 2) pp. LCCN QC16.S76 A4 2000.

Sengoopta:2002:BM

[Sen02a] Chandak Sengoopta. Back matter. British Journal for the History of Science, 35(1):128, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028422.

Sengoopta:2002:BRS

[Sen02b] Chandak Sengoopta. Book review: Simon A. Cole, Suspect Identities: A History of Fingerprinting and Criminal Identification. Cambridge, MA and London: Harvard University Press, 2001. Pp. 369. ISBN 0-674-00455-8. £23.95 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028417.

Sengoopta:2006:BRH

[Sen06a]  Chandak Sengoopta.  Book

review: Henry C. Lee and Frank Tirnady, Blood Evidence: How DNA is Revolutionizing the Way We Solve Crimes. Oxford: Perseus Publishing, 2003. Pp. xxx + 418. ISBN 0-7382-0602-4. £19.99 (hardback). British Journal for the History of Science, 39 (1):152–153, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028575.

Sengoopta:2006:MSQ

[Sen06b] Chandak Sengoopta. The most secret quintessence of life: sex, glands, and hormones, 1850– 1950. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-74863-4 (cloth). xii + 354 pp. LCCN QP572.S4 S46

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0666/2005027321-b.html;
http://www.loc.gov/catdir/
enhancements/fy0666/
2005027321-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0519/2005027321.html.

Seth:2004:BRE

[Set04] Suman Seth. Book review:

Erhard Scholz (ed.), Hermann Weyl’s Raum-Zeit-Materie and a General Introduction to His Scientific Work. DMV Seminar,

30. Basel, Boston and Berlin: Birkhäuser, 2001. Pp. vii + 403. ISBN 3-7643-6476-9. No price given (paperback). British Journal for the History of Science, 37(3):361–362, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028141;
http://www.jstor.org/
stable/4028443.

Seth:2007:CCM

[Set07] Suman Seth. Crisis and the construction of modern theoretical physics. British Journal for the History of Science, 40 (1):25–51, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500681.

Scott:2000:FWO

[SFS^{+}00] Andrew C. Scott, David Freedberg, Francesco Solinas, Jo Taylor, and Francis Haskell. Fossil woods and other geological specimens. The paper museum of Cassiano dal Pozzo. Series B, natural history. Harvey Miller, Turnhout, Belgium, 2000. ISBN 1-872501-91-5. 424 pp. LCCN NC255 .S386 2000.

                                        URL http://www.gbv.de/dms/
                                        goettingen/324414242.pdf.

Shapin:2000:DDR

[Sha00] Steven Shapin. Descartes the doctor: rationalism and its therapies. British Journal for the History of Science, 33(2): 131–154, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027919.

Shapin:2002:BRS

[Sha02] Steven Shapin. Book review: Stephen Gaukroger, Francis Bacon and the Transformation of Early-Modern Philosophy. Cambridge: Cambridge University Press, 2001. Pp. xii + 249. ISBN 0-521-80536-8. £14.95, $21.95 (paperback). British Journal for the History of Science, 35 (1):97–123, March 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028418.

Shamdasani:2003:JMM

[Sha03a] Sonu Shamdasani. Jung and the making of modern psychology: the dream of a science. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-53909-9 (paperback), 0-521-83145-8 (hard). xvi + 387 pp. LCCN BF173 .S485

2003. URL http://www.loc.
gov/catdir/description/
cam032/2003048562.html;
http://www.loc.gov/catdir/
enhancements/fy0732/
2003048562-b.html;
http://www.loc.gov/catdir/
toc/cam032/2003048562.html.

Shapin:2003:BRD

[Sha03b]  Steven Shapin.  Book re-

view: David N. Livingstone, Science, Space and Hermeneutics. Hettner-Lectures, 5. Heidelberg: Department of Geography, University of Heidelberg, 2002. Pp. 116. ISBN 3-88570-505-2. No price given (paperback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028322.

Shapin:2007:BRM

[Sha07] Steven Shapin. Book reviews:

Matthew L. Jones, The Good Life in the Scientific Revolution: Descartes, Pascal, Leibniz, and the Cultivation of Virtue. Chicago and London: University of Chicago Press, 2006. Pp. xvii + 384. ISBN 0-226-40954-6. $65.00, £46.00 (hardback). ISBN 0-226-40955-4. $27.50, £17.50 (paperback). British Journal for the History of Science, 40(4): 602–603, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shank:2008:NWB

[Sha08a]  John Bennett Shank.  The

Newton wars and the beginning of the French Enlightenment. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-74945-2 (cloth). xv + 571 pp. LCCN Q127.F8 S53

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0827/2008006414-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0827/
                                        2008006414-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0810/2008006414.html.

Shapin:2008:SLM

[Sha08b] Steven Shapin. The scientific life: a moral history of a late modern vocation. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-75024-8 (paperback). xvii + 468 pp. LCCN Q180.55.M67 S53

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip0810/
                                        2008004588.html.

Shapin:2009:BRN

[Sha09]  Steven Shapin.  Book re-

view: Noretta Koertge (ed.), New Dictionary of Scientific Biography. 8 vols. Detroit, MI: Charles Scribner’s Sons, Gale/ Cengage Learning, 2008. ISBN 978-0684313207. $995.00 (hardback). British Journal for the History of Science, 42(1):116– 117, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shea:2000:SVI

[She00] William R. Shea, editor. Science and the visual image in the Enlightenment, volume 4 of European studies in science history and the arts. Science History Publications/USA, Canton, MA, USA and Sagamore Beach, MA, USA, 2000. ISBN 0-88135-285-3. viii + 232 pp. LCCN Q222 .S35 2000.

Shermer:2001:BSW

[She01] Michael Shermer. The borderlands of science: where sense meets nonsense. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514326-4. viii + 360 pp. LCCN Q173 .S56

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0639/00063673-d.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
00063673-b.html.

Shea:2007:BRM

[She07] William Shea. Book review:

Mario Biagioli, Galileo’s Instruments of Credit: Telescopes, Images, Secrecy. Chicago and London: University of Chicago Press, 2006. Pp. xi + 302. ISBN 0-226-04561-7. $25.00, £22.50 (hardback). British Journal for the History of Science, 40(4): 601–602, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shipley:2003:BM

[Shi03a] Brian C. Shipley. Back matter. British Journal for the History of Science, 36(4):497– 512, December 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028228.

Shipley:2003:BRN

[Shi03b] Brian C. Shipley. Book review: Naomi Oreskes (ed.) with Homer L E Grand, Plate Tectonics: An Insider’s History of the Modern Theory of the Earth. Boulder and Oxford: Westview Press, 2001. Pp. xxiv + 424. ISBN 0-8133-3981-2. $35.00 (hardback). British Journal for the History of Science, 36(4):487–488, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4028225;   http://www.
                                        jstor.org/stable/4028285.

Shotter:2003:BRT

[Sho03]  David Shotter.  Book re-

view: Tim Murray (ed.), Encyclopaedia of Archaeology: The Great Archaeologists. Santa Barbara: ABC–CLIO, 1999. Pp. xxii + 950. ISBN 1-57607-199-5. $150.00 (hardback). British Journal for the History of Science, 36(4):480, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028220;

http://www.jstor.org/
stable/4028280.

Shriar:2008:BRH

[Shr08] Efram Sera Shriar. Book review: Henrika Kuklick (ed.) A New History of Anthropology. Oxford: Blackwell, 2008. Pp. xiii + 402. ISBN 978-0-631-22600-0. £22.99 (paperback). British Journal for the History of Science, 41(3):452–453, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Shriar:2009:BRC

[Shr09] Efram Sera Shriar. Book review: Cristina Grasseni (ed.), Skilled Visions: Between Apprenticeship and Standards. New York and Oxford: Berghahn, 2007. Pp. viii + 226. ISBN 978-1845-45210-0. £45.00 (hardback). British Journal for the History of Science, 42(4):604– 606, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Siegfried:2002:EAH

[Sie02] Robert Siegfried. From elements to atoms: a history of chemical composition, volume 92, pt. 4 of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 2002. ISBN 0-87169-924-9 (paperback). x + 278 pp. LCCN QD461 .S463 2002.

Silverman:2001:TSP

[Sil01] Lisa Silverman. Tortured subjects: pain, truth, and the body in early modern France. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-75753-6 (cloth), 0-226-75754-4 (paperback). xv + 264 pp. LCCN HV8599.F7 S55 2001. URL

http://www.loc.gov/catdir/
bios/uchi051/00009586.html;
 http://www.loc.gov/catdir/
description/uchi051/
00009586.html;
http://www.loc.gov/catdir/
toc/uchi051/00009586.html.

Silverman:2003:LMA

[Sil03] Kenneth Silverman. Lightning man: the accursed life of Samuel F. B. Morse. Alfred A. Knopf, New York, NY, USA, 2003. ISBN 0-375-40128-8. vi + 503 pp. LCCN TK5243.M7 S55

2003. URL http://www.loc.
gov/catdir/bios/random054/
2002043613.html;
http://www.loc.gov/catdir/
description/random0414/
2002043613.html;   http://
www.loc.gov/catdir/samples/
random045/2002043613.html.

Simon:2002:AHN

[Sim02]  Jonathan Simon.  Analysis

and the hierarchy of nature in eighteenth-century chemistry. British Journal for the History of Science, 35(1):1–16, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028400.

Simoes:2004:TPL

[Sim04] Ana Simões. Textbooks, popular lectures and sermons: the quantum chemist Charles Alfred Coulson and the crafting of science. British Journal for the History of Science, 37(3): 299–342, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028426.

Simon:2005:CPR

[Sim05] Jonathan Simon. Chemistry, pharmacy and revolution in France, 1777–1809. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-7546-5044-8. vi + 189 pp. LCCN RS67.F8 S56 2005. URL http:/

/www.loc.gov/catdir/toc/
ecip057/2005004152.html.

Siraisi:2001:MIU

[Sir01] Nancy G. Siraisi. Medicine and the Italian universities, 1250– 1600, volume 12 of Education and society in the Middle Ages and Renaissance, 0926-6070. Brill, Leiden, The Netherlands, 2001. ISBN 90-04-11942-6. ix + 389 pp. LCCN R791.A6 S573 2001.

Sirotkina:2006:BRG

[Sir06] Irina Sirotkina. Book review:

Georges Didi-Huberman, Invention of Hysteria: Charcot and the Photographic Iconography of the Salpêtrière. Translated by Alisa Hartz. Cambridge, MA and London: MIT Press, 2003. Pp. xii + 373. ISBN 0-262-04215-0. £23.50 (hardback). British Journal for the History of Science, 39(2):303–305, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028628.

Sirotkina:2009:BRM

[Sir09] Irina Sirotkina. Book review: Margareta Tillberg, Tsvetnaia vselennaia: Mikhail Matiushin ob iskusstve i zrenii. Transl. from English by D. Dukhavina and M. Iarosh. Moscow: Novoe Literaturnoe Obozrenie, 2008. pp.

512. ISBN 978 5 867 93600 6. No price given (paperback). British Journal for the History of Science, 42(4):609–611, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sivasundaram:2005:NGE

[Siv05] Sujit Sivasundaram. Nature and the godly empire: science and evangelical mission in the Pacific, 1795–1850, volume 7 of Cambridge social and cultural histories. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-84836-9 (hardcover). xi + 244 pp. LCCN BV3670 .S57

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0632/2006295546-d.html;
http://www.loc.gov/catdir/
enhancements/fy0632/
2006295546-t.html;

http://www.loc.gov/catdir/
enhancements/fy0733/
2006295546-b.html.

Sivasundaram:2006:BRT

[Siv06] Sujit Sivasundaram. Book review: Tony Ballantyne (ed.), Science, Empire and the European Exploration of the Pacific. The Pacific World: Lands, Peoples and History of the Pacific, 1500–1900. Vol. 6. Aldershot: Ashgate, 2004. Pp. xxxv + 367. ISBN 0-7546-3562-7. £75.00 (hardback). British Journal for the History of Science, 39(4): 591–593, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028519.

Sivasundaram:2008:BRF

[Siv08] Sujit Sivasundaram. Book reviews: Felix Driver and Luciana Martins (eds.) Tropical Visions in an Age of Empire. Chicago and London: University of Chicago Press, 2005. Pp. xii + 279. ISBN 0-226-16472-1. $25.00, £16.00 (paperback). David Arnold, The Tropics and the Traveling Gaze: India, Landscape, and Science, 1800–1856. Seattle and London: University of Washington Press, 2006. Pp. xiv+298. ISBN 0-295-98581-X. $50.00, £32.95 (hardback). British Journal for the History of Science, 41(3):455–456, September 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Skelton:2004:BRR

[Ske04] Peter Skelton. Book review: Randal Keynes, Annie’s Box: Charles Darwin, His Daughter and Human Evolution. London: Fourth Estate, 2001. Pp. xiv + 331. ISBN 1-84115-060-6. £16.99 (hardback). British Journal for the History of Science, 37(3): 353–354, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028135;
                                        http://www.jstor.org/
                                        stable/4028437.

Shteir:2006:FIS

[SL06] Ann B. Shteir and Bernard V. Lightman, editors. Figuring it out: science, gender, and visual culture. Interfaces, studies in visual culture. Dartmouth College Press, Hanover, NH, USA, 2006. ISBN 1-58465-602-6 (cloth), 1-58465-603-4 (paperback). xxx + 386 pp. LCCN Q223 .F545

                                        2006. URL http://www.loc.
                                        gov/catdir/toc/ecip0617/
                                        2006022093.html.

Sleigh:2001:BRP

[Sle01] Charlotte Sleigh. Book review: Paul Lawrence Farber, Finding Order in Nature: The Naturalist Tradition from Linnaeus to E. O. Wilson. Johns Hopkins Introductory Studies in the History of Science. Baltimore and London: Johns Hopkins University Press, 2000. Pp. x + 136. ISBN 0-8018-6390-2. £12.50 (paperback). British Journal for the History of Science, 34(4): 453–481, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028395.

Sleigh:2002:BRM

[Sle02] Charlotte Sleigh. Book review:

Michael T. Ghiselin and Alan E. Leviton (eds.), Cultures and Institutions of Natural History. Essays in the History and Philosophy of Science. Memoirs of the California Academy of Sciences, 25. San Francisco: California Academy of Sciences, 2000. Pp. 363. ISBN 0-940228-48-3. $40.00 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028203.

Sleigh:2003:BRM

[Sle03]  Charlotte Sleigh.  Book re-

view: Michael H. Whitworth, Einstein’s Wake: Relativity, Metaphor, and Modernist Literature. Oxford: Oxford University Press, 2001. Pp. xi + 254. ISBN 0-19-818640-1. £45.00 (hardback). British Journal for the History of Science, 36 (2):238–239, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028240.

Sleigh:2005:BRH

[Sle05] Charlotte Sleigh. Book review: Helen Small and Trudi Tate (eds.), Literature, Science, Psychoanalysis, 1830–1970: Essays in Honour of Gillian Beer. Oxford: Oxford University Press, 2003. Pp. vii + 255. ISBN 0-19-96667-0. £50.00 (hardback). British Journal for the History of Science, 38(2):237–238, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028712.

Sleigh:2007:SLB

[Sle07] Charlotte Sleigh. Six legs better: a cultural history of myrmecology. Animals, history, culture. The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN 0-8018-8445-4 (hardcover). viii + 302 + 6 pp. LCCN QL568.F7 S57

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0664/2006019846-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0664/
                                        2006019846-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0617/2006019846.html.

Sleigh:2009:BRS

[Sle09a] Charlotte Sleigh. Book review: Steven McClean, The Early Fiction of H. G. Wells: Fantasies of Science. Basingstoke: Palgrave Macmillan, 2009. Pp. ix + 242. ISBN 978-0-230-53562-6. £50.00 (hardback). British Journal for the History of Science, 42(4):

623–624, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sleigh:2009:BRT

[Sle09b]  Charlotte Sleigh.  Book re-

view: Timothy Boon, Films of Fact: A History of Science in Documentary Films and Television. London: Wallflower Press, 2008. Pp. x + 312. ISBN 978-1-905674-37-4. £16.99 (paperback). British Journal for the History of Science, 42(1):148– 149, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Slotten:2000:RTR

[Slo00] Hugh Richard Slotten. Radio and television regulation: broadcast technology in the United States, 1920–1960. The Johns Hopkins University Press, Baltimore, MD, USA, 2000. ISBN 0-8018-6450-X. xv + 308 pp. LCCN HE8689.8 .S56 2000. URL

http://www.loc.gov/catdir/
bios/jhu051/00008153.html;
http://www.loc.gov/catdir/
description/jhu051/
00008153.html.

Slotten:2004:HDC

[Slo04] Ross A. Slotten. The heretic in Darwin’s court: the life of Alfred Russel Wallace. Columbia University Press, New York, NY, USA, 2004. ISBN 0-231-13010-4 (cloth). viii + 602 pp. LCCN QH31.W2 S535 2004. URL

http://www.loc.gov/catdir/
toc/fy051/2003068833.html.

Scott:2005:MPS

[SM05a] William T. (William Taussig) Scott and Martin X. Moleski. Michael Polanyi: scientist and philosopher. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517433-X (hardcover). xx + 364 pp. LCCN B945.P584 S365

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0636/2004010795-d.html;
http://www.loc.gov/catdir/
enhancements/fy0723/
2004010795-b.html.

Solomon:2005:FBR

[SM05b]  Julie  Robin  Solomon

and Catherine Gimelli Martin, editors. Francis Bacon and the refiguring of early modern thought: essays to commemorate The Advancement of Learning (1605–2005). Literary and scientific cultures of early modernity. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-7546-5359-5. ix + 257 pp. LCCN B1194 .F72 2005. URL http://

www.loc.gov/catdir/toc/
ecip057/2005002769.html.

Smith:2000:BRH

[Smi00a] Roger Smith. Book reviews: H. F. Augstein (ed.), Race: The Origins of an Idea, 1760– 1850. Key Issues, 14. Bristol: Thoemmes Press, 1996. Pp. xxxiii + 260. ISBN 1-85506-455-3, £45.00 (hardback); 1-85506-454-5, £14.95 (paperback). Graham Richards, ‘Race’, Racism and Psychology: Towards a Reflexive History. London: Routledge, 1997, Pp. xvii+372. ISBN 0-415-10140-7, £50.00 (hardback); 0-415-10141-7, £10.99 (paperback). British Journal for the History of Science, 33 (1):115–124, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028073.

Smith:2000:EVC

[Smi00b] Roger Smith. The embodiment of value: C. S. Sherrington and the cultivation of science. British Journal for the History of Science, 33(3):283–311, September 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4027954.

Smil:2001:EEF

[Smi01a] Vaclav Smil. Enriching the earth:

Fritz Haber, Carl Bosch, and the transformation of world food production. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-19449-X (hardcover). xvii + 338 pp. LCCN S651 .S56 2001.

Smith:2001:BRR

[Smi01b] Roger Smith. Book review:

Rick Rylance, Victorian Psychology and British Culture 1850– 1880. Oxford: Oxford University Press, 2000. Pp. x + 355. ISBN 0-19-812283-7. £45.00 (hardback). British Journal for the History of Science, 34(3): 341–373, September 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028119.

Smith:2002:BRJ

[Smi02a] A. Mark Smith. Book review: Johannes Kepler, Optics. Paralipomena to Witelo & Optical Part of Astronomy. English translation by William H. Donahue. Sante Fe: Green Lion Press, 2000. Pp. xv + 459. ISBN 1-888009-12-8. $55.00. British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028204.

Smith:2002:BRL

[Smi02b] Pamela H. Smith. Book review: Lyndy Abraham, A Dictionary of Alchemical Imagery. Cambridge: Cambridge University Press, 1998. Pp. xxii + 249. ISBN 0-521-00000-9. £16.95, $24.95 (paperback). Hans-Werner Schütt, Auf der Suche nach dem Stein der Weisen: Die Geschichte der Alchemie. München: C. H. Beck, 2000. Pp. 602. ISBN 3-406-46638-9. DM 68.50 (hardback). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2003:BRJ

[Smi03a] David F. Smith. Book review: Jeannette Ewin, Fine Wines and Fish Oils: The Life of Hugh Macdonald Sinclair. Oxford: Oxford University Press, 2001. Pp. xx + 338. ISBN 0-19-262927-1. £25.00 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028323.

Smith:2003:BRG

[Smi03b] Roger Smith. Book review:

Graham Richards, Putting Psychology in its Place: A Critical Historical Overview. Second edition. New York: Routledge and Hove: Psychology Press, 2002. Pp. xiv + 368. ISBN 1-84169-234-4. £16.60 (paperback). British Journal for the History of Science, 36(3): 374–375, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028167.

Smith:2004:BRR

[Smi04] Pamela H. Smith. Book review:

Rudolph Gamper and Thomas Hofmeier, Das Alchemiehandbuch des Appenzeller Wundarztes Ulrich Ruosch. Basel: Schwabe, 2002. Pp. 158. ISBN 3-7965-1975-X. EUR33.50, SFr 48.00 (hardback). British Journal for the History of Science, 37(3):348–349, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.

                                        jstor.org/stable/4028129;
                                        http://www.jstor.org/
                                        stable/4028431.

Smith:2005:BRR

[Smi05a] David Smith. Book review: Roger Cooter and John Pickstone (eds.), Companion to Medicine in the Twentieth Century. Routledge World Reference. London and New York: Routledge, 2003. Pp. xix + 756. ISBN 0-415-28603-4. £24.99 (paperback). British Journal for the History of Science, 38(4): 490–491, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028472.

Smith:2005:BRS

[Smi05b] Pamela H. Smith. Book review: Stanton J. Linden (ed.), The Alchemy Reader: From Hermes Trismegistus to Isaac Newton. Cambridge: Cambridge University Press, 2003. Pp. xxv + 260. ISBN 0-521-79662-8. £16.95, $24.00 (paperback). British Journal for the History of Science, 38(3):363, September 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028684.

Smith:2006:CDV

[Smi06a] Jonathan Smith. Charles Darwin and Victorian visual culture. Cambridge studies in nineteenthcentury literature and culture. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85690-6 (hardcover). xxiii + 349 pp. LCCN QH46.5 .S65

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0703/2006283483-d.html;
http://www.loc.gov/catdir/
enhancements/fy0703/
2006283483-t.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2006283483-b.html.

Smith:2006:PAG

[Smi06b] Justin E. H. Smith, editor. The problem of animal generation in early modern philosophy. Cambridge studies in philosophy and biology. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-84077-5 (hardcover). xiii + 456 pp. LCCN QP251 .P76

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0633/2005021617-d.html;
http://www.loc.gov/catdir/
enhancements/fy0733/
2005021617-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0518/2005021617.html.

Smith:2006:BRV

[Smi06c] Pamela H. Smith. Book review:

Volker Fritz Brüning, Bibliographie der alchemistischen Literatur. Band 1: Die alchemistischen Druckwerke von der Erfindung der Buchdruckerkunst bis zum Jahr 1690. München: K. G. Saur, 2004. Pp. xii + 500. ISBN 3-598-11603-9. EUR248.00 (hardback). British Journal for the History of Science, 39 (2):289–290, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028616.

Smith:2006:BRJ

[Smi06d] Roger Smith. Book review: John van Wyhe, Phrenology and the Origins of Victorian Scientific Naturalism. Science, Technology and Culture, 1700–1945. Aldershot: Ashgate, 2004. Pp. xvii + 282. ISBN 0-7546-3408-6. £49.50 (hardback). British Journal for the History of Science, 39(3):456–458, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028499.

Smith:2006:BRR

[Smi06e] Roger Smith. Book review: Roy Porter, Flesh in the Age of Reason: How the Enlightenment Transformed the Way We See Our Bodies and Souls. London: Penguin, 2004. Pp. xviii + 574. ISBN 0-140-16735-8. £9.99 (paperback). British Journal for the History of Science, 39(4): 609–610, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028531.

Smith:2006:BRS

[Smi06f] Roger Smith. Book review: Sonu Shamdasani, Jung and the Making of Modern Psychology: The Dream of a Science. Cambridge: Cambridge University Press, 2003. Pp. xvi + 387. ISBN 0-521-53909-9. £18.95, $28.00 (paperback). British Journal for the History of Science, 39 (2):310–312, June 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028634.

Smith:2006:BRT

[Smi06g] Roger Smith. Book review:

Thomas Dixon, From Passions to Emotions: The Creation of a Secular Psychological Category. Cambridge: Cambridge University Press, 2003. Pp. x + 287. ISBN 0-521-82729-9. £47.50, $60.00 (hardback). British Journal for the History of Science, 39(1):137–138, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028563.

Smith:2007:BRJ

[Smi07a] Robert W. Smith. Book reviews: Jane Gregory, Fred Hoyle’s Universe. Oxford: Oxford University Press, 2005. Pp. x + 406. ISBN 0-19-850791-7. £20.00 (hardback). Simon Mitton, Fred Hoyle: A Life in Science. With an Introduction by Paul Davies. London: Aurum, 2005. Pp. xi+369. ISBN 1-85410-961-8. £18.99 (hardback). British Journal for the History of Science, 40(1):152–153, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500710.

Smith:2007:BRM

[Smi07b] Roger Smith. Book review: Margaret Schabas, The Natural Origins of Economics. Chicago and London: University of Chicago Press, 2005. Pp. xi + 231. ISBN 0-226-73569-9. $40.00, £25.50 (hardback). British Journal for the History of Science, 40(3): 445–446, September 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500754.

Smith:2008:BRD

[Smi08a]  Crosbie  Smith.

Book review: David N. Livingstone and Charles W. J. Withers (eds.), Geography and Revolution. Chicago and London: The University of Chicago Press, 2005. Pp. viii + 433. ISBN 0-226-48733-4. £45.00 (hardback). British Journal for the History of Science, 41(1):117–119, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:BRA

[Smi08b] Pamela H. Smith. Book review: Adam Mosley, Bearing the Heavens: Tycho Brahe and the Astronomical Community of the Late Sixteenth Century. Cambridge: Cambridge University Press, 2007. Pp. xiv + 354. ISBN 978-0-521-83866-5. £55.00 (hardback). British Journal for the History of Science, 41(4): 605–606, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:BRS

[Smi08c] Pamela H. Smith. Book review: Sachiko Kusukawa and Ian Maclean (eds.) Transmitting Knowledge: Words, Images, and Instruments in Early Modern Europe. Oxford: Oxford University Press, 2006. Pp. xvi + 274. ISBN 0-19-928878-X. £100.00 (hardback). British Journal for the History of Science, 41(1): 137–138, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:BRG

[Smi08d] Roger Smith. Book review:

Gregory Radick, The Simian Tongue: The Long Debate about Animal Language. Chicago and London: University of Chicago Press, 2007. Pp. xiv + 577. ISBN 978-0-226-70224-7. $45.00, £23.50. British Journal for the History of Science, 41(3): 449–450, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:BRR

[Smi08e] Roger Smith. Book review:

Roberta Bivins and John V. Pickstone (eds.) Medicine, Madness and Social History: Essays in Honour of Roy Porter. Basingstoke and New York: Palgrave Macmillan, 2007. Pp. x + 295. ISBN 978-0-203-52549-8.

£55.00 (hardback). British Journal for the History of Science, 41(2):281–282, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:BRP

[Smi08f] Roger Smith. Book reviews: Peter Becker and Richard F. Wetzell (eds.), Criminals and Their Scientists: The History of Criminology in International Perspective. Cambridge: Cambridge University Press and Washington, DC: German Historical Institute, 2006. Pp. xiii + 492. ISBN 978-0-521-81012-8. £60.00, $85.00 (hardback). Cesare Lombroso, Criminal Man. Translated and with a new Introduction by Mary Gibson and Nicole Hahn Rafter. Durham and London: Duke University Press, 2006. Pp. xvii+424. ISBN 0-8223-3723-1. £15.95 (paperback). British Journal for the History of Science, 41(4):619–621, December 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:ORMa

[Smi08g] Russell Smith. Optical reflection and mechanical rebound: the shift from analogy to axiomatization in the seventeenth century. Part 1. British Journal for the History of Science, 41(1):1– 18, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2008:ORMb

[Smi08h] Russell Smith. Optical reflection and mechanical rebound: the shift from analogy to axiomatization in the seventeenth century. Part 2. British Journal for the History of Science, 41(2):187– 207, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2009:BRL

[Smi09a] Pamela H. Smith. Book reviews: Lawrence M. Principe (ed.), Chymists and Chymistry: Studies in the History of Alchemy and Early Modern Chemistry. Sagamore Beach, MA: Science History Publications/USA, 2007. Pp. xiii + 274. ISBN 978-0-88135-396-9. $45.00 (hardback). Anna Marie Roos, The Salt of the Earth: Natural Philosophy, Medicine, and Chymistry in England, 1650– 1750. Leiden and Boston: Brill, 2007. Pp. xvi+293. ISBN 978-90-04-16176-4. $129.00 (hardback). British Journal for the History of Science, 42(1):130–132, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smith:2009:BRG

[Smi09b] Roger Smith. Book reviews: G.

E. R. Lloyd, Cognitive Variations: Reflections on the Unity and Diversity of the Human Mind. Oxford: Clarendon Press, 2007. Pp. viii + 201. ISBN 978-0-19-921461-7. £27.50 (hardback). 2009. ISBN 978-0-19-956625-9.

£14.99 (paperback). British Journal for the History of Science, 42(3):448–450, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Smocovitis:2008:BRG

[Smo08] Vassiliki Betty Smocovitis. Book review: Geoffrey C. Bowker, Memory Practices in the Sciences. Cambridge, MA and London: MIT Press, 2006. Pp. xi + 261. ISBN 0-262-02589-2. £22.95 (hardback). British Journal for the History of Science, 41 (1):154–155, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Snowden:2006:CMI

[Sno06] Frank M. (Frank Martin) Snowden. The conquest of malaria: Italy, 1900–1962. Yale University Press, New Haven, CT, USA, 2006. ISBN 0-300-10899-0. viii + 296 pp. LCCN RC163.I8 S56

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0643/2005013483-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0643/
                                        2005013483-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0512/2005013483.html.

Snyder:2006:RPV

[Sny06] Laura J. Snyder. Reforming philosophy: a Victorian debate on science and society. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-76733-7 (hardcover). x + 386 pp. LCCN B1607 .S69

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005035508-b.html;
http://www.loc.gov/catdir/
enhancements/fy0623/
2005035508-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip064/2005035508.html.

Soderqvist:2007:HPS

[Söd07] Thomas Söderqvist, editor. The History and Poetics of Scientific Biography. Science, technology and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2007. ISBN 0-7546-5181-9. xiv + 270 pp. LCCN Q141 .H575 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0620/2006029937.html.

Sokal:2008:BHS

[Sok08] Alan D. Sokal. Beyond the hoax:

science, philosophy and culture. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-923920-7 (hardcover). xxi + 465 pp. LCCN Q175 S6395 2008. URL http:

//www.loc.gov/catdir/toc/
fy0804/2008297511.html.

Sommer:2004:AAC

[Som04] Marianne Sommer. ‘an amusing account of a cave in Wales’: William Buckland (1784–1856) and the Red Lady of Paviland. British Journal for the History of Science, 37(1):53–74, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028256.

Sommer:2007:BOC

[Som07] Marianne Sommer. Bones and ochre: the curious afterlife of the Red Lady of Paviland. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02499-0. xii + 398 pp. LCCN GN286.3 .S66 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0719/2007021357.html.

Spary:2000:UGF

[Spa00] E. C. (Emma C.) Spary. Utopia’s garden: French natural history from Old Regime to Revolution. University of Chicago Press, Chicago, IL, USA, 2000. ISBN 0-226-76862-7 (cloth), 0-226-76863-5 (paper). xv + 321 pp. LCCN QH147 .S62

                                        2000. URL http://www.loc.
                                        gov/catdir/bios/uchi051/
                                        00029897.html;
                                        http://www.loc.gov/catdir/
                                        description/uchi051/
                                        00029897.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0610/
                                        00029897-t.html.

Spary:2002:BRJ

[Spa02] E. C. Spary. Book review: Jeff Loveland, Rhetoric and Natural History: Buffon in Polemical and Literary Context. Studies on Voltaire and the Eighteenth Century, 2001[ratio]03. Oxford: Voltaire Foundation, 2001. Pp. ix + 214. ISBN 0-7294-0747-0. £45.00, $67.50 (paperback).

British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028205.

Spary:2003:BRL

[Spa03] E. C. Spary. Book review:

Louise E. Robbins, Elephant Slaves and Pampered Parrots: Exotic Animals in Eighteenth-Century Paris. Animals, History, Culture. Baltimore and London: Johns Hopkins University Press, 2002. Pp. xiv + 349. ISBN 0-8018-6753-3. £33.00 (hardback). British Journal for the History of Science, 36(3): 367–368, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028161.

Spary:2005:BRR

[Spa05] Emma Spary. Book review: Roger L. Williams, French Botany in the Enlightenment: The Illfated Voyages of La Pérouse and His Rescuers. International Archives of the History of Ideas,

182. Dordrecht, Boston and London: Kluwer Academic Publishers, 2003. Pp. 240. ISBN 1-4020-1109-1. £57.00, $85.00 (hardback). British Journal for the History of Science, 38(3):369– 371, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028688.

Spary:2006:BRS

[Spa06] Emma Spary. Book review: Steven B. Karch, A History of Cocaine: The Mystery of Coca Java and the Kew Plant. London: Royal Society of Medicine Press, 2003. Pp. xi + 224. ISBN 1-85315-547-0. £24.95, $39.95 (paperback). British Journal for the History of Science, 39 (1):143–144, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028568.

Sponsel:2002:CRS

[Spo02] Alistair Sponsel. Constructing a “revolution in science”: the campaign to promote a favourable reception for the 1919 solar eclipse experiments. British Journal for the History of Science, 35(4): 439–467, December 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028276.

Spranzi:2006:BRF

[Spr06] Marta Spranzi. Book review: Fernand Hallyn, Les Structures rhétoriques de la science. De Kepler à Maxwell. Collection ≪Des travaux≫. Paris: Éditions du Seuil, 2004. Pp. 323. ISBN 2-02-063249-7. EUR24.00. British Journal for the History of Science, 39(2):285–286, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028613.

Selin:2000:AAC

[SS00] Helaine Selin and Xiaochun Sun, editors. Astronomy across cultures: the history of non-Western astronomy, volume 1 of Science across cultures. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6363-9. xxiii + 665 pp. LCCN QB16 .A75 2000.

Siegmund-Schultze:2001:RIM

[SS01] R. (Reinhard) Siegmund-Schultze. Rockefeller and the internationalization of mathematics between the two World Wars: documents and studies for the social history of mathematics in the 20th Century, volume 25 of Science networks historical studies. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 3-7643-6468-8, 0-8176-6468-8. xiii + 341 pp. LCCN QA13 .S54

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0812/2001025227-d.html.

Schiebinger:2005:CBS

[SS05] Londa L. Schiebinger and Claudia Swan, editors. Colonial botany: science, commerce, and politics in the early modern world. University of Pennsylvania Press, Philadelphia, 2005. ISBN 0-8122-3827-3 (cloth). vi + 346 pp. LCCN SB108.E85 C64

Staubermann:2001:MSP

[Sta01a] Klaus B. Staubermann. Making stars: projection culture in nineteenth-century German astronomy. British Journal for the History of Science, 34(4): 439–451, December 2001. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028373.

Staubermann:2001:TKS

[Sta01b] Klaus B. Staubermann. Tying the knot: skill, judgement and authority in the 1870s Leipzig spiritistic experiments. British Journal for the History of Science, 34(1):67–79, March 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028039.

Staley:2003:BRI

[Sta03] Richard Staley. Book reviews: Interdisciplinary atomism? Exploring twentieth-century culture through Einstein Marcia Bartusiak, Einstein’s Unfinished Symphony: Listening to the Sounds of Space–Time. Washington, DC: Joseph Henry Press, 2000. Pp. xii + 249. ISBN 0-309-06987-4. £17.95 (hardback). Alice Calaprice (ed.), The Expanded Quotable Einstein. With a foreword by Freeman Dyson. Princeton and Oxford: Princeton University Press, 2000. Pp. xliii+407. ISBN 0-691-07021-0. £11.95, $18.95 (cloth). Klaus Hentschel (ed.), Physics and National Socialism: An Anthology of Primary Sources. Ann M. Hentschel, Editorial Assistant and Translator. Erwin Hiebert and Hans Wussing (eds.), Science Networks: Historical Studies, 18. Basel, Boston and Berlin: Birkhäuser, 1996. Pp. ci+406+civ. ISBN 3-7643-5312-0. DM 178.00, SFR 148.00, EUR98.00 (cloth). Gerald Holton, Einstein, History, and Other Passions: The Rebellion Against Science at the End of the Twentieth Century. Cambridge and London: Harvard University Press, 2000. Pp. xii+240. ISBN 0-674-00433-7. £12.50 (paperback). Don Howard and John Stachel (eds.), Einstein: The Formative Years, 1879–1909. Einstein Studies, 8. Boston, Basel and Berlin: Birkhäuser, 2000. Pp. xi+258. ISBN 0-8176-4030-4. DM 128.00, EUR83.00, SFR 124.00. Arthur I. Miller, Einstein, Picasso: Space, Time and the Beauty that Causes Havoc. New York: Basic Books, 2001. Pp. x+357. ISBN 0-465-01859-9. $30.00, CAN$44.95 (cloth). British Journal for the History of Science, 36(2):221–230, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028234.

Staley:2004:BRM

[Sta04]  Richard  Sta-

ley. Book review: Michael Eckert and Karl Märker (eds.), Arnold Sommerfeld: Wissenschaftlicher Briefwechsel. Band 1: 1892–1918. Berlin, Diepholz and München: Deutsches Museum and Verlag für Geschichte der Naturwissenschaften und der Technik, 2000. Pp. 694. ISBN 3-928186-49-3. EUR68.00 (hardback). British Journal for the History of Science, 37(3):360– 361, September 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028140;
                                        http://www.jstor.org/
                                        stable/4028442.

Stanford:2006:EOG

[Sta06a] P. Kyle Stanford. Exceeding our grasp: science, history, and the problem of unconceived alternatives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517408-9 (cloth). xiv + 234 pp. LCCN Q175.32.I54 S73

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0637/2005051299-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0724/
                                        2005051299-b.html;
                                        http://www.
                                        oxfordscholarship.com/oso/
                                        public/content/philosophy/
                                        0195174089/toc.html.

Stanley:2006:BRI

[Sta06b] Matthew Stanley. Book review: Iwan Rhys Morus, When Physics Became King. Chicago and London: University of Chicago Press, 2005. Pp. xii + 303. ISBN 0-226-54202-5. $25.00 (paper). British Journal for the History of Science, 39(4):610– 611, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028532.

Stanley:2006:BRM

[Sta06c] Matthew Stanley. Book review: Mary Jo Nye (ed.), The Cambridge History of Science. Volume 5: The Modern Physical and Mathematical Sciences. Cambridge: Cambridge University Press, 2003. Pp. xxix + 678. ISBN 0-521-57199-5. £65.00, $95.00 (hardback). British Journal for the History of Science, 39(1): 123–125, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028552.

Staley:2007:BRA

[Sta07a] Kent W. Staley. Book review:

Allan Franklin, No Easy Answers: Science and the Pursuit of Knowledge. Pittsburgh: University of Pittsburgh Press, 2005. Pp. xii + 258. $29.95 (hardback). Michela Massimi, Pauli’s Exclusion Principle: The Origin and Validation of a Scientific Principle. Cambridge: Cambridge University Press, 2005. Pp. xiv+211. $80.00 (hardback). British Journal for the History of Science, 40(3):455–457, September 2007.

CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500761.

Stanley:2007:BRA

[Sta07b] Matthew Stanley. Book review: Arthur I. Miller, Empire of the Stars: Obsession, Friendship, and Betrayal in the Quest for Black Holes. Boston and New York: Houghton Mifflin, 2005. Pp xx + 364. ISBN 0-618-34151-X. $26.00 (hardback). British Journal for the History of Science, 40(3):457–458, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500762.

Stanley:2007:BRG

[Sta07c] Matthew Stanley. Book review: Geoffrey Cantor, Quakers, Jews, and Science: Religious Responses to Modernity and the Sciences in Britain, 1650–1900. Oxford: Oxford University Press, 2005. Pp. xii + 420. ISBN 0-19-927668-4. £50.00 (hardback). British Journal for the History of Science, 40(1): 120–121, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4500689.

Stanley:2007:BRR

[Sta07d] Matthew Stanley. Book review: Robert Fox and Graeme Gooday (eds.), Physics in Oxford, 1839–1939: Laboratories, Learning, and College Life. Oxford: Oxford University Press, 2005. Pp. vii + 363. ISBN 0-19-856792-8. £60.00 (hardback). British Journal for the History of Science, 40(2):298–299, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500735.

Stanley:2007:PMR

[Sta07e] Matthew Stanley. Practical mystic: religion, science, and A. S. Eddington. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-77097-4 (cloth). x + 313 pp. LCCN BL240.3 .S725

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0809/2007005482-d.html;
http://www.loc.gov/catdir/
enhancements/fy0809/
2007005482-t.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007005482-b.html.

Stanley:2007:STS

[Sta07f] Matthew Stanley. So simple a thing as a star: the Eddington– Jeans debate over astrophysical phenomenology. British Journal for the History of Science, 40(1):53–82, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500682.

Staley:2008:EGO

[Sta08a] Richard Staley, editor. Einstein’s generation: the origins of the relativity revolution. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-77056-7 (cloth), 0-226-77057-5 (paperback). LCCN QC7 .S777 2008.

                                        URL  http://www.loc.gov/
                                        catdir/enhancements/fy0829/
                                        2008014155-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0815/2008014155.html.

Stanford:2008:BRH

[Sta08b] P. Kyle Stanford. Book review: Hanne Andersen, Peter Barker and Xiang Chen, The Cognitive Structure of Scientific Revolutions. Cambridge: Cambridge University Press, 2006. Pp. xvii + 199. ISBN 978-0-521-85575-4. £45.00 (hardback). British Journal for the History of Science, 41 (1):116–117, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Staley:2009:BRS

[Sta09a] Kent W. Staley. Book review: Steven French and Décio Krause, Identity in Physics: A Historical, Philosophical, and Formal Analysis. New York: Oxford University Press, 2006. Pp. xv + 422. ISBN 0-19-927824-5. £55.00 (hardback). British Journal for the History of Science, 42(1): 145–146, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stanley:2009:BRR

[Sta09b] Matthew Stanley. Book review: Richard Staley, Einstein’s Generation: The Origin of the Relativity Revolution. Chicago and London: University of Chicago Press, 2009. Pp. x + 494. ISBN 978-0-226-77057-4. £26.00 (paperback). British Journal for the History of Science, 42(3): 470–471, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Steele:2000:OPE

[Ste00a] John M. Steele. Observations and predictions of eclipse times by early astronomers, volume 4 of Archimedes. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6298-5. xi + 321 pp. LCCN QB175 .S83

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0813/00037058-d.html;
http://www.loc.gov/catdir/
enhancements/fy0813/
00037058-t.html.

Stevenson:2000:MMB

[Ste00b] Christine Stevenson. Medicine and magnificence: British hospital and asylum architecture, 1660–1815. Yale University Press, New Haven, CT, USA, 2000. ISBN 0-300-08536-2 (cloth). viii + 312 pp. LCCN RA967 .S725 2000. URL http:/

/www.loc.gov/catdir/toc/
fy031/00043489.html.

Stepan:2001:PTN

[Ste01] Nancy Leys Stepan. Picturing tropical nature. Picturing history. Reaktion Books, London, UK, 2001. ISBN 1-86189-084-2. 283 pp. LCCN BD581 .S719 2001. US$25.00.

URL http://www.gbv.de/dms/
goettingen/321114647.pdf.

Stedall:2004:BRH

[Ste04a] Jackie Stedall. Book review: H. L. L. Busard, Johannes de Tinemue’s Redaction of Euclid’s Elements, the So-called Adelard III Version. Volume I: Introduction, Sigla and Descriptions of the Manuscripts, Editorial Remarks, Euclides, Elementa. Volume II: Conspectus Siglorum, Apparatus Criticus, Addenda. Boethius, 45. Stuttgart: Franz Steiner Verlag, 2001. Pp. 632. ISBN 3-515-07975-0. EUR100.00 (paperback). British Journal for the History of Science, 37(4): 468–469, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028646.

Steinle:2004:BRE

[Ste04b] Friedrich Steinle. Book review: Ernest B. Hook (ed.), Prematurity in Scientific Discovery: On Resistance and Neglect. Berkeley, Los Angeles and London: University of California Press, 2002. Pp. xx + 378. ISBN 0-520-23106-6. £55.00, $80.00 (hardback). British Journal for the History of Science, 37 (2):235–236, June 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028362.

Stedall:2005:BRA

[Ste05a] Jackie Stedall. Book review:

Amir R. Alexander, Geometrical Landscapes: The Voyages of Discovery and the Transformation of Mathematical Practice. Writing Science. Stanford: Stanford University Press, 2002. Pp. xvii + 293. ISBN 0-80473-260-4. £46.95 (hardback). British Journal for the History of Science, 38 (1):108–109, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028590.

Stedall:2005:BRM

[Ste05b] Jackie Stedall. Book review:

Menso Folkerts, Essays on Early Medieval Mathematics: The Latin Tradition. Variorum Collected Studies Series, CS751. Aldershot: Ashgate, 2003. Pp.

382. ISBN 0-86078-895-4. £59.50 (hardback). British Journal for the History of Science, 38(3): 358–359, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028679.

Stenhouse:2006:BRS

[Ste06] John Stenhouse. Book review:

Sujit Sivasundaram, Nature and the Godly Empire: Science and Evangelical Mission in the Pacific, 1795–1850. Cambridge Social and Cultural Histories. Cambridge and New York: Cambridge University Press, 2005. Pp. xii + 244. ISBN 0-521-84836-9. £48.00 (hardback). British Journal for the History of Science, 39(4):593–595, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028520.

Steigerwald:2007:BRP

[Ste07] Joan Steigerwald. Book review: Peter Hanns Reill, Vitalizing Nature in the Enlightenment. Berkeley, Los Angeles and London: University of California Press, 2005. Pp. x + 388. ISBN: 0-520-24135-5. $55.00, £35.95 (hardback). British Journal for the History of Science, 40 (2):292–293, June 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500730.

Stegenga:2009:BRA

[Ste09a] Jacob Stegenga. Book review: Angela N. H. Creager, Elizabeth Lunbeck and M. Norton Wise (eds.), Science without Laws: Model Systems, Cases, Exemplary Narratives. Durham, NC and London: Duke University Press, 2007. ISBN 978-0-8223-4068-3. £12.99 (paperback). British Journal for the History of Science, 42(4):626– 628, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stegenga:2009:BRJ

[Ste09b] Jacob Stegenga. Book review:

Jessica Riskin (ed.), Genesis Redux: Essays in the History and Philosophy of Artificial Life. Chicago and London: University of Chicago Press, 2007. Pp. xvii + 389. ISBN 978-0-226-72081-4. £16.00, $25.00 (paperback). British Journal for the History of Science, 42(3):437–438, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stewart:2009:SCB

[Ste09c] John Stewart. The scientific claims of British child guidance, 1918–1945. British Journal for the History of Science, 42(3): 407–432, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Stokes:2000:CST

[Sto00] Raymond G. Stokes. Constructing socialism: technology and change in East Germany 1945– 1990. Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 2000. ISBN 0-8018-6391-0 (hardcover). xi + 260 pp. LCCN HC290.795.T4 S76 2000. URL

http://www.loc.gov/catdir/
bios/jhu051/00008336.html;
http://www.loc.gov/catdir/
description/jhu051/
00008336.html.

Strick:2000:SLD

[Str00] James Edgar Strick. Sparks of life: Darwinism and the Victorian debates over spontaneous generation. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00292-X (cloth). xi + 283 pp. LCCN QH325 .S85 2000.

Sturdy:2006:BRC

[Stu06] Steve Sturdy. Book review: Chris Feudtner, Bitter Sweet: Diabetes, Insulin, and the Transformation of Illness. Studies in Social Medicine. Chapel Hill and London: University of North Caroline Press, 2003. Pp. xxiiii + 290. ISBN 0-8087-2791-6. £22.95, $29.95 (hardback). British Journal for the History of Science, 39(2):309–310, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028633.

Sturdy:2009:BRM

[Stu09] Steve Sturdy. Book reviews: Mark Jackson, Allergy: The History of a Modern Malady. London: Reaktion Books, 2007. Pp. 288. ISBN 978-1-86189-333-8. £12.25 (paperback). Gregg Mitman, Breathing Space: How Allergies Shape Our Lives and Landscapes. London: Yale University Press, 2007. Pp. xv + 312. ISBN 978-0-300-11035-7. £20.00 (hardback). British Journal for the History of Science, 42(2):282–284, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sufian:2007:HLN

[Suf07] Sandra M. (Sandra Marlene) Sufian. Healing the land and the nation: malaria and the Zionist project in Palestine, 1920– 1947. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-77935-1 (cloth). xviii + 385 pp. LCCN RA644.M2 S84

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0732/2007015180-d.html;
http://www.loc.gov/catdir/
enhancements/fy0827/
2007015180-b.html;   http:/
/www.loc.gov/catdir/toc/
ecip0715/2007015180.html.

Sumner:2001:JRS

[Sum01] James Sumner. John Richardson, saccharometry and the poundsper-barrel extract: the construction of a quantity. British Journal for the History of Science, 34(3):255–273, September 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028098.

Sumner:2005:BRW

[Sum05] James Sumner. Book review:

William J. Ashworth, Customs and Excise: Trade, Production and Consumption in England 1640–1845. Oxford: Oxford University Press, 2003. Pp. xiii + 396. ISBN 0-19-925921-6. £55.00 (hardback). British Journal for the History of Science, 38(4): 480–481, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028465.

Sumner:2006:BRA

[Sum06] James Sumner. Book review: Alex Roland with Philip Shiman, Strategic Computing: DARPA and the Quest for Machine Intelligence, 1983–1993. History of Computing. Cambridge, MA and London: MIT Press, 2002. Pp. xxvi + 427. ISBN 0-262-18226-2. £33.50 (hardback). British Journal for the History of Science, 39(4):622–624, December 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028541.

Sutton:2003:BRD

[Sut03] John Sutton. Book review: Dennis Des Chene, Spirits and Clocks: Machine and Organism in Descartes. Ithaca and London: Cornell University Press, 2001. Pp. xiii + 181. ISBN 0-8014-3764-4. £25.95 (hardback). British Journal for the History of Science, 36(2):233–235, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028236.

Suzuki:2005:BRR

[Suz05]  Akihito Suzuki.  Book re-

view: Roy Porter, Madness: A Brief History. Oxford: Oxford University Press, 2002. Pp. xii + 241. ISBN 0-19-280266-6. £11.99, $22.00 (hardback). British Journal for the History of Science, 38(1):120–121, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028600.

Suzuki:2006:MHP

[Suz06] Akihito Suzuki. Madness at home: the psychiatrist, the patient, and the family in England, 1820–1860, volume 13 of Medicine and society. University of California Press, Berkeley, CA, USA, 2006. ISBN 0-520-24580-6 (cloth). xii + 259 pp. LCCN RC450.G7 S89

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0622/2005010547-b.html;
http://www.loc.gov/catdir/
enhancements/fy0641/
2005010547-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0511/2005010547.html.

Sweeney:2001:FGC

[Swe01] Gerald Sweeney. Fighting for the good cause: reflections on Francis Galton’s legacy to American hereditarian psychology, volume 91, pt. 2 of Transactions of the American Philosophical Society, 0065-9746. American Philosophical Society, Philadelphia, PA, USA, 2001. ISBN 0-87169-912-5 (paperback). x + 136 pp. LCCN BF701 .S93 2001.

Sweet:2004:ADP

[Swe04]  Rosemary Sweet.  Antiquar-

ies: the discovery of the past in Eighteenth-Century Britain. Hambledon and London, London, UK, 2004. ISBN 1-85285-309-3. xxi + 473 + 24 pp. LCCN DA129.5 .S94 2004. URL http:

                                        //www.loc.gov/catdir/bios/
                                        hol055/2004478649.html;
                                        http://www.loc.gov/catdir/
                                        description/hol054/
                                        2004478649.html.

Taub:2008:AME

[Tau08] Liba Chaia Taub. Aetna and the moon: explaining nature in ancient Greece and Rome. The OSU Press Horning Visiting Scholars publication series. Oregon State University Press, Corvallis, OR, USA, 2008. ISBN 0-87071-196-2. xiv + 138 pp. LCCN Q223 .T38

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0808/2007039213-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0808/
                                        2007039213-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0727/2007039213.html.

Taylor:2004:VLN

[Tay04] William M. Taylor. The vital landscape: nature and the built environment in Nineteenth-Century Britain. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-7546-3069-2. xvii + 252 pp. LCCN SB451.36.G7 T38 2004.

Taylor:2006:UAW

[Tay06] Georgette Taylor. Unification achieved: William Cullen’s theory of heat and phlogiston as an example of his philosophical chemistry. British Journal for the History of Science, 39(4): 477–501, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Thornton:1990:TMB

[TBbTc90] John Leonard Thornton, Alain Besson, John Leonard Medical books Thornton, libraries, and collectors. Thornton’s medical books, libraries, and collectors /a study of bibliography and the book trade in relation to the medical sciences. Gower Pub. Co., Aldershot, Hampshire, England, 3rd rev. edition, 1990. ISBN 0-566-05481-7. xxi + 417 pp. LCCN Z286.M4 T47 1990.

Turchetti:2008:AOH

[TDNS08] Simone Turchetti, Katrina Dean, Simon Naylor, and Martin Siegert. Accidents and opportunities: a history of the radio echo-sounding of Antarctica, 1958–1979. British Journal for the History of Science, 41(3): 417–444, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Teich:2000:BWW

[Tei00]  Mikuláš Teich.  Bier, Wis-
senschaft  und  Wirtschaft

in Deutschland 1800–1914: ein Beitrag zur deutschen Industrialisierungsgeschichte. (German) [Beer, science and economics in Germany 1800–1914: a contribution to the history of German industrialization]. Böhlau, Wien, Austria, 2000. ISBN 3-205-99239-3. 353 pp. LCCN TP577 .T42 2000.

Terrall:2008:BRJ

[Ter08] Mary Terrall. Book review: Judith P. Zinsser and Julie Candler Hayes (eds.) Emilie du Châtelet: Rewriting Enlightenment Philosophy and Science. Oxford: Voltaire Foundation, 2006. Pp. xi + 325. ISBN 0-7294-0872-8. £60.00, EUR99.00, $122.00 (paperback). British Journal for the History of Science, 41(1):141– 142, March 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Fan:2000:BRM

[tF00] Fa ti Fan. Book review: Masao Watanabe, Science and Cultural Exchange in Modern History: Japan and the West. Tokyo: Hokusen-Sha, 1997. Pp. xi + 406. ISBN 4-938424-73-8. No price given. British Journal for the History of Science, 33 (2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027931.

Tyrrell:2002:CWF

[TF02] D. A. J. (David Arthur John) Tyrrell and Michael Fielder. Cold wars: the fight against the common cold. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-263285-X. xiv + 253 pp. LCCN RC150.1 .T975

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0613/2002074887-d.html;
http://www.loc.gov/catdir/
toc/fy038/2002074887.html.

Fan:2004:BNQ

[tF04] Fa ti Fan. British naturalists in Qing China: science, empire, and cultural encounter. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01143-0. xi + 238 pp. LCCN QH51 .F36 2004. URL http://

www.loc.gov/catdir/toc/
fy045/2003056554.html.

teHeesen:2002:WBS

[tH02] Anke te Heesen. The world in a box: the story of an eighteenth-century picture encyclopedia. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-32286-6 (cloth), 0-226-32287-4 (paperback). xii + 237 pp. LCCN AG27.W433 H44

2002. URL http://www.h-net.
org/review/hrev-a0d0u3-aa;
http://www.loc.gov/catdir/
bios/uchi051/2001052763.
html;   http://www.loc.gov/
catdir/description/uchi051/
2001052763.html;
http://www.loc.gov/catdir/
toc/fy031/2001052763.html.

Thagard:1999:HSE

[Tha99]  Paul Thagard.  How scien-

tists explain disease. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00261-4 (cloth). xviii + 263 pp. LCCN RB151 .T47

1999. URL http://www.loc.
gov/catdir/description/
prin021/98034717.html;
http://www.loc.gov/catdir/
toc/prin032/98034717.html.

Thompson:2002:SMA

[Tho02a] Emily Ann Thompson. The soundscape of modernity: architectural acoustics and the culture of listening in America, 1900–1933. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-20138-0 (hardcover). ix + 500 pp. LCCN NA2800 T48

2002. URL http://www.loc.
gov/catdir/toc/fy0705/
2001044533.html.

Thorpe:2002:BRJ

[Tho02b] Charles Thorpe. Book review: John S. Rigden, Rabi: Scientist and Citizen. Cambridge, MA and London: Harvard University Press, 2000. Pp. xxix + 302. ISBN 0-674-00435-3. £13.95 (paperback). British Journal for the History of Science, 35 (2):213–250, June 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028206.

Thomson:2003:BRD

[Tho03a] Mathew Thomson. Book review: David Wright, Mental Disability in Victorian England: The Earlswood Asylum 1847–1901. Oxford Historical Monographs. Oxford: Clarendon Press, 2001. Pp. xii + 244. ISBN 0-19-924639-4. £40.00 (hardback). British Journal for the History of Science, 36(2):246–247, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028244.

Thorpe:2003:BRE

[Tho03b] Charles Thorpe. Book review: Edward Teller with Judith Shoolery, Memoirs: A Twentieth-Century Journey in Science and Politics. Oxford: Perseus Press, 2001. Pp. xii + 628. ISBN 1-903985-12-9. £24.99 (hardback). British Journal for the History of Science, 36 (2):250–251, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028247.

Thorpe:2003:BRR

[Tho03c] Charles Thorpe. Book review:

Richard Polenberg (ed.), In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing. Ithaca and London: Cornell University Press, 2002. Pp. xxxii + 409. ISBN 0-8014-8661-0. $13.50 (paperback). British Journal for the History of Science, 36(3):378– 379, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028170.

Thorpe:2006:OTI

[Tho06] Charles Thorpe. Oppenheimer:

the tragic intellect. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-79845-3. xx + 413 + 12 pp. LCCN QC16.O62 T56

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0666/2006015223-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2006015223-d.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2006015223-t.html.

Thomas:2007:HWS

[Tho07] William Thomas. The heuristics of war: scientific method and the founders of operations research. British Journal for the History of Science, 40(2):251–274, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500719.

Thode:2009:BWN

[Tho09] Simon Thode. Bones and words in 1870s New Zealand: the moahunter debate through actor networks. British Journal for the History of Science, 42(2):225– 244, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tiles:2004:BRC

[Til04]  Mary  Tiles.

Book review: Cristina Chimisso, Gaston Bachelard: Critic of Science and the Imagination. Routledge Studies in Twentieth-Century Philosophy, 9. London and New York: Routledge, 2001. Pp. xi + 285. ISBN 0-415-26905-9. £71.99 (hardback). British Journal for the History of Science, 37(3):366–367, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028146;   http://www.
jstor.org/stable/4028448.

Timmermann:2002:BRR

[Tim02] Carsten Timmermann. Book review: Robert N. Proctor, The Nazi War on Cancer. Princeton and Oxford: Princeton University Press, 2000. Pp. x + 380. ISBN 0-691-07051-2. £10.50, $16.95 (paperback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028207.

Timmermann:2003:BRJ

[Tim03a] Carsten Timmermann. Book review: Jordan Goodman and Vivien Walsh, The Story of Taxol: Nature and Politics in the Pursuit of an Anti-Cancer Drug. Cambridge: Cambridge University Press, 2001. Pp. xiii + 282. £18.95, $27.95 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Timmermann:2003:BRK

[Tim03b] Carsten Timmermann. Book review: Kirk Jeffrey, Machines in our Hearts: The Cardiac Pacemaker, the Implantable Defibrillator, and American Health Care. Baltimore and London: Johns Hopkins University Press, 2001. Pp. xiii + 370. ISBN 0-8018-6579-4. £33.00 (hardback). British Journal for the History of Science, 36(2):251–252, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028248.

Timmermann:2003:BRN

[Tim03c] Carsten Timmermann. Book review: Nikolai Krementsov, The Cure: A Story of Cancer and Politics from the Annals of the Cold War. Chicago and London: University of Chicago Press, 2002. Pp. xvi + 261. ISBN 0-226-45284-0. £16.50, $26.00 (hardback). British Journal for the History of Science, 36(3):377– 378, September 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028169.

Timmermann:2003:JGV

[Tim03d] Carsten Timmermann. Jordan Goodman and Vivien Walsh, The Story of Taxol: Nature and Politics in the Pursuit of an Anti-Cancer Drug. Cambridge: Cambridge University Press, 2001. Pp. xiii + 282. £18.95, US$27.95 (hardback). British Journal for the History of Science, 36(1):87–127, March 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028325.

Timmermann:2004:BRA

[Tim04] Carsten Timmermann. Book review: Allen B. Weisse, Heart to Heart: The Twentieth-Century Battle Against Cardiac Disease. New Brunswick and London: Rutgers University Press, 2002. Pp. xvii + 415. ISBN 0-8135-3157-8. $35.00 (hardback). British Journal for the History of Science, 37(3):363–364, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028143;   http://www.
jstor.org/stable/4028445.

Timmermann:2005:BRH

[Tim05] Carsten Timmermann. Book review: Harry M. Marks, The Progress of Experiment: Science and Therapeutic Reform in the United States, 1900– 1990. Cambridge History of Medicine. Cambridge: Cambridge University Press, 2000. Pp. xii + 258. ISBN 0-521-78561-8. £14.95, $19.95 (paperback). British Journal for the History of Science, 38(1):118–119, March 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028599.

Timmermann:2008:BRA

[Tim08] Anke Timmermann. Book review: Allen G. Debus, The Chemical Promise: Experiment and Mysticism in the Chemical Philosophy, 1550–1800 (Selected Essays of Allen G. Debus). Sagamore Beach, MA: Science History Publications/USA, 2006. Pp. xxv + 548. ISBN 0-88135-296-9. $89.95 (hardback). British Journal for the History of Science, 41(1):134–135, March 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Torrey:2001:IPR

[TM01] E. Fuller (Edwin Fuller) Torrey and Judy Miller. The invisible plague: the rise of mental illness from 1750 to the present. Rutgers University Press, New Brunswick, NJ, USA, 2001. ISBN 0-8135-3003-2. xiv + 416 pp. LCCN RC455.2.E64

                                        T67 2001. URL http://www.e-
                                        streams.com/es0510/es0510_
                                        2136.htm;
                                        http://www.loc.gov/catdir/
                                        toc/fy031/2001019840.html.

Topham:2000:BSS

[Top00] Jonathan R. Topham. BJHS special section: book history and the sciences introduction. British Journal for the History of Science, 33(2):155–158, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027920.

Topham:2002:BRJ

[Top02] Jon Topham. Book review:

James A. Secord, Victorian Sensation: The Extraordinary Publication, Reception, and Secret Authorship of Vestiges of the Natural History of Creation. Chicago and London: University of Chicago Press, 2001. Pp. xix + 624. ISBN 0-226-74410-8. £22.50, $35.00 (hardcover). British Journal for the History of Science, 35(3):347–379, September 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Topham:2008:BRP

[Top08] Jonathan R. Topham. Book review: Peter Broks, Understanding Popular Science. Maidenhead: Open University Press, 2006. Pp. xi +183. ISBN 0-335-21548-3. £17.99 (paperback). British Journal for the History of Science, 41(4): 617–619, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Torrens:2002:PBG

[Tor02] H. S. Torrens. The practice of British geology, 1750–1850, volume CS736 of Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2002. ISBN 0-86078-876-8. xiv + 356 pp. LCCN QE261 .T67 2002.

Tosh:2009:BRR

[Tos09]  Nick Tosh.  Book review:

Robert N. Proctor and Londa Schiebinger (eds.), Agnotology: The Making and Unmaking of Ignorance. Stanford: Stanford University Press, 2008. Pp. viii + 298. ISBN 978-0-8047-5901-4. $24.95 (paperback). British Journal for the History of Science, 42(4):615–616, December 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Totelin:2004:BRM

[Tot04] Laurence M. V. Totelin. Book review: Maria Michela Sassi, The Science of Man in Ancient Greece. Translated by Paul Tucker. With a Foreword by Sir Geoffrey Lloyd. Chicago and London: University of Chicago Press, 2001. Pp. xxx + 224. ISBN 0-226-73530-3. £21.50, $34.00 (hardback). British Journal for the History of Science, 37(4):467–468, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028645.

Trevor-Roper:2006:EPV

[TR06] H. R. (Hugh Redwald) Trevor-Roper. Europe’s physician: the various life of Sir Theodore de Mayerne. Yale University Press, New Haven, CT, USA, 2006. ISBN 0-300-11263-7. xii + 438 + 8 pp. LCCN R507.M39 T74

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0643/2006010595-b.html;
http://www.loc.gov/catdir/
enhancements/fy0643/
2006010595-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip0611/2006010595.html.

Traister:2001:NAP

[Tra01] Barbara Howard Traister. The notorious astrological physician of London: works and days of Simon Forman. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-81140-9 (cloth). xviii + 250 pp. LCCN R489.F585 T73

2001. URL http://www.h-net.
org/review/hrev-a0c7o0-aa;
http://www.loc.gov/catdir/
bios/uchi051/00009585.html;
http://www.loc.gov/catdir/
description/uchi051/
00009585.html;
http://www.loc.gov/catdir/
toc/uchi051/00009585.html.

Traister:2007:BRL

[Tra07] Barbara H. Traister. Book review: Lauren Kassell, Medicine and Magic in Elizabethan London: Simon Forman: Astrologer, Alchemist, and Physician. Oxford Historical Monographs. Oxford: Clarendon Press, 2005. Pp. xviii + 281. ISBN 0-19-9279055. £50.00 (hardback). British Journal for the History of Science, 40 (1):132–133, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500697.

Tredwell:2006:BRE

[Tre06] Katherine A. Tredwell. Book review: Eric H. Ash, Power, Knowledge, and Expertise in Elizabethan England. Baltimore and London: Johns Hopkins University Press, 2004. Pp. ix + 265. ISBN 0-8018-7992-2. £30.00 (hardback). British Journal for the History of Science, 39(4): 606–607, December 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028529.

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001.

US$24.99. URL http://alsos.
wlu.edu/information.aspx?
id=784;
http://www.gbv.de/dms/
bowker/toc/9781903985120;
http://www.loc.gov/catdir/
enhancements/fy0833/
2001097880-d.html;
http://www.loc.gov/catdir/
toc/fy02/2001097880.html.

Tsou:2005:BRM

[Tso05]  Jonathan Y. Tsou.  Book

review: Michael Heidelberger and Friedrich Stadler (eds.), History of Philosophy of Science: New Trends and Perspectives. Vienna Circle Institute Yearbook, 9. Dordrecht, Boston and London: Kluwer Academic Publishers, 2002. Pp. x + 440. ISBN 1-4020-0509-1. £87.00, EUR138.00, $127.00 (hardback). British Journal for the History of Science, 38(3):

355–356, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028677.

Tsou:2007:BRG

[Tso07] Jonathan Y. Tsou. Book review: George A. Reisch, How the Cold War Transformed Philosophy of Science: To the Icy Slopes of Logic. Cambridge: Cambridge University Press, 2005. Pp. xiv + 418. ISBN 0-521-54689-3. £17.99, $26.99 (paperback). British Journal for the History of Science, 40(1): 153–155, March 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500711.

Tucker:2005:NEP

[Tuc05] Jennifer Tucker. Nature exposed: photography as eyewitness in Victorian science. The Johns Hopkins University Press, Baltimore, MD, USA, 2005. ISBN 0-8018-7991-4 (hardcover). ix + 294 pp. LCCN TR692 .T83

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0642/2004023568-b.html;
http://www.loc.gov/catdir/
enhancements/fy0642/
2004023568-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip052/2004023568.html.

Turner:2000:EIM

[Tur00] Gerard L’Estrange Turner. Elizabethan instrument makers: the origins of the London trade in precision instrument making. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-856566-6. xiv + 305 pp. LCCN Q185 .T854

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0611/2001268813-d.html;
http://www.loc.gov/catdir/
enhancements/fy0611/
2001268813-t.html;
http://www.loc.gov/catdir/
enhancements/fy0723/
2001268813-b.html.

Turchetti:2002:BRJ

[Tur02a] Simone Turchetti. Book review: Jean Medawar and David Pyke, Hitler’s Gift: Scientists who Fled Nazi Germany. London: Piatkus Books, 2001. Pp. xx + 268. ISBN 0-7499-22-397. £8.99 (paperback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turner:2002:BPR

[Tur02b] Stephen P. Turner. Brains/ practices/relativism: social theory after cognitive science. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-81739-3 (cloth), 0-226-81740-7 (paperback). ix + 214 pp. LCCN B831.3 .T86 2002. URL http:

//www.loc.gov/catdir/bios/
uchi051/2001037609.html;
http://www.loc.gov/catdir/
description/uchi052/
2001037609.html;   http://

www.loc.gov/catdir/toc/
uchi051/2001037609.html.

Turchetti:2003:ASG

[Tur03a] Simone Turchetti. Atomic secrets and governmental lies: nuclear science, politics and security in the Pontecorvo case. British Journal for the History of Science, 36(4):389–415, December 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028213.

Turner:2003:RAT

[Tur03b] Gerard L’Estrange Turner. Renaissance astrolabes and their makers, volume 766 of Variorum collected studies series. Ashgate/ Variorum, Aldershot, Hampshire, England, 2003. ISBN 0-86078-903-9. xii + 294 pp. LCCN QB85 .T85 2003. URL http://

www.loc.gov/catdir/toc/
fy042/2002043963.html.

Turchetti:2006:BRJ

[Tur06a] Simone Turchetti. Book review: John Krige and Dominique Pestre (eds.), Companion to Science in the Twentieth Century. Routledge World Reference. London and New York: Routledge, 2003. Pp. xxxv + 941. ISBN 0-415-286065-9. £26.99 (paperback). British Journal for the History of Science, 39(2):307–308, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028631.

Turner:2006:BRA

[Tur06b] Frank M. Turner. Book review: Aileen Fyfe, Science and Salvation: Evangelical Popular Science Publishing in Victorian Britain. Chicago and London: University of Chicago Press. 2004. Pp. xiv + 325. ISBN 0-226-27648-1. $25.00 (paperback). British Journal for the History of Science, 39(3): 458–459, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028500.

Turner:2008:BRG

[Tur08] Frank M. Turner. Book review: Gowan Dawson, Darwin, Literature and Victorian Respectability. Cambridge: Cambridge University Press, 2007. Pp. xii + 282. ISBN 978-0-521-87249-2 £50.00, $90.00 (hardback). British Journal for the History of Science, 41(3):469– 470, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turchetti:2009:BRJ

[Tur09a] Simone Turchetti. Book review: Jacob Darwin Hamblin, Poison in the Well: Radioactive Waste in the Oceans at the Dawn of the Nuclear Age. New Brunswick, NJ and London: Rutgers University Press, 2008. Pp. x + 311. ISBN 978-0-8135-4220-1. $49.95 (hardback). British Journal for the History of Science, 42(1): 149–150, March 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Turner:2009:BRT

[Tur09b] Frank M. Turner. Book review:

Thomas Dixon, The Invention of Altruism: Making Moral Meanings in Victorian Britain. Oxford: British Academy for the Oxford University Press, 2008. Pp. xiv + 420. ISBN 978-0-19-726426-3. £60.00 (hardback). British Journal for the History of Science, 42(2):312–313, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tweney:2009:BRA

[Twe09] Ryan D. Tweney. Book review:

Alice Jenkins (ed.), Michael Faraday’s Mental Exercises: An Artisan Essay Circle in Regency London. Liverpool: Liverpool University Press, 2008. Pp. xii + 250. ISBN 978-1-84631-140-6. £47.50, $85.00 (hardback). British Journal for the History of Science, 42(2):306–308, June 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tybjerg:2002:BRA

[Tyb02]  Karin Tybjerg.  Book re-

view: Andrew Barker, Scientific Method in Ptolemy’s Harmonics. Cambridge: Cambridge University Press, 2000. Pp. viii + 281. ISBN 0-521-55372-5. £45.00 (hardback). British Journal for the History of Science, 35(3): 347–379, September 2002. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Tybjerg:2004:BRJ

[Tyb04] Karin Tybjerg. Book review: J. Lennart Berggren and Alexander Jones, Ptolemy’s Geography : An Annotated Translation of the Theoretical Chapters. Princeton and Oxford: Princeton University Press, 2000. Pp. xiii + 192. ISBN 0-691-01042-0. £24.95, $39.50 (hardback). British Journal for the History of Science, 37(2):193–196, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028330.

Urry:2007:BRA

[Urr07]  James Urry.  Book review:

Amiria Henare, Museums, Anthropology and Imperial Exchange. Cambridge: Cambridge University Press, 2005. Pp. xix + 323. ISBN 0-521-83591-7. £48.00, $80.00 (hardback). British Journal for the History of Science, 40(2):280–282, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500723.

Unwin:2007:DES

[UU07] Patrick Unwin and Robert Unwin. ‘A devotion to the experimental sciences and arts’: the subscription to the great battery at the Royal Institution 1808–9. British Journal for the History of Science, 40(2):181–203, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500716.

Valente:2005:FUR

[Val05] K. G. Valente. ‘A finite universe?’ Riemannian geometry and the modernist theology of Ernest William Barnes. British Journal for the History of Science, 38(2):197–217, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

VanWyhe:2002:AHN

[Van02a] John Van Wyhe. The authority of human nature: the Schädellehre of Franz Joseph Gall. British Journal for the History of Science, 35(1):17–42, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028401.

VandenBroecke:2002:BRA

[Van02b] Steven Vanden Broecke. Book review: Anthony Grafton and Nancy Siraisi (eds.), Natural Particulars: Nature and the Disciplines in Renaissance Europe. Dibner Institute Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 1999. Pp. xi + 426. ISBN 0-262-07193-2. £33.50 (hardback). British Journal for the History of Science, 35(2):213–250, June 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028181.

VanRiper:2004:BRC

[Van04a] A. Bowdoin Van Riper. Book review: Claudine Cohen, The Fate of the Mammoth: Fossils, Myth, and History. Translated by William Rodarmor with a foreword by Stephen Jay Gould. Chicago and London: University of Chicago Press, 2002. Pp. xxxiv + 297. ISBN 0-226-11292-6. £19.00, $30.00 (hardback). British Journal for the History of Science, 37(1):107–108, March 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028265.

VanWyhe:2004:POV

[Van04b] John Van Wyhe. Phrenology and the origins of Victorian scientific naturalism. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-7546-3408-6. xvii + 282 pp. LCCN BF879 .V36 2004.

VanWyhe:2005:BRS

[Van05] John Van Wyhe. Book review: Stanley Finger, Origins of Neuroscience: A History of Explorations into Brain Function. Oxford: Oxford University Press, 2001. Pp. xviii + 462. ISBN 0-19-514694-8. £29.50 (paperback). British Journal for the History of Science, 38 (2):222–223, June 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028700.

Varcoe:2000:CPP

[Var00] Ian Varcoe. Comment: practical proposals by scientists for reforming the machinery of scientific advice, 1914–17. British Journal for the History of Science, 33 (1):109–114, March 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028068.

vanCleempoel:2002:CRS

[vC02] Koenraad van Cleempoel, editor.

A catalogue raisonné of scientific instruments from the Louvain School, 1530 to 1600, volume 65 (n.s. 28) of De diversis artibus. Brepols, Turnhout, Belgium, 2002. ISBN 2-503-51218-6. LCCN Q185 .C58 2002.

vanCleempoel:2005:AGC

[vCA^{+}05] Koenraad van Cleempoel, Silke Ackermann, et al. Astrolabes at Greenwich: a catalogue of the astrolabes in the National Maritime Museum, Greenwich. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-853069-2 (hardcover). ix + 339 + 18 pp. LCCN QB85 .C65

2005. URL http://www.loc.

gov/catdir/enhancements/
fy0628/2006296703-t.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2006296703-b.html;
http://www.loc.gov/catdir/
enhancements/fy0725/
2006296703-d.html.

vanDalen:1999:MGI

[vD05] D. (Dirk) van Dalen. Mystic, geometer, and intuitionist: the life of L. E. J. Brouwer. Clarendon Press, New York, NY, USA, 1999–2005. ISBN 0-19-850297-4 (vol. 1), 0-19-851620-7 (vol. 2). xv + 440 (vol. 1), x + 442–946 (vol. 2) pp. LCCN QA29.B697 D35

1999. URL http://www.loc.
gov/catdir/enhancements/
fy0606/99214056-d.html;
http://www.loc.gov/catdir/
toc/fy0602/99214056.html.

Vermij:2002:CCR

[Ver02] R. H. Vermij. The Calvinist Copernicans: the reception of the new astronomy in the Dutch Republic, 1575–1750, volume 1 of History of science and scholarship in the Netherlands. Koninklijke Nederlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2002. ISBN 90-6984-340-4 (hardcover). x + 433 pp. LCCN QB33.N4 V47

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0628/2005419549-d.html.

Vermij:2003:FNP

[Ver03] Rienk Vermij. The formation of the Newtonian philosophy: the case of the Amsterdam mathematical amateurs. British Journal for the History of Science, 36(2):183–200, June 2003. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028232.

Vernon:2004:BRC

[Ver04] Keith Vernon. Book review:

Colin Divall and Andrew Scott, Making Histories in Transport Museums. Making Histories in Museums. London and New York: Leicester University Press, 2001. Pp. x + 221. ISBN 0-7185-0106-3. £60.00 (hardback). British Journal for the History of Science, 37(2):238–239, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028365.

Vermeir:2005:MML

[Ver05] Koen Vermeir. The magic of the magic lantern (1660–1700): on analogical demonstration and the visualization of the invisible. British Journal for the History of Science, 38(2):127–159, June 2005. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vicedo:2007:BRP

[Vic07] Marga Vicedo. Book review: P.

Kyle Stanford, Exceeding Our Grasp: Science, History, and the Problem of Unconceived Alternatives. Oxford: Oxford University Press, 2006. Pp. xiv + 234. ISBN 0-19-517408-9. £26.99 (hardback). British Journal for the History of Science, 40(4): 619–621, December 2007. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Vidal:2005:BRB

[Vid05]  Fernando Vidal.  Book re-

view: Bernadette Bensaude-Vincent and Bruno Bernardi (eds.), Rousseau et les sciences. Paris, Budapest and Turin: L’Harmattan, 2003. Pp. 316. ISBN 2-7475-5100-8. EUR25.90 (paperback). British Journal for the History of Science, 38(3): 368–369, September 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028687.

vonLinne:2001:ND

[vLP01] Carl von Linné and Michael John Petry. Nemesis divina, volume 177 of Archives internationales d’histoire des idées = International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6820-7 (hardcover). xviii + 483 pp. LCCN BT135 .L4813

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0823/2001033539-d.html;
http://www.loc.gov/catdir/
enhancements/fy0823/
2001033539-t.html.

Wade:2003:DDO

[Wad03] Nicholas J. Wade. Destined for distinguished oblivion: the scientific vision of William Charles Wells (1757–1817). History and philosophy of psychology. Kluwer Academic… Plenum Publishers, New York, NY, USA, 2003. ISBN 0-306-47385-2. xi + 310 pp. LCCN BF241 .W319

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0818/2002034182-d.html;
http://www.loc.gov/catdir/
enhancements/fy0818/
2002034182-t.html.

Wagner:2005:BRH

[Wag05] Donald B. Wagner. Book review:

H. T. Huang, Science and Civilisation in China. Volume 6: Biology and Biological Technology. Part V: Fermentations and Food Science. Joseph Needham: Science and Civilisation in China. Cambridge: Cambridge University Press, 2000. Pp. xviii + 741. ISBN 0-521-65270-7. £95.00 (hardback). British Journal for the History of Science, 38(1): 103–104, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028586.

Waller:2003:BRG

[Wal03] John C. Waller. Book review:

Gerald Sweeney, ‘Fighting for the Good Cause’: Reflections on Francis Galton’s Legacy to American Hereditarian Psychology. Transactions of the American Philosophical Society, 91, Part 2. Philadelphia: American Philosophical Society, 2001.

Pp. x + 136. ISBN 0-87169-912-5. $18.00 (paperback). British Journal for the History of Science, 36(2):247–248, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028245.

Wallace:2004:DSE

[Wal04] William A. Wallace. Domingo de Soto and the early Galileo: essays on intellectual history. Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2004. ISBN 0-86078-964-0. xii + 1v. (various pgs.) pp. LCCN QC133 .W33 2004.

Wall:2006:PTT

[Wal06a] Cynthia Wall. The prose of things: transformations of description in the Eighteenth Century. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-87158-4. xiii + 316 pp. LCCN PR769 .W35

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0666/2005027014-b.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0666/
                                        2005027014-d.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        ecip0519/2005027014.html.

Waller:2006:BRA

[Wal06b] John C. Waller. Book review: Angelique Richardson, Love and Eugenics in the Late Nineteenth Century: Rational Reproduction and the New Woman. Oxford and New York: Oxford University Press, 2003. Pp.

250. ISBN 0-1981-8700-9. £51.00 (hardback). British Journal for the History of Science, 39(3): 464–466, September 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028504.

Waller:2006:BRD

[Wal06c] John C. Waller. Book review: Diana Wyndham, Eugenics in Australia: Striving for National Fitness. London: Diana Wyndham and the Galton Institute, 2003. Pp. 406. ISBN 0-9504066-7-8. £5.00 (paperback). British Journal for the History of Science, 39(1):141–142, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028566.

Warwick:2003:MTC

[War03] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-87374-9 (cloth), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN QC19.6 .W37 2003. URL

http://www.loc.gov/catdir/
bios/uchi051/2002153732.
html;   http://www.loc.gov/
catdir/description/uchi051/
2002153732.html;
http://www.loc.gov/catdir/
toc/fy037/2002153732.html.

Warwick:2008:BRJ

[War08] Andrew Warwick. Book review: Jeffrey Crelinsten, Einstein’s Jury: The Race to Test Relativity. Princeton and Oxford: Princeton University Press, 2006. Pp. xxix + 397. ISBN 0-691-12310-1. £22.99 (hardback). British Journal for the History of Science, 41(2):310–313, June 2008. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Watanabe:1997:SCE

[Wat97]  Masao Watanabe.  Science

and cultural exchange in modern history, Japan and the West. Hokusen-sha, Tokyo, Japan, 1997. ISBN 4-938424-73-8. xi + 406 pp. LCCN Q125 .W335 1997.

Waters:2000:ODH

[Wat00] D. W. Waters. Obituary: Derek Howse (1919–1998). British Journal for the History of Science, 33(2):223–225, June 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4027924.

Wallace:2002:ITA

[WB02] Alfred Russel Wallace and Andrew Berry. Infinite tropics: an Alfred Russel Wallace anthology. Verso, London, UK, 2002. ISBN 1-85984-652-1. xvii + 430 pp. LCCN QH366.2 .W342 2002.

Wallace:2002:ARW

[WC02]  Alfred  Russel  Wal-

lace and Jane R. Camerini. The Alfred Russel Wallace reader: a selection of writings from the field. Center books in natural history. The Johns Hopkins University Press, Baltimore, MD, USA, 2002. ISBN 0-8018-6781-9, 0-8018-6789-4 (paperback). xix + 219 pp. LCCN QH375 .W332

2002. URL http://www.loc.
gov/catdir/bios/jhu052/
2001000539.html;
http://www.loc.gov/catdir/
description/jhu051/
2001000539.html;
http://www.loc.gov/catdir/
toc/jhu051/2001000539.html.

Wear:2000:KPE

[Wea00] A. (Andrew) Wear. Knowledge and practice in early modern English medicine, 1550–1680. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-55226-5 (hardcover), 0-521-55827-1 (paperback). viii + 496 pp. LCCN R486 .W43 2000.

URL  http://www.loc.gov/
catdir/description/cam021/
00022600.html;   http://
www.loc.gov/catdir/samples/
cam031/00022600.html;
http://www.loc.gov/catdir/
toc/cam021/00022600.html.

Weart:2003:DGW

[Wea03] Spencer R. Weart. The discovery of global warming. New histories of science, technology, and medicine. Harvard University Press, Cambridge, MA, USA, 2003. ISBN 0-674-01157-0. x + 228 pp. LCCN QC981.8.G56 W43 2003.

Weber:2006:BRR

[Web06] Thomas P. Weber. Book reviews: Ross A. Slotten, The Heretic in Darwin’s Court: The Life of Alfred Russel Wallace. New York: Columbia University Press, 2004. Pp. x + 602. ISBN 0-231-13010-4. £39.50 (hardback). Martin Fichman, An Elusive Victorian: The Evolution of Alfred Russel Wallace. Chicago and London: University of Chicago Press, 2004. Pp. x+382. ISBN 0-226-24613-2. $40.00 (hardback). British Journal for the History of Science, 39(3):461–462, September 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028502.

Weber:2007:BRS

[Web07] Thomas P. Weber. Book reviews: Sharon E. Kingsland, The Evolution of American Ecology. Baltimore: The Johns Hopkins University Press, 2005. Pp. x + 313. ISBN 0-8018-8171-4. £33.50 (hardback). Kim Cuddington and Beatrix Beisner (eds.), Ecological Paradigms Lost. Routes of Theory Change. London: Academic Press, 2005. Pp. xxiv+435. ISBN 0-12-088459-3. £49.95 (hardback). British Journal for the History of Science, 40(3):453–455, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500760.

Weber:2008:BRH

[Web08] Thomas P. Weber. Book review: Horst Bredekamp, Darwins Korallen. Frühe Evolutionsmodelle und die Tradition der Naturgeschichte. Berlin: Wagenbach Verlag, 2005. Pp. 111. ISBN 978-3803151735. EUR22.50 (hardback). Olaf Breidbach, Visions of Nature: The Art and Science of Ernst Haeckel. Munich: Prestel Verlag, 2006. Pp. 299. ISBN 978-3791336640. $100.00 (hardback). British Journal for the History of Science, 41(2):306– 307, June 2008. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weber:2009:BRM

[Web09] Thomes P. Weber. Book reviews: Marcos Cueto, Cold Wars, Deadly Fevers: Malaria Eradication in Mexico 1955– 1975. Baltimore: Johns Hopkins University Press, 2007. Pp. xv + 264. ISBN 978-0-8018-8645-4. $45.00, £30.00 (hardback). Frank M. Snowden, The Conquest of Malaria: Italy, 1900– 1962. New Haven: Yale University Press, 2006. Pp. viii+296. ISBN 0-300-10899-0. $43.00, £25.00 (hardback). Sandra M. Sufian, Healing the Land and the Nation: Malaria and the Zionist Project in Palestine, 1920– 1947. Chicago and London: University of Chicago Press, 2007. Pp. xviii+385. ISBN 978-0-226-77935-5. $40.00, £21.00 (hardback). British Journal for the History of Science, 42(2):291– 293, June 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weeks:2007:BRJ

[Wee07] Sophie Weeks. Book review: Julie Robin Solomon and Catherine Gimelli Martin (eds.), Francis Bacon and the Refiguring of Early Modern Thought: Essays to Commemorate The Advancement of Learning (1605–2005). Literary and Scientific Cultures of Early Modernity. Aldershot: Ashgate, 2005. Pp. xiii + 257. ISBN 0-7546-5359-5. £47.50 (hardback). British Journal for the History of Science, 40(2):284–285, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4500725.

Weeks:2008:BRK

[Wee08]  Sophie Weeks.  Book re-

view: Katharine Park and Lorraine Daston (eds.) The Cambridge History of Science, Volume 3: Early Modern Science. Cambridge: Cambridge University Press, 2006. Pp. xxvii + 865. ISBN 0-521-57244-4. £90.00 (hardback). British Journal for the History of Science, 41(1): 136–137, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weeks:2009:BRB

[Wee09a] Sophie Weeks. Book review: Bernadette Bensaude-Vincent and William R. Newman (eds.), The Artificial and the Natural: An Evolving Polarity. Cambridge, MA: MIT Press, 2007. Pp. viii + 331. ISBN 978-0-262-02620-8. £25.95 (hardback). British Journal for the History of Science, 42(3):438–440, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weeks:2009:BRD

[Wee09b] Sophie Weeks. Book review:

Deborah E. Harkness, The Jewel House: Elizabethan London and the Scientific Revolution. New Haven: Yale University Press, 2007. Pp. xviii + 349. ISBN 978-0-300-11196-5. £19.99 (hardback). British Journal for the History of Science, 42(1):133– 135, March 2009. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Weiner:2000:MNE

[Wei00] Douglas R. Weiner. Models of nature: ecology, conservation, and cultural revolution in Soviet Russia. Pitt series in Russian and East European studies. University of Pittsburgh Press, Pittsburgh, PA, USA, 2000. ISBN 0-8229-5733-7. xii + 324 pp. LCCN QH77.S626 W45

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0743/2001266370-d.html.

Weinberg:2001:FSC

[Wei01] Steven Weinberg, editor. Facing up: science and its cultural adversaries. Harvard University Press, Cambridge, MA, USA, 2001. ISBN 0-674-00647-X. LCCN Q171 .W419 2001. URL

                                        http://www.loc.gov/catdir/
                                        toc/fy02/2001024219.html.

Weindling:2002:BRD

[Wei02a]  Paul Weindling.  Book re-

view: Daniel Gasman, Haeckel’s Monism and the Birth of Fascist Ideology. Studies in Modern European History, 33. New York: Peter Lang, 1998. Pp. vii + 482. ISBN 0-8204-4108-2. $69.95 (hardback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028419.

Weindling:2002:BRS

[Wei02b] Paul Weindling. Book review: Susanne Zimmermann, Die Medizinische Fakultät der Universität Jena während der Zeit des Nationalsozialismus. Ernst-Haeckel-Haus-Studien: Monographien zur Geschichte der Biowissenschaften und Medizin, Band 2. Berlin: VWB, 2000. Pp. 223. ISBN 3-86135-481-0. DM 48.00 (paperback). British Journal for the History of Science, 35(1):97–123, March 2002. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028420.

Weintraub:2002:HEB

[Wei02c] E. Roy Weintraub. How economics became a mathematical science. Science and cultural theory. Duke University Press, Durham, NC, USA, 2002. ISBN 0-8223-2856-9 (cloth), 0-8223-2871-2 (paperback). xiii + 313 pp. LCCN HB135 .W437

2002. URL http://www.h-net.
org/review/hrev-a0c5j0-aa.

Weisse:2002:HHT

[Wei02d] Allen B. Weisse. Heart to heart: the Twentieth Century battle against cardiac disease: an oral history. Rutgers University Press, New Brunswick, NJ, USA, 2002. ISBN 0-8135-3157-8 (cloth). xvii + 415 pp. LCCN RC666.5 .W456 2002. URL

http://www.loc.gov/catdir/
toc/fy034/2002020161.html.

Werrett:2001:WNC

[Wer01] Simon Werrett. Wonders never cease: Descartes’s météores and the rainbow fountain. British Journal for the History of Science, 34(2):129–147, June 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028007.

Werrett:2004:BRG

[Wer04] Simon Werrett. Book review:

Gerhard W. Kramer, The Firework Book: Gunpowder in Medieval Germany. The Journal of the Arms and Armour Society, 17(1). London: Arms and Armour Society, 2001. Pp. 89. ISSN 00004-2439. £10 (paperback). British Journal for the History of Science, 37(4):469– 470, December 2004. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028647.

Werrett:2009:BRB

[Wer09]  Simon Werrett.  Book re-

view: Bernadette Bensaude-Vincent and Christine Blondel (eds.), Science and Spectacle in the European Enlightenment. Aldershot: Ashgate, 2008. Pp. xi + 164. ISBN 978-0-7546-6370-6. £55.00 (hardback). British Journal for the History of Science, 42(3):457–459, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Withers:2008:HGP

[WHF08] Charles Withers, Rebekah Higgitt, and Diarmid Finnegan. Historical geographies of provincial science: themes in the setting and reception of the British Association for the Advancement of Science in Britain and Ireland, 1831–c. 1939. British Journal for the History of Science, 41(3): 385–415, September 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Whitworth:2001:EWR

[Whi01] Michael H. Whitworth. Einstein’s wake: relativity, metaphor, and modernist literature. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-818640-1. ix + 254 pp. LCCN PR478.S26 W47

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0613/2002277601-d.html;
http://www.loc.gov/catdir/
toc/fy033/2002277601.html.

White:2003:THM

[Whi03] Paul White. Thomas Huxley: making the “man of science”. Cambridge science biographies. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-64019-9, 0-521-64967-6 (paperback). xiv + 205 pp. LCCN QH31.H9 W56

2003. URL http://www.loc.
gov/catdir/description/
cam022/2002067654.html;
http://www.loc.gov/catdir/
samples/cam033/2002067654.
html;
http://www.loc.gov/catdir/
toc/cam031/2002067654.html.

Whitworth:2004:BRG

[Whi04] Michael H. Whitworth. Book review: George Levine, Dying to Know: Scientific Epistemology and Narrative in Victorian England. Chicago and London: University of Chicago Press, 2002. Pp. xi + 326. ISBN 0-226-47536-0. £31.50, $45.00 (hardback). British Journal for the History of Science, 37(2):212–213, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028342.

White:2005:BRS

[Whi05] Paul White. Book review: Sofia Åkerberg, Knowledge and Pleasure at Regent’s Park: The Gardens of the Zoological Society of London during the Nineteenth Century. Idéhistoriska skrifter,

36. Umeå: Department of Historical Studies, 2001. Pp. 254. ISBN 91-7305-147-0. No price given (paperback). British Journal for the History of Science, 38 (1):113–114, March 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028594.

White:2006:BRM

[Whi06] Paul White. Book review: Margaret E. Derry, Bred for Perfection: Shorthorn Cattle, Collies, and Arabian Horses since 1800. Animals, History, Culture. Baltimore and London: Johns Hopkins University Press, 2003. Pp. xvi + 198. ISBN 0-8018-7344-4. £27.50 (hardback). British Journal for the History of Science, 39(2):294–295, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

                                        tronic).  URL http://www.
                                        jstor.org/stable/4028620.

Wigelsworth:2008:BPP

[Wig08] Jeffrey R. Wigelsworth. Bipartisan politics and practical knowledge: advertising of public science in two London newspapers, 1695–1720. British Journal for the History of Science, 41(4): 517–540, December 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Willmoth:1997:FSN

[Wil97] Frances Willmoth, editor. Flamsteed’s stars: new perspectives on the life and work of the first Astronomer Royal, 1646–1719. Boydell Press, Woodbridge, Suffolk, UK, 1997. ISBN 0-85115-706-8. xiv + 271 pp. LCCN QB36.F6 F53 1997.

Williams:2002:RHC

[Wil02] Rosalind H. Williams. Retooling:

a historian confronts technological change. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-23223-5. xv + 252 pp. LCCN T173.8 .W55 2002. URL http:/

/www.loc.gov/catdir/toc/
fy036/2001060365.html.

Williams:2003:FBE

[Wil03] Roger Lawrence Williams.

French botany in the Enlightenment: the ill-fated voyages of La Pérouse and his rescuers, volume 182 of Archives internationales d’histoire des idées == International archives of the history of ideas. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1109-1 (hardcover). 240 pp. LCCN QK21.F8 W57

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0813/2002044855-d.html;
http://www.loc.gov/catdir/

                                        enhancements/fy0813/
                                        2002044855-t.html.

Williams:2006:BRW

[Wil06] Rosalind Williams. Book review: Walter A. Rosenblith (ed.), Jerry Wiesner: Scientist, Statesman, Humanist: Memories and Memoirs. With a Foreword by Edward M. Kennedy. Cambridge, MA and London: MIT Press, 2003. Pp. xxiv + 612. ISBN 0-262-18232-7. £22.95 (hardback). British Journal for the History of Science, 39(2):312–313, June 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

                                        jstor.org/stable/4028635.

Wilding:2008:RTS

[Wil08]  Nick Wilding.  The return

of Thomas Salusbury’s life of galileo (1664). British Journal for the History of Science, 41 (2):241–265, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2009:BRLa

[Wil09a]  Adrian Wilson.  Book re-

view: Lianne McTavish, Childbirth and the Display of Authority in Early Modern France. Aldershot: Ashgate, 2005. Pp. xiv + 257. ISBN 0-7546-3619-4. £45.00 (hardback). British Journal for the History of Science, 42(2):285–287, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2009:BRR

[Wil09b] Daniel C. S. Wilson. Book review: Richard G. Olson, Science and Scientism in Nineteenth-Century Europe. Urbana and Chicago: University of Illinois Press, 2008. Pp. 349. ISBN 978-0-252-07433-2. £14.99 (paperback). British Journal for the History of Science, 42(3): 469–470, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2009:BRM

[Wil09c] Duncan Wilson. Book review:

Mike Fortun, Promising Genomics: Iceland and deCODE Genetics in a World of Speculation. Berkeley: University of California Press, 2008. Pp. x + 330. ISBN 978-0-520-24751-2. £17.95 (paperback). British Journal for the History of Science, 42(4): 631–632, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wilson:2009:BRLb

[Wil09d] Emily Wilson. Book review:

Liba Taub, Aetna and the Moon: Explaining Nature in Ancient Greece and Rome. Corvallis, OR: Oregon State University Press, 2008. Pp. xiv + 138. $24.95. ISBN 978-0-87071-196-1 (hardback). British Journal for the History of Science, 42(3):454– 456, September 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Winterburn:2008:BRL

[Win08] Emily Winterburn. Book review: Liba Taub and Frances Willmoth (eds.), The Whipple Museum of the History of Science: Instruments and Interpretations, to Celebrate the 60th Anniversary of R. S. Whipple’s Gift to the University of Cambridge. Cambridge: Whipple Museum of the History of Science, 2006. Pp. xx + 492. ISBN 0-521-86609-X. £35.00 (hardback). British Journal for the History of Science, 41(2):313–314, June 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Winterbottom:2009:PUH

[Win09a] Anna Winterbottom. Producing and using the historical relation of Ceylon: Robert Knox, the East India Company and the Royal Society. British Journal for the History of Science, 42(4): 515–538, December 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wintroub:2009:HII

[Win09b] Michael Wintroub. The Heavens Inscribed: the instrumental poetry of the Virgin in early modern France. British Journal for the History of Science, 42(2):161–185, June 2009. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Withers:2001:GSN

[Wit01] Charles W. J. Withers, editor. Geography, science, and national identity: Scotland since 1520, volume 33 of Cambridge studies in historical geography. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-64202-7 (hardcover). LCCN DA757.7 .W58

2001. URL http://www.h-net.
org/review/hrev-a0c6j9-aa;
 http://www.loc.gov/catdir/
description/cam021/
2001025556.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2001025556-b.html;
http://www.loc.gov/catdir/
samples/cam031/2001025556.
html;   http://www.loc.gov/catdir/
toc/fy02/2001025556.html.

Withers:2007:PET

[Wit07] Charles W. J. Withers. Placing the Enlightenment: thinking geographically about the age of reason. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-90405-9 (cloth). xiii + 330 + 12 pp. LCCN B802 .W58

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0707/2006032563-b.html;
http://www.loc.gov/catdir/
enhancements/fy0707/
2006032563-d.html;   http:/
/www.loc.gov/catdir/toc/
ecip072/2006032563.html.

Wallis:2008:BRA

[WM08] Max K. Wallis and Trevor W.

Marshall. Book review: Aant Elzinga, Einstein’s Nobel Prize: A Glimpse behind Closed Doors. Sagamore Beach, MA: Science History Publications/USA, 2006. Pp xii + 128. ISBN 0-88135-283-7. $39.95 (hardback). British Journal for the History of Science, 41 (1):148–149, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Wood:2001:GPS

[WO01] Roger J. Wood and Vítězslav Orel. Genetic prehistory in selective breeding: a prelude to Mendel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850584-1 (hardcover). xvii + 323 pp. LCCN SF376.2 .W66

                                        2001. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0612/2001021144-d.html;
                                        http://www.loc.gov/catdir/
                                        toc/fy02/2001021144.html.

Wood:2006:MCK

[Woo06] Karen Wood. Making and circulating knowledge through Sir William Hamilton’s campi phlegraei. British Journal for the History of Science, 39(1):67–96, March 2006. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Worboys:2000:SGD

[Wor00] Michael Worboys. Spreading germs: disease theories and medical practice in Britain, 1865– 1900. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77302-4. xvi + 327 pp. LCCN RA427 .W68

2000. URL http://www.loc.
gov/catdir/description/
cam021/00023501.html;
http://www.loc.gov/catdir/
toc/cam026/00023501.html.

Worboys:2007:BRN

[Wor07] Michael Worboys. Book reviews: Nadja Durbach, Bodily Matters: The Anti-vaccination Movement in England, 1853– 1907. Durham, NC and London, Duke University Press, 2005. Pp. xiii + 276. ISBN 0-8223-3412-7. $84.95 (hardback). ISBN 0-8223-3423-2 (paperback). British Journal for the History of Science, 40(2):301–302, June 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4500737.

Webster:2005:PRH

[WPM05] Charles  Webster,  Mar-

garet Pelling, and Scott Mandelbrote, editors. The practice of reform in health, medicine, and science, 1500–2000: essays for Charles Webster. Ashgate, Aldershot, Hampshire, England, 2005. ISBN 0-7546-3933-9. xv + 376 pp. LCCN R486 .P73

2005. URL http://www.loc.
gov/catdir/toc/ecip0513/
2005015471.html.

Whewell:2001:PW

[WR01] William Whewell and Michael Ruse. Of the plurality of worlds. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-89435-5 (cloth), 0-226-89436-3 (paper). v + 510 pp.

LCCN QB54 .W47 2001. URL

                                        http://www.loc.gov/catdir/
                                        bios/uchi051/00067667.html;
                                         http://www.loc.gov/catdir/
                                        description/uchi051/
                                        00067667.html;
                                        http://www.loc.gov/catdir/
                                        toc/uchi051/00067667.html.

Wiesner:2004:JWS

[WR04] Jerome B. (Jerome Bert) Wiesner and Walter A. Rosenblith, editors. Jerry Wiesner: scientist, statesman, humanist: memories and memoirs. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-18232-7 (hardcover). xxiv + 612 pp. LCCN TK140.W54 J47 2004.

Wright:2000:TST

[Wri00] David Wright. Translating science: the transmission of Western chemistry into Late Imperial China, 1840–1900, volume 48 of Sinica Leidensia, 0169-9563. Brill, Leiden, The Netherlands, 2000. ISBN 90-04-11776-8. xxvi + 558 pp. LCCN QD49.C58 W75 2000.

Wright:2001:MDV

[Wri01]  David Wright.  Mental dis-

ability in Victorian England: the Earlswood Asylum, 1847– 1901. Oxford historical monographs. Clarendon Press, New York, NY, USA, 2001. ISBN 0-19-924639-4. vii + 244 pp. LCCN HV3008.G72 E278

                                        2001. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0612/2001021791-d.html.

Walters:2001:DMJ

[WS01a] S. M. (Stuart Max) Walters and E. A. (Elizabeth Anne) Stow. Darwin’s mentor: John Stevens Henslow, 1796–1861. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-59146-5 (hardcover). xx + 338 pp. LCCN QK31.H35 W35

2001. URL http://www.loc.
gov/catdir/description/
cam021/2001025196.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
2001025196-b.html;
http://www.loc.gov/catdir/
samples/cam031/2001025196.
html;
http://www.loc.gov/catdir/
toc/fy02/2001025196.html.

Weyl:2001:HWR

[WS^{+}01b] Hermann Weyl, Erhard Scholz, et al., editors. Hermann Weyl’s Raum–Zeit–Materie and a general introduction to his scientific work, volume 30 of DMV Seminar. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 3-7643-6476-9, 0-8176-6476-9. vi + 403 pp. LCCN QC173.58 .H37 2001.

Yamalidou:2000:BRA

[Yam00] Maria Yamalidou. Book review:

Aleksandr Yakovlevich Kipnis, Boris Efimovich Yavelov and John Shipley Rowlinson, Van der Waals and Molecular Science. Oxford: Clarendon Press, 1996. Pp. ix + 313. ISBN 0-19-855210-6. £60.00. British Journal for the History of Science, 33(2):231–254, June 2000. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4027940.

Yamalidou:2004:BRJ

[Yam04a] Maria Yamalidou. Book review: J. S. Rowlinson, Cohesion: A Scientific History of Intermolecular Forces. Cambridge: Cambridge University Press, 2002. Pp. viii + 333. ISBN 0-521-81008-6. £65.00 (hardback). British Journal for the History of Science, 37(3):358–359, September 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028139;   http://www.
jstor.org/stable/4028441.

Yamalidou:2004:BRP

[Yam04b] Maria Yamalidou. Book reviews: Peter Harman and Simon Mitton (eds.), Cambridge Scientific Minds. Cambridge: Cambridge University Press, 2002. Pp. viii + 343. ISBN 0-521-78612-6. £14.95 (paperback). David Millar, Ian Millar, John Millar and Margaret Millar, The Cambridge Dictionary of Scientists. Second edition. Cambridge: Cambridge University Press, 2002. Pp. xii+428. ISBN 0-521-00062-9. £14.95, $20.00 (paperback). British Journal for the History of Science, 37(4):466–467, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http:

//www.jstor.org/stable/
4028152;   http://www.
jstor.org/stable/4028644.

Yandell:2002:HCH

[Yan02]  Ben Yandell.  The honors

class: Hilbert’s problems and their solvers. A. K. Peters, Ltd., Wellesley, MA, USA, 2002. ISBN 1-56881-141-1, 1-56881-216-7 (paperback). ix + 486 pp. LCCN QA26 .Y36 2002.

Yeo:2001:EVS

[Yeo01a] Richard R. Yeo. Encyclopaedic visions: scientific dictionaries and enlightenment culture. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-65191-3 (hardcover). xxi + 336 pp. LCCN AE1 .Y46

2001.  URL http://www.e-
streams.com/es0805/es0805_
2064.htm;   http://www.loc.
gov/catdir/description/
cam021/00063005.html;
http://www.loc.gov/catdir/
enhancements/fy0731/
00063005-b.html;
http://www.loc.gov/catdir/
toc/cam021/00063005.html.

Yeo:2001:SPS

[Yeo01b] Richard R. Yeo. Science in the public sphere: natural knowledge in British culture, 1800– 1860, volume 726 of Variorum collected studies series. Ashgate, Aldershot, Hampshire, England, 2001. ISBN 0-86078-865-2. xviii + 302 pp. LCCN Q175.52.G7 Y46 2001.

Yeo:2004:BRA

[Yeo04] Richard Yeo. Book review: Ankete Heesen, The World in a Box: The Story of an Eighteenth-Century Picture Encyclopedia. Translated by Ann M. Hentschel. Chicago and London: University of Chicago Press, 2002. Pp. xii + 237. ISBN 0-226-32287-4. £13.00, $20.00 (paperback). British Journal for the History of Science, 37(2):208–209, June 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028338.

Yost:2000:BRC

[Yos00] Robinson M. Yost. Book reviews: Crosbie Smith, The Science of Energy: A Cultural History of Energy Physics in Victorian Britain. London and Chicago: The Athlone Press and University of Chicago Press, 1998. Pp. xi + 404. ISBN 0-485-11431-3, £55.00 (hardback); ISBN 0-485-12145-X; £19.95 (paperback). British Journal for the History of Science, 33(1):115–124, March 2000. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4028074.

Zagorin:2001:FBC

[Zag01] Perez Zagorin. Francis Bacon’s concept of objectivity and the idols of the mind. British Journal for the History of Science, 34(4):379–393, December 2001. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028370.

Zenderland:1998:MMH

[Zen98] Leila Zenderland. Measuring minds: Henry Herbert Goddard and the origins of American intelligence testing. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-44373-3 (hardcover). ix + 466 pp. LCCN BF431 .Z46

1998. URL http://www.loc.
gov/catdir/description/
cam028/97006101.html;
http://www.loc.gov/catdir/
toc/cam026/97006101.html.

Zinsser:2006:EDC

[ZH06] Judith P. Zinsser and Julie Candler Hayes, editors. Emilie Du Châtelet: rewriting Enlightenment philosophy and science, volume 2006:01 of SVEC, 0435-2866. Voltaire Foundation, Oxford, UK, 2006. ISBN 0-7294-0872-8. xi + 325 pp. LCCN PQ1981.D55 Z66 2006.

Ziche:2000:MWW

[Zic00] Paul Ziche, editor. Monismus um 1900: Wissenschaftskultur und Weltanschauung. (German) [Monismus around 1900: scientific culture and world view], volume 4 of Ernst-Haeckel-Haus-Studien. VWB, Verlag für Wissenschaft und Bildung, Berlin, Germany, 2000. ISBN 3-86135-483-7. 200 pp. LCCN B827 .M66

Zigman:2004:BRP

[Zig04] Peter Zigman. Book review: Paul Ziche (ed.), Monismus um 1900: Wissenschaftskultur und Weltanschauung. Ernst-Haeckel-Haus Studien: Monographien zur Geschichte der Biowissenschaften und Medizin, 4. Berlin: Verlag für Wissenschaft und Bildung, 2000. Pp. 200. ISBN 3-86135-483-7. EUR24.00 (paperback). British Journal for the History of Science, 37(4):484, December 2004. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (elec-

tronic).  URL http://www.
jstor.org/stable/4028660.

Zilsel:2003:SOM

[Zil03] Edgar Zilsel. The Social Origins of Modern Science, volume 200 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 0-7923-6457-0, 94-011-4142-8 (e-book). ISSN 0068-0346. lix + 267 pp. LCCN Q174 .B67 vol. 200;

Q175.5. URL http://www.loc.
gov/catdir/enhancements/
fy0813/00042657-d.html;
http://www.loc.gov/catdir/
enhancements/fy0813/
00042657-t.html;
https://link.springer.com/
book/10.1007/978-94-

011-4142-0. Edited 59 years after the author’s death by Diederick Raven, Wolfgang Krohn, and R. S. (Robert Sonné) Cohen.

Ziman:2000:TIE

[Zim00a] J. M. (John M.) Ziman, editor. Technological innovation as an evolutionary process. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-62361-8. xvii + 379 pp. LCCN T173.8 .T363 2000. URL http:

//www.loc.gov/catdir/
description/cam029/
99015474.html;   http://
www.loc.gov/catdir/samples/
cam032/99015474.html;
http://www.loc.gov/catdir/
toc/cam021/99015474.html.

Zimmermann:2000:MFU

[Zim00b] Susanne Zimmermann. Die Medizinische Fakultät der Universität Jena während der Zeit des Nationalsozialismus. (German) [The medical faculty of the University of Jena during the time of National Socialism], volume 2 of Ernst-Haeckel-Haus-Studien. Verlag für Wissenschaft und Bildung, Berlin, Germany, 2000. ISBN 3-86135-481-0. 223 pp. LCCN R786.F77 Z55 2000.

Zinsser:2005:MWB

[Zin05] Judith P. Zinsser, editor. Men, women, and the birthing of modern science. Northern Illinois University Press, DeKalb, IL, USA, 2005. ISBN 0-87580-340-7 (hardcover). viii + 215 pp. LCCN Q130 .M46

2005. URL http://www.loc.
gov/catdir/toc/ecip055/
2004030000.html.

Zuidervaart:2005:ECM

[Zui05]  Huib  J.

Zuidervaart. An eighteenthcentury medical–meteorological society in the Netherlands: an investigation of early organization, instrumentation and quantification. part 1. British Journal for the History of Science, 38(4): 379–410, December 2005. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Zuidervaart:2006:ECM

[Zui06]  Huib  J.

Zuidervaart. An eighteenthcentury medical–meteorological society in the Netherlands: an investigation of early organization, instrumentation and quantification. part 2. British Journal for the History of Science, 39(1):49–66, March 2006. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).